A Complete Bibliography of Publications in the Journal of Cell Biology: 2000–2004

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

24 May 2021 Version 1.01

Title word cross-reference

1 [TPW^{+}04, XJW^{+}04]. 118 [TSY^{+}02]. 12 [CRP^{+}04]. 2 [CRS^{+}03]. 3 [BSR^{+}03, FBG^{+}01, HBG^{+}02, IFP^{+}03, KMH^{+}04, LeB03-83, LeB03-85, MMG^{+}04, OSWG02, OMB^{+}01, PLC^{+}02, VBR^{+}01a, WSR03, WE02, ZCH^{+}02]. 5 [DOB^{+}01, NHG^{+}03, PWY^{+}03]. 6 [AHA^{+}04, GGGK03, TST^{+}03]. 9 [SHE^{+}02]. ^{+} [AZB^{+}00, BPMG00, GF03, LeB04-67, MAG^{+}04, PSE^{+}03, RLTC^{+}02, TMK^{+}00, TTP^{+}01, YpHRL03]. − [RLTC^{+}02, SV03]. ^{−/}− [DPO^{+}04]. ^{1} [CYC^{+}04]. ^{2} [EBWC01, HCD^{+}00]. ^{2+} [ACP^{+}02, BPMG00, BKZ^{+}03, CRH02, DC02a, DVE^{+}00, DH02, Dov02-69, DFYL00, DFZ^{+}03, FFKC00, FYI^{+}03, GMY^{+}03, HTPC04, HDP^{+}01, pHYpXL00, IKA^{+}03, JJM^{+}02, LeB02-28, LeB03-58, LeB04-75, LRA^{+}02, LPPT^{+}02, LC04, MH02, MCH^{+}00, MZ00, MW04, MPV^{+}01, NMH^{+}04, OTB03, PDMO03, PFM^{+}00, PLW^{+}04, PMB^{+}00, RPS^{+}02, RLC00, RMMP04, SJW^{+}04, SMT^{+}03, SSP^{+}03, SZZ^{+}00, UTH^{+}02, WLW^{+}04, YpHRL03]. _{cyt}^{2+} [SM04b]. ^{Cdh1} [HPE^{+}01]. ^{Cip1}/W AF 1 [TYA^{+}02]. ^{ctn} [TAD^{+}00]. ^{Glued} [LVWA01, VMH^{+}02]. ^{INK4a} [OBG^{+}03, PDW^{+}00]. ^{KIP1} [TJS^{+}04, MLS^{+}01, SCTM00]. ^{NTR} [LS02]. ^{Shc} [VCDHD03]. _{(1,2)} [HHOP02]. _{0} [PQF^{+}00, WFQ^{+}00, YKW^{+}00, vdLBK^{+}04]. _{1} [HBG^{+}02, vdLBK^{+}04]. _{1−−42} [ZMGL02]. _{180} [RCL^{+}00]. _{1S} [FKG00]. _{2} [ARK^{+}00, BRM^{+}00, ELO^{+}01, HP00, HGC02, HSW00, HWBD^{+}01, HCD^{+}00, LFT^{+}00, POW^{+}01a, PBD^{+}04, ST03, SHA^{+}03, Tum04z, ZCW^{+}03]. _{2α} [HP03]. _{3} [CAW^{+}04, HGC00, MZR^{+}04, PGSE^{+}01, SWG^{+}03, TFAM^{+}04, WLW^{+}04]. _{3/6} [vdHvODML^{+}00]. _{4} [CAW^{+}04, PWC^{+}01]. _{5} [RDNB02]. _{B} [NYT^{+}03]. _{CRAC} [MZ00]. _{i} [BFH^{+}01, MTV^{+}00]. _{L} [HOvD^{+}00, KSS^{+}03, KCY^{+}04, RRL^{+}00]. M [WGF^{+}00]. _{o} [ELO^{+}01]. _{SOC} [WSL^{+}01]. α [AKK^{+}00, AMF01, ALJ00, ABF^{+}03, ANC^{+}02, BSD^{+}01, BS04a, BIC^{+}03, BFG^{+}04, BMKA01, BMF00, BSL^{+}01, BBDM02, BLPP01, BWN^{+}01, CWA^{+}03, CYC^{+}04, CLB^{+}02, CBL^{+}02, DSB^{+}02, DWD^{+}00, EPH^{+}02, FKG00, FNKH02, GPDvH^{+}03, GMRYM^{+}02, GPL^{+}02, GHK^{+}03, GAC^{+}03, HBSQ01, HGC02, HFK^{+}03, HYMS^{+}02, JW04, JLJD03, JON^{+}03, KNI^{+}04, KVC^{+}03, KYS^{+}02, KVM03, KCY^{+}04, KKL^{+}01, KSM^{+}01, LWDH01, LCS^{+}02, LPT03, MKD^{+}01, MO01, MDT^{+}01, MGLPM00, MSY^{+}04, MCF^{+}02, NM02, NZHR01, ONM00, OEM^{+}02, PUK02, PWC^{+}01, PPA^{+}03, RSG00, RRM^{+}00, RBD^{+}01, RDS02, SMS^{+}04a, SJIM01, SWH^{+}02, SHPY02, SFSD00, SGO^{+}00, SKH03, TSMS01, TCH^{+}02, TTS00, TE01, WPM^{+}00, WGF^{+}00, WMG^{+}04, WCGT^{+}00, YSK^{+}04, YCK^{+}03, ZLR^{+}03, ZSY^{+}03, ZLVS02, ZTK^{+}03, dVKS04, vWJK^{+}03]. α_{2}β_{1} [ISID^{+}03]. α_{3}β_{1} [WFI^{+}04]. α_{V} β_{3} [ELO^{+}01, BBR04, CBK^{+}00, MZT^{+}03, NLRD01, SHN^{+}01]. α_{v}β_{5} [CBK^{+}00]. α_{V} β_{6} [ACMR04]. α_{v}/β_{3} [vWJK^{+}03]. β [ACMR04, APLB00, BCG03, BEG01, BAD^{+}00, BMD^{+}00, BKD^{+}00, BFG^{+}04, BDR^{+}03, BSL^{+}01, BFSO^{+}04, BWN^{+}01, CLG^{+}03, CSIK03, CMC^{+}02, CWA^{+}03, CGY^{+}01b, CSD00, CLB^{+}02, CBL^{+}02, DCC^{+}02, DSB^{+}02, DZT^{+}00, DH02, DWD^{+}00, ESS^{+}00, EKH^{+}03, EOJ^{+}03, EPH^{+}02, EPH^{+}03, FNZ^{+}03, FHM^{+}03, FPP^{+}02, GPDvH^{+}03, GMRYM^{+}02, GHK^{+}03, GWG01, GG04, GPZ^{+}02, HDJ00, HIE^{+}01, HCK^{+}00, HBK^{+}02, HCC02a, HBSQ01, HRB^{+}01, HKE^{+}04, HDL02, HVB^{+}00, HKO03, IAG^{+}00, IEJ^{+}01, IYT^{+}04, JLK^{+}02, JW04, KLE^{+}02, KH04, KHH^{+}01, KMH^{+}00, KWSK^{+}04, KS02, LGGS^{+}04, LZC^{+}03, LeB04-100, LSK01, LDP02, LXL^{+}04b, LSMS03, LCI^{+}01, LCS^{+}02, LCG^{+}04, LTD^{+}01, MKD^{+}01, MSJ^{+}02, MEFC03, MDT^{+}01, MGLPM00, MCB00a, MCF^{+}02, MWC^{+}02, ONM00, OEM^{+}02, PWY^{+}03, PGSE^{+}01, PTM^{+}01, PUK02, PLC^{+}02, RBM^{+}00, RWCC01, RM01, RSG01, RP03a, RJA^{+}03, RRJ^{+}01, RDS02, SMS^{+}04a, SYW^{+}03]. β [SHPY02, SSH^{+}04, SDL^{+}03, SKF^{+}01, SGO^{+}00, SKH03, SEW^{+}01, TM00, TWBV^{+}01, TJC^{+}03, THK^{+}00, TMA^{+}04, WNM^{+}03, Wel02-67, WBT^{+}03, WHP^{+}02, WMG^{+}04, WCGT^{+}00, YCX^{+}01, YCK^{+}03, ZLR^{+}03, ZLVS02, ZMGL02, ZTK^{+}03, dVKS04, vdFKK^{+}02]. β_{1} [GFM^{+}04b, HBB^{+}02]. β_{2} [GCO^{+}04, KRA^{+}01, LSMS03, REK^{+}03, KCY^{+}04, WGF^{+}00]. β_{3} [NH03b, XBL^{+}03]. β_{Heavy} [MWK^{+}02]. βγ [BSR^{+}03, NB03a]. Δ [CM01, IEJ^{+}01, MC02, YRC^{+}04, FYI^{+}03, JKB^{+}03, MOMK03, PLR03]. ϵ [AKH^{+}04, BMD^{+}00, DWFAdL^{+}02, LUB^{+}02, LSK01, MZ00, WPO00, WPS^{+}01]. γ [BRS^{+}01, BSR^{+}03, CS03a, CBC^{+}01, GPAS^{+}01, GBY^{+}03, GMC^{+}00, GRSL^{+}04b, HOK^{+}02, HLB^{+}00, HF03, ISID^{+}03, JKB^{+}03, KKW^{+}03, KJK^{+}03, LUB^{+}02, LDI^{+}03, MBS^{+}01, MPR^{+}01, NLRD01, SLG02, WLW^{+}04, ZKW^{+}00]. ι [MJY^{+}04]. κ [BLPP01, CSIK03, Ern00, GV00, GPLS02, KSR^{+}04, MKK^{+}00b, RRM^{+}00, SLR^{+}03, SRSW04, SDS^{+}04, ZLR^{+}03]. λ [KMH^{+}04]. μ [CWA^{+}03, HLB^{+}02, JFS^{+}03, RCS^{+}02, RWH02, SMW^{+}03, VYC^{+}00]. ω [MWC^{+}02]. σ [DGB^{+}00, JKB^{+}03, LGGS^{+}04]. θ [VBH^{+}02]. ζ [KMH^{+}04, LPT03, PSD^{+}04b, TP01].

-1 [ZN01]. -2 [IWG^{+}01]. -3 [DCC^{+}02, MWC^{+}02]. -4 [FWP^{+}00]. -acetyltransferase [CYC^{+}04]. -actin [FHM^{+}03, GPZ^{+}02, RJA^{+}03]. -Actinin [LWDH01, TTS00, YSK^{+}04]. -activated [CSIK03, FNKH02]. -Adaptin [HLB^{+}00]. -Adducin [BIC^{+}03]. -Adenosylmethionine [PSKK^{+}00]. -adrenergic [REK^{+}03]. -Amyloid [SKF^{+}01, EKH^{+}03]. -Arrestin [DZT^{+}00, PLC^{+}02]. -barrel [IYT^{+}04]. -bisphosphate [RWH02]. -Catenin [CGY^{+}01b, ESS^{+}00, IEJ^{+}01, LCG^{+}04, PTM^{+}01, SKF^{+}01, CLG^{+}03, CMC^{+}02, GG04, HBK^{+}02, LZC^{+}03, LSK01, LDP02, MSJ^{+}02, MOMK03, PLR03, PGSE^{+}01, SEW^{+}01, TJC^{+}03, WBT^{+}03, GWG01, HVB^{+}00, RSG01, TM00, THK^{+}00, MO01]. -Catenin-induced [WHP^{+}02]. -Catenin/Lymphoid [ESS^{+}00]. -Catenin/T [CGY^{+}01b, SKF^{+}01]. -cell [BFSO^{+}04]. -cells [TMA^{+}04]. -Cl [RLTC^{+}02]. -controlled [LRA^{+}02]. -Coupled [MTV^{+}00]. -Dependent [MCH^{+}00, DFYL00, HDP^{+}01, LC04, ZCH^{+}02]. -dystrobrevin [GAC^{+}03]. -Dystroglycan [BMF00, KSM^{+}01]. -Factor [CD00]. -Formyl [GHS00]. -G [MCB00a]. -globin [LCI^{+}01]. -Glucosidase [GPL^{+}02]. -Glycanase [SPH^{+}00]. -handling [PDMO03]. -Helical [KKL^{+}01]. -herpesvirus [TE01]. -independent [HP04, PWS^{+}01]. -induced [HKE^{+}04, SMS^{+}04a, SFSD00, SMT^{+}03]. -Integrin [PWC^{+}01]. -integrin-cytoskeleton [vWJK^{+}03]. -kinase [BSR^{+}03, FBG^{+}01, HBG^{+}02, IFP^{+}03, KMH^{+}04, OMB^{+}01, PWY^{+}03, PLC^{+}02, WE02]. -kinases [VBR^{+}01a, WSR03]. -like [STJ^{+}01]. -Mediated [ELO^{+}01, ACMR04, BFH^{+}01, WLW^{+}04]. -Netrin [KMH^{+}00]. -p38-Signaling [vdHvODML^{+}00]. -peptide [KH04]. -permeable [IKA^{+}03]. -phosphatase [DOB^{+}01]. -phosphate [AHA^{+}04, GGGK03, TPW^{+}04, TST^{+}03]. -phosphorylated [TSY^{+}02]. -RB [OBG^{+}03]. -residue [CRP^{+}04]. -secretase [CBC^{+}01, GRSL^{+}04b, HF03, SYW^{+}03]. -sensitive [ACP^{+}02]. -Signaling [BPMG00]. -smooth [HGC02]. -Spectrin [HDJ00, MWK^{+}02]. -Syntrophin [AKK^{+}00, AMF01]. -synuclein [WMG^{+}04]. -Terminal [ARK^{+}00, HSW00, HCD^{+}00, BRM^{+}00, HGC02, PBD^{+}04, ZCW^{+}03]. -Toxin [BSD^{+}01]. -tubulin [DWFAdL^{+}02, HOK^{+}02, SLG02, GMC^{+}00, ZKW^{+}00].

/Calcium [HGC00]. /Calmodulin [pHYpXL00]. /CXCR4 [KVC^{+}03]. /DER [HCC02b]. /DER-independent [HCC02b]. /H [AZB^{+}00]. /K [GF03]. /myristoyl [OTB03]. /Nuclear [RRM^{+}00]. /p101 [BSR^{+}03]. /Rho/mDia [ABF^{+}03]. /SLP [KSK^{+}00]. /SLP-76-Associated [KSK^{+}00].

1 [AS00, AKT01, ASMW01, ASAJ01, ABF^{+}03, APLB00, BMG^{+}01, BGM03, BWRT03, BYMS^{+}02, BMC00, CSIK03, CGY01a, COB01, CGY^{+}01b, CBRBM04, DM02, DBB^{+}02a, DM00, ERS^{+}04, EKdM^{+}04, FWP^{+}00, FSK^{+}00, FPSM01, GK03a, GJB^{+}00, IWG^{+}01, IBS^{+}02, ITF^{+}02, IFP^{+}03, JKB^{+}03, JOF^{+}02, JWJJ00, JTK^{+}02, KLE^{+}02, KWH^{+}00, KMiM^{+}01, KCL^{+}00, KMS00, KFO04, KJK^{+}03, LGGS^{+}04, LeB02-65, LeB03-62, LeB04-55, LAF^{+}00, LM00a, LCT^{+}04, LPL^{+}04, MLLA00, MWF02, MDQ^{+}03, MMH^{+}00, MTW^{+}04, MCF^{+}02, NB03a, PGS^{+}01, PMKM03, PCB^{+}03, PSE^{+}03, PPA^{+}03, RBE^{+}02, RWK^{+}04, RGK^{+}03, SBMB^{+}04, STA^{+}01, SSM^{+}04, SMSF00, TDhL^{+}02, TFF03, TTP^{+}01, TGD^{+}03, WSW^{+}00, WND^{+}00, ZN01, KVC^{+}03, PUK02, HYMS^{+}02]. 1-dependent [GHK^{+}03]. 1-expressing [PPA^{+}03]. 1-Integrins [APLB00]. 1-Mediated [PTM^{+}01]. 1-Syntrophin [HYMS^{+}02]. 1/ [LeB03-40]. 1/CD148 [LZC^{+}03]. 10 [HMAM01, TGM^{+}01]. 11 [KWH^{+}00]. 112 [RMG^{+}00]. 12 [AOJ^{+}04, HBSJ04]. 13F [YCK^{+}03]. 14 [GGGN02]. 14-3-3 [BKS^{+}02, DGB^{+}00, MBLCE03, PSD^{+}04b, Wel02a]. 14-Kilodalton [WFT^{+}01]. 145 [BMM^{+}01]. 145-Kilodalton [BMM^{+}01]. 15.5-kD [GBOL03]. 16 [MSN^{+}02]. 160 [MMR^{+}00]. 170 [LKM^{+}04, LeB04-64, PPGN^{+}02, TDFV02]. 175 [GTR^{+}03]. 19 [JJM^{+}02, PRS^{+}04]. 1A [FPM^{+}03, TM04]. 1B [AFK^{+}03, DLK^{+}02, FPM^{+}03, FTD^{+}01, RKF^{+}04]. 1c [DLY^{+}02, SCL^{+}01]. 1R [WDL^{+}04].

2 [AAM^{+}04, BFC^{+}02, BGA^{+}04, BBDM02, CNHK02, CA00, CVZ^{+}04, CQH^{+}00, CS03b, DZT^{+}00, Dov03c, Dov03-30, DOB^{+}01, EHM^{+}00, FSD00, FKI^{+}01, FFST01, HFG^{+}04, HOvD^{+}00, HBSQ01, HKO03, ISID^{+}03, JON^{+}03, KHvOD00, KSS^{+}03, LMGM^{+}02, MMR^{+}00, MEV^{+}04, MTB^{+}02, MHK^{+}01, MBSR03, ONS^{+}00, OMB^{+}01, PWY^{+}03, PFM^{+}00, PLC^{+}02, RBW^{+}02, SNF^{+}02, SKH03, SKO04, TSMS01, WSL^{+}03, ZVPK03, KKW^{+}03, BFG^{+}04]. 2- [VYC^{+}00]. 2-adaptin [RWH02]. 2-adrenergic [PLC^{+}02]. 2/ [AAD03, SSW^{+}01a]. 2/Raftk [SNL^{+}00]. 2/S100A10 [BGA^{+}04]. 20 [Dov04i, PBD^{+}04]. 205 [MGL^{+}00]. 26s [HKK^{+}00]. 28 [MSN^{+}02]. 29 [Wel02-51]. 2A [NCMO^{+}02]. 2b [LC04]. 2h3 [HGC00, MF01a].

3’ [ARM02, BFC^{+}02, BMD^{+}00, BMLU02, BRS^{+}01, BCM04, DCC^{+}02, DBB02b, Dov03-31, EWD02, EKT^{+}00, FFY^{+}00, FWP^{+}00, FCF^{+}01, FSM^{+}01, GRBD01, GTPMU00, HCK^{+}00, HP00, HBSQ01, HC02, HBD^{+}02, IOLA^{+}00, ISF^{+}01, JKB^{+}03, JRW^{+}01, KWH^{+}00, KNIO01, KSK^{+}00, LBP00, Mil02, OEM^{+}02, PRLR02, POH^{+}04, RRM^{+}00, SPW00, SAG^{+}01, TJC^{+}03, WKS^{+}00, WSL^{+}03, WHS00, dVKS04, CWA^{+}03, HBSQ01, KYS^{+}02, SKH03, ZTK^{+}03]. 3- [KRU^{+}04]. 3-activating [LJK^{+}01]. 3-Kinase [MBS^{+}01, BMD^{+}00, GTPMU00, HBSQ01, BFC^{+}02]. 3-kinase/AKT [BFC^{+}02]. 3-P-dependent [BMLU02]. 3’-UTR [ARM02]. 35 [SJB^{+}03]. 38k [ONS^{+}00]. 3A [EWD02]. 3B [NOM^{+}04]. 3K [LCT^{+}04]. 3t3 [CDWB01, TYY^{+}00, BKZ^{+}03, SRSW04, TIO^{+}02].

4 [AFR01, BTD^{+}00, FYI^{+}03, HBSQ01, HEW^{+}01, JPGR^{+}00, KWSK^{+}04, LCGR00, MR01b, RyHHK00, YHF^{+}01, GHK^{+}03, LCS^{+}02, SHPY02]. 4-Associated [HBSQ01]. 45 [AP00]. 4e [DLS00, CBL^{+}02]. 4q [TSL04].

5 [AKM^{+}04, BTD^{+}00, BKD^{+}04, BM02, KGC^{+}00, LCT^{+}04, LHvdH00, NCUJ^{+}00, OBG^{+}03, RyHHK00, TNK^{+}00, YHF^{+}01, ALJ00, DSB^{+}02, KVM03, LWDH01]. 5-biphosphate [MHS01]. 5-Bisphosphate [AZB^{+}00, BRM^{+}00, BND^{+}02, BRY^{+}01, BTD^{+}00, RyHHK00, YHF^{+}01]. 5-MCDE [LCT^{+}04]. 5-mediated [ZLVS02]. 5-trisphosphate [CVZ^{+}04]. 5/pipetail [WBT^{+}03]. 53bp1 [RIDC01, SCMH00, KMG^{+}03, SRL^{+}04, WRSMO^{+}04]. 58 [GKM^{+}01]. 58-Kd [GKM^{+}01]. 59 [PPGN^{+}02]. 5a [TJC^{+}03].

6 [ADL^{+}03, BMC00, VCGB^{+}02, WHM^{+}02, CBL^{+}02, GPDvH^{+}03, HBSQ01, KYS^{+}02, MKD^{+}01, MGLPM00, ONM00, SGO^{+}00, ZLR^{+}03]. 6-Bisphosphatase [BMC00]. 6-O [ADL^{+}03]. 60- [SMW^{+}03]. 60S [GSP^{+}02a]. 63 [KKL^{+}01].

7 [PAR^{+}04, RDS02, WTG01, ANC^{+}02, BWN^{+}01]. 70s [ASP^{+}00]. 76 [KSK^{+}00].

8 [CHM04, SZZ^{+}00, BSL^{+}01]. 87 [WLW^{+}04].

9 [SRD^{+}02, CYC^{+}04]. 95 [BKD^{+}04, EHCC^{+}00, RPZ^{+}02, TMK^{+}00].

A-kinase [ASS02]. A-mediated [EGC^{+}03]. A/C [KMP02b]. A1 [vdHvODML^{+}00]. A11 [SMT^{+}03, SMS^{+}04a]. A2G [MPG^{+}03]. A5 [WBC^{+}00]. Aaa [Val00, ASK^{+}03, TR00]. AAK1 [CS02, RCS^{+}02]. Abelson [GLJP01]. Aberrant [FSCF^{+}03, HYMS^{+}02, XB04, AKK^{+}00]. Abi1 [IFP^{+}03]. Abl [WMJ^{+}04, MWMK04, WLO^{+}02]. Abl-related [MWMK04]. Ablation [HDL^{+}00, ZCH^{+}01, EKdM^{+}04, FHL^{+}03, KPA^{+}03, MKR01, SGK^{+}02b, DEG00, RBM^{+}00]. able [AvdWM^{+}01]. Abnormal [XRH^{+}03, CBG^{+}01, VMS^{+}02, HGK^{+}01]. abnormalities [DCC^{+}02, YIS^{+}03, LWS^{+}00]. Abolishes [ZBB^{+}00]. aborted [FSCF^{+}03]. abound [LeB04-98]. Abp1 [KEGDQ01]. Abp1p [GRBD01, OM01]. abrogates [EPH^{+}03]. Absence [AKK^{+}00, LKLD04, BGBG03, SPO^{+}02, WLR01, EHZ^{+}01, MCG^{+}00, SKJ^{+}00]. absolutely [VBH^{+}02]. absorber [Dov02v]. ABU [LeB02a]. abundance [HvdBP^{+}01]. abundant [HBV^{+}01]. ACA [JBK04]. Acaps [JBN^{+}00]. accelerate [ERS^{+}04]. accelerated [DH02]. accelerates [FLWMG02]. Acceptor [CB00]. access [Mel04b]. accessory [VLL^{+}03]. accompanied [CBG^{+}01]. Accompanies [KCL^{+}00]. Accumulate [BHBJ00]. accumulates [YRC^{+}04]. Accumulation [HSMB02, MBS^{+}01, OPZ^{+}01, PTM^{+}01, LL02, LKLD04, MSY^{+}04, SKM^{+}02, SLD^{+}02, HK00, MPSM00a, SHS^{+}00, TWS^{+}00a]. Ace2p [WKZ^{+}02]. acetylation [KLT^{+}03]. Acetylcholine [DLXP00, JLJD03, MWF02, SHW01, BL01, BF01, JCR^{+}01, KWO^{+}00, WYHP00]. acetylcholinesterase [CSG^{+}04]. acetyltransferase [CYC^{+}04]. Achr [BL01]. Acid [BPMG00, CK03, Dov02a, Dov02-66, HvdHG^{+}03, JBA^{+}01, MWC^{+}02, SAS^{+}02, SCD02, UGKT^{+}02, WK02, HLK01, OKSH00, vRTvdB^{+}00]. acidic [GGT^{+}02, WMA^{+}04a, WMT^{+}01]. acids [LeB02-71, TWS^{+}00b]. acinar [DWB03, MMG^{+}04, PRJK01]. Acquisition [CSJ00]. acrosome [Dov03-41, FYI^{+}03]. across [FBH03, SKN^{+}03, SEP^{+}01, Vin04, Wel04-61]. act [BHW02c, CC02, DBL^{+}02, Dov02-53, Dov03g, FGR^{+}04, GGGK03, KWS^{+}02, SSM^{+}04]. ActA [SPW00]. Actin [CEGZ^{+}04, DLXP00, DT00, Dov04a, EKT^{+}00, GOL^{+}01, GYL02, GRBD01, HP00, HMN^{+}00, KEGDQ01, KSK^{+}00, LeB02b, LBP00, PZP^{+}01, QK00, RRM^{+}03, SPW00, Sre04a, TRC^{+}00, TCV^{+}00, Tum04a, WKS^{+}00, Wel04a, ZBB^{+}00, BIC^{+}03, BFG^{+}04, BET^{+}03, BGW^{+}04, CRP^{+}04, CG03, CFC^{+}01, CTE^{+}04, DSH^{+}03, DPB03, Dov02b, Dov02-61, DOB^{+}01, EWTW02, EGWK^{+}01, FHM^{+}03, FCF^{+}01, GOV^{+}02, GOV^{+}03, GAT^{+}03, GCT02, GGNK04, GCT^{+}04, GFGP03, GPZ^{+}02, GCR^{+}03, HZS^{+}01, HKP^{+}04, HGC02, HvdBP^{+}01, HGP^{+}04, HC02, HPG^{+}02, HBD^{+}02, HKBH03, IUK04, ISB^{+}04, KHC02, KYW^{+}04, KHLW02, KKTP03, KB04, KGT^{+}02, LCM02, LeB02o, LeB02-36, LeB03-91, LRB^{+}03, LZS^{+}03, LP04b, MSJ^{+}04, MSA^{+}03, MWMK04, NYT^{+}03, OM01, OO02, RWSV03, RJA^{+}03, RSD^{+}04, SAWS02, SRK^{+}03, SKF02, SPB^{+}02, STA^{+}01, SKGC^{+}03, SSO^{+}03, SMW^{+}03, SMZ^{+}03, SAG^{+}01, Sre04b, TT04, VBR^{+}01b, WVY^{+}01]. actin [WXD^{+}03, WSF^{+}01, Wel01e, Wel01-56, Wel02e, Wel02-44, Wel03q, Wel04d, Wel04-27, WHD^{+}03, WBWS03, WWGK02, WLO^{+}02, YSK^{+}04, YCB04, ZWSC02, CBZ^{+}00, FNK^{+}00, FLX^{+}00, FWY01, GOL^{+}01, GHS00, GKM^{+}01, HHSV00, JPM^{+}00a, KYF00, KEGDQ01, LFT^{+}00, LSSL00, MKK^{+}00a, NT00, NLBK00, OSN^{+}00, OSMF00, Ono01, QKK00, RPTNM01, RS00b, RRB^{+}01, RPE00, TTS00, TN00b, VDMH01, WSDW^{+}00, Wel01a, WBM^{+}00, YHF^{+}01, YHD^{+}00, qZC01, Zig00, POW^{+}01a]. actin-based [RWSV03, SRK^{+}03, WHD^{+}03]. actin-binding [EGWK^{+}01, GOV^{+}02]. actin-bundling [WSF^{+}01]. Actin-Dependent [TRC^{+}00, HPG^{+}02, WVY^{+}01]. Actin-Driven [DLXP00]. Acting [RS00b, MMM^{+}04]. actinin [YSK^{+}04, LWDH01, TTS00]. action [CLB^{+}03, SMH^{+}02, Wel03s, YF00]. activate [GCH03, VCDHD03, TM00].

Activated [BMD^{+}00, DSSY00, KY00, LTD^{+}01, NGKH02, RRL^{+}00, SNL^{+}00, WFT^{+}01, CFC^{+}01, CSM03, DLPB03, DWM03, GRSL^{+}04b, JOF^{+}02, LXL04a, MZ00, MDT^{+}01, NMH^{+}04, NCGD^{+}03, PBB^{+}04, PBP^{+}01, PSD^{+}04b, SSRX04, WVBY^{+}03, XWL03, YK03a, ZLVS02, BGFJ01, CDEM00, CAP00, DZT^{+}00, FFKC00, LR00, MMG^{+}01, SPC00, CSIK03, FNKH02]. activates [BSD^{+}01, BPPFM^{+}03, IFP^{+}03, LP04b, MWE^{+}03, NH03b, RLZ^{+}03, SSP^{+}03, Wel03-30, ZLR^{+}03, LBWH^{+}00, MTV^{+}00, OKSH00, RHM00, RM00, Zwe00]. Activating [CGY^{+}01b, Dov02b, HGC00, GMY^{+}03, KKW^{+}03, LJK^{+}01, LXL^{+}04b, HOvD^{+}00, JBN^{+}00, MBSB00, WBC^{+}00, GQI^{+}02]. Activation [ALWR01, BLPP01, CK00, CDWB01, ESS^{+}00, FHJW^{+}01, FSK^{+}04, FSM^{+}01, GJB^{+}00, GRBD01, GPLS02, HP00, HEN^{+}01, JZ02, KY00, PMBC^{+}00, PTM^{+}01, SC01a, WFF^{+}01, Wel03a, WWGK02, ACMR04, AKH^{+}04, BIC^{+}03, BWA^{+}04, BSR^{+}03, CDW^{+}03, CWG^{+}02, CK02, DWB03, DH02, DRC^{+}02, DMLK04, DWM03, FGR^{+}04, GWL03, GJS^{+}03a, HdEV^{+}02, HLB^{+}02, ITF^{+}02, ISID^{+}03, JFS^{+}03, KAC^{+}04, Kel03, KWS^{+}02, KWSK^{+}04, KSR^{+}04, KFO04, LMGM^{+}02, LJK^{+}01, LG02b, MTW^{+}02, MLZ^{+}01, MTM^{+}03, MNC^{+}03, MCF^{+}02, MFF^{+}03, NMH^{+}04, NOOG^{+}04, OSB04, PKR^{+}02, PAR^{+}04, PSK^{+}03, RBE^{+}02, RLTC^{+}02, RDS02, RDNB02, STA^{+}01, SKN^{+}03, SDEZ^{+}03, SZvBuH^{+}04b, SKO04, SXD^{+}03, TNMM03, TNM^{+}03, TFAM^{+}04, TIO^{+}02, VCGB^{+}02, Wel03-48, WDL^{+}04, XRI^{+}04, YPN^{+}04, BKD^{+}00, CDEM00, FYHH00, GV00, HGB^{+}00, HVM^{+}00, KSN^{+}01, KWH^{+}00, MHE^{+}00, MKK^{+}00b, PSKK^{+}00, RyHHK00, SFSD00, TP01, WMS00, WGF^{+}00]. Activation [ZN01]. Activation/Phosphorylation [GJB^{+}00]. Activator [AKDS00, WSL^{+}00, BBDK^{+}04, BND^{+}02, CACL03, MSS^{+}01, BMG^{+}01, KH01, WTG01]. Activator-Mediated [AKDS00]. activators [JRL^{+}03]. Active [BDKM04, LeB02c, MMDC00, NSLSK02, WBU03, vMZM^{+}00, BPD^{+}04, Dov03y, JKG^{+}02, LMHJ02, LeB04-89, MPG^{+}02, MAG^{+}04, TRMI^{+}04, WSS^{+}04, AFN00, BKD^{+}00, FMP^{+}00]. activities [CKL^{+}03, GK04, HF03, MSN^{+}02, SDL^{+}03, WSL^{+}03, RSG00, UAZG00]. Activity [DFYL00, ESS^{+}00, HES00, HGP^{+}00, HPE^{+}01, LCRS01, RRM^{+}00, RGG00, SHN^{+}01, SZZ^{+}00, YOK^{+}03, BBR04, BJM^{+}02, CBL^{+}02, DFZ^{+}03, GMRYM^{+}02, GAT^{+}03, GGT^{+}02, HBD^{+}02, KSC^{+}04, KvHB^{+}01, LRF^{+}02a, LBS^{+}02, MHS01, MWMK04, MNT^{+}03, MHNSM03, MRT^{+}01, MB03c, OTY^{+}04, OWW02, POW^{+}01a, PHS^{+}03, PL01, SKS^{+}04, STA^{+}01, ST03, SHE^{+}02, SSRX04, SEW^{+}01, SF01, TYA^{+}02, TSY^{+}02, Tum04q, UIY^{+}01, UTH^{+}02, UGKT^{+}02, WKYC02, WCBC04, WBT^{+}03, WAOC^{+}03, WLPD04, ZCW^{+}03, BL01, CMMP00, DCM00, FHP00, GMZ^{+}00, Gum00, HLU00, PMK^{+}00, PVjL^{+}00, PCC^{+}00, RMMC01, RM00, SDDS00, SBC01, ZEtK^{+}00]. Activity- [DFYL00]. Activity-Dependent [RGG00, LCRS01]. activity-induced [OTY^{+}04]. actomyosin [GYL02, MB01, WSDW^{+}00]. Actopaxin [NT00, CBLT04]. acts [CGF^{+}04, Dov04e, GCG^{+}01, MP04, TMG03, WSWM04b, WKZ^{+}02, vWJK^{+}03, FWM^{+}01]. Acute [pHYpXL00, TEB^{+}03, MRM^{+}00]. Acyl [TNW^{+}02, BFC^{+}02]. Acyl-CoA [TNW^{+}02]. ADAM [SGC^{+}02, IAG^{+}00]. ADAM10 [YSW02, SWK^{+}04]. ADAM17 [SWK^{+}04]. adapter [GBY^{+}03, MKM04, RML^{+}02, HKJ00]. Adaptin [HLB^{+}00, RWH02]. Adapting [Dov02c]. Adaptor [SEM^{+}00, CS02, FPM^{+}03, JFS^{+}03, NOM^{+}04, POH^{+}04, TONN02, ZWAH03, FPSM01]. adaptor-associated [CS02]. ADAR1 [DJ03]. additional [GI02]. Addresses [LGK00]. Adducin [BIC^{+}03]. Adenine [BPMG00, SNF^{+}02]. adenocarcinoma [JZ02]. Adenomatous [MKST00, MTM^{+}02, RSG01]. Adenosine [DAC00]. Adenosylmethionine [PSKK^{+}00]. Adenovirus [KSK01, LCGR00, MBH^{+}02, Wel02b, RMC^{+}02]. adenylyl [CMS^{+}02, ZFH^{+}04]. ADF [AOJ^{+}04, GOV^{+}03, OO02, ZBB^{+}00]. ADF/ [GOV^{+}03]. ADF/cofilin [AOJ^{+}04, OO02, ZBB^{+}00]. ADF/cofilin-dependent [OO02]. Adherens [HP04, MCA^{+}03, Wel04-32, HW00b]. adherent [BYMS^{+}02, SPB^{+}01]. adhering [KVM03]. Adhesins [RVB^{+}01]. Adhesion [ASAJ01, DSSY00, FSGDN^{+}00, KRA^{+}01, RSBE00, SHN^{+}01, WG03, AMBW04, AJ01, BWA^{+}04, BSEB04, CGM^{+}02, CFWH^{+}01, CWA^{+}03, CPG^{+}03, CTE^{+}04, EPH^{+}02, ELNA^{+}03, FSTC02, FSK^{+}04, FKH^{+}04, GSP^{+}02b, GGNK04, HV03, HEW^{+}01, HKO03, ISF^{+}01, ICK^{+}04, KKK^{+}02, LeB04-76, LSMS03, LPL^{+}04, MGAL^{+}01, MWMK04, MPAP^{+}03, MCA^{+}03, NG02, OPP^{+}03, OMiKF02, REK^{+}03, SLB^{+}01, SPB^{+}02, SSO^{+}03, SLD^{+}02, TWS^{+}04, TSY^{+}02, XSK^{+}01, YWW^{+}04, YMK^{+}04, ZTK^{+}03, AKFB00, CP01, FSK^{+}00, GPAS^{+}01, GWG01, GTPMU00, HZ00, IAG^{+}00, KHH^{+}01, NT00, RRK^{+}00, RMG^{+}00, SMSF00, SMSM00, TNM^{+}00, TGMC^{+}00, TAD^{+}00, TNK^{+}00, THK^{+}00, THZ^{+}01, ZGB01, ZN01]. Adhesion-independent [WG03]. adhesions [BFG^{+}04, DSB^{+}02, GHS^{+}03, KAK^{+}03, BDK^{+}01, BKD^{+}00, LWDH01, PO00, TYY^{+}00]. adhesive [GG04, HPG^{+}02, Gum00]. adhesiveness [KCY^{+}04]. Adipocyte [CSD00]. Adipogenesis [ASMW01]. adjacent [ZAE^{+}04]. adjustor [KOS^{+}04]. adopt [ZGN^{+}04]. ADP [SGdM^{+}01, BLC00, HVM^{+}00, BKZ^{+}03, MZ00, NCGD^{+}03, POW^{+}01a, VCGB^{+}02]. ADP-actin [POW^{+}01a]. ADP-ribose [SGdM^{+}01, BKZ^{+}03]. ADP-ribosylation [VCGB^{+}02]. adrenergic [PLC^{+}02, REK^{+}03]. adult [DDV^{+}03, HViV^{+}02, HAK^{+}04, KFR^{+}04, LCRS01, OMB^{+}01, Pro03, SGK^{+}02b, Tum04o, BHY^{+}00]. Advance [KSN^{+}01]. Advanced [FXPT00]. aequorin [MPV^{+}01]. affect [CSJ03a, RGM^{+}02, vdFKK^{+}02, NBWB^{+}00]. Affecting [GJB^{+}00, GPZ^{+}02, PGV^{+}00]. Affects [ACE^{+}01, XHG^{+}00, KLE^{+}02, SM03a, SAP03, EMW^{+}01, LGM^{+}01, vRTvdB^{+}00]. affinity [KCY^{+}04, RCS^{+}02, SHF^{+}03, BEG01]. Affixin [YSK^{+}04, YSS^{+}01]. African [HDH02]. After [AL03, AKDS00, DB00, DWM03, FP02, HBB^{+}02, LMG04, PSW^{+}02, SGK^{+}02b, Stu04, US04, WGvA^{+}01, vARP^{+}00]. again [Wel04-37]. against [AKDS00, BHL^{+}03, NDM^{+}03]. age [LeB03-54, ZCW^{+}03]. age-dependent [ZCW^{+}03]. Age2 [PNSJ01]. agent [LeB02-32]. Aggregate [BMS^{+}00b]. aggregated [TNM^{+}03]. aggregates [MBN^{+}01, RBD^{+}01, BL01, CLT^{+}01, SSW^{+}01b]. Aggregation [DDW^{+}01, MZH^{+}02, NGK^{+}03, MMG^{+}01, THK^{+}00]. aggresome [LFM^{+}04]. aggresome-like [LFM^{+}04]. Aggresomes [HWW01]. aging [Bir04e, EKdM^{+}04, FBV^{+}04, GCT^{+}04, HJL^{+}04, Wel03c]. Agrin [BF01, BSS00, Dov03a, GRCF02, WYHP00, HHS03, BHFL01, BL01]. Agrin-Induced [BF01, WYHP00]. AHNAK [BGA^{+}04, SSOS01]. aid [BPKK01]. AIF [WLH^{+}04]. Aim [KMiM^{+}01]. Aim-1 [KMiM^{+}01]. AIP [YTM03]. AIR [RBW^{+}02]. AIR-2 [RBW^{+}02]. Akap [GDRS01]. Akap149 [SMTC00]. AKAP79 [OGD03]. Akap95 [SCLC00]. Akr1p [RFCD02]. Akt [DWB03, ZLMP00, FAF^{+}04, FGSW03, LCT^{+}04, Man04a, OMB^{+}01, WGF^{+}00, BFC^{+}02]. ALB3 [MGMH03]. albicans [WBAS04]. Aldrich [BYLA^{+}01, FSM^{+}01, HP00]. Algal [Wel01b]. alignment [DJT^{+}03, MSL^{+}02, TBRG01, AMEC01, YMM^{+}00, dEP01]. aligns [XB04]. alive [Les01m]. alkaline [CGF^{+}04]. Alkalinization [SSA^{+}00]. Allele [ALWR01, YHZ^{+}01]. allele-specific [YHZ^{+}01]. alleles [EAD^{+}02]. allows [PG02]. alone [Wel04o]. along [ABOS^{+}02, BEG01, HHF^{+}00, HKBH03, MKST00, Wel03-45, Wel04-52, XB04]. alphoid [iONOM02]. also [LeB03-39, LeB04-64, WBAS04, PWU00]. Alteration [SGW^{+}02]. Alterations [EHZ^{+}01, PTH^{+}04]. Altered [BRM^{+}00, SWB03, GJ00, PHWK^{+}00]. Alternate [KLF^{+}00]. alternately [RPNM03]. Alternating [CSJ03a]. Alternative [HBH^{+}04, vdHvODML^{+}00, AvdWM^{+}01, LHR04]. Alters [vdHvODML^{+}00, SBS02a, TRW^{+}00]. altruistic [FBV^{+}04]. alveolar [HDP^{+}01, MSS^{+}01]. Aly [GI02]. Alzheimer [ABRA03, CLM^{+}03, EKH^{+}03, LeB04r, LeB04s, SIBG01, SYW^{+}03]. Amassin [HV03]. ameliorates [WST01]. Ameloblastin [FKH^{+}04, LeB04a]. ameloblasts [FKH^{+}04]. amidation [LeB02d]. AMIGO [KPKY^{+}03]. AMIGOs [Dov03-52]. amino [CK03, SAS^{+}02, BSS00, HLK01, WHS00]. among [FFY^{+}00, THK^{+}00]. AMP [REK^{+}03]. AMPA [Dov03-35, TCK^{+}03]. amphiphysin [TSL^{+}03]. amplification [MWE^{+}03]. Amplified [BMS^{+}00b]. amplifier [Dov01f]. Amputation [WXQ^{+}00]. AMSH [MCU04]. Amylin [DDL^{+}04]. Amyloid [KB03, SKF^{+}01, ARLC^{+}04, ABRA03, CKS02, CBC^{+}01, EKH^{+}03, KH04, LeB02-71, SYW^{+}03, Tum03m, Wel03-48, ZMGL02, LXL^{+}04b, SIBG01]. Amyloid- [LXL^{+}04b]. Amyloidogenic [DDW^{+}01, EKH^{+}03]. amyloids [LeB03-74]. anabolic [AOH^{+}02]. analogues [BSW^{+}04]. Analysis [LS02, LRF^{+}02b, WS00, BBBS04, CKZ^{+}02, KSM^{+}01, LDP02, LL02, LGM^{+}04, LVD^{+}04, RDP03, SKGC^{+}03, XWL03, ZJM^{+}02, GKSR00, Lit00, OSMF00, ODR^{+}01, PGSL00, SWJ^{+}00]. Anaphase [GKG^{+}01, DJT^{+}03, GNH^{+}04, HPE^{+}01, KSK01, LP04a, MRC^{+}02, SHP01, XB04, GSW^{+}00, JSCR01, RM00, SHHH01, dEP01]. Anaphase-Promoting [GKG^{+}01, HPE^{+}01, KSK01, MRC^{+}02]. Anatomical [FLX^{+}00]. anchor [Wel04c, KSF^{+}00, SDDS00]. anchorage [DKA01, ZLR^{+}03, ZWB04, GHC01]. anchorage-independent [ZLR^{+}03]. Anchored [FSGDN^{+}00, IWG^{+}01, Wel03b, BCP03, SOH^{+}04, SKT^{+}03, SWE^{+}03, TSL04, PVL^{+}00, MPBR03, PSP^{+}04]. Anchoring [SCLC00, ASS02, CSG^{+}04, DB02, HTT^{+}02, LCM02, Wel04-48, BF01, GDRS01, SMTC00]. Anchors [PS00]. and/or [PBB^{+}04, TFAM^{+}04]. Androgen [CLB^{+}03, BMS^{+}00b, YHK^{+}00]. Androgen-stimulated [CLB^{+}03]. aneuploidy [WBP^{+}03, CHM^{+}01]. angiogenesis [HFK^{+}03, MTB^{+}02, Sre04b, XJW^{+}04, XRP^{+}01, MCB00b]. angiogenic [CPA^{+}03, GGF^{+}03]. Angiomotin [TLM^{+}01]. angiotensin [HBG^{+}02]. angles [ALEH02, LeB04z]. anillin [TMG03, OSMF00]. animal [MRK04]. anion [RPS^{+}02, SMS^{+}01b]. Ankyrin [Dov01a, GSB^{+}03, JB01, BBG^{+}03, BZSC00, KS02, MKR01, NYT^{+}03]. ankyrin-1 [BBG^{+}03]. Ankyrin-G [JB01]. Annexin [TFM04, BGA^{+}04, NOS^{+}01, WK02, PLL^{+}00]. annexin-mediated [WK02]. Annexins [BD00]. anoikis [JW04, VMK^{+}03]. antagonism [SDD04]. antagonist [SDS^{+}04]. Antagonistic [Gla01a]. Antagonize [GJB^{+}00, TIO^{+}02, WM03]. antagonizes [CKL^{+}03]. Antagonizing [LeB04c]. Antenna [RBB00]. Anterior [HVB^{+}00]. anterograde [MMPO^{+}01, PPR^{+}00]. Anthrax [ALC^{+}03, Dov03b, Kur03, ALP^{+}04]. Anti [Sre04b, SHE^{+}02]. anti-actin [Sre04b]. Anti-angiogenesis [Sre04b]. anti-apoptotic [SHE^{+}02]. antiapoptotic [CKL^{+}03]. antibody [LeB03-99]. Anticancer [LeB02e]. antideath [LeB02-32]. Antifibrillogenic [DDW^{+}01]. antigen [ALP^{+}04, KRS^{+}01, Les01m, MGL^{+}00, SDDS00]. Antigens [LeB04d]. Antioxidant [TSMS01]. antiparallel [Néd02]. antiviral [LeB03-52]. Any [LeB04e]. AP [AFK^{+}03, CS03b, Dov03c, FPM^{+}03, GK03a, JKB^{+}03, LCT^{+}04, MBSR03, NOM^{+}04, PWY^{+}03, PRLR02, POH^{+}04, PLC^{+}02, RWK^{+}04, FPSM01, GJB^{+}00]. AP-1 [GK03a, JKB^{+}03, LCT^{+}04, RWK^{+}04, FPSM01, GJB^{+}00]. AP-1A [FPM^{+}03]. AP-1B [FPM^{+}03]. AP-1B-dependent [AFK^{+}03]. AP-2 [CS03b, Dov03c, MBSR03, PWY^{+}03, PLC^{+}02, KKW^{+}03]. AP-3 [JKB^{+}03, PRLR02, POH^{+}04]. AP-3B [NOM^{+}04]. AP2 [JFS^{+}03, JK01, RCS^{+}02]. Apaf [HOvD^{+}00, ERS^{+}04, RBE^{+}02, SSM^{+}04]. Apaf-1 [HOvD^{+}00, ERS^{+}04, RBE^{+}02, SSM^{+}04]. APAF1 [FCM^{+}01]. apart [LeB04j, Wel01d, Wel04-67, BGFJ01]. APC [MKST00, FFY^{+}00, GK03b, HdEV^{+}02, HPE^{+}01, LeB03a, MKT01, SHP01, SCY01, Wel02-45]. APC-dependent [SHP01]. APC/C [HdEV^{+}02, SCY01]. Apc8 [MRC^{+}02]. apCAM [SF01]. Apg1 [KFS^{+}00]. Apg5 [MYH^{+}01]. Apg5-Deficient [MYH^{+}01]. Apg8 [KIO^{+}00]. Apg8/Aut7 [KIO^{+}00]. Apg9p [NKH^{+}00]. Apg9p/ [NKH^{+}00]. Apical [Dov01c, HGP^{+}00, PLL^{+}00, RSG01, TSMT00, MBMMA^{+}03, MWK^{+}02, NOS^{+}01, PSP^{+}04, TNBH01, Wel03-62, HW00b, TCS01, VHLS00]. aPKC [IOIF^{+}04, RAS^{+}03]. apoptose [Wel04u]. Apoptosis [CGY^{+}01b, CK00, CFM^{+}02, ESH^{+}01, Gre00, KY00, LeB03b, RRM^{+}00, SEM^{+}00, SPB^{+}01, APM^{+}02, BLU^{+}04, BMS^{+}03, BSW^{+}04, BPKR^{+}02, CSL^{+}03, CLM^{+}03, CNMS02, DC03, DMLK04, DWM03, ERS^{+}04, GLDM01, GCH03, HJL^{+}04, HKE^{+}04, KLG^{+}02, KAC^{+}04, LLP^{+}02, LeB02d, LeB04-111, LGRP^{+}02, LSS^{+}02, MEV^{+}04, MO01, MWE^{+}03, MTPT02, PPP^{+}01, PSK^{+}03, PHM^{+}02, RDP03, RKR^{+}03, RGG03, SISO01, SMS^{+}01a, Tum03j, WKYC02, Wel01-60, Wel03h, WLH^{+}04, WLPD04, YSC^{+}02, ZRDG02, ZLH^{+}03, AV01, Bur01, CWMO00, CGL^{+}01, CA00, FWM^{+}01, GLK^{+}00, HOvD^{+}00, IOLA^{+}00, LCGR00, LSA^{+}00, MMDC00, MMR^{+}00, MBSB00, NBWB^{+}00, NSL^{+}01, RMMC01, SME^{+}00, SMS^{+}01b, TRW^{+}00, WGF^{+}00, ZLMP00, YSC^{+}21]. Apoptosis-Associated [RRM^{+}00]. Apoptosis-inducing [CFM^{+}02, APM^{+}02]. apoptosome [WLPD04]. Apoptotic [Dov04b, GMRS00, HOvD^{+}00, LVWA01, PBB^{+}04, RBD^{+}01, WGvA^{+}01, AAM^{+}04, DPO^{+}04, HdO^{+}01, HBC^{+}03, LXL04a, SKS^{+}04, SHE^{+}02, HLU00, PDJ00]. APP [SIBG01, CLM^{+}03, KLE^{+}02, KH04, Lai03a, LKL^{+}03, SIBG01, SVT^{+}02]. APP-BP1 [CLM^{+}03]. APP-induced [CLM^{+}03]. Apparatus [HMN^{+}00, AZ03, BPKK01, BS02, CNMS02, CSM03, FR01, KMLS04, MMFS01, Pel01, DOL^{+}01, HMRH01, LHvdH00, PPM^{+}00, SSN01, vdB00]. appears [MCBB^{+}04]. appetite [LeB02r]. APPL [LeB04-30]. apposition [DW02]. approach [Wel04-38]. AQP4 [CPN^{+}01]. AQP4-expressing [CPN^{+}01]. aquaporin [AMF01, KBK^{+}03, KHvOD00]. aquaporin-2 [KBK^{+}03, KHvOD00]. aquaporin-4 [AMF01]. Arabidopsis [ARK^{+}00, GPL^{+}02]. Architectural [RBB00]. Architecture [WCA^{+}03, DWB03, Mis01, RCY^{+}03, RSD^{+}04, Tum03e, WDL^{+}04, HF01]. arcs [SKF02]. areas [LFM^{+}04]. ARF [VCGB^{+}02, RAD^{+}02, VCGB^{+}02]. ARF-GAP-mediated [RAD^{+}02]. Arf1p [GTBM04]. ARF6 [AM03, KKW^{+}03, PDR^{+}03, SC01a, BRY^{+}01, JBN^{+}00]. Arf6-regulated [BRY^{+}01]. ArfGAP [PNSJ01]. ARFGAP1 [YLG^{+}02]. Arg [MWMK04, Wel04d]. arginine [BBDK^{+}04, MC02, YF00]. argosomes [Wel01-66]. arising [MHK04]. ARK [ZRDG02]. Arl2 [BLC00]. arm [ATG^{+}03, ALEH02, BVH04, Dov04o, KRMB03, LC01, SCB02, GPAS^{+}01, PS00]. Armadillo [CdBB^{+}01, HCD^{+}00]. Arms [Wel01d, LeB04-100, TCH^{+}02]. ARNO [Dov01d, SC01a]. Arp [Gla01b]. Arp2 [BCM04, DBB02b, Dov02f, Dov03-31, EKT^{+}00, FCF^{+}01, FSM^{+}01, GRBD01, HP00, HC02, HBD^{+}02, JRW^{+}01, KSK^{+}00, LeB04f, LJK^{+}01, LBP00, Mil02, SPW00, SAG^{+}01, WKS^{+}00, ZCH^{+}02]. Arp2/ [Dov02f, LeB04f, ZCH^{+}02]. Arp2/3 [BCM04, DBB02b, Dov03-31, EKT^{+}00, FCF^{+}01, FSM^{+}01, GRBD01, HP00, HC02, HBD^{+}02, JRW^{+}01, KSK^{+}00, LJK^{+}01, LBP00, Mil02, SPW00, SAG^{+}01, WKS^{+}00]. Arp2/3-activating [LJK^{+}01]. Arpless [LeB04g]. arrangements [CKW^{+}03, ZAE^{+}04]. array [RTFW02]. arrays [CPN^{+}01, MTM^{+}02, NSLSK02, TTS00]. Arrest [LRB^{+}03, ABRA03, DLPB03, FHJW^{+}01, FP02, FBG^{+}01, KSK01, LRS^{+}02, RP03b, SRL^{+}04, SM03b, THG^{+}04, TEC^{+}03, WFI^{+}04, WK04, DTO^{+}01, GBM^{+}00]. Arrested [LeB02f, KRS^{+}02]. Arrestin [DZT^{+}00, PLC^{+}02]. arrestin3 [SGK02a]. Arresting [Dov02g]. Arteries [LeB02g, LeB03-45]. Articular [YCX^{+}01]. ARVCF [FJK^{+}04]. Arx [NMG04]. Ase1p [SLP03]. Asf [KLF^{+}00]. ASH1 [LGM^{+}01]. Ash1p [LGM^{+}01]. Ashen [HWHH01]. ASK1 [CKL^{+}03]. Asn2 [PHWK^{+}00]. Asp [WBG01]. aspartate [UTH^{+}02]. Aspects [NSW00]. Aspergillus [EM00]. assay [WHD^{+}03]. assemble [ECK^{+}03, MB01]. assembles [PMKM03, TCZ^{+}03]. assemblies [BHK^{+}02, BMM^{+}01, HK00]. Assembling [LeB04h]. Assembly [CFC^{+}00, DM02, EKT^{+}00, GKG^{+}01, GDHS01, HMN^{+}00, LBP00, ONS^{+}00, PRLR02, SGYD^{+}01, SKT^{+}00, WKS^{+}00, AR03, BDKM04, BKD^{+}04, BGBG03, CRP^{+}04, CMM^{+}02, CK02, DK04, DC03, DLT^{+}02, DPB03, Dov01-33, DWFAdL^{+}02, ERMT^{+}04, EGWK^{+}01, FJM^{+}04, FSCF^{+}03, GC04, GBZ^{+}02, GV04, HOK^{+}02, HCL^{+}03, HMVG02, HG03, HF03, HGP^{+}04, IKS^{+}02, IYT^{+}04, JB01, JGR^{+}04, KLZ04, KYM04, LMG04, LMVW03, LHC^{+}02, LSMS^{+}01, MRK04, MDF01, MLKH04, MC02, MSN^{+}02, iONOM02, PBD^{+}02, PG02, PBT^{+}02, RRSV02, RFLT02, RMW03, SERB03, SKF02, SPB^{+}02, SRC^{+}01, SBS^{+}02b, SMC^{+}02, SRKN03, SN04, SML^{+}04, TDL03, TLS^{+}01, THG^{+}04, UJK^{+}02, UJL^{+}03, WMCW03, WDMH03, YKT^{+}04, YEG01, YK03b, AFB^{+}01, BKDH01, CKS^{+}00, GKSR00, HHSV00, HWW01, KMCM00, LCS^{+}01, ODR^{+}01, Ono01, PCK^{+}00, PDV^{+}00, RBV00, SCM^{+}00, SMS00, SLSR^{+}00, SMTC00, TGMC^{+}00, TCR00]. Assembly [TYY^{+}00, VDMH01, WPJ^{+}00, WHS00, ZKW^{+}00]. assistance [MGP^{+}02a]. associate [RWCC01, WSS^{+}04]. Associated [FSGDN^{+}00, KLK^{+}01, KSK^{+}00, RRM^{+}00, SCLC00, BPPFM^{+}03, BSL^{+}01, CS02, DBB^{+}02a, DLK^{+}02, EPH^{+}03, FBH03, FTD^{+}01, GSP^{+}02a, GK03b, GSN^{+}04, HBAF^{+}02, KNR^{+}04, iKFH^{+}04, MDP02, MCU04, MPV^{+}01, PFB^{+}03, PMU^{+}02, RRB^{+}03, SWBE^{+}04, SKK^{+}04, THO^{+}04, TIS^{+}01, TVF^{+}03, WBP^{+}03, WHP^{+}02, YHT02, vBDH03, CFC^{+}00, EHZ^{+}01, FKSJ01, HK00, HBSQ01, IdCAS^{+}00, MAAZ^{+}00, MRM^{+}00, NHI^{+}00, NSL^{+}01, PZP^{+}01, SMLM00, TNM^{+}00, VMD^{+}01, WDLK01, BGM03, CPC^{+}02b, NOS^{+}01, TCZ^{+}03]. Associates [CS01, SHN^{+}01, CiKBG03, EGWK^{+}01, NCMO^{+}02, DSP^{+}01, JWJJ00, PLL^{+}01, SFSD00, SGO^{+}00, WRGK00]. Associating [TYS^{+}00]. Association [FMP^{+}00, KLK^{+}01, LPT03, TGD^{+}03, AKM^{+}04, DAV^{+}03, GBOL03, IIN^{+}01, KLG^{+}02, LSMS03, LYL^{+}04, NB03a, RP03a, SJ01b, TBRG01, MPSM00b, PBL^{+}00, PLL^{+}00, SKT^{+}00, SCS^{+}00, WGP^{+}00]. assurance [LeB02-77]. Asters [PMBC^{+}00, HOK^{+}02]. astic [Jay01]. astral [GMD^{+}02, MHlW02]. astrocyte [SK04]. Astrocytes [Tum04b, RGG00]. Asymmetric [CB00, FYHH00, BMS^{+}03, IOIF^{+}04, Tum04-29, YCK^{+}03, LGM^{+}01, SLT^{+}01]. asymmetrically [LMG04, KPB^{+}00]. asymmetry [Tum03a, ZW04]. Atelp [ARK^{+}00]. Atf [MHK^{+}01]. Atf-2 [MHK^{+}01]. ating [LeB04g]. Atomic [VDMH01]. ATP [CPC^{+}02a, DBLG02, HBH^{+}04, HM00, HSB00, LeB04-108, RDNB02, SHM02, WCIN04, WGvA^{+}01, Wel01-58, vdLBK^{+}04]. ATP-dependent [DBLG02, WCIN04]. ATPase [CKFH00, CRH02, FAT^{+}02, GF03, MHNSM03, DWD^{+}00, SKR^{+}00]. Atpases [TR00]. atrial [SYY^{+}03]. Atrophin [WND^{+}00]. Atrophin-1 [WND^{+}00]. Atrophy [WND^{+}00, MPG^{+}03, RJA^{+}03, TBJ^{+}01]. AtSNAP33 [HGS^{+}01]. Attached [BL01]. attachment [BWV^{+}01, BKD^{+}04, DMH^{+}02, FNZ^{+}03, FHM^{+}03, GES04, HCL^{+}03, HIG^{+}01, LdCK^{+}01, MSL^{+}02, Wel02x, AV01, GZY^{+}00, KHN00, SD00, WFF^{+}01]. attachments [GK03b, HPG^{+}02]. Attack [Wel02c, Wel02d, Dov01r]. attempt [Dov02-57]. attenuates [IIN^{+}01, RXS^{+}03]. Attenuation [SJIM01]. atToc159 [BHW^{+}02a, SHKS02, SRW^{+}04]. attract [CSL^{+}03]. attraction [BHNG01]. attractive [Dov01l, Dov03-37]. attracts [LeB04-64]. Atypical [KMH^{+}04, SYH^{+}01, NCMO^{+}02, WRGK00]. aureus [BSD^{+}01]. aurora [RBW^{+}02, AMEC01, BTVB03, DJT^{+}03, GG01, GMD^{+}02, GNH^{+}04, HKHO01, HCL^{+}03, KSK^{+}02, LeB03c, TUK03, ZSS01]. Aurora-A [HKHO01]. Aut1p [KHK01]. Aut2p [KHK01]. Aut7 [KIO^{+}00]. Aut7p [KHK01]. Autocrine [LLH^{+}01, MWL01, WDL^{+}04, ZLR^{+}03, MGAL^{+}01, CSD00, FRO01]. autoimmune [CKF^{+}03, DPO^{+}04, CdBB^{+}01, SEP^{+}01]. autoimmunity [LeB04-45]. autointegration [STKCW02]. Autonomous [WWK^{+}00, KHB^{+}04, MN03]. autonomously [CC02, MYO^{+}04]. Autophagic [MSF^{+}00, EOB^{+}04, IBS^{+}02, TOM01]. Autophagosome [ADKK00, MYH^{+}01]. Autophagy [KFS^{+}00, KIO^{+}00, HPFG03, Wel02s, WE02, KNIO01, KHK01, NKH^{+}00]. autophosphorylation [ZSY^{+}03]. Autoregulation [CSSBY^{+}03]. Autosomal [CNBWN00]. Autotaxin [UGKT^{+}02, Dov02d]. autotypic [PMU^{+}02]. Auxin [LeB03d]. Average [KSN^{+}01]. Avidity [DSM^{+}03]. awaken [Wel04w]. Awakening [LeB04i]. awareness [LeB03-56]. away [LeB02d]. axial [EOJ^{+}03]. Axin [FFY^{+}00, HSC01]. axis [KVC^{+}03, WBT^{+}03, FFY^{+}00, HVB^{+}00]. Axo [TK00, MDP02, TGMC^{+}00]. Axo-Glial [TK00, MDP02]. Axon [ACE^{+}01, HDJ00, ABF^{+}03, BHNG01, CS03a, Dov03-40, GYL02, KS02, KPKY^{+}03, LGGS^{+}04, LeB03-30, LeB04f, MMBB03, RJA^{+}03, SWBE^{+}04, Wel01-33, BAD^{+}00, MHW^{+}00, Rut00]. Axonal [AKDS00, Bro03, ATG^{+}03, DMBS02, EMY^{+}04, FMF^{+}04, GLS^{+}03, KEHAM^{+}02, KL04, LRBH02, LeB04-68, LDK^{+}03, NH03a, PSW^{+}02, RGM^{+}02, TE01, WWD03, WAC^{+}03, AGB^{+}00, TTHH00]. Axoneme [PS00]. Axonin [FSK^{+}00]. Axonin-1 [FSK^{+}00]. axons [Dov02-63, FPP^{+}02, LeB02s, LeB02-87, LeB02-99, LeB03-28, MTT^{+}04, PSE^{+}03, RCY^{+}03, SWH^{+}02, FSK^{+}00, SHS^{+}00]. Axotomy [FXPT00]. Axs [LeB03-88].

B [iONOM02, BLPP01, BLPP01, RRM^{+}00, AMEC01, AP00, CSIK03, DSN^{+}01, DJT^{+}03, DRG^{+}03, FWM^{+}01, FGSW03, GG01, GNH^{+}04, GPLS02, HCL^{+}03, KSR^{+}04, LeB03c, LeB03r, MWE^{+}03, MHE^{+}00, MKR01, NHS00, PBB^{+}04, PVL^{+}00, RJyH02, SLR^{+}03, SRSW04, SDS^{+}04, SC01b, SSOS01, TST^{+}03, ZLR^{+}03, ZSS01, vdFKK^{+}02, Ern00, GV00, MKK^{+}00b]. B/Akt [FGSW03]. B/C [DSN^{+}01]. B1 [Dov03w, DOL^{+}01, KJY04, MHH^{+}03, MYKG00, SKO04]. B1/ [SKO04]. B2 [DOL^{+}01]. B3 [MCB00b]. baby [Sod02]. BACE1 [SYW^{+}03]. Back [BWRT03, Dov01-35, Dov02-40, LeB03-27, LeB03-62, LeB04o, Wel02b, Wel03v, Wel03-47, Wel04-34, Wel04-75]. Backward [Wel04e, Dov03h, LeB02-103]. backwards [Wel03-49]. Bacteria [LeB02h, Wel04f, GGB^{+}04]. Bacterial [LBWH^{+}00, Wel03c, Wel01-52]. bad [LeB03-103, Wel01-39]. BAF [Dov02h]. BAFfling [Wel02e]. Bak [NSL^{+}01, ZLH^{+}03]. balances [MR01a]. Balancing [GFGP03, Man04a]. Balbiani [PZP^{+}01]. Band [BMM^{+}01]. bands [SJB^{+}03]. BAP31 [BSMS03]. barbed [KKTP03]. Barentsz [vEPP^{+}01, WHAH03]. Barr [OBG^{+}03]. barrel [IYT^{+}04]. Barrier [STKCW02, BD02, CBG^{+}01, FHF^{+}02, JKW^{+}03, MB03b, NHG^{+}03, WCBC04, FL00, SEP^{+}01, Ste00]. Barrier-to-autointegration [STKCW02]. Barriers [Wel03d, Dov03e]. BARS [LeB04j]. basal [LeB04-110, MLKH04, HW00b, SLT^{+}01, TSMT00]. based [COB01, CWG^{+}02, FHF^{+}02, GV03, JKB^{+}03, JB01, RWSV03, SRK^{+}03, STA^{+}01, WHD^{+}03, WWGK02, YMK^{+}04, YOK^{+}03, RPTNM01, TMD^{+}01]. basement [Bir04b, KVM03, SML^{+}04, JCR^{+}01, LCS^{+}01, ME00]. Basic [KY00]. Basis [AB00, SJA^{+}00]. basket [FGS^{+}02]. basolateral [AFK^{+}03, TNBH01]. Bassoon [TRMI^{+}04]. Battle [Wel01e]. Bax [NSL^{+}01, ZLH^{+}03, GLK^{+}00, GMRS00, KLG^{+}02, KAC^{+}04, PPP^{+}01, SKS^{+}04, VMK^{+}03, ZMGL02, vARP^{+}00]. Bax-Mediated [vARP^{+}00]. Baz [IOIF^{+}04]. Bazooka [WRGK00]. Bcl [Dov03d, BLPP01, AAM^{+}04, CA00, CVZ^{+}04, CQH^{+}00, HOvD^{+}00, KSS^{+}03, MEV^{+}04, OMB^{+}01, PFM^{+}00, RRL^{+}00, TGM^{+}01, WAOC^{+}03]. Bcl-10 [TGM^{+}01]. Bcl-2 [AAM^{+}04, CA00, CVZ^{+}04, CQH^{+}00, HOvD^{+}00, KSS^{+}03, OMB^{+}01, PFM^{+}00]. Bcl-2-regulated [MEV^{+}04]. Bcl-w [Dov03d, WAOC^{+}03]. bcl-x [BLPP01, KSS^{+}03]. Bcl10 [GV00]. Bcl10-Induced [GV00]. BDNF [RLTC^{+}02, SNS^{+}04]. BDNF-induced [RLTC^{+}02, SNS^{+}04]. be [LeB02-30, LB03, MCBB^{+}04, PBP^{+}01, Ped04, Wel01t, Wel02-34]. beans [Dov01-36]. become [Dov04f, HAK^{+}04, Wel04-37, BKD^{+}00]. bed [Wel03-32]. Bee1p [EKT^{+}00]. before [CM01, Dov01t, EOJ^{+}03, KPB^{+}00, LeB02-69, PR00, VAHV00, Wel04-45]. Beginning [GKG^{+}01]. behavior [LCS^{+}02, MVL^{+}02, PMK^{+}00]. behind [BHIWH01]. Being [Wel04g]. belonging [PPGN^{+}02]. Bend [MN04, Wel01g]. bent [Gla01b]. Best [LeB04k]. better [Dov03-53]. between [APLB00, BBG^{+}03, BDR^{+}03, CS04, CH01, CBC^{+}01, DB00, DMBS02, EAD^{+}02, GYS02, HLB^{+}00, MLC^{+}01, MRC^{+}02, MKK^{+}00a, NY00, NKP^{+}01, OSN^{+}00, PCR^{+}01, PPM^{+}00, QKK00, RAD^{+}02, RMHM00, RS00c, RPE00, SMI^{+}00, SRL^{+}04, WGP^{+}00, dVKS04]. Beware [Wel01f]. beyond [Dov04d, Mil02, DKJ00]. Bfa1p [PMGS02]. BH3 [WAOC^{+}03]. BH3-only [WAOC^{+}03]. bHLH [PRJK01]. bias [WHB04]. Bicarbonate [ZFH^{+}04]. Bicarbonate-responsive [ZFH^{+}04]. bicoid [Wel01-38]. Bid [vARP^{+}00]. Bid- [vARP^{+}00]. Bidirectional [RH00, DSV^{+}03]. Big [LeB02i, Dov01q, Tum03j]. Bigger [Bir04a]. Biglycan [BMF00]. Bik1 [LdCK^{+}01]. Bilateral [HSB00]. bilayer [FRT^{+}01]. Bilayered [WSW^{+}00]. bim [GCH03]. binary [Dov04h]. Bind [Wel01g, ABOS^{+}02, CM01, SS01]. Binding [BBG^{+}03, BRM^{+}00, Bir04b, ESS^{+}00, FSM^{+}01, GRBD01, KIO^{+}00, KSK^{+}00, LeB03e, LHW^{+}01, PLR03, SHB^{+}03, SBH00, vMZM^{+}00, ACMR04, AOJ^{+}04, BHW02b, BCG03, BM02, CPC^{+}02a, CFWH^{+}01, DCC^{+}02, DSB^{+}02, DMBS02, EBWC01, EGWK^{+}01, FGR^{+}04, GOV^{+}02, GBOL03, GK03a, GSB^{+}03, GSSP03, HZS^{+}01, HDL02, HBD^{+}02, HKO03, JK01, JKG^{+}02, Kil03, KYW^{+}04, KFO04, LKM^{+}04, LeB02-55, LRF^{+}02a, LRA^{+}02, LRF^{+}02b, MKS^{+}02, MGAL^{+}01, MKJ^{+}02, MHNSM03, NG02, REM^{+}02, RCS^{+}02, RWH02, SRC^{+}01, SSGLS01, Tum04e, VMH^{+}02, VT04, VCGB^{+}02, WYR^{+}02, WHAH03, ZTK^{+}03, vBDH03, vdFKK^{+}02, AERD^{+}01, AKFB00, EHM^{+}00, GZY^{+}00, HvdKDS01, JC01, KEGDQ01, KHN00, KH01, LGM^{+}01, MKK^{+}00a, MÖS^{+}00, OSMF00, OKSH00, RSG00, RIDC01, RMMC01, SIBG01, SCMH00, WDLK01, YSS^{+}01, DJ03, GGNK04, WE02]. Binds [RB01, CDK04, CFC^{+}01, CYC^{+}04, DBB^{+}02a, DKAH04, DOB^{+}01, GOV^{+}02, ISF^{+}01, XB04, YTM03, AERD^{+}01, AHMJ01, BMF00, NT00, SJS^{+}00, YSGS00]. bioavailability [DCC^{+}02]. biochemical [SHCM03, CS00]. biochemically [FPM^{+}03]. Biogenesis [GDHS01, JPRR00, KRR^{+}01, BTH^{+}03, FBG^{+}01, GGB^{+}04, HCCB03, LMM^{+}02, RH04, SKM03, Tum04-30, vdLBK^{+}04, ADKK00, CH01, MLC^{+}01, NSW00]. biogenic [YRC^{+}04]. biological [SRG^{+}04, SGF^{+}00]. biology [EA03, Gal02, Mel01, MM03, PAG02, Mel00a]. biomimetic [WHD^{+}03]. Listeria [SBG^{+}04]. biorientated [Wel02z]. biosensor [CWG^{+}02]. biosynthesis [CK03, GPL^{+}02, NHI^{+}00, TWS^{+}00b]. Biosynthetic [AFB^{+}01]. Biotechnology [Wel01h]. BiP [Dov02-35, HJ02, KOS^{+}04]. Biphasic [BTD^{+}00]. biphosphate [MHS01]. bipolar [GC04, KM01]. bipolarity [MSL^{+}02]. Birth [Dov02i, Gla01d]. Bisphosphatase [BMC00]. Bisphosphate [AZB^{+}00, BRM^{+}00, AM03, BND^{+}02, BRY^{+}01, RWH02, BTD^{+}00, RyHHK00, YHF^{+}01]. bistability [MHK04]. Blastoderm [HW00b]. Bld10p [MLKH04]. blebbing [IBS^{+}02]. BLM [LKLD04, SRL^{+}04]. Block [SSL^{+}00, Dov03-28, LeB03i, MP04, MSY^{+}04, PSKK^{+}00]. blocked [HS02, KAC^{+}04, UCY^{+}02]. Blocking [LeB02j, SDDS00]. blocks [BBBS04, DMH^{+}02, DW02, LdVV^{+}02, OBG^{+}03, SVT^{+}02, TBRG01, CKS^{+}00, MBSB00, MSHG00]. Blood [LeB02k, CPA^{+}03, LeB04-85, NHG^{+}03, SEP^{+}01]. blood-brain [NHG^{+}03]. bloody [MF01b]. Bloom [BKI^{+}01, DC03, FP02]. blow [Dov04b]. BMP [FLS^{+}03, KPA^{+}03, KWSK^{+}04, LeB03f, PAR^{+}04, WRCU00]. BMP-4 [KWSK^{+}04]. BMP-7 [PAR^{+}04]. BMPs [RPNM03]. Bni1p [MMM^{+}04]. bodies [BCB^{+}02, FM01, FNKH02, Jeg01d, KRS^{+}01, LBS^{+}02, OL02, PMGS02, SRKN03, SN04, TBJ^{+}01, WSS^{+}04, BHBJ00, PLL^{+}01, SLT^{+}01]. body [BBMS03, CMM^{+}02, DHK^{+}04, HMG03, JBK04, JGW02, JJM^{+}02, Kil03, LeB02-106, LO04, MLKH04, NM02, Tum02b, BKB^{+}01, LCS^{+}01, PGSL00]. boisterous [Wel03e]. bond [LeB04-39, ØTW04]. bonds [DSM^{+}03, GJS^{+}03a, YWW^{+}04]. Bone [DCC^{+}02, SCD02, Wel01i, ZHH^{+}02, DDL^{+}04, Dov01u, Dov04h, JBA^{+}01, LeB02-44, LeB04-34, PPK^{+}01, BMKA01, CKS^{+}00, IHK^{+}00, LQPC^{+}00, MSV^{+}00]. bones [Dov02y, Dov04f, Wel03-57, ECV^{+}00]. Boning [Dov04c]. boost [Dov03n]. border [TM04]. Borealin [GCH^{+}04]. Borg [Wel01f]. Both [MHlW02, CS04, DB02, Dov02-29, Kel03, LKLD04, MEV^{+}04, SC01a, SHE^{+}02, TFF03, WHAH03, YTM03, HSKG00]. botox [Dov03-28]. Botryllus [Wel03-29]. botulinum [DRG^{+}03]. bound [CSG01, EWTW02, HSMB02, PG02, MDJF00, PZP^{+}01, AQHO03]. boundary [KH04, Tum02e]. bouquet [CC02, TSDS00]. box [JBK04, iONOM02, CP01, WAPB^{+}00, BGM03]. boxes [MRC^{+}02, DBS^{+}01]. BP1 [CLM^{+}03]. BPAG1 [LZPL01]. BPAG1n4 [LDK^{+}03]. Brain [DFYL00, ICK^{+}04, KWS^{+}02, Wel04h, BHW^{+}03, CLM^{+}03, Dov03-30, Dov03-52, Jeg01b, LeB03-44, LeB03-94, LeB04-67, LeB04-85, NHG^{+}03, SST^{+}01, VFN^{+}04, Wel01c, Wel02-43, SEP^{+}01]. Brain-Derived [DFYL00, KWS^{+}02]. branch [HS02, Les01o, SAG^{+}01]. branch-formation [SAG^{+}01]. branched [YCB04]. branches [Tum04a]. Branching [LeB02l, LeB03-29, MNC^{+}03, WSL^{+}03]. Brave [Wel02f]. BRCA1 [DC03, YHZ^{+}01]. BRCA1-induced [YHZ^{+}01]. break [Dov02e, Dov02-69, Les01b]. breakage [Wel01i]. breakdown [Dov02-57, LRD^{+}03, LGM^{+}04, SERB03]. Breaking [Dov03e, LeB03h]. breaks [LeB04-47, LKLD04, Wel03w, SCMH00]. breast [LRWB04, PSS^{+}04, Wel01-63, WDS^{+}03, BZSC00, HPQ^{+}00]. breath [LeB04-109]. Breathe [LeB04l]. Breathing [Wor03]. breed [LeB04l]. Brief [Wel00]. bright [Dov03-49]. brightest [LeB02m]. Bring [Wel02g]. Bringing [Jeg01a]. brings [Dov01-29]. Bristles [TCV^{+}00, GCR^{+}03, LeB03-38]. Brn3a [WWD^{+}04]. Bro1 [LO04]. Broad [CDM^{+}02, CS00]. brucei [HHM^{+}04]. brush [TM04]. BTG1 [BBDK^{+}04]. Bub1 [SBC01]. Bub2p [PMGS02]. BUB3 [SCY01, BJB^{+}03, SBC01]. Bub3p [HJSM00]. bubbly [LeB04n]. BUBR1 [SCY01, Che02, DJT^{+}03, LeB03g]. bud [AMG^{+}01, LeB02n, LeB03-80, SMH^{+}02, YITe03]. Bud6p [HBB^{+}04]. Budding [CFB^{+}03, PMSB01, ARQ^{+}04, BTVB03, CNT03, GES04, HGP^{+}04, Kil03, LOC03, SHP01, VT04, WCA^{+}03, YITe03, vBDH03, FK01, HLRG01, JSCR01, OAR^{+}00, PWU00, RHM00, SCM^{+}00]. Bugs [Wel04i, Wel04-70]. Build [LeB02o, Dov03-47]. Building [Dov02j, Hua02, LeB03h, Sre04c, Wel02h, BBBS04, LeB03i, TYS^{+}00]. builds [Dov03-53, LeB03-38, LeB03-98]. buildup [Dov01x]. Built [PN00]. bulk [ZW04]. bullet [Wel01r]. Bullosa [CLWR01]. BuMPy [Gla01a]. bunch [LeB02m]. Bundle [LeB02q, LeB02p, TCV^{+}00, SMW^{+}03, MMH^{+}00, VDMH01]. bundles [GCR^{+}03, LZS^{+}03, MTM^{+}02, VYW^{+}03, ZWSC02, DT00, qZC01]. bundling [BMY^{+}01, MKJ^{+}02, WSF^{+}01]. buries [Dov03d].

C [PSKK^{+}00, WMS00, APM^{+}02, BSMS03, DKJ00, DRC^{+}02, DMLK04, MH01, MEV^{+}04, MNT^{+}03, OSWG02, PKR^{+}02, WMJ^{+}04, YSC^{+}21, vdLBK^{+}04, vARP^{+}00, ACBG04, CdLvM^{+}04, DSN^{+}01, FCL^{+}00, FYI^{+}03, HdEV^{+}02, HLB^{+}02, HRB^{+}01, KSN^{+}01, KMH^{+}04, KMP02b, LTB^{+}00, LRA^{+}02, LPT03, MHT^{+}04, MBLCE03, MSV^{+}00, MST^{+}00, MSI^{+}03, MR01b, MSD^{+}04, MWC^{+}02, MJY^{+}04, NLRD01, PLW^{+}04, RDC^{+}04, SBZ^{+}00, SGF^{+}00, SCY01, SYH^{+}01, STJ^{+}01, VBH^{+}02, VZTN03, WGvA^{+}01, WRGK00, WSE00, YpHRL03, CAGK^{+}03, ESS^{+}00, FSK^{+}04, GSB^{+}01, MNT^{+}03, PBD^{+}04, VCDHD03, WLO^{+}02, YSC^{+}02]. c-Abl [WMJ^{+}04]. C-dependent [MSI^{+}03]. c-erbB- [OSWG02]. c-Fos [ESS^{+}00]. C-Jun [PSKK^{+}00, PKR^{+}02]. C-like [ACBG04, MR01b]. c-Myc-induced [YSC^{+}21]. C-type [CdLvM^{+}04]. C. [AOJ^{+}04, BKD^{+}04]. C/EBP [HRB^{+}01]. C2 [DKAH04, EBWC01, PLL^{+}00]. C2b [DVE^{+}00]. C2C12 [ARM02, CMC^{+}02]. Ca [YpHRL03, ACP^{+}02, BPMG00, BKZ^{+}03, CRH02, DC02a, DVE^{+}00, DH02, Dov02-69, DFYL00, DFZ^{+}03, EBWC01, FFKC00, FYI^{+}03, GMY^{+}03, HTPC04, HDP^{+}01, pHYpXL00, IKA^{+}03, JJM^{+}02, LeB02-28, LeB03-58, LeB04-75, LRA^{+}02, LPPT^{+}02, LC04, MH02, MCH^{+}00, MZ00, MW04, MPV^{+}01, NMH^{+}04, OTB03, PDMO03, PFM^{+}00, PLW^{+}04, PMB^{+}00, RPS^{+}02, RLC00, RMMP04, SJW^{+}04, SMT^{+}03, SSP^{+}03, SZZ^{+}00, SM04b, WLW^{+}04, YpHRL03]. CA3 [UTH^{+}02]. CAAX [MHH^{+}03]. cable [HGP^{+}04]. cable-dependent [HGP^{+}04]. cables [HKBH03, LeB04-73]. CAD [LXL04a, CKL^{+}03, LDS^{+}00]. Cadherin [GWG01, KIK^{+}00b, LeB02-34, NG02, OTY^{+}04, CFWH^{+}01, CMC^{+}02, CWA^{+}03, CiKBG03, CSSBY^{+}03, DIR03, HT01, LZC^{+}03, LeB03x, LeB04-29, PCB^{+}03, SEW^{+}01, TT04, XAB^{+}03, YMK^{+}04, YK03a, AML00, APLB00, Gum00, HPQ^{+}00, LLH^{+}00, TNM^{+}00, TAD^{+}00, THK^{+}00, CTE^{+}04, CSSBY^{+}03, HKP^{+}04, IDvH^{+}02, ISB^{+}04, LeB04-49, PLR03, PGSE^{+}01, SBG^{+}04, WG03, XLH^{+}01]. Cadherin- [THK^{+}00]. cadherin-activated [YK03a]. cadherin-based [YMK^{+}04]. Cadherin-mediated [NG02]. Cadherins [Wel02i, LCM02, STP^{+}00]. Caenorhabditis [ACE^{+}01, AP00, BWV^{+}01, BHL^{+}03, COB01, GSW^{+}00, GGGN02, HKHO01, HOK^{+}02, HEW^{+}01, HMAM01, JJM^{+}02, LHvdH00, MR01b, MCB00a, NM02, OMM^{+}03, ODR^{+}01, Ono01, PCB^{+}03, RMG^{+}00, RBW^{+}02, RDS02, SB03, UCY^{+}02]. CagA [CAGK^{+}03]. Caged [RRJ^{+}01]. Cajal [Dov03f, OL02, Tum02b, BCB^{+}02, DHK^{+}04, FM01, HMG03, JBK04, Jeg01d, PGSL00, SRKN03, SN04, TBJ^{+}01]. Cal [AKM^{+}04]. Calcineurin [FHP00, MRM^{+}04, PC01, DCM00, WHM^{+}02]. Calcineurin-dependent [PC01]. Calcineurin-GATA-6 [WHM^{+}02]. calcitonin [DDL^{+}04]. Calcium [GTPG03, HGC00, LeB02r, LeB03k, LeB03j, LeB04m, MSGS02, Riz03, SRSW04, Sre04d, Tum03f, Tum04c, VCGB^{+}02, WGP^{+}00, Wel01j, Wel03f, ARMB04, BSMS03, Cam03, CGF^{+}04, CVZ^{+}04, CDW^{+}03, CMS^{+}02, Dov04n, JAS02, Lai03c, NGK^{+}03, OTB03, RLZ^{+}03, SUT^{+}01, TFAM^{+}04, Tum04n, VMS^{+}02, WDW03, YWH04, CGL^{+}01, FKG00, LAF^{+}00, MF01a, SBI^{+}00, Zwe00]. Calcium-coated [Wel01j]. Calcium-dependent [SRSW04, JAS02]. Calcium-independent [MSGS02]. calcium-inhibited [CMS^{+}02]. Calcium-regulated [VCGB^{+}02]. calcium-signaling [CDW^{+}03]. caliber [RGM^{+}02]. Calmodulin [pHYpXL00, EES^{+}01, HN03, SUT^{+}01, YDRS01]. calmodulin-dependent [SUT^{+}01]. Calnexin [RLC00]. calpain [DH02, BKD^{+}00, RRM^{+}00]. Calreticulin [HBL^{+}01, LPPT^{+}02, NZA^{+}01, OPP^{+}03]. calsequestrin [GTM^{+}01]. Cam [SMSM00, DMBS02, NYT^{+}03, SLB^{+}01, WWK^{+}00]. CamKII [SSP^{+}03]. Camp [ONM00, LeB02-28, TLS^{+}01, ZFH^{+}04]. CAMs [JB01]. Can [Ped04, ATF^{+}04, DBL^{+}02, GCG^{+}01, KWS^{+}02, LRBH02, LeB02-30, LeB03-100, MEV^{+}04, MBN^{+}01, Wel01t, Wel03z, ZLH^{+}03, BMD^{+}00, GOL^{+}01, MGL^{+}00, dEP01]. canal [Dov02-40]. cancer [LeB02-110, Wel01-63, YHZ^{+}01, HPQ^{+}00]. cancer-predisposing [YHZ^{+}01]. Candida [WBAS04]. Canine [HGP^{+}00, AFK^{+}03, MBMMA^{+}03, FFST01, VHLS00]. cannabinoids [Dov03-51]. canonical [CKS^{+}04, TJC^{+}03]. cap [LeB02-76]. Cap23 [FLX^{+}00, LFT^{+}00]. Capable [LQPC^{+}00]. Capacitative [LAF^{+}00, PFM^{+}00, Zwe00]. CapG [WLKS01]. capillaries [KVM03]. Capillary [MCB00b]. Capping [JRW^{+}01, KYW^{+}04, FFSF03, KKTP03, KB04, POW^{+}01a, SGdM^{+}01]. caps [Dov03-44]. capture [DKA01, HBB^{+}04, KCG^{+}03]. Carbohydrate [BSEB04]. carboxy [NMHH03]. carboxy-terminal [NMHH03]. carboxyl [MHH^{+}03, RCY^{+}03, RFLT02, WS00]. carboxyl-terminal [MHH^{+}03, RCY^{+}03, RFLT02]. carboxymethylation [MHH^{+}03]. Carboxypeptidase [KNIO01]. Carcinoembryonic [SDDS00]. carcinogenesis [MWC^{+}02, MJY^{+}04, PSS^{+}04]. carcinoma [BBR04, MSD^{+}04, PGSE^{+}01, SSRX04, SGPL^{+}00]. carcinomas [JW04, WFF^{+}01]. Card [GV00]. cardiac [AKM^{+}04, GGH^{+}04, GPLS02, LPPT^{+}02, MEK^{+}04, MPR^{+}03, PWC^{+}01, RMHM00, TTS00, THK^{+}00]. cardiomyocyte [AKH^{+}04, HCK^{+}00, MHK^{+}01]. Cardiomyocytes [MWN^{+}04a, BFC^{+}02, BBDM02]. cardiomyogenesis [PDL^{+}03]. Cargo [GWBW00, Sea04, VMD^{+}01, YB01, BPKK01, DKAH04, GTBM04, JFS^{+}03, LGB^{+}02, MJS02, MBN^{+}01, QDG^{+}04b, AFB^{+}01, KKS^{+}01, MNHR00]. Cargo-selective [Sea04]. cargoes [Bro03, RGM^{+}02, CDTW00]. cargos [Dov02-28]. Carmil [JRW^{+}01]. carnation [SKM03]. Carrier [LCM00]. carriers [LS02, PPM^{+}00]. carry [BHPN04]. carrying [MPG^{+}03]. cartilage [HBC^{+}03, HIO^{+}04, LeB03-73, SDML04, YCX^{+}01]. CAS [CK00, CK02]. CAS/Crk [CK00, CK02]. Cascade [vdHvODML^{+}00, GLS^{+}03, JBA^{+}01, OMWSN02, dSAH02]. cascades [Dov04h, MHK04, PDJ00]. Casein [DSSWW00]. Caspase [BSMS03, DKJ00, ESH^{+}01, GV00, HLU00, LLP^{+}02, RGG03, CDM^{+}02, CDK04, CJ02, CSJ03a, CNMS02, CSJ^{+}03b, CFM^{+}02, DRC^{+}02, DMLK04, Dov03-61, ERS^{+}04, HKE^{+}04, KLD^{+}03, KSR^{+}04, LeB02j, LeB02-29, LXL04a, MEV^{+}04, PSK^{+}03, RBE^{+}02, SSM^{+}04, ST03, SHE^{+}02, SPB^{+}01, SZvBuH^{+}04b, TNMM03, Wel04j, WK04, KWH^{+}00, MMR^{+}00]. Caspase-1 [KWH^{+}00]. Caspase-11 [KWH^{+}00]. caspase-12 [KLD^{+}03]. caspase-2 [MEV^{+}04, RBE^{+}02, MMR^{+}00]. caspase-3 [SZvBuH^{+}04b, KWH^{+}00]. caspase-4 [HKE^{+}04]. caspase-8 [KSR^{+}04, SPB^{+}01]. caspase-9 [ERS^{+}04, MEV^{+}04, SSM^{+}04]. caspase-activated [LXL04a]. Caspase-Dependent [HLU00]. caspase-independent [CJ02, CSJ03a, CSJ^{+}03b, CFM^{+}02, LeB02j, ST03]. Caspase-mediated [LLP^{+}02, RGG03]. Caspases [FL00, LeB03l, BSD^{+}01, MSP^{+}01, MMDC00]. Caspr [GSSP03, FSGDN^{+}00, GSP^{+}02b]. Caspr2 [PSE^{+}03, TGD^{+}03]. Cast [OTRI^{+}02, Dov02k, TRMI^{+}04]. catabolic [AOH^{+}02]. catalysis [SBS^{+}02b]. catalytic [RZB^{+}03, PHWK^{+}00]. catalyzed [ESC^{+}01, WBM^{+}00]. cataract [MKR01]. Cataracts [Dov01e, Dov03-65, MWSL^{+}03]. Catastrophe [LeB03m, Wel03-28]. Catch [YWW^{+}04, Dov02u, Wel03-55]. catches [Dov03-58]. Catching [Dov03g]. Catenin [ESS^{+}00, IEJ^{+}01, LCG^{+}04, PTM^{+}01, Wel03g, WHP^{+}02, CLG^{+}03, CMC^{+}02, CWA^{+}03, CiKBG03, DIR03, GG04, HBK^{+}02, IDvH^{+}02, LZC^{+}03, LSK01, LDP02, MSJ^{+}02, MOMK03, MO01, PLR03, PGSE^{+}01, PCB^{+}03, SEW^{+}01, TJC^{+}03, WBT^{+}03, XAB^{+}03, GWG01, HVB^{+}00, NLBK00, RSG01, TM00, THK^{+}00]. Catenin/Lymphoid [ESS^{+}00]. Catenin/T [CGY^{+}01b, SKF^{+}01]. catenins [FJK^{+}04]. Caterpillar [Wel04k]. Cathepsin [FWM^{+}01]. Cation [SSGLS01, AHA^{+}04, FAT^{+}02, JOF^{+}02]. cation-independent [AHA^{+}04]. caught [Dov02-53]. cause [Dov04p, LZS^{+}03, MSY^{+}04]. caused [GK03b, HSC01, LSW^{+}03, ML04, MRT^{+}01, XRH^{+}03, UIK^{+}01, yZCKA01]. Causes [WFQ^{+}00, ASK^{+}03, BFG^{+}04, CLG^{+}03, HPFG03, HIO^{+}04, KBK^{+}03, iKFH^{+}04, Lai03a, LTF^{+}01, MMBB03, NMHH03, PRS^{+}04, ZCH^{+}01, ZWSC02, CDWB01, LWS^{+}00, PQF^{+}00, PDJ00, RRJ^{+}01]. Causing [AC00]. caution [Wel02v]. Cautious [Wel03h]. caveolae [PMP^{+}03, NL03, VHLS00]. caveolae-mediated [PMP^{+}03]. Caveolae/ [NL03]. Caveolin [vM01, KNI^{+}04, ORZ^{+}04, FKI^{+}01, MLLA00, OPZ^{+}01, PLL^{+}01]. caveolin-1 [ORZ^{+}04, MLLA00]. Caveolin-2 [FKI^{+}01]. caveosomes [LeB03-70]. Cbfa1 [JBA^{+}01, KPL^{+}02, LTF^{+}01, UIK^{+}01]. Cbfa1-independent [KPL^{+}02]. Cbk1p [WKZ^{+}02]. Cbl [SHN^{+}01]. Cbp [BKB^{+}01, TM00, vMZM^{+}00]. Cbp/P300 [vMZM^{+}00]. cCAF [FLWMG02]. Ccr5 [SPMM^{+}00]. Ccz1 [WSK^{+}03]. CD14 [DPO^{+}04]. CD148 [LZC^{+}03, LeB03-90, LW03]. CD151 [CWA^{+}03, KYS^{+}02, SSO^{+}03, SGO^{+}00]. CD31 [JOF^{+}02, RRK^{+}00]. Cd31-Induced [RRK^{+}00]. Cd34 [BHY^{+}00, TTP^{+}01]. CD40 [BHW02b]. CD44 [CP03, Dov01b, NMH^{+}04, OKM^{+}01, PAR^{+}04, SRD^{+}02, DLR^{+}01, KIC^{+}01, OKSH00, SRKR00]. CD46 [LBH^{+}02]. CD47 [LCS^{+}02]. CD81 [TTM^{+}03]. CD9 [TTM^{+}03, MGLPM00, SFSD00]. Cd9-Dependent [MGLPM00]. CD95 [BLU^{+}04]. CD95-induced [BLU^{+}04]. Cdc12p [KKTP03]. Cdc14p [PMGS02, Wel02j]. Cdc15p [CG03]. Cdc1p [RRB^{+}01]. cdc2 [MZT^{+}03, KCWF02, KTY^{+}00]. CDC20 [SCY01, Wel03r, HdEV^{+}02, KBWG02, RJyH02, RM00]. Cdc20-Dependent [RM00]. Cdc20p [HJSM00, HPE^{+}01]. Cdc24p [NA00]. Cdc28 [RHM00, RM00]. Cdc31p [JGW02]. Cdc34 [TBRG01]. Cdc34/Ubc3 [TBRG01]. Cdc37 [LRF^{+}02a]. cdc42 [KHLW02, AMG^{+}01, AS00, BS04b, CTE^{+}04, FSK^{+}04, FSM^{+}01, GHS00, HP00, HS04, HGC00, HN03, ISB^{+}04, LSSL00, LJK^{+}01, RDH^{+}01, RyHHK00, SSO^{+}03, SWG^{+}03, WYHP00, WSR03, YHD^{+}00]. Cdc42-Dependent [LSSL00, SSO^{+}03]. Cdc42-GEF [FSK^{+}04]. Cdc42-Induced [YHD^{+}00]. cdc42-mediated [KHLW02]. Cdc42hs [GKM^{+}01]. Cdc42p [GBZ^{+}02, NA00]. Cdc45 [VTH^{+}02]. Cdc48p [FNFL03]. CDC5 [SL01]. Cdc6 [FKSJ01, OSB04]. Cdh1 [HdEV^{+}02, RJyH02]. CDK [FNFL03, HLU00]. Cdk1 [DOL^{+}01]. Cdk5 [NB03b, TSL^{+}03]. CDX2 [BvdWD^{+}04]. CeCDC [GGGN02]. CeCDC-14 [GGGN02]. CeGLC [RDS02]. CeGLC-7 [RDS02]. Cell [ACE^{+}01, BYLA^{+}01, CS01, CW02, CGY^{+}01b, CK00, CQH^{+}00, DSSY00, DB02, HSB00, KRA^{+}01, KSK^{+}00, LA04, LYMC00, Mel01, NMH^{+}04, PFW^{+}00, PTM^{+}01, Pow01b, Qua00, RS00a, RSBE00, RDS02, SNL^{+}00, SHN^{+}01, SM00, SKF^{+}01, US04, VCG^{+}00, WWK^{+}00, XHG^{+}00, Zam04, ZEW^{+}01, Zwe00, ADL^{+}03, ARQ^{+}04, AZ03, AJ01, AQC04, BGM03, BFC^{+}02, BSD^{+}01, BS04a, BGA^{+}04, BJM^{+}02, BMS^{+}03, BSW^{+}04, BWA^{+}04, BFG^{+}04, BPKR^{+}02, BFSO^{+}04, BHK^{+}02, CDM^{+}02, CGM^{+}02, CMC^{+}02, CWA^{+}03, CS03a, CYC^{+}04, CPG^{+}03, CARG^{+}04, CPA^{+}03, CTE^{+}04, CRS^{+}03, CMS^{+}02, CBLT04, CFM^{+}02, CKW^{+}03, DSB^{+}02, DMC^{+}03, DMH^{+}02, DBB^{+}02a, DFJ^{+}02, DBL^{+}02, Dov01y, Dov01-27, Dov03d, Dov03-60, DSB^{+}01, DBS^{+}03, EA03, FAT^{+}02, FNZ^{+}03, FHL^{+}03, FFSF03, FCM^{+}01, FNFL03, FRM^{+}02, FGSW03, FSK^{+}04, FKH^{+}04, Gal02, GMRYM^{+}02, GWL03, GGF^{+}03, GSB^{+}03, GHK^{+}03, GSP^{+}02b, GMLM^{+}04, GCT^{+}04]. cell [GQI^{+}02, HP04, HPFG03, HKE^{+}04, HRV^{+}01, HEW^{+}01, HIT^{+}02, HT01, HMH^{+}03, HGP^{+}04, HDH02, IBS^{+}02, JLK^{+}02, JW04, KMK^{+}02, KVC^{+}03, KHB^{+}04, KNR^{+}04, KI04, KDH^{+}04, KFO04, Lai03e, LeB02-91, LeB03n, LeB03v, LeB03-94, LeB04-61, LRF^{+}02a, LFM^{+}04, LSMS03, Les01k, LAL^{+}03, LCS^{+}02, LHC^{+}02, LW03, LPL^{+}04, MB03a, MWL01, MZT^{+}03, MC01, MTB^{+}02, MWN^{+}04a, MTM^{+}03, MGAL^{+}01, MEK^{+}04, MM03, MYO^{+}04, MAG^{+}04, MGP^{+}02b, MPAP^{+}03, MSD^{+}04, MCA^{+}03, NMG04, NG02, OMiKF02, PBB^{+}04, PBD^{+}04, PMGS02, PRJK01, PAG02, PSD^{+}04b, PMPH03, QS02, REK^{+}03, RZB^{+}03, RDP03, RP03a, RMC^{+}02, RAS^{+}03, SOH^{+}04, SLB^{+}01, SC01a, SCTF04, SCPP02, SRSW04, SP03, SM03a, SSO^{+}03, SKN^{+}03, ST03, SDMC^{+}04, SKH03, SEW^{+}01, SAP03, SDD04, SLD^{+}02, TNM^{+}03, TT04, TMG03, TOTC01, TWS^{+}04, TST^{+}03, TSY^{+}02, Tum03b, Tum03d, Tum04c, Tum04t]. cell [Tum04-29, UJL^{+}03, UGKT^{+}02, WFI^{+}04, WNM^{+}03, WSWSL04, WCBC04, WKZ^{+}02, Wel02v, Wel03-27, Wel04-63, WZB^{+}01a, WPS^{+}01, WSR03, WML^{+}03, WG03, WLO^{+}02, WMJ^{+}04, WDS^{+}03, WSL^{+}01, WK04, XLH^{+}01, YSK^{+}04, YMK^{+}04, YK03a, YOK^{+}03, ZHDB04, ZSY^{+}03, ZLVS02, ZTK^{+}03, ZAE^{+}04, AC00, ALJ00, AEL^{+}00, BD00, BMD^{+}00, BT00, BZSC00, CP01, DB00, FSK^{+}00, FWM^{+}01, FCL^{+}00, FKI^{+}01, GPAS^{+}01, GVT^{+}00, GBM^{+}00, HK00, HPQ^{+}00, HSB00, HAP^{+}00, HZ00, IAG^{+}00, IHN^{+}01, IOLA^{+}00, JBN^{+}00, JPM^{+}00a, KIC^{+}01, KR01, KGE^{+}00, KH01, KGC^{+}00, KGvdG^{+}00, LFT^{+}00, LR00, MGL^{+}00, MST^{+}00, MKK^{+}00a, Mel00a, MTV^{+}00, ME00, NBWB^{+}00, NKP^{+}01, NT00, Nob00, PGS^{+}01, PO00, PBL^{+}00, PVjL^{+}00, RMG^{+}00, RCL^{+}00, RBBA00, RRJ^{+}01, SIBG01, SFSD00, SMSF00, SMSM00, SEP^{+}01, SGPL^{+}00, TNM^{+}00, TGMC^{+}00, TRW^{+}00, TB00b, TSK^{+}00, THZ^{+}01, WFF^{+}01, WTG01, Wel02-56, WPO00]. Cell [YSS^{+}01, YKW^{+}00, ZN01, ZLMP00]. cell-autonomous [KHB^{+}04]. Cell-Free [SM00]. Cell-to-cell [Zam04]. Cell/Lymphoid [PTM^{+}01]. Cells [ALWR01, Dov03h, HGP^{+}00, HGC00, IWG^{+}01, LeB04n, LeB04o, RTM^{+}01, SGYD^{+}01, TW02, Wel02k, WL04, APBC^{+}02, ACBG04, ABRA03, AAM^{+}04, AFK^{+}03, ATF^{+}04, ASGGR02, BFC^{+}02, BBP02, BBR04, BCA^{+}03, BRB^{+}01, BMS^{+}03, BSW^{+}04, BLP^{+}02, BGW^{+}04, BKZ^{+}03, BNSR03, CSP^{+}04, CdLvM^{+}04, CS04, CPN^{+}01, CLG^{+}03, CSIK03, CSL^{+}03, CK02, CLB^{+}02, CPA^{+}03, CGBL^{+}02, CBS04, CACL03, DBH^{+}01, DC03, DME^{+}04, DDV^{+}03, DRBF03, DMH^{+}02, DPO^{+}04, DRG^{+}03, DRC^{+}02, DMLK04, Dov01x, Dov02r, Dov03-47, Dov03-49, Dov03-44, Dov03-48, Dov03-54, Dov03-55, ERS^{+}04, ERMT^{+}04, FPP^{+}02, FCF^{+}01, FLS^{+}04, GMRYM^{+}02, GNS^{+}04, GGT^{+}02, GBJ01, GK03b, HDP^{+}01, HViV^{+}02, HEN^{+}01, HdO^{+}01, HP03, HN03, HAK^{+}04, HBSJ04, ITM^{+}04, JAS02, JOF^{+}02, JON^{+}03, KBWG02, KRS^{+}02, KVM03, KSC02, KRS^{+}01, Kur03, LUB^{+}02, LeB02c, LeB02f, LeB02k, LeB02-30, LeB02-74, LeB02-86]. cells [LeB03x, LL03a, LeB03-33, LeB03-42, LeB03-47, LeB03-49, LeB03-51, LeB03-50, LeB03-55, LeB03-60, LeB03-72, LeB04c, LeB04-67, LeB04-94, LeB04-105, LeB04-110, LXL04a, LBH^{+}02, LRA^{+}02, LGM^{+}04, LCT^{+}04, LP04a, LRWB04, MRM^{+}04, MBMMA^{+}03, MO01, MP04, MVL^{+}02, MF01b, MMFS01, MPV^{+}01, MTM^{+}02, MAG^{+}04, MBSR03, MMG^{+}04, NDM^{+}03, NGKH02, OTB03, OER^{+}02, OGD03, PGSE^{+}01, PDL^{+}03, PPP^{+}01, PRLR02, PMU^{+}02, PPA^{+}03, PSW^{+}02, PPK^{+}01, Pro03, QPDJ^{+}02, RPNM03, RWSV03, RDNB02, SAWS02, SGdM^{+}01, SCB02, SPB^{+}02, SRG^{+}04, SRSW04, SMS^{+}01a, SK04, SBG^{+}04, SSRX04, SRB^{+}04, SRD^{+}02, Sre04o, SPB^{+}01, SCY01, SwZK^{+}02, TAA04, TAA^{+}02, TUK03, TCP^{+}03, TOM01, TMA^{+}04, TKS^{+}02, TEB^{+}03, TNBH01, Tum02c, Tum04h, Tum04g, VFN^{+}04, VBR^{+}01b, VBH^{+}02, WSC^{+}03, WNM^{+}03, Wel01u, Wel01-39, Wel01-59, Wel03-63, Wel04e, Wel04-61, XAB^{+}03, YGWN01, Yum01, ZGN^{+}04]. cells [ZRDG02, dMMBK^{+}02, dVKS04, AV01, AEL^{+}00, BHY^{+}00, CNBWN00, CH01, CHM^{+}01, CDWB01, DBS^{+}01, DLR^{+}01, DWD^{+}00, FRO01, FPSM01, FFKC00, FFST01, HPQ^{+}00, HVT^{+}00, HFM^{+}01, HHF^{+}00, JLS^{+}01, KTY^{+}00, KMiM^{+}01, KC01, KHH^{+}01, KMG^{+}01, LWDH01, LSA^{+}00, LQPC^{+}00, LRW^{+}00, LM00b, LBWH^{+}00, MF01a, MCG^{+}00, McI01, MSM^{+}01, MKST00, MYH^{+}01, OAR^{+}00, PBL^{+}00, PMK^{+}00, PFM^{+}00, PGSL00, PKF^{+}00, RSG01, SRKR00, SMS^{+}01b, SWJ^{+}00, STE^{+}01, SSW^{+}01b, TTR00, TTP^{+}01, TW02, VHLS00, WL03, YHF^{+}01, YYM^{+}01, YW00]. Cellular [BRM^{+}00, MSY^{+}04, XAB^{+}03, BSEB04, ELNA^{+}03, HTS02, KYS^{+}02, KLF04, KSM^{+}01, LeB04-50, Mis01, SHM02, TOM01, YLQ^{+}02, ABP^{+}00, GWG01, KYF00, MR02, SCMH00, TGM^{+}01, UAZG00]. Cellularization [SFV^{+}00]. cellulose [GPL^{+}02]. CENP [GKYY03, LeB03g, iONOM02, TBRG01, Wel01k, YMM^{+}00, ZSS01, DJT^{+}03, MR01b, SBC01]. CENP-A [ZSS01].

CENP-B [iONOM02]. CENP-E [TBRG01, Wel01k]. CENP-meta [YMM^{+}00]. CENs [Wel01d]. centers [SLG02]. Central [ANC^{+}02, BHNG01, DWM03, GNH^{+}04, HHS03, ISS^{+}04, MN04, SWBE^{+}04, SR03, Wel02-27, BAD^{+}00, CdBB^{+}01, DR00a, GG01, MHW^{+}00]. centric [GBMA04]. centrin [JGW02, Kil03, MKA^{+}00]. centrin-binding [Kil03]. centriolar [DWFAdL^{+}02]. centriole [GJS^{+}03b, LeB02-82]. centrioles [Dov02i, Sre04c, PMK^{+}00]. Centromere [SK01, WBP^{+}03, Wel03i, CDB02, GKYY03, sKCK^{+}01, MB03c, iONOM02, AMEC01, DTO^{+}01, MPSM00b, WK01]. Centromere-associated [WBP^{+}03]. Centromeres [AH01, GNH^{+}04, MR01b]. centromeric [LeB02-88]. centrosomal [DM02, HOK^{+}02, OER^{+}02, MST^{+}00]. Centrosome [BNSR03, APBC^{+}02, CPC^{+}02b, HKHO01, OMM^{+}03, TSH^{+}04, MST^{+}00, MKA^{+}00, PMK^{+}00, ZKW^{+}00]. Centrosomes [KR01, LeB03n, Tum04d, Dov03v, GMD^{+}02, KBWG02, KRS^{+}02, QS02, RME^{+}00, SCS^{+}00]. centrosomin [TUK03]. Cep135 [OER^{+}02]. ceramide [BMS^{+}03, BSW^{+}04, FR01, LXL^{+}04b, TRW^{+}00]. cerebellar [ABF^{+}03, Dov03i, FHL^{+}03, PPK^{+}01, SLB^{+}01]. cerebellum [Dov01u, Wel01s]. cerevisiae [HBB^{+}04, AQHO03, AC00, BT00, EKC^{+}03, FBV^{+}04, HSKG00, HvdBP^{+}01, HGP^{+}04, HKBH03, JRL^{+}03, JC01, KKPB03, KNIO01, KSK01, LVD^{+}04, MB03c, NPL04, PLH^{+}01, PGV^{+}00, RRB^{+}01, SKR^{+}00, SL01, STE^{+}01, VYC^{+}00, VTGT^{+}03, WKZ^{+}02, WAPB^{+}00, WK01, YF00, vRTvdB^{+}00]. CFTR [SPN^{+}04, ZAE^{+}04, YRC^{+}04]. cGKI [SWH^{+}02]. cGMP [FHL^{+}03, LeB02s, LeB03-67, MRT^{+}01, MFF^{+}03, SWH^{+}02]. cGMP-dependent [FHL^{+}03]. cGMP-mediated [SWH^{+}02]. CH [FGSW03]. CH-ILKBP [FGSW03]. Chain [AP00, DSV^{+}00, ATG^{+}03, CWG^{+}02, HZS^{+}01, JLJD03, RL03, RFLT02, UHR^{+}03, XRH^{+}03, FK01, GPAS^{+}01, KMH^{+}00, LVWA01, PVjL^{+}00, TWS^{+}00b, YDRS01, vRTvdB^{+}00]. Chains [DDW^{+}01, LVD^{+}04, NKM^{+}03, MDJF00]. Change [GOL^{+}01]. Changes [KLT^{+}03, NBWB^{+}00, CLB^{+}03, DH02, FNZ^{+}03, LKM^{+}04, RRS03, TNM^{+}03, VMK^{+}03, WSC^{+}03, AML00, BTD^{+}00]. Channel [HGP^{+}00, DBY01, DC02a, IKA^{+}03, JOF^{+}02, JJM^{+}02, KS02, LeB02-51, PPP^{+}01, RPS^{+}02, RWCC01, RLZ^{+}03, RDS02, SAS^{+}02, Wel02-34, Wel04-36, ATE^{+}01, EHCC^{+}00, FKG00, FFKC00, SMS^{+}01b, TMK^{+}00]. Channeling [Wel02l]. Channels [Wel04l, CLM^{+}04, Dov02-64, Dov03r, Dov03-63, Jeg01a, JB01, LeB02-109, PSE^{+}03, RPZ^{+}02, SLB02, Wel04-77, Coo01, FLLE^{+}01, FFKC00, SBI^{+}00]. chaperone [BXR^{+}02, Dov02-35, Dov02l, JMG04b, LeB02-53, LeB04p, LeB04m, SATA^{+}02, WAV^{+}04, NHI^{+}00]. chaperones [MGP^{+}02a, iNFK^{+}01]. chaperonin [KKPB03, MDJF00]. characteristic [ALEH02, RKF^{+}04]. Characterization [AZP^{+}02, GMC^{+}00, KIK^{+}00a, KSS^{+}03, LYKH00, ML04, OER^{+}02, PO00, RVB^{+}01, SSK^{+}03, SVG^{+}00, UOB^{+}02, BSL^{+}01, HGS^{+}01, SwZK^{+}02]. Charcot [PQF^{+}00, RKF^{+}04]. charge [Dov03b, Tum03a, HSW00]. charged [KNI^{+}04]. charging [PPWM04]. Checking [LeB04q, Wel04m]. Checkpoint [GKG^{+}01, SCY01, AAM^{+}04, BJB^{+}03, Che02, CNT03, Dov01w, Dov01-33, Dov02w, GES04, HdEV^{+}02, HCL^{+}03, HMC^{+}01, Hoy01, IKS^{+}02, KCWF02, LPPT^{+}02, MSR04, OSB04, SVLM02, THG^{+}04, TEC^{+}03, VN04, VLL^{+}03, WBP^{+}03, Wel03j, YDPK04, AOC01, HJSM00, SBC01]. checks [LeB03g]. chemical [SKGC^{+}03]. chemoattractant [ZVPK03]. chemokine [FLWMG02, SKN^{+}03, Wel02m]. chemokines [Wel02-40]. chemotactic [HTRK02, vEWS^{+}01]. chemotaxis [GMLM^{+}04, SWG^{+}03, SWE^{+}03, VCDHD03, DBS^{+}01, FMMF01]. chew [LeB03l]. chewed [Wel03i]. Chfr [MP04, KCWF02]. Chick [GBD^{+}00]. chimaera [MPV^{+}01]. Chinese [LCI^{+}01]. CHIP [YRC^{+}04, Cha00]. Chk1 [Dov01f, FHJW^{+}01, SRL^{+}04]. Chk1-mediated [SRL^{+}04]. Chk2 [Dov01f]. Chlamydomonas [PDV^{+}00, AZP^{+}02, IBKSP01, MLKH04, MN04, RP03b, SLT^{+}01, TDL03]. chloroplast [AZP^{+}02, BHW^{+}02a, HBV^{+}01, JMB^{+}04, MGMH03, SHKS02, SRW^{+}04, Tum03k, RBB00, VMO01]. chloroplasts [SSK^{+}03, KKR^{+}00, SS00]. CHO1 [KGT^{+}02]. choice [RPNM03, Wel04-40, Wel04-39, BMS^{+}00b]. Cholera [MZ00, SHWH00, TR02]. Cholesterol [CSJ00, LeB04r, LeB04s, PKF^{+}00, ARLC^{+}04, LeB03-45, SKM^{+}02, BSW^{+}00, PLL^{+}01, SKT^{+}00, vM01]. Cholesterol- [SKT^{+}00]. chondrocyte [DLPB03, FAF^{+}04, HIO^{+}04, LRS^{+}02, NGS^{+}01, TPA^{+}03, WK02, ZZM^{+}03, YCX^{+}01]. chondrocytes [SDML04, TPA^{+}03, IHK^{+}00, UIK^{+}01]. chondrodysplasia [TPA^{+}03]. chondroitin [YPN^{+}04, LLH^{+}00]. Choose [LeB02t, LeB02c, LeB02-61, LeB03-65]. Choosing [LeB03o]. chop [Wel02-28]. Chopping [Wel01l]. Choreographing [LeB03p]. Christmas [Pow01c]. Chromaffin [JLS^{+}01]. chromatid [SAH^{+}03, HSKG00]. Chromatin [SMS00, BC03, CW02, CKW^{+}03, DBLG02, GKYY03, LeB02-45, LeB02-88, LeB03b, LeB03-81, LeB03-75, LWZ03, MPB02, MEFC03, MWHM01, NW04, iONOM02, SSGLS01, Tum02e, YHZ^{+}01, CW01, FKSJ01, JWJJ00, SKJ^{+}00, TWS^{+}00a, Wel01-28]. chromokinesin [LC01, MSM04b]. Chromosomal [Dov02m, LGK00, DJ03, LKLD04, SGdM^{+}01, HSKG00, JWJJ00, WWJ^{+}00]. Chromosome [GHC01, GK03b, SCLC00, WSC^{+}03, BJB^{+}03, BVH04, CC02, CW02, CH03, DJT^{+}03, GES04, JKG^{+}02, sKCK^{+}01, KLK^{+}04, LHK02, LC01, MPB02, MPG^{+}02, MSL^{+}02, MSM04b, PTH^{+}04, RKKP02, RBW^{+}02, SWB03, TCH^{+}02, WBP^{+}03, Wel02x, WDMH03, YK03b, AMEC01, CW01, CNJ01, FAAS00, GG01, LHvdH00, PN00, PMSB01, TB00b, WK01, YMM^{+}00, dEP01]. Chromosome-Associated [SCLC00]. Chromosomes [CGBL^{+}02, Wel03k, BGBG03, CSM04, LeB03c, LeB04b, Les02e, NW04, PRS^{+}04, SSGLS01, SWB03, TBRG01, VLL^{+}03, Wel01j, HMAM01, PN00, RME^{+}00]. Chronological [HJL^{+}04]. Cilia [Bir04c, Tum03a, ITM^{+}04, JGR^{+}04, MCBB^{+}04, Tum04r, PDV^{+}00]. cilia/flagella [MCBB^{+}04]. ciliary [SK04, YLY^{+}02]. cilium [LeB02-85]. circle [LeB04-41, VYC^{+}00]. circuit [Vin04]. circuits [UTH^{+}02]. circular [Wel04n]. Circulating [KMG^{+}01]. circulation [LeB02-40]. Cis [PSWU00, GTBM04, LGW00, HMRH01, STP^{+}00]. cis-Golgi [GTBM04, LGW00]. Cis-Snare [PSWU00]. CISK [XLGS01]. cisternae [MBN^{+}01, SHF^{+}03, HMRH01, OAR^{+}00]. cisternal [MMPO^{+}01]. City [Wel02-51]. Civil [Wel03l]. CK1 [LSK01]. Cl [SV03, RLTC^{+}02]. Cl41 [LCT^{+}04]. Cla4 [VT04]. Clamps [HSW00]. Clash [LeB02u]. class [BHW02c, CSG01, FMJG04, GGD^{+}04, JMG04b, KRS^{+}01, KJB^{+}02, PPWM04, VBR^{+}01a, WRSMO^{+}04, MGL^{+}00, RPTNM01, WSE00]. classical [PLP02]. Clathrin [JFS^{+}03, MBSR03, PWS^{+}01, SDS00, Wel04o, WZB^{+}01b, ALC^{+}03, CEGZ^{+}04, CNH^{+}02, CHM04, CS02, CS03b, EGWK^{+}01, FPM^{+}03, GK03a, LeB02-92, MBMMA^{+}03, MBH^{+}02, NOM^{+}04, RWH02, SM03a, TSL^{+}03, TST^{+}03, Wel01x, Wel04-54, FPSM01, NS01, KKW^{+}03]. clathrin-coated [CEGZ^{+}04, GK03a, SM03a]. Clathrin-dependent [PWS^{+}01, ALC^{+}03, CHM04]. Clathrin-Mediated [SDS00, MBSR03, CS02, CS03b, MBMMA^{+}03, MBH^{+}02, RWH02, TSL^{+}03, WZB^{+}01b]. clathrin/AP [KKW^{+}03]. clathrin/AP-2 [KKW^{+}03]. clathrin/dynamin [TST^{+}03]. clathrin/dynamin- [TST^{+}03]. Claudin [FHF^{+}02, TWBV^{+}01, NHG^{+}03, WSH^{+}04, FFST01]. claudin- [NHG^{+}03]. Claudin-11 [TWBV^{+}01]. Claudin-2 [FFST01]. Claudin-based [FHF^{+}02]. claudins [PMU^{+}02]. Clb5p [SCM^{+}00]. ClC [RDS02]. clean [Wel02w]. clean-up [Wel02w]. clearance [HdO^{+}01]. clears [LeB04-57]. Cleavage [Glo04, AZ03, BTH^{+}03, BSMS03, CWG^{+}02, CNMS02, Dov03-45, GRSL^{+}04b, ISS^{+}04, LLP^{+}02, US04, LVWA01, LW00, LHvdH00]. Cleaved [MMDC00]. Cleaves [KIC^{+}01, MMR^{+}00]. Cleaving [Wel01m]. Climp [KKL^{+}01]. Climp-63 [KKL^{+}01]. CLIP [LKM^{+}04, LeB04-64, PPGN^{+}02, TDFV02]. CLIP-170 [LKM^{+}04, LeB04-64, PPGN^{+}02, TDFV02]. CLIPR [PPGN^{+}02]. CLIPR-59 [PPGN^{+}02]. CLIPs [Dov02n]. cliques [Dov02m]. Cln3 [ARQ^{+}04]. clogs [LeB03-45]. Clonal [DGB^{+}00, LQPC^{+}00, SwZK^{+}02, MEK^{+}04, ZWB04, LMW^{+}00]. clones [Dov03-27, LeB04-43]. Cloning [BHW^{+}03, Wel02n]. Close [MWP^{+}00, Dov03i, Dov03-50, DW02, MB01, Wel03g]. close-ups [Dov03-50]. Closed [NW04, LeB04-71, YSGS00]. Closet [McF00]. closure [FLWMG02, KGE^{+}00]. clotting [LeB03j]. Clumping [Dov01g]. cluster [MWF02]. Clustering [CRP^{+}04, RPZ^{+}02, SJB^{+}03, Tum04e, CHA^{+}01, FCLSN03, KEHAM^{+}02, KCY^{+}04, KS02, PSE^{+}03, SDD04, TBW^{+}04, VBR^{+}01b, BF01, CDWB01, EHCC^{+}00, TMK^{+}00, WYHP00, WGF^{+}00]. Clusters [DLXP00, Dov02-53, JCPWS01, SBS02a, SHW01, WKJS^{+}04, BKD^{+}00, BGFJ01, JCR^{+}01, NSL^{+}01]. CMT4B1 [BBP^{+}04]. CMT4B1-like [BBP^{+}04]. CNS [CRS^{+}03, Dov03a, RPNM03]. CoA [TNW^{+}02, CKFH00]. Coactivator [BKB^{+}01, TM00, YHK^{+}00]. Coalesce [NSL^{+}01]. Coalignment [CLM^{+}04]. coat [RAD^{+}02, RMR02, SDL02, YLG^{+}02]. Coated [SDS00, CEGZ^{+}04, GK03a, JFS^{+}03, SM03a, Wel01j, BSW^{+}00, HvdKDS01, NS01, MJS02]. Coatomer [SSN01]. cochlear [iKFH^{+}04]. cocktail [Tum02d]. Cod1p [CKFH00, CRH02]. Cod1p/ [CRH02]. Cod1p/Spf1p [CKFH00]. code [WMA^{+}04a]. coelomocytes [HV03]. Cofactor [HRE^{+}01, LHW^{+}01, HF03, TNW^{+}02, BLC00]. cofactor-containing [TNW^{+}02]. Cofilin [ZBB^{+}00, BGR^{+}01, GOV^{+}03, MSD^{+}04, NOOG^{+}04, CBZ^{+}00, AOJ^{+}04]. cofilin-dependent [OO02]. cofilin-phosphatase [NOOG^{+}04]. COG [UOB^{+}02, Wel02-27]. Cohesin [LYKH00, Wel03m, Wel04p, ALEH02, EOJ^{+}03, MYC^{+}02, LGK00, SVG^{+}00, Wel01n]. cohesins [EOJ^{+}03]. cohesion [RBW^{+}02, SAH^{+}03, WHB04, Wel01n, HSKG00, MST^{+}00]. Coil [EM00, DRP^{+}03, MWM^{+}02, OER^{+}02, YLY^{+}02]. Coiled [EM00, DRP^{+}03, OER^{+}02, YLY^{+}02]. Coiled-Coil [EM00, DRP^{+}03, OER^{+}02, YLY^{+}02]. coilin [Jeg01d, OL02, TBJ^{+}01]. coincident [LCI^{+}01, qZC01]. Col1a1 [JPM^{+}00b]. Cold [Wel01o]. Cold-induced [Wel01o]. coli [MTM^{+}02, MKST00, RSG01, ASP^{+}00, HSMB02, KM00, UHR^{+}03]. Collagen [ACE^{+}01, Dov01h, KLGC^{+}01, LeB02v, LeB04t, BS04a, ELNA^{+}03, ISID^{+}03, SH02, SML^{+}04, WDS^{+}03, XRP^{+}01, ZZM^{+}03, ECO^{+}00, GJ00, HAP^{+}00, NHI^{+}00]. collagenase [SOH^{+}04]. Collapse [FNK^{+}00, WBC^{+}00, qZC01]. collar [VT04]. Colocalization [BET^{+}03, TN00b, FNK^{+}00]. Colocalize [vMZM^{+}00]. Colocalizes [AP00, vEPP^{+}01, DLS00]. colon [MWC^{+}02, MJY^{+}04, PGSE^{+}01]. colonocyte [TKB^{+}04]. Colony [NLRD01]. Colorado [Wel03-38]. colorectal [GK03b]. columnar [SDML04, TAA04]. combination [KEHAM^{+}02]. come [Dov02-33, Lai03f, LL03a, LeB03-63, Wel02-41, Wel04-34]. come-hither [LL03a]. comes [LeB04w, Mel00a]. comet [LeB04-106]. comets [LeB04g]. Coming [LeB02w]. Comm [Wel02o]. Commissural [FSK^{+}00]. commit [LeB02-75]. Commitment [WWK^{+}00, FNFL03, DKJ00, FXPT00]. committed [CSJ^{+}03b]. Committing [Wel03n]. common [JBK04, LS02, MBN^{+}01, LFT^{+}00, MHK^{+}01, SL00]. communicates [LeB02-35]. communication [LM01, LWCKL01, NGK^{+}03, RDC^{+}04, HSB00, LTB^{+}00, LGP00, RGG00]. commute [CRE02]. compaction [Wel04k, YK03b]. company [LeB04-62]. Comparisons [WLKS01]. Compartment [WFT^{+}01, ASYL04, HLB^{+}02, HIG^{+}01, KSNS^{+}04, XH04, NHS00]. Compartmentalization [GPDvH^{+}03, WSC^{+}01, GH00, WJG^{+}00]. compartmentalize [GGT^{+}02]. Compartmentalized [CSJ^{+}03b, Wel01p, FRM^{+}02]. compartments [Dov03f, GMRYM^{+}02, KMP02b, BHKL01, KLF^{+}00, MGL^{+}00]. Compensation [WML^{+}03, JWJJ00]. Compensatory [SBI^{+}00]. Competent [WWG^{+}00]. competition [EAD^{+}02]. competitive [LRA^{+}02]. Complementation [RS00c]. complete [OSB04]. completion [LeB02b, TFM04, ZSS01]. Complex [CS01, CENMR^{+}01, EKT^{+}00, FSM^{+}01, GKG^{+}01, GRBD01, HP00, JRW^{+}01, KFS^{+}00, KSK^{+}00, LBP00, PSWU00, SZZ^{+}00, SPW00, WKS^{+}00, ASK^{+}03, BGA^{+}04, BBBS04, BBSF01, BCM04, CWA^{+}03, CBW^{+}01, CDB02, CM01, CSM03, CKZ^{+}02, CLSK02, DBB02b, DMA^{+}01, Dov01-33, ERMT^{+}04, FVC04, FCF^{+}01, FBH03, FP02, FGS^{+}02, GYS02, GI02, GNDLS^{+}01, GSP^{+}02b, HPE^{+}01, HC02, HBD^{+}02, IFP^{+}03, IKS^{+}02, JHS^{+}02, JGR^{+}04, KLE^{+}02, KSK01, KYM04, KMLS04, LeB03q, LJK^{+}01, LRF^{+}02b, MOMK03, MRC^{+}02, MYC^{+}02, Mil02, MSN^{+}02, NKM^{+}03, NCMO^{+}02, POH^{+}04, PMGS02, Pfe01, PTH^{+}04, PLC^{+}02, PDR^{+}03, PBT^{+}02, RZB^{+}03, RBE^{+}02, RP03b, RMW03, SPK^{+}01, SAG^{+}01, SV03, SDC^{+}01, SCY01, SST^{+}01, SYVB03, SDL02, TBTN01, TNW^{+}02, UOB^{+}02, WCIN04, WPC^{+}02, WSK^{+}03, WKZ^{+}02, WDFNN04, WDMH03, WWS^{+}03, YMR03, YCK^{+}03, ZW04, AML00, CDM^{+}02, CFC^{+}00, DB00, DSP^{+}01]. Complex [DR00a, FAAS00, sGS01, GZY^{+}00, GMC^{+}00, HLK01, HKK^{+}00, JCR^{+}01, JWJJ00, KHK01, KRR^{+}01, LM00b, MGL^{+}00, MMR^{+}00, MDJF00, MNHR00, NKP^{+}01, PCR^{+}01, PWU00, RSG00, RSG01, RAS^{+}00, RM00, RPE00, SHWH00, SLSR^{+}00, SYH^{+}01, TWBV^{+}01, THK^{+}00, WK01, WMHB^{+}00, WHS00, WSE00, ZKW^{+}00, yZCKA01]. Complex-Mediated [SPW00]. Complex/Cyclosome [GKG^{+}01, KSK01]. Complex/Cyclosome-Dependent [GKG^{+}01]. Complexed [NCUJ^{+}00]. Complexes [LYKH00, SZZ^{+}00, BHW02b, BWV^{+}01, BWK^{+}03, CiKBG03, DBH^{+}01, DC03, EOJ^{+}03, FPM^{+}03, GG04, HSMB02, LMHJ02, LeB04h, LG01, MTG^{+}02, MGP^{+}02a, NSLSK02, OGD03, PRLR02, PG02, XRI^{+}04, YKT^{+}04, YGWN01, ZCW^{+}03, BKD^{+}00, DSH^{+}00, FPSM01, KNIO01, MRM^{+}00, SVG^{+}00, ZLG00]. Complexity [SL00, Ste00]. complicated [Dov02-58]. Component [AP00, CQH^{+}00, MMS00, CMM^{+}02, CWA^{+}03, CK03, GSP^{+}02a, JGW02, MCA^{+}03, WWS^{+}03, YLG^{+}02, YLY^{+}02, dMMBK^{+}02, ATE^{+}01, BKB^{+}01, HJSM00, KHN00, KPB^{+}00, LGK00, LHvdH00, PLH^{+}01, RMG^{+}00, RBB00, SS00, WSE00]. Components [HLK01, DHK^{+}04, GSB^{+}03, HMG03, HBAF^{+}02, IKS^{+}02, PBT^{+}02, PL01, RRM^{+}03, TOM01, CH01, WK01]. composition [SWB03]. Compressions [GJ00]. comprised [ISS^{+}04]. Compromised [CNBWN00]. compromising [ARM02]. Computational [EA03, MM03]. Computer [Néd02]. concentrate [Wel01-35]. Concentrates [MKST00]. Concentration [CGL^{+}01]. Concentrative [MJS02]. concept [Mis01]. concert [KSC^{+}04]. Concomitant [SMSM00]. Condensation [LeB04u, SCLC00, CH03, Dov03z, KLK^{+}04, LHK02, MSM04b, Tum04k, FAAS00, GG01, HSKG00, SKJ^{+}00]. Condensin [ALEH02, CSM04, Wel02p, WDMH03, YK03b, FAAS00, GG01]. Conditional [FPP^{+}02, RBM^{+}00, SGK^{+}02b, SMR^{+}02, CLG^{+}03, KPA^{+}03]. Conditions [KWH^{+}00]. conductance [FS03, PPP^{+}01, WMA^{+}04a, BHKL01]. Conducting [ATE^{+}01]. conducts [Dov04i]. cone [DLY^{+}02, Dov01n, MTW^{+}04, SF01, ZWSC02, FNK^{+}00, qZC01]. cones [Dov02-52, EWD02, GWL03, NK02, RJA^{+}03, SKF02, KSN^{+}01, WBC^{+}00]. Conference [Wel01-28, Wel02-51]. Configuration [MMH^{+}00]. confines [CSO^{+}04]. Conformation [GMRS00, BWW^{+}02, YSGS00]. Conformational [LKM^{+}04, SKS^{+}04, TH00, GHS^{+}03]. conformations [ABCK^{+}03, ALEH02]. confusion [Wel02n]. Congenital [WFQ^{+}00, ZCH^{+}01]. congression [LC01]. conjugates [BHW02c]. conjugating [PPWM04, TBRG01]. Conjugation [KHK01]. Conjunction [Dov04d]. Conjunctions [Dov03j]. Connect [Wel01q]. Connected [HSMJ01, Dov01u, Les01n, LCS^{+}01]. Connection [FRK^{+}01, LeB03a, LeB04-78, Wel03-56]. connectivity [SWH^{+}02]. connects [Tum03i]. Connexin [LeB02x, Dov03-65, LWCKL01, XLH^{+}01]. connexin43 [RDC^{+}04, TLS^{+}01, LTB^{+}00, LWS^{+}00]. connexins [MWSL^{+}03]. conoid [HRM02]. conquers [Wel04-51]. Consequences [Dov02o]. Conserved [HES00, YF00, BRB^{+}01, CLSK02, Kil03, KMS^{+}04, TE01, WCA^{+}03, dSAH02, vBDH03, PHWK^{+}00, SPH^{+}00, SYH^{+}01, WK01]. Conserving [Tum03g]. consistent [MMPO^{+}01]. Consisting [IBKSP01]. constituent [AZ03, SGO^{+}00]. constitute [FMF^{+}04]. Constitutive [HRE^{+}01, JON^{+}03, PWY^{+}03, CSP^{+}00, SGAS00, TSK^{+}00]. Constrains [SBI^{+}00]. constraints [BVH04]. Constricted [SDS00]. constructed [GCR^{+}03]. Constructing [Pfe01]. construction [Wel04h]. Contact [LZC^{+}03, CMC^{+}02, HKP^{+}04, TT04, WKJS^{+}04, ZTK^{+}03, SMI^{+}00, Zwe00]. Contactin [FSGDN^{+}00, GSSP03]. Contactin-Associated [FSGDN^{+}00]. contacts [HPG^{+}02, SLD^{+}02, KHH^{+}01, RZB^{+}01]. contain [EBWC01, MMPO^{+}01]. Containing [HRE^{+}01, KRA^{+}01, RB01, RDH^{+}01, Wel01r, MSMK04, TNW^{+}02, WMG^{+}04, ARK^{+}00, BPS^{+}00, FMMF01, KHH^{+}01, PPR^{+}00, DOB^{+}01, NOM^{+}04]. contains [BCG03, SLG02, WK01]. Content [LAF^{+}00]. continues [LeB04-111]. continuous [GCR^{+}03, MR01a]. contractile [DSG04, Yum01]. contractility [GYL02, WDS^{+}03, RS00c]. Contraction [KKA^{+}01, MB01, IKS^{+}01, RMMP04]. Contrasting [LXL04a]. contribute [ACBG04, LP04a, CLT^{+}01, KGE^{+}00, TSMT00]. contributes [KCG^{+}03, LeB02-37, MRK04, PG02, ZZM^{+}03]. contribution [HBB^{+}04]. contributions [YLQ^{+}02, PMK^{+}00]. Control [ABF^{+}03, HRB^{+}01, YWH04, CHM04, CAW^{+}04, DSH^{+}03, Dov02i, EPH^{+}03, FAT^{+}02, GLSG^{+}02, HCC02b, KRMB03, Las03, MSM^{+}04a, PBD^{+}04, PY03, SKS^{+}04, SMZ^{+}03, SMR^{+}02, SR03, SPA^{+}04, VKB^{+}01, VN04, WSL^{+}01, WSH^{+}04, YDPK04, JPM^{+}00a, KWOP00, LR00, LYMC00, MLLA00, MCG^{+}00, NSW00, PS00, RS00c, RGG00, SHS^{+}00, WGF^{+}00]. controlled [EPN^{+}03, FNZ^{+}03, HdEV^{+}02, LRA^{+}02, LRWB04, MMBM04, SOH^{+}04, SKK^{+}02, DR00b]. Controlling [Dov03k, TV00, MPAP^{+}03, ZHDB04, FMMF01]. Controls [SDS00, ARQ^{+}04, AOJ^{+}04, BS04a, DDL^{+}04, DFJ^{+}02, DKA01, HFG^{+}04, JJM^{+}02, LJK^{+}01, MJV^{+}03, MYO^{+}04, MSS^{+}01, NCGD^{+}03, PDR^{+}03, WBAS04, BMG^{+}01, FWY01, RLC00, WRGK00]. conundrum [DRF02]. Conventional [HTPC04, MSY^{+}04, SLG02]. Convergence [NLRD01, MYO^{+}04]. convergent [KMH^{+}04]. Conversion [FFST01, TAA04]. convertase [BTH^{+}03]. convertases [DMC^{+}03]. Cooh [BHKL01, BCG03, EWSN00, HHS03, XBL^{+}03, CFC^{+}00, FKG00, OLB^{+}00, TMHP00]. Cooh-Terminal [BHKL01, BCG03, HHS03, XBL^{+}03]. cooperate [BWK^{+}03, BGW^{+}04, EBWC01, GCT02, GTBM04, MB01, SCD02, SKR^{+}00]. cooperates [BJB^{+}03]. Cooperation [Wel01s, sKCK^{+}01, LeB03e]. cooperatively [JLK^{+}02]. Coordinate [LR00, LLH^{+}00, OEM^{+}02, CG03]. Coordinated [NGS^{+}01, SCM^{+}00]. coordinately [CDE^{+}03]. coordinates [ITF^{+}02, JB01, LO04, AFB^{+}01]. Coordinating [Dov02p, KLF04]. Coordination [GWBW02, LeB02y, DAV^{+}03, SBS02a]. Cophosphorylation [TSL^{+}03]. COPI [LOS^{+}01, LeB02-41, RAD^{+}02, SSL^{+}00, SDL02, YLG^{+}02, BSW^{+}00]. COPII [HCCB03, MJS02, SSL^{+}00, SMH^{+}02, WMA^{+}04a, XH04, CB00, SKR^{+}00, FS03]. COPII-coated [MJS02]. COPII-dependent [WMA^{+}04a]. Coping [Dov03l]. COPs [LeB02z]. cord [DWM03, GLSG^{+}02, LeB02-47, PMKV01]. core [DIR03, LMHJ02, MPV^{+}01, VCGB^{+}02, WCA^{+}03, XJW^{+}04, KC01, KWO^{+}00, LM00b, WMT^{+}01]. coreceptor [KPL^{+}02, OPP^{+}03, AKFB00, SPMM^{+}00]. corner [LeB02-27]. Corneum [EMW^{+}01]. coronin [HBD^{+}02]. corpse [EC03]. Corralled [Wel04q]. Correct [SHP01, FP02, LeB02-83, REM^{+}02, ZH04]. correcting [HCL^{+}03]. Correction [GFM^{+}04a, GRSL^{+}04a, KBGG04a, LeB04v, Man04b, MWN^{+}04b, QDG^{+}04a, SZvBuH^{+}04a, TR04a, WSWM04a, WMA^{+}04b, WRGK04, YSC^{+}21]. Corrects [LMW^{+}00]. Correlate [AKH00]. Correlates [BRM^{+}00]. correlation [LMVW03]. Correlative [PPM^{+}00]. Cortactin [BGW^{+}04, HKP^{+}04, WKS^{+}00, MOMK03, OM01, MKK^{+}00a]. cortex [YITe03, AC00, HCOC00, LFT^{+}00]. Cortical [FK01, RSK02, EKC^{+}03, FCF^{+}01, IUK04, KHLW02, MB03a, SKGC^{+}03, Yum01, HCOC00, PGV^{+}00]. Cosignaling [NDS^{+}02]. Costameric [RPE00]. costs [LeB02-63]. cotransporter [RLTC^{+}02]. Cotyledon [TOM01, HMRH01]. Counteracted [YF00]. Counterparts [FSBH00]. counters [BMM^{+}02]. couple [WWD03, GHS00]. coupled [CEGZ^{+}04, JBA^{+}01, SGK02a, SDL^{+}03, YSW02, ZJM^{+}02, BC00, MTV^{+}00]. couples [DJT^{+}03, TMA^{+}04, KMiM^{+}01, WSE00]. Coupling [CK00, SGF^{+}00, EPH^{+}02, FAF^{+}04, LWZ03, NYT^{+}03, SAWS02, TSH^{+}04, WSWSL04, AKT01, RMHM00]. covalent [MGP^{+}02a]. Covered [KF00]. cpTatC [CM01]. cPTPRO [SSW^{+}01a]. Crac [FFKC00]. cracked [TEB^{+}03]. Cranial [SC04]. Craniofacial [LWS^{+}00]. Cre1 [RRS03]. created [BC03, PDMO03]. Creatine [HSW00]. Creb [vMZM^{+}00]. Creb-Binding [vMZM^{+}00]. CReP [JON^{+}03]. crest [DRBF03, FLS^{+}04, HBK^{+}02, MRM^{+}04, SC04, XLH^{+}01]. crew [Wel02w]. Cripto [PDL^{+}03]. crisis [Lai03a, Wel02q]. Cristae [Wel02q]. Crithidia [DE01]. Critical [GBY^{+}03, PSK^{+}03, WLW^{+}04, XBL^{+}03, CS03a, DBL^{+}02, GJS^{+}03a, KVC^{+}03, KvHB^{+}01, KFR^{+}04, LPPT^{+}02, SR03, Wel03p, DCM00, FRK^{+}01, MHK^{+}01, SYH^{+}01]. Crk [BGW^{+}04, CK00, CK02, PBL^{+}00, SEM^{+}00]. CRM1 [MSM^{+}04a, RXS^{+}03, BHL^{+}01, LHW^{+}01, Tum03h]. CRM1-dependent [MSM^{+}04a]. Crm1-Mediated [BHL^{+}01, LHW^{+}01]. Crm1p [HKJ00]. Cross [BDR^{+}03, MDJF00, NY00, TCV^{+}00, Wel01t, ESC^{+}01, LZS^{+}03, MSA^{+}03, MWMK04, APLB00, TV00, VDMH01]. Cross-linking [TCV^{+}00, Wel01t, ESC^{+}01, MSA^{+}03, MWMK04]. cross-links [LZS^{+}03]. Cross-Talk [NY00, BDR^{+}03]. Crossed [Wel02r]. crossover [Wel03-43]. Crouching [Tum04f]. crowding [EKB^{+}03]. crucial [FHF^{+}02, SPB^{+}02]. Crumbs [Dov02q, LeB02-27, MWK^{+}02, RML^{+}02, SDMC^{+}04]. CRYP [SSW^{+}01a]. CRYP-2 [SSW^{+}01a]. CRYP-2/ [SSW^{+}01a]. crypt [BvdWD^{+}04, Gla01e]. cryptic [XRP^{+}01]. Crystal [HWBD^{+}01, TDhL^{+}02]. Crystalloid [JPRR00]. Crz1p [PC01]. CSB [vdBCH^{+}04]. Csc [RGK^{+}03]. Csc-1 [RGK^{+}03]. CSF [THG^{+}04, WDL^{+}04]. CSF-1R [WDL^{+}04]. Csk [OEM^{+}02]. Csl [PGS^{+}01]. CSX [AKM^{+}04]. CSX/NKX2 [AKM^{+}04]. CSX/NKX2-5 [AKM^{+}04]. Cta4p [FAT^{+}02]. Cte [HRE^{+}01]. CTL [PBB^{+}04]. CTL-mediated [PBB^{+}04]. cue [SSW^{+}01a]. cues [NH03a, TSMT00]. culprit [Wel02-63, Wel03-63]. Culture [TTR00]. cultured [ABF^{+}03, RL03, ZMGL02]. Cup [Dov03m, Las03, WHAH03, CS04]. Cup-ling [Las03]. cures [LeB04-51]. Current [MZ00]. curvature [FRT^{+}01]. curves [LeB04-88]. cushion [EGC^{+}03, LCG^{+}04]. Cut [Dov03n]. Cut23 [MRC^{+}02]. Cutaneous [THE^{+}00]. Cv1 [YHF^{+}01]. Cvt [NKH^{+}00]. Cvt7p [NKH^{+}00]. Cvt9 [KKS^{+}01]. Cvt9/ [KKS^{+}01]. cxc [FLWMG02]. CXCR4 [KVC^{+}03]. cyclase [CMS^{+}02, ZFH^{+}04]. Cycle [CS01, DSSY00, LGW00, ARQ^{+}04, BMLU02, CW02, CARG^{+}04, FNFL03, KSC02, Lai03e, MWN^{+}04a, MKTW01, PMGS02, QS02, RDS02, RMW03, SKK^{+}02, SCPP02, SRSW04, SAP03, TOTC01, Tum03b, Tum03d, Tum04c, UJL^{+}03, US04, ASP^{+}00, BT00, CP01, CH01, FCL^{+}00, IBKSP01, KR01, LYMC00, MST^{+}00, NS01, PVjL^{+}00]. cycle-dependent [CW02, QS02]. cycles [BS04b, BJ03, GBZ^{+}02, LeB03-68, LeB04-102, MMFS01]. Cyclic [REK^{+}03, BKZ^{+}03]. Cyclin [FKSJ01, GKG^{+}01, LeB03r, MLS^{+}01, SHVR02, YHK^{+}00, dEP01, CARG^{+}04, ERMT^{+}04, KFR^{+}04, RJyH02, Tum04o]. Cyclin-Dependent [MLS^{+}01, SHVR02]. cycling [DAV^{+}03, Pow01b, SSG^{+}02, NKP^{+}01]. Cyclins [DOL^{+}01]. cyclosome [KSK01]. Cyclosome-Dependent [GKG^{+}01]. Cyclosporin [CJ02]. Cyk [JPGR^{+}00]. Cyk-4 [JPGR^{+}00]. Cypher [ZCH^{+}01]. cysteine [RFCD02, SGC^{+}02, IAG^{+}00, NY00, TOM00]. cysteine-rich [RFCD02, SGC^{+}02]. cystic [FS03, LeB04-33, WMA^{+}04a, BHKL01]. cytoarchitecture [BGA^{+}04, CNBWN00]. Cytochrome [WGvA^{+}01, vARP^{+}00, APM^{+}02, BSMS03, DRC^{+}02, DMLK04, MH01, MEV^{+}04, MNT^{+}03, DKJ00, PVL^{+}00]. cytodomain [HBB^{+}02]. Cytokeratins [ToIB^{+}00, MMDC00]. Cytokine [MHE^{+}00, DBL^{+}02, LeB03s, RBD^{+}01, Sre04l, XLGS01, Ern00]. cytokine-independent [XLGS01]. Cytokine-Induced [MHE^{+}00]. cytokinesis [BPC03, BGBG03, DSG04, GJS^{+}03b, GGGN02, HGS^{+}01, KKTP03, KGT^{+}02, LeB02q, LeB02-100, MHlW02, NPS^{+}03, OKC02, Stu04, SAP03, TMG03, TFM04, ZSS01, AHMJ01, EU00, GG01, KNM^{+}00, KR01, Mit01, OSMF00, PSB00, RS00c, SL01, WBG01]. cytokinetic [CG03]. cytokinin [RRS03]. cytomegalovirus [AAM^{+}04]. Cytoplasm [KIO^{+}00, KHK01]. Cytoplasmic [BMS^{+}00b, CBS04, DKA01, HMC^{+}01, KAiK^{+}02, PMBC^{+}00, TCS01, TYA^{+}02, UIY^{+}01, ALP^{+}04, CYC^{+}04, DBB^{+}02a, DW02, DBS^{+}03, GTR^{+}03, GSP^{+}02b, GPZ^{+}02, HMVG02, HT01, KYM04, LKL^{+}03, LYL^{+}04, MTPT02, PPGN^{+}02, RWK^{+}04, RTFW02, SBMB^{+}04, WPC^{+}02, WSF^{+}01, XBL^{+}03, XSK^{+}01, BMD^{+}00, FL00, FK01, GV00, HOvD^{+}00, LVWA01, LW00, MMDC00, RSG00, SJS^{+}00, TNM^{+}00]. cytoskeletal [BFH^{+}01, CLB^{+}03, DB02, FHM^{+}03, FLS^{+}03, HDH02, KSNS^{+}04, KLF04, RCY^{+}03, SPB^{+}02, SSO^{+}03, ZSY^{+}03, BL01, BF01, GG00, GKM^{+}01, KH01, SF01]. Cytoskeleton [KSK^{+}00, PFW^{+}00, AJ01, BN02, BFG^{+}04, CTE^{+}04, CBRBM04, FNZ^{+}03, GJS^{+}03a, GSB^{+}03, GGNK04, GCT^{+}04, LCM02, OEM^{+}02, RRM^{+}03, SKGC^{+}03, VBR^{+}01b, WLO^{+}02, vWJK^{+}03, DBS^{+}01, sGS01, KEGDQ01, KKR^{+}00, MMWC00, NLBK00, OSN^{+}00, QKK00, ZGB01]. Cytoskeleton-like [KKR^{+}00]. cytoskeletons [SS02]. cytosol [BSMS03, ØTW04, YMR03, GSB^{+}00]. Cytosolic [DH02, NHB00, RLC00, HBV^{+}01, JMG04b, RZB^{+}03, SHA^{+}03, TMA^{+}04, VM02, KKS^{+}01, PR00]. cytostatic [TEC^{+}03]. Cytotoxic [SNL^{+}00, SBM^{+}01, Zwe00]. cytotoxicity [IIN^{+}01, ZMGL02].

D [MA00, UTH^{+}02, BLC00, GMD^{+}02, HHOP02, LeB03-83, LeB03-85, MMG^{+}04, PGSE^{+}01, SC01a, SRW^{+}02, UGKT^{+}02]. D-mediated [HHOP02]. D-Titin [MA00]. D1 [DAV^{+}03]. D2 [KFR^{+}04, SCLC00]. D2/Eg7 [SCLC00]. Dad1 [NSLSK02]. DAG [BHW^{+}03]. Dam1 [sKCK^{+}01]. Dam1p [CBW^{+}01, CENMR^{+}01]. Damage [Wel03o, Dov03q, GGT^{+}02, KMG^{+}03, LeB03t, PSD^{+}04a, SCK04, SM03b, SZvBuH^{+}04b, WK04, vdBCH^{+}04, BKI^{+}01, RIDC01, TWS^{+}00a]. damage-induced [SM03b]. damaged [SVLM02]. Dance [Dov03o, Dov03-49]. dancers [SP02]. danger [Wel02-67]. DAP [IBS^{+}02, WKYC02]. DAP-kinase [WKYC02]. DaPKC [SDMC^{+}04]. DaPKC-dependent [SDMC^{+}04]. Darby [HGP^{+}00, AFK^{+}03, FFST01, MBMMA^{+}03, VHLS00]. dark [Dov02x]. data [RO04]. daughter [BPD^{+}04, WKZ^{+}02, Wel03p, PMK^{+}00]. day [LeB02d, Wel04-44]. Dbl [BHNG01]. DC [CdLvM^{+}04]. DC3 [DMLK04]. DC4 [DMLK04]. DCC [MTW^{+}04]. de-inhibition [LeB04-83]. Deacetylase [LeB03t, KMG^{+}03, MSR04, SNF^{+}02]. Dead [Wel01u]. deafness [iKFH^{+}04]. Deamidation [PHWK^{+}00]. Death [CQH^{+}00, Dov01i, Dov03p, FXPT00, LeB04w, LeB04x, LeB04y, Les01a, Wel01v, Wel02s, BFC^{+}02, BSD^{+}01, BMS^{+}03, CDM^{+}02, CJ02, CSJ03a, CSJ^{+}03b, CFM^{+}02, DMH^{+}02, DBY01, DBL^{+}02, Dov03-61, FCM^{+}01, GCT^{+}04, Gri03, HPFG03, HKE^{+}04, IBS^{+}02, LeB02j, LeB02-29, LeB03-46, LeB04-61, LXL^{+}04b, Les01g, LAL^{+}03, MWAM01, NMG04, PKR^{+}02, PBD^{+}04, RMC^{+}02, RBD^{+}01, ST03, SDS^{+}04, TST^{+}03, Wel01-52, Wel02-68, Wel03j, WK04, DKJ00, HLU00, MSO^{+}00, ME00, VCG^{+}00]. debate [LeB04-111]. Debcl [CQH^{+}00]. Dec [MGL^{+}00]. Dec-205 [MGL^{+}00]. Decarboxylase [PSKK^{+}00]. decay [CBS04, Wel04-30]. Decisions [BYLA^{+}01]. Decisive [SEP^{+}01]. decline [BMM^{+}02]. Decoding [LeB02-28]. decondensation [MWHM01]. decrease [KPL^{+}02, ZCW^{+}03]. Decreased [MSV^{+}00, WLPD04]. decreases [CSIK03, HKO03]. DEDD [LeB02-29, LSS^{+}02]. dedicated [LFM^{+}04]. dedifferentiation [LeB04-58]. deep [SKM03, Wel03-66]. deep-orange [SKM03]. defeats [Wel02-68]. defect [NDM^{+}03]. Defective [GSW^{+}00, HWHH01, HBB^{+}02, NOM^{+}04, TDL03, XJW^{+}04, CLG^{+}03, DSG04, GNDLS^{+}01, LFM^{+}04]. Defects [TTHH00, CBG^{+}01, GK03b, KHB^{+}04, MMBB03, WLR01, YIS^{+}03, KHH^{+}01, NHI^{+}00]. Deficiency [IKS^{+}01, ASK^{+}03, iKFH^{+}04, CKS^{+}00, LWS^{+}00]. deficient [HYMS^{+}02, KPL^{+}02, MO01, NHG^{+}03, SGdM^{+}01, WLR01, AKK^{+}00, CMW^{+}01, GBM^{+}00, HWHH01, KdVS^{+}00, MYH^{+}01, CKF^{+}03]. Deficits [LAF^{+}00]. define [FPM^{+}03, SKM03, TCK^{+}03]. defined [NK02]. Defines [ALWR01, KPKY^{+}03, POH^{+}04, ZFH^{+}04, BHY^{+}00]. Defining [BGH^{+}03, KPA^{+}03, RSBE00, Rut00]. Degeneration [AKDS00, IKA^{+}03, iKFH^{+}04, KL04, SZvBuH^{+}04b, HGB^{+}00]. Degradation [CKFH00, DDW^{+}01, Dov02w, FNFL03, LeB03-48, LSS^{+}02, LdVV^{+}02, MGP^{+}02a, SPN^{+}04, TJC^{+}03, TOM01, VM02, BHKL01, GSB^{+}00, iNFK^{+}01]. Degradative [WMG^{+}04, EOB^{+}04]. degrees [KRMB03]. Dek [MRM^{+}00]. delay [MSR04, Wel03-66, dEP01]. delayed [DC03, LWS^{+}00]. delaying [BPKR^{+}02]. delays [HIO^{+}04]. Deleting [Dov01j]. Deletion [CDFT^{+}01, LSW^{+}03]. Delivers [Hol01]. delivery [ALP^{+}04, KBK^{+}03, MHlW02, NCGD^{+}03]. Delta [AMBW04, MGRPVAT02]. demonstrate [SWB03]. Demonstration [CS00]. demoted [Dov03c]. Demyelinating [CMW^{+}01]. Demyelination [AKDS00, ZEW^{+}01]. Dendrimers [GBD^{+}00]. dendrite [HTT^{+}02]. dendrites [HTT^{+}02]. Dendritic [LeB02-30, LFM^{+}04, Tum04g, AAM^{+}04, CdLvM^{+}04, HIN^{+}03, KRS^{+}01, SBC^{+}03, TTH^{+}01, VYW^{+}03, EHM^{+}00, MGL^{+}00, TNK^{+}00]. Dense [MPV^{+}01, VCGB^{+}02, KWO^{+}00, WMT^{+}01]. dense-core [VCGB^{+}02]. Density [CSJ00, BGM03, CKW^{+}03, MTW^{+}02, MPB02, OPP^{+}03, REM^{+}02]. Dentato [WND^{+}00]. DEP [LZC^{+}03]. DEP-1 [LZC^{+}03]. DEP-1/CD148 [LZC^{+}03]. depend [HBH^{+}04]. dependence [HBG^{+}02, PDJ00]. Dependent [AFN00, CWMO00, DZT^{+}00, DFYL00, GKG^{+}01, GV00, HLU00, pHYpXL00, HM00, KLGC^{+}01, MRH^{+}01, MHS^{+}00, MLS^{+}01, RGG00, TRC^{+}00, ALC^{+}03, AFK^{+}03, BS04a, BHNG01, BMLU02, BVH04, CDM^{+}02, CPC^{+}02a, CW02, CHM04, CMC^{+}02, COB01, CSO^{+}04, CM01, DC03, DWB03, DPB03, DBLG02, FHL^{+}03, GCO^{+}04, GHS^{+}03, GJS^{+}03a, GHK^{+}03, HDP^{+}01, HOK^{+}02, HP04, HCC02b, HBC^{+}03, HP03, HNK^{+}03, HGP^{+}04, HPG^{+}02, JAS02, JJM^{+}02, KYS^{+}02, KHC02, KSK01, KJB^{+}02, LCM02, LRF^{+}02b, LC04, LCRS01, LRB^{+}03, MH01, MTT^{+}04, MSM^{+}04a, MHH^{+}03, MMM^{+}04, MSI^{+}03, MCF^{+}02, NL03, NK02, NYT^{+}03, OSNG04, OO02, PDL^{+}03, PC01, PWS^{+}01, QS02, RPS^{+}02, RTFW02, SHB^{+}03, SUT^{+}01, SNF^{+}02, SHM02, SRSW04, SM03a, SHP01, SB03, SSO^{+}03, SHVR02, SDMC^{+}04, SGW^{+}02, SKH03, SEW^{+}01, SC04, TKHR03, TIO^{+}02, UIY^{+}01, VBR^{+}01b, VBH^{+}02, WCIN04, WVY^{+}01, WMA^{+}04a, WSF^{+}01, WW02]. dependent [WWGK02, WDL^{+}04, YK03b, ZCH^{+}02, ZWB04, ZCW^{+}03, ZAE^{+}04, BMS^{+}00a, BHKL01, BKD^{+}00, CP01, FWY01, HCOC00, HBSQ01, HKJ00, HFM^{+}01, IAG^{+}00, LVWA01, LSSL00, MCH^{+}00, MRM^{+}00, MGLPM00, PSKK^{+}00, PGS^{+}01, RBBA00, RM00, SCM^{+}00, SMS^{+}01b, WSE00]. depends [DM02, EKH^{+}03, HKBH03, LTD^{+}02, MZH^{+}02, PSE^{+}03, PL01, SCB02, SDS^{+}04, WPC^{+}01, KHvOD00]. dephosphorylates [SSH^{+}04]. Dephosphorylation [NZHR01, MMBM04, SKK^{+}02]. deplete [LeB03-100]. depleted [MBSR03]. depletion [MH02, SHM02, TFAM^{+}04]. Depolarization [DKJ00]. Depolymerization [qZC01]. Depolymerizes [WSW^{+}00]. depolymerizing [GK04, OWW02, GOL^{+}01]. deprivation [JMB^{+}04]. deprived [ERS^{+}04, YJS^{+}03, CJ02, CSJ03a]. DER-independent [HCC02b]. Deregulation [ERMT^{+}04]. Derivatives [DT00]. Derived [DFYL00, BSW^{+}04, KVC^{+}03, KWS^{+}02, WNM^{+}03, XJW^{+}04, GBM^{+}00, LQPC^{+}00, MNHR00, TTP^{+}01, TOM00]. dermopathy [HvdHG^{+}03]. des-acyl [BFC^{+}02]. desensitization [MRT^{+}01]. Desensitizing [LeB02-31]. Design [Vin04]. designed [Dov02-70]. Desmin [MMWC00]. desmocollin [CBG^{+}01]. Desmoglein [EMW^{+}01]. desmosomal [BGH^{+}03]. Desmosome [HHSV00]. Desmosomes [Gla01c, LeB03u, GSN^{+}04]. despite [SGdM^{+}01]. destabilizer [vBDH03]. destabilizing [SHCM03]. Destined [KKS^{+}01]. Destroy [Wel04r]. Destroyed [dEP01]. destruction [LeB04-36, Pow01g, RJyH02, Wel02-58, Wel03-60, RSG01, Wel01v]. Desulfation [LeB03v]. desumoylating [ZWB04]. detaches [CACL03]. detachment [Wel04-66, RRL^{+}00]. detectable [MPB02]. Detection [SN04]. detector [Kel03]. detergent [NK02]. detergent-resistant [NK02]. Determinants [LDS^{+}00, SV03]. Determination [SMSF00]. determine [ERS^{+}04, RFLT02, vdFKK^{+}02, BMS^{+}00b]. determined [NG02, SK01, PN00]. determines [CBRBM04, HTRK02, KYS^{+}02, Tum03e, DOL^{+}01, DAC00]. determining [RJA^{+}03]. deubiquitinating [WSWM04b]. deubiquitination [LO04]. develop [SDML04]. Developing [CBK^{+}00, MDP02, MEK^{+}04, SYY^{+}03, SwZK^{+}02, ECV^{+}00, ECO^{+}00, MHE^{+}00]. development [AR03, CDE^{+}03, CROfC04, Dov02g, DSB^{+}01, FJK^{+}04, FSCF^{+}03, GYS02, GQI^{+}02, HBK^{+}02, HT01, HSC01, ICK^{+}04, JBA^{+}01, KPKY^{+}03, LCG^{+}04, NH03b, PMKV01, SLB^{+}01, STA03, SGK^{+}02b, SHPY02, WBU03, Zam04, IEJ^{+}01, IHK^{+}00, LLH^{+}01, MHW^{+}00, ME00, Nob00, ToIB^{+}00, TSMT00, VAHV00]. developmental [DRF02, DFJ^{+}02, MWAM01, WVY^{+}01, YIS^{+}03]. Developmentally [PTH^{+}04]. DGDG [JMB^{+}04]. DHHC [RFCD02]. di-acidic [WMA^{+}04a]. Dia [LeB02-81]. Diabetes [Wel04s, BFSO^{+}04, KBK^{+}03]. diabetes-linked [BFSO^{+}04]. Diablo [ESH^{+}01, RDP03]. Diacylglycerol [TP01, LRA^{+}02, LPT03, SJIM01]. diapedesis [CS04]. Diaphanous [GCT02]. Dibasic [SMSM00]. Dictyostelium [Dov02-38, FSTC02, FMMF01, JRW^{+}01, LAL^{+}03, RS00c, Yum01, vEWS^{+}01]. did [Wel02-29]. dies [Dov04p]. Differences [NWT^{+}01]. Different [HF03, KRMB03, PR02, YB01, vdFKK^{+}02, ALP^{+}04, ALEH02, CFB^{+}03, Dov01-31, EAD^{+}02, GOV^{+}02, GFGP03, LOS^{+}01, MTPT02, MSMK04, NMH^{+}04, NGKH02, PMU^{+}02, SLG02, UML^{+}03, BGFJ01, CDTW00, DSH^{+}00]. Differential [CAB00, CS03b, ECO^{+}00, HFK^{+}03, HMH^{+}03, HBB^{+}04, HBG^{+}02, IOIF^{+}04, LWDH01, MSM^{+}01, MB03c, SCTM00, TPW^{+}04, WGF^{+}00, SGK^{+}02b]. differentially [DSB^{+}02, LCS^{+}02, SLG02, TNBH01]. differentiate [EGS^{+}04]. differentiated [BLP^{+}02, CSP^{+}04, NDM^{+}03, PMKV01, GH00, MCG^{+}00]. Differentiation [WSF^{+}01, AKM^{+}04, ARM02, ANC^{+}02, BBDK^{+}04, BPKR^{+}02, CNHK02, CS03a, CBG^{+}01, DLPB03, DLT^{+}02, EPH^{+}03, FLWMG02, FLS^{+}04, FAF^{+}04, FKH^{+}04, GCG^{+}01, GLSG^{+}02, GSB^{+}01, HViV^{+}02, HNK^{+}03, HPS^{+}04, JZ02, JBA^{+}01, LHC^{+}02, LA04, MRM^{+}04, MSJ^{+}02, MTB^{+}02, MSS^{+}01, PGSE^{+}01, RDNB02, SK04, SCD02, SSRX04, TJS^{+}04, WVBY^{+}03, Wel02-49, WCTU02, WDS^{+}03, WLPD04, BMKA01, CSD00, FHP00, HVT^{+}00, HLZW00, IEJ^{+}01, MSV^{+}00, MHK^{+}01, PCC^{+}00, SDDS00, SEI^{+}00, WRCU00, YCX^{+}01]. Differentiation- [WSF^{+}01]. Differently [SKT^{+}03]. differs [PLP02]. difficulties [LeB02i]. diffuse [NSLSK02, PKF^{+}00]. diffusion [FRM^{+}02]. digest [Wel04v]. digits [Jeg01c]. dilation [DW02]. dileucine [JKB^{+}03]. dileucine-based [JKB^{+}03]. dimensional [BS04a, CSO^{+}04, HZS^{+}01, WDS^{+}03, HAP^{+}00]. dimer [ABCK^{+}03]. dimerization [WFC^{+}02, ZN01]. dimerizes [DLT^{+}02]. dimethylarginine [BCB^{+}02]. Dinucleotide [BPMG00, SNF^{+}02]. dinucleotide-dependent [SNF^{+}02]. diphtheria [RZB^{+}03]. Direct [FLS^{+}03, HBD^{+}02, LMVW03, LSSL00, MSC^{+}03, MFF^{+}03, NHB00, PMPH03, SD00, TIS^{+}01, YK03a, CROfC04, FRT^{+}01, GRSL^{+}04b, VT04, FSK^{+}00]. Directed [RSBE00, DBS^{+}03, GLS^{+}03, NOS^{+}01, CDFT^{+}01, KGvdG^{+}00]. direction [Dov02-66, HTRK02, LeB02-61]. directional [NH03a, SCTF04, SWE^{+}03, vEWS^{+}01]. directionality [MSD^{+}04]. directionally [MWL01]. directly [CDK04, GCH03, SVI^{+}04, ZZM^{+}03]. directs [DKAH04, KSS^{+}03, KIK^{+}00a]. Dis1 [vBDH03]. Dis1/XMAP215 [vBDH03]. disaccharidase [TM04]. disassemble [ITM^{+}04, DOL^{+}01]. Disassembly [SBS02a, BFG^{+}04, BTVB03, HGP^{+}04, LRD^{+}03, MSN^{+}02, OPP^{+}03, SKK^{+}04, UJL^{+}03, KTY^{+}00, LWDH01]. Disc [BMM^{+}01]. discoideum [RS00c]. discontinuous [HDP^{+}01]. Discovered [RS00c]. Discovery [WLR01]. discrepancy [CBC^{+}01]. Discs [XRI^{+}04, RML^{+}02]. disease [BXR^{+}02, CLM^{+}03, DPO^{+}04, Dov02l, Dov02-49, Wel02-39, Wel03q, CdBB^{+}01, CNBWN00, FSBH00, PDV^{+}00]. diseases [KDH^{+}04, Wel02-63, SJA^{+}00]. Disgorging [Dov01k]. Dishevelled [SSP^{+}03, KMS00, TN00b]. Dishevelled-1 [KMS00]. Disjunction [AMEC01]. Disk [EHZ^{+}01]. Dislocation [VM02]. disorder [LeB04-51, RKF^{+}04]. Disorders [ALWR01, LdVV^{+}02]. disorganized [MKR01]. Dispensable [MCB00a, GI02, MCBB^{+}04, WPC^{+}02, ZW04]. displacement [BFH^{+}01, KRF^{+}03, MKTW01]. display [ALEH02, NGKH02, SKT^{+}03]. Disrupt [HDJ00, ARM02, GOV^{+}03, FL00, HHHJ00, NS01]. disrupted [LKLD04, TJS^{+}04, TTHH00]. disrupting [WKYC02]. Disruption [BFH^{+}01, BBP^{+}04, KWSK^{+}04, MSHG00, FPP^{+}02, HvdHG^{+}03, WCBC04, WDL^{+}04, XRH^{+}03, LCM00, WPM^{+}00]. disrupts [DWB03, DW02, KFO04, LWCKL01, MKD^{+}01, RWK^{+}04, LM00a, TAD^{+}00]. Dissecting [KM00]. Dissection [ADKK00, MYH^{+}01, LHK02]. dissociates [BIC^{+}03]. dissolution [SAH^{+}03]. Distal [DSP^{+}01, SLT^{+}01]. Distantly [CW01]. Distinct [ABCK^{+}03, DJ03, GG04, GBJ01, OLB^{+}00, PMU^{+}02, QS02, RTM^{+}01, RMC^{+}02, SWK^{+}04, SDN^{+}00, TYY^{+}00, TWS^{+}04, VKB^{+}01, VBR^{+}01a, YLQ^{+}02, YCK^{+}03, ZTK^{+}03, ASYL04, CROfC04, DPB03, FPM^{+}03, GMRYM^{+}02, GFGP03, GBMA04, KEHAM^{+}02, LRWB04, MSN^{+}02, OMM^{+}03, PMPH03, SKT^{+}03, SWG^{+}03, SKM03, TOM01, YPN^{+}04, ADKK00, BKD^{+}00, BSS00, CDTW00, HOvD^{+}00, HW00b, JMG01, KNIO01, LR00, NM00, TMK^{+}00]. Distribution [BRM^{+}00, FPSM01, GJB^{+}00, vdHvODML^{+}00, ACP^{+}02, BMS^{+}03, LCSG03, LCRS01, MHT^{+}04, MSM^{+}04a, PDMO03, REM^{+}02, SM03a, TCK^{+}03, VTGT^{+}03, YWH04, vdFKK^{+}02, DKMW00, EMW^{+}01, HCR^{+}01, MMWC00, PHWK^{+}00, RSG01, SPC00, TSDS00]. disulfide [ØTW04, TR02, NWT^{+}01, TV00]. divergent [CKS^{+}04, ZGN^{+}04]. Diversification [IHK^{+}00]. diversify [Wel04f]. diverts [SPN^{+}04]. divide [Sre04m]. divided [LeB03-39, SCS^{+}00]. dividing [LXL04a]. Division [LeB04z, XHG^{+}00, GKN^{+}03, IOIF^{+}04, LeB03w, LeB03-40, LeB04-69, Les01k, LG02a, YOK^{+}03, RCL^{+}00, TN00a, VMO01, vdB00]. divisions [CSM04, Tum04-29, YCK^{+}03]. Dlg1 [XRI^{+}04]. DMPK [ARM02]. DNA [BKI^{+}01, CLB^{+}03, CLM^{+}03, CLB^{+}02, CNJ01, CH03, DB00, Dov03q, EKdM^{+}04, FHJW^{+}01, FNKH02, FKSJ01, GBD^{+}00, HSMJ01, KMG^{+}03, KKPB03, LeB03w, LeB03t, LeB03-59, LeB03-85, LeB04-40, LDS^{+}00, LRW^{+}00, LKLD04, MPG^{+}02, MCG^{+}00, MEFC03, MN03, MSHG00, iONOM02, RIDC01, SCMH00, SBB00, SMS00, SCK04, SVLM02, SM03b, TWS^{+}00a, Tum04u, VTH^{+}02, Wel02-50, WK04, YDPK04, vdBCH^{+}04]. DNA-free [LeB03w]. DNase [LeB04x, LXL04a]. Dnm1p [MMS00, TN00a, TONN02]. DNup214 [RXS^{+}03]. DNup88 [RXS^{+}03]. Do [Dov02r, SSL^{+}00, Wel04u, ERS^{+}04, SSM^{+}04, Sre04n, Tum04i, BHBJ00]. Doa4 [LO04]. docked [FJM^{+}04]. Docking [PSWU00, BYMS^{+}02, FMJG04, GQI^{+}02, WMCW03, WSK^{+}03, PWU00, WSE00]. Does [Wel02t, RGM^{+}02, VBH^{+}02, WKJS^{+}04, CLT^{+}01, CFC^{+}00, SRHV00].

Doing [Wel03r]. Dok [GQI^{+}02, GSB^{+}01]. dok-4 [GSB^{+}01]. dok-5 [GSB^{+}01]. Dok-related [GQI^{+}02]. Dok1 [WMJ^{+}04]. Domain [ACE^{+}01, BRM^{+}00, EM00, KLGC^{+}01, RB01, RDH^{+}01, AMF01, BWRT03, BDR^{+}03, BMY^{+}01, BKD^{+}04, CYC^{+}04, DBB^{+}02a, DKAH04, FMF^{+}04, FJM^{+}04, GOV^{+}02, GTR^{+}03, GGNK04, HIE^{+}01, HHS03, HT01, HDL02, KNI^{+}04, KVM03, KJK^{+}03, LRF^{+}02a, MHT^{+}04, OKM^{+}01, ORZ^{+}04, RWCC01, RZB^{+}03, RFCD02, SGC^{+}02, SHKS02, WMCW03, XBL^{+}03, XSK^{+}01, XLGS01, ZCH^{+}01, AKW00, CDTW00, CFC^{+}00, DVE^{+}00, FKI^{+}01, FMMF01, GV00, HCD^{+}00, IAG^{+}00, MMG^{+}01, NCUJ^{+}00, PLL^{+}00, RMMC01, SGG01, SME^{+}00, SMSM00, TNM^{+}00, VAHV00, BHW^{+}02a, CHM04]. domain-associated [DBB^{+}02a]. domain-selective [FMF^{+}04]. Domains [HLB^{+}00, JRW^{+}01, NHB00, SDN^{+}00, ZHDB04, vMZM^{+}00, ASYL04, BBBS04, BC03, DJ03, EBWC01, FHM^{+}03, FPM^{+}03, GMLM^{+}04, LMGM^{+}02, MAG^{+}04, NZA^{+}01, RCY^{+}03, SLR^{+}03, SJB^{+}03, TNBH01, UML^{+}03, WPS^{+}01, AH01, BMD^{+}00, DSH^{+}00, MLLA00, SKT^{+}00]. Dominant [CMS^{+}02, MWSL^{+}03, NS01, GK03b, HOK^{+}02, KBK^{+}03, CNBWN00, FWM^{+}01, GBON00, PLL^{+}01, UIK^{+}01]. Dominant-Interfering [NS01]. dominos [Wel03-52]. don [Dov03-44]. Don’t [Wel04v, Wel04-34]. door [Wel02b]. doors [Tum04j]. Dormant [Wel04w]. dorsal [SLB02, KGE^{+}00]. Dosage [JWJJ00]. dots [LeB02m, LeB02x]. double [BM02, DJ03, SCMH00]. double-stranded [BM02, DJ03]. doublecortin [TSH^{+}04]. doublet [MR01a]. doublets [CTE^{+}04]. Down [MGRPVAT02, BJM^{+}02, Dov02e, Dov03e, LeB02h, LeB02v, LeB03-54, LeB03-91, LeB03-92, LeB04x, LW03, RLTC^{+}02, SRD^{+}02, Wel01-43, Wel03w, Wor03]. down-regulates [LW03, RLTC^{+}02]. down-regulating [SRD^{+}02]. Down-regulation [MGRPVAT02, BJM^{+}02]. Downregulates [ZEtK^{+}00]. Downregulation [DGB^{+}00, KMiM^{+}01, LM00a, PFM^{+}00, RRL^{+}00]. Downs [Krä00]. downstream [APM^{+}02, CGM^{+}02, KMH^{+}04, LRWB04, PDR^{+}03, SC01a, WHP^{+}02, WBWS03, FMMF01]. dp115 [KR03]. dragging [BHIWH01]. dragon [Tum04f]. Dral [SME^{+}00]. drive [SH02]. Driven [DLXP00, GYL02, KB04, MRK04]. drivers [Gla01a]. drives [HMC^{+}01, MN04, MPB^{+}04, MSN^{+}02]. dronc [CDK04, CDM^{+}02]. droplets [ORZ^{+}04, FKI^{+}01, OPZ^{+}01, vM01]. Drosophila [BC00, CA00, DWD^{+}00, GG01, GMD^{+}02, KHC02, LMDS^{+}03, MCA^{+}03, RRSV02, RAS^{+}03, WRGK00, ZGB01, AMEC01, AW00, BHNG01, BYLA^{+}01, BGBG03, BC03, CDM^{+}02, CDK04, CNH^{+}02, CHM04, CME^{+}02, CQH^{+}00, DTO^{+}01, DRC^{+}02, DMLK04, FSBH00, GF03, GG00, GLJP01, GCR^{+}03, GMC^{+}00, HES00, HP04, HKK^{+}00, HC02, HW00b, IdCAS^{+}00, IOIF^{+}04, KRU^{+}04, KGE^{+}00, KR03, KNK^{+}01, Las00, LW00, LM00b, Lit00, MDF01, MCBB^{+}04, MWK^{+}02, ML04, MMC00, MS00, OSMF00, PLR03, PFB^{+}03, RKKP02, RBV00, RRM^{+}03, RWSV03, RXS^{+}03, RSK02, STA03, SBB00, SFV^{+}00, SDMC^{+}04, SKM03, SK01, TSMT00, TUK03, TCV^{+}00, VCG^{+}00, WVY^{+}01, WBG01, WS00, WMHB^{+}00, WSH^{+}04, WWBGG03, YCK^{+}03, YMM^{+}00, ZCH^{+}02, ZHDB04, ZRDG02]. DRP [IBS^{+}02]. DRP-1 [IBS^{+}02]. Drp1 [KLG^{+}02]. drug [ERS^{+}04, Wel04-68, CS00]. drug-treated [ERS^{+}04]. drum [Dov04n]. Dual [DSSY00, DRBF03, EHCC^{+}00, KWH^{+}00, MJG03, MOMK03, SAWS02, SRW^{+}02, SHPY02, HIG^{+}01, LeB02-32, RKJL03, GVT^{+}00]. dual-functional [RKJL03]. Dual-wavelength [SAWS02]. duct [HViV^{+}02, LLH^{+}01]. Duct-like [LLH^{+}01]. ductal [JZ02]. due [BFH^{+}01, KHB^{+}04, WBP^{+}03]. Duo1p [CENMR^{+}01]. Duo1p/Dam1p [CENMR^{+}01]. Duox [ESC^{+}01]. duplication [HHM^{+}04, Kil03]. during [AMG^{+}01, ASP^{+}00, ASMW01, AML00, ATF^{+}04, BIC^{+}03, BPC03, BET^{+}03, BLP^{+}02, BGW^{+}04, BSW^{+}00, BJ03, BGBG03, BKDH01, CDM^{+}02, CS01, CK00, CAB00, CDEM00, DSG04, DSH^{+}00, DBB^{+}02a, EGWK^{+}01, FHJW^{+}01, FBH03, FNK^{+}00, GG01, GMLM^{+}04, HW00a, HFK^{+}03, HT01, HYMS^{+}02, IUK04, IBS^{+}02, IOLA^{+}00, IHK^{+}00, JBA^{+}01, JKG^{+}02, sKCK^{+}01, KLG^{+}02, KAC^{+}04, KSBE03, KRS^{+}02, KLT^{+}03, KLZ04, KFO04, LLP^{+}02, LMVW03, LWDH01, LW00, LSS^{+}02, LOC03, LGM^{+}04, LCS^{+}01, LCSG03, LKLD04, LCG^{+}04, MH02, MB03a, MRK04, MMR^{+}00, MJG03, MDF01, MBS^{+}01, MWM^{+}02, MWMK04, MYO^{+}04, MGP^{+}02a, MMBM04, NH03b, NSL^{+}01, Nel00, NGS^{+}01, NCGD^{+}03, PMSB01, PCB^{+}03, PVjL^{+}00, PSK^{+}03, PDJ00, PHM^{+}02, RME^{+}00, RH00, RDP03, RKR^{+}03, RGG03, RDC^{+}04, RRM^{+}03, RCL^{+}00, RMMP04, SHS^{+}00, SHB^{+}03, SD00, SGK^{+}02b, SRC^{+}01, SRL^{+}04, SWG^{+}03, SWE^{+}03, SDD04]. during [SXD^{+}03, TSMT00, TMG03, TRW^{+}00, TONN02, TN00b, TWS^{+}04, TBW^{+}04, UJL^{+}03, VMK^{+}03, VAHV00, WBG01, WWJ^{+}00, WSC^{+}03, WSK^{+}03, WSWM04b, WGvA^{+}01, WHB04, WWGK02, WC03, WMJ^{+}04, WLWB01, WZB^{+}01b, WWBGG03, WK04, YOK^{+}03, Yum01, Zam04]. duty [MKTW01]. dwarfism [HIO^{+}04, Wel01o]. dye [Dov03-58]. Dying [LL03a, LeB03-103, LeB04-109, Wel04-45]. Dynacortin [RS00c]. Dynactin [DSV^{+}03, Del03, RH00, GK04, HMVG02, HMC^{+}01, LKM^{+}04, LeB04-64, QS02, TDFV02]. Dynactin-Dependent [MHS^{+}00]. Dynamic [BYMS^{+}02, BT00, FNZ^{+}03, GMLM^{+}04, JdDD03, KCL^{+}00, RME^{+}00, SKGC^{+}03, SCK04, TNM^{+}03, Bir04e, BKS^{+}02, DBLG02, GWL03, HC02, MPV^{+}01, SKF02, WSWSL04, Wel03x, dVKS04, AKH00, BKB^{+}01, YYM^{+}01]. dynamically [BHK^{+}02, DSP^{+}01, SSN01]. Dynamics [CLB^{+}02, KNR^{+}04, LCM02, LRW^{+}00, OTB03, ASS02, CEGZ^{+}04, CKS^{+}04, DLY^{+}02, FAT^{+}02, GCT02, HMG03, IUK04, KBWG02, KAC^{+}04, KKK^{+}02, LXL04a, LG01, MSC^{+}03, MSA^{+}03, MWMK04, OO02, RRSV02, SKT^{+}03, TT04, TWS^{+}04, VGL03, WBWS03, Yum01, BD00, DE01, DMO00, FWY01, HLRG01, HHF^{+}00, LWDH01, LFT^{+}00, PCK^{+}00, PMSB01, SEI^{+}00, TB00b]. Dynamin [Eri00, LeB04-27, SDS00, TSL^{+}03, WWS^{+}03, FSD00, HvdKDS01, KEGDQ01, MKK^{+}00a, OSN^{+}00, TN00a, WWG^{+}00]. dynamin- [TST^{+}03]. Dynamin-2 [FSD00]. dynamin-related [WWS^{+}03]. dynamite [LeB04-28]. Dynein [GWBW00, KHN00, LeB02-33, MHS^{+}00, RH00, DBS^{+}03, HMVG02, HMC^{+}01, LYL^{+}04, QS02, RP03b, RTFW02, TDFV02, TSH^{+}04, Xia03, FK01, GK04, HCOC00, LVWA01, TCS01, YDRS01]. dynein-dependent [RTFW02]. Dynein-Mediated [GWBW00, Xia03]. Dynein/ [GK04]. Dynein/Dynactin [MHS^{+}00, HMC^{+}01]. Dynein/Dynactin-Dependent [MHS^{+}00]. Dyneins [PS00]. Dysfunction [LWS^{+}00]. Dysregulation [LCGR00]. dystrobrevin [GAC^{+}03]. dystroglycan [Jeg01b, KSM^{+}01, SST^{+}01, BMF00, JCR^{+}01]. Dystrophic [Wel04x, VMS^{+}02, BMF00, BWN^{+}01, LMW^{+}00]. Dystrophies [CZBH00]. Dystrophin [CFC^{+}00, RPE00, TTP^{+}01]. Dystrophy [RRM^{+}00, Dov01-30, KSNS^{+}04, TCP^{+}03, WST01, BWN^{+}01, Cha00, CDFT^{+}01, vdVWS^{+}00].

E-Cadherin [GWG01, LeB02-34, LeB03x, LeB04-29, SEW^{+}01, CSSBY^{+}03, IDvH^{+}02, ISB^{+}04, PLR03, PGSE^{+}01, WG03]. E-cadherin- [SBG^{+}04]. E-cadherin-mediated [CTE^{+}04, HKP^{+}04]. E10 [TGM^{+}01]. E1A [BLP^{+}02]. E2f [GH00]. E2F4 [OBG^{+}03]. E2F4/5 [OBG^{+}03]. E3 [CSG01, YRC^{+}04]. E3karp [RB01]. E4 [LCGR00]. E4orf4 [KSK01]. E5 [SSA^{+}00]. each [CGBL^{+}02, BMD^{+}00]. early [DLT^{+}02, Dov03q, EGWK^{+}01, HCC02a, HBG^{+}02, KLK^{+}04, KHLW02, LeB02-56, LCI^{+}01, RRM^{+}03, RMC^{+}02, RSK02, RKF^{+}04, SYVB03, WSC^{+}03, LGP00, MTG^{+}02, MSH^{+}00, NS00, RCL^{+}00, SCMH00, TCR00, WAPB^{+}00, WJG^{+}00, YSS^{+}01]. Early/ [MTG^{+}02]. East [LeB04-107]. eat [Dov01-36]. EB1 [RRSV02, TB00b]. EBA [GTR^{+}03]. EBA-175 [GTR^{+}03]. EBP [HRB^{+}01]. Ebp50 [RB01]. Ecdysone [CDM^{+}02, CDK04]. Ecdysone-induced [CDM^{+}02]. ECM [GMRYM^{+}02, LeB02-35]. Ectodomain [MGAL^{+}01, THE^{+}00, BBR04, SWK^{+}04, YIS^{+}03]. Ectodomains [FWP^{+}00]. EDA [MCG^{+}03]. Edge [ZBB^{+}00, EWD02, Mil02, SCTF04, WBWS03, qZC01]. editing [DJ03]. Editor [Mel00b]. EEA1 [HCC02a]. EEA1-enriched [HCC02a]. effector [DSH^{+}03, DRP^{+}03, DBL^{+}02, KSC^{+}04, PDR^{+}03, STA^{+}01, SPK^{+}01, SDL02, WE02, HLU00, NCUJ^{+}00]. effectors [FBG^{+}01, HS04, TEB^{+}03, DCM00]. Effects [BHFL01, KMK^{+}02, PDMO03, TNMM03, TTH^{+}01, YHZ^{+}01, GLA00]. efficiency [BWK^{+}03]. efficient [BPKK01, GAT^{+}03, GWL03, Hua02, TSDS00, WJG^{+}00]. efflux [MW04, CS00]. Eg5 [KMCM00, KM01, MPB^{+}04]. Eg7 [SCLC00]. EGF [STJ^{+}01, DFJ^{+}02, KYBS03, LBS^{+}02, MKD^{+}01, MSD^{+}04, NMHH03, SHB^{+}03, STJ^{+}01, YIS^{+}03, YSW02, SFSD00, ZBB^{+}00]. Egf-Induced [ZBB^{+}00]. EGF-like [SHB^{+}03]. EGF-R [MKD^{+}01]. EGFR [Jeg01e, SWK^{+}04]. Egg [PMBC^{+}00, SEM^{+}00, MPB^{+}04, SHCM03, SVLM02, SM03b, WDMH03, AKH00, BGFJ01, BKDH01, WSDW^{+}00]. Eggs [LeB03y, PMBC^{+}00, HTPC04, Wel02f, MGLPM00]. EGL [JJM^{+}02]. EGL-19 [JJM^{+}02]. Egr1 [WXQ^{+}00]. eIF [CBL^{+}02]. eIF-4E [CBL^{+}02]. eIF2 [NZHR01]. Eif4e [DLS00, WHAH03]. ejection [PRS^{+}04]. elastic [SMW^{+}03]. Electrical [Wel02u]. electrolyte [TKB^{+}04]. Electromechanical [RMHM00]. Electron [McI01, MÖS^{+}00, LeB03l, PPM^{+}00]. electrotaxis [ZJM^{+}02]. elegans [ACE^{+}01, AOJ^{+}04, AP00, BWV^{+}01, BHL^{+}03, BKD^{+}04, COB01, GSW^{+}00, GGGN02, HKHO01, HOK^{+}02, HEW^{+}01, HMAM01, JJM^{+}02, LHvdH00, MR01b, MCB00a, NM02, OMM^{+}03, ODR^{+}01, Ono01, PCB^{+}03, RMG^{+}00, RBW^{+}02, RDS02, SB03, UCY^{+}02]. Element [HRE^{+}01, MKS^{+}02, TFAM^{+}04]. elements [EOJ^{+}03, RRB^{+}01]. Elevated [EGC^{+}03, LAF^{+}00]. Elevating [TBRG01]. elevation [BMS^{+}03]. elicit [MW04]. Elimination [Wel04y, FMF^{+}04]. Elk [ASAJ01, SBMB^{+}04]. Elk-1 [ASAJ01, SBMB^{+}04]. elongate [Dov02-34, MDF01]. elongation [ABF^{+}03, CME^{+}02, HTT^{+}02, LZS^{+}03, SHP01, vdBCH^{+}04, ASP^{+}00, JSCR01, MSHG00, SHHH01, TTHH00]. elusive [Tum03b, LS00]. embryo [GGGN02, SSM^{+}04, GBD^{+}00, LGP00, LW00]. Embryoid [LCS^{+}01]. Embryonic [NHI^{+}00, BSL^{+}01, COB01, CDE^{+}03, FSCF^{+}03, HViV^{+}02, KPL^{+}02, LHC^{+}02, PDL^{+}03, SPB^{+}02, SSRX04, WNM^{+}03, XJW^{+}04, MYH^{+}01, ME00, RCL^{+}00]. embryos [KHLW02, RSK02, SSP^{+}03, MSFH00, WS00]. emerging [AMG^{+}01]. Emp24 [MNHR00]. Ena [KSK^{+}00]. Ena/Vasodilator [KSK^{+}00]. Ena/Vasodilator-Stimulated [KSK^{+}00]. Enabling [LeB02-36]. Encephalomyelitis [SEP^{+}01]. encode [VTGT^{+}03]. encoded [AZP^{+}02, CSG01, SRW^{+}04, VZTN03, RBB00]. encodes [BWV^{+}01, AHMJ01, HJSM00, Ono01, RMG^{+}00, SJ01a]. Encoding [AW00, SPH^{+}00]. end [FVC04, FFSF03, GK03b, KCG^{+}03, KKTP03, LeB03-57, NOS^{+}01, SGdM^{+}01, Xia03, vBDH03, SC01b]. Endo180 [ELNA^{+}03, SWE^{+}03]. endocannabinoid [BHW^{+}03, WWD03]. endocardial [EGC^{+}03]. endochondral [NGS^{+}01]. endocytic [ALP^{+}04, LS02, PBP^{+}01, PDR^{+}03, SGK02a, SKGC^{+}03, BMS^{+}00a, SPS^{+}00, VPP^{+}01, WAPB^{+}00]. Endocytosing [Gri03]. Endocytosis [CSJ00, ISB^{+}04, QK00, SDS00, SPMM^{+}00, Wel04z, ALC^{+}03, CNH^{+}02, CHM04, CS02, CS03b, CSG01, EGWK^{+}01, GRSL^{+}04b, HHOP02, JK01, LeB02-48, LeB04-29, LeB04-68, LeB04-95, MBMMA^{+}03, MBSR03, NL03, NB03b, PWY^{+}03, PLC^{+}02, RWH02, SKK^{+}02, TSL^{+}03, TNBH01, Wel04-69, WZB^{+}01b, dSAH02, CSP^{+}00, DZT^{+}00, FNK^{+}00, KEGDQ01, MGL^{+}00, QKK00, SBI^{+}00]. Endocytotic [FMF^{+}04]. endoderm [SSRX04]. endogenous [BMS^{+}03, GYL02, LMHJ02, PSK^{+}03, CDWB01]. Endophilin [KJY04, FRT^{+}01]. Endoplasmic [GSB^{+}00, BSMS03, EWM^{+}04, FAT^{+}02, HSMB02, HKE^{+}04, HIT^{+}02, JCPWS01, KOS^{+}04, MJG03, NSLSK02, PDMO03, SRB^{+}04, VKB^{+}01, WH03, WWGK02, ZLH^{+}03, AFB^{+}01, CGL^{+}01, DR00b, KSF^{+}00, KKL^{+}01, MNHR00, NBWB^{+}00, NSW00, iNFK^{+}01, OPZ^{+}01, PVL^{+}00, PR00, PGV^{+}00, SSN01, TOM00, TH00, WGP^{+}00]. endoprotease [MHH^{+}03]. Endoproteolytic [yZCKA01]. Endorepellin [BFG^{+}04]. endoribonuclease [TCZ^{+}03]. Endosomal [WFT^{+}01, BHW02c, Dov02s, DKAH04, MBH^{+}02, POH^{+}04, PMP^{+}03, PFSG03, RRM^{+}03, Sea04, SKM03, SSG^{+}02, SYVB03, WWBGG03]. Endosomal/Lysosomal [WFT^{+}01]. Endosome [NHB00, BWK^{+}03, MCU04, SYVB03, MSH^{+}00, PMB^{+}00, WAPB^{+}00]. endosome-associated [MCU04]. Endosome-to-Golgi [NHB00]. Endosomes [LeB03z, LeB04-30, SDN^{+}00, TRC^{+}00, ATF^{+}04, BBMS03, BBP02, Dov03t, FCBH01, HCC02a, HGP^{+}04, KSBE03, LeB03-78, LO04, PMKM03, SATA^{+}02, SPO^{+}02, SKM^{+}02, SV03, BP00, NS00, NCUJ^{+}00, SG00, WJG^{+}00, MTG^{+}02]. endosomes/lysosomes [SKM^{+}02]. Endostatin [ACE^{+}01, iHGK^{+}02, KLGC^{+}01, Dov02-44]. Endothelial [IWG^{+}01, MTV^{+}00, PFW^{+}00, SEP^{+}01, WCBC04, BFC^{+}02, BYMS^{+}02, BFG^{+}04, CS04, CLG^{+}03, CSIK03, CPA^{+}03, EPH^{+}02, FFSF03, GMRYM^{+}02, GGF^{+}03, HEN^{+}01, ITM^{+}04, JOF^{+}02, LZC^{+}03, LCG^{+}04, MTB^{+}02, OMiKF02, TAA^{+}02, WNM^{+}03, WSR03, WSL^{+}01, XAB^{+}03, AML00, AV01, ECV^{+}00, HGK^{+}01, ZN01]. endothelial-mesenchymal [LCG^{+}04]. endothelium [PS03, SKN^{+}03]. ends [Dov02-50, HKBH03, MDF01, MTM^{+}02, NW04, OMM^{+}03, RKKP02, VMH^{+}02, AKH00, MKST00, SLT^{+}01]. enemies [LeB03-47]. Energy [OSNG04, CWG^{+}02, HDL02, Lai03a, LeB02-63, LGB^{+}02, SP03, SMW^{+}03, SN04]. Energy- [OSNG04]. energy-independent [HDL02]. engage [EBWC01, WW02]. engaged [SRB^{+}04]. Engineered [SWB03]. Engineering [TUV00]. engulfment [LeB02-37]. enhance [CSL^{+}03, KR01, MGL^{+}00]. Enhanced [BWN^{+}01, DSG04, DRBF03]. enhancement [DSM^{+}03, WXQ^{+}00]. Enhancer [ESS^{+}00, PTM^{+}01, SKF^{+}01]. enhances [ARLC^{+}04, BMY^{+}01, CYC^{+}04, CAGK^{+}03, DWB03, FAF^{+}04, KWSK^{+}04, RPS^{+}02, SRK^{+}03, SVT^{+}02, WNM^{+}03, YPN^{+}04, FNK^{+}00, MSV^{+}00, SRKR00, YHD^{+}00]. enhancing [BSMS03, LQPC^{+}00]. enlargement [KDH^{+}04, LeB02-107]. enough [Wel03-43, MWP^{+}00]. enriched [SSG^{+}02, HCC02a]. enriches [BSW^{+}04]. ensuring [OSB04]. Enter [Wel01w]. enteric [FLS^{+}04]. entering [IFP^{+}03]. Enth [HCD^{+}00]. Enthoprotin [WLGP^{+}02]. entire [MMFS01]. Entities [PKF^{+}00]. entrance [ARQ^{+}04]. entry [BLP^{+}02, CdLvM^{+}04, DRG^{+}03, GJS^{+}03b, LeB04-68, MH02, MP04, RZB^{+}03, SBG^{+}04, SSRX04, Sod02, TFAM^{+}04, CMMP00, LAF^{+}00, Zwe00]. envelope [GMY^{+}03, LRD^{+}03, LSMS^{+}01, MKM04, MEFC03, RXS^{+}03, SERB03, SSK^{+}03, MLC^{+}01, SMWG00, SMTC00]. envelopes [SCP02]. environment [Cam03, GTPG03, NBWB^{+}00]. environmental [TNMM03]. Envoplakin [DKMW00]. enzyme [CGF^{+}04, DJ03, MRT^{+}01, TBRG01, WSWM04b, ZWB04, DSSWW00, DOL^{+}01]. enzymes [PPWM04, SRW^{+}02]. EP [TOM00]. Epac [REK^{+}03]. Eph [WKJS^{+}04]. EphA [MNC^{+}03]. EphB [HIN^{+}03]. EphB1 [VCDHD03]. ephrin [WKJS^{+}04, WBC^{+}00]. Ephrin-A5 [WBC^{+}00]. Epi64 [RB01]. Epiblast [GNS^{+}04]. epicardial [SHPY02]. Epidermal [BRG^{+}00, STJ^{+}01, BD02, CNHK02, CBG^{+}01, FHF^{+}02, LCT^{+}04, LSW^{+}03, MWL01, PCB^{+}03, SMT^{+}03, CBZ^{+}00, CDWB01, KYF00, MPR^{+}01, THE^{+}00]. epidermis [BHL^{+}03]. Epidermolysis [CLWR01]. Epimorphin [HRB^{+}01, LLH^{+}01]. epithelia [RAS^{+}03, TAA04, HSB00, SYH^{+}01, WRGK00]. Epithelial [ESS^{+}00, MCB00a, BGM03, BWV^{+}01, BGW^{+}04, CBRBM04, DFJ^{+}02, EPH^{+}03, GLJP01, HP04, HEW^{+}01, JLK^{+}02, KSC^{+}04, LBH^{+}02, LRWB04, MWL01, MNC^{+}03, MTM^{+}02, MCF^{+}02, MMG^{+}04, NH03b, NCMO^{+}02, RB03, SC01a, SRG^{+}04, SMS^{+}01a, SSO^{+}03, SDMC^{+}04, TKS^{+}02, WDS^{+}03, WDL^{+}04, WSH^{+}04, YIW^{+}04, CWMO00, CLAC00, DWD^{+}00, FPSM01, HVT^{+}00, KIK^{+}00b, KGC^{+}00, LSA^{+}00, LM00a, LCS^{+}01, LBWH^{+}00, MKST00, RSG01, Ste00, YYM^{+}01, ZEtK^{+}00]. epithelialization [CAW^{+}04]. epithelium [ANC^{+}02, GLDM01, MSS^{+}01, FCL^{+}00, Krä00, TSMT00]. Epitope [PQF^{+}00]. Epitope-Tagged [PQF^{+}00]. EPLIN [MSA^{+}03]. Eps15 [CSP^{+}00, HCD^{+}00]. Eps8 [IFP^{+}03, STA^{+}01]. Epsin [HCD^{+}00]. EpsinR [MPV^{+}03]. Epstein [OBG^{+}03]. equal [GKYY03, PDMO03]. equals [Bir04a]. Equatorial [RS00c]. Equine [TGM^{+}01]. ErbB [SKO04, HBSQ01, SGPL^{+}00, OSWG02]. ErbB-2 [SKO04, HBSQ01]. ErbB2 [HEN^{+}01, SGPL^{+}00, GVT^{+}00, PMKV01]. ERBB4 [WAV^{+}04]. ERBB4/HER4 [WAV^{+}04]. erects [Dov03-57]. Erg [IHK^{+}00]. Ergosterol [UN03]. ERK [AFN00, ASAJ01, HFK^{+}03, MKK^{+}00b, RGGL00, SDS^{+}04, SPA^{+}04, SZvBuH^{+}04b, WGF^{+}00, YPN^{+}04, VCDHD03]. ERK1 [BFC^{+}02, BBDM02, DZT^{+}00]. ERK1/2 [BFC^{+}02, BBDM02, DZT^{+}00]. Erk2 [CAP00]. ERM [DMBS02, GWL03, LKH^{+}04]. Ero1 [TR02]. ERp57 [LC04]. Err [BMKA01]. errors [Wel03i]. Erv14p [GTBM04]. Erv41p [OBH^{+}01]. Erv46p [OBH^{+}01]. Erythroid [WWK^{+}00, BBDK^{+}04]. Erythrokeratoderma [SJA^{+}00]. escape [CSL^{+}03, LeB02-46, LeB02-68, VLL^{+}03]. Escherichia [ASP^{+}00, HSMB02, KM00, UHR^{+}03]. Escorter [RVB^{+}01]. Escorting [LeB04-31]. ESCRT [BBMS03, BWK^{+}03, KSBE03]. Esophageal [SML^{+}04]. Esp1 [JSCR01]. Espin [LZS^{+}03]. Essential [AMEC01, BHW^{+}02a, BPMG00, IUK04, KIO^{+}00, MCB00a, SBZ^{+}00, SMS^{+}01b, SBH00, WVBY^{+}03, WSL^{+}01, XHG^{+}00, BJB^{+}03, BLP^{+}02, Che02, CDE^{+}03, CPG^{+}03, ELNA^{+}03, FYI^{+}03, GLS^{+}03, GF03, GGB^{+}04, GTR^{+}03, GKYY03, HIE^{+}01, KBG^{+}03, KSD04, Kil03, LDK^{+}03, LMDS^{+}03, LVD^{+}04, MSL^{+}02, MCBB^{+}04, MLKH04, MKJ^{+}02, MCG^{+}03, MCA^{+}03, NM02, PKR^{+}02, PBD^{+}02, PNSJ01, REM^{+}02, RCY^{+}03, SK04, SSGLS01, SXD^{+}03, TOTC01, TGD^{+}03, TVF^{+}03, THG^{+}04, UJK^{+}02, WBP^{+}03, dMMBK^{+}02, vEPP^{+}01, DVE^{+}00, EU00, ECV^{+}00, HSKG00, HCOC00, IdCAS^{+}00, JC01, RBV00, RMMC01, WCGT^{+}00, YMM^{+}00]. established [LCI^{+}01, WSWSL04]. establishing [MTM^{+}02, SYH^{+}01]. Establishment [MCB00a, HP04, MB03c]. ester [IBP^{+}04]. Estrogen [ESS^{+}00, BMKA01]. Etf1 [WE02]. Ethylmaleimide [MSH^{+}00]. Eto [WND^{+}00]. Eto/Mtg8 [WND^{+}00]. euchromatin [PFB^{+}03]. eukaryotes [TW04]. eukaryotic [JR02, ØTW04, DLS00]. evades [LeB04-29]. Evanescent [TSK^{+}00]. evasion [Dov01y]. event [SHP01]. events [DLPB03, SMR^{+}02, ZCH^{+}02, GCN00, HLU00, IAG^{+}00]. Everyone [LeB02-38]. everywhere [Dov01z, LeB03d, Der01]. Evidence [BKD^{+}00, BSW^{+}00, CGY01a, FSD00, MN03, MMH^{+}00, MMFS01, OMiKF02, PPR^{+}00, PL01, OMWSN02, WWJ^{+}00]. evil [LeB03-74]. evoked [WWBGG03]. Evolution [DGB^{+}00, HES00, LeB03-27, HZ00]. evolutionarily [BRB^{+}01, SYH^{+}01]. Evolutionary [MM01, Mel04a]. EWI [SKH03]. EWI-2 [SKH03]. Ex [CMMP00]. Examined [MDJF00]. Exchange [AZB^{+}00, LeB02-39, AQC04, BPPFM^{+}03, DBLG02, OMWSN02, TCH^{+}02, WZB^{+}01b, YEG01, NM00, NA00, WSFN00, WBM^{+}00, WSE00]. exchanger [ARMB04, DB02, Sre04e, YpHRL03]. Exchanging [Wel01x]. excitement [Dov02-30, LeB02-57]. excluded [LW03, CW01]. Exclusion [OAR^{+}00]. exclusive [LeB03-90]. Exclusively [LGM^{+}01]. Execution [FWM^{+}01, TRW^{+}00]. executioners [Dov03-58]. Exhibit [TMK^{+}00]. exhibiting [RBD^{+}01]. exhibits [HYMS^{+}02, MKTW01, AKW00]. Existing [GOL^{+}01]. exit [CFB^{+}03, HS04, LeB03-77, LP04a, MPBR03, POH^{+}04, Sre04e, UHR^{+}03, WMA^{+}04a, WKZ^{+}02, Wel02w, YITe03, RHM00]. Exo70p [BHPN04]. exocrine [PRJK01]. exocyst [BHPN04, PDR^{+}03]. exocytic [BHPN04, HS02]. Exocytosis [HGC00, Tum03c, Wel03s, AMG^{+}01, AM03, EBWC01, GAT^{+}03, JAS02, THO^{+}04, TMA^{+}04, VCGB^{+}02, YGWN01, AEL^{+}00, BPS^{+}00, DVE^{+}00, GCN00, HWHH01, MF01a, MCH^{+}00, SGAS00, SBI^{+}00]. Exocytotic [SBI^{+}00]. Exogenous [HPQ^{+}00, WHS00, LMHJ02]. exon [GI02, MCG^{+}03, WCIN04, CDFT^{+}01, LMW^{+}00, MRM^{+}00]. exosome [Wel03-65]. Expaaaaanded [LeB04-32]. expand [LeB02k]. Expanded [MMG^{+}01]. expansion [HDP^{+}01, NMG04, ADKK00, GVT^{+}00, LMW^{+}00]. experience [Dov01x]. Experimental [SEP^{+}01]. explains [WLR01]. explanation [Les01i]. Exploits [CSJ00]. Export [BHL^{+}01, GDHS01, HRE^{+}01, LeB04-33, LHW^{+}01, SBH00, YB01, BM02, FGS^{+}02, GI02, GNDLS^{+}01, HWBD^{+}01, KBG^{+}03, MSM^{+}04a, NMHH03, OBG^{+}03, PBT^{+}02, RXS^{+}03, SK04, SCP02, VSO^{+}01, WMA^{+}04a, WPC^{+}01, Wel02-61, Wel03d, AFN00, AFB^{+}01, HKJ00, HBL^{+}01, PLH^{+}01]. Exported [SHWH00]. Exportin [BM02]. Exportin-5 [BM02]. exposure [XRP^{+}01, LBWH^{+}00]. express [GNS^{+}04, TFF03]. expressed [CPA^{+}03, PSW^{+}02, BMKA01]. expressing [ACBG04, BCA^{+}03, EAD^{+}02, HViV^{+}02, PRLR02, PPA^{+}03, SJA^{+}00, CPN^{+}01]. Expression [BHY^{+}00, CZBH00, CBK^{+}00, DFYL00, MO01, AR03, BMM^{+}02, CDM^{+}02, CDK04, CWA^{+}03, CSSBY^{+}03, GCH03, HSC01, HPE^{+}01, JW04, JLJD03, KLD^{+}03, LHR04, MZT^{+}03, MSI^{+}03, NGS^{+}01, PPP^{+}01, PFB^{+}03, RPS^{+}02, RKR^{+}03, SCPP02, SMR^{+}02, TMA^{+}04, WAV^{+}04, XAB^{+}03, ZZM^{+}03, BLPP01, BWN^{+}01, ECO^{+}00, FLLE^{+}01, FFKC00, HPQ^{+}00, HAM01, HW00b, LYMC00, MSV^{+}00, SCTM00, SSW^{+}01b, TM00, TMK^{+}00, WHS00]. extended [Wel01y]. Extending [Mil02]. Extends [DKJ00]. Extension [ZBB^{+}00, HBSJ04, MYO^{+}04, SMW^{+}03, TPW^{+}04, Wel02-59, WLO^{+}02, RS00a]. external [BHL^{+}03]. extra [SLR^{+}03]. extra- [SLR^{+}03]. Extracellular [CGF^{+}04, CS01, CK00, MSMK04, PSD^{+}04a, WSL^{+}00, ASGL^{+}01, BHW02b, BSL^{+}01, BBDM02, CFWH^{+}01, CSO^{+}04, CACL03, ESC^{+}01, KYS^{+}02, LM01, NMH^{+}04, RWCC01, SOH^{+}04, SMS^{+}01a, GPAS^{+}01, GKSR00, KIK^{+}00b, Qua00, RRL^{+}00]. Extracellular-Regulated [CK00]. Extract [SCLC00, MSC^{+}03, MPB^{+}04, SHCM03, BGFJ01]. Extracts [PMBC^{+}00, SEM^{+}00, SVLM02, SM03b, WDMH03, AKH00, BKDH01, KTY^{+}00, WSDW^{+}00]. extravasation [WCBC04]. extrinsic [PHM^{+}02]. extrusion [RLTC^{+}02]. eye [IBP^{+}04, KPL^{+}02]. eyes [Dov02y]. Ezrin [BRM^{+}00, BYMS^{+}02, FGR^{+}04, GLA00].

F [DT00, FNK^{+}00, vdLBK^{+}04, BIC^{+}03, CG03, CP01, DSH^{+}03, FWY01, GOL^{+}01, GOV^{+}02, GOV^{+}03, HZS^{+}01, HP03, KEGDQ01, KGT^{+}02, MWMK04, RKJL03, WSDW^{+}00, WKS^{+}00, WAPB^{+}00, WWGK02, WLO^{+}02, YSK^{+}04, vdLBK^{+}04]. F-Actin [DT00, WKS^{+}00, BIC^{+}03, CG03, DSH^{+}03, GOV^{+}02, GOV^{+}03, HZS^{+}01, KGT^{+}02, MWMK04, WLO^{+}02, YSK^{+}04]. F-actin-dependent [WWGK02]. F3 [FSGDN^{+}00]. F3/Contactin [FSGDN^{+}00]. F508 [YRC^{+}04]. Fab1p [BND^{+}02]. face [Sil02]. faces [LeB04-101]. Facilitate [HLB^{+}00, FHM^{+}03, WKZ^{+}02, SMSM00]. facilitates [CiKBG03, GKM^{+}01, HW00a]. facilitating [FGSW03]. Factor [CGY^{+}01b, DSSY00, DSV^{+}00, DFYL00, ESS^{+}00, IWG^{+}01, KY00, LDS^{+}00, MZ00, NLRD01, PMBC^{+}00, PTM^{+}01, RM01, RRM^{+}00, SKF^{+}01, ZEW^{+}01, ALP^{+}04, APM^{+}02, AZP^{+}02, BGM03, BMM^{+}02, BPPFM^{+}03, BvdWD^{+}04, BPKR^{+}02, BFSO^{+}04, COB01, CFM^{+}02, ERS^{+}04, EPH^{+}02, FMJG04, GCT02, GFM^{+}04b, HIE^{+}01, HW04, HTS02, HHOP02, HNK^{+}03, IIN^{+}01, IKA^{+}03, JKG^{+}02, JON^{+}03, KWS^{+}02, KMP02b, Les01j, MH02, MWL01, NCGD^{+}03, OMB^{+}01, OMWSN02, PRJK01, PC01, PBT^{+}02, RZB^{+}03, RBD^{+}01, SPB^{+}02, STKCW02, SK04, SHCM03, SRKN03, STJ^{+}01, TIS^{+}01, TEC^{+}03, UJK^{+}02, UHR^{+}03, VCGB^{+}02, WKZ^{+}02, WHP^{+}02, YEG01, BRG^{+}00, BLC00, BLPP01, CWMO00, CBZ^{+}00, CD00, CMMP00, CDWB01, DLS00, ECV^{+}00, Ern00, FRO01, FWM^{+}01, GOL^{+}01, GBM^{+}00, HOvD^{+}00, HCD^{+}00, IHK^{+}00, KLF^{+}00, LR00, LCS^{+}01, MBSB00, MPR^{+}01, MHE^{+}00, NM00, NA00, ONS^{+}00, PGS^{+}01]. Factor [PDJ00, TSMS01, THE^{+}00, GPLS02, KVC^{+}03, SATA^{+}02]. factor- [HIE^{+}01, RBD^{+}01, Ern00, GPLS02]. Factor-1 [ESS^{+}00, SKF^{+}01, TIS^{+}01, KVC^{+}03]. Factor-2 [ONS^{+}00]. factor-dependent [HNK^{+}03]. factor-deprived [ERS^{+}04]. Factor-like [BLC00]. Factor-Mediated [CGY^{+}01b, NLRD01]. Factor-Stimulated [DSSY00]. Factor/Lymphoid [SKF^{+}01]. Factor/Neuregulin [ZEW^{+}01]. factor/TrkA [SATA^{+}02]. Factories [LRW^{+}00]. factors [AQC04, BN02, BPD^{+}04, FNZ^{+}03, GCH03, HNK^{+}03, Tum03h, ABP^{+}00, AW00, DLS00, MBSB00, MS00, SCTM00]. factory [Hua02]. FADD [KSR^{+}04]. fads [Sub02]. Fail [Pow01d, TBJ^{+}01]. Fail-safe [Pow01d]. failure [Dov03s, Stu04, US04]. FAIM [SDS^{+}04]. FAK [HMH^{+}03, MSI^{+}03, Sch04, SPA^{+}04, YPN^{+}04, ZSY^{+}03]. falciparum [GTR^{+}03, KBGG04b]. fallout [RRM^{+}03]. falls [Wel01k]. Families [NY00, OLB^{+}00]. Family [HLB^{+}00, PFW^{+}00, BHNG01, CG03, CROfC04, DK04, DBB^{+}02a, DMBS02, GGB^{+}04, GSB^{+}01, HW04, KPKY^{+}03, PPGN^{+}02, SF01, TCK^{+}03, Wel04b, dMMBK^{+}02, vBDH03, GH00, KMH^{+}00, LCGR00, NLBK00, SGPL^{+}00, TR00, TWBV^{+}01, TWS^{+}00b, WDLK01, YOK^{+}03]. famine [LeB03-31]. Far1 [ARQ^{+}04]. Far1p [FNFL03]. Farnesylation [RGGL00]. FAS [AV01, KY00, MSO^{+}00, GLDM01, KSR^{+}04]. FAS-Mediated [KY00, GLDM01]. fasciculata [DE01]. Fascin [BCM04, LeB04g, AS00, AKT01]. fascin-ating [LeB04g]. Fascin-mediated [BCM04]. FASEB [Wel01-28, Wel02-51]. Fashion [KO00]. Fast [Sre04f, EMY^{+}04, JLJD03, LeB03k, Tum04y, Wel04-69, HVM^{+}00]. faster [Bir04a, Dov02t]. Fat [LeB04-34]. Fat1 [TT04]. fatal [Les01b, XJW^{+}04]. Fate [BYLA^{+}01, LRBH02, LeB02-99, PDL^{+}03, RPNM03]. fates [ZGN^{+}04]. Fatp [HvdHG^{+}03]. fatter [LeB03-28]. fatty [HvdHG^{+}03, MWC^{+}02, TWS^{+}00b, vRTvdB^{+}00]. favorites [Dov03-38]. Fc [GBY^{+}03, LUB^{+}02, MBS^{+}01, MZ00, WPO00, WPS^{+}01]. Fe65 [SIBG01]. features [HvdHG^{+}03]. feed [LeB04-67]. Feedback [NZHR01, Wel03b]. feeds [Wel03-47]. feels [Dov01-31, LeB04p]. Fences [Dov02u, Wel03-40]. fenestration [LRD^{+}03]. Fer [APLB00]. fertilization [HTPC04, MGLPM00]. fetal [GBN^{+}01]. Feudal [LeB04-35]. Fewer [LeB03-28]. Ffa [CGY01a]. Ffa-1 [CGY01a]. FG [ZW04, SBH00]. FG-repeat [ZW04]. Fgf [MBSB00, AFR01, AR03, DLPB03, LRS^{+}02, LM01, MTB^{+}02, NDS^{+}02, WWD03, ZVPK03]. FGF-2 [MTB^{+}02, ZVPK03]. FGF-4 [AFR01]. FGFR1 [WFC^{+}02]. Fgfr2 [MBSB00]. FHL2 [MSJ^{+}02]. Fiber [WSL^{+}00, UTH^{+}02, DCM00, YHF^{+}01]. Fibers [KKA^{+}01, Dov03n, JLJD03, KCG^{+}03, LeB03-66, LeB04-37, LCRS01, MRK04, MKR01, PR02, SH02, TGD^{+}03, VMS^{+}02, BHFL01, BL01, HHF^{+}00, LMW^{+}00, MÖS^{+}00, PO00, TYY^{+}00]. fibril [SRC^{+}01]. Fibrillarin [SWJ^{+}00]. Fibrillin [BKZ^{+}01, GKSR00]. fibrillogenesis [DSB^{+}02, ECO^{+}00]. fibrillogenic [BTH^{+}03]. fibrinogen [Wel03-31]. Fibrinolysis [AKDS00]. fibripositors [CLM^{+}04]. Fibroblast [CMMP00, HTS02, KY00, LCS^{+}01, RM01, COB01, ELNA^{+}03, HKO03, MSI^{+}03, MWMK04, PBT^{+}02, SSM^{+}04, FRO01, MBSB00]. Fibroblasts [KY00, CLB^{+}03, FLWMG02, GPZ^{+}02, KI04, LRB^{+}03, SH02, TWS^{+}04, TIO^{+}02, BDK^{+}01, HCTM00, MCH^{+}00, SMI^{+}00, SPC00, TYY^{+}00]. Fibromodulin [ECO^{+}00]. Fibronectin [DSSY00, LeB03-29, RSBE00, DSB^{+}02, Dov03-46, HKO03, MCG^{+}03, SRC^{+}01, SDEZ^{+}03, AKFB00, KHH^{+}01, SMSM00]. fibronectin-binding [DSB^{+}02]. Fibronectin-like [SMSM00]. fibronectin-stimulated [SDEZ^{+}03]. fibrosis [FS03, LeB04-33, WMA^{+}04a, BHKL01]. Fidelity [KR01]. fight [LeB03-31]. Fighting [LeB04-36]. Figure [Ros02]. filaggrin [LSW^{+}03]. Filament [GRBD01, HMN^{+}00, TCV^{+}00, BMY^{+}01, HMVG02, HPG^{+}02, KKTP03, LeB02-62, LP04b, MKS^{+}02, OO02, RFLT02, sGS01, HHSV00, RBV00, STE^{+}01]. Filamentous [GKM^{+}01]. Filaments [AP00, BMY^{+}01, FCF^{+}01, GOV^{+}03, GCR^{+}03, ISB^{+}04, KK02, LSS^{+}02, MDF01, MSA^{+}03, MPR^{+}03, WPC^{+}02, WC03, YCB04, CG00, GV00, Ono01, TTS00, VDMH01, WPJ^{+}00, YYM^{+}01]. Filamin [FCF^{+}01, TCH^{+}00, DOB^{+}01, vdFKK^{+}02]. filled [LeB02-41]. Filopodia [SKF02, Dov03-57, GGF^{+}03, SBC^{+}03, VYW^{+}03, WMJ^{+}04]. filopodia-like [VYW^{+}03]. Fimbrin [VDMH01]. final [GJS^{+}03b, GCN00]. finally [Dov03y]. find [Dov01j, Dov02-66, Dov03-37]. Finding [Dov01l, Dov02w, Dov02x, Dov02v, LeB03-30]. finds [Tum02b]. finger [AKH^{+}04, BKD^{+}04, MKS^{+}02, Wel03v, Wel04-35, HCD^{+}00, NCUJ^{+}00, SEI^{+}00]. finger-1 [MKS^{+}02]. Fingers [LeB04-37]. Fip200 [UAZG00]. Fire [BGFJ01]. firefighter [Dov03-39]. First [RB01, BHW^{+}03, Wel03o]. Fis1p [MMS00, TONN02]. fish [Dov02-31, MPB02]. Fishing [SS01]. Fission [LBP00, MMS00, BN02, BPC03, BSMS03, KLG^{+}02, KKTP03, LeB04-27, LP04b, MRC^{+}02, MZR^{+}04, MMBM04, TMG03, TONN02, TBW^{+}04, BHM^{+}00, FSY00, FSKS00, TMD^{+}01]. fits [Wel02-43]. FKHR [DBL^{+}02]. FKHR-L1 [DBL^{+}02]. flagella [MCBB^{+}04, MN04, IBKSP01, PDV^{+}00]. Flagellar [GDRS01, Tum04i, Dov03k, TDL03, YDRS01]. Flat [Wel04-27]. Flat-out [Wel04-27]. fleet [Dov01a]. flexible [LeB02-86]. Flicking [Wel02x]. Flies [LeB03-31, Dov03s]. Flight [Dov01m, HES00, KNK^{+}01, MDF01, RBV00]. Flightin [RBV00]. FLIP [KSR^{+}04]. Flipping [Wel04-28, Wel02y]. FLN2 [TCH^{+}00]. Flow [Wel04-29, DSM^{+}03, NYT^{+}03, SM03a, SKN^{+}03, Wel03u, YLQ^{+}02, Yum01, AML00, SD00]. Fluorescence [SP03, LMVW03, SN04, MPR^{+}01, SGAS00]. fluorescent [CWG^{+}02, SAWS02, SV03, VZTN03, Zam04]. flux [LeB04-35, MPB^{+}04, SSP^{+}03]. fly [Wel02-31]. Flying [GG00]. Focal [ALWR01, BPS^{+}00, RZB^{+}01, ZWSC02, BFG^{+}04, EPH^{+}02, GYS02, GHS^{+}03, GGNK04, KAK^{+}03, LPL^{+}04, MPAP^{+}03, OPP^{+}03, SPB^{+}02, TWS^{+}04, AML00, BDK^{+}01, BKD^{+}00, GTPMU00, KHH^{+}01, NT00, TYY^{+}00, THZ^{+}01]. Foci [Wel04-30, CBS04]. focus [KMP^{+}02a]. Fodrin [TYS^{+}00]. Fold [LeB03-32]. foldable [YRC^{+}04]. folding [DK04, LeB04u, NSK04, Pow01i, SWB03, TW04, VKB^{+}01, Wel04m, BLC00]. follicle [KPA^{+}03]. Follicular [TSMT00]. follow [LeB02g]. followed [LRD^{+}03]. food [KBGG04b]. foot [Dov02-56]. Force [CTE^{+}04, SS02, AMBW04, BSEB04, GYS02, GJS^{+}03a, HGC02, JdDD03, KB04, LeB04-38, PRS^{+}04, SF01, vWJK^{+}03]. force-dependent [GJS^{+}03a]. force-driven [KB04]. Forced [LeB04-39]. forces [BHL^{+}03, LMG04, SB03, BDK^{+}01, KGE^{+}00]. Forcing [LeB03-33]. foreigners [LeB04-36]. forget [Dov03-27, LeB03y]. fork [FP02]. forkhead [HNK^{+}03]. forks [DC03]. form [DBH^{+}01, GBMA04, HBV^{+}01, HPE^{+}01, LMHJ02, LeB02-71, PMU^{+}02, RKF^{+}04, ZCH^{+}01, BKD^{+}00, LM00b, SJA^{+}00, TTS00, UIK^{+}01, WSW^{+}00]. Formation [DLXP00, DB00, MHS^{+}00, SDS00, SHF^{+}03, VYW^{+}03, BBMS03, BSW^{+}04, BHK^{+}02, CG03, CEGZ^{+}04, CSO^{+}04, CMS^{+}02, DDL^{+}04, DRP^{+}03, EGC^{+}03, FHM^{+}03, FCBH01, GWL03, GRCF02, GGH^{+}04, HKP^{+}04, HYMS^{+}02, IBS^{+}02, JBA^{+}01, KYZ^{+}04, KRS^{+}02, KDH^{+}04, KMLS04, LLGB03, LeB02-41, LKH^{+}04, LMDS^{+}03, MSGS02, MRK04, MGP^{+}02a, MKR01, MPAP^{+}03, NM02, ØTW04, RRM^{+}03, RWSV03, SHVR02, SAG^{+}01, SDD04, SYVB03, TBTN01, TFM04, VT04, VBR^{+}01a, WXD^{+}03, WBT^{+}03, WBAS04, WWGK02, YLG^{+}02, Yum01, ZWAH03, ZHH^{+}02, AKT01, BPS^{+}00, BMS^{+}00b, BKD^{+}00, BMKA01, BSW^{+}00, CLT^{+}01, DR00b, HVB^{+}00, HW00b, IHN^{+}01, JCR^{+}01, LWDH01, MSV^{+}00, MTV^{+}00, MYH^{+}01, NKH^{+}00, ONM00, SM00, TSDS00, VMO01, WBG01, WWJ^{+}00, YHF^{+}01]. Formed [SD00]. formerly [Pow01e]. Formin [DPB03, Wel01z, Wel04-32, Wel04-31, FVC04, KKTP03, KB04, MMM^{+}04]. Formin-dependent [DPB03]. Forming [ATE^{+}01]. formins [Dov03-31]. forms [GG04, HRM02, JGR^{+}04, NMG04, ZCW^{+}03, MRM^{+}00, RPE00, SMS^{+}01b, TWBV^{+}01]. Formyl [GHS00]. Fortune [Bre00]. forward [Dov03h, Wel04e]. Fos [ESS^{+}00]. found [CSM03, Dov04k, LeB02-105, LdVV^{+}02]. Four [Dov03r, IBKSP01, PP00, SME^{+}00]. FOXO [GCH03, HNK^{+}03]. FoxO3a [BBDK^{+}04]. fractures [LTF^{+}01]. fragility [CLG^{+}03, CBG^{+}01]. fragment [BTH^{+}03, CNMS02, NMHH03, SHB^{+}03, TMA^{+}04, ZCW^{+}03]. Fragmentation [LDS^{+}00, CNMS02, CSM03, ECK^{+}03, LLP^{+}02, CDEM00]. fragments [WSWM04b, DSSWW00]. Frame [LCGR00]. frames [LHR04]. FRAP [RPNM03, WCIN04]. Free [SM00, DH02, Dov02u, LeB03w, Wel02a, AEL^{+}00, PWC^{+}01]. free-ranging [Dov02u]. Freed [Les01c]. Freedom [Wel04-33]. frequent [BCM04]. FRET [SP03, LMVW03, LeB03-34, OGD03, YOK^{+}03]. FRET-based [YOK^{+}03]. FRET-ing [LeB03-34]. FRG [FSK^{+}04]. Frizzled [THK^{+}00]. Frizzled-2 [THK^{+}00]. frog [Wel03-53]. front [BHIWH01, Sre04g, Wel01-44, Wel04-29]. Fructose [BMC00]. Fructose-1 [BMC00]. FSHD [TSL04]. FtsY [MGMH03]. Ftsz [VMO01, LeB02-69, Eri00, KKR^{+}00]. fucosylation [SMR^{+}02]. fuels [LeB03-35]. Function [CENMR^{+}01, CBK^{+}00, GMRS00, HES00, MCB00a, YMR03, vARP^{+}00, AOH^{+}02, ABRA03, BDKM04, CGM^{+}02, CNH^{+}02, CBW^{+}01, CRH02, DIR03, DRBF03, Dov04d, EWM^{+}04, FCF^{+}01, FBH03, FS03, GF03, HPG^{+}02, HBD^{+}02, IDvH^{+}02, JKW^{+}03, JJM^{+}02, KKPB03, Kil03, KWSK^{+}04, LRF^{+}02b, MKD^{+}01, MCBB^{+}04, MDW^{+}04, NB03a, NOM^{+}04, OL02, PLR03, PRLR02, PCB^{+}03, PNSJ01, PSK^{+}03, RGG03, RL03, SRG^{+}04, SLP03, SMS^{+}01a, SGC^{+}02, SMC^{+}02, TDFV02, TTM^{+}03, TSH^{+}04, Tum02b, UOB^{+}02, WPC^{+}01, Wel02e, WW02, WC03, XAB^{+}03, YLG^{+}02, YKT^{+}04, CGY01a, EMW^{+}01, FPSM01, HHSV00, HMAM01, JBN^{+}00, KNIO01, LLH^{+}00, MMWC00, MLS^{+}01, ONM00, PDW^{+}00, Rut00, SL01, Ste00, TNB01, VDMH01, WK01]. Functional [DMBS02, DWFAdL^{+}02, HGS^{+}01, sKCK^{+}01, MGMH03, NZA^{+}01, NWT^{+}01, OSMF00, ODR^{+}01, SRL^{+}04, STP^{+}00, TCK^{+}03, WDFNN04, YD00, BCA^{+}03, BBBS04, CSJ^{+}03b, GC04, GBMA04, HBH^{+}04, LDP02, LVD^{+}04, MBM^{+}04, RKJL03, RSD^{+}04, TRMI^{+}04, FLLE^{+}01, IHK^{+}00, LCM00, LMW^{+}00, MR02, OSN^{+}00, PCR^{+}01]. functionally [CVZ^{+}04, FPM^{+}03, IKS^{+}02, WPM^{+}00]. functioning [MDQ^{+}03, WAV^{+}04]. Functions [LBP00, AMG^{+}01, DMA^{+}01, Dov03-46, FMJG04, GCO^{+}04, GFM^{+}04b, Hem01, HEW^{+}01, IOIF^{+}04, Kel03, KPA^{+}03, LHC^{+}02, LYL^{+}04, MSM04b, MKS^{+}02, RBW^{+}02, RML^{+}02, SLB^{+}01, SHPY02, SKH03, TONN02, WVBY^{+}03, WBT^{+}03, WE02, DWD^{+}00, FSD00, KKS^{+}01, KM00, LS00, WTG01, YF00]. funeral [EC03]. furrow [Glo04, ISS^{+}04, RRM^{+}03, Sre04k, Wel03-33]. furrows [CWG^{+}02, CW04]. Further [Pro03]. Fus3 [KMS^{+}04]. Fus3p [MMM^{+}04]. fuse [Dov04b, MWN^{+}04a, Dov01-28]. Fusion [Dov03s, Dov03u, HDP^{+}01, LeB03-35, LeB03-36, PSWU00, Sre04g, TCR00, TSK^{+}00, BRB^{+}03, BMLU02, BBSF01, CHA^{+}01, CDE^{+}03, Dov03t, Dov03-36, EWTW02, FJM^{+}04, KAC^{+}04, LeB04-27, LeB04-59, LeB04-102, LdVV^{+}02, LMDS^{+}03, MSGS02, MLZ^{+}01, MWM^{+}02, MSN^{+}02, PBP^{+}01, PG02, SVI^{+}04, SYVB03, TTM^{+}03, Tum03f, UJK^{+}02, WMCW03, WSK^{+}03, Wel04-28, WWS^{+}03, XH04, AKW00, CB00, CAB00, Coo01, DM00, DR00b, FLLE^{+}01, FRK^{+}01, Gau00, GBON00, HW00a, MMH^{+}00, MSH^{+}00, PGS^{+}01, PWU00, PMB^{+}00, SJ01a, TR00, WWG^{+}00, WSE00]. FX [SMR^{+}02]. Fxfg [SBH00]. Fyb [KSK^{+}00]. Fyn [CGM^{+}02, KSK^{+}00, MWF02, MKD^{+}01, MTW^{+}04]. Fyn-Binding [KSK^{+}00]. Fyve [NCUJ^{+}00]. Fzo1 [FRK^{+}01]. Fzr [RJyH02]. Fzr/Cdh1 [RJyH02]. Fzy [RJyH02]. Fzy/Cdc20 [RJyH02].

G [JB01, LeB03-37, SGK02a, BHW^{+}02a, BFH^{+}01, BSR^{+}03, BC00, ELO^{+}01, IOIF^{+}04, ISB^{+}04, KVM03, KS02, MTV^{+}00, MCB00a, NB03a, SDL^{+}03, YSW02, YCK^{+}03, ZJM^{+}02, JLJD03, YCK^{+}03]. G-domain [BHW^{+}02a]. G-proteins [LeB03-37]. G1 [BPKR^{+}02, FNFL03, LCI^{+}01, LCSG03, LRB^{+}03, WSC^{+}03]. G1-CDK [FNFL03]. G1-S [BPKR^{+}02]. G2 [CS01, KCWF02, MSR04]. G2/M [CS01, MSR04]. G3BP [TCZ^{+}03]. Gab1 [CNHK02, SBZ^{+}00, SGF^{+}00]. Gab2 [GBY^{+}03]. GABA [NOM^{+}04]. GABA-containing [NOM^{+}04]. GAD65 [KEHAM^{+}02]. GADD34 [NZHR01, SSH^{+}04]. GADD34-Mediated [NZHR01]. Gag [RKKP02, LMVW03, PHS^{+}03]. Gag3p [FSY00]. gains [Wel03-34]. galactolipid [JMB^{+}04]. galactolipids [MDP02]. Galactosylceramide [SKT^{+}00]. Galectin [HVT^{+}00]. game [Les02d]. ganglion [SLB02]. Gap [LGP00, CMS^{+}02, Dov02z, LM01, LWCKL01, RDC^{+}04, SAS^{+}02, TLS^{+}01, XLH^{+}01, GAT^{+}03, LeB02-41, RAD^{+}02, SDL^{+}03, KGvdG^{+}00, LTB^{+}00, QRLL00, RGG00]. Gap43 [FLX^{+}00, LFT^{+}00]. Garbage [Pow01f]. gastric [CPN^{+}01]. gastrointestinal [BPKR^{+}02]. gastrulation [DMC^{+}03, IUK04]. GATA [WHM^{+}02]. GATE [MSN^{+}02]. GATE-16 [MSN^{+}02]. gated [JB01, KKTP03]. gatekeeper [Dov02-35]. Gazing [Dov02z]. Gbp [FFY^{+}00]. Gcs1 [PNSJ01]. GDNF [PUK02, PPA^{+}03, YJS^{+}03]. GDNF-deprived [YJS^{+}03]. GDP [HWBD^{+}01]. GEF [FSK^{+}04, LeB03q, LeB03-75, STA^{+}01]. GEFs [GCO^{+}04]. gels [SH02]. Gelsolin [CKS^{+}00, WXD^{+}03, WLKS01]. gelsolin-null [WLKS01]. Gem [PFW^{+}00, WYR^{+}02]. Gem1p [FMC^{+}04].

Gemin4 [CPP^{+}00]. Geminin [LeB04-40, MBM^{+}04]. Gene [BYLA^{+}01, MPB02, RGM^{+}02, RKR^{+}03, WND^{+}00, BWV^{+}01, BXR^{+}02, CLG^{+}03, CKW^{+}03, EAD^{+}02, GCH03, HvdHG^{+}03, JLJD03, LHR04, MWMK04, MPG^{+}03, NH03b, NDM^{+}03, PMKV01, PFB^{+}03, PPK^{+}01, RJA^{+}03, RP03b, SGK^{+}02b, SJB^{+}03, TBJ^{+}01, Wel04-59, WCTU02, WAV^{+}04, ZZM^{+}03, AHMJ01, BLPP01, Cha00, DEG00, FSBH00, HAM01, HDL^{+}00, LYMC00, Ono01, PDV^{+}00, PZP^{+}01, RMG^{+}00, SME^{+}00, SRHV00, SPH^{+}00, TM00, TWS^{+}00b]. gene-marked [PPK^{+}01]. gene-rich [SJB^{+}03]. general [JRL^{+}03, AW00, HLK01, KRR^{+}01]. generalist [Wel04-37]. generate [BCA^{+}03, XH04, BHM^{+}00]. generated [WDS^{+}03]. generates [CSJ^{+}03b, KRF^{+}03]. generating [Néd02]. Generation [FRT^{+}01, WGvA^{+}01, ARLC^{+}04, ACBG04, HGC02, KH04, MGP^{+}02b, RGG03, BDK^{+}01]. Genes [AW00, BvdWD^{+}04, FM01, KMP^{+}02a, LLGB03, LeB04-44, MPG^{+}02, SJB^{+}03, SMC^{+}02, ZAE^{+}04, CA00, TTHH00, WCGT^{+}00]. Genetic [ZJM^{+}02, SKGC^{+}03, RS00c]. genetically [VZTN03]. genome [OSB04, FSBH00, GG00, Las00, MMC00, MS00]. Genomes [MSFH00]. genomic [WSS^{+}04, Lit00]. Genomics [Mel00a]. geometry [LeB03-37, TKHR03]. geometry-dependent [TKHR03]. Geranylgeranylated [GBON00]. Germ [KPB^{+}00]. germinal [HMG03]. Germinating [TOM00]. germline [Wel04-53, KPB^{+}00]. Ges [PFW^{+}00]. Get [Wel04-34, Dov02-46, Dov03-45, Dov03-55, LeB02-56, LeB02-57, LeB03-64, Wel01u, Wel03i, Wel04x]. getaway [LeB03k]. gets [Dov01d, Dov01c, Dov01n, Dov01o, Dov01-33, Dov02f, Dov02-58, Dov02-59, Dov02-67, Dov03c, Dov03-29, LeB03-36, LeB03-88, Mel01, Sre04a, Wel01-61, Wel02-28, Wel03-42, Wel04p, Wel04z]. Getting [Dov02-27, Dov03v, Dov03w, Dov03x, LeB02-42, LeB02-43, LeB04-41, Vin04, Wel02z, LeB04-89]. GFP [NSLSK02]. GFR [PPA^{+}03, PUK02]. Gga [BP00]. Ggas [DPM^{+}00, GGGK03]. Ghrelin [BFC^{+}02]. Giant [HES00, YEG01]. giantin [PL01]. Gic1 [HS04]. Gic2 [HS04]. Gin4 [WBAS04]. Gip1p [TBTN01]. GIT1 [ZWAH03]. Give [SRHV00]. Giving [Wel04-35]. gla [MSMK04, NGS^{+}01]. gla-containing [MSMK04]. gland [HSC01, JZ02, SHB^{+}03, WSL^{+}03, ASMW01, IEJ^{+}01, MLS^{+}01]. glands [WVBY^{+}03]. Glc7p [TBTN01]. Glfg [SBH00]. glia [SWBE^{+}04]. Glial [TK00, ZEW^{+}01, MDP02, TGMC^{+}00]. gliding [LP04b]. glioma [Tum04t]. Gliotactin [STA03, GF03]. global [SH02, SMR^{+}02, RS00c]. globally [Dov04e]. globin [LeB02-93, LCI^{+}01]. Globoid [IHN^{+}01]. glomerular [KVM03]. glucocorticoid [MKS^{+}02, MMG^{+}04]. Glucocorticoids [GJB^{+}00]. glucokinase [RP03a]. glucose [EGC^{+}03]. Glucosidase [GPL^{+}02]. glue [Dov02-28, LeB03-38, Tum04w, Wel04-41]. Gluing [Tum02c]. GLUT4 [WSC^{+}01]. Glutamate [AGB^{+}00, LL03b, TNB01]. glutamine [CRE02]. Glycanase [SPH^{+}00]. glycans [XJW^{+}04]. Glycine [KRA^{+}01]. Glycogen [HCK^{+}00]. Glycophosphatidyl [SDDS00]. Glycoprotein [WFQ^{+}00, BTH^{+}03, MDP02, YHT02, CS00, PQF^{+}00, TH00].

Glycosphingolipids [SDC^{+}01]. Glycosylation [Dov03y]. Glycosylphosphatidyl [FSGDN^{+}00]. glypican [DMC^{+}03]. glypican-3 [DMC^{+}03]. GM130 [PSD^{+}04b, PL01, LGW00]. Gnawing [LeB02-44]. go [Dov02-36, Dov03-62, LeB02-60, LeB03-89, LeB03-101, Pow04, Sod02, Wel03m, Wel03-62]. goes [Dov02h, Dov02-29, Dov03-41, Dov04d, Wel03f, Wel04h]. Going [Les01d, MKT01]. Golgi [ATF^{+}04, BBP02, BPKK01, Bar04, BHKL01, BP00, CEGZ^{+}04, CS01, CNMS02, CDEM00, CSM03, DRP^{+}03, DSSWW00, DOL^{+}01, EWSN00, FR01, sGS01, GTBM04, HLB^{+}02, HCC02b, HHM^{+}04, HGP^{+}00, HMRH01, HLB^{+}00, JCPWS01, KTY^{+}00, KNI^{+}04, LLP^{+}02, LOS^{+}01, LeB03-39, LeB04j, LeB04-52, LeB04-102, LRF^{+}02b, LGW00, MMR^{+}00, MMPO^{+}01, MMFS01, MBN^{+}01, MSN^{+}02, Nel00, NKP^{+}01, NHB00, OAR^{+}00, Pel01, PG02, Pfe01, PPM^{+}00, PNSJ01, PSD^{+}04b, PWU00, PWS^{+}01, PL01, RAD^{+}02, RRB^{+}01, SSN01, SSA^{+}00, Sea04, SPK^{+}01, SFV^{+}00, SV03, SDC^{+}01, SDL02, UJK^{+}02, UJL^{+}03, UOB^{+}02, WDLK01, WSWM04b, WPC^{+}01, Wel01-31, Wel02-27, WJG^{+}00, XTN^{+}02, XH04, YGWN01, yZCKA01]. Golgi-localized [UOB^{+}02]. Golgi/TGN [PPGN^{+}02]. golgin [DRP^{+}03, GTBM04, MMR^{+}00]. Golgin-160 [MMR^{+}00]. golgin-84 [DRP^{+}03]. Golgins [SBS^{+}02b]. gonadal [COB01]. gondii [ARMB04, DSN^{+}01, GGD^{+}04, HRM02, JR02, SJ01b, SCS^{+}00, CSJ00, RVB^{+}01]. gone [MF01b, Wel01-39]. gonorrhoeae [LBH^{+}02]. good [Les02a, Wel01t]. Gordian [PSB00]. GOS [MSN^{+}02]. GOS-28 [MSN^{+}02]. Got [KF00]. govern [SHVR02, YWW^{+}04]. Governs [FAAS00]. gp210 [DW02]. Gp41 [MMH^{+}00]. Gp60 [MTV^{+}00]. gp91 [ESC^{+}01]. GPI [MPBR03, PSP^{+}04, SWE^{+}03]. GPI-anchored [MPBR03, PSP^{+}04, SWE^{+}03]. GPR [LeB03-40]. GPR-1 [LeB03-40]. GPR-1/ [LeB03-40]. Gradient [BEG01, MC02]. gradients [HTRK02, vEWS^{+}01]. granular [GSP^{+}02a]. granule [ABF^{+}03, FHM^{+}03, SBS02a, SLB^{+}01, TMA^{+}04, CAB00, HWHH01, JLS^{+}01]. Granules [HGC00, DSH^{+}03, SS04, THO^{+}04, TCZ^{+}03, TOM01, KCL^{+}00]. Granzyme [MWE^{+}03, PBB^{+}04, TST^{+}03]. GRASP55 [SPK^{+}01]. GRASP55-rab2 [SPK^{+}01]. GRASP65 [LLP^{+}02, Wel02-28]. graveyard [Wel03t]. Grb2 [SGF^{+}00]. Greater [Les02b]. Greatwall [YFW^{+}04]. grip [Dov01n, LeB02-42]. groove [LeB03-36]. ground [KC02, SL00]. Group [DBS^{+}01]. Grow [LeB03-41, Les02c, Dov03-52, LeB03-93, Wel03u]. Grow-your-own [Les02c]. Growing [Les01e, MKST00]. Growth [DSSY00, GJ00, IWG^{+}01, KY00, RM01, ZEW^{+}01, qZC01, AMG^{+}01, BMM^{+}02, BS04b, BBDM02, Bre03, CNHK02, CSL^{+}03, COB01, DLPB03, DLY^{+}02, Dov01n, Dov02-52, EWD02, EPH^{+}02, ECK^{+}03, FNZ^{+}03, GLS^{+}03, GWL03, GFM^{+}04b, HIE^{+}01, HTS02, HP03, HNK^{+}03, IKA^{+}03, JZ02, KSK01, LRS^{+}02, MWL01, MEK^{+}04, MTW^{+}04, NK02, NGK^{+}03, PBT^{+}02, RCY^{+}03, RJA^{+}03, RP03b, SMT^{+}03, SMS^{+}04a, SKF02, SRC^{+}01, SATA^{+}02, SDML04, SEW^{+}01, SF01, STJ^{+}01, TIS^{+}01, UGKT^{+}02, WK02, WNM^{+}03, WKZ^{+}02, WHP^{+}02, WWD03, XRP^{+}01, ZWSC02, BRG^{+}00, BHM^{+}00, BLPP01, CBZ^{+}00, CLAC00, CMMP00, CDWB01, DCM00, ECV^{+}00, FRO01, FSK^{+}00, FNK^{+}00, FWY01, GBM^{+}00, HFM^{+}01, KSN^{+}01, LR00, LCS^{+}01, MBSB00, MPR^{+}01, PGS^{+}01, SHS^{+}00, THE^{+}00, WBC^{+}00, WTG01]. Gsa9 [KKS^{+}01]. GSK [EWD02, TJC^{+}03, FFY^{+}00]. GSK-3 [EWD02, TJC^{+}03, FFY^{+}00]. Gsp1p [DMA^{+}01]. GTP [DMA^{+}01, AQHO03, DSB^{+}02, GBZ^{+}02, SDS00, SHKS02, VT04, WYR^{+}02]. GTP-bound [AQHO03]. GTPase [AFK^{+}03, Dov02-29, FMC^{+}04, GAT^{+}03, GTBM04, GGNK04, HN03, KMK^{+}02, LeB04-42, MSJ^{+}04, MMS00, PFW^{+}00, RMW03, SHKS02, WSWSL04, WWS^{+}03, AFB^{+}01, CB00, FSD00, FSKS00, FWY01, JBN^{+}00, KEGDQ01, RRK^{+}00, RDH^{+}01, TN00a, WWG^{+}00, WSE00]. GTPase-Mediated [MMS00]. GTPases [CMC^{+}02, DBB^{+}02a, WW02, YOK^{+}03, AS00, MSM^{+}01, NLBK00, Rid00]. guanine [AQC04, BPPFM^{+}03, OMWSN02, YEG01, WSFN00]. Guanosine [WYHP00]. guards [Tum04f]. Guidance [ACE^{+}01, MTW^{+}04, SWH^{+}02, Sre04i, SSW^{+}01a, FSK^{+}00, MHW^{+}00, Rut00]. guide [SKF02]. guides [GGF^{+}03]. gut [FLS^{+}04]. gypsy [BC03].

H [AZB^{+}00, LeB03-42, Nel00, DB02, HMN^{+}00, JBK04, NGKH02]. H-Warts [HMN^{+}00]. H-Warts/Lats1 [HMN^{+}00]. H/ACA [JBK04]. H1 [DBLG02]. H2a [CW01, MJG^{+}01]. H2AX [FCLSN03]. H2b [GZY^{+}00, MJG^{+}01]. H3 [AMEC01, GG01]. HA95 [MEFC03]. hair [KPA^{+}03, WLR01, CLAC00]. hairy [LeB03f]. Half [SME^{+}00, WHS00]. halts [LeB04-40]. hampers [SRD^{+}02]. hamster [LCI^{+}01]. hand [OM01]. hand-holds [OM01]. handle [Dov03m, Dov03v]. handling [PDMO03]. Haptotactic [HBSQ01, KHH^{+}01]. harboring [TSL04]. hard [Tum04y]. Harvesting [RBB00]. hastens [LeB04r]. HB [YIS^{+}03]. Hcap [SCLC00]. Hcf106 [CM01]. Hcp [MR01b]. Hcp-4 [MR01b]. Head [GTM^{+}01, KRF^{+}03, NH03a]. Head-to-tail [GTM^{+}01]. heads [Dov03-62, MF01b]. healing [KHLW02, LeB03-43, MDT^{+}01, MCG^{+}03, Wel01-67, Wel02u, Wel02-52, LQPC^{+}00, THE^{+}00]. healthy [Dov02y]. Heart [LeB04-43, FSCF^{+}03, GGH^{+}04, LCG^{+}04, Sod02, Wel01-45]. Heart-shaped [LeB04-43]. Hearts [Wel03u, LeB04-47]. heat [Dov02v, JKG^{+}02, LeB04p, Wel03-60, BMC00, HCD^{+}00]. heat-shock [Dov02v]. heavily [RGM^{+}02]. Heavy [AP00, ATG^{+}03, JLJD03, RGM^{+}02, RL03, RFLT02, XRH^{+}03]. Hec [LeB02f]. hedgehog [FLS^{+}04]. hedgehogs [Wel02l]. HeLa [DMH^{+}02, MMFS01, SCY01, WSC^{+}03]. Held [Wel03v, LeB02z]. Helical [KKL^{+}01]. helicase [FNKH02, SRL^{+}04]. helices [KNI^{+}04]. Helicobacter [CAGK^{+}03, Wel03w]. helix [Wel02-30, MMH^{+}00]. help [Dov03-27, LeB02-68, LeB04-89]. helped [LB03]. hemagglutinin [MLZ^{+}01, SKT^{+}03, AKW00]. hematopoiesis [KRZ^{+}04, GJ00]. Hematopoietic [DLR^{+}01, KMiM^{+}01]. Hemichannels [QRLL00]. hemicomplexes [JKB^{+}03]. hemidesmosomes [MKD^{+}01, SGO^{+}00]. Hemifusion [AKW00]. hemolysin [GGT^{+}02]. hemorrhage [XJW^{+}04]. hensin [TAA04, HVT^{+}00]. Heparan [Dov01o, SYW^{+}03, ADL^{+}03, AFR01, AR03, KRU^{+}04]. Heparin [LeB03-44]. hepatic [TNBH01]. Hepatocellular [WFF^{+}01]. hepatoma [dMMBK^{+}02]. HepG2 [dMMBK^{+}02]. HER4 [WAV^{+}04]. Here [Der01, Wel02-46]. hereditary [ASK^{+}03, EMY^{+}04]. Herpes [CDW^{+}03, Dov02-30, SLB02]. herpesvirus [CSG01, TE01, TGM^{+}01]. herpesvirus-encoded [CSG01]. HES6 [GCG^{+}01]. Hesperadin [HCL^{+}03]. Heterochromatin [PFB^{+}03, GBMA04, JKG^{+}02, LeB04i, LCI^{+}01, TSL04, TBW^{+}04, AH01]. Heterochromatingets [Wel03x]. Heterodimer [SBH00]. Heterodimeric [LDS^{+}00, KRF^{+}03]. Heterodimers [STP^{+}00]. Heterogeneity [TCS01]. heterogeneous [PDMO03, PZP^{+}01]. heteropentameric [TNW^{+}02]. heterosynaptic [UTH^{+}02]. heterotrimeric [NB03a, YCK^{+}03]. Heterotypic [PMB^{+}00, SGG01]. Heterozygously [CMW^{+}01]. HGF [MNC^{+}03, SBG^{+}04]. HGF-induced [MNC^{+}03]. HGF-R [SBG^{+}04]. HGF-R/Met-mediated [SBG^{+}04]. hGle1 [KBG^{+}03]. hidden [Dov01h, Dov01p]. Hide [Wel03y]. Hiding [LeB02-45]. Hierarchical [Dov03z, SWB03]. Hierarchy [WMCW03, HTRK02]. High [BJM^{+}02, WPS^{+}01, FRT^{+}01, MKTW01, PPP^{+}01, RCS^{+}02, DBS^{+}01, MF01a]. higher [CKW^{+}03]. Highly [HES00, MTM^{+}02, SPH^{+}00, Zwe00]. highway [LeB04-28]. Hijacking [BS02]. Him [HMAM01]. Him-10 [HMAM01]. hinge [ALEH02]. Hip [Sre04a, Wel01-29]. Hip1R [EGWK^{+}01]. HIPK2 [WWD^{+}04]. Hippi [Wel02-29]. hippocalcin [OTB03]. Hippocampal [DFYL00, WSL^{+}00, EOB^{+}04, LRBH02, UTH^{+}02, WAC^{+}03, WXQ^{+}00]. hippocampus [ACBG04, HIN^{+}03]. Histamine [CPN^{+}01]. Histocompatibility [MGL^{+}00]. Histone [KMG^{+}03, KLT^{+}03, MSR04, Wel03e, AMEC01, CW01, CNJ01, GG01, GZY^{+}00, MJG^{+}01, RME^{+}00]. Histones [Dov03-27, Wel01-28, KC01]. Hitched [LeB04-44]. Hitchhiking [Sub02]. hither [LL03a]. HIV [AL03, DM00, JKB^{+}03, MVL^{+}02, MMH^{+}00, PMKM03, PHS^{+}03, RWK^{+}04, SPMM^{+}00, Sod02, WSW^{+}00]. HIV-1 [DM00, JKB^{+}03, MMH^{+}00, PMKM03, RWK^{+}04, WSW^{+}00]. Hmg [DBS^{+}01, CKFH00]. Hmg1 [DBS^{+}01]. HMGB1 [PSD^{+}04a]. hMis12 [GKYY03]. Hnrnp [vdHvODML^{+}00]. hnRNPI [HIG^{+}01]. hNuf2 [DMH^{+}02]. Hold [Zet01, LeB02-94]. holds [OM01]. hole [LeB02-70]. holes [Wel01-65]. Holey [MB03b]. holiday [LeB04y]. holocentric [PRS^{+}04, HMAM01]. holocentrics [LeB04-38]. home [Lai03c, LeB02s, LeB02-38]. Homeobox [MCB00b]. homeostasis [CRH02, MCF^{+}02]. Homer [SDL^{+}03]. homing [TCP^{+}03]. Homologue [BYLA^{+}01, CQH^{+}00, COB01, DC02a, JGW02, LRF^{+}02b, PCB^{+}03, RP03b, SNF^{+}02, TMG03, VLL^{+}03, WHB04, OSMF00, ONS^{+}00, Ono01, PDV^{+}00, TGM^{+}01, TSDS00]. homologues [RML^{+}02]. Homology [RDH^{+}01, ESC^{+}01, KJK^{+}03, XLGS01, FMMF01, HCD^{+}00, KMH^{+}00]. Homomultimerization [SMSM00]. homophilic [CFWH^{+}01, ICK^{+}04, TNK^{+}00]. homotypic [BMLU02, BBSF01, WSK^{+}03, PWU00]. hood [Wel03k]. Hook3 [WDLK01]. hooks [Dov01s]. hop [FRM^{+}02]. hormone [CDM^{+}02, LeB04-81, SJW^{+}04, SXD^{+}03]. hormone-dependent [CDM^{+}02]. hormone-sensitive [SXD^{+}03]. Host [CSJ00, GGT^{+}02, RZB^{+}03, SBG^{+}04, SJ01b]. Hox [LeB03-46]. HP1 [PFB^{+}03, Kel03]. Hr12 [RGGL00]. HRD [HCC02b]. Hrp36 [PZP^{+}01]. Hrs [BBMS03, SYVB03, AL03, PHS^{+}03]. Hsc70 [CNH^{+}02, CHM04, NS01, YRC^{+}04]. Hse1 [BWK^{+}03]. hSIRT3 [SNF^{+}02]. Hsl1 [WBAS04]. Hsl7 [YDPK04]. Hsp40s [JC01]. Hsp47 [NHI^{+}00]. Hsp60 [KKPB03]. Hsp70 [BHW02b, DDW^{+}01]. Hsp90 [YMH01]. hugs [LeB04-88]. Human [CHA^{+}01, DGB^{+}00, FLLE^{+}01, FCBH01, GKG^{+}01, GZY^{+}00, GKYY03, HdEV^{+}02, JBK04, KSK^{+}02, LYKH00, MSM04b, PFW^{+}00, PDT^{+}02, SMLM00, SCLC00, XB04, BBR04, BSW^{+}04, BXR^{+}02, BFSO^{+}04, CPN^{+}01, CLB^{+}02, CGBL^{+}02, CBS04, DDV^{+}03, DH02, ERMT^{+}04, FCF^{+}01, GMRYM^{+}02, GJS^{+}03b, HViV^{+}02, ITM^{+}04, JKG^{+}02, LGM^{+}04, LP04a, iONOM02, PSS^{+}04, PPWM04, PHS^{+}03, RDC^{+}04, RKF^{+}04, SMT^{+}03, SMS^{+}04a, SNF^{+}02, SMS^{+}01a, STJ^{+}01, TSL04, ZMGL02, ZAE^{+}04, DLR^{+}01, DOL^{+}01, FFKC00, FSBH00, GHS00, IAG^{+}00, KMiM^{+}01, KC01, KKK^{+}00, LLH^{+}01, OSMF00, PGSL00, SMI^{+}00, SWJ^{+}00]. hunchback [Wel02-46]. Huntingtin [MZH^{+}02, ZCW^{+}03, CLT^{+}01]. Huntington [Dov02-49]. Hur [BGS00]. Hvps45 [NCUJ^{+}00]. Hyaluronan [SRD^{+}02]. Hyaluronan-CD44 [SRD^{+}02]. Hyaluronic [OKSH00]. Hybrid [PMB^{+}00]. Hydrogen [CAP00, ARMB04, JOF^{+}02]. hydrolysis [GBZ^{+}02, Gre00, TRW^{+}00]. hydrophobic [ORZ^{+}04]. Hydroxylase [HAM01]. hypercontraction [ML04]. hypercontraction-induced [ML04]. Hyperphosphorylation [DSV^{+}00]. Hyperplasia [HGK^{+}01]. hypertonic [DC02a]. hypertrophic [BBDM02, SDML04, ZZM^{+}03, SCTM00, YCX^{+}01]. hypertrophy [AKH^{+}04, GPLS02, HIO^{+}04, HYMS^{+}02, PR02, HCK^{+}00, HGB^{+}00]. Hypoacetylation [CNJ01]. Hypomyelination [WFQ^{+}00]. Hypophosphorylated [BPD^{+}04].

I-activated [MZ00]. i-SNAREs [VVR^{+}04]. IA [KPA^{+}03]. IAP [WLPD04]. Ib [DLT^{+}02]. ICA512 [TMA^{+}04]. Icad [LDS^{+}00]. Icad/Cad [LDS^{+}00]. ICAM [BYMS^{+}02]. ICAM-1 [BYMS^{+}02]. ICAP [DBB^{+}02a]. ICAP-1 [DBB^{+}02a]. Identification [BPPFM^{+}03, BSL^{+}01, CKL^{+}03, CS02, GGD^{+}04, IHN^{+}01, ICN^{+}03, LYKH00, LVD^{+}04, MPSM00a, QPDJ^{+}02, RB01, RVB^{+}01, SHCM03, SAS^{+}02, TMHP00, TAA^{+}02, TCP^{+}03, TEB^{+}03, SwZK^{+}02]. identified [DMBS02, LeB02-74, VMD^{+}01, WS00]. identify [SMC^{+}02]. Identities [SSW^{+}01b]. identity [PRJK01, SK01]. Idiosyncratic [LMW^{+}00]. If [Dov02-33, Tum04-28]. IFN [FNKH02]. IFN- [FNKH02]. IFT [QDG^{+}04b, PDV^{+}00]. IFT88 [PBD^{+}02]. Igcam [Rut00]. IGF [EPN^{+}03, HAK^{+}04]. IGF-I [HAK^{+}04]. IGF-II [EPN^{+}03]. II [AW00, AAD03, CLB^{+}02, CH03, DSSWW00, DRG^{+}03, EPN^{+}03, FLS^{+}03, FNKH02, HDP^{+}01, pHYpXL00, JKB^{+}03, KSC02, KRS^{+}01, KBGG04b, KI04, KMP02b, LeB02-101, LP04b, LVD^{+}04, MGL^{+}00, MSR04, MMBM04, MWC^{+}02, PCR^{+}01, PBL^{+}00, RH00, RSK02, SGW^{+}02, Yum01, vMZM^{+}00, KSF^{+}00, CLB^{+}02, TCH^{+}02]. II-dependent [SGW^{+}02]. II-like [DSSWW00]. IIa [SMZ^{+}03]. IIa-mediated [SMZ^{+}03]. IIB [DLY^{+}02, OEM^{+}02, dVKS04]. III [BVH04, Cor02, HDL^{+}00, JMG^{+}04a, Mar01, PPWM04, VBR^{+}01a]. IIIb [DLT^{+}02]. IL-1 [CSIK03]. ILK [WD01]. ILKBP [FGSW03]. image [HZS^{+}01, RY04]. Imaging [OGD03, SGAS00, SDN^{+}00, EA03, HGP^{+}04, IBP^{+}04, SP03, MPR^{+}01]. Imbalance [PLL^{+}01]. immaturity [Wel02-39]. immobile [DBH^{+}01]. immobilizes [LeB03x]. Immortalization [DGB^{+}00]. Immune [LeB03-47, LeB04-45, Wel04-39, CSG01, LeB04-105, MM01, SR03, Wel03l]. immunity [LeB02-102, LeB04-80]. immunoglobulin [RWCC01]. immunoglobulin-like [RWCC01]. immunologic [LW03]. immunological [SDD04]. Impact [EKdM^{+}04]. Impacts [HW04]. Impaired [CKF^{+}03, HSC01, MDT^{+}01, BBP^{+}04]. Impairment [FHL^{+}03]. impairs [ABRA03, BFH^{+}01, PSW^{+}02, RLTC^{+}02]. impedes [FPP^{+}02]. Imperfecta [JPM^{+}00b]. implicated [EWD02, GWL03, KPKY^{+}03, LSMS03, MEFC03, JWJJ00]. Implication [CBW^{+}01]. Implications [VDMH01]. Import [GDHS01, HM00, RM01, YB01, BHW^{+}02a, CLSK02, FMJG04, HBH^{+}04, HBV^{+}01, HDL02, JMG04b, LeB02i, LGB^{+}02, MSY^{+}04, PPWM04, PC01, SRW^{+}04, TVF^{+}03, WPC^{+}02, YTM03, BEG01, BMC00, CAP00, DSH^{+}00, KRR^{+}01, LCM00, MJG^{+}01, NM00, YF00]. Importance [SLR^{+}03, Wel03-28, WBM^{+}00]. important [BCG03, LMDS^{+}03, RRSV02, UJL^{+}03, UIY^{+}01, SGG01, TYS^{+}00]. imported [TNW^{+}02]. Importin [BCG03, RM01, HDL02, MSY^{+}04, BEG01]. importin- [HDL02]. inactivates [LP04a]. inactivating [CACL03]. inactivation [BS04a, CLG^{+}03, HMC^{+}01, CNJ01, RME^{+}00]. inactive [CW02, EWD02, KRF^{+}03, MPG^{+}02, MAG^{+}04, PTH^{+}04, CW01, RME^{+}00]. Inad [LM00b]. INCENP [sKCK^{+}01, AMEC01]. inch [Wel04-52]. including [ASYL04]. inclusions [LeB04-32, NMG04, MMDC00]. incongruous [MWSL^{+}03]. incorporation [JFS^{+}03, GPAS^{+}01]. increase [BWK^{+}03, BND^{+}02, DBLG02]. increased [ASK^{+}03, CLG^{+}03, CLM^{+}03, HBAF^{+}02, MRT^{+}01, NMG04, SGdM^{+}01, CKS^{+}00, KIK^{+}00b, KH01, THK^{+}00, ZEtK^{+}00]. increases [LeB02r, MO01, SMS^{+}01a, MPSM00a]. increasing [DKA01, BEG01, KSN^{+}01]. Independent [FSM^{+}01, HM00, AHA^{+}04, BCM04, CJ02, CSJ03a, CSJ^{+}03b, CFM^{+}02, FS03, GKYY03, HP04, HCC02b, HDL02, JHS^{+}02, KBWG02, KPL^{+}02, KMLS04, LeB02j, LeB04-44, MSGS02, OMB^{+}01, PWS^{+}01, PL01, RPZ^{+}02, SHW01, ST03, TST^{+}03, VLL^{+}03, WG03, WK04, XLGS01, YpHRL03, ZLR^{+}03, GWG01, CSP^{+}04, RBE^{+}02, TJC^{+}03]. Independently [SKF^{+}01, BSD^{+}01, JCPWS01, LG02a, MEV^{+}04, NMH^{+}04, SZvBuH^{+}04b, DR00b, DWD^{+}00, SBC01]. indicate [DCC^{+}02]. Indications [vdVWS^{+}00]. indicator [TNMM03]. indirect [KNK^{+}01, LeB04-56, MDF01]. individual [CS04, KAC^{+}04]. induce [AvdWM^{+}01, DLPB03, GCG^{+}01, LLGB03, LRS^{+}02, MSY^{+}04, NH03b, SCD02, AML00, CGL^{+}01, DBS^{+}01]. Induced [BF01, MHE^{+}00, RGGL00, WYHP00, WSL^{+}00, ZBB^{+}00, BSD^{+}01, BDKM04, BGR^{+}01, BND^{+}02, CDM^{+}02, CRS^{+}03, DBL^{+}02, EES^{+}01, GPLS02, HKE^{+}04, IIN^{+}01, JMG^{+}04a, LCM02, LZC^{+}03, LFM^{+}04, MKM04, MMM^{+}04, MO01, ML04, MRT^{+}01, NOOG^{+}04, OTY^{+}04, PKR^{+}02, PFB^{+}03, RKR^{+}03, RLTC^{+}02, RDS02, SMT^{+}03, SSM^{+}04, SMS^{+}01a, SK04, SM03b, SNS^{+}04, TFAM^{+}04, Wel01o, WHP^{+}02, WSR03, YSK^{+}04, YHZ^{+}01, YSC^{+}21, ZRDG02, AV01, BKD^{+}00, CWMO00, Ern00, FRO01, FWM^{+}01, Gau00, GLA00, GLK^{+}00, GV00, HGB^{+}00, HVM^{+}00, KH01, LCGR00, MMDC00, PSKK^{+}00, PWC^{+}01, RRK^{+}00, SFSD00, SSW^{+}01b, TN00b, WXQ^{+}00, YHD^{+}00, BLU^{+}04, CLM^{+}03, HKE^{+}04, MNC^{+}03, SMS^{+}04a, TPW^{+}04, TST^{+}03, YSC^{+}02]. Induces [DSV^{+}00, FSM^{+}01, ZEW^{+}01, BBDK^{+}04, BJM^{+}02, BSMS03, BHK^{+}02, CPN^{+}01, CFC^{+}01, CNMS02, DMH^{+}02, FAF^{+}04, HdO^{+}01, HN03, JMB^{+}04, KSK01, LXL^{+}04b, MH01, MSI^{+}03, MBM^{+}04, MTPT02, PSD^{+}04a, PUK02, REK^{+}03, SHW01, WKYC02, WDW03, HPQ^{+}00, IEJ^{+}01, LSA^{+}00, LBWH^{+}00, MBSB00, MPSM00b, MMH^{+}00, OKSH00, PLL^{+}01, RBBA00, SME^{+}00, TGM^{+}01, TNK^{+}00, THK^{+}00, WBC^{+}00, WFF^{+}01, YHF^{+}01]. Inducible [WFC^{+}02, BRR^{+}02, HSC01, CLWR01]. Inducing [LeB03-48, APM^{+}02, CFM^{+}02, HPFG03]. Induction [HVT^{+}00, HBAF^{+}02, KFS^{+}00, MH02, VHLS00, AZ03, LCT^{+}04, MBLCE03, PGSE^{+}01, SGPL^{+}00]. inductive [WSL^{+}03]. infect [Wel03o]. infected [GGT^{+}02, SLB02]. infection [AAM^{+}04, LBH^{+}02, PAG02]. infections [Gal02]. Infectious [PMKM03]. infectivity [KSD04]. inflammation [DPO^{+}04, LeB03-69]. inflammatory [LTD^{+}01]. Inflexibility [LeB04-46]. Influence [LGB^{+}02, MPB02]. Influenza [HGP^{+}00, MLZ^{+}01, SKT^{+}03, AKW00]. influx [DFZ^{+}03, NMH^{+}04, VMS^{+}02, YpHRL03, PFM^{+}00]. influx-independent [YpHRL03]. information [SNF^{+}02]. ing [LeB03-34, Sre04k]. ingression [ISS^{+}04]. Inheritance [CKW^{+}03, Bar04, EKC^{+}03, HvdBP^{+}01, JCPWS01, RRB^{+}01]. Inhibit [DDW^{+}01, BFC^{+}02]. inhibited [CMS^{+}02, Dov02f, KSR^{+}04, BMD^{+}00]. Inhibiting [GJB^{+}00, SYVB03, TYA^{+}02, YSC^{+}02, YSC^{+}21, GWG01]. Inhibition [JON^{+}03, KNM^{+}00, KRA^{+}01, NZHR01, PDJ00, DMH^{+}02, HIE^{+}01, JHS^{+}02, LZC^{+}03, LeB04-83, LRB^{+}03, MWC^{+}02, PGSE^{+}01, SMT^{+}03, SMS^{+}04a, SHM02, SCY01, TEB^{+}03, Wel04-47, YRC^{+}04, CAP00, DKJ00, SMI^{+}00]. Inhibitor [RGGL00, CACL03, Dov03t, Dov03u, FNFL03, iHGK^{+}02, LeB03-72, LeB04-59, WNM^{+}03, BMG^{+}01, KMCM00, MLS^{+}01, RMMC01, WTG01]. inhibitor-1 [CACL03]. Inhibitors [IWG^{+}01, SSL^{+}00, MSR04, GBON00].

Inhibitory [HBSQ01, SSW^{+}01a, WSL^{+}03]. Inhibits [CGY^{+}01b, DSV^{+}00, HGP^{+}00, PTM^{+}01, ZEW^{+}01, AKH^{+}04, CJ02, CARG^{+}04, DDL^{+}04, EGC^{+}03, GSB^{+}03, GSSP03, HGC02, KCWF02, LTF^{+}01, OO02, PBD^{+}04, SGW^{+}02, TJC^{+}03, TBRG01, ALJ00, YHF^{+}01]. initial [FMF^{+}04, GJS^{+}03a, JB01, KS02, LGGS^{+}04, YSK^{+}04, BAD^{+}00]. Initiate [Dov02-34, BTH^{+}03, ISS^{+}04, OEM^{+}02, ZLH^{+}03]. initiates [KYM04, MRM^{+}04, PWS^{+}01]. initiation [BS04b, CME^{+}02, JON^{+}03, MEFC03, SM04b, SBC^{+}03, DB00, DLS00, FHP00, TSMS01]. Injection [TTP^{+}01]. Injury [AKDS00, DWM03, PSW^{+}02, SGK^{+}02b, PWC^{+}01]. innate [MM01, SR03]. inner [CLSK02, sKCK^{+}01, LeB04-41, LBS^{+}02, MJV^{+}03, OSNG04, Tum04-27, AMEC01, FRK^{+}01, PS00]. innovation [DV04]. Inositol [FSGDN^{+}00, CVZ^{+}04, DOB^{+}01, SDDS00]. Inositol-Anchored [FSGDN^{+}00]. Inoue [DV04]. insensitive [TNMM03, MAAZ^{+}00]. insertion [ALP^{+}04, DK04, OWW02, LW00, SBI^{+}00]. inside-out [XBL^{+}03]. insight [WHD^{+}03]. Insights [YYM^{+}01, AG03, GKSR00, MR02]. insipidus [KBK^{+}03]. insoluble [NOS^{+}01]. InsP [TFAM^{+}04, WLW^{+}04]. instability [GOV^{+}03, GK03b, JdDD03, SGdM^{+}01]. instructs [HAK^{+}04]. insulator [BC03]. Insulin [AOH^{+}02, LeB04-47, LeB04-48, BMM^{+}02, Dov04e, GFM^{+}04b, HNK^{+}03, KMH^{+}04, KJK^{+}03, LeB03-84, THO^{+}04, TIS^{+}01, TMA^{+}04, Tum04m, UML^{+}03, WSC^{+}01, Wel01p, GKM^{+}01, LR00, PGS^{+}01, yZCKA01, HFG^{+}04]. insulin-like [BMM^{+}02, GFM^{+}04b, HNK^{+}03, TIS^{+}01, LR00, PGS^{+}01]. insulin-stimulated [KMH^{+}04]. Intact [NDM^{+}03, OAR^{+}00]. Integral [MMS00, CMM^{+}02, OSNG04, VTGT^{+}03, JMG01, NKH^{+}00, SKCS00]. Integrase [KM00]. integrated [LHR04]. integrates [SERB03, KYF00]. integration [DK04, WAOC^{+}03, XTN^{+}02]. Integrin [ACMR04, ASAJ01, BS04a, CRS^{+}03, CBL^{+}02, GMRS00, ISID^{+}03, KHH^{+}01, KRA^{+}01, LeB03-49, LG02b, MPAP^{+}03, NLRD01, PCK^{+}00, SHN^{+}01, WD01, ZN01, BBR04, BFG^{+}04, BSL^{+}01, CYC^{+}04, DBB^{+}02a, DH02, Dov01i, EPH^{+}02, EPH^{+}03, FNZ^{+}03, FPP^{+}02, GPDvH^{+}03, GCO^{+}04, GJS^{+}03a, GHK^{+}03, HBB^{+}02, HFK^{+}03, JW04, KSC^{+}04, KYS^{+}02, KCY^{+}04, LMGM^{+}02, LCS^{+}02, LDI^{+}03, MZT^{+}03, MKD^{+}01, MCF^{+}02, OEM^{+}02, REK^{+}03, SHPY02, SKN^{+}03, SKH03, SDD04, TPA^{+}03, Tum04e, WKYC02, WFI^{+}04, XBL^{+}03, YSK^{+}04, YKT^{+}04, YMK^{+}04, ZLR^{+}03, ZSY^{+}03, ZLVS02, ZTK^{+}03, dVKS04, vdFKK^{+}02, vWJK^{+}03, AKFB00, ALJ00, BMD^{+}00, BKD^{+}00, BWN^{+}01, HBSQ01, HLZW00, IAG^{+}00, LWDH01, LLH^{+}00, MGLPM00, ONM00, PWC^{+}01, RBM^{+}00, RRK^{+}00, RMG^{+}00, RBBA00, SMSM00, SGO^{+}00, TWBV^{+}01, THZ^{+}01, WGF^{+}00, YSS^{+}01, ZGB01, CWA^{+}03]. Integrin- [NLRD01]. integrin-dependent [GCO^{+}04, KYS^{+}02, SKH03]. Integrin-linked [CRS^{+}03, WD01, TPA^{+}03]. Integrin-Mediated [ASAJ01, GMRS00, KRA^{+}01, SHN^{+}01, LMGM^{+}02, REK^{+}03]. Integrin-specific [MPAP^{+}03]. integrin-stimulated [ZSY^{+}03]. Integrins [CBK^{+}00, CROfC04, ELO^{+}01, LeB04-49, Wel04-40, WW02, CACL03, DSB^{+}02, GMRYM^{+}02, GFM^{+}04b, MGAL^{+}01, SC04, SPB^{+}01, Tum04s, APLB00]. Integrity [CENMR^{+}01, BWV^{+}01, BFH^{+}01, RB03, SSGLS01, WPC^{+}01, WNM^{+}03, FSKS00]. intensity [SDL^{+}03]. interact [BPKK01, DSH^{+}00, TNM^{+}00]. Interacting [HMN^{+}00, RDH^{+}01, TCH^{+}00, GSN^{+}04, SHE^{+}02, TONN02, WXD^{+}03, ABP^{+}00, Ono01, CARG^{+}04, HGS^{+}01, LTD^{+}02]. Interaction [FFY^{+}00, NHB00, RH00, TUK03, WWD^{+}04, BYMS^{+}02, BGA^{+}04, BDR^{+}03, BSEB04, DMBS02, Dov03-29, DSB^{+}01, FM01, KSM^{+}01, KMS^{+}04, KJK^{+}03, LMVW03, MEFC03, MHH^{+}03, MRC^{+}02, MKS^{+}02, MGMH03, MPR^{+}03, NMH^{+}04, NH03a, RPZ^{+}02, RAD^{+}02, SRL^{+}04, SKT^{+}03, SRD^{+}02, TRMI^{+}04, TIS^{+}01, WZB^{+}01a, YSK^{+}04, vdBCH^{+}04, BMG^{+}01, BLC00, BZSC00, BGS00, FSK^{+}00, sGS01, HSW00, MPR^{+}01, MDJF00, MRM^{+}00, PCR^{+}01, RLC00, YSS^{+}01]. Interactions [EKT^{+}00, GTD^{+}02, LCS^{+}02, POW^{+}01a, WKS^{+}00, BGH^{+}03, CSSBY^{+}03, FRT^{+}01, FPP^{+}02, FSK^{+}04, GSB^{+}03, LeB02-31, MDP02, MW04, Néd02, NGKH02, OEM^{+}02, SWBE^{+}04, SHF^{+}03, SPA^{+}04, TDFV02, YIW^{+}04, dVKS04, AC00, HCOC00, MKK^{+}00a, SGG01, TNB01]. Interactor [HCD^{+}00, MPSM00a]. interactors [WD01]. Interacts [EM00, ELO^{+}01, SZZ^{+}00, WND^{+}00, WFT^{+}01, CHM04, CVZ^{+}04, IKS^{+}02, JGW02, KMG^{+}03, KK02, KSK01, KGT^{+}02, MSJ^{+}02, MKS^{+}02, MWK^{+}02, SDL02, YSK^{+}04, ZLVS02, ALJ00, FK01, HJSM00, HCD^{+}00, SKCS00, TM00, TN00a, THZ^{+}01]. Intercalated [EHZ^{+}01]. Intercellular [NGK^{+}03, TNK^{+}00, CiKBG03, HV03, HPG^{+}02, KvHB^{+}01, LM01, LeB04-28]. interchromatin [SBS02a]. Interdependent [FJM^{+}04]. interendothelial [CMS^{+}02]. Interference [DW02, Pow01h]. interferes [ERMT^{+}04, SDD04]. Interfering [NS01]. interleukin [LTD^{+}01]. Intermediate [CG00, HPG^{+}02, BMY^{+}01, HMVG02, LSS^{+}02, WC03, XH04, YRC^{+}04, FK01, sGS01, LVWA01, WPJ^{+}00, YYM^{+}01]. intermediates [ATF^{+}04, SN04, DE01]. Intermembrane [DSH^{+}00, WWG^{+}00]. Internal [DC02a, YSN^{+}01, BHL^{+}03, FCBH01, CMMP00, SGAS00]. internalization [DFZ^{+}03, GMRYM^{+}02, HCC02a, JK01, PWS^{+}01, SLB02, UIY^{+}01]. internalized [HBG^{+}02, SG00]. interneuron [ACBG04]. interphase [BNSR03, CGBL^{+}02, MMFS01, WSC^{+}03, CH01]. Interplay [UHR^{+}03, FJK^{+}04, KRZ^{+}04]. interstitial [TAA^{+}02]. Intestinal [FCL^{+}00, LSA^{+}00]. intestine [BvdWD^{+}04]. Intimacy [LeB04-50]. intolerant [Wel04g]. intra [MMPO^{+}01, SDL02]. intra-Golgi [MMPO^{+}01, SDL02]. Intraarterial [TTP^{+}01]. Intracellular [AZB^{+}00, DDW^{+}01, pHYpXL00, MMG^{+}01, PHWK^{+}00, yZCKA01, BWRT03, BDR^{+}03, HTRK02, MVL^{+}02, OKM^{+}01, PBT^{+}02, SJW^{+}04, SLR^{+}03, Sod02, ZMGL02, DZT^{+}00, HLK01, MLLA00, PFM^{+}00, PLL^{+}01, RLC00]. Intraflagellar [MR01a, QDG^{+}04b, JGR^{+}04, PBD^{+}02]. Intramitochondrial [AERD^{+}01, DE01, WWS^{+}03]. intranuclear [CPC^{+}02a, LCRS01, NMG04, SMI^{+}00]. Intraorganellar [PMB^{+}00]. Intrinsic [PHM^{+}02, BSD^{+}01, GOV^{+}03]. intrinsically [BCA^{+}03].

Introducing [WCGT^{+}00, FFST01]. invading [Dov01t]. Invadolysin [MKY^{+}04, Tum04j]. Invagination [HvdKDS01, SM00]. Invasion [CK00, BGW^{+}04, Dov01-35, Dov02a, GTR^{+}03, HEN^{+}01, HMH^{+}03, JBA^{+}01, UIY^{+}01, Wel02-33, WG03, BZSC00, HPQ^{+}00, HAP^{+}00, SGPL^{+}00]. invasion-promoting [UIY^{+}01]. invasiveness [KvHB^{+}01]. Inventories [AG03]. Involucrin [DEG00]. involution [SHB^{+}03, ASMW01]. involved [CSG01, CFM^{+}02, CRH02, DK04, HBSJ04, KGT^{+}02, MTM^{+}02, MHlW02, OER^{+}02, RRB^{+}03, SMC^{+}02, VTGT^{+}03, WHM^{+}02, YEG01, ARK^{+}00, CH01, MHW^{+}00, NS00, RIDC01, SMWG00, SYH^{+}01, TOM00, WBG01, WAPB^{+}00, WMHB^{+}00, WWKV00, YSS^{+}01]. Involvement [HKE^{+}04, SGG01, VMS^{+}02, ARMB04, LSSL00, PPR^{+}00]. Involves [ESS^{+}00, GDHS01, NHB00, GBZ^{+}02, HCC02b, KNI^{+}04, MTG^{+}02, NYT^{+}03, SKO04, UML^{+}03, AFN00, BKD^{+}00, GZY^{+}00, KRR^{+}01, LCGR00]. involving [ARQ^{+}04, IKA^{+}03, JBA^{+}01, KMS^{+}04, LL02]. Inward [FLLE^{+}01]. ion [DB02, SSGLS01, Coo01, DAC00, EHCC^{+}00]. Ionophore [ARMB04]. Ionophore-resistant [ARMB04]. Ip [HGC00]. Ipl1 [BTVB03, sKCK^{+}01]. Ipl1/Aurora [BTVB03]. Iqg1p [OKC02]. IQGAP1 [ISB^{+}04]. Ire1 [KOS^{+}04]. Ires [CMMP00]. IRS [HFG^{+}04]. islets [BFSO^{+}04]. Isoenzyme [HSW00]. Isoenzyme-Specific [HSW00]. isoenzymes [LRA^{+}02]. Isoform [UML^{+}03, BBG^{+}03, MTG^{+}02, WLR01, EMW^{+}01]. Isoform-specific [UML^{+}03]. Isoforms [QK00, TTS00, BHL^{+}03, GAC^{+}03, RFLT02, BSS00, LLH^{+}01, PVjL^{+}00, SEP^{+}01]. Isolated [KKA^{+}01, SCP02, BMD^{+}00, DSSWW00]. Isolation [WMHB^{+}00, LQPC^{+}00]. isomerase [TR02]. Isomerases [NWT^{+}01]. isopeptidase [LH03, MCU04]. isozymes [PLW^{+}04]. ither [Nel00]. IV [KS02, LGGS^{+}04, BAD^{+}00, WWK^{+}00, XRP^{+}01]. IV-spectrin [KS02].

J [CHM04, TVF^{+}03]. J-domain [CHM04]. J-protein [TVF^{+}03]. JAC [PCB^{+}03]. JAC-1 [PCB^{+}03]. JAK [Wel02-33]. JAK/STAT [Wel02-33]. JAM [ISF^{+}01, Dov02o, LeB02-102]. JCB [Mel04a, Ros03, RO04, SM04a]. JEM [SM04a]. Jil [JWJJ00]. Jil-1 [JWJJ00]. Jip [VMD^{+}01]. JNK [KSC^{+}04, MMG^{+}04, GJB^{+}00, KCL^{+}01]. job [Dov02-42]. Joined [Wel01-29]. Joining [PGSL00]. joins [LeB02-62]. journal [Mel01]. July [Pow01c, Wel01-28, Wel02-51]. jump [Wel02i]. Jun [PSKK^{+}00, PKR^{+}02, PBD^{+}04]. JunB [Dov04c, KHB^{+}04]. Junctate [TFAM^{+}04]. junction [BPPFM^{+}03, CSG^{+}04, GI02, GF03, GSP^{+}02b, GGH^{+}04, HP04, IKS^{+}01, KLZ04, LM01, MCA^{+}03, NCMO^{+}02, STA03, SAS^{+}02, TLS^{+}01, WCIN04, Wel02-36, WSH^{+}04, TGMC^{+}00, WHS00, ZLG00]. junction-dependent [HP04]. Junctional [ISF^{+}01, LWCKL01, RDC^{+}04, LTB^{+}00, LGP00, QRLL00, RGG00, SYH^{+}01, WHS00]. junctions [BD02, CiKBG03, Dov03j, Dov04d, FHF^{+}02, HYMS^{+}02, KvHB^{+}01, PMU^{+}02, STA03, Wel01e, Wel03w, Wel04-32, XLH^{+}01, HW00b, JCR^{+}01, KYF00, LM00a, TK00]. junctophilin [IKS^{+}01]. June [MR02, Wel02-51]. just [LeB02-87, Wel04t]. juxtamembrane [HT01]. Juxtaparanodal [PSE^{+}03, TGD^{+}03].

K-loop [OWW02]. K6 [WLR01, WC03, WCGT^{+}00, WCGT^{+}00]. K6a [WLR01]. K6b [WLR01]. Kap142p [YB01]. Kap142p/Msn5p [YB01]. Kappa [MHE^{+}00]. Kar9 [LeB03-80]. Kar9p [HBB^{+}04]. Karyopherin [YB01, BM02]. Karyopherins [LMM^{+}02, MJG^{+}01]. Kazrin [GSN^{+}04]. KCC2 [RLTC^{+}02]. Kd [GKM^{+}01, GBOL03, SSG^{+}02]. Kdel [TOM00]. keepers [LeB04-98]. Keeping [LeB02-47, LeB03-50, Pow01g]. keeps [Dov02-61, LeB02d, LeB03-96, LeB04-48, Wel03g]. Kelch [KHC02]. Keratin [CWMO00, IIN^{+}01, JKW^{+}03, LeB03-51, BMY^{+}01, WLR01, WC03, WPJ^{+}00, YYM^{+}01]. Keratin-Dependent [CWMO00]. Keratinocyte [BD02, CGM^{+}02, CAW^{+}04, GSN^{+}04, HBSQ01, THE^{+}00]. Keratinocytes [DGB^{+}00, SMT^{+}03, SMS^{+}04a]. Keratins [Bir04d, KO00, TKB^{+}04, Dov01j, GLDM01, KFO04]. Keratocytes [RRJ^{+}01]. Kes1p [LRF^{+}02b]. Kettin [HES00, KNK^{+}01]. KEULE [AHMJ01]. key [FCM^{+}01, LLGB03, SMT^{+}03, TFAM^{+}04]. KH [FHM^{+}03]. kick [LeB02-48]. Kicking [LeB03-52]. Kid [LC01]. Kidney [HGP^{+}00, AFK^{+}03, BSL^{+}01, LeB02-72, MBMMA^{+}03, CNBWN00, FFST01, PDV^{+}00, VHLS00]. Kif2a [GC04, LeB04-51]. Kif3 [TYS^{+}00]. Kif4 [PPR^{+}00]. KIF4A [MSM04b]. KIF5A [XRH^{+}03]. KIFC3 [NOS^{+}01, XTN^{+}02]. kill [LeB03-94, LeB03-55]. Killer [SNL^{+}00, Wel03-29]. Killing [SNL^{+}00, Dov03-43, LeB03-35, Wel01-30]. Kilobases [BGFJ01]. Kilodalton [WFT^{+}01, BMM^{+}01]. Kinase [CS01, CK00, DSSY00, Dov04h, pHYpXL00, KFS^{+}00, KKA^{+}01, KY00, MBS^{+}01, RDH^{+}01, SNL^{+}00, SHN^{+}01, SCLC00, TYY^{+}00, WWK^{+}00, WFT^{+}01, BGR^{+}01, BND^{+}02, BSR^{+}03, BTVB03, CGM^{+}02, CMM^{+}02, CFC^{+}01, CPC^{+}02b, CWG^{+}02, CRS^{+}03, CS02, Dov03g, DFZ^{+}03, EPH^{+}02, FHL^{+}03, FBG^{+}01, FGSW03, GCT02, GJS^{+}03a, GMD^{+}02, GLJP01, GSB^{+}01, HKHO01, HZS^{+}01, HLB^{+}02, HEN^{+}01, HG03, HBG^{+}02, IBS^{+}02, IFP^{+}03, JFS^{+}03, JHS^{+}02, KMH^{+}04, KNI^{+}04, KI04, KKW^{+}03, KSK^{+}02, LM01, LeB02-48, LeB03-53, LRF^{+}02a, LRA^{+}02, LDI^{+}03, LPT03, MTT^{+}04, MKD^{+}01, MHK04, MTB^{+}02, MRC^{+}02, MNC^{+}03, MWC^{+}02, MJY^{+}04, NPS^{+}03, OMB^{+}01, PWY^{+}03, PLW^{+}04, PPA^{+}03, PLC^{+}02, RDC^{+}04, RBW^{+}02, SRG^{+}04, SSOS01, SF01, SKO04, TYA^{+}02, TPA^{+}03, VZTN03, WKYC02, WVBY^{+}03, WXD^{+}03, WYR^{+}02, WNM^{+}03, WKZ^{+}02, Wel03-30, WZB^{+}01a, WAV^{+}04, WLO^{+}02, WD01, WE02, XLGS01, YpHRL03, YFW^{+}04, ZSS01, ZLVS02, dSAH02]. Kinase [AFN00, APLB00, BMD^{+}00, BRS^{+}01, CDEM00, CAP00, DSSWW00, FCL^{+}00, FMMF01, GDRS01, GLK^{+}00, GG01, GTPMU00, HCK^{+}00, HBSQ01, HSW00, HLZW00, JWJJ00, KSN^{+}01, KTY^{+}00, KMiM^{+}01, KMB^{+}01, KNIO01, KHH^{+}01, LTB^{+}00, MMG^{+}01, MTV^{+}00, MLS^{+}01, PHWK^{+}00, PVjL^{+}00, SJIM01, SCM^{+}00, SBC01, SMTC00, SYH^{+}01, TP01, THZ^{+}01, UAZG00, WBC^{+}00, WRGK00, YSS^{+}01, YF00, ZGB01, ASS02, LRWB04, OGD03, PBD^{+}04]. Kinase-3 [HCK^{+}00]. Kinase-Anchoring [SCLC00]. kinase/AKT [BFC^{+}02]. kinase/c [PBD^{+}04]. kinase1 [BBDM02]. kinase1/2 [BBDM02]. Kinases [LeB04-52, CFB^{+}03, CROfC04, HIN^{+}03, ISID^{+}03, KMS^{+}04, MP04, OEM^{+}02, SHVR02, VBR^{+}01a, WBAS04, WSR03, dVKS04, LCGR00, MMC00, PS00, SGPL^{+}00]. Kinesin [DSV^{+}00, Hol01, RH00, TYS^{+}00, TB00a, ABCK^{+}03, CiKBG03, GC04, KKK^{+}02, KGT^{+}02, LDC^{+}04, LeB02-42, MPB^{+}04, MHNSM03, NH03a, NPS^{+}03, Tum04k, XRH^{+}03, BHM^{+}00, KMCM00, RCL^{+}00, TUV00, TV00, VMD^{+}01, YMM^{+}00]. kinesin-1 [KKK^{+}02]. kinesin-like [KGT^{+}02, NPS^{+}03, BHM^{+}00]. Kinesin-Mediated [DSV^{+}00]. Kinetic [XWL03, LGM^{+}04]. kinetically [HOK^{+}02]. Kinetics [KC01, DHK^{+}04]. Kinetochore [CENMR^{+}01, MRK04, CBW^{+}01, CDB02, Che02, DMH^{+}02, Dov02-65, HCL^{+}03, HMC^{+}01, KCG^{+}03, LeB03-66, LeB04-82, LdCK^{+}01, MSL^{+}02, MB03c, SERB03, TDFV02, Wel01-27, WCA^{+}03, AMEC01, CHM^{+}01, HMAM01, KHN00, ODR^{+}01, SBC01, SMLM00, YD00, YMM^{+}00]. Kinetochore-driven [MRK04]. kinetochores [BRB^{+}01, DJT^{+}03, Dov03-59, JTK^{+}02, KBWG02, MSC^{+}03, TCH^{+}02, TBRG01, HHF^{+}00, SMS00]. Kinetoplast [AERD^{+}01, DE01]. KinI [GC04, MHNSM03]. Kir [PFW^{+}00]. Kir2.1 [FLLE^{+}01]. Kiss [Les01g]. KLP [PRS^{+}04]. KLP-19 [PRS^{+}04]. Knock [Dov01t]. Knockdown [MMBB03]. Knocked [LeB03-54]. Knockin [LAF^{+}00]. knockout [TTH^{+}01, TBJ^{+}01, Wel01-40, NHI^{+}00]. KNOLLE [HGS^{+}01, AHMJ01]. KNOLLE-interacting [HGS^{+}01]. Knot [PSB00]. Knowing [Dov02-36]. knows [LeB03r, LeB03-53]. Krox [Dov04i, PBD^{+}04]. Krox-20 [Dov04i, PBD^{+}04]. Krp [RCL^{+}00]. Kss1 [KMS^{+}04]. Ku86 [EKdM^{+}04]. Kuzbanian [YSW02].

L [DKA01, DSM^{+}03, JJM^{+}02, Tum03a, YWW^{+}04]. L-selectin [DKA01, DSM^{+}03, YWW^{+}04]. L-type [JJM^{+}02]. L1 [DBL^{+}02, DMBS02, ICK^{+}04, JB01, LeB02-49, MGAL^{+}01, NYT^{+}03, PPR^{+}00, SLB^{+}01, SKK^{+}02, SMSM00]. L1-CAM [NYT^{+}03]. L1-Containing [PPR^{+}00]. L1-dependent [NYT^{+}03]. L1CAM [COB01, GSB^{+}03, LeB03-89]. L23 [UHR^{+}03]. labeled [NSLSK02]. Lack [AQHO03, HGK^{+}01]. lacking [CBG^{+}01, EKdM^{+}04, GBN^{+}01, IKS^{+}01, ICK^{+}04, KHB^{+}04, NOM^{+}04, TPA^{+}03, XJW^{+}04, MGLPM00, WS00]. Lactotrophs [CAB00]. LAD [COB01]. LAD-1 [COB01]. laevis [IOLA^{+}00, MPB^{+}04, TMHP00, VAHV00]. laid [LeB04-55]. Lake [Wel02-51]. Lamas [Lai03b]. lamella [CWG^{+}02, ECK^{+}03, RWSV03]. Lamellae [ONM00]. Lamellipod [ZBB^{+}00, LeB04-87, CBZ^{+}00]. lamellipodia [NOOG^{+}04, OKSH00, WKS^{+}00]. lamellipodial [DLY^{+}02, GHK^{+}03, MWMK04, RS00a]. Lamin [KMP02b, Dov03w, MHH^{+}03, SGW^{+}02, MSHG00, SGG01]. Lamina [SMTC00]. laminar [ITM^{+}04]. Laminin [CS03a, SFH04, BJM^{+}02, CRS^{+}03, KVM03, LHC^{+}02, SHB^{+}03, SKH03, TAA04, WFI^{+}04, Wel01-64, ZLR^{+}03, Dov03-30, GPAS^{+}01, KMH^{+}00, KGC^{+}00, MWF02, SEP^{+}01]. laminin- [CRS^{+}03]. Laminin-1 [MWF02]. laminin-2 [BJM^{+}02, Dov03-30]. laminin-5 [SHB^{+}03, SKH03, WFI^{+}04, ZLR^{+}03, KGC^{+}00]. Laminins [SEP^{+}01]. Lamins [Bur01, LSMS^{+}01, MYKG00, SC01b]. Lamp [SFV^{+}00]. landmark [SCPP02]. landmarks [OKC02]. LAP2 [MEFC03]. Large [MWHM01, MBN^{+}01, MKTW01, NSLSK02, SWB03, XRI^{+}04, YHZ^{+}01, HKJ00, WMT^{+}01]. Large-scale [MWHM01, SWB03, YHZ^{+}01]. Largely [CW01]. Lasp [LPL^{+}04]. Lasp-1 [LPL^{+}04]. LAT [TNM^{+}03, WPS^{+}01, HK00]. Latching [Wel03-31]. Late [WFT^{+}01, AMG^{+}01, CSM03, HBG^{+}02, PMKM03, SV03, SKM03, YGWN01, BP00, PMB^{+}00, RRB^{+}01, SG00]. Latent [HKO03, ACMR04, DCC^{+}02, MSM^{+}04a]. later [Dov03-64, Wel03o]. lateral [DK04, TSMT00]. Lats1 [HMN^{+}00]. lattice [MHNSM03]. launches [Tum03d]. Lava [SFV^{+}00]. lay [LeB04-37]. Lc8 [YDRS01]. LCMV [KSM^{+}01]. Ld [NT00]. LD4 [WZB^{+}01a]. LDH [LeB02-109]. lead [DSG04, LeB03n, NDM^{+}03, IAG^{+}00]. leader [LeB02g]. Leading [ZBB^{+}00, EWD02, ECK^{+}03, Mil02, MKR01, SCTF04, UGKT^{+}02, WBWS03, qZC01]. Leads [DGB^{+}00, RRM^{+}00, FSCF^{+}03, HJL^{+}04, JZ02, LKLD04, SKM^{+}02, AKK^{+}00, CDFT^{+}01, HGK^{+}01, ZEtK^{+}00]. Leaf [Les01h]. leak [Dov03-63]. leakage [MBH^{+}02, HDL^{+}00]. Leaky [FFST01]. Learning [LeB02-50, Dov03i, FHL^{+}03]. least [TOM01, NM00]. leave [LeB03-32]. leaves [LeB03b, Gil01]. leaving [MBN^{+}01]. lectin [CdLvM^{+}04]. LEF1 [NH03b]. left [BVH04, LeB04-53, Wel04-75]. leg [Dov01u, Wel02-31]. legislation [Hel03]. Leishmania [Wel01-47]. length [GK04, SGdM^{+}01, SCB02, TDL03, AKW00, TUV00]. lens [LM01, MWSL^{+}03, MKR01]. Lesion [HGB^{+}00]. Less [Dov02-37, Tum04k, RT02, Sre04n, Wel04-76]. Lessons [KdVS^{+}00]. Let [Wel02-34, Dov02-36, LeB02-52, RO04, Sod02, Wel03m, Wel01-31, Wel02-35]. lethal [ALP^{+}04, HvdHG^{+}03, JWJJ00]. lethality [FSCF^{+}03, ZWB04, NHI^{+}00]. Lethargic [LeB04-53]. lets [Dov03d]. Letter [Mel00b]. Leucine [KRA^{+}01, KPKY^{+}03, BMF00]. leucine-rich [KPKY^{+}03]. leukemia [TIS^{+}01, WSS^{+}04, BHBJ00, BKB^{+}01, HCD^{+}00, MRM^{+}00]. leukemia-associated [TIS^{+}01]. Leukocyte [KRA^{+}01, CS04, DKA01, SM03a, LSA^{+}00]. leukocytes [BYMS^{+}02]. level [TBRG01, WTG01, ZLMP00]. levels [XAB^{+}03]. lever [KRMB03, SCB02]. Leydig [GBM^{+}00, DME^{+}04]. LF3 [TDL03]. liberated [SHB^{+}03]. liberates [BTH^{+}03]. liberation [CP03]. Library [RS00c]. licence [Wel01-61]. License [Wel04-41]. licensing [OSB04]. lid [Dov03m, Pow01g]. lies [Wel02-67]. Life [Lai03c, Wel03-32, LeB02c, LeB03q, Krä00]. lifespan [MCG^{+}03]. Lifetimes [KSN^{+}01]. Ligand [EKB^{+}03, GTR^{+}03, HIG^{+}01, Jeg01b, KVC^{+}03, LCM02, SJW^{+}04, SHW01, SMR^{+}02, ZTK^{+}03, BMS^{+}00b, CSP^{+}00, DLR^{+}01]. ligand-independent [SHW01]. ligand-induced [LCM02]. ligands [SWK^{+}04, SV03, STJ^{+}01, BGS00, THE^{+}00]. ligase [DKAH04, KCWF02, YRC^{+}04, HLK01]. ligases [CSG01]. ligates [ZLR^{+}03]. ligation [BHK^{+}02]. Light [DSV^{+}00, DDW^{+}01, CWG^{+}02, DV04, Dov02-52, HZS^{+}01, LeB02-60, Les01c, LVD^{+}04, Pow01i, Wel02-41, Wel04-72, PPM^{+}00, PVjL^{+}00, RBB00, YDRS01]. lights [Tum04c]. like [ACBG04, BMM^{+}02, BLC00, BBP^{+}04, BHM^{+}00, DSSWW00, EGS^{+}04, GFM^{+}04b, HNK^{+}03, KKR^{+}00, KK02, KGT^{+}02, LR00, LLH^{+}01, LFM^{+}04, MCBB^{+}04, MZH^{+}02, MR01b, NPS^{+}03, PGS^{+}01, PSE^{+}03, PQF^{+}00, RWCC01, RS00b, SHB^{+}03, SMSM00, SRD^{+}02, STJ^{+}01, TIS^{+}01, VYW^{+}03]. LIM [AKM^{+}04, BGR^{+}01, BKD^{+}04, GCT02, MSJ^{+}02, ZCH^{+}01, EHZ^{+}01, SME^{+}00]. LIM-only [MSJ^{+}02]. limb [ASGL^{+}01, LeB04-86, IHK^{+}00]. limelight [Moo02]. limited [LeB03-41]. Limiting [SDS00]. limits [KH04, Wel01b]. LIMK1 [FLS^{+}03]. LIMP [JKB^{+}03]. Limulus [SMW^{+}03]. LIN [LHvdH00]. LIN-5 [LHvdH00]. line [DFJ^{+}02, LeB03-50, Wel04-46, HSW00, OLB^{+}00, Zet01, MKS^{+}02, ZCH^{+}01]. Lineage [HBK^{+}02, WWK^{+}00, GNS^{+}04]. Lineage-specific [HBK^{+}02]. lines [Tum04t]. ling [Las03]. lining [SAS^{+}02]. Link [KSK^{+}00, BBG^{+}03, DSH^{+}03, Pow01e, Tum03b, Wel01z, Wel02-61, MLC^{+}01, OSN^{+}00, RS00c, RPE00, ZGB01]. linkages [SF01, YK03b, vWJK^{+}03]. Linked [CENMR^{+}01, BFSO^{+}04, CRS^{+}03, Dov02-57, PPP^{+}01, TPA^{+}03, WD01, HLZW00, MMWC00, THZ^{+}01, VDMH01, YSS^{+}01, ZGB01]. linker [HDH02, KAiK^{+}02, PPGN^{+}02, WZB^{+}01a]. Linking [Dov03-31, Wel03-33, CH03, ESC^{+}01, MSA^{+}03, MWMK04, NKM^{+}03, RML^{+}02, SPK^{+}01, TCV^{+}00, Wel01t, MDJF00, TV00]. Links [JRW^{+}01, LZS^{+}03, OKC02, Wel04d, KEGDQ01, QKK00]. lipase [SXD^{+}03]. lipases [BHW^{+}03]. Lipid [Lai03d, WSC^{+}01, ALC^{+}03, BND^{+}02, EKH^{+}03, GPDvH^{+}03, KH04, KvHB^{+}01, NDS^{+}02, ORZ^{+}04, SBG^{+}04, SNS^{+}04, TNM^{+}03, VBR^{+}01b, EU00, FKI^{+}01, KNM^{+}00, NKP^{+}01, OPZ^{+}01, PLL^{+}01, vM01]. Lipids [LeB02-51, Dov02u, FJM^{+}04, RMR02]. Lipolysis [Bir03]. lipolytic [SXD^{+}03]. lipolytica [TCR00, TNW^{+}02, GKN^{+}03]. Lipoprotein [CSJ00, MTW^{+}02, OPP^{+}03]. LIS1 [DBS^{+}03, TDFV02, Xia03, EM00, LYL^{+}04, TSH^{+}04]. Lis1-Related [EM00]. lissencephaly [LOC03]. listen [LeB02-109]. Listeria [BGR^{+}01, BCM04, GGT^{+}02, PAG02, SPW00, SAG^{+}01]. Listeria-induced [BGR^{+}01]. Listeriolysin [DC02b]. Little [LeB02-52]. Live [HGP^{+}04, MSH^{+}00, DBH^{+}01, Dov01v, LGM^{+}04, NGKH02, SP03, MSM^{+}01, PGSL00]. liver [SwZK^{+}02, Wel02-35]. lives [Pow01j]. Living [SGYD^{+}01, BBP02, CLB^{+}02, KSC02, LRA^{+}02, MVL^{+}02, OTB03, OGD03, SCB02, dVKS04, HHF^{+}00, KC01, LRW^{+}00, SWJ^{+}00, YYM^{+}01]. LMP1 [OBG^{+}03]. loaded [Tum04-28]. loading [DSB^{+}02, GBZ^{+}02, GKYY03, PFM^{+}00]. Local [GWL03, KYBS03, RRJ^{+}01, SHS^{+}00, SH02]. Localization [GMRS00, LDS^{+}00, POH^{+}04, PVjL^{+}00, RTM^{+}01, WKS^{+}00, YDRS01, AMF01, BN02, BCB^{+}02, CW02, CMC^{+}02, Che02, CBC^{+}01, FHM^{+}03, GMRYM^{+}02, GBOL03, GNDLS^{+}01, GPZ^{+}02, HLB^{+}02, JBK04, KYM04, KJB^{+}02, LKM^{+}04, Las03, LeB02-90, LJK^{+}01, LL02, MPB02, POW^{+}01a, PLP02, RJA^{+}03, SCPP02, SKS^{+}04, STA^{+}01, SSOS01, TE01, TM04, UJL^{+}03, VMK^{+}03, WKZ^{+}02, vEPP^{+}01, AERD^{+}01, BMS^{+}00b, BKI^{+}01, BT00, DOL^{+}01, DAC00, HOvD^{+}00, KHvOD00, KKL^{+}01, KKK^{+}00, LGM^{+}01, LYMC00, MSM^{+}01, RSG01, RMG^{+}00, SBC01, WMHB^{+}00]. localizations [SP03]. localize [BPD^{+}04, GMD^{+}02, Wel04r, ZLH^{+}03, GBD^{+}00]. Localized [ACP^{+}02, BTD^{+}00, SCTF04, TSY^{+}02, BDKM04, CRP^{+}04, FGS^{+}02, BAD^{+}00, FKG00, FWY01, MMR^{+}00, PO00, SSN01, UOB^{+}02]. localizes [MTM^{+}02, RXS^{+}03, NHS00, SLT^{+}01]. localizing [AQC04, GKM^{+}01]. locally [Dov04e, MSMK04]. located [BRR^{+}02, CGBL^{+}02]. location [TKS^{+}02, DE01]. lock [Wel04-54]. locks [Tum03h]. Locomoting [RRJ^{+}01]. locus [LZPL01, LCI^{+}01, SMR^{+}02, TSL04]. Lon [LeB02-53]. Lonely [LeB03-55]. Long [Dov01v, GCR^{+}03, PLW^{+}04, TTR00, Dov03-32, Wel01-46, ECV^{+}00, IHK^{+}00, TWS^{+}00b]. Long-Term [TTR00, PLW^{+}04]. longitudinally [ABOS^{+}02]. look [Dov02-67, Dov03i, LeB04n, Mel01]. loop [MWE^{+}03, OWW02]. loops [Lai03b, NW04]. Loopy [LeB04-54]. loose [LeB04e]. loosening [NHG^{+}03]. Loricrin [KdVS^{+}00, SJA^{+}00]. Losing [LeB03-56]. Loss [ASK^{+}03, LSW^{+}03, MBM^{+}04, PSS^{+}04, PRS^{+}04, RRM^{+}00, Tum03d, WC03, ARLC^{+}04, LeB02p, RGG03, WBP^{+}03, Wel03-34, ZWSC02, GBM^{+}00, PDJ00, qZC01]. Lost [RML^{+}02]. love [Tum03n, Wel01-30, Wel03-56]. loves [LeB04-62]. Low [CSJ00, OPP^{+}03, Cam03, DBH^{+}01, GTPG03, Lai03c, MTW^{+}02, SHF^{+}03, Wel01-53]. Low-Density [CSJ00]. Lrp5 [KPL^{+}02]. Lst1p [SKR^{+}00]. LST8 [CK03]. Ltbp [DCC^{+}02, HKO03, KWSK^{+}04, DCC^{+}02]. LTBP-2 [HKO03]. Ltbp-3 [DCC^{+}02]. LTBP-4 [KWSK^{+}04]. LTD [FHL^{+}03]. Lte1 [YITe03]. lubrication [LeB02e]. lumen [CBRBM04, MW04, MBN^{+}01, GSB^{+}00]. Lumens [LeB04-55]. Lumican [ECO^{+}00]. Luminal [HRB^{+}01, KKL^{+}01, LAF^{+}00, NBWB^{+}00]. lung [KWSK^{+}04, RL03]. lupus [Les01i]. Luysian [WND^{+}00]. ly [LeB04-84]. Lymphatic [PS03]. Lymphocyte [BPMG00, LKH^{+}04, XRI^{+}04]. Lymphocytes [HWHH01, SBM^{+}01, Zwe00]. Lymphoid [ESS^{+}00, PTM^{+}01, SKF^{+}01, HSC01]. lymphomagenesis [SSM^{+}04]. Lyn [KNI^{+}04]. Lysine [HSW00]. lysis [PBB^{+}04]. lysophosphatidic [UGKT^{+}02]. lysophospholipase [UGKT^{+}02]. Lysophospholipids [Moo02]. Lysosomal [WFT^{+}01, POH^{+}04, MGL^{+}00]. Lysosome [HLB^{+}00, CHA^{+}01, PMB^{+}00]. Lysosomes [And02, RTM^{+}01, TRC^{+}00, Dov01g, JAS02, SKM^{+}02, MCH^{+}00, PMB^{+}00]. Lysyl [HAM01].

M [GLS^{+}03, SMW^{+}03, VYC^{+}00, CS01, CAW^{+}04, HSW00, KCWF02, MKS^{+}02, MSR04, PMBC^{+}00, ASK^{+}03]. M-line [MKS^{+}02]. M2 [HGP^{+}00]. machine [STA^{+}01]. machineries [TOM01]. Machinery [CQH^{+}00, SGYD^{+}01, FS03, KSBE03, LHK02, SGK02a, SKGC^{+}03, SCK04, dMMBK^{+}02, vdBCH^{+}04, PLH^{+}01, VAHV00]. machines [MLZ^{+}01, NG01]. macroH2A [CW02]. Macroh2a1 [RME^{+}00]. Macrophage [ECK^{+}03, NLRD01, Dov01r, HPFG03]. macrophages [CSL^{+}03, HHJS04, NCGD^{+}03, PMKM03, TFF03, WLKS01, CMW^{+}01]. macropinocytosis [HdO^{+}01, MBH^{+}02]. Mad [LeB02f, LeB02-95, LeB04-84]. Mad-ly [LeB04-84]. Mad-ness [LeB02-95]. Mad1 [Che02, SBC01, TEC^{+}03]. MAD2 [SCY01, DJT^{+}03, HHF^{+}00, LeB04-101, SBC01, TEC^{+}03]. Mad2p [HJSM00]. MAD3 [HJSM00]. made [LeB03-102]. Madin [HGP^{+}00, AFK^{+}03, FFST01, MBMMA^{+}03, VHLS00]. Madrid [MR02]. Mads [LeB02-54]. Magi [SCL^{+}01]. Magi-1c [SCL^{+}01]. Maguk [RML^{+}02]. mainly [LYL^{+}04]. maintain [MKJ^{+}02, PRJK01, RSD^{+}04, SDML04, iNFK^{+}01]. maintained [WSC^{+}03]. Maintaining [RB03, FKH^{+}04, HCL^{+}03, CNJ01, YCX^{+}01]. Maintains [WGvA^{+}01, AOH^{+}02, BJM^{+}02]. Maintenance [WPC^{+}01, HG03, HPS^{+}04, KJY04, KHLW02, MSL^{+}02, PBD^{+}02, SWBE^{+}04, SLG02, SHVR02, SAH^{+}03, YHB^{+}04, GH00, WWG^{+}00, YMM^{+}00]. maize [CC02, YD00]. major [JAS02, KNK^{+}01, CHM^{+}01, DLR^{+}01, MGL^{+}00]. Majority [BHY^{+}00]. Make [Wel02-36, Bir04d, Lai03d, LeB02-78, LeB03z, LeB03-77, LeB04d, Sre04n, Tum03j, Wel01-31, Wel02-30, Wel02-35, Wel03-57, Wel04-40, Wel04-39]. makes [Dov02-68, Dov03-35, LeB02p, LeB03k, Wel01-65, Wel02-65, Wel04-53, Wel04-77]. Making [Dov02-38, Dov02-39, Tum04l, Wel01-32, Wel01-33, Wel03-35, Wel04-42, LeB02-72]. MAL [MBMMA^{+}03, SWBE^{+}04, dMMBK^{+}02]. MAL2 [dMMBK^{+}02]. Malaria [Cam03, LeB04-56, KSD04, KBGG04b]. male [BGBG03, VLL^{+}03, JWJJ00]. Malformations [UIK^{+}01]. maligned [Les02a]. Mammalian [BHW02c, CBRBM04, GGGK03, IWG^{+}01, KEGDQ01, SZZ^{+}00, TT04, AHA^{+}04, BRB^{+}01, CKZ^{+}02, DBH^{+}01, FHJW^{+}01, FHF^{+}02, KBWG02, KI04, KGT^{+}02, LCSG03, LRB^{+}03, MP04, OER^{+}02, PPP^{+}01, RML^{+}02, RDNB02, SRW^{+}02, SRB^{+}04, TCH^{+}02, TUK03, UOB^{+}02, VFN^{+}04, WBP^{+}03, YGWN01, CHM^{+}01, KCL^{+}00, PHWK^{+}00, PKF^{+}00, SMS^{+}01b, TWS^{+}00b]. Mammary [HRB^{+}01, BBR04, DWB03, DFJ^{+}02, HSC01, JZ02, MWL01, MSS^{+}01, MMG^{+}04, SHB^{+}03, SMS^{+}01a, WVBY^{+}03, WSL^{+}03, WDL^{+}04, ZLR^{+}03, ASMW01, IEJ^{+}01, MLS^{+}01]. Man [Wel04-43]. Managing [Wel03-36]. manipulation [Ros02, RY04]. Manner [WWK^{+}00, DWB03, NK02, SHW01, SC04, GWG01, PGS^{+}01]. mannose [AHA^{+}04, GGGK03, TST^{+}03]. mannose- [TST^{+}03]. many [RMR02, Tum04j]. map [Dov01-27, DSB^{+}01, HG03, KMS^{+}04, LeB02-55, LeB04-57, MTB^{+}02, AFN00, GLK^{+}00, WWKV00]. MAP-muddled [LeB04-57]. map1b [TTHH00, TTH^{+}01, MHW^{+}00]. MAP2 [ABOS^{+}02, HTT^{+}02, TTH^{+}01]. MAPK [LeB04-70, LTD^{+}02, MMM^{+}04, MSR04, SRSW04, SDEZ^{+}03, SSRX04, VCDHD03, CNHK02]. MAPK-interacting [LTD^{+}02]. MAPK-NF- [SRSW04]. MAPK/ERK [VCDHD03]. Mapping [LeB02-55, WPS^{+}01]. Marching [BHIWH01]. MARCKS [IUK04, LFT^{+}00]. Marie [RKF^{+}04]. mark [LeB03b, LeB04-57, MTT^{+}04, Gil01]. MARK/ [MTT^{+}04]. marked [BHPN04, PPK^{+}01]. marker [CPA^{+}03, DLT^{+}02, SISO01, STA03, Wel01-34]. Markers [NKP^{+}01]. marks [EOJ^{+}03]. marrow [PPK^{+}01]. Mass [LeB04-58, TOM00, MSJ^{+}04, RCY^{+}03, CKS^{+}00]. Massive [LMW^{+}00, HV03, MH01]. Mast [HGC00, ISS^{+}04, WPS^{+}01, MSL^{+}02, MF01a, WPO00]. MAST/ [MSL^{+}02]. mat [GSW^{+}00]. Matching [DCM00]. maternal [GJS^{+}03b]. Math [Dov03-33, Tum03i]. Mathematical [BLU^{+}04, CNT03]. Mating [BVH04, Dov04o, HW00a]. matrices [CACL03]. Matrilysin [LBWH^{+}00]. Matriptase [LSW^{+}03]. Matriptase/MT [LSW^{+}03]. Matriptase/MT-SP1 [LSW^{+}03]. Matrix [AIH^{+}00, AV01, CK00, ECV^{+}00, Jeg01c, LHC^{+}02, LGW00, WND^{+}00, BPKK01, BSL^{+}01, CSO^{+}04, DSB^{+}02, ESC^{+}01, JBA^{+}01, MSMK04, NMH^{+}04, NGS^{+}01, PSD^{+}04b, SOH^{+}04, SRS01, SRC^{+}01, SMS^{+}01a, Sre04n, UIY^{+}01, WKYC02, Wel01-48, WDS^{+}03, ZTK^{+}03, GPAS^{+}01, HAP^{+}00, IOLA^{+}00, KIC^{+}01, PCK^{+}00, Qua00, RMG^{+}00, RRL^{+}00, WWJ^{+}00]. matter [Wel02t]. Maturation [PMBC^{+}00, FBG^{+}01, HKHO01, HHJS04, LeB03-99, LTF^{+}01, MH02, SM04b, VBR^{+}01a, PDJ00]. maturation-promoting [MH02]. Maturation/M [PMBC^{+}00]. mature [GBN^{+}01, BHKL01, GMZ^{+}00]. maximal [KRF^{+}03, MWE^{+}03]. may [MKS^{+}02, SGO^{+}00]. Mba1 [PLH^{+}01]. Mbl [LeB03-87]. MCAK [GC04, OWW02]. Mcd1p [LGK00]. MCDE [LCT^{+}04]. Mcf [WTG01]. Mcf-7 [WTG01]. MDCK [ATF^{+}04, PPA^{+}03, KTY^{+}00]. mDia [ABF^{+}03]. mDia1 [TIO^{+}02]. Mdm33 [MJV^{+}03]. Mdv1p [TN00a, TONN02]. mdx [BMM^{+}02, TTP^{+}01, VMS^{+}02, WST01]. Me [Wel04-44, LeB02-52, LeB03-32, Sod02, Wel02o, Wel02a, Wel02-51]. meals [Dov04f]. Meanwhile [Dov02-40]. measure [Tum04v]. measured [SV03]. measurement [LMVW03]. measurements [CTE^{+}04]. Measuring [Tum04m]. Mechanical [LRBH02, BWV^{+}01, BMY^{+}01, DKA01, vWJK^{+}03, HSB00]. Mechanically [RPE00]. Mechanism [HM00, KWOP00, PVL^{+}00, RyHHK00, SBC^{+}03, BLU^{+}04, CSP^{+}04, CKS02, FMF^{+}04, FGR^{+}04, GV04, HCC02b, IKA^{+}03, KB04, LJK^{+}01, LWZ03, MBN^{+}01, SHM02, SDS^{+}04, VN04, Wel04n, WW02, WHD^{+}03, WML^{+}03, WG03, AFN00, AOC01, CHM^{+}01, LFT^{+}00, Rut00, TP01, TMD^{+}01]. Mechanisms [ITF^{+}02, CFB^{+}03, MTPT02, TKHR03, VKB^{+}01, YPN^{+}04, KMCM00, NM00, TMK^{+}00, WSDW^{+}00, WGF^{+}00]. Mechanosensors [RZB^{+}01]. mechanosignal [OMiKF02]. Mediate [pHYpXL00, CFWH^{+}01, DRG^{+}03, GSB^{+}01, IBS^{+}02, IYT^{+}04, JCPWS01, KMG^{+}03, KMP02b, LTD^{+}01, MEFC03, MRC^{+}02, TSH^{+}04, WFI^{+}04, ZLR^{+}03, GV00]. Mediated [AKDS00, ASAJ01, BYLA^{+}01, BHL^{+}01, CGY^{+}01b, CSJ00, DSV^{+}00, ELO^{+}01, FWP^{+}00, GMRS00, GWBW00, GDHS01, KFS^{+}00, KKA^{+}01, KY00, KRA^{+}01, LHW^{+}01, MBS^{+}01, MMS00, NLRD01, NZHR01, PTM^{+}01, QK00, RM01, RGGL00, SHN^{+}01, SDS00, SPW00, SEM^{+}00, SSL^{+}00, vARP^{+}00, ALC^{+}03, ACMR04, BS04a, BBP02, BFH^{+}01, BDR^{+}03, BCM04, CTE^{+}04, CS02, CS03b, DC02a, EPH^{+}02, EGC^{+}03, FRT^{+}01, FM01, GLDM01, GBY^{+}03, HKP^{+}04, HBB^{+}02, HdO^{+}01, HFK^{+}03, HHOP02, JHS^{+}02, KHLW02, KSR^{+}04, LMGM^{+}02, LLP^{+}02, LUB^{+}02, LM01, LSMS03, MWAM01, MBMMA^{+}03, MBH^{+}02, MZH^{+}02, MBSR03, MSN^{+}02, MWC^{+}02, NG02, PBB^{+}04, PMP^{+}03, PPA^{+}03, REK^{+}03, RAD^{+}02, RGG03, RBD^{+}01, RWH02, RXS^{+}03, RDS02, SWH^{+}02, SVI^{+}04, SRL^{+}04, SBG^{+}04, SHA^{+}03, SMZ^{+}03, SAG^{+}01, SHKS02, SCY01, SKO04, TJS^{+}04, TSL^{+}03, Tum04u, TEC^{+}03, UJK^{+}02, VBH^{+}02, WK02, WLW^{+}04, WSWM04b, WCBC04, WCTU02]. mediated [WSR03, WZB^{+}01b, Xia03, XSK^{+}01, YLQ^{+}02, YpHRL03, ZLVS02, AS00, CAP00, HCTM00, HSW00, KKL^{+}01, MJG^{+}01, NM00, PLL^{+}00, WYHP00]. Mediates [HRB^{+}01, YB01, BM02, CLM^{+}03, EOB^{+}04, GLS^{+}03, GSB^{+}03, HV03, ISID^{+}03, KEHAM^{+}02, MWAM01, MTW^{+}04, MCF^{+}02, NYT^{+}03, RCS^{+}02, SMS^{+}04a, SRG^{+}04, SJ01b, WWD^{+}04, WWS^{+}03, YSK^{+}04, YSW02, APLB00, EHCC^{+}00, FCL^{+}00, GLK^{+}00, KIK^{+}00b, KH01, PWC^{+}01, RRK^{+}00, SGF^{+}00]. mediating [CAW^{+}04, KWH^{+}00, MAAZ^{+}00]. Mediation [MZ00, YCK^{+}03]. mediator [BSD^{+}01, CGM^{+}02, FBV^{+}04, SMT^{+}03, YTM03]. mediators [CPG^{+}03, LCT^{+}04]. Medium [vRTvdB^{+}00]. Meet [DA01]. meets [Wel02-45, Cha00, LeB04-107]. megakaryocyte [CSJ^{+}03b]. Megakaryocytic [KMiM^{+}01]. meiosis [BGBG03, KLT^{+}03, RBW^{+}02, TBW^{+}04, VLL^{+}03, PN00, TSDS00]. Meiotic [EOJ^{+}03, LTD^{+}02, TSDS00, YK03b, CC02, CSM04, FCLSN03, WHB04, YD00]. Mek [AFN00, KTY^{+}00, DOL^{+}01, RGGL00]. Mek/ERK [RGGL00]. Mek1 [CDEM00, BS04a, CSM03, SDEZ^{+}03]. MEK1-dependent [BS04a]. Mekk1 [MMG^{+}01]. Melanocytes [HCR^{+}01]. Melanocytic [RTM^{+}01]. melanogaster [BC00, AW00, CQH^{+}00, DRC^{+}02, HKK^{+}00, IdCAS^{+}00, Las00, SFV^{+}00, SKM03]. Melanoma [YPN^{+}04, BS04a, FCF^{+}01, WHP^{+}02]. melanophores [GTD^{+}02]. melanosomal [SDC^{+}01]. melanosome [BTH^{+}03]. Melanosomes [RTM^{+}01, HCR^{+}01]. melt [LeB02v]. melt-down [LeB02v]. member [LeB02-62, vBDH03, KMH^{+}00, TWBV^{+}01, WDLK01]. members [DK04, DMBS02, GSB^{+}01, GH00]. Membrane [ALP^{+}04, HGP^{+}00, HSMJ01, IWG^{+}01, JAS02, KIC^{+}01, KSF^{+}00, KHK01, KIO^{+}00, LeB04-59, Mel00c, MMS00, NHB00, RH04, SDN^{+}00, WGvA^{+}01, ARLC^{+}04, AM03, And02, ATF^{+}04, BHW^{+}02a, BRB^{+}03, BGA^{+}04, BWA^{+}04, BRR^{+}02, BSR^{+}03, BRY^{+}01, CLM^{+}04, CKF^{+}03, CSG01, CLSK02, DK04, DDV^{+}03, DRP^{+}03, DAV^{+}03, DKAH04, FMJG04, FPM^{+}03, FBH03, FJM^{+}04, FRM^{+}02, FGSW03, GGD^{+}04, GGB^{+}04, GSN^{+}04, HSMB02, HBV^{+}01, HPG^{+}02, IBS^{+}02, IYT^{+}04, JB01, JMG04b, KLE^{+}02, KSS^{+}03, KSD04, KVM03, LYL^{+}04, MH01, MSGS02, MSJ^{+}04, MWK^{+}02, MN03, MWM^{+}02, MJV^{+}03, MPV^{+}01, MSN^{+}02, NSK04, NCGD^{+}03, OSNG04, OGD03, PWY^{+}03, POH^{+}04, PUK02, PWS^{+}01, RCS^{+}02, SOH^{+}04, SVI^{+}04, SPK^{+}01, SJ01b, SHKS02, SZvBuH^{+}04b, SML^{+}04, TE01, TWS^{+}04, TR02, Tum03f, Tum04-27, UJK^{+}02, UIY^{+}01, UML^{+}03, UN03, VCGB^{+}02, VTGT^{+}03, WAOC^{+}03, WSL^{+}01, YGWN01, YHB^{+}04, vdLBK^{+}04]. Membrane [ATE^{+}01, BD00, BRG^{+}00, CAB00, DM00, EU00, FSY00, FKI^{+}01, Gau00, GLA00, GZY^{+}00, GBON00, HF01, HK00, HW00a, HAP^{+}00, JCR^{+}01, JLS^{+}01, JMG01, KIK^{+}00a, KRR^{+}01, LVWA01, LW00, LCS^{+}01, Lit00, MAAZ^{+}00, MMH^{+}00, MSH^{+}00, NS00, NSW00, NKH^{+}00, PLL^{+}00, PPM^{+}00, PKF^{+}00, RS00a, RSG01, RMG^{+}00, SJS^{+}00, SSN01, SD00, SJ01a, SKR^{+}00, SKT^{+}00, SKCS00, TGM^{+}01, TR00, TSK^{+}00, VHLS00, WAPB^{+}00, WPO00, YHF^{+}01, ZGB01]. Membrane-Anchored [IWG^{+}01, SOH^{+}04]. Membrane-Binding [KIO^{+}00]. membrane-bound [CSG01, HSMB02]. membrane-proximal [MWM^{+}02]. Membrane-Type [KIC^{+}01, UIY^{+}01]. membranes [BJ03, DW02, FRT^{+}01, HSMB02, KKW^{+}03, LMHJ02, LRD^{+}03, LBS^{+}02, NK02, NOS^{+}01, SHA^{+}03, WPS^{+}01, CB00, FRK^{+}01, ME00]. membranous [Bro03]. memories [Dov03-35, Wel01-32]. memory [Wel03-64]. MEN [Wel02j, DG01]. meningitidis [HEN^{+}01]. merge [Dov03-57]. Merotelic [CHM^{+}01]. Mesangial [KVM03]. Mesenchymal [ESS^{+}00, BJM^{+}02, DDV^{+}03, LCG^{+}04, NH03b, KIK^{+}00b, TN00b, YIW^{+}04, ZEtK^{+}00]. Mesenchymal-epithelial [YIW^{+}04]. mesoangioblast [PSD^{+}04a]. message [Vin04, Wel04-45, Wel02-51]. messages [Wel01-51]. Messenger [MS00, CPC^{+}02a, CGF^{+}04, Wel04v]. Met-mediated [SBG^{+}04]. meta [YMM^{+}00]. metabolic [CSJ03a, LeB04-35, KWOP00]. Metabolism [Wel04-46, CKS02, LGRP^{+}02, LG02a, ONM00, TSMS01]. Metabolite [KNM^{+}00]. metalloprotease [YSW02, KGC^{+}00]. Metalloproteinase [FWP^{+}00, UIY^{+}01, ECV^{+}00, IOLA^{+}00, KIC^{+}01]. metalloproteinases [JBA^{+}01, NMH^{+}04, HAP^{+}00]. Metaphase [DTO^{+}01, GSW^{+}00, GK04, GNH^{+}04, SHP01, SWB03, TBRG01, TEC^{+}03, AMEC01, YMM^{+}00]. metaphase/anaphase [SHP01]. metastasis [JLK^{+}02, WFI^{+}04, WCBC04, HPQ^{+}00]. Metastasizing [Wel03-37]. Metastatic [BZSC00]. Methionyl [KKK^{+}00]. methyl [BBDK^{+}04, UTH^{+}02]. methylation [BCB^{+}02, CNJ01, YF00]. Mex67p [SBH00]. Mex67p/Mtr2p [SBH00]. Mfn1 [CDE^{+}03]. Mfn2 [CDE^{+}03, KLG^{+}02]. Mgm1 [HBH^{+}04]. Mgm1p [WWG^{+}00, WWS^{+}03]. MHC [KRS^{+}01, RWK^{+}04, SDD04, ML04]. Mice [CBG^{+}01, HvdHG^{+}03, KHB^{+}04, RRM^{+}00, YIS^{+}03, BMM^{+}02, CKF^{+}03, DCC^{+}02, DPO^{+}04, EKdM^{+}04, FSCF^{+}03, HSC01, ICK^{+}04, KPL^{+}02, MDT^{+}01, MPG^{+}03, NOM^{+}04, NHG^{+}03, QPDJ^{+}02, RGM^{+}02, SGdM^{+}01, SMR^{+}02, Sre04h, TOTC01, TPA^{+}03, TBJ^{+}01, WST01, WLR01, XJW^{+}04, BWN^{+}01, CMW^{+}01, CMMP00, DG01, GBM^{+}00, GJ00, HWHH01, HVB^{+}00, KdVS^{+}00, LAF^{+}00, NHI^{+}00, PQF^{+}00, SJA^{+}00, TTHH00, TTP^{+}01, WFF^{+}01]. Micro [GS02]. Micro-RNAs [GS02]. Microautophagic [SM00]. microbial [Gal02]. microbicide [LeB02-75]. microdomain [WSC^{+}01, ZFH^{+}04]. Microdomains [CdLvM^{+}04, MF01a, PMPH03, RPS^{+}02]. microfibrils [ASGL^{+}01, BKZ^{+}01, GKSR00]. microfilament [HIG^{+}01, SB03]. microfilament-dependent [SB03]. micron [MYC^{+}02]. Microphthalmia [WHP^{+}02]. Microphthalmia-associated [WHP^{+}02]. MicroRNAs [Tum03j]. microscopy [DV04, OGD03, SAWS02, SP03, SSGLS01, McI01, PPM^{+}00, SGAS00, TSK^{+}00]. Microsomes [SHWH00]. microtubular [BN02]. Microtubule [APBC^{+}02, AC00, BMS^{+}00a, DLK^{+}02, FTD^{+}01, PMBC^{+}00, ABOS^{+}02, CFC^{+}01, CKS^{+}04, CBRBM04, DM02, DMH^{+}02, Dov02-28, Dov02-62, DWFAdL^{+}02, FAT^{+}02, FVC04, GES04, GV03, GK03b, GWBW02, HCL^{+}03, HBB^{+}04, HKBH03, ISS^{+}04, KBWG02, KAiK^{+}02, KKK^{+}02, LdCK^{+}01, MSC^{+}03, MSL^{+}02, MTT^{+}04, MWMK04, MPB^{+}04, MTM^{+}02, MKJ^{+}02, MHNSM03, Néd02, NOS^{+}01, OMM^{+}03, OER^{+}02, OWW02, RRB^{+}03, RTFW02, SAWS02, SLG02, SHCM03, TTH^{+}01, TUK03, VMH^{+}02, WBWS03, Xia03, ZWSC02, vBDH03, AKH00, FSKS00, GHC01, IdCAS^{+}00, KSN^{+}01, KYF00, KHN00, KMS00, SEI^{+}00, TB00b, TMD^{+}01, WBG01, WDLK01, YW00]. Microtubule-associated [DLK^{+}02, FTD^{+}01, RRB^{+}03, vBDH03]. microtubule-based [GV03]. microtubule-dependent [MTT^{+}04]. microtubule-destabilizing [SHCM03]. microtubule-independent [KBWG02]. microtubule-nucleating [TUK03]. microtubule-organizing [SLG02]. Microtubules [AZ03, KSW00, LeB02-56, LeB02-57, NH03a, RH00, WSDW^{+}00, Wel01-35, GMD^{+}02, HRV^{+}01, JdDD03, KAK^{+}03, LKM^{+}04, LeB02h, MR01a, MHlW02, SKF02, SRS01, Tum04i, Tum04w, Wel01z, Wel01-56, Wel04d, XB04, DR00b, FK01, HCOC00, MKST00, WSW^{+}00]. microvascular [XAB^{+}03]. microvesicles [Wel02c]. Microvillar [RB01]. microvilli [ZH04, DT00]. microvillus [LZS^{+}03]. microvillus-type [LZS^{+}03]. Mid1 [MMBM04]. Mid2p [BPC03, TMG03]. midbody [TFM04]. middle [KC02, RCY^{+}03, Tum04x]. midline [BHNG01]. midzone [MKJ^{+}02, MHlW02, XB04]. migrate [Wel03-46]. migrating [APBC^{+}02, SAWS02, Wel01m, BDK^{+}01, LWDH01]. Migration [ACE^{+}01, CK00, LeB04-60, NK02, CSIK03, CYC^{+}04, CAW^{+}04, CBLT04, DRBF03, DB02, FLS^{+}04, FAF^{+}04, GHK^{+}03, HTRK02, KSC^{+}04, KVC^{+}03, LeB02-39, LeB02-40, LeB02-82, LOC03, LPL^{+}04, MWL01, MZT^{+}03, MSI^{+}03, MGAL^{+}01, MPAP^{+}03, PSD^{+}04a, PSD^{+}04b, SC01a, SRD^{+}02, TSH^{+}04, TTH^{+}01, TWS^{+}04, TSY^{+}02, Wel04-55, WML^{+}03, WLWB01, Xia03, ZSY^{+}03, ZLVS02, ALJ00, AML00, BZSC00, HPQ^{+}00, HBSQ01, IOLA^{+}00, JPM^{+}00a, KIC^{+}01, KGC^{+}00, Mor00, ONM00, PBL^{+}00, Qua00, TTHH00, TWBV^{+}01, THE^{+}00]. Miha [ESH^{+}01]. Miha/Xiap [ESH^{+}01]. mimics [PHS^{+}03]. Minds [CW04]. mineralization [MSMK04, WK02, MBSB00]. Mineralized [Sre04h]. minerals [Dov02-68, Wel03-57]. Mini [Wel04-47]. Minicircle [AERD^{+}01, DE01]. Minus [KCG^{+}03, LeB03-57, NOS^{+}01]. Minus-end [KCG^{+}03, LeB03-57]. Miro [FMC^{+}04]. Misfolded [VN04]. Misfolding [SPN^{+}04]. mislocalized [WMG^{+}04]. misorientation [PRS^{+}04]. MISS [LTD^{+}02]. missegregation [BJB^{+}03, Dov03-34, PRS^{+}04]. Missense [BXR^{+}02, MPG^{+}03]. missing [Pow01e]. Mist1 [PRJK01]. mistake [Les01a]. Mistargeting [SC01b, TKB^{+}04]. Mitochondria [BAZA03, DSV^{+}00, LeB03-58, Tum04n, ASK^{+}03, Dov04b, GGB^{+}04, HIT^{+}02, JMB^{+}04, LeB02-70, LeB02-78, LeB04-53, PPP^{+}01, RPS^{+}02, Riz03, RMMP04, TVF^{+}03, ATE^{+}01, DSH^{+}00, NSL^{+}01, WGP^{+}00, WWG^{+}00]. Mitochondrial [ABRA03, APM^{+}02, HSMJ01, MMS00, WGvA^{+}01, vARP^{+}00, ASS02, BSMS03, CDE^{+}03, CLSK02, FMC^{+}04, HBH^{+}04, IYT^{+}04, KLG^{+}02, KJY04, KAC^{+}04, KKPB03, KSS^{+}03, MH01, MSP^{+}01, MN03, MJV^{+}03, MWE^{+}03, PLW^{+}04, RGG03, SNF^{+}02, TONN02, Wel02-37, WW02, WWS^{+}03, YpHRL03, YTM03, YWH04, YHB^{+}04, DKJ00, FSY00, FRK^{+}01, KIK^{+}00a, KRR^{+}01, MMWC00, PLH^{+}01, SJ01a, TN00a, vEWS^{+}01, vdB00]. Mitofusins [CDE^{+}03]. Mitogen [KY00, WFT^{+}01, CDEM00, CAP00]. Mitogen-Activated [DSSY00, KY00, WFT^{+}01]. mitogenic [SSRX04, SCTM00]. Mitosis [GKG^{+}01, SRS01, Wel01-36, AAD03, CSM03, Dov02p, Dov03-33, GV03, LeB02-57, LeB04-84, LGM^{+}04, MRK04, MCBB^{+}04, MP04, MRC^{+}02, MMBM04, SKK^{+}04, TCH^{+}02, TBRG01, WSC^{+}03, Wel03n, WWGK02, CDEM00, GHC01, IdCAS^{+}00, Nel00, PMSB01, RHM00, SC01b, WWJ^{+}00]. Mitotic [CENMR^{+}01, HMN^{+}00, LGW00, SKJ^{+}00, SCLC00, ARQ^{+}04, BJB^{+}03, BMS^{+}03, BLP^{+}02, BTVB03, CFB^{+}03, DMH^{+}02, HS04, HG03, HMC^{+}01, JCPWS01, JTK^{+}02, LMG04, LeB03-101, LC01, LP04a, MB03a, MKJ^{+}02, NPS^{+}03, OMM^{+}03, RRB^{+}03, RRSV02, SB03, WSWM04b, WBP^{+}03, WKZ^{+}02, YITe03, BT00, CHM^{+}01, DSSWW00, FAAS00, KTY^{+}00, KMCM00, SHHH01, SRHV00]. Mixed [Sre04i]. mixing [MWSL^{+}03]. Mkk [vdHvODML^{+}00]. MKlp2 [GNH^{+}04]. Mlc [TYY^{+}00]. MLCK [TWS^{+}04, Wel02-38, TYY^{+}00]. Mlp [ZWB04]. Mlp-dependent [ZWB04]. Mmm1p [HSMJ01]. Mmm2p [YHB^{+}04]. MMP [CSO^{+}04, GMRYM^{+}02, HBC^{+}03, LeB02-35, MCF^{+}02, SOH^{+}04, SHB^{+}03, SRD^{+}02, WSL^{+}03, KGC^{+}00]. MMP-2 [WSL^{+}03]. MMP-3 [WSL^{+}03]. MMP-9 [SRD^{+}02]. MMP-dependent [SHB^{+}03]. Mob2p [WKZ^{+}02]. mobile [Les01j]. mobility [BVH04, CPC^{+}02a, LWZ03, TNM^{+}03, DBS^{+}01, KLF^{+}00, SWJ^{+}00]. mobilization [FYI^{+}03, Riz03, WCIN04, TOM00]. mobilizes [GMY^{+}03, LeB03-42]. mode [GOV^{+}03]. model [BXR^{+}02, CNT03, Dov03-34, EMY^{+}04, Hel03, LeB04-61, LDP02, MMPO^{+}01, MZH^{+}02, RDN^{+}03, TCP^{+}03, CLWR01, VDMH01]. modeled [LeB04-42]. Modeling [VGL03, BLU^{+}04]. models [Dov03-33]. modes [DPB03, GOV^{+}02, LeB04-42, PR02]. Modification [KIO^{+}00, YSN^{+}01]. Modifier [WTG01]. Modifies [PGS^{+}01]. modular [BBBS04]. modulate [DSB^{+}02, LCS^{+}02, TKB^{+}04, BGS00, LFT^{+}00, MKK^{+}00a]. Modulated [KO00]. Modulates [vdHvODML^{+}00, GGNK04, MZT^{+}03, MPG^{+}03, PSP^{+}04, PAR^{+}04, PCB^{+}03, RJA^{+}03, TPW^{+}04, UTH^{+}02]. Modulating [Dov03-35, DCC^{+}02, GMRYM^{+}02, SEW^{+}01]. Modulation [DFYL00, KKK^{+}02, SSG^{+}02, WLO^{+}02, XLH^{+}01, DMC^{+}03, EMY^{+}04, GLDM01, GFM^{+}04b, HTS02, LC04, PLW^{+}04, SJW^{+}04, SGK02a, SPA^{+}04]. modulatory [MKS^{+}02]. Module [DVE^{+}00]. Modules [OLB^{+}00]. moesin [BYMS^{+}02, MWK^{+}02]. Molecular [AB00, KSM^{+}01, iNFK^{+}01, NG01, QKK00, RWSV03, BBG^{+}03, GG04, GTD^{+}02, MGP^{+}02a, RCY^{+}03, RSD^{+}04, SLP03, TONN02, TGD^{+}03, Vin04, WCA^{+}03, YLQ^{+}02, IHN^{+}01, NHI^{+}00, SJA^{+}00]. Molecule [FSGDN^{+}00, FKH^{+}04, HCL^{+}03, ISF^{+}01, LeB02-58, LSMS03, MGAL^{+}01, OMiKF02, SLD^{+}02, FSK^{+}00, KMCM00, MNHR00, SMSM00, SGO^{+}00, TGMC^{+}00, TNK^{+}00, ZN01]. molecule-1 [OMiKF02]. Molecule-5 [TNK^{+}00]. molecules [FSK^{+}04, SLB^{+}01, YMK^{+}04, TNM^{+}00, VMD^{+}01]. Molotov [Tum02d]. Mon1 [WSK^{+}03]. Monastrol [KMCM00]. Monitoring [ØTW04, Wel02-38, RMMP04]. Monoamine [WMT^{+}01]. monocyte [WLWB01]. Monocytes [AML00, CSL^{+}03]. monocytogenes [BCM04, GGT^{+}02, PAG02, SBG^{+}04, SPW00, SAG^{+}01]. monomer [ABCK^{+}03]. Monomeric [Wel04-48]. monomers [POW^{+}01a, GLA00, KHvOD00]. mononuclear [TTM^{+}03]. monopolar [LeB04-51]. Monoubiquitination [GRSL^{+}04b]. Morphogenesis [HRB^{+}01, KLGC^{+}01, COB01, CNT03, Dov03l, Dov03-33, GPL^{+}02, GLJP01, GGH^{+}04, KSD04, KCG^{+}03, MOMK03, MEK^{+}04, MNC^{+}03, NM02, PCB^{+}03, SML^{+}04, WVBY^{+}03, WSL^{+}03, ZH04, HGK^{+}01, HAP^{+}00, KGE^{+}00, LCS^{+}01, MLS^{+}01, MCB00b, TN00b]. morphogenetic [HT01, SCD02, TOTC01, ZHH^{+}02, LYMC00]. Morphogenic [GLA00]. morphological [ZCH^{+}02]. Morphologically [OMM^{+}03]. morphology [FCF^{+}01, FMC^{+}04, HBH^{+}04, KJY04, KYS^{+}02, MJV^{+}03, UOB^{+}02, TRW^{+}00]. Mosaic [ALWR01]. Mossy [UTH^{+}02, WSL^{+}00]. most [BHPN04, TSL04]. mother [Wel03-39, PMK^{+}00]. mothers [Wel03z]. Motif [KRA^{+}01, WZB^{+}01a, WMT^{+}01]. Motif-Containing [KRA^{+}01]. motifs [DAV^{+}03, NT00]. Motile [YW00]. Motility [KLGC^{+}01, Mel00c, SHN^{+}01, SPW00, XHG^{+}00, FFSF03, HMH^{+}03, HDH02, KSD04, KI04, SRK^{+}03, SKN^{+}03, SAG^{+}01, SC04, SPA^{+}04, UGKT^{+}02, WZB^{+}01a, WHD^{+}03, WWGK02, XLH^{+}01, YWH04, KIK^{+}00b, KH01, KHH^{+}01, LVWA01, PS00]. motion [DSH^{+}03, LeB02-84, LeB04-46, JLS^{+}01]. motogenic [CAGK^{+}03]. motoneuron [BXR^{+}02, Dov02l]. motoneurons [RJA^{+}03, SGK^{+}02b]. motor [ABCK^{+}03, BXR^{+}02, Dov02-32, Dov02-41, Dov03h, GV03, HBH^{+}04, HvdBP^{+}01, LS02, LeB03-57, LJK^{+}01, MMBB03, NH03a, Néd02, NOS^{+}01, TVF^{+}03, Val03, XTN^{+}02, BMS^{+}00a, PCR^{+}01, TYS^{+}00, TUV00]. Motoring [Dov01w, Wel04c]. motorized [Dov01c]. motors [Dov03-56, GWBW02, GTD^{+}02, RPTNM01]. mounting [Jay01]. Mouse [GBMA04, AR03, BSW^{+}04, BXR^{+}02, CMC^{+}02, Dov02-48, EMY^{+}04, HTPC04, KLT^{+}03, LCT^{+}04, LCG^{+}04, MSJ^{+}02, MEK^{+}04, NH03b, RDN^{+}03, RMMP04, RKF^{+}04, SSM^{+}04, SR03, TKS^{+}02, THG^{+}04, VMS^{+}02, VLL^{+}03, XLH^{+}01, CLWR01, ECO^{+}00, GKSR00, GMZ^{+}00, KSF^{+}00, LMW^{+}00, MSFH00, MYH^{+}01, MLS^{+}01, PDV^{+}00, ToIB^{+}00, WCGT^{+}00]. mouths [Dov02r]. move [Dov03h, LeB02-101, Tum04n, Wel01-42, Wel04e, CG00]. Movement [DLXP00, Wel03-38, Dov02-47, DBS^{+}03, GSB^{+}03, HRV^{+}01, HGP^{+}04, LeB02-42, SCTF04, Sre04j, CDTW00, GHC01, KGvdG^{+}00, PGSL00, SIBG01]. movements [DMC^{+}03, Dov02-39, IUK04, KLF04, SAWS02, SH02, AC00, ASP^{+}00]. mover [Dov04l]. moves [LeB02-91, MKST00]. Moving [Wel04-49, Dov01d, OL02]. Mp1 [WFT^{+}01]. Mpf [PMBC^{+}00]. MPR [Dov03-43]. Mps1p [CMM^{+}02]. Mps3p [JGW02]. MRE11 [FP02]. mRNA [AOJ^{+}04, AZP^{+}02, BCB^{+}02, CBS04, Dov03m, FHM^{+}03, GI02, GBD^{+}00, GPZ^{+}02, HBAF^{+}02, KCL^{+}00, KBG^{+}03, LGM^{+}01, ONS^{+}00, PZP^{+}01, RJA^{+}03, SBS02a, SBH00, VSO^{+}01, Wel03t, WMHB^{+}00, vEPP^{+}01]. Mrnas [BGS00]. Mrnps [MÖS^{+}00]. MS [Wel02-39]. Msb3p [GAT^{+}03]. Msb4p [GAT^{+}03]. Msc2 [EWM^{+}04]. Msl [JWJJ00]. Msn2 [JRL^{+}03]. Msn4 [JRL^{+}03]. Msn5p [YB01]. MT-SP1 [LSW^{+}03]. MT1 [CSO^{+}04, GMRYM^{+}02, HBC^{+}03, MCF^{+}02, SOH^{+}04, KGC^{+}00]. MT1-MMP [GMRYM^{+}02, SOH^{+}04, KGC^{+}00]. MT1-MMP-dependent [CSO^{+}04, HBC^{+}03, MCF^{+}02]. Mtdna [HSMJ01, LeB03-102, YHB^{+}04]. Mtg8 [WND^{+}00]. Mtmr2 [BBP^{+}04]. mTOR [DWB03, EPN^{+}03, JHS^{+}02]. mTOR-dependent [DWB03]. Mtr2p [SBH00]. mua [BWV^{+}01]. mua-3 [BWV^{+}01]. MUC2 [BvdWD^{+}04]. Mucosa [WCGT^{+}00]. Mucosal [LBWH^{+}00, MF01a]. muddled [LeB04-57]. Multi [MMS00]. Multi-Step [MMS00]. Multicolor [SDN^{+}00]. multidomain [ESC^{+}01]. Multimembrane [Dov03-36]. multimeric [MLZ^{+}01]. Multinucleated [HFM^{+}01]. Multiple [CFWH^{+}01, FWP^{+}00, HIN^{+}03, JMG01, KGE^{+}00, MLLA00, DAV^{+}03, LTF^{+}01, SJB^{+}03, WAC^{+}03, BMM^{+}01, NS01, NSW00, PS00, SJS^{+}00, SKCS00]. multipotent [BCA^{+}03, HAK^{+}04]. multiprotein [CBW^{+}01]. multisite [MHK04]. Multispanning [HW00a]. Multistep [TCR00]. multivesicular [BBMS03, FCBH01, KRS^{+}01, LBS^{+}02, LO04]. Munc18 [BJ03]. munching [Wel01-34]. mungo [TOM01]. MUP [HEW^{+}01]. MUP-4 [HEW^{+}01]. Murf [SEI^{+}00]. murine [AAM^{+}04, WLR01, ZHH^{+}02, CDFT^{+}01, DEG00]. muscarinic [CAW^{+}04, JLJD03]. Muscle [AP00, AKH^{+}04, BMM^{+}02, BL01, Dov01x, HES00, LeB04-62, MKS^{+}02, TCH^{+}00, ZGN^{+}04, AOJ^{+}04, ASGGR02, BJM^{+}02, BKD^{+}04, CSP^{+}04, CSL^{+}03, DDV^{+}03, Dov03-63, GBN^{+}01, GNS^{+}04, GBJ01, HGC02, HP03, HYMS^{+}02, IKS^{+}01, JLJD03, JJM^{+}02, KSNS^{+}04, KK02, KNK^{+}01, LCRS01, MRM^{+}04, MDF01, NDM^{+}03, NM02, PR02, QPDJ^{+}02, RLZ^{+}03, RFLT02, RMMP04, RDNB02, SML^{+}04, TAA^{+}02, TCP^{+}03, VMS^{+}02, WHM^{+}02, BHY^{+}00, BHFL01, BL01, BMF00, CDFT^{+}01, DBS^{+}01, DCM00, EHZ^{+}01, FRO01, FKG00, FHP00, GH00, HSW00, HFM^{+}01, LR00, LQPC^{+}00, LMW^{+}00, MCG^{+}00, MMWC00, ONS^{+}00, Ono01, PGS^{+}01, RMHM00, RMG^{+}00, SEI^{+}00, TTS00, TTP^{+}01]. Muscle-specific [BMM^{+}02, MKS^{+}02, TCH^{+}00, AOJ^{+}04, WHM^{+}02]. muscles [BBG^{+}03, Dov01z, Dov03-47, LeB04-65, Wel04x, RBV00]. Muscular [Cha00, MCB00a, RRM^{+}00, Dov01-30, Dov02-42, KSNS^{+}04, MPG^{+}03, RJA^{+}03, TCP^{+}03, TBJ^{+}01, WST01, Wel04-50, BWN^{+}01, CZBH00, CDFT^{+}01, vdVWS^{+}00]. MuSK [CSG^{+}04, LLGB03, SHW01]. must [PBP^{+}01]. Mutagenesis [BRM^{+}00, MHT^{+}04, RSBE00]. Mutant [ARM02, ALWR01, LdVV^{+}02, ARMB04, BXR^{+}02, IKS^{+}01, KRF^{+}03, KBK^{+}03, MDT^{+}01, OWW02, PRLR02, CLT^{+}01, LCM00, LAF^{+}00, PLL^{+}01, RS00c, SJA^{+}00]. Mutants [GSW^{+}00, PGV^{+}00, GNDLS^{+}01, HS02, SKT^{+}03, TDL03, DTO^{+}01, NS01]. mutation [BXR^{+}02, EPH^{+}03, GK03b, HBB^{+}02, MPG^{+}03, RKF^{+}04, SHE^{+}02, SMR^{+}02, GKSR00, LMW^{+}00, WCGT^{+}00]. Mutational [SWJ^{+}00]. Mutations [DSG04, HDJ00, HHHJ00, ISS^{+}04, XSK^{+}01, ML04, YHZ^{+}01, MBSB00]. mutual [SKT^{+}03]. MVB [Dov01k, KSBE03]. My [Wel03-39, Sod02]. Myc [Wel04-51, YSC^{+}02, YSC^{+}21, SSM^{+}04]. myc-induced [SSM^{+}04]. mycobacterial [FBG^{+}01]. Mycobacterium [RMR02]. Myelin [MDP02, Sre04j, BBP^{+}04, GLS^{+}03, RKF^{+}04, SKM^{+}02, YHT02, SKT^{+}00]. Myelin-associated [MDP02, YHT02]. myelin-directed [GLS^{+}03]. myelinated [PSE^{+}03, TGD^{+}03]. myelinating [PMU^{+}02, SHS^{+}00]. Myelination [Sch02, ZEW^{+}01, CS03a, CRS^{+}03, Wel03-61, XSK^{+}01, GVT^{+}00, YKW^{+}00]. Myeloid [MRM^{+}00]. Myf5 [BHY^{+}00]. Myo1p [LBP00, LVD^{+}04]. Myo2p [HvdBP^{+}01, RPTNM01, RRB^{+}01]. Myo4p [EKC^{+}03, KJB^{+}02, RPTNM01]. myoblast [HNK^{+}03, LMDS^{+}03, Coo01, FLLE^{+}01]. myoblasts [CMC^{+}02, GCG^{+}01, GBN^{+}01, MSJ^{+}02]. myocardium [MEK^{+}04, SYY^{+}03]. myocyte [IKA^{+}03, PGS^{+}01]. myocytes [MPR^{+}03, PWC^{+}01, THK^{+}00]. MyoD [ARM02, GBN^{+}01, GNS^{+}04, GBD^{+}00]. MyoD-positive [GBN^{+}01]. myofibril [FSCF^{+}03, LMDS^{+}03, MDF01]. myofibrillar [KNK^{+}01]. myofibrillogenesis [LPPT^{+}02]. myofibrils [BBG^{+}03]. myofibroblast [HGC02]. myofibroblasts [FLWMG02, RL03]. myogenesis [BJM^{+}02, EPN^{+}03, vdFKK^{+}02, PCC^{+}00]. Myogenic [ASGGR02, MGP^{+}02b, ARM02, GCG^{+}01, HPS^{+}04, TAA^{+}02, HLZW00, SEI^{+}00]. myogenic-endothelial [TAA^{+}02]. Myopalladin [BMM^{+}01]. myopathy [ML04, ZCH^{+}01]. myopodin [WSF^{+}01]. Myosin [AP00, Cra00, DLY^{+}02, EKT^{+}00, GAB^{+}01, HZS^{+}01, KKT04, LBP00, MKTW01, MMBM04, REM^{+}02, SB03, Val03, WH03, Wel01-37, Yum01, BWW^{+}02, CWG^{+}02, Dov02p, Dov03-62, GGD^{+}04, ICN^{+}03, JLJD03, KRF^{+}03, KK02, KRMB03, KI04, KJB^{+}02, LVD^{+}04, RL03, RFLT02, RSK02, Wel04-48, WWGK02, CDTW00, NS00, ONS^{+}00, PVjL^{+}00, LP04b, SCB02, TM04]. myosin-1A [TM04]. Myosin-I [GAB^{+}01, EKT^{+}00, LBP00]. Myosin-II [MMBM04]. Myosins [JRW^{+}01, DSN^{+}01, RSD^{+}04, LSSL00, Mac00, RPTNM01]. myositis [CKF^{+}03]. Myostatin [MTM^{+}03]. myotome [HT01]. myotrophin [GPLS02]. myotrophin-induced [GPLS02]. myotube [KYZ^{+}04]. Myotubes [EGS^{+}04]. MyRIP [DSH^{+}03]. myristoyl [OTB03]. Mystery [Min01, NB03b].

N [CMC^{+}02, KIK^{+}00b, APLB00, CYC^{+}04, FSM^{+}01, GHS00, HIE^{+}01, HPQ^{+}00, LeB04-49, MSH^{+}00, RyHHK00, STP^{+}00, SPH^{+}00, TRC^{+}00, UTH^{+}02, XLH^{+}01]. N- [STP^{+}00]. N-Cadherin [KIK^{+}00b, LeB04-49, XLH^{+}01]. N-cadherin-dependent [CMC^{+}02]. N-methyl-d-aspartate [UTH^{+}02]. N-Wasp [TRC^{+}00]. n89 [IEJ^{+}01]. NAADP [GMY^{+}03, BPMG00]. NADPH [KWS^{+}02, SHA^{+}03]. nail [WLR01]. name [Dov02d]. Nanofibers [LeB04-63]. Nanometer [KAK^{+}03]. nanoscopes [Wel04-43]. NAP [LeB04-91]. Nap1 [MST^{+}00]. NAPO [SISO01]. Narcissistic [LeB04-64]. Nascent [BDK^{+}01, EKB^{+}03, NKM^{+}03, Tum04f, UHR^{+}03, Mar01, MDJF00]. native [LMHJ02, BLC00]. Natural [SNL^{+}00, KB03, MWHM01, Wel03-29]. NBS1 [DC03, FP02]. Nbt [PBL^{+}00]. Nc1 [ACE^{+}01, KLGC^{+}01]. Nc1/Endostatin [ACE^{+}01, KLGC^{+}01]. NCAM [LSMS03, LSMS03, NDS^{+}02]. NCAM-mediated [LSMS03]. Nck [CRP^{+}04, GQI^{+}02]. Ndc80p [WK01]. Ndj1p [TSDS00]. near [JLS^{+}01, LCI^{+}01, UHR^{+}03]. Nebulin [OLB^{+}00]. Necessary [BHL^{+}01, GPLS02, HKP^{+}04, KI04, LeB02-39, LC01, STA03, TE01, ZHH^{+}02, JCR^{+}01, SGPL^{+}00]. neck [ABCK^{+}03, Wor03]. neckless [OWW02]. Necrosis [DSV^{+}00, IIN^{+}01, RBD^{+}01, CWMO00, FWM^{+}01]. Necrotic [MSO^{+}00]. Nectin [MNK^{+}02, TNM^{+}00]. nectins [FSK^{+}04]. Nedd4 [PLL^{+}00]. NEDD8 [TOTC01]. need [Dov03h, LeB04m, Wel04-76]. needed [SDML04, PWU00]. needles [Wel03-32]. needs [Dov02-43, LeB02-38, LeB02-93, LeB04-89]. Nef [JKB^{+}03, RWK^{+}04]. Negative [KMB^{+}01, KCL^{+}01, PSS^{+}04, WYR^{+}02, WBT^{+}03, HCK^{+}00, PLL^{+}01, UIK^{+}01]. Negatively [SKF^{+}01, CK03, FFSF03, GQI^{+}02, MDQ^{+}03, MTB^{+}02, MTM^{+}03, SM04b, WBAS04]. neighborhood [Dov02-27]. Neisseria [HEN^{+}01, LBH^{+}02]. Neither [DOL^{+}01]. Neogenesis [PPK^{+}01]. neogenin [KYZ^{+}04]. Neoplasia [IEJ^{+}01]. neovessel [CSO^{+}04]. nephrogenic [KBK^{+}03]. nephronectin [BSL^{+}01]. Nerve [AKDS00, KMLS04, CS03a, NK02, PSW^{+}02, SGK^{+}02b, SATA^{+}02, BLPP01]. Nerve-independent [KMLS04]. Nerveless [LeB04-65]. Nerves [WFQ^{+}00, Wel01q]. Nervous [Wel02-40, Wel03-40, BHNG01, DWM03, HHS03, SWBE^{+}04, SR03, BAD^{+}00, CMW^{+}01, MHW^{+}00]. NES [RXS^{+}03]. NES-mediated [RXS^{+}03]. Nesca [MKM04]. ness [LeB02-95]. Netrin [KMH^{+}00, MTW^{+}04]. Netrin-1 [MTW^{+}04]. Netrins [KYZ^{+}04]. NETs [Wel04-70]. Network [HGP^{+}00, HLB^{+}00, Tum03e, ARQ^{+}04, CEGZ^{+}04, DLPB03, HMVG02, KMS^{+}04, PNSJ01, SBC^{+}03, Tum03i, VYW^{+}03, VGL03, WKZ^{+}02, Wel04b, WWGK02, YGWN01, KKR^{+}00, MJG^{+}01, PS00, WJG^{+}00]. networks [CFC^{+}01, DBH^{+}01, YCB04, KYF00, STE^{+}01, WSDW^{+}00]. Neural [Dov03-37, Dov03-38, FSM^{+}01, LeB03-60, LSMS03, SLD^{+}02, BMS^{+}03, BSW^{+}04, DRBF03, FLS^{+}04, HBK^{+}02, HAK^{+}04, MRM^{+}04, PDL^{+}03, SK04, SC04, VFN^{+}04, Wel04-35, XLH^{+}01, ZVPK03, BHFL01]. Neuregulin [ZEW^{+}01, NOOG^{+}04, SRKR00]. neuregulin-induced [NOOG^{+}04]. neurexins [SST^{+}01]. neurite [HN03, HBSJ04, LSMS03, MKM04, SDS^{+}04, SSW^{+}01a, TYA^{+}02, TPW^{+}04, WLO^{+}02, FLX^{+}00, GKM^{+}01, MAAZ^{+}00, TYS^{+}00]. neuritogenesis [NDS^{+}02, NYT^{+}03, SMZ^{+}03]. neuroblast [IOIF^{+}04, YCK^{+}03]. neuroblasts [RAS^{+}03, WRGK00]. Neurocan [LLH^{+}00]. neurodegenerative [LdVV^{+}02]. neuroendocrine [HViV^{+}02, MPV^{+}01]. neurofascin [GSSP03, RWCC01]. Neurofilament [ATG^{+}03, RGM^{+}02, REM^{+}02, RCY^{+}03, XRH^{+}03, SHS^{+}00]. neurofilaments [ATG^{+}03, Dov02-41, REM^{+}02, SVT^{+}02, AGB^{+}00]. neurogenesis [KFR^{+}04, Tum04o]. neurogenin [HViV^{+}02]. Neuroglian [GF03]. neuromuscular [CSG^{+}04, HYMS^{+}02, LLGB03, AKK^{+}00, JCR^{+}01]. neuron [Dov03-51, JB01, MWAM01, MMBB03, SSG^{+}02, SLD^{+}02, WWD^{+}04, PCR^{+}01]. neuron-enriched [SSG^{+}02]. neuron-to-neuron [SLD^{+}02]. Neuronal [ARLC^{+}04, EES^{+}01, MCB00a, ABRA03, CFM^{+}02, DLY^{+}02, DFZ^{+}03, FCM^{+}01, GSB^{+}01, LeB03-46, MDQ^{+}03, MOMK03, MHS01, PKR^{+}02, PSK^{+}03, PHM^{+}02, RPZ^{+}02, RLTC^{+}02, SKF02, SRG^{+}04, SZvBuH^{+}04b, SNS^{+}04, TSH^{+}04, TTH^{+}01, Tum04b, WLPD04, XRH^{+}03, DKJ00, Ern00, EHM^{+}00, KSN^{+}01, PDJ00, RGG00, TTHH00]. neuronopathy [BXR^{+}02]. Neurons [DFYL00, LeB02-60, LeB02-59, LeB03-61, Wel03-41, Wel04-52, ABF^{+}03, BCA^{+}03, CJ02, CSJ03a, Dov02-45, EOB^{+}04, GCH03, GPZ^{+}02, HHS03, Lai03f, LS02, LRBH02, LeB03-62, LDK^{+}03, OMB^{+}01, PUK02, PPK^{+}01, RKR^{+}03, RBD^{+}01, SVT^{+}02, SSW^{+}01a, SLB02, WMG^{+}04, WAC^{+}03, YJS^{+}03, ZMGL02, AGB^{+}00, FXPT00, GLK^{+}00, HGB^{+}00, MHE^{+}00]. neuropathy [BBP^{+}04, PQF^{+}00]. Neuropilin [SMSF00]. Neuropilin-1 [SMSF00]. neurotoxic [RDN^{+}03]. neurotoxin [DRG^{+}03, MAAZ^{+}00]. Neurotransmitter [pHYpXL00, TRMI^{+}04, WWBGG03, Lit00]. Neurotrophic [DFYL00, KWS^{+}02, OMB^{+}01, SK04]. Neurotrophin [pHYpXL00, LS02, MKM04]. Neurotrophin-3 [pHYpXL00]. neurotrophin-induced [MKM04]. neurotrophins [EES^{+}01]. neutral [LXL^{+}04b]. neutralize [WAOC^{+}03]. neutrophil [Dov03-39, SWG^{+}03, PWC^{+}01, SD00]. Neutrophils [LeB02-61, DH02, GCO^{+}04, GHS00]. never [LeB02-81, LeB03y]. newly [LFM^{+}04, yZCKA01]. Next [LeB04-66]. NF [SRSW04, CSIK03, GLS^{+}03, KSR^{+}04, SLR^{+}03, SDS^{+}04, ZLR^{+}03, GV00, MKK^{+}00b]. NF- [SDS^{+}04, GV00, MKK^{+}00b]. NF-M [GLS^{+}03]. Nfat [HP02, SYY^{+}03]. NFATc [LCRS01]. NFATC2 [HP03, HFM^{+}01]. NFATC2-dependent [HP03, HFM^{+}01]. NG2 [ACBG04, BCA^{+}03, LeB04-67]. NG2-expressing [ACBG04]. NgCAM [Dov03-40, FSK^{+}00]. NGF [CJ02, CSJ03a, PKR^{+}02, TPW^{+}04]. NGF-deprived [CJ02, CSJ03a]. NGF-induced [TPW^{+}04]. NH [PBD^{+}04, HGC02, HWBD^{+}01, ZCW^{+}03, ARK^{+}00, BRM^{+}00, HSW00, HCD^{+}00]. NHE1 [DB02]. NHPX [GBOL03, LL02]. NHPX/15.5 [GBOL03]. NHPX/15.5-kD [GBOL03]. nicastrin [KLE^{+}02]. niche [Tum02c]. nick [Les01f]. nicotinamide [SNF^{+}02]. Nicotinic [BPMG00, ANC^{+}02]. nidulans [EM00]. Niemann [MHT^{+}04]. nil [Wel01-64]. Nim1 [WBAS04]. Nischarin [ALJ00]. nitric [MRT^{+}01, RP03a, WST01, GLK^{+}00]. Nitrosylation [MSP^{+}01, RP03a]. NITY [XBL^{+}03]. Nk [SNL^{+}00]. NKX2-5 [AKM^{+}04]. Nmd3p [HKJ00]. Nmd5p [PC01]. nNOS [AMF01]. No [LeB04-68, LeB04-69, EC03, Dov01-30, LeB03-67, LeB03-84]. noble [Dov04p]. Nodal [PDL^{+}03]. Nodal-dependent [PDL^{+}03]. node [Jeg01a]. nodes [Dov03-44, GWL03, KS02, LGGS^{+}04, BAD^{+}00]. Nogo [PSW^{+}02]. Nogo-A [PSW^{+}02]. nomenclature [GAB^{+}01, LDC^{+}04]. Non [WWK^{+}00]. Non-Cell-Autonomous [WWK^{+}00]. noncardiomyocytes [MWN^{+}04a]. noncell [MYO^{+}04]. noncell-autonomously [MYO^{+}04]. noncoding [MPG^{+}02]. nonconventional [VBH^{+}02]. nonhelical [BMY^{+}01]. Noninvasive [IBP^{+}04]. nonmembranous [Bro03]. nonmitochondrial [YJS^{+}03]. nonmuscle [KSNS^{+}04, RSK02]. nonphosphorylated [GAC^{+}03, KHvOD00]. Nonprocessive [RPTNM01]. nonraft [NGKH02]. Nonreceptor [APLB00]. nonsecretory [JAS02]. nonselective [JOF^{+}02]. Nonsense [LMW^{+}00]. Nontranscriptional [SJW^{+}04, CLB^{+}03]. nonvesicular [FR01]. Normal [HES00, MGLPM00, SGdM^{+}01, CKW^{+}03, KSD04, MCG^{+}03, SMS^{+}01a, UOB^{+}02, ZHH^{+}02, SMI^{+}00]. Notch [BYLA^{+}01, SKF^{+}01, Wel03-42, AMBW04, BDR^{+}03, DA01, Dov01b, GRSL^{+}04b, KRU^{+}04, LeB03-54, PSS^{+}04, WS00]. Notch1 [GLSG^{+}02]. Novel [CAW^{+}04, FSM^{+}01, GSB^{+}01, HS04, MMS00, RDH^{+}01, RGGL00, VSO^{+}01, WFT^{+}01, WE02, AvdWM^{+}01, AKM^{+}04, ABRA03, AAM^{+}04, BYMS^{+}02, BHNG01, BWV^{+}01, BSW^{+}04, BRR^{+}02, BSL^{+}01, BM02, CFC^{+}01, CBW^{+}01, CKS02, CKL^{+}03, CSG01, DRP^{+}03, FSTC02, FMC^{+}04, GJS^{+}03b, GSN^{+}04, HSMB02, HEW^{+}01, HRM02, IDvH^{+}02, IYT^{+}04, IKA^{+}03, JGW02, KSNS^{+}04, KK02, KR03, KPKY^{+}03, LLGB03, LM01, LL02, MKM04, MEFC03, MLKH04, MZR^{+}04, OER^{+}02, PPP^{+}01, POH^{+}04, QPDJ^{+}02, RRB^{+}03, RBE^{+}02, SISO01, STA03, SRC^{+}01, SRSW04, UJK^{+}02, WSF^{+}01, WLR01, YLY^{+}02, YJS^{+}03, dMMBK^{+}02, vdLBK^{+}04, ABP^{+}00, ALJ00, BMM^{+}01, CW01, EHM^{+}00, sGS01, HJSM00, IdCAS^{+}00, JSCR01, KMH^{+}00, LHvdH00, MPSM00a, NHS00, NSL^{+}01, NCUJ^{+}00, PO00, PDW^{+}00, PLH^{+}01, RMG^{+}00, RBB00, SFV^{+}00, SL01, SGO^{+}00, TWBV^{+}01, TP01, WDLK01, WWJ^{+}00, WWKV00, YSS^{+}01, vEWS^{+}01, vdVWS^{+}00]. novo [KRS^{+}02, iONOM02]. Np95 [BLP^{+}02]. NPC [BRB^{+}01, LeB03h, LRD^{+}03, SKK^{+}04]. NPFX [HHOP02]. Npl4 [YMR03]. Nr [SLB^{+}01]. Nrcam [FSK^{+}00, Dov01e, MKR01]. NRG [LeB03-62, BWRT03]. NRG-1 [LeB03-62, BWRT03]. Nsec1 [YSGS00]. NSF [MSN^{+}02, WPM^{+}00]. Nuclear [AFN00, ASAJ01, BHL^{+}01, DBH^{+}01, DLS00, Dov02-46, ESS^{+}00, Ern00, FNKH02, Gil01, HM00, HRE^{+}01, KYM04, LeB02-63, LDS^{+}00, LRD^{+}03, LHW^{+}01, LG01, MYKG00, MTPT02, Mor00, MJG^{+}01, NM00, PTM^{+}01, RM01, RRM^{+}00, SS04, SK04, SBH00, TMA^{+}04, WND^{+}00, YB01, BCG03, BBBS04, BRR^{+}02, BM02, CSIK03, CKZ^{+}02, DMA^{+}01, Dov01x, Dov02h, DW02, FGS^{+}02, GI02, GMY^{+}03, GNDLS^{+}01, GPZ^{+}02, GPLS02, HDL02, IKS^{+}02, LeB02-54, LeB03-101, LXL04a, LOC03, LCRS01, LSMS^{+}01, LMM^{+}02, LGB^{+}02, MKM04, MSM^{+}04a, MEFC03, MKS^{+}02, MSY^{+}04, NMHH03, OSNG04, OBG^{+}03, PPWM04, PC01, RXS^{+}03, RMW03, SERB03, SHM02, ST03, SCP02, SSRX04, SGW^{+}02, SRKN03, Tum02e, Tum04-27, WCIN04, WPC^{+}02, WSS^{+}04, WSF^{+}01, Wel01-27, Wel03f, WAV^{+}04, Xia03, ZW04, ZFH^{+}04, AC00, ABP^{+}00, BEG01, BHBJ00, BKB^{+}01, BMKA01, BLPP01, CAP00, DBS^{+}01]. Nuclear [DSP^{+}01, FL00, HKJ00, HBL^{+}01, LCM00, LGM^{+}01, MLC^{+}01, MHE^{+}00, MSHG00, PZP^{+}01, RRM^{+}03, RAS^{+}00, SGG01, SL00, SMWG00, SMTC00, TMD^{+}01, YSN^{+}01, YF00]. Nuclear-fallout [RRM^{+}03]. nucleating [TUK03]. Nucleation [HP00, SPW00, ZBB^{+}00, BCM04, CG03, SAG^{+}01, ADKK00, GHS00, YHD^{+}00]. nucleator [Dov03-42]. Nuclei [PMBC^{+}00, BPD^{+}04, CKW^{+}03, ZAE^{+}04]. nucleo [SBMB^{+}04]. nucleo-cytoplasmic [SBMB^{+}04]. Nucleocytoplasmic [MSM^{+}04a, NA00, VPP^{+}01, AJ01, BKS^{+}02, JRL^{+}03, KBG^{+}03, ZW04]. Nucleocytosolic [HCD^{+}00]. Nucleoids [HSMJ01, KKPB03, YHB^{+}04]. Nucleolar [CH01, KKK^{+}00, SGYD^{+}01, BPD^{+}04, GBOL03, LL02, LGM^{+}04, FMP^{+}00, VAHV00]. Nucleoli [LeB03-63, YSN^{+}01]. Nucleolin [CPA^{+}03, DB00]. Nucleolus [GDHS01, GSP^{+}02a, PLP02, SHVR02, CH01, DMO00, PP00]. Nucleoplasm [CH01, KLF^{+}00, MÖS^{+}00]. nucleoporin [BCG03, DMA^{+}01, RXS^{+}03]. Nucleoporins [SBH00, MSM^{+}04a, VSO^{+}01, BEG01, MLC^{+}01, SMWG00]. Nucleosomal [SKJ^{+}00]. Nucleosomes [LeB03-64]. nucleotide [AQC04, BPPFM^{+}03, OMWSN02, VCGB^{+}02, YEG01, WSFN00, WBM^{+}00, WSE00]. Nucleotides [LeB03-65]. Nucleus [vMZM^{+}00, AZP^{+}02, BAZA03, BKS^{+}02, Dov03p, LeB03-79, LeB04-70, MAG^{+}04, SRW^{+}04, TSH^{+}04, Wel01-62, HLRG01]. nucleus-encoded [AZP^{+}02, SRW^{+}04]. Nude [EM00]. Nude/Ro11 [EM00]. Nudel [LYL^{+}04]. Nudf [EM00]. null [FSCF^{+}03, TDL03, WMG^{+}04, WLKS01, GJ00, KHH^{+}01, WCGT^{+}00, DCC^{+}02]. nullo [HW00b]. Num1p [FK01, HCOC00]. NuMA [Wel01-43, MHS^{+}00]. Numb [SPS^{+}00, PSS^{+}04]. number [VTGT^{+}03, PN00, vRTvdB^{+}00]. Numbers [Les02d, LeB02k, Wel02-41]. Nup133 [BBBS04, VSO^{+}01]. Nup153 [MSM^{+}04a]. Nup160 [VSO^{+}01]. Nup214 [MSM^{+}04a]. Nup2p [DMA^{+}01, DSP^{+}01]. Nup358 [SERB03]. Nup60p [DMA^{+}01]. Nup82p [GNDLS^{+}01]. Nup84p [SLSR^{+}00]. nurse [Wel01-38]. NuSAP [RRB^{+}03]. nutrients [EPN^{+}03]. Nxt1 [BHL^{+}01].

O [ADL^{+}03, DC02b, KRU^{+}04, XJW^{+}04]. O-glycans [XJW^{+}04]. OAPs [CPN^{+}01]. oblivion [LeB02-46]. Obscurin [YEG01, BBG^{+}03]. observation [MSC^{+}03, SD00]. observed [VMS^{+}02, TSK^{+}00]. Observing [WPO00]. Occludentes [FFST01]. Occludin [LM00a]. occupancy [Wel03-30]. occur [ALP^{+}04, MEV^{+}04]. occurs [APM^{+}02, GLDM01, DR00b, MGLPM00, YSN^{+}01]. off [BJ03, Wel01k, Wel02j, Wel04o]. offspring [Wel03z]. OH [JHS^{+}02]. oils [Dov02-31]. Old [Wel01-39, Wel02-42, Dov02-45, Wel03z]. olfactomedin [HV03]. OLIG2 [SK04]. oligodendrocyte [CRS^{+}03, CROfC04, GLSG^{+}02, LXL^{+}04b, Nob00, TGMC^{+}00, TTR00]. oligodendrocytes [EAD^{+}02, HAK^{+}04, PMKV01, TWBV^{+}01]. Oligodendroglial [EMY^{+}04]. oligomer [Tum03m]. Oligomerization [KLGC^{+}01, GTM^{+}01, PSP^{+}04, ZN01]. Oligomerization-Dependent [KLGC^{+}01]. oligomers [Tum04l, GLA00]. oligosaccharyltransferase [NKM^{+}03]. Omp85 [GGB^{+}04]. Oncogenic [FRO01, KY00, LM00a, ZEtK^{+}00, BDKM04, YSC^{+}02, YSC^{+}21]. Oncoprotein [SSA^{+}00]. One [Tum04o, Tum04p, Wel02-43, Wel03-43, Dov01j, Dov02-48, Dov02-47, Dov02-56, Dov02-64, Dov03r, Dov03-56, Dov04j, Dov04o, HS02, KRF^{+}03, LeB03-61, Tum03c, Wel01l, Wel01-64, Wel04-44]. ongoing [MWAM01]. Only [Dov02-48, Dov02-47, AZ03, Dov03r, LeB03-41, LVD^{+}04, Wel04-57, MSJ^{+}02, WAOC^{+}03]. ons [Dov02-63]. onset [RKF^{+}04, SKK^{+}04]. onto [Wel03-31]. oocyte [KLT^{+}03, MH02, MB01, SM04b]. oocytes [LRD^{+}03, THG^{+}04, WMHB^{+}00, YSN^{+}01]. Oogenesis [TSMT00]. Op18 [BKDH01]. open [LeB02-51, LeB02-109, LHR04, Wel04-36, LCGR00]. opener [VCGB^{+}02]. Opening [LeB04-71]. opens [Tum04j]. operated [MH02]. Operating [PPM^{+}00]. opposing [DLY^{+}02, HTRK02]. opposite [GWBW02]. opposite-polarity [GWBW02]. optimally [EGS^{+}04]. optimum [GGT^{+}02]. Oral [WCGT^{+}00]. orange [SKM03]. Orbit [IdCAS^{+}00, MSL^{+}02, Sre04k, ISS^{+}04, KSW00]. orbit/mast [ISS^{+}04]. orchestrate [Tum04b, WSL^{+}03]. ORD [WHB04]. order [CKW^{+}03, LeB03-59, WSC^{+}03]. Ordered [LP04a]. Ordering [GCN00]. ORF [RMC^{+}02]. Organellar [FRK^{+}01]. Organelle [RH00, Bre03, DSV^{+}03, Del03, Dov01q, EWTW02, ICN^{+}03, SJ01b, VMH^{+}02, Wel01-40]. organelle-bound [EWTW02]. organelle-specific [ICN^{+}03]. organelles [KAC^{+}04, LeB03i, SVT^{+}02, SLD^{+}02, WMG^{+}04, PMB^{+}00]. organismal [EKdM^{+}04]. Organization [DR00a, MMG^{+}04, QK00, DSG04, DM02, DSB^{+}02, GAT^{+}03, HMVG02, KHC02, KR03, KMP02b, MPG^{+}02, MSJ^{+}04, Mis01, OER^{+}02, PRJK01, RRB^{+}03, SH02, SPB^{+}02, SRB^{+}04, SGW^{+}02, TBTN01, TTH^{+}01, TGD^{+}03, WBAS04, WSH^{+}04, BKZ^{+}01, BL01, HHSV00, JPM^{+}00a, MSHG00, SGG01, SGO^{+}00, WBG01, WWKV00]. organize [CC02, Dov02-47, KVM03, TMG03, GG01]. organized [Dov03w, Ono01]. organizer [Dov04a]. organizing [BN02, BPD^{+}04, CBRBM04, GRCF02, GMLM^{+}04, SLG02]. organs [GBN^{+}01]. orientation [CC02, HBB^{+}04, HKBH03, LeB03n, LC01, CHM^{+}01, SCM^{+}00]. Oriented [MEK^{+}04]. origins [LeB03o, LeB03-65, LCSG03, VTH^{+}02, Wel01a, AERD^{+}01, BGFJ01, HLRG01]. Orphan [BMKA01, Mit01]. orthogonal [CPN^{+}01]. orthologue [WLH^{+}04, CGY01a]. Oscillating [Wel03-44]. oscillation [LC01]. oscillations [HTPC04, WDW03, RLC00]. Oscillatory [JRL^{+}03, VZTN03]. oskar [Las03, WHAH03, vEPP^{+}01]. osmosensor [RRS03]. Osmotic [BND^{+}02]. Osp [TWBV^{+}01]. Osp/ [TWBV^{+}01]. ossification [NGS^{+}01, LWS^{+}00]. osteoblast [FAF^{+}04, KPL^{+}02, KHB^{+}04, LTF^{+}01, SCD02, TJS^{+}04, BMKA01, LWS^{+}00, MSV^{+}00]. osteoblasts [AOH^{+}02, LTF^{+}01, MBSB00]. Osteoclast [SHN^{+}01, FNZ^{+}03, KHB^{+}04, SRD^{+}02, ECV^{+}00]. osteoclast-like [SRD^{+}02]. osteoclastic [WXD^{+}03]. Osteoclasts [NLRD01, MNT^{+}03, MKK^{+}00b]. osteogenesis [KRZ^{+}04, JPM^{+}00b]. osteogenic [JBA^{+}01]. osteopenia [HIO^{+}04, KPL^{+}02, LTF^{+}01]. osteopenic [KHB^{+}04]. osteosarcoma [TJS^{+}04]. other [Che02, CGBL^{+}02, LeB03-66, MPBR03]. our [MF01b, CW04]. Outer [HSMJ01, WGvA^{+}01, ABOS^{+}02, BHW^{+}02a, CBW^{+}01, GGB^{+}04, HBV^{+}01, IYT^{+}04, KSS^{+}03, MH01, MR01a, SSK^{+}03, SHKS02, YHB^{+}04, ATE^{+}01, FSY00, FRK^{+}01, KIK^{+}00a, KRR^{+}01, SJ01a, TMHP00]. outfolding [BBP^{+}04]. Outgrowth [HDJ00, WSL^{+}00, HN03, LSMS03, MKM04, MMBB03, SDS^{+}04, TTH^{+}01, FLX^{+}00, GKM^{+}01, MAAZ^{+}00, OKSH00, TNK^{+}00]. outside-in [GLS^{+}03, ISID^{+}03, SKK^{+}02, XBL^{+}03]. Overall [HES00]. overcomes [BHNG01]. overexcited [Wel04x]. Overexpressing [PFM^{+}00]. Overexpression [LTF^{+}01, SKM^{+}02, WFQ^{+}00, HIO^{+}04, PSKK^{+}00, UIK^{+}01]. overhang [LeB03-103]. Overlapping [SLB^{+}01, GCR^{+}03, PNSJ01]. own [Les02c]. oxidase [ESC^{+}01, KWS^{+}02, MNT^{+}03, SHA^{+}03, TNW^{+}02]. oxidase/peroxidase [ESC^{+}01]. Oxidation [Wel04-53, TR02, vRTvdB^{+}00]. Oxidative [TW04, ASK^{+}03, SVT^{+}02]. oxide [MRT^{+}01, RP03a, WST01, GLK^{+}00]. oxygen [CPG^{+}03, RGG03, Wel01-53, Wel03-37, WK04]. oxysterol [LRF^{+}02b].

P [MZR^{+}04, SWG^{+}03, LFT^{+}00, BMLU02, CS00, CKFH00, CRH02, FAT^{+}02, PQF^{+}00, SD00, SM03a, SG00, TFF03, WFQ^{+}00, WE02, YKW^{+}00, LKH^{+}04]. P-binding [WE02]. P-selectin [SM03a, TFF03]. P-Type [CKFH00, CRH02, FAT^{+}02]. P0 [CMW^{+}01, RKF^{+}04, XSK^{+}01]. p101 [BSR^{+}03]. p107 [LRS^{+}02, VFN^{+}04]. p110 [BSR^{+}03, SRKN03]. p115 [CNMS02, LeB02-64, PL01, SBS^{+}02b, Wel01-41, DSSWW00]. p120 [CiKBG03, DIR03, FJK^{+}04, IDvH^{+}02, PLR03, PCB^{+}03, XAB^{+}03, KGvdG^{+}00, NLBK00, TAD^{+}00]. p120-catenin [DIR03]. p120/ [PLR03]. p120catenin [MCA^{+}03]. p130 [LRS^{+}02]. p150 [VMH^{+}02, LVWA01]. p16 [OBG^{+}03, PDW^{+}00]. p160ROCK [CPC^{+}02b]. p190RhoGAP [SAP03]. p21 [CFC^{+}01, WVBY^{+}03, ZLVS02, MCG^{+}00, TYA^{+}02]. p21-activated [CFC^{+}01, WVBY^{+}03, ZLVS02]. p24 [BPKK01]. p27 [TJS^{+}04, MLS^{+}01, SCTM00]. P2X [RDNB02]. P2Y [BKZ^{+}03, ELO^{+}01]. P300 [vMZM^{+}00]. p38 [FYHH00, GLK^{+}00, MTB^{+}02, MP04, MSR04, vdHvODML^{+}00]. p38MAPK [HBSJ04]. p47 [UJL^{+}03, WSWM04b]. p47-mediated [UJK^{+}02, WSWM04b]. p52 [VCDHD03]. p53 [BS04a, BRR^{+}02, DC03, FCM^{+}01, LRB^{+}03, MBM^{+}04, SMS^{+}01a, Stu04, Tum04q, ZMGL02, PCC^{+}00, RIDC01, SME^{+}00, SCMH00]. p53-dependent [DC03]. p53-inducible [BRR^{+}02]. P53-Responsive [SME^{+}00]. p62dok [GSB^{+}01]. P63 [KKL^{+}01]. p75 [PKR^{+}02, YHT02, LS02]. p75NTR [CARG^{+}04, MWAM01]. p75NTR-interacting [CARG^{+}04]. p75NTR-mediated [MWAM01]. p97 [UJK^{+}02, WSWM04b, YMR03]. p97/p47 [UJK^{+}02]. p97/p47-mediated [UJK^{+}02]. Pac1 [LOC03]. Pac11p [FK01]. packages [LeB04w, Wel04-56]. packs [LeB03s]. PAI [BMG^{+}01]. PAI-1 [BMG^{+}01]. pain [LeB02-97]. pair [MN04]. pairing [VLL^{+}03, TSDS00]. PAK [CFB^{+}03, VT04]. PAK1 [SDEZ^{+}03, SPC00]. palate [NH03b]. Palindromic [LeB04-72]. Palladin [PO00]. Pallido [WND^{+}00]. palmitoyl [RFCD02]. Palmitoylation [YKT^{+}04, LeB02-50, EHCC^{+}00]. Pals1 [RML^{+}02]. Pancreas [CBK^{+}00, KVC^{+}03, PRJK01]. pancreatic [HViV^{+}02, LLH^{+}01]. Papillomavirus [SSA^{+}00]. PAR-1 [CBRBM04, LeB04-55]. PAR-3 [ISF^{+}01]. PAR-4 [BMS^{+}03]. PAR-dependent [TKHR03]. PAR1 [MTT^{+}04]. paracellular [BPPFM^{+}03]. Parade [Wel01h, WL04]. Paradoxically [PSKK^{+}00]. parallel [LZS^{+}03, MTM^{+}02]. paralleled [MRT^{+}01]. parallelism [CLM^{+}04]. parameters [SKF02]. paramyosin [LMDS^{+}03]. paramyxovirus [RKJL03]. paranodal [GSP^{+}02b, TGMC^{+}00]. Paranodin [FSGDN^{+}00]. Paranodin/Contactin [FSGDN^{+}00]. Paranodin/Contactin-Associated [FSGDN^{+}00]. paraplegia [ASK^{+}03, EMY^{+}04]. parasite [JR02, KBGG04b]. parasites [Cam03]. parasitophorous [SJ01b]. Parental [MSFH00]. PARP [EKdM^{+}04]. PARP-1 [EKdM^{+}04]. Parsing [Tum04q]. part [BRB^{+}01]. partial [LRD^{+}03]. Participant [SCMH00]. Participate [QK00]. participates [ASS02, BPKR^{+}02, BKS^{+}02, COB01]. Particle [GDHS01, KLK^{+}01, HSMB02, MJG03, PLP02, UHR^{+}03]. particles [CPN^{+}01, IBKSP01]. Partitioning [VYC^{+}00, LMGM^{+}02, PSP^{+}04]. partner [LeB02-73, LeB02-93]. Partners [HvdKDS01]. party [LeB02-62]. pass [MAG^{+}04]. passage [LeB03-71]. passenger [LeB04b, SMLM00]. Passing [Wel03-45]. Passive [SMWG00]. pasts [Dov03-27]. Patches [LeB04-73, HGP^{+}04, YCB04]. path [Dov02-44, Wel04-50]. pathfinding [MMBB03]. pathogens [Dov04a]. Pathological [KWH^{+}00]. Pathology [RKF^{+}04]. Pathway [HGC00, KCL^{+}01, KIO^{+}00, NHB00, RGGL00, RRM^{+}00, SDN^{+}00, ALP^{+}04, AAM^{+}04, ABF^{+}03, BSD^{+}01, BS04a, BvdWD^{+}04, COB01, CK03, CKS^{+}04, CLSK02, EOB^{+}04, FMC^{+}04, Gla01a, GKYY03, HOK^{+}02, HS02, HP03, HBAF^{+}02, LLGB03, LXL^{+}04b, LL02, LO04, MSR04, MZR^{+}04, NOOG^{+}04, OBG^{+}03, OKM^{+}01, PBP^{+}01, PMP^{+}03, PAR^{+}04, PFSG03, PHM^{+}02, RPNM03, SRSW04, SRL^{+}04, SSRX04, TOTC01, TJC^{+}03, TCP^{+}03, TST^{+}03, UCY^{+}02, VBH^{+}02, WHM^{+}02, WYR^{+}02, YJS^{+}03, vdLBK^{+}04, BEG01, Gau00, HFM^{+}01, KRR^{+}01, MSO^{+}00, MTV^{+}00, RBBA00, TCR00, Zwe00, vdVWS^{+}00]. pathway-dependent [COB01]. Pathways [FWP^{+}00, BDR^{+}03, CG03, CSJ03a, CAW^{+}04, CDW^{+}03, HFK^{+}03, IOIF^{+}04, KMH^{+}04, LRWB04, MPAP^{+}03, PBB^{+}04, PWS^{+}01, RMC^{+}02, WAC^{+}03, GHS00, KHK01, LR00, MKK^{+}00b, NKH^{+}00, RIDC01, WRCU00]. pattern [CLG^{+}03, HTPC04, SLT^{+}01]. patterning [ASGL^{+}01, Jeg01c, SDD04]. patterns [DWM03]. PAX4 [BFSO^{+}04]. paxillin [HBSJ04, Sch04, TSY^{+}02, WZB^{+}01a, LWDH01, NT00, PBL^{+}00, SPA^{+}04]. PC [HBSJ04]. PC-12 [HBSJ04]. PC12 [HN03, TEB^{+}03, AEL^{+}00]. PCH [CG03]. PCM [DM02]. PCM-1 [DM02]. PDE5 [LeB03-67, MFF^{+}03]. Pdgf [GBM^{+}00]. Pds5p [HSKG00, SAH^{+}03]. PDZ [AMF01, LM00b, PMU^{+}02, RB01, TNB01, ZCH^{+}01]. Pea [HMRH01]. Pecam [ZN01, JOF^{+}02, LeB02-65]. PECAM-1 [JOF^{+}02, LeB02-65]. peeled [LeB03-64]. Pemphigus [CdBB^{+}01]. peptide [ARLC^{+}04, HGC02, KH04, LXL^{+}04b, MC02, SRW^{+}04, ZMGL02, SPH^{+}00, BHW02b]. Peptides [DDW^{+}01, KRA^{+}01]. perfect [EC03]. perform [LeB02a]. Peri [MMPO^{+}01]. Peri-Golgi [MMPO^{+}01]. pericentric [GBMA04]. Pericentrin [JGR^{+}04, Tum04r, MCBB^{+}04]. pericentrin-like [MCBB^{+}04]. perichondrial [JBA^{+}01]. Pericytes [HGK^{+}01]. Perilipin [SXD^{+}03]. perinuclear [SPO^{+}02]. Peripheral [WFQ^{+}00, CS03a, LCI^{+}01, PSW^{+}02, RKF^{+}04, TSL04, BAD^{+}00, CMW^{+}01, HCR^{+}01, MHW^{+}00, SFV^{+}00]. peripherin [RBD^{+}01, RDN^{+}03]. Periphery [JBN^{+}00]. periplakin [GSN^{+}04, DKMW00]. periplakin-interacting [GSN^{+}04]. Peritoneal [TFF03]. permeability [BPPFM^{+}03, SMWG00]. Permeabilization [WGvA^{+}01, MH01]. Permeabilized [GHS00]. permeable [IKA^{+}03]. permease [UN03, Wel02y, HLK01]. peroxidase [ESC^{+}01]. Peroxide [LeB02-66, JOF^{+}02, CAP00]. Peroxin [SKCS00]. Peroxisomal [TR00, FMJG04, JMG04b, VTGT^{+}03, JMG01, SJS^{+}00, SKCS00, TCR00]. Peroxisome [GKN^{+}03, SSL^{+}00, HvdBP^{+}01, LG02a, MSJ^{+}04, MDT^{+}01, SMC^{+}02, TR04b, VTGT^{+}03, SJS^{+}00, TCR00, vRTvdB^{+}00]. Peroxisomes [Wel01-42, NPL04, Sub02, TNW^{+}02]. Persistence [DPO^{+}04]. persistent [KPL^{+}02, MWL01]. persists [HPE^{+}01]. perspective [MM01]. Perturbation [RRM^{+}00]. perturbations [LeB02-66]. Perturbs [WHS00]. PEX11 [LG02a]. Pex11p [vRTvdB^{+}00]. PEX19 [FMJG04, JMG04b, SJS^{+}00]. Pex19p [SKCS00]. Pex1p [TR00]. PEX3 [FMJG04, SSL^{+}00]. PEX3-Mediated [SSL^{+}00]. Pex6p [TR00]. Pex7p [NPL04]. Ph [AZB^{+}00, CM01, GGT^{+}02, MC02]. pH-dependent [CM01]. pH/Tat [MC02]. phagocyte [ESC^{+}01]. phagocytes [SHA^{+}03, TTM^{+}03]. Phagocytosis [MBS^{+}01, Mel00c, BGR^{+}01, DH02, GSS01, GBY^{+}03, LUB^{+}02, NCGD^{+}03, BTD^{+}00]. phagosomal [FBG^{+}01, MSH^{+}00]. phagosome [Dov04j, FBG^{+}01, HHJS04, VBR^{+}01a, BPS^{+}00, GDK^{+}01]. Phase [CS01, PMBC^{+}00, ARQ^{+}04, BLP^{+}02, BPKR^{+}02, GJS^{+}03b, HPE^{+}01, KAC^{+}04, KRS^{+}02, KGT^{+}02, LCI^{+}01, LCSG03, OSB04, TFM04, MSHG00, TRW^{+}00, SRL^{+}04]. phases [CROfC04, HBG^{+}02, LKLD04, IBKSP01]. PHD [Dov01y]. Phenotype [RGGL00, BJM^{+}02, MGP^{+}02b]. phenotypic [MPG^{+}03]. Pheromone [MMM^{+}04, HW00a]. Pheromone-induced [MMM^{+}04]. phosphatase [DOB^{+}01, GGGN02, JON^{+}03, KvHB^{+}01, LZC^{+}03, LW03, MDQ^{+}03, NOOG^{+}04, NCMO^{+}02, PMGS02, TBTN01, Tum02a, BT00, SMTC00, OGD03]. phosphatases [RDS02, MMC00, PS00]. Phosphate [BPMG00, JMB^{+}04, AHA^{+}04, GGGK03, KKW^{+}03, LGRP^{+}02, LDI^{+}03, PWY^{+}03, TPW^{+}04, TST^{+}03]. phosphates [LeB03-28]. phosphatidylinositide [JHS^{+}02]. phosphatidylinositide-3-OH [JHS^{+}02]. Phosphatidylinositol [AZB^{+}00, BRM^{+}00, BRY^{+}01, MBS^{+}01, PWY^{+}03, YHF^{+}01, AM03, BND^{+}02, FBG^{+}01, KMH^{+}04, KKW^{+}03, LDI^{+}03, MHS01, RWH02, VBR^{+}01a, BRS^{+}01, BTD^{+}00, FMMF01, KNIO01, RyHHK00, MZR^{+}04]. Phosphatidylinositol-4 [BTD^{+}00]. Phosphatidylserine [HdO^{+}01]. phosphodiesterase [MRT^{+}01]. phosphohydrolase [LGRP^{+}02]. Phosphoinositide [FGR^{+}04, IFP^{+}03, PLC^{+}02, BSR^{+}03, XWL03, HBSQ01]. Phosphoinositides [GSS01, DKAH04, SRW^{+}02, XLGS01, HCTM00]. Phospholipase [FYI^{+}03, MSD^{+}04, NLRD01, DAV^{+}03, SC01a, SRW^{+}02, SHA^{+}03, VBH^{+}02]. Phospholipid [TRW^{+}00]. Phospholipids [FRM^{+}02]. phosphopeptide [CRP^{+}04]. Phosphoprotein [IWG^{+}01, KSK^{+}00]. phosphoregulation [SKGC^{+}03]. Phosphorylated [CS01, RBBA00, GAC^{+}03, LBH^{+}02, RGM^{+}02, SKK^{+}04, ZSS01, KHvOD00, TSY^{+}02]. phosphorylates [SSOS01, WMJ^{+}04]. Phosphorylation [ASAJ01, CBLT04, DSSWW00, DBLG02, GJB^{+}00, HBSJ04, LTB^{+}00, LeB03-68, LGW00, MRH^{+}01, MTW^{+}04, NPS^{+}03, PBL^{+}00, RCS^{+}02, RM00, ZBB^{+}00, ATG^{+}03, Che02, FGR^{+}04, GK03a, GHK^{+}03, KHC02, KI04, LKL^{+}03, LWCKL01, LDI^{+}03, MWF02, MHK04, MRT^{+}01, PPA^{+}03, RLZ^{+}03, RDC^{+}04, SMS^{+}04a, SKK^{+}02, SDEZ^{+}03, SDMC^{+}04, UJL^{+}03, VT04, VZTN03, XSK^{+}01, AMEC01, BF01, BLPP01, CSP^{+}00, GLA00, GG01, KWO^{+}00, KO00, RLC00, TGM^{+}01, TYY^{+}00]. Phosphorylation-Dependent [MRH^{+}01]. phosphorylation-dephosphorylation [SKK^{+}02]. phosphotyrosine [COB01]. phosphotyrosine-based [COB01]. Photo [MDJF00]. Photocross [NKM^{+}03]. Photocross-linking [NKM^{+}03]. photolabeling [KAC^{+}04]. photon [IBP^{+}04]. photoreceptor [PBD^{+}02, LM00b]. Photorelease [RRJ^{+}01]. phox [ESC^{+}01, XLGS01]. Physical [TRMI^{+}04, WDS^{+}03]. Physiological [LSK01, QRLL00, SD00]. physiology [SML^{+}04]. PI [BFC^{+}02, HBG^{+}02, LCT^{+}04, OMB^{+}01, SWG^{+}03, WSR03, BMD^{+}00, GTPMU00, LFT^{+}00]. PI-3K [LCT^{+}04]. PI3 [LRWB04, VBH^{+}02]. PI3-K- [VBH^{+}02]. PI3-kinase [LRWB04]. PI3K [FAF^{+}04, HFG^{+}04, SCTF04, JPM^{+}00a]. PI3P [PFSG03]. pick [LeB04-75, MHT^{+}04]. picket [Wel03-40]. Picking [Wel04-54]. picture [RY04]. Pigmented [RTM^{+}01]. pile [LeB02-80]. PilT [LeB02-67]. PINCH [KSC^{+}04]. Pincher [SATA^{+}02]. PINCHing [Tum04s]. Ping [LeB04-74]. Pinning [Wel01-43]. pinocytic [SATA^{+}02]. Pinosomes [LeB02-68]. Pins [Wel04-65]. PIP [LeB02-52, POW^{+}01a, Tum04z, BRM^{+}00, HP00]. PIP5KI [WLW^{+}04]. pipetail [WBT^{+}03]. pit [Dov02c, LeB04d, HvdKDS01]. Pits [SDS00, JFS^{+}03, SM03a]. Pituitary [CAB00]. Pitx2 [HAM01]. Pivotal [SAG^{+}01, Tum04r, GPAS^{+}01]. PKA [HTT^{+}02, KLZ04, RLZ^{+}03]. PKC [HLB^{+}02, KMH^{+}04, AKH^{+}04, BBDM02, LUB^{+}02, LSMS03, NMH^{+}04, NCMO^{+}02, SSP^{+}03, SSO^{+}03, VBH^{+}02, XSK^{+}01, BMD^{+}00, SMS^{+}04a]. PKC- [LUB^{+}02, SSO^{+}03]. PKC-mediated [XSK^{+}01]. PKCs [HTPC04, LeB04-75, EKdM^{+}04]. Pkh [dSAH02]. PKL [WZB^{+}01a]. PLA [ST03]. place [LeB02o, LeB02-47, LeB03c, LeB03r, Ped04]. placental [JKW^{+}03, ToIB^{+}00]. places [Dov02-54, LL03b, LG01]. plain [Dov04k]. Plakin [KYF00]. Plakoglobin [CLAC00, KHLW02, CdBB^{+}01]. plakophilin [BGH^{+}03, GGH^{+}04, HHSV00]. plakophilin-3 [BGH^{+}03]. planar [MYO^{+}04]. Plane [LHvdH00]. plant [HGS^{+}01, RRS03, Wel04-77, Zam04, ARK^{+}00]. plants [LeB03-86, VMO01]. plaque [DLT^{+}02, KDH^{+}04]. Plaques [GTPMU00, HDL^{+}00]. Plasm [KPB^{+}00]. plasma [AM03, And02, ATF^{+}04, BWA^{+}04, CLM^{+}04, GSN^{+}04, KLE^{+}02, LMHJ02, OGD03, PWY^{+}03, PUK02, PWS^{+}01, SHA^{+}03, SZvBuH^{+}04b, UML^{+}03, UN03, VCGB^{+}02, YGWN01, CAB00, GLA00, HW00a, JLS^{+}01, PPM^{+}00, PKF^{+}00, SKR^{+}00, VHLS00, ZGB01]. Plasmalemmal [MBS^{+}01, LFT^{+}00]. plasmamembrane [SUT^{+}01]. plasmepsin [KBGG04b]. Plasmid [Wel02-44, MYC^{+}02, VYC^{+}00]. Plasmin [SMSM00, WSL^{+}00]. Plasmin-Sensitive [SMSM00]. Plasminogen [AKDS00, CACL03, WTG01, WSL^{+}00, BMG^{+}01, KH01]. plasmodesmata [LeB03-71, Zam04]. Plasmodium [GTR^{+}03, KBGG04b]. plasticity [JLK^{+}02, Pro03, FLX^{+}00, TNB01]. Plastid [SCS^{+}00]. plate [SDML04, WK02, GJ00]. platelet [BIC^{+}03, ISID^{+}03, NGK^{+}03, OMiKF02, GBM^{+}00, ZN01]. platelets [CSJ^{+}03b, Dov01p, LTD^{+}01, SD00]. platform [GMLM^{+}04, ZLG00]. platforms [Lai03d]. play [Dov03-38, LeB02g, SR03, SWG^{+}03, VSO^{+}01, MHK^{+}01, SEP^{+}01]. played [YMK^{+}04]. Playing [LeB02-69]. plays [GRCF02, LeB02-53, MCA^{+}03, PSD^{+}04b, ZSS01, GPAS^{+}01, SYH^{+}01, vRTvdB^{+}00]. PLC [ISID^{+}03, KJK^{+}03, SDL^{+}03, MPR^{+}01]. Pleckstrin [RDH^{+}01, KJK^{+}03, FMMF01, RBBA00]. Plexin [SKO04]. PLIC [NB03a]. PLIC-1 [NB03a]. Plk1 [KCWF02, LP04a]. Plod [HAM01]. Plp [EAD^{+}02]. plug [Wel04-58]. pluripotent [BSW^{+}04, GNS^{+}04, SwZK^{+}02]. plus [FVC04, GK03b, LeB03-57, MTM^{+}02, VMH^{+}02, Xia03, vBDH03]. plus-end [GK03b]. Plzf [HCD^{+}00, NMHH03]. Pml [BHBJ00, FNKH02]. PMN [CSIK03]. PNG1 [SPH^{+}00]. pocket [LRB^{+}03]. pockets [Lai03e]. Podosome [CKS^{+}00]. podosomes [BDKM04, Dov03o, ECK^{+}03, Wel04-64, OSN^{+}00]. point [FNFL03, XAB^{+}03, DKJ00, FXPT00]. Pointed [FFSF03, MDF01]. Pointed-end [FFSF03]. points [BHW^{+}03, Wel01-57]. Poised [Wel03-46]. Poking [LeB02-70]. polar [PRS^{+}04, TTS00]. polarisome [BS04b]. Polarity [LeB04-76, LRWB04, MCB00a, Wel03-47, BN02, CBRBM04, Dov01-27, DSB^{+}01, GWBW02, HP04, LKH^{+}04, MYO^{+}04, OKC02, RAS^{+}03, SCTF04, SDMC^{+}04, SWG^{+}03, WSWSL04, Wel02-54, Wel03-45, ZHDB04, MR02, TB00b, WRGK00]. Polarization [Wel04-55, MMM^{+}04, SKN^{+}03, Wel04-71, WSR03, TSMT00]. polarize [KLF04, SB03, DWD^{+}00]. Polarized [Bre03, HGP^{+}00, LW00, Mel00c, AMG^{+}01, AFK^{+}03, ACP^{+}02, BS04b, CK02, Dov03x, HKBH03, KBK^{+}03, MTM^{+}02, NH03a, TNBH01, WKZ^{+}02, Wel04-61, BHM^{+}00, CNBWN00, FPSM01, HSB00, RSG01]. pole [CMM^{+}02, JGW02, Kil03, PMGS02, Wel02h, vEPP^{+}01, WWKV00]. Poles [MHS^{+}00, GK04, HMC^{+}01, GHC01, HHF^{+}00, RCL^{+}00]. poleward [MPB^{+}04]. polices [Del03]. Pollen [LeB04-77, FWY01]. Polo [MRC^{+}02, Wel02-45, DTO^{+}01, NPS^{+}03]. polo-like [NPS^{+}03]. Poly [MAG^{+}04, HVM^{+}00, SGdM^{+}01]. polyalanine [NMG04]. Polyamines [LeB03-69, SR03]. Polyamino [LeB02-71]. Polycomb [PTH^{+}04]. Polycystic [CNBWN00, PDV^{+}00]. polycystin [JGR^{+}04]. polycystin-2 [JGR^{+}04]. polyglutamine [MZH^{+}02, BMS^{+}00b, MMG^{+}01, SSW^{+}01b]. Polyglycylation [XHG^{+}00]. polymer [HRM02]. Polymerase [vMZM^{+}00, KSC02, KMP02b, SGdM^{+}01, SGW^{+}02, AW00, FMP^{+}00, HVM^{+}00, Mar01, PCR^{+}01, VAHV00]. polymerization [BGW^{+}04, GFGP03, KB04, MB01, CBZ^{+}00, HVT^{+}00, LSSL00, YHD^{+}00, Zig00]. polymerizing [STA^{+}01]. Polymorphonuclear [LSA^{+}00]. Polypeptide [MDJF00]. Polypeptides [MMG^{+}01]. polyphemus [SMW^{+}03]. polyphosphate [DOB^{+}01]. Polyploidization [KMiM^{+}01]. Polyploids [LdCK^{+}01]. Polyploidy [MPSM00b]. polyposis [MTM^{+}02, MKST00, RSG01]. polysome [NSLSK02]. Polysomes [PZP^{+}01]. polytopic [CLSK02, NSK04]. Polyubiquitination [HLK01]. pombe [CG03]. pong [LeB04-74]. pool [AM03, EWD02, Wel02i, GVT^{+}00]. pools [Tum04p]. Pop [Dov03-41]. population [ASGGR02, QPDJ^{+}02]. populations [ISS^{+}04, SKF02]. pore [BBBS04, CKZ^{+}02, DBH^{+}01, DMA^{+}01, Dov01v, Dov02j, Dov02-37, Dov02-43, DW02, FGS^{+}02, HDP^{+}01, IKS^{+}02, LeB02-54, LeB04-74, LeB04-78, LMM^{+}02, LG01, OSNG04, RMW03, SAS^{+}02, Tum02e, WPC^{+}02, Wel01-27, ZW04, ATE^{+}01, DSP^{+}01, KRR^{+}01, RAS^{+}00]. Pores [TF00, LeB04-71]. Porin [KRR^{+}01]. portals [CdLvM^{+}04]. position [LMG04, SB03, AOC01, Gil01]. positioning [BNSR03, CPC^{+}02b, Glo04, RRSV02, TKHR03, XTN^{+}02, HLRG01, TMD^{+}01]. Positions [RCL^{+}00]. positive [GBN^{+}01, MGL^{+}00]. possessing [DFJ^{+}02]. Possible [MZ00, LeB02p]. post [EWSN00, PDR^{+}03, BHKL01]. post-endocytic [PDR^{+}03]. post-Golgi [EWSN00, BHKL01]. posterior [vEPP^{+}01, HVB^{+}00]. Postmitochondrial [ZLMP00]. postmitotic [CSP^{+}04, DMO00, GH00]. Postnatal [BCA^{+}03, ACBG04, ZHH^{+}02, TTR00]. Postreplicative [TWS^{+}00a]. postsynaptic [ASYL04, KMLS04, MWF02, EHCC^{+}00]. Posttranscriptional [MÖS^{+}00]. Posttranslational [PR00]. potassium [RPZ^{+}02]. Potential [Dov01z, WGvA^{+}01, DFJ^{+}02, FMF^{+}04, iHGK^{+}02]. potentiates [SMH^{+}02, WCBC04]. Potentiation [pHYpXL00, MZ00, Tum04-30, YpHRL03]. power [Dov01p, LeB03-43]. powers [SMW^{+}03]. PP1c [SSH^{+}04]. PP2A [KSK01]. PP2A-dependent [KSK01]. PPAR [MDT^{+}01, MDT^{+}01]. Prb [PCC^{+}00, CSP^{+}04, LRS^{+}02]. pRb-independent [CSP^{+}04]. pRb-related [LRS^{+}02]. PRC1 [MKJ^{+}02]. Pre [MS00, AOJ^{+}04, XH04, MÖS^{+}00, ONS^{+}00, BCB^{+}02, GSP^{+}02a, SBS02a]. pre-Golgi [XH04]. Pre-Messenger [MS00]. pre-mRNA [AOJ^{+}04, ONS^{+}00, BCB^{+}02, SBS02a]. pre-rRNA [GSP^{+}02a]. preadipocytes [SCD02]. precedes [LeB04-76, PBT^{+}02]. Precise [YKW^{+}00]. precocious [JZ02, IEJ^{+}01]. Precursor [Nob00, SKF^{+}01, ABRA03, CKS02, CM01, CBC^{+}01, EKH^{+}03, SYW^{+}03, VFN^{+}04, GVT^{+}00, SIBG01, SRHV00, TTR00]. precursors [BSW^{+}04, LeB03-46, Wel04-35]. predisposing [YHZ^{+}01]. predominance [AOH^{+}02]. predominantly [CGBL^{+}02, GPZ^{+}02, JMG04b, SZvBuH^{+}04b, SJS^{+}00]. prefer [LeB03-70]. preformed [KCG^{+}03]. Prefusion [NLRD01]. Preimplantation [MSFH00]. preirradiation [MGP^{+}02b]. premature [HPE^{+}01]. Premelanosomes [RTM^{+}01]. Prenylation [AB00]. preparation [CSM04]. prepower [BWW^{+}02]. preprotein [SHKS02]. preproteins [YTM03]. prereplication [ERMT^{+}04]. Prerequisite [SMTC00]. preribosomes [GSP^{+}02a, GNDLS^{+}01]. presence [MBM^{+}04]. Presenilin [EOB^{+}04, KCL^{+}01, SKF^{+}01, CBC^{+}01, WMG^{+}04, KLE^{+}02, LAF^{+}00, MPSM00a]. presenilin-1 [WMG^{+}04, KLE^{+}02, LAF^{+}00]. present [GBN^{+}01]. presentation [KRS^{+}01, MGL^{+}00]. presequence [TVF^{+}03]. Preservation [vARP^{+}00]. preserves [CSP^{+}04]. Presomitic [GBD^{+}00]. pressure [LeB02-65, LeB03m, RRS03]. Pressured [LeB02-72]. presynaptic [KEHAM^{+}02, LeB04-95, MHS01, MDW^{+}04]. pretty [Wel04-44]. Prevent [KY00, BJB^{+}03, GYL02, GGT^{+}02, TTM^{+}03, WBP^{+}03, ZWB04]. Prevents [DGB^{+}00, AKH^{+}04, BSW^{+}04, HPE^{+}01, LeB04s, LeB04-36, DB00, RRL^{+}00]. prickly [Dov03-48]. primacy [KCY^{+}04]. Primary [DGB^{+}00, ITM^{+}04, JGR^{+}04, PMKM03, SGdM^{+}01, Tum04r, WPS^{+}01, ZCH^{+}02, ZMGL02, vRTvdB^{+}00]. primitive [Dov03-55]. primordial [Dov04l, MHT^{+}04]. principles [Vin04]. Prion [STE^{+}01, Dov02-49, MZH^{+}02, PMP^{+}03, Wel02-63]. prion-like [MZH^{+}02]. Prions [LeB03-70, Wel03-48, Wel04-56, Wel04-57]. prior [PBT^{+}02]. PRK2 [CGM^{+}02]. Prm1p [HW00a]. Pro [HOvD^{+}00]. Pro-Apoptotic [HOvD^{+}00]. Pro168 [SKS^{+}04]. Proapoptotic [CQH^{+}00, WAOC^{+}03, APM^{+}02, BRR^{+}02, RMMC01]. probes [YOK^{+}03]. Probing [KMCM00]. problem [Cam03]. procaspase [MWE^{+}03]. procaspase-3 [MWE^{+}03]. proceeds [LRD^{+}03]. Process [MMS00, WGvA^{+}01, ALC^{+}03, DBLG02, HDP^{+}01, JBA^{+}01, vRTvdB^{+}00]. Processed [ESH^{+}01, Dov03w, LSW^{+}03]. processes [WVY^{+}01]. Processing [DMC^{+}03, SGYD^{+}01, SKF^{+}01, CBC^{+}01, EKH^{+}03, GSP^{+}02a, GSSP03, KLE^{+}02, KLD^{+}03, LKL^{+}03, MGRPVAT02, SYW^{+}03, FMP^{+}00, MS00, SRHV00, VAHV00, yZCKA01]. Processive [KB04, TUV00]. processivity [LeB04-79, TB00a]. Procollagen [HAM01]. produces [BBP^{+}04, CKS^{+}00, LMW^{+}00]. producing [DME^{+}04]. Product [WND^{+}00, BSMS03, KB03, RJA^{+}03, RP03b, TBJ^{+}01, MRM^{+}00, SME^{+}00]. production [LWZ03, UGKT^{+}02, Wel04-46]. Products [MBS^{+}01]. Profilin [LeB04-79, YHD^{+}00, KKTP03, SMZ^{+}03, WBM^{+}00]. profiling [SMC^{+}02, CZBH00]. Proform [TOM00]. Progenitor [DME^{+}04, BCA^{+}03, BMS^{+}03, DFJ^{+}02, HAK^{+}04, KVC^{+}03, TAA^{+}02, WCTU02, DLR^{+}01, WRCU00]. Progenitors [WWK^{+}00, Dov03-37, Dov03-38, ZVPK03]. program [Dov02-64, Dov03r, Dov03-42, FBV^{+}04, GCG^{+}01, LCI^{+}01, FCL^{+}00]. programmed [CDM^{+}02, IBS^{+}02, ME00]. Programs [LYMC00]. Progression [DSSY00, CARG^{+}04, HPS^{+}04, MMPO^{+}01, TOTC01, US04, CP01, KR01]. progressive [BXR^{+}02, iKFH^{+}04, SJA^{+}00]. proHB [NMHH03]. proinflammatory [RBD^{+}01]. proliferate [WMG^{+}04]. proliferation [BGM03, BFSO^{+}04, DWB03, FLS^{+}04, HBB^{+}02, KPL^{+}02, LZC^{+}03, LRWB04, MTB^{+}02, MSS^{+}01, MGP^{+}02b, PSD^{+}04a, PBD^{+}04, RAS^{+}03, SEW^{+}01, TPA^{+}03, ZHDB04, CLAC00, FRO01, SFSD00, TWBV^{+}01]. proliferation-dependent [SEW^{+}01]. proliferator [MDT^{+}01]. prolific [Dov03-42]. Proline [RDH^{+}01, WXD^{+}03]. Proline-Rich [RDH^{+}01, WXD^{+}03]. prolonged [LW03]. prometaphase [RTFW02, RCL^{+}00, dEP01]. Promiscuous [LeB02-73]. promise [LeB02-107]. Promote [DDW^{+}01, SPW00, ADL^{+}03, CNHK02, GCH03, KYZ^{+}04, KAiK^{+}02, MSI^{+}03, SAH^{+}03, SC04, TMG03, VCDHD03, WMJ^{+}04, BHKL01]. promoted [ZCW^{+}03]. promoter [CDK04, Dov03u, MWHM01]. Promotes [ESH^{+}01, PFW^{+}00, AKM^{+}04, AQC04, BHNG01, BMY^{+}01, BFSO^{+}04, CHA^{+}01, ELNA^{+}03, EGWK^{+}01, HdO^{+}01, JFS^{+}03, JON^{+}03, LG02a, MSJ^{+}02, MGAL^{+}01, PGSE^{+}01, PDL^{+}03, PPA^{+}03, SDS^{+}04, SZvBuH^{+}04b, SLD^{+}02, WKZ^{+}02, WDL^{+}04, YLG^{+}02, YSC^{+}02, YSC^{+}21, BZSC00, HGB^{+}00, KSN^{+}01, KIC^{+}01, MHE^{+}00, MSH^{+}00, MCB00b, PGS^{+}01, SHHH01, WTG01]. Promoting [GKG^{+}01, LeB03-71, PMBC^{+}00, HPE^{+}01, KSK01, MH02, MRC^{+}02, OBG^{+}03, TJC^{+}03, UIY^{+}01, RSG00, RM00]. Promyelocytic [BHBJ00, WSS^{+}04, BKB^{+}01, HCD^{+}00]. pronecrotic [KWS^{+}02]. proof [Pro03]. propagate [ISS^{+}04]. propagation [BKZ^{+}03, MN04, MB03c]. Propeptide [ARK^{+}00]. proper [HTT^{+}02, LP04a, MCG^{+}03, RRSV02, SWBE^{+}04, STA^{+}01, WDMH03, YK03b]. properties [BMY^{+}01, DFJ^{+}02, Néd02, RFLT02, WDS^{+}03, AKH00, FFKC00, PN00, YYM^{+}01]. prophase [CSM03, RTFW02, SVG^{+}00]. prophase/prometaphase [RTFW02]. Proposed [Hel03]. Proprotein [BTH^{+}03, DMC^{+}03]. Propulsion [TRC^{+}00, BCM04]. Propulsive [BDK^{+}01]. prosper [Wel01-46]. Prostaglandin [HP03]. Prostate [LeB02-74, BMS^{+}03, TKS^{+}02]. prosurvival [WAOC^{+}03]. protease [ASK^{+}03, LeB03-72, LeB04-56, SGC^{+}02, Wel03-50, FWM^{+}01, HOvD^{+}00, NY00]. Proteases [LeB02-75, BMG^{+}01]. proteasomal [FS03, LdVV^{+}02, LBS^{+}02]. proteasome [KL04, ZCW^{+}03, BHKL01]. Proteasomes [HKK^{+}00]. protect [BHL^{+}03]. Protecting [Dov02-50]. protection [GLDM01, JKW^{+}03, LeB04-72, NDM^{+}03]. protective [ALP^{+}04]. Protects [AKDS00, JW04, HGB^{+}00]. Protein [BGS00, DSSY00, EM00, FSGDN^{+}00, FCL^{+}00, FSM^{+}01, GRBD01, HES00, HP00, HSMJ01, IBKSP01, JRW^{+}01, KSN^{+}01, KFS^{+}00, KY00, KLK^{+}01, KSK^{+}00, LeB04-80, LHW^{+}01, LGW00, MMC00, MWC^{+}02, MJY^{+}04, NZHR01, NCMO^{+}02, PSP^{+}04, PSWU00, RTM^{+}01, RB01, RDH^{+}01, RDC^{+}04, SEM^{+}00, SKF^{+}01, SCLC00, SSOS01, TCH^{+}00, Wel04-57, Wel04-58, WFT^{+}01, vMZM^{+}00, AvdWM^{+}01, AKM^{+}04, ASS02, ABRA03, ACMR04, AOJ^{+}04, AKH^{+}04, BBDK^{+}04, BHW^{+}02a, BWV^{+}01, BHW02c, BRR^{+}02, BSL^{+}01, BKD^{+}04, BJ03, BTVB03, CG03, CMM^{+}02, CHM04, CKS02, CPC^{+}02b, CARG^{+}04, CKL^{+}03, CAGK^{+}03, CBC^{+}01, DCC^{+}02, DBB^{+}02a, DRP^{+}03, DLK^{+}02, Dov03n, DSB^{+}01, EKH^{+}03, EWM^{+}04, EGWK^{+}01, FMF^{+}04, FVC04, FHL^{+}03, FBH03, FP02, FTD^{+}01, FGS^{+}02, FGSW03, GBOL03, GTR^{+}03, GKYY03, GJS^{+}03b, GSN^{+}04, GPZ^{+}02, GQI^{+}02, HS02, HLB^{+}02, HPFG03, HvdHG^{+}03, HV03, HBV^{+}01, HEW^{+}01, HMC^{+}01]. protein [HBD^{+}02, HDH02, HIG^{+}01, HKO03, IYT^{+}04, IOIF^{+}04, JPH^{+}01, JZ02, sKCK^{+}01, KCWF02, KMH^{+}04, KSD04, KK02, KYW^{+}04, KOS^{+}04, KSK01, KKTP03, KPKY^{+}03, KGT^{+}02, KMS^{+}04, LLP^{+}02, LRF^{+}02a, LOC03, LTD^{+}02, LRA^{+}02, Les02a, LRF^{+}02b, LP04b, LPT03, LGB^{+}02, MTW^{+}02, MHT^{+}04, MJG03, MHK04, MSJ^{+}02, MCBB^{+}04, MLKH04, MKS^{+}02, MMBB03, MZH^{+}02, MJV^{+}03, MPV^{+}01, MTM^{+}02, MGP^{+}02a, MKJ^{+}02, MPBR03, MB03c, NB03a, NPS^{+}03, NGS^{+}01, OTB03, OER^{+}02, OPP^{+}03, POW^{+}01a, PBD^{+}02, PG02, PPGN^{+}02, PFB^{+}03, PLW^{+}04, PHS^{+}03, Pow01i, PSD^{+}04b, RRB^{+}03, RCS^{+}02, RMC^{+}02, RML^{+}02, RH04, RFCD02, RKF^{+}04, RKJL03, SGK02a, SWBE^{+}04, SYW^{+}03, SRG^{+}04, SP03, SDL^{+}03, SBS^{+}02b, SCP02, SKM^{+}02, SJ01b, SCD02, SHF^{+}03, SSG^{+}02, TONN02, TKB^{+}04, TE01, TVF^{+}03, TR02, TW04, Tum04p, TBW^{+}04, UOB^{+}02, UCY^{+}02, VTH^{+}02, VKB^{+}01, VZTN03, WCIN04, WPC^{+}02]. protein [WMCW03, WHB04, WSF^{+}01, WHAH03, WWS^{+}03, WE02, XTN^{+}02, XB04, YSW02, YLY^{+}02, YpHRL03, YEG01, YHB^{+}04, YCK^{+}03, ZHH^{+}02, ZJM^{+}02, ZCH^{+}01, dMMBK^{+}02, dVKS04, vdLBK^{+}04, AERD^{+}01, AMEC01, ABP^{+}00, ATE^{+}01, ALJ00, AHMJ01, BMM^{+}01, BLC00, BKI^{+}01, BT00, BHBJ00, BC00, BMC00, BHM^{+}00, CdBB^{+}01, CP01, CW01, CGY01a, CDEM00, CFC^{+}00, CAP00, DB00, DBS^{+}01, DSSWW00, DR00a, EHZ^{+}01, EHM^{+}00, FSY00, FMMF01, GDRS01, HK00, HSKG00, HCOC00, HW00a, HKJ00, HCD^{+}00, IdCAS^{+}00, JPRR00, KEGDQ01, KSF^{+}00, KPB^{+}00, LTB^{+}00, LM00b, LGM^{+}01, LYMC00, MPSM00a, MCG^{+}00, Mar01, MAAZ^{+}00, MST^{+}00, MRM^{+}00, MKK^{+}00a, MKST00, MR01b, MSH^{+}00, NHS00, NSW00, NT00, NKH^{+}00, NWT^{+}01, OSMF00, ONS^{+}00, Ono01, OPZ^{+}01, PGS^{+}01, PO00, PVL^{+}00, PHWK^{+}00, PLH^{+}01, RIDC01, RSG01, RMMC01, RCL^{+}00, RBB00, SIBG01, SPS^{+}00, SME^{+}00, SCMH00, SJ01a].

Protein [SBC01, SLT^{+}01, SKT^{+}00, SFV^{+}00, SMLM00, SS00, SEI^{+}00, SMTC00, SYH^{+}01, TYS^{+}00, TGM^{+}01, TN00a, TNB01, TOM00, TSDS00, THZ^{+}01, WBG01, WDLK01, WWJ^{+}00, WAPB^{+}00, WHS00, WRGK00, WWG^{+}00, YSS^{+}01, YKW^{+}00, YF00, ZLG00, vEWS^{+}01, TVF^{+}03, WBP^{+}03]. protein-1 [ACMR04, AKH^{+}04, MKS^{+}02]. Protein-only [Wel04-57]. Proteinase [DZT^{+}00, TOM00]. Proteins [GDHS01, HGP^{+}00, HM00, HLB^{+}00, KSK^{+}00, NHB00, PSWU00, PWU00, YB01, BPKK01, BWK^{+}03, BHW02c, BFH^{+}01, BC04, BCP03, BM02, CPC^{+}02a, Che02, CM01, CSG01, CLSK02, DK04, DM02, Dov02-46, EKdM^{+}04, FMJG04, FJM^{+}04, GI02, GGB^{+}04, GES04, GV03, HFG^{+}04, HIG^{+}01, IYT^{+}04, ISB^{+}04, JMG04b, JGR^{+}04, KNR^{+}04, KAiK^{+}02, LOS^{+}01, LRS^{+}02, LeB02i, LeB03-37, LeB03-52, LeB03-101, LeB04-97, LFM^{+}04, LRB^{+}03, MJS02, MWF02, MMPO^{+}01, MBN^{+}01, MPBR03, MSMK04, NSK04, OSNG04, OMB^{+}01, PSP^{+}04, PDMO03, POH^{+}04, PMP^{+}03, PTH^{+}04, PMU^{+}02, PNSJ01, PMPH03, RKKP02, SS01, SCPP02, SB03, SDL^{+}03, SRW^{+}04, SDC^{+}01, SMH^{+}02, SDL02, TRMI^{+}04, TE01, TCK^{+}03, Tum04f, TEC^{+}03, Val03, VN04, VTGT^{+}03, WYR^{+}02, WAOC^{+}03, WC03, Zam04, vBDH03, ARK^{+}00, BL01, BP00, CB00, CGL^{+}01, HMRH01, JBN^{+}00, JMG01, iNFK^{+}01]. Proteins [PR00, SJS^{+}00, SSN01, SKCS00, SSW^{+}01b, TNM^{+}00, TB00b, Val00, VPP^{+}01, VMD^{+}01, WDLK01]. Proteoglycan [WSL^{+}00, BCA^{+}03, YPN^{+}04, BMF00, LLH^{+}00]. proteoglycans [ADL^{+}03, LS00]. proteolipid [SKM^{+}02, SKT^{+}00]. Proteolysis [GKG^{+}01, HdEV^{+}02, LP04a]. Proteolytic [NMHH03, OKM^{+}01, RRM^{+}00, WSL^{+}00, XRP^{+}01, MGRPVAT02]. proteolytically [LSW^{+}03]. Proteome [GDK^{+}01]. Proteomic [CKZ^{+}02]. protofilaments [ABOS^{+}02]. Proton [HGP^{+}00]. Protrusion [CBZ^{+}00]. protrusions [CLM^{+}04, TWS^{+}04]. proves [Tum03m]. provide [GLDM01, NH03a, PNSJ01]. provides [BSEB04, WHD^{+}03, WWJ^{+}00]. Providing [Mel04b]. Proximal [TKS^{+}02, dVKS04, JAS02, MWM^{+}02]. Prx1 [MSI^{+}03]. PS [HdO^{+}01, HdO^{+}01]. psbC [AZP^{+}02]. PSD [RPZ^{+}02, TMK^{+}00, EHCC^{+}00]. PSD-95 [RPZ^{+}02, TMK^{+}00, EHCC^{+}00]. pseudohypha [WBAS04]. pseudopodia [CK02]. Psgap [RDH^{+}01]. PSGL [TFF03]. PSGL-1 [TFF03]. Psychosine [Mit01, IHN^{+}01, KNM^{+}00]. PTB [HIG^{+}01]. PTB/hnRNPI [HIG^{+}01]. PtdIns [BMLU02, WE02]. PTEN [KvHB^{+}01, PTM^{+}01]. PTP [CWA^{+}03, ZSY^{+}03]. publishing [Hel03]. Pulling [Sre04l, Wel03-49, Wel03-50, Wel04-59, Wel04-60, Tum04-29]. pulls [Wel04-58]. pulmonary [WFI^{+}04]. pump [SUT^{+}01, CGL^{+}01, CS00]. pumps [LeB04m]. punch [LeB03s]. Purification [CK02, SHCM03]. Purified [BHFL01, TTR00]. purinergic [BKZ^{+}03]. Purkinje [FHL^{+}03, JB01, PPK^{+}01]. purloins [MYC^{+}02]. purpose [LeB02-32]. pushed [Wel04p]. pushing [Wel04-27, TMD^{+}01]. Put [Wel02-46, Ped04]. putative [KNR^{+}04, MHT^{+}04, RBE^{+}02, TCP^{+}03]. puts [Dov03m, LeB03c]. Putting [Dov01-27, LeB02-76]. Pyk [SNL^{+}00]. Pyk-2 [SNL^{+}00]. Pyk-2/Raftk [SNL^{+}00]. Pyk2 [KHH^{+}01, SHN^{+}01, UAZG00]. pylori [CAGK^{+}03, LeB03-42].

q [JLJD03]. QSulf1 [ADL^{+}03]. quality [HCC02b, VKB^{+}01, VN04, NSW00]. Quantitation [KAC^{+}04]. Quantitative [LGM^{+}04, MSJ^{+}04]. Quantity [LeB02-77]. quarters [Dov01-34]. question [LB03]. Quick [LeB03-73, LeB04-81, LeB03-77]. quiescence [Wel04-50]. quiescent [SRB^{+}04, BHY^{+}00]. quiet [LeB04-110]. quiets [LeB04-94]. Quo [Mel00c].

R [MKD^{+}01, BMD^{+}00, GBY^{+}03, LUB^{+}02, STP^{+}00, SJB^{+}03, Tum03a]. R-bands [SJB^{+}03]. R-mediated [GBY^{+}03, LUB^{+}02]. R/Met [SBG^{+}04]. rab [DRP^{+}03, AB00, CB00, GAT^{+}03, OMWSN02, PSWU00]. Rab/Ypt [PSWU00]. Rab11 [RRM^{+}03, SDN^{+}00, WJG^{+}00]. Rab13 [KLZ04]. rab2 [SPK^{+}01]. Rab27A [DSH^{+}03, BAM^{+}01, HWHH01, HCR^{+}01, SBM^{+}01]. Rab32 [ASS02]. Rab4 [SDN^{+}00]. Rab5 [FBG^{+}01, NCUJ^{+}00, SDN^{+}00, WWBGG03]. Rab6 [MTG^{+}02]. Rab8 [AFK^{+}03]. Rab9 [BBP02]. Rab9-mediated [BBP02]. Rabenosyn [NCUJ^{+}00]. Rabenosyn-5 [NCUJ^{+}00]. Rabgap [RB01]. Rabies [Gau00]. Rabs [LeB02-78]. Rac [KH01, SWG^{+}03, VBR^{+}01b, ZLR^{+}03, AS00, AQC04, BKD^{+}00, CK02, CTE^{+}04, DSG04, FJK^{+}04, HBB^{+}02, HGC00, HN03, ITF^{+}02, IFP^{+}03, ISB^{+}04, KMK^{+}02, Pow01b, STA^{+}01, TIO^{+}02, WYHP00, WSR03, ZEtK^{+}00]. Rac-based [STA^{+}01]. Rac-dependent [VBR^{+}01b]. Rac-mediated [HBB^{+}02]. Rac-specific [STA^{+}01]. Rac/Cdc42 [HN03]. Rac1 [BMD^{+}00, BZSC00, GGNK04, LCM02, MTW^{+}02, OKSH00, SC01a, WBWS03]. Rac1-dependent [LCM02]. race [Wel04-68]. RacGAP [Tum03o]. Rad [PFW^{+}00, WYR^{+}02]. Rad/Gem/Kir [PFW^{+}00]. RAD50 [FP02]. RAD50/MRE11/NBS1 [FP02]. Rad51 [TWS^{+}00a]. radial [GLS^{+}03, RCY^{+}03, GDRS01, SHS^{+}00, YDRS01]. radiation [Les02b]. Radical [PWC^{+}01]. Radixin [iKFH^{+}04]. Rae1 [BJB^{+}03]. Raf [KFO04, LM00a]. Raf-1 [KFO04, LM00a]. RAF1 [CSM03]. RAF1-activated [CSM03]. raft [ALC^{+}03, GMLM^{+}04, KNR^{+}04, LMGM^{+}02, NL03, PSP^{+}04, SWBE^{+}04, Tum04l, VBR^{+}01b, WSC^{+}01, dMMBK^{+}02, HK00, NKP^{+}01]. raft-associated [KNR^{+}04, SWBE^{+}04]. raft-dependent [NL03]. Rafting [Wel04-61, Dov03-50, Sre04f]. Raftk [SNL^{+}00]. Rafts [LeB03-74, Wel01-44, EKH^{+}03, GPDvH^{+}03, Kur03, Lai03d, LeB04-60, LeB04-96, NDS^{+}02, SBG^{+}04, SKT^{+}03, SNS^{+}04, TNM^{+}03, LFT^{+}00, PKF^{+}00]. Ran [HM00, LeB03-75, FSKS00, HDL02, LeB02i, LHW^{+}01, LGB^{+}02, NM00, RMW03, SHM02, Wel02-65]. Ran- [HDL02]. Ran-Binding [LHW^{+}01]. Ran-dependent [SHM02]. RanBP2 [SKK^{+}04]. randomly [CGBL^{+}02]. RanGAP1 [JTK^{+}02]. RanGAP1*SUMO1 [SKK^{+}04]. ranging [Dov02u]. RanGTP [LWZ03]. Ranvier [BAD^{+}00, GWL03, KS02, LGGS^{+}04]. Rap1 [AQC04, BWA^{+}04, REK^{+}03, RRK^{+}00, SKN^{+}03]. Rapid [KBWG02, MRT^{+}01, NKP^{+}01, SG00, TCH^{+}02, CTE^{+}04, MH01, SH02, SC04, SWE^{+}03]. rapidly [LRD^{+}03]. Rappaport [CW04]. rapsyn [MWF02]. RAR [WCTU02]. RAR-mediated [WCTU02]. ras [RRL^{+}00, BMD^{+}00, CNHK02, DSSY00, FRO01, HGB^{+}00, HFK^{+}03, ITF^{+}02, JLK^{+}02, KSC^{+}04, KY00, KGvdG^{+}00, LeB02-73, LeB02-79, MJY^{+}04, NGKH02, PMPH03, RGGL00, SCTF04, SSRX04, TP01, YITe03, ZEtK^{+}00]. Ras-Induced [RGGL00]. Ras/MAPK [CNHK02]. Ras/Mitogen [DSSY00]. Ras/Mitogen-Activated [DSSY00]. RasGAP [TCZ^{+}03]. RasGAP-associated [TCZ^{+}03]. RasGTPase [GQI^{+}02]. RasGTPase-activating [GQI^{+}02]. rat [BFSO^{+}04, EOJ^{+}03, SLB02, CAB00, DBS^{+}01]. ratchet [Wel02-37]. Rate [SDS00, AMBW04, RCY^{+}03, RS00a]. rates [PWY^{+}03]. Raver1 [HIG^{+}01]. RAW [LUB^{+}02]. RB [Lai03e, OBG^{+}03, HPS^{+}04, Tum03d, Wel01b, Wel02-60]. Rbl [HGC00, MF01a]. Rbl-2h3 [HGC00, MF01a]. RCC1 [LWZ03, NM00]. Rce1 [MHH^{+}03]. Rcy1p [WAPB^{+}00]. Rdna [FAAS00]. reaction [FYI^{+}03, DR00b]. reactivation [BLP^{+}02]. Reactive [CPG^{+}03, RGG03, WK04]. readily [LMHJ02]. reading [LHR04, LCGR00]. Ready [Wel03-51, LeB04-30, Dov01-28]. Real [MPR^{+}01, RDP03, SKGC^{+}03, Wel02-47]. Real-time [RDP03, SKGC^{+}03]. realistic [Mel04b]. rear [Dov01-29, Wel01-44]. rearrangement [HT01, ZSY^{+}03]. rearrangements [CPN^{+}01, HC02, KH01]. reassembly [LGM^{+}04, WSWM04b, DSSWW00, DMO00]. rebinding [BMLU02]. REC8 [EOJ^{+}03]. Recapitulation [HViV^{+}02]. Rec’d [LeB03-76]. Receptor [CSJ00, DLXP00, ESS^{+}00, ELO^{+}01, KSK^{+}00, MBS^{+}01, MZ00, QK00, RM01, ALC^{+}03, AOH^{+}02, AFR01, AR03, AHA^{+}04, ANC^{+}02, BSD^{+}01, BHW^{+}02a, BHW02b, BHK^{+}02, CDK04, CLB^{+}03, COB01, CAGK^{+}03, DDL^{+}04, DFZ^{+}03, FSTC02, FLS^{+}03, GGD^{+}04, GLDM01, GTBM04, GFM^{+}04b, GSB^{+}01, GRSL^{+}04b, HCC02a, HIN^{+}03, HSMB02, HBV^{+}01, HEN^{+}01, HdO^{+}01, HHS03, HBG^{+}02, ICN^{+}03, JMG04b, JLJD03, KVC^{+}03, KYBS03, KPA^{+}03, KSM^{+}01, KJK^{+}03, LS02, LeB02-105, LBS^{+}02, MTW^{+}02, MJG03, MDT^{+}01, NDS^{+}02, OPP^{+}03, PFSG03, PPA^{+}03, PLC^{+}02, REK^{+}03, RLZ^{+}03, RRS03, RDNB02, SJW^{+}04, SHW01, SGK02a, SHB^{+}03, SRG^{+}04, SPO^{+}02, SSH^{+}04, SHKS02, SRW^{+}04, SDS^{+}04, SSG^{+}02, SDD04, SNS^{+}04, SKO04, STJ^{+}01, TIS^{+}01, TCK^{+}03, TST^{+}03, UTH^{+}02, UML^{+}03, WNM^{+}03, WWD03, WAV^{+}04, XWL03, YHT02, YSW02, YTM03, ZHH^{+}02, ZJM^{+}02, ARK^{+}00, BMS^{+}00b, BL01, BMKA01, BF01, CD00, CSD00]. Receptor [DZT^{+}00, FNK^{+}00, FKSJ01, GHS00, GKM^{+}01, HK00, HBL^{+}01, JCR^{+}01, KMB^{+}01, KH01, MGL^{+}00, MBSB00, MPR^{+}01, MSO^{+}00, Rap00, SBZ^{+}00, SJIM01, SSN01, SFSD00, SGPL^{+}00, THE^{+}00, TNB01, WFF^{+}01, WYHP00, YHK^{+}00]. receptor-activated [XWL03]. Receptor-Mediated [CSJ00, MBS^{+}01, QK00]. receptor-targeting [GLDM01]. receptor/ [PLC^{+}02]. receptor/arrestin3 [SGK02a]. Receptors [DFYL00, LeB03-77, AJ01, BPKK01, BKZ^{+}03, CVZ^{+}04, CAW^{+}04, Dov03-35, GMY^{+}03, GGGK03, LeB03-55, LHC^{+}02, MWF02, SDL^{+}03, TPW^{+}04, TFAM^{+}04, WBU03, WKJS^{+}04, YSW02, BC00, CDWB01, KRR^{+}01, MRH^{+}01, Mit01]. recipient [MGP^{+}02b]. Reciprocal [MKK^{+}00b, CAW^{+}04]. reckon [LeB04-38]. Recognition [GDHS01, JKB^{+}03, SNL^{+}00, AFR01, BSEB04, CSG01, HSMB02, HHOP02, MJG03, MM01, PLP02, UHR^{+}03]. recognize [BHW02c]. Recombinant [RPS^{+}02, BHFL01]. recombination [WRSMO^{+}04, WHB04, YK03b]. recombination-dependent [YK03b]. recondensation [MWHM01]. Reconstituting [Tum03k]. Reconstitution [SCP02, XH04, GMC^{+}00, SM00]. reconstruction [HZS^{+}01]. Reconstructs [TCR00]. recruit [GNS^{+}04, GTBM04, TBJ^{+}01]. recruited [FNKH02, SSH^{+}04, NCUJ^{+}00, VAHV00]. recruiting [LO04, RWK^{+}04, PHS^{+}03]. Recruitment [DBB02b, SMTC00, TRC^{+}00, WKJS^{+}04, BBMS03, BSR^{+}03, GHS^{+}03, KKW^{+}03, PUK02, PLC^{+}02, RSK02, SPB^{+}01, SNS^{+}04, VTH^{+}02, WHAH03, ECV^{+}00, GG01, GV00, KCL^{+}00, KHK01, SMSM00, SEP^{+}01, TGM^{+}01, ZKW^{+}00]. Recruits [SCLC00, CG03, Dov01a, KSBE03, OMWSN02, PWY^{+}03, VCDHD03, YITe03, MSH^{+}00, MNHR00]. Rectifier [FLLE^{+}01]. recycle [SC04, MGL^{+}00]. Recycling [ATF^{+}04, CD00, LeB03-78, Mel00c, SDN^{+}00, BET^{+}03, HBG^{+}02, LeB04-90, MTG^{+}02, PDR^{+}03, RRM^{+}03, SPN^{+}04, SPO^{+}02, SV03, HHHJ00, NS00, SPMM^{+}00, WAPB^{+}00]. redirects [PDL^{+}03, OPZ^{+}01]. redistributes [BRB^{+}01, MWF02, WWJ^{+}00]. redistribution [CK02, GMLM^{+}04, THO^{+}04, WSF^{+}01, ZWSC02, TN00b]. Redox [Wel02-48, LC04]. Reduce [CGL^{+}01]. Reduced [KLF^{+}00, PFM^{+}00, TPA^{+}03]. reduces [KWSK^{+}04, BWN^{+}01]. reducing [BPKR^{+}02]. Reductase [CKFH00]. reduction [Wel03a, PN00]. Redundancy [Ste00, YD00]. redundant [EBWC01]. reelin [Wel02-47]. reenter [MWN^{+}04a]. REF1 [GI02]. REF1/ [GI02]. Reflection [SGAS00]. Reflux [HDL^{+}00]. refolding [MLZ^{+}01]. Reformation [PMB^{+}00]. regeneration [CS03a, HYMS^{+}02, PSW^{+}02, LQPC^{+}00]. Region [YW00, BCG03, GSP^{+}02b, GBJ01, RMC^{+}02, STA^{+}01, DR00a, JC01, OLB^{+}00]. Region-Specific [YW00]. regional [CWG^{+}02]. Regions [SPW00, BPD^{+}04, GBJ01, SWB03, TGD^{+}03, WSS^{+}04, CDTW00, DSP^{+}01]. Reglucosylation [TH00]. regular [Bir04d]. Regulate [CK00, GMRS00, SHN^{+}01, ABCK^{+}03, CGF^{+}04, CDE^{+}03, CVZ^{+}04, CFB^{+}03, CROfC04, GK04, GCT02, GMD^{+}02, HTPC04, HPG^{+}02, JLK^{+}02, LKM^{+}04, MKS^{+}02, MWMK04, POW^{+}01a, RPNM03, WBAS04, YGWN01, ZTK^{+}03, BL01, ECO^{+}00, LFT^{+}00, SIBG01, SGO^{+}00, TB00b]. Regulated [CS01, CK00, FWP^{+}00, HW00b, MKK^{+}00a, MCG^{+}03, Rap00, RH00, SEM^{+}00, AM03, BvdWD^{+}04, BLP^{+}02, BBDM02, BKD^{+}04, BHK^{+}02, CDM^{+}02, DPB03, GK03a, GKN^{+}03, IKA^{+}03, ISB^{+}04, JdDD03, LM01, LKL^{+}03, LTD^{+}01, MEV^{+}04, MWHM01, MSMK04, NMH^{+}04, PTH^{+}04, SCPP02, SHKS02, SAP03, SF01, VM02, VMH^{+}02, VCGB^{+}02, ZWAH03, AEL^{+}00, CSP^{+}00, FFKC00, HW00a, KHN00, MST^{+}00, ONM00, RS00a, SBM^{+}01, ZN01, AAM^{+}04, BRY^{+}01]. Regulates [ASAJ01, KIO^{+}00, SNL^{+}00, SZZ^{+}00, SKF^{+}01, WWK^{+}00, WSL^{+}00, ATG^{+}03, AM03, AR03, AFK^{+}03, AAD03, BBMS03, BBDK^{+}04, BGM03, BBR04, BPPFM^{+}03, BGA^{+}04, BWA^{+}04, BBDM02, BTVB03, CMC^{+}02, CWA^{+}03, CK03, CKS^{+}04, CBLT04, CSG01, DBB^{+}02a, DOB^{+}01, FCLSN03, FFSF03, FMC^{+}04, FLS^{+}04, FGSW03, GMRYM^{+}02, GHS^{+}03, GHK^{+}03, GSSP03, GLJP01, GQI^{+}02, JOF^{+}02, KSC^{+}04, KHC02, KRS^{+}01, KLZ04, KS02, LSS^{+}02, LW03, LDI^{+}03, LHR04, MKM04, MZT^{+}03, MDQ^{+}03, MBMMA^{+}03, MTB^{+}02, MSA^{+}03, MTM^{+}03, MNC^{+}03, NB03a, NGK^{+}03, NCMO^{+}02, PWY^{+}03, PFSG03, PLC^{+}02, RDC^{+}04, RLTC^{+}02, RWH02, RAS^{+}03, RDNB02, SBMB^{+}04, SCTF04, SYW^{+}03, SSO^{+}03, SDEZ^{+}03, SGC^{+}02, SAH^{+}03, SKH03, SEW^{+}01, SPA^{+}04, SYVB03, SM04b, SSOS01, TBTN01, TYA^{+}02, TT04, TSL^{+}03, TNBH01, VFN^{+}04, WZB^{+}01a, WAV^{+}04, WLH^{+}04, WDS^{+}03, ZSY^{+}03, ZLVS02, ASMW01, AV01, BLC00, BMKA01]. Regulates [BF01, HAM01, HCR^{+}01, KWO^{+}00, KMS00, LTB^{+}00, MCH^{+}00, NSW00, NT00, NLBK00, PBL^{+}00, PCC^{+}00, SFSD00, TWBV^{+}01, TP01, THZ^{+}01, WGP^{+}00, WJG^{+}00, ZLMP00, Zig00]. regulating [CDK04, HDP^{+}01, PSK^{+}03, RJyH02, RCY^{+}03, SDL^{+}03, SRD^{+}02, SWG^{+}03, TMK^{+}00, TNB01]. Regulation [AZB^{+}00, AJ01, ASGL^{+}01, BS04b, BMS^{+}03, BKI^{+}01, BC04, BKDH01, CKFH00, DAV^{+}03, DFZ^{+}03, FVC04, Gum00, HFM^{+}01, HAP^{+}00, HNK^{+}03, KLD^{+}03, KRU^{+}04, KCL^{+}01, KLGC^{+}01, LPL^{+}04, MTW^{+}02, MWM^{+}02, MR02, MHS01, MNT^{+}03, MHNSM03, ME00, PMGS02, RDH^{+}01, RP03a, SBT^{+}04, SSRX04, SEI^{+}00, TSMS01, TCV^{+}00, WXD^{+}03, WRCU00, WBWS03, XLGS01, vdHvODML^{+}00, ARMB04, BJM^{+}02, BHW^{+}03, BWA^{+}04, CNHK02, CKS02, CBL^{+}02, CMS^{+}02, CFM^{+}02, FCM^{+}01, GTD^{+}02, HFG^{+}04, HS04, HMH^{+}03, HBD^{+}02, ISID^{+}03, KCY^{+}04, LGRP^{+}02, LSK01, LRF^{+}02b, LHC^{+}02, MOMK03, Mil02, MGRPVAT02, PSS^{+}04, SRW^{+}02, SRSW04, SMZ^{+}03, TDL03, TWS^{+}04, VTGT^{+}03, Wel02-48, Wel04-76, WLPD04, ZZM^{+}03, CAB00, CMMP00, FLX^{+}00, HLZW00, JWJJ00, KMB^{+}01, LLH^{+}00, MRH^{+}01, SCTM00, SVG^{+}00, TCS01, TYY^{+}00, VCG^{+}00, WGF^{+}00]. Regulator [BPMG00, HMN^{+}00, AFR01, BJB^{+}03, CS02, FLS^{+}03, FS03, JK01, MTT^{+}04, SNF^{+}02, WMA^{+}04a, WBT^{+}03, YITe03, ZCH^{+}02, BHKL01, HCK^{+}00]. regulators [BS04b, WYR^{+}02]. regulatory [FJM^{+}04, JBA^{+}01, LVD^{+}04, RP03b, RKJL03, TCK^{+}03, HKK^{+}00, JPM^{+}00a]. reinforcement [GJS^{+}03a]. reinhardtii [AZP^{+}02, MN04]. reinitiation [LeB02-76, LHR04]. Rejection [Dov03-43]. Related [EM00, CKW^{+}03, GQI^{+}02, MTW^{+}02, NB03a, OPP^{+}03, WWS^{+}03, BMKA01, CW01, FSK^{+}00, RCL^{+}00, TN00a, WWG^{+}00, sKCK^{+}01, LRS^{+}02, MHT^{+}04, MWMK04, OMB^{+}01]. Relationship [SMI^{+}00, Dov02-65, GYS02, GC04, LeB02-34]. relative [KM01]. Release [pHYpXL00, HGP^{+}00, vARP^{+}00, APBC^{+}02, APM^{+}02, ACP^{+}02, BMLU02, BSMS03, CVZ^{+}04, DC02a, DSH^{+}03, GK03a, LeB02-56, MH01, MEV^{+}04, NMHH03, OKM^{+}01, PBT^{+}02, RDP03, RLZ^{+}03, Riz03, RBW^{+}02, TRMI^{+}04, Wel01y, WWBGG03, DKJ00, FFKC00, HSB00, Lit00, PR00, SRHV00]. released [TR02]. relevant [PSS^{+}04]. relief [LeB02-97]. relocalization [FP02, RME^{+}00]. Relocation [GNH^{+}04]. remaining [BHPN04]. remains [SKK^{+}04]. Remaking [Wel01-45]. remarkable [NDM^{+}03]. remember [Wel04-73]. Remodel [WSDW^{+}00]. Remodeling [EWTW02, LeB03-79, CTE^{+}04, HBC^{+}03, OTY^{+}04, RRM^{+}03, TYA^{+}02, IOLA^{+}00]. remodels [ADL^{+}03]. removal [BHW02c]. Removing [ESH^{+}01]. Ren [GZA^{+}02]. renew [Wel03-27]. renewal [MTM^{+}03, RSD^{+}04]. renewing [SwZK^{+}02]. Reorganization [KRS^{+}01, PFW^{+}00, RTFW02, HKP^{+}04, MMBM04, SSO^{+}03, SBC^{+}03, VBR^{+}01b, YSK^{+}04, DBS^{+}01, GKM^{+}01]. reorganize [Dov03-60]. Repair [Wel01-46, DC03, DDV^{+}03, EKdM^{+}04, Pro03, RDC^{+}04, WC03, SBB00]. repaired [LeB03-76]. Repeat [SBH00, TONN02, BMF00, KIK^{+}00b, NHS00, TN00a, ZW04]. repeats [CFWH^{+}01, GBMA04, JMG^{+}04a, KPKY^{+}03, STJ^{+}01, TDhL^{+}02, HCD^{+}00]. replacement [RGM^{+}02]. replicate [Wel03-51]. Replicating [Wel03-52]. Replication [BGFJ01, CME^{+}02, CH03, DC03, FHJW^{+}01, FP02, LeB02y, LeB03o, LCI^{+}01, LCSG03, MEFC03, Pow01d, SCK04, SVLM02, SM03b, SK01, Tum03h, VTH^{+}02, YDPK04, AERD^{+}01, AH01, CGY01a, DB00, DE01, FKSJ01, Gil01, HLRG01, LRW^{+}00, MCG^{+}00, MSHG00, SMS00]. replisome [MN03]. Repopulating [Tum04t]. reporter [VZTN03]. repositioning [LCI^{+}01]. Repress [YCX^{+}01]. Represses [WND^{+}00, BvdWD^{+}04]. Repressing [Wel02-49]. Repression [JLJD03, CARG^{+}04, Kel03, Tum04v, WCTU02, WBU03, WWD^{+}04, WHAH03, SKJ^{+}00]. repressive [PTH^{+}04]. repressor [GCG^{+}01]. Reproduction [MKA^{+}00]. Reprogrammed [Wel03-53]. Reprogramming [Wel04-62]. repulsion [BHNG01, DRBF03]. repulsive [SSW^{+}01a]. require [LdCK^{+}01, SML^{+}04, VBH^{+}02, WKJS^{+}04, CFC^{+}00, GLA00]. Required [BYLA^{+}01, ELO^{+}01, FSGDN^{+}00, HSMJ01, KLK^{+}01, AM03, AZ03, AZP^{+}02, BWV^{+}01, BDKM04, BMLU02, BCB^{+}02, BBSF01, BKD^{+}04, CSG^{+}04, CMM^{+}02, CNH^{+}02, CPC^{+}02b, CRS^{+}03, CSM03, DMC^{+}03, DSV^{+}03, DH02, DRP^{+}03, DMLK04, EWTW02, EWM^{+}04, EKC^{+}03, FAT^{+}02, FP02, FNFL03, FYI^{+}03, FKH^{+}04, FCBH01, GSP^{+}02a, GCO^{+}04, GYL02, GC04, GAT^{+}03, GMD^{+}02, GPL^{+}02, GJS^{+}03b, GGGN02, HKHO01, HTT^{+}02, HPS^{+}04, HDH02, JGR^{+}04, KJY04, KHLW02, KSK^{+}02, LLP^{+}02, LCG^{+}04, LBS^{+}02, MB03a, MC01, MPBR03, MSD^{+}04, MJY^{+}04, NPS^{+}03, NGS^{+}01, NDS^{+}02, iONOM02, OTY^{+}04, PLR03, PMKV01, PRJK01, RMW03, SWBE^{+}04, SWH^{+}02, SRC^{+}01, SLG02, ST03, SDMC^{+}04, SDC^{+}01, SVLM02, SDL02, TFM04, TEC^{+}03, TBW^{+}04, UJK^{+}02, UN03, UOB^{+}02, VTH^{+}02, VKB^{+}01, VBR^{+}01b, WVY^{+}01, WRSMO^{+}04, WSC^{+}01, WHB04, WDMH03, WHAH03, WLWB01, WSH^{+}04, XRP^{+}01, YHB^{+}04, YK03b]. required [ZH04, ZWB04, dSAH02, BMC00, BHM^{+}00, CDTW00, CDEM00, CSP^{+}00, DZT^{+}00, DSSWW00, FSY00, FSKS00, FHP00, FKSJ01, GG01, GH00, HSKG00, HMAM01, HW00b, KTY^{+}00, KR01, LM00b, LHvdH00, MF01a, MST^{+}00, MHW^{+}00, MR01b, MLS^{+}01, NKH^{+}00, Ono01, PDV^{+}00, PWU00, RMG^{+}00, SJS^{+}00, SJ01a, SBC01, SFV^{+}00, SBM^{+}01, THE^{+}00, TSDS00, WJG^{+}00, WWG^{+}00, YCX^{+}01, YMM^{+}00, ZGB01, vEWS^{+}01]. Requirement [GGH^{+}04, HT01, HVB^{+}00, IOLA^{+}00, SYY^{+}03, WCTU02, AMF01, HMVG02, HG03, AKW00]. Requirements [HES00, HRE^{+}01, CS03b, HBK^{+}02, HLB^{+}02, LGB^{+}02, MB03c, RWSV03, SGK^{+}02b, CB00, TH00]. Requires [AZB^{+}00, CKFH00, PSWU00, WKS^{+}00, ACMR04, BND^{+}02, Che02, DB02, EES^{+}01, FJK^{+}04, GBOL03, GSP^{+}02b, GNH^{+}04, HC02, KKK^{+}02, LZC^{+}03, LeB02l, MWF02, MWMK04, MMG^{+}04, NK02, PC01, RZB^{+}03, RRM^{+}03, RTFW02, Sea04, SWE^{+}03, VT04, VCGB^{+}02, WSK^{+}03, WHP^{+}02, BP00, BLPP01, DOL^{+}01, FRO01, FLLE^{+}01, GSB^{+}00, GHC01, HVT^{+}00, KHK01, SCM^{+}00, TR00, YKW^{+}00]. Requiring [MMS00]. Requisite [HLU00]. Rer1p [SSN01]. rereplication [MBM^{+}04]. rescue [Dov02n, KAiK^{+}02, Wel04i, Ern00]. Rescues [KHH^{+}01]. resealing [CKF^{+}03]. Resemble [HWW01]. Reservoir [LW00]. Residence [PVL^{+}00]. resident [BTH^{+}03, LOS^{+}01]. Residents [OAR^{+}00]. Resides [SDDS00]. Residual [TBJ^{+}01]. residue [CRP^{+}04]. residues [KNI^{+}04, SAS^{+}02, PBL^{+}00]. resistance [GLDM01, MO01, CWMO00]. resistant [ARMB04, NK02, WPM^{+}00]. resolution [HWBD^{+}01, WPS^{+}01, YK03b, MR01b]. resolves [Wel03m]. resonance [SP03, SN04]. resonant [CWG^{+}02]. resorption [DDL^{+}04]. respect [CGBL^{+}02]. Respective [PMK^{+}00]. Respiratory [MMWC00]. respond [RRS03]. Response [HRE^{+}01, NZHR01, SNL^{+}00, APM^{+}02, BMS^{+}03, CVZ^{+}04, CAGK^{+}03, FNZ^{+}03, GCT02, HTS02, JRL^{+}03, Jay01, KMG^{+}03, LRB^{+}03, LHR04, MSD^{+}04, MRT^{+}01, NMH^{+}04, RH04, SPB^{+}02, SR03, SVLM02, UCY^{+}02, Wel04s, WWD03, WDS^{+}03, BKI^{+}01, HSB00, NSW00, SCMH00, WTG01]. Responses [SGF^{+}00, vEWS^{+}01]. responsible [GBJ01, HDL02, JAS02, KI04, STA^{+}01, BDK^{+}01]. responsive [ZFH^{+}04, KEGDQ01, SME^{+}00]. responsiveness [PUK02]. Restore [BMD^{+}00]. restores [OWW02, BWN^{+}01, TTP^{+}01]. Restricted [MBS^{+}01, GSP^{+}02a]. restricting [SSRX04]. Restriction [JLS^{+}01, GHK^{+}03]. restrictive [HvdHG^{+}03]. restructures [CSM04]. Restructuring [GTPMU00]. result [MWSL^{+}03]. results [KSNS^{+}04, LeB04-81, MKR01, NMG04, FSK^{+}00, HDL^{+}00, TRW^{+}00]. Ret [GSB^{+}01, SRG^{+}04]. retained [SHE^{+}02]. Retains [WMT^{+}01]. Retention [GSP^{+}02b, BPKK01, VKB^{+}01, LM00b, yZCKA01]. reticulum [BBG^{+}03, BSMS03, EWM^{+}04, FAT^{+}02, HKE^{+}04, HIT^{+}02, JCPWS01, KOS^{+}04, MJG03, NSLSK02, PDMO03, SRB^{+}04, VKB^{+}01, WH03, WWGK02, ZLH^{+}03, AFB^{+}01, CGL^{+}01, DR00b, GSB^{+}00, KSF^{+}00, KKL^{+}01, MNHR00, NBWB^{+}00, NSW00, iNFK^{+}01, OPZ^{+}01, PVL^{+}00, PR00, PGV^{+}00, SSN01, TOM00, TH00, WGP^{+}00]. retinal [SSW^{+}01a]. retinoblastoma [JZ02, LRB^{+}03, MCG^{+}00]. Retinoic [WK02, JBA^{+}01, SCD02]. retinoid [WBU03, WRCU00]. Retinosomes [IGP04]. retinyl [IBP^{+}04]. Retraction [Dov04m, LMDS^{+}04, GYL02, MB03a, WLWB01]. retrieval [Sea04, VKB^{+}01, SSN01]. retrograde [GSB^{+}03, LS02, LDK^{+}03, MMPO^{+}01, NYT^{+}03, SDL02]. retromer [AHA^{+}04, Sea04]. retrotranslocation [YMR03, iNFK^{+}01]. retrotransposons [RKKP02]. Retroviral [HRE^{+}01]. retrovirus [LMVW03]. Rev [HRE^{+}01, WSW^{+}00]. reveal [CTE^{+}04, ISS^{+}04, Néd02, XSK^{+}01, SJA^{+}00]. revealed [AAM^{+}04, DME^{+}04, HZS^{+}01, MPV^{+}01, SKGC^{+}03, Sil02, TNMM03]. reveals [ARMB04, ACP^{+}02, BLU^{+}04, BBBS04, CWG^{+}02, CK02, Dov01h, HdEV^{+}02, HCL^{+}03, IBP^{+}04, LS02, LL02, LPPT^{+}02, MHT^{+}04, MSJ^{+}04, RGM^{+}02, SAWS02, SGK^{+}02b, VZTN03, WCIN04, WCA^{+}03, WPS^{+}01, WC03, CMMP00, MÖS^{+}00, PPM^{+}00, WCGT^{+}00]. Reversal [Bre00, RGGL00]. Reverse [GMZ^{+}00]. Reversed [KBK^{+}03]. Reversible [KIO^{+}00, MGP^{+}02b, MC02, AML00, TV00]. Revertant [LMW^{+}00]. reverting [KvHB^{+}01]. revolution [Wel02-69]. Rewiring [Wel04-63]. rewrapped [LeB03-64]. RGB [RO04]. Rgd [ELO^{+}01]. RGS [SDL^{+}03]. RGT [XBL^{+}03]. rhabdomere [ZH04]. Rho [ABF^{+}03, DWM03, KKA^{+}01, Rid00, TYY^{+}00, Wel04-64, BHNG01, BPPFM^{+}03, BDKM04, CGM^{+}02, CMC^{+}02, CPC^{+}02b, DSG04, DBB^{+}02a, DPB03, Dov01l, DWM03, LMGM^{+}02, LeB02-81, LeB03-83, LeB04-82, LKH^{+}04, MOMK03, MSM^{+}01, NLBK00, TYA^{+}02, TIO^{+}02, WBC^{+}00, WYR^{+}02, WMS00, WW02, WSR03, YHT02, YOK^{+}03, YEG01, ZEtK^{+}00]. Rho-associated [CPC^{+}02b]. Rho-dependent [BHNG01, TIO^{+}02]. Rho-family [YOK^{+}03]. Rho-Kinase [KKA^{+}01, TYA^{+}02]. Rho/PRK2 [CGM^{+}02]. Rho1p [AQHO03, MSJ^{+}04]. Rhoa [ONM00, BJM^{+}02, DSB^{+}02, Dov01-29, FJK^{+}04, MB03a, SMZ^{+}03, SKO04, TSY^{+}02, WLWB01, BKD^{+}00, DSSY00]. RhoA/ROCK [SMZ^{+}03]. Rhodopsin [TMHP00]. Rhogap [RDH^{+}01, Tum03o]. RhoGEF [BHNG01, SKO04, TIS^{+}01]. RhoGEF-mediated [SKO04]. RI [WPS^{+}01, WPO00]. ribbon [DRP^{+}03]. Ribonucleoprotein [KJB^{+}02, PZP^{+}01, WMHB^{+}00]. Ribonucleoprotein-dependent [KJB^{+}02]. Ribonucleoproteins [PP00]. Ribose [HVM^{+}00, SGdM^{+}01, BKZ^{+}03]. Ribosomal [KMP^{+}02a, HKJ00, KKK^{+}00, SRHV00, VAHV00]. ribosome [HSMB02, MJG03, PLP02, UHR^{+}03, ASP^{+}00, CH01, CMMP00, MDJF00, RSG00]. Ribosomes [LeB02-80, AR00]. Ribosylation [MZ00, NCGD^{+}03, BLC00, VCGB^{+}02]. Rich [RDH^{+}01, KPKY^{+}03, RFCD02, SJB^{+}03, SGC^{+}02, WXD^{+}03, BKZ^{+}01, BMF00, IAG^{+}00, SKT^{+}00, SML^{+}04]. ride [Dov04a, LeB04-73]. ridges [ABOS^{+}02]. Riding [LeB03-80]. Rif1 [XB04]. right [LeB02o, LeB02-61, LeB04z, LeB04-63]. rigidity [MB03a]. RII [MWC^{+}02]. Rik1 [TBW^{+}04]. RIM1 [TRMI^{+}04]. Rin [HN03]. Ring [LeB03-81, AKH^{+}04, CG03, DSG04, Dov02-40, FJM^{+}04, GBZ^{+}02, TMG03, VLL^{+}03, WMCW03, Yum01, BKD^{+}04, MKS^{+}02, GMC^{+}00, MDJF00, PZP^{+}01, SEI^{+}00, VMO01, ZKW^{+}00]. rings [BPC03, MB01, WXD^{+}03, Wel01-67, WSW^{+}00]. RIP [KSR^{+}04]. Rise [SRHV00]. risk [Les02b]. Rlp7p [GSP^{+}02a]. Rme [CHM04]. Rme-8 [CHM04]. RNA [AOJ^{+}04, AW00, BHBJ00, BM02, BS02, CPC^{+}02a, CNJ01, DJ03, FMP^{+}00, FM01, GDHS01, JBK04, JPM^{+}00b, Kel03, KLK^{+}01, KSC02, KPB^{+}00, KKK^{+}00, KYM04, KMP02b, Las03, LeB03-82, LeB04-83, LGM^{+}01, Mar01, MS00, PCR^{+}01, PLP02, SRHV00, SGW^{+}02, VAHV00, vMZM^{+}00]. RNA-binding [DJ03, AOJ^{+}04]. RNA-editing [DJ03]. RNA-mediated [FM01]. RNAi [Tum04u, Wel02-50]. RNAi-mediated [Tum04u]. RNAs [GS02, HRE^{+}01, MAG^{+}04]. RNF [BKD^{+}04]. RNF-5 [BKD^{+}04]. Rng3p [LP04b]. RNP [KYM04, SRKN03]. Rnq1 [MZH^{+}02]. Ro11 [EM00]. road [RT02, AR00]. roam [MAG^{+}04]. Robo [BHNG01]. Robust [WSWSL04]. ROCK [LeB02-81, LeB02-82, LeB03-83, LKH^{+}04, SBT^{+}04, SMZ^{+}03, TWS^{+}04, TIO^{+}02, WDS^{+}03, TYY^{+}00]. ROCK-generated [WDS^{+}03]. ROCK-steady [LeB02-82]. Rocking [Wel04-65]. rod [LeB03-87, SGG01]. Role [AFR01, AHA^{+}04, ALWR01, CP01, CBZ^{+}00, EWSN00, EKT^{+}00, FBG^{+}01, FMMF01, JPM^{+}00a, KGC^{+}00, KGvdG^{+}00, KJK^{+}03, MAAZ^{+}00, NSK04, OMB^{+}01, ORZ^{+}04, RSG00, SJIM01, SBG^{+}04, TDFV02, TWS^{+}00b, WXQ^{+}00, WWBGG03, XTN^{+}02, ANC^{+}02, BHW^{+}02a, BGR^{+}01, CLSK02, CH03, DCC^{+}02, DRC^{+}02, DWM03, DBS^{+}03, DWFAdL^{+}02, FGS^{+}02, GRCF02, GCT^{+}04, GBY^{+}03, HJ02, HCL^{+}03, HGS^{+}01, HCCB03, HvdBP^{+}01, HMC^{+}01, HWBD^{+}01, IUK04, IDvH^{+}02, KBG^{+}03, KOS^{+}04, KR03, KFR^{+}04, LUB^{+}02, LM01, LOC03, LSMS^{+}01, MSL^{+}02, MHT^{+}04, MSJ^{+}04, MDW^{+}04, MCA^{+}03, OSB04, OKM^{+}01, OMiKF02, PSD^{+}04b, Pro03, SCPP02, SRW^{+}02, SAG^{+}01, SR03, TM04, VSO^{+}01, VMH^{+}02, WLW^{+}04, Wel02-47, Xia03, XSK^{+}01, ZSS01, ZJM^{+}02, ZRDG02, AKT01, BRS^{+}01, CdBB^{+}01, CMW^{+}01, Coo01, EU00, GPAS^{+}01, GVT^{+}00, HvdKDS01, HBSQ01, IHN^{+}01, JSCR01, JC01, KWH^{+}00, KKK^{+}00, MKA^{+}00, MKK^{+}00b, MHK^{+}01]. Role [OSMF00, PMB^{+}00, QRLL00, RRB^{+}01, Rut00, SBZ^{+}00, SMS^{+}01b, SYH^{+}01, TCS01, WCGT^{+}00, ZKW^{+}00, vRTvdB^{+}00]. Roles [BN02, BSR^{+}03, CSD00, KDH^{+}04, LKH^{+}04, YMK^{+}04, DLY^{+}02, DJ03, GV03, MSMK04, QS02, RJyH02, SWK^{+}04, SWG^{+}03, TWS^{+}04, VBR^{+}01a, XBL^{+}03, YCK^{+}03, AMEC01, BMM^{+}01, FLX^{+}00, HLZW00, SEP^{+}01, TYY^{+}00, WGF^{+}00]. rolling [DKA01, SM03a, YLQ^{+}02]. Romeo [Wel01-47]. Ron [LeB03-43]. room [Wel01-32]. root [SLB02]. rootlet [YLY^{+}02]. Rootletin [YLY^{+}02]. roots [Tum03l]. ROP [FWY01]. ROP2 [SJ01b]. ROS [Tum03l, KMB^{+}01]. rotation [KRMB03, MN04]. Rotationally [SLT^{+}01]. rough [Dov02-27]. rounding [MB03a]. route [Dov03-40, LeB04-56, Sub02, BP00]. RPTP [vWJK^{+}03]. RPTP- [vWJK^{+}03]. Rre [HRE^{+}01]. Rrna [FMP^{+}00, SGYD^{+}01, GSP^{+}02a, GNDLS^{+}01]. Rsp5 [DKAH04]. Rsp5p [HCC02b]. Rsp5p-dependent [HCC02b]. Rubral [WND^{+}00]. Rud3p [GTBM04]. Ruffling [YHF^{+}01]. Rule [Wel01-28]. run [LeB03-60, Wel02-65, TUV00]. Running [LeB04-84, Pow01h]. runx2 [TJS^{+}04, FAF^{+}04, ZZM^{+}03]. rush [Wel04l]. ryanodine [GMY^{+}03, RLZ^{+}03, MRH^{+}01].

S [MSP^{+}01, Wel03-54, ARQ^{+}04, BLP^{+}02, BPKR^{+}02, GJS^{+}03b, HPE^{+}01, HBB^{+}04, KRS^{+}02, LKLD04, OSB04, PSKK^{+}00, RP03a, SRL^{+}04]. S-Nitrosylation [MSP^{+}01, RP03a]. S-phase [SRL^{+}04]. S100A10 [BGA^{+}04]. S100C [SMT^{+}03, SMS^{+}04a, SMI^{+}00]. S100C/ [SMT^{+}03]. S100C/A11 [SMS^{+}04a]. S2 [RWSV03]. S6 [JHS^{+}02]. S6K [Man04a]. Sa [LYKH00]. Sa/Scc3p [LYKH00]. Saccharomyces [AQHO03, AC00, BT00, EKC^{+}03, FBV^{+}04, HSKG00, HvdBP^{+}01, HGP^{+}04, HKBH03, JRL^{+}03, JC01, KKPB03, KNIO01, KSK01, LVD^{+}04, MB03c, NPL04, PLH^{+}01, PGV^{+}00, RRB^{+}01, SKR^{+}00, SL01, STE^{+}01, VYC^{+}00, VTGT^{+}03, WKZ^{+}02, WAPB^{+}00, WK01, YF00, vRTvdB^{+}00]. Saccular [PPM^{+}00]. SadA [FSTC02]. safe [Pow01d]. Salmonella [HPFG03, MSH^{+}00]. Salt [Wel02-51]. SAP97 [TMK^{+}00]. Sapk [KCL^{+}01]. Sapk/Jnk [KCL^{+}01]. Sar1 [AFB^{+}01, HWBD^{+}01]. Sar1-GDP [HWBD^{+}01]. Sar1p [FS03]. Sar1p/COPII [FS03]. Sarcoglycan [TCH^{+}00]. Sarcolemma [RPE00]. sarcolemmal [AMF01]. Sarcomere [HDJ00, YEG01, RBV00]. sarcomeric [MKS^{+}02, YEG01, BMM^{+}01, HSW00]. sarcoplasmic [BBG^{+}03]. SART3 [SRKN03]. SART3/p110 [SRKN03]. Satellite [CSL^{+}03, GBMA04, JMG^{+}04a, MTM^{+}03, NDM^{+}03, RDNB02, ZGN^{+}04, BHY^{+}00]. saturable [NGKH02]. Say [LeB03-84, LL03a]. Saying [Dov01-30]. says [LeB02-109, LeB03-67]. SC-35 [SJB^{+}03]. SC1 [CARG^{+}04]. Sca [TTP^{+}01]. Sca-1 [TTP^{+}01]. Scaffold [WFT^{+}01, CK02, DWFAdL^{+}02, LeB04-94, OGD03, SRG^{+}04]. scaffolding [GBY^{+}03, VMD^{+}01]. Scaffolds [LeB04-85, LeB03-34]. scale [MWHM01, SWB03, YHZ^{+}01]. SCAR [ZCH^{+}02]. scaRNAs [JBK04]. Scc3p [LYKH00]. scenic [Dov03-40]. Schizosaccharomyces [CG03]. Schwann [Dov03-44, YKW^{+}00, CS03a, Dov04i, FPP^{+}02, GVT^{+}00, GWL03, PBD^{+}04, PMU^{+}02, PSW^{+}02, SRKR00, ZEW^{+}01]. Sciatic [AKDS00]. Scission [HvdKDS01]. sclerosis [JHS^{+}02]. Scotin [BRR^{+}02]. Scrambling [TRW^{+}00]. Scribble [ZHDB04]. SDF [ABF^{+}03]. SDF-1 [ABF^{+}03]. sea [HV03, RCL^{+}00]. seals [LeB02-99]. seams [LeB02w]. search [Wel03-37]. Searching [KC02, LeB03-85, Wel01-48]. Sec1 [AHMJ01]. Sec14p [LRF^{+}02b]. Sec14p-dependent [LRF^{+}02b]. Sec16p [LeB02z, SMH^{+}02]. Sec1p [BJ03, SVI^{+}04]. Sec1p/Munc18 [BJ03]. Sec24p [SKR^{+}00]. Sec2p [EWSN00, OMWSN02]. Sec34 [SDL02]. Sec34/Sec35p [SDL02]. Sec35p [SDL02]. Sec3p [BHPN04]. Sec4p [GAT^{+}03, OMWSN02]. Sec6 [SZZ^{+}00, YGWN01]. Sec6/ [YGWN01]. Sec6/8 [SZZ^{+}00]. Sec61 [RSG00, RSG00]. Sec61p [SHWH00]. secondary [WPS^{+}01]. secretase [CBC^{+}01, GRSL^{+}04b, HF03, SYW^{+}03]. secretases [KH04]. Secreted [HGP^{+}00]. secretion [CGF^{+}04, Cor02, Dov02-55, HDP^{+}01, OKC02, TEB^{+}03, CAB00, SBM^{+}01]. Secretory [HGC00, JR02, RVB^{+}01, SCB02, AQHO03, DSH^{+}03, HBAF^{+}02, MJS02, MPV^{+}01, MPBR03, OMWSN02, JLS^{+}01, NSW00]. Secures [GZY^{+}00]. Securin [LeB02-83, HdEV^{+}02]. Secyeg [KM00]. Sed5p [PG02]. seedlings [TOM01]. Seeds [TOM00]. Seeing [LeB02-84, LeB02-85]. seek [Wel03y]. Seeking [SL00]. seeks [Dov02-65]. sees [LeB03-57]. segment [FMF^{+}04, RKJL03, DAC00, KKL^{+}01, TMHP00]. segments [JB01, KS02, LGGS^{+}04, BAD^{+}00, RSG00]. Segregates [KPB^{+}00]. segregation [Bre03, FMF^{+}04, GKYY03, sKCK^{+}01, MSM04b, Wel02-44, Wel02-62, WDMH03, AMEC01, BSW^{+}00, LHvdH00, WK01]. Seizure [WSL^{+}00]. Seizure-Induced [WSL^{+}00]. Selectin [LeB02-86, DKA01, SMR^{+}02, YLQ^{+}02, DLR^{+}01, SD00, SG00, DKA01, DSM^{+}03, SM03a, TFF03, YWW^{+}04]. selectin-mediated [YLQ^{+}02]. Selecting [Dov02-51]. Selectins [Wel03-55]. Selection [AFB^{+}01]. Selective [BSW^{+}04, GFM^{+}04b, HK00, LCM00, SLB02, TAD^{+}00, ZMGL02, FMF^{+}04, NHG^{+}03, Sea04, SRW^{+}04, WPC^{+}02, dVKS04, BP00, KKS^{+}01, Zwe00]. Selectively [HGP^{+}00, AFK^{+}03, PBB^{+}04]. Self [Dov01-31, LeB02-30, LeB03-56, LeB04-36, LeB04-77, MTM^{+}03, Mis01, RSD^{+}04, SH02, SwZK^{+}02]. self-awareness [LeB03-56]. self-destruction [LeB04-36]. self-organization [Mis01, SH02]. self-renewal [MTM^{+}03, RSD^{+}04]. self-renewing [SwZK^{+}02]. Self-stimulation [Dov01-31]. Semaphorin [EWD02]. Semaphorin3a [FNK^{+}00]. Semaphorins [LeB02-87]. semipolarized [CC02]. Send [Wel02-51]. sense [Bir04c, LeB02-58]. senses [LeB02-65]. Sensing [Wel04-66, KOS^{+}04, MHT^{+}04, SWE^{+}03, DVE^{+}00, HCTM00]. Sensitive [FWP^{+}00, SMSM00, ACP^{+}02, CPC^{+}02a, GMY^{+}03, LeB03-58, SXD^{+}03, Tum04v, MSH^{+}00, DMA^{+}01]. sensitivity [ASK^{+}03, OTB03, SMS^{+}01a, NBWB^{+}00]. Sensitization [CDWB01]. sensor [Wel01-58]. sensory [LDK^{+}03, SWH^{+}02, WWD^{+}04]. sentence [Gri03]. separate [Dov03-64, Les01d]. separates [KH04]. separation [LeB02-83, TMG03, Wel01-49, DTO^{+}01, MSFH00, PGSL00]. Separin [JSCR01]. septate [GF03, STA03, WSH^{+}04, TK00]. Septin [GBZ^{+}02, VT04, BPC03, TBTN01, TMG03, Wel04q, WBAS04]. Sequence [ELO^{+}01, EKB^{+}03, SWB03, AERD^{+}01]. Sequences [LeB02-88, XBL^{+}03, SMSM00]. Sequential [MGP^{+}02a, MSN^{+}02, SBS^{+}02b, THO^{+}04, VN04]. sequentially [FGR^{+}04, KTY^{+}00]. Sequestered [MMDC00, SSW^{+}01b]. Sequestering [KKS^{+}01]. Ser364 [TLS^{+}01]. SERCA [LC04]. Serca1 [CGL^{+}01]. Serca2b [RLC00]. serine368 [RDC^{+}04, LTB^{+}00]. Serum [SPB^{+}02, GCT02]. serve [ATF^{+}04, DLY^{+}02]. serves [HHOP02]. set [Dov01-28, Wel02a, XAB^{+}03]. sets [Wel01b, WLPD04]. several [APM^{+}02]. severe [MPG^{+}03, RKF^{+}04, YIS^{+}03, ZCH^{+}01, CDFT^{+}01]. severity [MPG^{+}03]. Sex [LeB02-89, Wel03-56]. Sfi1p [Kil03]. Sgo [Tum04w]. Sh [TOM00]. SH2 [DOB^{+}01]. SH2-containing [DOB^{+}01]. Sh3 [JRW^{+}01]. Shaker [PSE^{+}03]. Shaker-like [PSE^{+}03]. Shank3 [SRG^{+}04]. Shape [Dov01-32, TCV^{+}00, FAT^{+}02, Gla01b, HIN^{+}03, LeB04-65, LG02b, SB03, Wel04-60, YHB^{+}04, MKK^{+}00a, PGS^{+}01, RS00c, vdB00]. shaped [LeB04-43]. Shaping [HF01, LeB03-86]. share [Gla01a, JBK04]. Shared [FLX^{+}00, Tum03m]. Shares [LBP00, KMH^{+}00]. She3p [EKC^{+}03]. Shear [WSR03, ITM^{+}04, YWW^{+}04]. Shedding [FWP^{+}00, Pow01i, MGAL^{+}01, SWK^{+}04, Wel04y, YIS^{+}03, THE^{+}00]. Sheet [KGE^{+}00]. Shh [LeB04-86]. Shigella [BGW^{+}04]. shine [RO04]. Shining [Dov02-52]. Shinya [DV04]. ship [Dov01-32, DOB^{+}01]. SHIP-2 [DOB^{+}01]. shmoo [Dov04g, Wel04-31]. shock [DC02a, Dov02v, JKG^{+}02, BMC00]. Shocking [LeB02-90]. short [Dov02-65, GCR^{+}03, Dov03-32, GPAS^{+}01]. short-term [Dov02-65]. show [CBG^{+}01, HvdHG^{+}03, YIS^{+}03]. shown [SSGLS01]. shows [GNDLS^{+}01, KAC^{+}04]. SHP [CNHK02, LMGM^{+}02, MDQ^{+}03]. SHP-1 [MDQ^{+}03]. SHP-2 [CNHK02, LMGM^{+}02]. Shp2 [SGF^{+}00]. shrink [LeB03-93]. shrinkage [ST03]. shrinking [Dov03p]. shuffle [Wel03r]. shuns [LeB04-70]. shuttling [JRL^{+}03, KBG^{+}03, MSM^{+}04a, MTPT02, SBMB^{+}04, HCD^{+}00, KCL^{+}00, NA00, VPP^{+}01]. Sic [LeB02-43]. sick [LeB02-43]. side [ATG^{+}03, ASGGR02, Dov02x, LeB02t, LeB03-57, LeB03-66, Les02a]. sidekick [LeB02-98]. sides [Tum04d]. sight [Dov04k]. sightings [LeB04-106]. Sigma [Wel01-49]. SIGN [CdLvM^{+}04]. Signal [CS01, ELO^{+}01, GDHS01, MSS^{+}01, PLP02, SM03a, BBDM02, EKB^{+}03, GKN^{+}03, HTT^{+}02, HSMB02, HHOP02, JBK04, JK01, KSS^{+}03, LM01, LeB04-30, MWAM01, MJG03, MC02, PSD^{+}04a, RPNM03, Tum04y, UHR^{+}03, Wel02-53, WW02, YHT02, YWH04, FSD00, FKG00, HVM^{+}00, KIK^{+}00a, KEGDQ01, KSF^{+}00, TMHP00, Wel02-51]. Signal-dependent [SM03a]. Signal-Regulated [CS01, BBDM02]. Signaling [BYLA^{+}01, BPMG00, CGY^{+}01b, Dov02-53, Dov02-54, FWP^{+}00, IWG^{+}01, KCL^{+}01, KSK^{+}00, MBSB00, MHK04, SHN^{+}01, SZZ^{+}00, SKF^{+}01, Wel04-67, vdHvODML^{+}00, AMBW04, ADL^{+}03, BSD^{+}01, BWRT03, BHW^{+}03, BAZA03, BFH^{+}01, BDR^{+}03, BHK^{+}02, COB01, CAW^{+}04, CDW^{+}03, DLPB03, DMC^{+}03, DPB03, EWD02, EPH^{+}02, FLS^{+}03, FAF^{+}04, GPDvH^{+}03, GLS^{+}03, GTPG03, GFM^{+}04b, GMLM^{+}04, iHGK^{+}02, HIE^{+}01, HFG^{+}04, HTRK02, HFK^{+}03, IOIF^{+}04, JLJD03, KRU^{+}04, KMH^{+}04, KYBS03, KLZ04, KWSK^{+}04, LMGM^{+}02, LRS^{+}02, LeB04-100, LDP02, LKH^{+}04, LW03, LTD^{+}01, LRWB04, MWL01, MKD^{+}01, MBLCE03, MTW^{+}04, MPAP^{+}03, MMG^{+}04, NH03b, OEM^{+}02, OKM^{+}01, PGSE^{+}01, PDL^{+}03, PAR^{+}04, PFSG03, PHM^{+}02, SJW^{+}04, SLR^{+}03, SCTF04, SKK^{+}02, SWH^{+}02, SRG^{+}04, SGK^{+}02b, SATA^{+}02, SCD02, SDS^{+}04, Tum04s, UML^{+}03, Vin04, WLW^{+}04, WSC^{+}01, WSWSL04, Wel01p, WWD03, WKJS^{+}04, XBL^{+}03, YSK^{+}04, YMK^{+}04, YK03a]. signaling [YSC^{+}02, YSC^{+}21, ZWAH03, ZHH^{+}02, ZJM^{+}02, AIH^{+}00, BZSC00, CSD00, DA01, FRO01, FCL^{+}00, GSB^{+}00, GWG01, HK00, IAG^{+}00, KMB^{+}01, KHH^{+}01, LR00, LCS^{+}01, MSO^{+}00, MTV^{+}00, PDJ00, RIDC01, Rap00, SBZ^{+}00, SJIM01, SRKR00, THK^{+}00, VMD^{+}01, WTG01, WRCU00, WPO00, ZLG00, CKS^{+}04]. Signalplex [LM00b]. Signals [GMRS00, NLRD01, BSMS03, BKZ^{+}03, Dov03-65, Dov04e, GK03a, HHS03, JKB^{+}03, KEHAM^{+}02, LeB02-28, LeB03-46, MWE^{+}03, MHlW02, OPP^{+}03, PLW^{+}04, RCS^{+}02, SKN^{+}03, Sre04d, WKYC02, JMG01, YCX^{+}01]. silencers [Wel04-59]. silencing [MWAM01, Tum04u, SRHV00]. silent [SNF^{+}02]. Sim4 [PRA03]. Similar [HCD^{+}00]. Simple [CDB02, GLDM01, Dov03-45, LeB03-96]. simplex [CDW^{+}03, SLB02, CLWR01]. simulations [Néd02]. simultaneous [BMS^{+}03]. Single [FFKC00, RDP03, WBP^{+}03]. Sinuous [WSH^{+}04]. Sir [SNF^{+}02]. sister [SAH^{+}03, HSKG00, MR01b]. Site [BRM^{+}00, RSBE00, WSL^{+}03, Dov03y, GBOL03, HZS^{+}01, JK01, KYS^{+}02, MGP^{+}02b, PLR03, POH^{+}04, RWH02, SRC^{+}01, TUK03, UHR^{+}03, VCGB^{+}02, XRP^{+}01, AFB^{+}01, CMMP00, KWO^{+}00, TGMC^{+}00, VMO01]. Site-Directed [RSBE00]. Site-specific [WSL^{+}03]. Sites [WKS^{+}00, ACP^{+}02, BHPN04, CBS04, DSH^{+}03, Dov02-51, EBWC01, FNKH02, GFGP03, GQI^{+}02, KLG^{+}02, Kil03, KR03, NGKH02, PLP02, SHE^{+}02, SCK04, SLD^{+}02, ZTK^{+}03, BPS^{+}00, BTD^{+}00, FNK^{+}00, HWW01, SMSF00, SBI^{+}00, TWS^{+}00a, ZGB01]. Situ [CS00]. situation [LeB03f]. six [SWK^{+}04, MMH^{+}00]. Size [NHG^{+}03, KRMB03, LeB04q, LGB^{+}02, VTGT^{+}03, Wel01b, Wel02t, Wel03p, WSH^{+}04, vRTvdB^{+}00]. Size-selective [NHG^{+}03]. sizer [ARQ^{+}04]. Sizing [Les01k]. Skeletal [DDV^{+}03, RL03, UIK^{+}01, ASGGR02, CSP^{+}04, EPN^{+}03, GBN^{+}01, GNS^{+}04, GBJ01, HP03, HYMS^{+}02, IKS^{+}01, JLJD03, KSNS^{+}04, KSD04, LCRS01, NDM^{+}03, PR02, RLZ^{+}03, RMMP04, RDNB02, TAA^{+}02, VMS^{+}02, WCTU02, BHY^{+}00, BHFL01, BL01, CDFT^{+}01, DCM00, FRO01, FKG00, FHP00, KMG^{+}01, RMHM00, TTS00, WRCU00]. skeleton [JB01, KHLW02, MWK^{+}02, WSL^{+}01]. Skeletons [McF00]. Skeletor [WWJ^{+}00]. Skin [ALWR01, Dov02q, Dov02-48, Dov04j, GKSR00, MDT^{+}01, MCG^{+}03, Wel01c, YIW^{+}04, RBM^{+}00, SJA^{+}00]. Skipping [LMW^{+}00]. Skp2 [CP01]. Sla1p [HHOP02]. SLAP [KSK^{+}00]. Slc27a4 [HvdHG^{+}03]. sleep [Wel03-66]. Sli15 [sKCK^{+}01]. Slide [Les01l, LeB02-106]. Slingshot [LeB04-87, NOOG^{+}04]. slippery [Lai03d]. Slipping [Tum04x]. Slit2 [DRBF03]. Sln1 [RRS03]. Slow [LeB02-91, SH02, Wel04-68, Bir04e, HDP^{+}01, JLJD03, RGM^{+}02]. slowly [NSLSK02]. Slows [AGB^{+}00]. SLP-76-Associated [KSK^{+}00]. Sly1 [PG02]. SM [BJ03, GBOL03, TBJ^{+}01]. Smac [RDP03, RMMC01]. Smac/DIABLO [RDP03]. SMAD [RPNM03, CSD00]. Smad1 [PAR^{+}04]. Smad3 [YCX^{+}01, BDR^{+}03]. Smad6 [HIO^{+}04]. Smad6/ [HIO^{+}04]. Smad7 [HIE^{+}01, SSH^{+}04]. Smads [MHK^{+}01]. Small [HN03, MBN^{+}01, Wel03-57, Dov02-39, HCL^{+}03, ISB^{+}04, KMK^{+}02, LeB04w, SRKN03, BMF00, HDL^{+}00, KMCM00, PGS^{+}01, PKF^{+}00, RRK^{+}00, RBB00, TCR00, WYHP00, YSN^{+}01]. smart [LeB02-59]. SMC1 [EOJ^{+}03]. SMC3 [EOJ^{+}03]. Smitin [KK02]. Smn [MMBB03, RJA^{+}03, BCB^{+}02, CDFT^{+}01, MPG^{+}03, TBJ^{+}01, NDM^{+}03]. smooth [BJM^{+}02, HGC02, KK02, MRM^{+}04, RFLT02, WHM^{+}02, DBS^{+}01, ONS^{+}00, TTS00, WGP^{+}00]. Smurf1 [HIO^{+}04]. sn1 [BHW^{+}03]. sn1-DAG [BHW^{+}03]. snag [Dov02-70]. SNAP [WPM^{+}00]. SNAP25 [THO^{+}04]. SNAP’d [LeB04-66]. Snare [PSWU00, BBSF01, HGS^{+}01, MTG^{+}02, MWM^{+}02, MW04, MSN^{+}02, PBP^{+}01, PG02, SVI^{+}04, SYVB03, Wel04-69, CB00, WSE00]. SNARE-mediated [SVI^{+}04]. SNAREpin [MSGS02, SBS^{+}02b]. Snarepins [WPM^{+}00, MWM^{+}02]. SNAREs [Dov04n, EBWC01, LMHJ02, LeB02-64, LeB04-59, MPBR03, RAD^{+}02, SBT^{+}04, SDL02, VVR^{+}04, Wel02g, GBON00]. Snaring [Wel04-70]. Snrna [YSN^{+}01, FM01, GBOL03]. snRNP [SN04]. snRNPs [HDL02, TBJ^{+}01]. snurportin1 [HDL02]. SNX1 [LeB04-88]. Sodium [RWCC01, ARMB04, JB01, KS02, SLB02]. sodium/hydrogen [ARMB04]. Solo [Wel04-71]. Solubility [iNFK^{+}01]. soluble [ZFH^{+}04, LLH^{+}01]. solution [Fre02]. solve [Cam03]. solved [Min01]. song [Dov04i]. Sonic [FLS^{+}04]. sort [GGGK03, Gla01c]. sorted [VN04, HMRH01]. Sorting [Dov02-55, KSNS^{+}04, LOS^{+}01, LeB02-92, NHB00, RTM^{+}01, Tra03, AHA^{+}04, BWK^{+}03, Dov03x, FS03, GK03a, HS02, JKB^{+}03, KSBE03, MJS02, MPBR03, NG02, PSP^{+}04, PFSG03, RCS^{+}02, Sea04, SDC^{+}01, ARK^{+}00, BMS^{+}00a, EHCC^{+}00, KNIO01, KWO^{+}00, SKR^{+}00]. SOS [LeB04-89, ITF^{+}02, IFP^{+}03, STA^{+}01]. Sos-1 [ITF^{+}02, IFP^{+}03, STA^{+}01]. sound [Dov02-56]. source [KNK^{+}01, SDML04]. Sox [LeB02-89]. Sox5 [SDML04]. Sox6 [SDML04]. SOX9 [BvdWD^{+}04]. SP1 [LSW^{+}03]. space [RJyH02, DSH^{+}00, WWG^{+}00]. spaces [Dov02-60, TAA^{+}02]. Spain [MR02]. spanning [MN03]. Spare [LeB03-87]. spares [LeB04-77]. spastic [ASK^{+}03, EMY^{+}04]. Spatial [AR03, GHK^{+}03, HCTM00, KLG^{+}02, LCSG03, MPG^{+}02, MSFH00, VMK^{+}03, BHW^{+}03, KMP02b, OKC02, SCPP02, ZAE^{+}04, SPC00, SGO^{+}00, TYY^{+}00]. spatially [MMBM04, NK02, PMPH03, RDC^{+}04]. Spatio [BKZ^{+}03, TNMM03]. Spatio-temporal [BKZ^{+}03, TNMM03]. Spc42p [CMM^{+}02]. speak [LeB04-52]. special [Dov03-35]. specialization [NZA^{+}01, WDFNN04]. Specialized [AH01, HWW01]. species [CPG^{+}03, DMLK04, RGG03, WK04]. Specific [Hem01, HSW00, KKL^{+}01, LMGM^{+}02, SNL^{+}00, YW00, AFR01, AOJ^{+}04, BMM^{+}02, BXR^{+}02, CDK04, Dov01o, Dov03x, FHL^{+}03, FBH03, FJM^{+}04, HIE^{+}01, HBK^{+}02, HTS02, ICN^{+}03, JBK04, MKS^{+}02, MPAP^{+}03, RP03b, SJB^{+}03, TCH^{+}00, UML^{+}03, WHM^{+}02, WKZ^{+}02, WSL^{+}03, YHZ^{+}01, ZZM^{+}03, BRS^{+}01, CDEM00, CMMP00, JWJJ00, MÖS^{+}00, MNHR00, PGS^{+}01, SYH^{+}01, TSDS00, STA^{+}01]. specifically [AMG^{+}01, AZP^{+}02, CYC^{+}04, TSL04]. specification [ASGGR02, LCSG03, FFY^{+}00, KPB^{+}00, LHvdH00]. specificity [BSEB04, NG02, PG02, Wel01-54, CS00, DOL^{+}01, SDDS00]. specifies [CLM^{+}04]. specify [LRBH02, LeB02-88, HLK01]. speckle [MAG^{+}04, SAWS02, KLF^{+}00]. speckles [KMP02b, LL02, DLS00]. spectraplakin [BHL^{+}03]. Spectrin [HDJ00, MCB00a, BIC^{+}03, Dov02-42, JB01, KS02, LGGS^{+}04, LSMS03, MWK^{+}02, NM02, WSL^{+}01, BAD^{+}00, DWD^{+}00]. spectrin-based [JB01]. spectrometry [MSJ^{+}04]. spectroscopy [LMVW03]. Speed [LeB04-90].

Speedy [PDT^{+}02]. spells [LeB02-29]. Speract [WDW03]. sperm [FYI^{+}03, SMW^{+}03, Wel01-55, WDW03, GZY^{+}00]. spermatogenesis [BBP^{+}04]. Spermatogenic [GBM^{+}00]. Spermatozoa [GMZ^{+}00]. Spermidine [CYC^{+}04]. Spermidine/ [CYC^{+}04]. spermine [CYC^{+}04]. Spf1p [CRH02, CKFH00]. Sphingolipid [PKF^{+}00, LGRP^{+}02, MHT^{+}04]. sphingolipids [Dov04l, PWS^{+}01, TWS^{+}00b]. Sphingomyelin [TRW^{+}00, BSW^{+}00, Gre00]. sphingomyelinase [LXL^{+}04b]. sphingosine [MO01, TPW^{+}04, LGRP^{+}02]. sphingosine- [TPW^{+}04]. Sphingosine-1-phosphate [LGRP^{+}02]. sphingosine-induced [MO01]. Spike [AKT01]. Spikes [AS00]. spillover [UTH^{+}02]. spin [Wel02k]. spinal [DWM03, GLSG^{+}02, MPG^{+}03, PMKV01, RJA^{+}03, TBJ^{+}01]. Spindle [BGBG03, CENMR^{+}01, GKG^{+}01, GES04, HKBH03, MHS^{+}00, SBC01, TEC^{+}03, AZ03, BTVB03, CMM^{+}02, Che02, Dov01-33, GK04, GC04, GNH^{+}04, GV04, HdEV^{+}02, HCL^{+}03, HG03, HMC^{+}01, Hoy01, HBB^{+}04, ISS^{+}04, IKS^{+}02, JGW02, KCG^{+}03, Kil03, KSK^{+}02, LMG04, LeB03-37, LeB03-88, LeB03-92, LeB04q, LeB04-97, LTD^{+}02, MSL^{+}02, MRK04, MKJ^{+}02, PMGS02, RRB^{+}03, RRSV02, SHP01, TKHR03, THG^{+}04, Wel01-48, Wel02h, Wel03-33, Wel04-60, WDMH03, AOC01, BKDH01, GG01, GHC01, HJSM00, HHF^{+}00, JSCR01, KMCM00, LHvdH00, RCL^{+}00, SCM^{+}00, SHHH01, WBG01, WWJ^{+}00, WWKV00]. spindle-stabilizing [LTD^{+}02]. spindles [JTK^{+}02, KM01, LC01, MSC^{+}03, MPB^{+}04, SB03]. spine [OTY^{+}04]. spines [HIN^{+}03, EHM^{+}00]. splash [Tum03j]. splice [RDN^{+}03, vdFKK^{+}02]. spliceosomal [AvdWM^{+}01, HDL02]. Spliceosome [HM00]. Splicing [Dov03-46, LeB04-91, vdHvODML^{+}00, AvdWM^{+}01, AOJ^{+}04, BCB^{+}02, BPD^{+}04, KMP02b, LeB02-89, LeB04-80, LL02, MCG^{+}03, SBS02a, DLS00, KLF^{+}00, MRM^{+}00, ONS^{+}00]. spoil [LeB03-87]. spoils [Dov02-30]. Spoke [GDRS01]. Spokes [YDRS01]. Spontaneous [PR00]. spore [TBTN01]. sporozoites [KSD04]. sporulation [LeB02-69]. Spots [SMS^{+}04b]. spreading [AQC04, CRS^{+}03, CBLT04, DBB^{+}02a, ISID^{+}03, LRD^{+}03, MC01, WZB^{+}01a, WLO^{+}02, WMJ^{+}04, ZSY^{+}03, BMD^{+}00, IAG^{+}00, RS00a, RBBA00, THZ^{+}01]. springs [Wel04-36]. Sprouting [PFW^{+}00, GGF^{+}03]. Sprouty [IWG^{+}01, LeB03-48]. Sprouty-1 [IWG^{+}01]. squamous [ANC^{+}02, JW04, TAA04]. squeeze [Wel01u]. squeezed [Wel04z]. Squeezing [Wel01-50, Wel02-52]. SR [BPD^{+}04, YF00]. Src [BDKM04, CROfC04, EPH^{+}02, FSK^{+}04, ISID^{+}03, Jay01, LCGR00, LWCKL01, LDI^{+}03, MWF02, MSV^{+}00, MTV^{+}00, MNT^{+}03, OEM^{+}02, PPA^{+}03, RDH^{+}01, SHN^{+}01, SF01, VCDHD03, WCBC04, WDL^{+}04]. Src-astic [Jay01]. Src-dependent [WDL^{+}04]. Src-mediated [EPH^{+}02, PPA^{+}03]. Src64 [KHC02]. Src64-dependent [KHC02]. SRm160 [WCIN04]. SRP [MJG03, MGMH03, GDHS01]. SRP/FtsY [MGMH03]. Ssa2p [BMC00]. Stability [HSMJ01, KLK^{+}01, DKA01, GHK^{+}03, HBAF^{+}02, LSK01, LTD^{+}02, SMZ^{+}03, Tum03e, YDPK04, KMS00, RBV00, TB00b]. stabilization [CFC^{+}01, ZWB04, JCR^{+}01]. stabilize [MDP02]. stabilizes [BPC03, LGGS^{+}04, MPR^{+}03, Tum04w, vdBCH^{+}04, GOL^{+}01]. stabilizing [GBJ01, KvHB^{+}01, LeB02-34, LTD^{+}02, MSA^{+}03, YLQ^{+}02]. Stable [LeB02-93, DMH^{+}02, Néd02, WSW^{+}00]. stably [MWF02, WSC^{+}03]. stacked [SHF^{+}03]. stacking [LLP^{+}02]. Stage [PSWU00, Dov03-49, WSK^{+}03, YSS^{+}01]. stages [ALP^{+}04, EGWK^{+}01, GJS^{+}03b, SKM03, WSC^{+}03, YGWN01, NS01]. stairmaster [Wel04a]. stalled [DC03]. Stand [LeB02-94]. standardized [LDC^{+}04]. Staphylococcus [BSD^{+}01]. starch [TOM01]. starfish [LRD^{+}03]. start [Dov02-32, Dov02-51, FNFL03]. started [Dov03-29]. starter [Les01h]. Starts [GKG^{+}01, Dov03-61, Wel04o]. Stat [MSS^{+}01, RPNM03, Vin04, Wel02-33]. Stat1 [MSM^{+}04a]. STAT3 [MYO^{+}04, SGK^{+}02b]. STAT5A [LeB04-31, WAV^{+}04]. State [KIO^{+}00, CSP^{+}04, FKH^{+}04, GHS^{+}03, HMG03, WSWSL04, DAC00, GOL^{+}01, GH00, KLF^{+}00]. states [ADL^{+}03]. static [GSB^{+}03, KM01]. Stay [Bir04e]. Stayin [Les01m]. Staying [Les01n]. Steady [HMG03, LeB02-82, LeB02-91, Wel04-68, DAC00, KLF^{+}00]. Steady-state [HMG03]. steers [LeB02s]. Stem [ALWR01, Dov03-47, BSW^{+}04, DDV^{+}03, LeB02-74, LeB03n, LHC^{+}02, MRM^{+}04, PDL^{+}03, Pro03, QPDJ^{+}02, RPNM03, SPB^{+}02, SK04, SSRX04, SwZK^{+}02, TCP^{+}03, TKS^{+}02, WNM^{+}03, Wel03-27, KMG^{+}01, MYH^{+}01, TTP^{+}01]. Step [BHL^{+}01, MMS00, SDS00, AMG^{+}01, DLT^{+}02, KRMB03, LeB04o, Mel04a, PWU00, TCR00]. steps [HP04, LeB03-27, ADKK00]. stereocilia [iKFH^{+}04, RSD^{+}04]. sterol [MHT^{+}04]. sterol-sensing [MHT^{+}04]. Stick [Tum04y, Dov02-38, LeB03-33]. Sticking [Tum04z, Wel02-53, Wel04-49]. sticks [LeB02-49, LeB03-75]. Sticky [Dov03-48, LeB04-92, Wel01-51, Wel02-54, Wel03-58, DSG04, Wel03-54]. stiffness [EGS^{+}04, KNK^{+}01, LeB04k]. still [LeB04-44]. Stimulate [HGC00, PMBC^{+}00]. Stimulated [DSSY00, KSK^{+}00, BKZ^{+}03, CLB^{+}03, KMH^{+}04, SJW^{+}04, SKK^{+}02, SDEZ^{+}03, ZSY^{+}03, CBZ^{+}00, KHH^{+}01, MMG^{+}01, MTB^{+}02, ZVPK03]. Stimulates [HP00, LBP00, FLWMG02, HP03, KKW^{+}03, MWL01, SC01a, SVI^{+}04, WK02, WSFN00]. Stimulating [NLRD01]. Stimulation [AS00, DSSY00, Dov01-31, MSGS02, MSD^{+}04, REK^{+}03, LLH^{+}01, WXQ^{+}00]. stimuli [APM^{+}02]. stimulus [SDL^{+}03]. stoichiometric [SST^{+}01]. Stoichiometry [KHvOD00]. Stonin [MBAB01]. Stop [LeB03-89, Wel02-55, Wel02-56, Dov03-32, LeB04d]. Stops [LVWA01]. storage [IBP^{+}04, JPH^{+}01, LFM^{+}04, HMRH01, JPRR00, OPZ^{+}01, TOM00]. store [GMY^{+}03, MH02, MPV^{+}01, TFAM^{+}04]. store-operated [MH02]. Stored [SMW^{+}03]. stores [LeB03-58, PFM^{+}00]. Stories [WL03]. storing [Wel01l]. story [Wel01-63]. straight [Wel04h]. Strand [FFST01, SCMH00]. stranded [BM02, DJ03]. strange [Dov02-54]. straps [Dov02-39]. strategy [Wel02w]. stratified [ANC^{+}02]. Stratum [EMW^{+}01]. strength [HPG^{+}02, CKS^{+}00]. strengthened [CTE^{+}04]. Stress [JMG^{+}04a, KKA^{+}01, ASK^{+}03, BAZA03, BND^{+}02, Dov03l, HKE^{+}04, HIT^{+}02, ITM^{+}04, JRL^{+}03, JKG^{+}02, KOS^{+}04, KFO04, LeB02a, LHR04, MP04, RKR^{+}03, SS04, SVT^{+}02, TCZ^{+}03, VM02, WSF^{+}01, Wel04j, Wel04s, WSR03, ZRDG02, DB00, HSB00, KCL^{+}00, MMG^{+}01, PO00, TYY^{+}00, YHF^{+}01]. stress-dependent [WSF^{+}01]. Stress-induced [JMG^{+}04a, HKE^{+}04, ZRDG02]. stress-sensing [KOS^{+}04]. Stressed [Dov03-49, JON^{+}03]. stresses [MSY^{+}04]. stretch [Dov02-33, BMS^{+}00b]. striated [BBG^{+}03, SEI^{+}00]. striation [LeB04k]. strict [WLPD04]. Stringent [AKW00, CMMP00]. stripes [Wel03-35]. stripping [LeB04-103]. stroke [BWW^{+}02]. strokes [Dov03k]. stromal [KVC^{+}03]. Stromelysin [ASMW01, IOLA^{+}00]. Stromelysin-1 [ASMW01]. Stromelysin-3 [IOLA^{+}00]. strong [LeB03j, Wel03-57, BDK^{+}01, RPE00, TAD^{+}00]. Structural [AKH00, BBBS04, HLB^{+}02, LeB02-95, AZ03, SSGLS01, YLY^{+}02, WCGT^{+}00]. Structurally [AKK^{+}00, HCD^{+}00]. Structure [HES00, HDJ00, SLSR^{+}00, Wel04-72, vARP^{+}00, BYMS^{+}02, Dov04k, HWBD^{+}01, KKPB03, MKS^{+}02, PL01, SPK^{+}01, TDhL^{+}02, Tum03m, WPC^{+}01, EMW^{+}01, HMAM01, JWJJ00, PMSB01, PGV^{+}00, WPJ^{+}00]. structures [BKD^{+}04, IBP^{+}04, LeB04-104, LFM^{+}04, MMG^{+}04, BHBJ00, RMG^{+}00, SYH^{+}01]. STT3 [NKM^{+}03]. Stu2 [SHHH01]. Stu2p [vBDH03]. Stuck [Sre04m]. studies [TCK^{+}03]. stuff [LeB04-63]. Subcellular [DKMW00, GH00, GMRS00, GJB^{+}00, vdHvODML^{+}00, ASYL04, CBC^{+}01, DAV^{+}03, MHT^{+}04, MSM^{+}04a, SSOS01, VMK^{+}03, vdFKK^{+}02, KHvOD00, RSG01]. subcomplex [BRB^{+}01]. Subdomain [KKL^{+}01, MBS^{+}01, WGP^{+}00]. Subdomain-Specific [KKL^{+}01]. Subject [GKG^{+}01]. submembraneous [DOB^{+}01]. Subnuclear [NHS00]. subpopulations [LOS^{+}01]. Subsequent [ZBB^{+}00, WSC^{+}03]. subset [HS02, HC02, YMK^{+}04]. Substrate [FBH03, WSL^{+}00, AOH^{+}02, COB01, KKK^{+}02, LTD^{+}02, LdVV^{+}02, NK02, PSD^{+}04b, Wel04-33, YSK^{+}04, vdLBK^{+}04, DOL^{+}01, GKM^{+}01, JC01, YSS^{+}01]. substrate-2 [AOH^{+}02]. substrate-dependent [NK02]. Substrate-specific [FBH03]. substrates [EGS^{+}04, GSB^{+}01, LeB02-29, PR00]. substratum [SC04]. substratum-dependent [SC04]. Substructure [FMP^{+}00]. subtelomere [TSL04]. subunit [CYC^{+}04, ESC^{+}01, NKM^{+}03, RGM^{+}02, RCY^{+}03, RCS^{+}02, RP03b, TVF^{+}03, vdLBK^{+}04, ALJ00, FKG00, HKJ00, JPM^{+}00a, ONS^{+}00, PHWK^{+}00, TCS01]. Subunits [LYKH00, BHPN04, PRLR02, RWCC01, YCK^{+}03, vdFKK^{+}02, GOL^{+}01, GMC^{+}00]. subventricular [ACBG04]. Sudden [Wel01-52]. sufficient [CRP^{+}04, LRF^{+}02a]. sugar [LeB04-92, Wel01-50]. suggesting [YLG^{+}02]. suggests [BBG^{+}03, KB04, OSMF00]. Suicidal [LeB02-96]. Suicide [Dov02-57]. sulfate [ADL^{+}03, AFR01, AR03, Dov01o, KRU^{+}04, SYW^{+}03, YPN^{+}04, LLH^{+}00]. sulfation [ADL^{+}03]. sulfotransferase [KRU^{+}04]. SUMO [AAD03, JTK^{+}02, LH03]. SUMO-1 [JTK^{+}02]. SUMO-2 [AAD03]. SUMO-2/ [AAD03]. sumoylated [SAH^{+}03]. SUMOylation [SBMB^{+}04].

SUP [AOJ^{+}04]. SUP-12 [AOJ^{+}04]. Super [TWBV^{+}01]. Superfamily [TYS^{+}00]. Superoxide [FBV^{+}04, MH01]. supple [LeB03-51]. support [CSL^{+}03, LeB03-40, AKW00]. supporting [MCA^{+}03]. Supports [DSSY00, Hel03, HEN^{+}01, YKT^{+}04, IAG^{+}00]. Suppress [CK00]. suppressed [DSG04]. suppresses [LeB02-102, CLAC00, GWG01, WMS00]. suppressing [WKYC02]. Suppression [SMS^{+}01a, UAZG00, JZ02, TSY^{+}02, WG03]. Suppressor [HMN^{+}00, PTM^{+}01, HFG^{+}04, JHS^{+}02, KSC^{+}04, ZHDB04, PDW^{+}00, RIDC01, RSG01]. suppressors [SSM^{+}04]. Supramolecular [BKZ^{+}01]. Surface [DFYL00, FSGDN^{+}00, ADL^{+}03, CPA^{+}03, GSB^{+}03, HRV^{+}01, KNR^{+}04, LeB03v, MBMMA^{+}03, NGKH02, PMPH03, SM03a, SC04, KGC^{+}00, MTV^{+}00, NKP^{+}01, TRW^{+}00, TMK^{+}00, TSK^{+}00]. surfactant [Wel01y]. surrounding [MWN^{+}04a]. Surveillance [AOC01]. Survey [FSBH00]. Survival [Dov02-58, EAD^{+}02, GMRS00, WWK^{+}00, Wel03-59, BS04a, BFSO^{+}04, DMC^{+}03, EES^{+}01, FGSW03, HBB^{+}02, JON^{+}03, KVC^{+}03, LHC^{+}02, LPL^{+}04, MWAM01, MDQ^{+}03, MTB^{+}02, MMBB03, NGS^{+}01, OMB^{+}01, SGK^{+}02b, SPA^{+}04, UCY^{+}02, WKYC02, Wel02-57, WWD^{+}04, XLGS01, ZLR^{+}03, AIH^{+}00, BLPP01, Ern00, LR00, MHE^{+}00, MKK^{+}00b, PCR^{+}01, WGF^{+}00, ZLMP00]. survive [Dov01r]. Survivin [SMLM00]. Surviving [Wel01-53, Wel03-60, Wel04-73]. Sustained [LSA^{+}00, Wel01-54]. sustaining [LeB02-30, OMB^{+}01]. sustains [LeB03-44, WFF^{+}01]. Swallowing [LeB02-97, LeB04t]. sweat [Dov03-59]. Sweet [Fre02, Wel03-59]. swell [Wel03-41]. Swelling [LeB04-93, RDS02]. swelling-induced [RDS02]. swept [LeB02-57]. Swimming [Wel01-55]. Swiss [CDWB01, SRSW04, TIO^{+}02]. Switch [JW04, Wel02-57, LDI^{+}03, OTB03, RKJL03, WRSMO^{+}04, Wel02x, Wel04-28]. switchboard [Dov02s]. switches [HdEV^{+}02, MHK04]. Switching [Dov04o]. Syk [OEM^{+}02]. Symmetric [SJA^{+}00]. Symmetrical [BCB^{+}02]. sympathetic [CJ02, CSJ03a, GCH03, GRCF02, MWAM01, OMB^{+}01, PKR^{+}02, PSK^{+}03, YJS^{+}03, FXPT00]. Synapse [LeB03-90, ZWAH03, Dov02k, Dov02-51, LLGB03, LeB02-98, LW03, SDD04, WWBGG03]. synapses [Dov03a, GRCF02, Les02c, Wel04-39, AKK^{+}00]. synapsin [BET^{+}03]. synaptic [BET^{+}03, HHS03, KL04, KKW^{+}03, LeB02-106, LeB04-94, NOM^{+}04, NB03b, Riz03, TSL^{+}03, Tum03g, YpHRL03, HHHJ00, JCR^{+}01, WXQ^{+}00]. Synaptojanin [HHHJ00]. synaptonemal [EOJ^{+}03, DR00a]. Synaptotagmin [MCH^{+}00, CKF^{+}03, EBWC01, JK01, MSGS02, TEB^{+}03, DVE^{+}00, KSF^{+}00]. Synaptotagmins [ASYL04, DRG^{+}03]. Synbindin [EHM^{+}00]. SynCAM [LeB02-98]. synchronicity [LeB03-95]. synchronize [LeB02-98]. Synchronized [MLZ^{+}01]. synchrony [Tum04b]. syncytial [RSK02]. Syndapin [QK00]. Syndecan [Rap00, BBR04, HTS02, AKT01, EHM^{+}00, FWP^{+}00]. syndecan-1 [BBR04, AKT01, FWP^{+}00]. Syndecan-2 [EHM^{+}00]. syndecan-4 [HTS02].

Syndecan-Regulated [Rap00]. Syndecans [IAG^{+}00]. Syndrome [BYLA^{+}01, FSM^{+}01, HP00, DC03, FP02, Wel01-52, BKI^{+}01, CGY01a, SJA^{+}00]. Synergism [CNJ01]. Synergistic [TTH^{+}01]. synergistically [HPG^{+}02]. Synergy [RSBE00]. synovial [DDV^{+}03]. syntaxin [PG02, AHMJ01]. Syntaxin1a [YSGS00]. synthase [LeB04-108, RP03a, WST01, vdLBK^{+}04, HCK^{+}00]. Synthesis [SSL^{+}00, CLB^{+}03, CLM^{+}03, LeB02-98, LeB04-40, LTD^{+}01, PLP02, FMP^{+}00, KKK^{+}00, MLC^{+}01, SJS^{+}00]. synthesized [LFM^{+}04, yZCKA01]. Synthetase [KKK^{+}00]. Syntrophin [HYMS^{+}02, AMF01, AKK^{+}00]. synuclein [WMG^{+}04]. System [WSL^{+}00, BHNG01, Dov03q, Dov03-48, DWM03, HHS03, KL04, LeB04-35, LDP02, SWBE^{+}04, SR03, TOTC01, Tum03k, Wel03l, WWD03, AEL^{+}00, BAD^{+}00, CMW^{+}01, MHW^{+}00]. Systemic [Wel02-58]. systems [Pow04].

T [BHK^{+}02, SDD04, BPMG00, BWA^{+}04, CGY^{+}01b, Dov01y, FFKC00, GQI^{+}02, HK00, KSK^{+}00, LeB04c, LeB04-94, LCS^{+}02, LW03, PTM^{+}01, SEP^{+}01, SKF^{+}01, SBM^{+}01, TNM^{+}03, VBR^{+}01b, VBH^{+}02, Zwe00, EBWC01, HGS^{+}01, MWM^{+}02, PBP^{+}01]. T-Lymphocyte [BPMG00]. T84 [LSA^{+}00]. TACC [GMD^{+}02]. TAG [PSE^{+}03, TGD^{+}03]. TAG-1 [PSE^{+}03, TGD^{+}03]. Tagged [Gla01d, PQF^{+}00]. tail [BCP03, BMY^{+}01, Dov03d, DW02, GTM^{+}01, HT01, RGM^{+}02, RCY^{+}03, RWK^{+}04, UIY^{+}01, WDW03, WLWB01, CDTW00, PVL^{+}00]. tail-anchored [BCP03]. tail-dependent [UIY^{+}01]. Tailed [TOM00]. Tailless [MDJF00]. Tails [Mac00]. Tak1 [MHK^{+}01]. take [LeB02-71, LeB02-94, LeB03-47, Tum03a, Tum04d]. takes [Dov02t, Dov03b, Dov03-40, Dov04a, LeB02-33, LeB03e, LeB04j, LeB04y, LeB04-56, LeB04-91, LG02b, Wel02b, Wel02-50, vdB00]. Taking [Dov03-50, Wel01-56]. tale [BCP03]. talent [Dov01h]. Tales [Gla01e]. Talin [LeB04-95, LDI^{+}03, MDW^{+}04]. Talin1 [GJS^{+}03a]. Talk [NY00, BDR^{+}03, LeB04-49, APLB00]. talks [LeB03-62]. tandem [EBWC01]. tangle [LeB03u]. Tangled1 [SGHL01]. Target [SNL^{+}00, Zwe00, FCM^{+}01, GLS^{+}03, KMH^{+}04, LeB02-110, PBB^{+}04, WHP^{+}02, BGS00, IHN^{+}01, MHK^{+}01]. Target-Cell [Zwe00]. Targeted [DEG00, MKR01, GG04, HvdHG^{+}03, HBB^{+}02, KLE^{+}02, RKKP02, SMR^{+}02, XRH^{+}03, FKI^{+}01]. Targeting [HBV^{+}01, KIO^{+}00, KHH^{+}01, SRKN03, Tum04-27, ABRA03, BHW^{+}02a, DJT^{+}03, DJ03, GK04, GLDM01, GBJ01, HHOP02, KEHAM^{+}02, KKK^{+}02, KAK^{+}03, KSK^{+}02, LPL^{+}04, MJG03, ORZ^{+}04, PWS^{+}01, SHA^{+}03, SHKS02, SCK04, UN03, WAC^{+}03, DZT^{+}00, EHCC^{+}00, FKG00, JMG01, KHK01, OLB^{+}00, PLL^{+}00, TMHP00, YKW^{+}00]. Targets [HMN^{+}00, CAGK^{+}03, JTK^{+}02, LRS^{+}02]. Tat [MC02]. tau [ABOS^{+}02, Dov02x, MBLCE03, TTHH00, SVT^{+}02]. Tbc [RB01]. Tbc/Rabgap [RB01]. Tbc2 [AZP^{+}02]. Tbce [BXR^{+}02]. TC10 [KMH^{+}04, WSC^{+}01]. Tcf3 [LSK01]. Tcp1 [MDJF00]. Tcr [KSK^{+}00]. tea [Wel01-56]. tea1p [BN02, FVC04]. Tea2p [BHM^{+}00]. tear [LeB02h].

Tearing [LeB03-91, LeB03-92]. technology [LeB03-100]. Teeny [LeB04-96]. telencephalin [EOB^{+}04]. tell [LeB02-51, Wel04-38]. Telomerase [KLK^{+}01, EKdM^{+}04, JBK04, LeB02-45, Ped04]. Telomerase/Vault [KLK^{+}01]. Telomerase/Vault-Associated [KLK^{+}01]. Telomere [VLL^{+}03, FCLSN03, SGdM^{+}01, TBW^{+}04, GZY^{+}00, TSDS00]. Telomere-independent [VLL^{+}03]. Telomeres [CC02, Wel04-74, LeB03p, LeB04-54, TSL04, XB04]. Telomeric [LeB04-97, RKKP02]. telophase [BPD^{+}04]. temperature [CPC^{+}02a, OSNG04]. temperature-dependent [OSNG04]. Temporal [SPC00, AR03, BHW^{+}03, BKZ^{+}03, HTPC04, KLG^{+}02, TNMM03, VMK^{+}03, ZH04]. temporally [MMBM04, RDC^{+}04]. temptation [RY04]. Tenascin [Jeg01e, WMS00, MBLCE03, MSI^{+}03, STJ^{+}01]. Tenascin-C [WMS00, MBLCE03]. tendon [CLM^{+}04]. Tendons [ECO^{+}00]. Tensegrity [Pow01j]. Tension [LeB02-99, Jay01, KMK^{+}02, LRBH02, Wel03-36, KHN00, RS00a]. Tep1 [KLK^{+}01]. tER [KR03]. teratoma [BSW^{+}04]. teratomas [Tum04h]. Term [TTR00, Dov02-65, PLW^{+}04]. Terminal [BRM^{+}00, BHKL01, BHL^{+}01, TJS^{+}04, HGC02, HPS^{+}04, KGT^{+}02, MHH^{+}03, NMHH03, PBD^{+}04, RCY^{+}03, RFLT02, TFM04, ZCW^{+}03, ARK^{+}00, CFC^{+}00, HVT^{+}00, HSW00, HCD^{+}00, OLB^{+}00, WHS00, BCG03, HHS03, XBL^{+}03]. terminally [BLP^{+}02, CSP^{+}04, PMKV01, GH00, MCG^{+}00]. terminate [Dov04g]. termination [BS04b]. Termini [BSS00]. Terminus [EWSN00, HWBD^{+}01, FKG00, TMHP00, WS00]. territories [MPB02, MPG^{+}02]. Tes [GHS^{+}03]. testosterone [DME^{+}04]. testosterone-producing [DME^{+}04]. Tetanic [WXQ^{+}00]. tetanus [LS02, MAAZ^{+}00]. tether [DMA^{+}01, DKA01, PL01]. tethered [LMG04]. Tethering [SB01, FMF^{+}04, SBS^{+}02b, WSK^{+}03, WDFNN04, DSSWW00]. tethering/docking [WSK^{+}03]. Tethers [SD00]. Tetrahymena [SLG02, XHG^{+}00]. Tetramer [KHvOD00]. Tetraspan [SGO^{+}00]. tetraspanin [Hem01, KYS^{+}02, YKT^{+}04, SFSD00, TWBV^{+}01]. Tetraspanins [TTM^{+}03]. Tg [PDV^{+}00]. TGF [ACMR04, BDR^{+}03, CSD00, DCC^{+}02, HCC02a, HKO03, JLK^{+}02, KWSK^{+}04, LeB04-100, MCF^{+}02, MWC^{+}02, NH03b, PUK02, SMS^{+}04a, SFSD00, SSH^{+}04, WNM^{+}03, YCX^{+}01]. TGF- [ACMR04, BDR^{+}03, CSD00, DCC^{+}02, HKO03, KWSK^{+}04, MCF^{+}02, SFSD00, WNM^{+}03, YCX^{+}01]. TGN [BBSF01, MTG^{+}02, PPGN^{+}02, Sre04a, SLD^{+}02, SG00]. Tha4 [CM01]. Tha4-dependent [CM01]. thaliana [ARK^{+}00]. thapsigargin [GMY^{+}03]. thapsigargin-sensitive [GMY^{+}03]. Their [JRW^{+}01, NHB00, SZZ^{+}00, BPKK01, BDKM04, DSH^{+}03, Dov03-27, FJK^{+}04, FAF^{+}04, HBAF^{+}02, LeB02k, LeB02-78, LeB03c, LeB04-109, MDF01, PPWM04, Pro03, PBT^{+}02, RPZ^{+}02, RWCC01, Wel04-38, AKH00, MKST00, MHK^{+}01, PR00, SVG^{+}00, TNM^{+}00, WCGT^{+}00]. them [CS04]. there [Wel02-34, Der01]. Thereby [WTG01]. thermophila [SLG02, XHG^{+}00]. Thick [AP00, MKS^{+}02, RFLT02, RBV00]. Thin [MDF01, MPR^{+}03]. things [Dov01d, Dov03-29]. third [CGF^{+}04, SMSM00]. those [ISS^{+}04]. Three [Dov03-52, SPW00, BS04a, CSO^{+}04, HZS^{+}01, KEHAM^{+}02, WDS^{+}03, HAP^{+}00]. three-dimensional [BS04a, CSO^{+}04, HZS^{+}01, WDS^{+}03]. threshold [BLU^{+}04, WLPD04, YWW^{+}04]. Thrombin [CDWB01]. Thrombospondin [AS00, AKT01]. Thrombospondin-1 [AS00, AKT01]. thrombospondins [LCS^{+}02]. thrombus [NGK^{+}03]. Throughout [BMKA01]. Thylakoid [CM01, MC02]. thylakoids [MGMH03]. Thymosin [RRJ^{+}01]. thyroid [LeB04-81, SJW^{+}04]. Ti [MAAZ^{+}00]. Tia [KCL^{+}00]. Tia-1 [KCL^{+}00]. Tiam1 [BZSC00]. tickling [SB01]. tied [Dov04o, LeB04x]. tiered [LJK^{+}01]. Tight [ZLG00, BPPFM^{+}03, Dov01-34, Dov02-60, FHF^{+}02, KLZ04, NCMO^{+}02, PMU^{+}02, FFST01, GKSR00, LM00a, WHS00]. Tilt [Wel04-75, GOL^{+}01]. Tim [DSH^{+}00]. Tim13p [CLSK02]. Tim23p [DSH^{+}00]. Tim8p [CLSK02]. Time [LeB04-98, LMG04, LeB02o, LeB02-100, LeB03-41, LeB04o, LeB04-48, Les01f, RJyH02, RDP03, SKGC^{+}03, Tum03c, MPR^{+}01]. Timely [YKW^{+}00]. Times [BGFJ01]. timing [LeB02n, LeB02-58, LCI^{+}01, SK01, Gil01]. Timp [FWP^{+}00]. Timp-3 [FWP^{+}00]. Tiny [SP02, LeB04-96]. tip [GGF^{+}03, FWY01]. tips [GWL03]. Tir [CRP^{+}04]. Tissue [AKDS00, AKFB00, CLT^{+}01, WSL^{+}00, AFR01, BWV^{+}01, EGS^{+}04, PSD^{+}04a, Pro03, CHM^{+}01, CMMP00, IOLA^{+}00]. tissue-like [EGS^{+}04]. tissue-specific [AFR01]. tissues [CDK04, GBD^{+}00]. Titin [MA00, KK02, LeB02u, MKS^{+}02, SS01]. titin-like [KK02]. titinic [Wel02-59]. titins [Dov01m]. Tlg [BBSF01]. TLR1 [SLR^{+}03]. TLR2 [SLR^{+}03]. Tmod [Wel01-57]. Toc159 [HBV^{+}01]. Toc64 [SS00]. together [GGGK03, Les01e, MBH^{+}02, Tum04s, Wel01d, Wel02-53, Wel04-67, Wel04-74]. Tom [ATE^{+}01, KRR^{+}01]. Tom20 [KIK^{+}00a, YTM03]. Tom40 [ATE^{+}01]. Tomography [MÖS^{+}00]. tomosyn [SBT^{+}04]. too [LeB03-90]. Tooth [PQF^{+}00, RKF^{+}04]. Tooth-like [PQF^{+}00]. Topo [LeB02-101]. topogenesis [HBH^{+}04, KSF^{+}00]. Topoisomerase [Dov02-59, MSR04, AAD03, CH03, TCH^{+}02]. topoisomerases [CLB^{+}02]. topologically [KMLS04]. TOR [CK03, KFS^{+}00]. Tor-Mediated [KFS^{+}00]. TORrific [Wel01-58]. Tortoise [vEWS^{+}01]. Total [SGAS00]. touch [Dov03-44, Wel04-72]. Touching [Wel01-59]. tough [Dov02-61, Tum03n]. Toxic [Wel01-60, Tum03m, ZCW^{+}03]. toxicity [MZH^{+}02]. Toxin [BSD^{+}01, MZ00, ALC^{+}03, Dov03b, Kur03, LS02, RZB^{+}03, TR02, SHWH00]. Toxoplasma [DSN^{+}01, Tum03n, ARMB04, CSJ00, Dov01s, GGD^{+}04, HRM02, JR02, RVB^{+}01, SJ01b, SCS^{+}00]. Tpa [WSL^{+}00]. Tpa/Plasmin [WSL^{+}00]. Tpr [FGS^{+}02]. TPX2 [KSK^{+}02, WWKV00]. tracers [Zam04]. track [Dov02t]. Tracking [JPM^{+}00b, LeB04-99]. tracks [KLF04, LeB03z]. tract [KDH^{+}04, KPKY^{+}03, NMG04]. traction [SF01]. TRADD [IIN^{+}01, MTPT02]. Traffic [FS03, LeB02-102, PY03, Pel01, BRY^{+}01, Del03, Dov02o, JR02, LeB04-57, LYL^{+}04, SOH^{+}04, SPK^{+}01, SVT^{+}02, MLLA00, TSK^{+}00, WAPB^{+}00].

Traffic-independent [FS03]. trafficked [ASYL04]. Trafficking [HLB^{+}00, KNI^{+}04, KBGG04b, PMP^{+}03, AJ01, GTR^{+}03, KLE^{+}02, KEHAM^{+}02, RWK^{+}04, SDL02, WWBGG03, BRG^{+}00, CNBWN00, HLK01, Lit00, Mar01, MTV^{+}00, ZLG00]. Trail [MMDC00]. TRAILs [Wel02r]. traitorous [Tum04g]. Trans [HGP^{+}00, HLB^{+}00, MW04, BBP02, CEGZ^{+}04, FSK^{+}04, PPGN^{+}02, PNSJ01, YGWN01, WJG^{+}00]. Trans-Golgi [HGP^{+}00, HLB^{+}00, BBP02, CEGZ^{+}04, PPGN^{+}02, PNSJ01, YGWN01, WJG^{+}00]. trans-Golgi/TGN [PPGN^{+}02]. Trans-SNARE [MW04]. transactivation [TPW^{+}04, YSW02, CDWB01]. Transcribed [LeB03-95]. Transcriptase [GMZ^{+}00]. Transcription [CGY^{+}01b, Dov01-34, Dov02-60, IHK^{+}00, LeB03-96, Wel01-61, WND^{+}00, ZAE^{+}04, BvdWD^{+}04, BFSO^{+}04, EPN^{+}03, FNKH02, GCH03, HW04, HNK^{+}03, JKG^{+}02, JMG^{+}04a, KSC02, KMP02b, LeB04-78, Les01j, MPB02, MSS^{+}01, MWHM01, NH03b, PRJK01, PC01, SBS02a, SYY^{+}03, SGW^{+}02, Tum03b, WHM^{+}02, WKZ^{+}02, Wel02-61, WHP^{+}02, vdBCH^{+}04, ABP^{+}00, AW00, BKB^{+}01, HCD^{+}00, SRHV00, SKJ^{+}00, Wel01-28]. Transcription-dependent [ZAE^{+}04]. Transcriptional [ESS^{+}00, PTM^{+}01, Wel02-62, CARG^{+}04, DLPB03, FCM^{+}01, GCG^{+}01, GG04, JRL^{+}03, KWS^{+}02, SEW^{+}01, WWD^{+}04, TM00]. Transcriptionally [vMZM^{+}00, MAG^{+}04, WSS^{+}04]. Transcriptome [SMC^{+}02]. transcripts [ARM02, GBJ01, LeB04-99, Mar01]. transcytosis [dMMBK^{+}02]. transdifferentiation [Wel02-32]. transducer [MSS^{+}01, vWJK^{+}03]. transduces [YHT02]. Transducing [FSD00]. Transduction [ELO^{+}01, HTT^{+}02, OMiKF02, SS02, WW02, Wel02-51]. transendothelial [CSIK03, WLWB01, AML00]. Transfected [AGB^{+}00]. Transfer [PSWU00, CWG^{+}02, JMB^{+}04, RPS^{+}02, SP03, SN04]. transferase [BBDK^{+}04, RFCD02]. transferred [TR02]. Transferrin [SPO^{+}02]. transform [Wel01-59]. transformation [LeB03-73, LCG^{+}04, MJY^{+}04, NH03b, SSM^{+}04, YSC^{+}02, YSC^{+}21, GWG01, PSKK^{+}00]. Transformed [RGGL00, Wel03-42]. transforming [HIE^{+}01]. transgene [MPG^{+}03, WST01]. Transgenic [HGB^{+}00, SJA^{+}00, HSC01, CMMP00, PQF^{+}00, TMHP00, WFF^{+}01]. Transglutaminase [AKFB00, CLT^{+}01]. transient [CWG^{+}02, LL02, MGP^{+}02a, KHN00]. transiently [BPD^{+}04]. transit [RGM^{+}02, SRW^{+}04]. Transition [ESS^{+}00, GSW^{+}00, ABCK^{+}03, BPKR^{+}02, GNH^{+}04, KCWF02, MSR04, SHP01, AKW00, GLA00, KIK^{+}00b, MMH^{+}00, ZEtK^{+}00]. Transitions [AKH00, Nob00]. transits [BKS^{+}02]. translates [SKN^{+}03]. Translation [LHR04, Wel01-62, AZP^{+}02, BS02, CBL^{+}02, JON^{+}03, PP00, TSMS01]. translational [KWS^{+}02, Las03, WHAH03]. translationally [LeB02-59]. translocase [MGMH03, MC02, TVF^{+}03, KM00]. Translocated [SNL^{+}00]. translocates [MKM04, NPL04]. Translocation [ASAJ01, ESS^{+}00, VBH^{+}02, DB02, FBH03, FGSW03, LCRS01, PBT^{+}02, RZB^{+}03, SXD^{+}03, TMA^{+}04, WSC^{+}01, Wel04-58, GLK^{+}00, MPR^{+}01, RSG00].

Translocon [LeB03-97, Dov03-53, FBH03, NSLSK02, SSK^{+}03, KM00, SS00]. translocon-associated [FBH03]. translocons [SRB^{+}04]. Transmembrane [WGvA^{+}01, ABRA03, BWV^{+}01, FS03, HEW^{+}01, KPKY^{+}03, TCK^{+}03, WMA^{+}04a, AKW00, BHKL01, DAC00, SFSD00]. transmembranes [Wel04-38]. Transmigration [LSA^{+}00]. transmigratory [CS04]. Transmission [HIT^{+}02, SF01, Dov03-56, KKPB03, Wel03-58, FAAS00]. transmitter [Riz03]. transplants [BSW^{+}04]. Transport [DSV^{+}00, EWSN00, FSGDN^{+}00, GWBW00, HRE^{+}01, MHS^{+}00, RH00, Tum04-28, ATG^{+}03, AFK^{+}03, ATF^{+}04, BBP02, BCG03, Bro03, BKS^{+}02, BJ03, CiKBG03, CM01, DSV^{+}03, Dov01c, EMY^{+}04, FR01, HCC02b, HvdHG^{+}03, HMC^{+}01, HKBH03, JGR^{+}04, LS02, LeB03l, LRF^{+}02b, LDK^{+}03, LBS^{+}02, MTG^{+}02, MTT^{+}04, MR01a, MMPO^{+}01, NH03a, NOS^{+}01, OSNG04, PBD^{+}02, PFSG03, PNSJ01, QDG^{+}04b, RGM^{+}02, REM^{+}02, RCY^{+}03, RTFW02, SKF02, SCB02, SHM02, Sod02, SMH^{+}02, TKB^{+}04, Tum03k, VMH^{+}02, WSWSL04, Wel03q, XRH^{+}03, Zam04, ZW04, AGB^{+}00, AFN00, ARK^{+}00, BP00, IBKSP01, LCM00, MAAZ^{+}00, NKH^{+}00, OAR^{+}00, PPR^{+}00, PR00, SG00, TCS01, TOM00, WSDW^{+}00, WJG^{+}00, YW00]. transporter [EWM^{+}04, KWO^{+}00, WMT^{+}01]. Transporting [DAC00]. Transports [TYS^{+}00]. trap [Dov03-58, Dov03-53]. Trapp [WSFN00]. trapped [Dov02-27]. travel [LeB02-63, Wel02-36]. traveled [RT02]. Traveling [LeB02-103]. travels [LeB03-82]. traverse [MBN^{+}01]. treadmill [RSD^{+}04]. treated [ERS^{+}04]. treatment [CPN^{+}01, LeB02-110]. trefoil [BPKR^{+}02]. triad [IKS^{+}01, FKG00]. tribute [DV04]. Tric [MDJF00]. tricellular [STA03]. Trichocyte [WPJ^{+}00]. tricks [Dov02-45]. trigger [BGW^{+}04, CRP^{+}04, EBWC01, MC02, UHR^{+}03, Wel03-61, WAOC^{+}03, TN00a]. triggered [DC02a, MGP^{+}02b, RMC^{+}02, RRL^{+}00]. Triggering [Stu04, MSR04, PCC^{+}00]. triggers [ALC^{+}03, CDW^{+}03, LMGM^{+}02, MBH^{+}02, PPWM04, IAG^{+}00]. trip [Dov03-50, LeB02-104]. Triphosphatase [DAC00]. Triphosphatases [WYHP00]. trisphosphate [CVZ^{+}04]. Triton [NOS^{+}01, SS02]. Triton-insoluble [NOS^{+}01]. TrkA [MWAM01, MDQ^{+}03, SATA^{+}02]. TrkB [DFZ^{+}03, RLTC^{+}02, SNS^{+}04]. Trna [ASP^{+}00, KKK^{+}00]. Trophic [PDJ00]. tropomodulin [MPR^{+}03]. tropomodulin1 [FSCF^{+}03]. tropomodulin3 [FFSF03]. Tropomyosin [Dov02-61, OO02, KSNS^{+}04, MPR^{+}03]. trouble [Wel02-32]. TRP [DC02a, LM00b]. TRPC [VMS^{+}02]. Truncated [CGL^{+}01]. Truncations [BHKL01]. trunk [DRBF03]. Trypanin [HDH02]. Trypanosoma [HHM^{+}04]. Trypanosomatid [AERD^{+}01]. trypanosomes [HDH02]. tryptophan [UN03]. TSC1 [GGNK04, HFG^{+}04]. TSC1-2 [HFG^{+}04]. TSC1-binding [GGNK04]. TSC2 [GGNK04]. Tsg101 [PHS^{+}03]. Tsg101-recruiting [PHS^{+}03]. TSP [TDhL^{+}02]. TSP-1 [TDhL^{+}02]. TSR [Sil02]. tube [WSH^{+}04]. Tuberous [JHS^{+}02]. Tubes [Wel03-62, FWY01, MSF^{+}00]. Tubular [PPM^{+}00]. Tubules [FSY00]. Tubulin [Dov02-62, XHG^{+}00, BXR^{+}02, DWFAdL^{+}02, HOK^{+}02, HRM02, KM01, SLG02, BLC00, GMC^{+}00, ZKW^{+}00]. tubulin-specific [BXR^{+}02]. tubulogenesis [PPA^{+}03]. tug [Wel02m]. tug-of-war [Wel02m]. Tugging [Tum04-29]. Tumor [DSV^{+}00, Dov03-54, Dov03-55, HMN^{+}00, PTM^{+}01, RIDC01, SOH^{+}04, WFI^{+}04, CKW^{+}03, EPH^{+}03, GK03b, HFG^{+}04, IIN^{+}01, JHS^{+}02, LRWB04, RBD^{+}01, SSM^{+}04, UGKT^{+}02, WCBC04, WML^{+}03, WG03, XRP^{+}01, ZHDB04, BMG^{+}01, BZSC00, CWMO00, FWM^{+}01, PDW^{+}00, RSG01, LeB03a]. tumor-associated [EPH^{+}03]. tumorigenesis [EKdM^{+}04]. tumorigenic [MGP^{+}02b, Wel03-63, PSKK^{+}00]. Tumors [Sre04n, Wel04-76, LeB02-46, LeB04-93, Wel04g, Wel04w, ZLR^{+}03]. tunes [SDL^{+}03]. turgor [RRS03]. Turn [Dov02-63, LeB03-60, LeB04-50, Tum04g, Wel04-75, KO00]. turn-on [LeB04-50]. Turn-ons [Dov02-63]. Turning [Dov01-35, Tum03o, LeB02p, LeB04f, ZWSC02, RRJ^{+}01]. turnover [DBH^{+}01, DIR03, EOB^{+}04, GYL02, MR01a, XWL03]. turns [Jeg01e, LeB02-27, LeB02-73, Wel02j]. twice [Dov01t]. twin [MC02]. Twinfilin [WVY^{+}01, POW^{+}01a]. twist [Tum04i]. Two [CDTW00, DSH^{+}00, Dov01-36, Dov02-64, Dov03-56, FHM^{+}03, GHS00, IYT^{+}04, KNIO01, LeB03-98, LeB04-100, RVB^{+}01, TNM^{+}00, BHPN04, CG03, CFB^{+}03, Del03, DMLK04, GOV^{+}02, HFK^{+}03, IBP^{+}04, ISS^{+}04, LeB03i, LeB04-101, LJK^{+}01, MTG^{+}02, MN03, MSMK04, PWS^{+}01, RKKP02, SKF02, TOM01, Tum04p, Wel03-34, Mac00, NY00, NM00, TR00, WSDW^{+}00]. two-photon [IBP^{+}04]. two-tiered [LJK^{+}01]. two-way [Del03]. Type [CKFH00, JRW^{+}01, KIC^{+}01, ZZM^{+}03, ACBG04, BVH04, CdLvM^{+}04, CMS^{+}02, Cor02, CRH02, FAT^{+}02, FLS^{+}03, GFM^{+}04b, HDP^{+}01, IKS^{+}01, JJM^{+}02, KKW^{+}03, LDI^{+}03, LZS^{+}03, LVD^{+}04, MWSL^{+}03, MPG^{+}03, MRT^{+}01, RKF^{+}04, SSH^{+}04, TDhL^{+}02, UIY^{+}01, XRP^{+}01, ACE^{+}01, BT00, FFST01, HSW00, HAP^{+}00, KSF^{+}00, KH01, LSSL00, SC01b]. type-dependent [BVH04]. types [Dov04o, GFGP03]. Tyr247 [LWCKL01]. Tyr265 [LWCKL01]. Tyr31 [TSY^{+}02]. Tyr31/ [TSY^{+}02]. Tyrosine [CS01, CSP^{+}00, ESC^{+}01, GAC^{+}03, LDI^{+}03, RDH^{+}01, SNL^{+}00, CGM^{+}02, DFZ^{+}03, GJS^{+}03a, GSB^{+}01, HIN^{+}03, HEN^{+}01, KHC02, LBH^{+}02, LW03, MWF02, PPA^{+}03, SRG^{+}04, SF01, SKO04, WXD^{+}03, WAV^{+}04, WLO^{+}02, dVKS04, APLB00, BLPP01, KMB^{+}01, PBL^{+}00, SGPL^{+}00]. Tyrosine-Phosphorylated [CS01, GAC^{+}03].

U [HDL02]. U1a [HM00]. U2 [FM01, YSN^{+}01]. U2b [HM00]. U4 [GBOL03, SRKN03]. U4/U6 [SRKN03]. U6 [GBOL03, SRKN03]. Ubc3 [TBRG01]. Ubinuclein [ABP^{+}00]. Ubiquilin [MPSM00a]. ubiquitin [BHW02c, CSG01, DKAH04, KL04, LdVV^{+}02, MCU04, NB03a, PPWM04, TBRG01, YRC^{+}04, HLK01, LeB02-105, LeB04-102, LeB04-103, PPWM04]. ubiquitin-conjugating [PPWM04, TBRG01]. ubiquitin-related [NB03a]. ubiquitinated [BWK^{+}03]. ubiquitinates [KCWF02]. Ubiquitination [LBS^{+}02, DKAH04, HCC02b, LFM^{+}04, KO00]. UCS [LP04b]. Ufd1 [YMR03]. UGO1 [SJ01a]. Ulp1 [LH03]. Ultrastructural [GNDLS^{+}01]. Ultrastructure [PPM^{+}00]. umbrella [KDH^{+}04]. Unable [CGL^{+}01]. unambiguously [MTM^{+}02]. Unattached [Dov02-65]. unattracted [Wel02-42]. unc [Ono01, BKD^{+}04, RMG^{+}00, AP00]. UNC-112 [RMG^{+}00]. Unc-45 [AP00]. unc-78 [Ono01]. UNC-95 [BKD^{+}04]. Unc104 [ABCK^{+}03, Wor03]. Unc112 [Sch00]. uncaging [ACP^{+}02]. Uncertainty [LeB04-104]. Unchain [Sod02]. uncoating [Wel04n]. Uncooperative [LeB04-105]. Uncoupled [SMS00]. uncouples [MH02]. Uncoupling [TAD^{+}00]. Uncovering [WAC^{+}03]. undergo [DC03, FRM^{+}02, IBKSP01]. Undergoes [HCD^{+}00]. underlies [MEK^{+}04]. undifferentiated [BJM^{+}02]. undoes [LeB03t]. undone [LeB02w, LeB03-63]. unequal [LeB03-40]. unexpected [LG01, ZSS01]. Unfolded [NZHR01, TR02, RH04, UCY^{+}02, NSW00]. unfolding [Wel01-63, YHZ^{+}01]. uniformity [HHJS04]. Unique [SHA^{+}03, HZS^{+}01, SKF02, FLX^{+}00]. unitary [MKTW01]. unite [Dov03-60]. united [LeB03-39]. Universal [AERD^{+}01]. universe [Tum03i]. unknown [LeB03-47]. unleash [LeB03-49]. unliganded [WBU03]. unligated [Dov01i, SPB^{+}01]. unlike [TSL04]. unlikely [Wel02-63]. unmineralized [HBC^{+}03]. unperturbed [LKLD04]. unsticks [LeB03-72]. untested [Hel03]. untethered [Wel01-41]. until [HPE^{+}01, LMG04]. untranslated [GBJ01]. Untying [PSB00]. unusual [LL03b]. unwinding [LeB04-108]. up-regulates [MZT^{+}03]. up-regulation [BWA^{+}04]. uPAR [LeB03-49, SWE^{+}03]. uPARAP [ELNA^{+}03]. uPARAP/ [ELNA^{+}03]. upon [LBH^{+}02, MPBR03, REK^{+}03, SNL^{+}00, SHE^{+}02, SKK^{+}04, TNM^{+}03, TR02, YRC^{+}04]. Upregulated [BMF00]. Upregulation [ESS^{+}00, TLS^{+}01]. ups [Dov03-50, Krä00]. Upstream [DCM00]. uptake [BHW02b, ELNA^{+}03, MBH^{+}02, RMMP04]. urchin [HV03, RCL^{+}00]. Ure3 [STE^{+}01]. Urge [Dov03-57]. urinary [KDH^{+}04]. Urokinase [KH01, WTG01]. Uroplakin [DLT^{+}02, HDL^{+}00]. uroplakins [KDH^{+}04]. uropod [LKH^{+}04]. urothelial [DLT^{+}02, HDL^{+}00]. use [CSL^{+}03, GOV^{+}03, TOM01]. uses [Tum03n, WMA^{+}04a, WWD03, FKSJ01]. Using [Dov02-66, OGD03, SV03, MDJF00, MYH^{+}01, TMHP00]. UT [Wel02-51]. utilizes [HRV^{+}01]. utilizing [GGF^{+}03]. Utp [HSB00]. UTR [ARM02]. Utrophin [Dov02-67, GOV^{+}02, GBJ01, Wel01-51, AKK^{+}00].

v [EPH^{+}02, LWCKL01, BWW^{+}02, CDTW00, ICN^{+}03, KKT04, KJB^{+}02, MKTW01, RPTNM01, SCB02, Val03, WH03, Wel01-37, WWGK02, MPBR03, RAD^{+}02, BS04a, DSB^{+}02, GMRYM^{+}02, HFK^{+}03, JW04, MCF^{+}02, ZLVS02]. V-based [WWGK02]. v-mediated [BS04a]. v-Src [LWCKL01]. Va [REM^{+}02]. vab [BHL^{+}03]. vab-10 [BHL^{+}03]. Vac14p [BND^{+}02]. Vaccinia [HRV^{+}01]. Vacuolar [HMRH01, HS02, ARK^{+}00, WSE00]. Vacuole [BRB^{+}03, HLB^{+}00, KIO^{+}00, PSWU00, BMLU02, EWTW02, JPH^{+}01, KBGG04b, MW04, SJ01b, WMCW03, WSK^{+}03, JPRR00, KKS^{+}01, KHK01, PWU00, SM00]. Vacuole/Lysosome [HLB^{+}00]. vacuoles [FJM^{+}04, TOM00]. Vadis [Mel00c]. vagal [DRBF03]. Vam6p [CHA^{+}01]. Vam7p [BMLU02]. Vamp [MAAZ^{+}00]. Vamp3 [BPS^{+}00]. Vamp3-Containing [BPS^{+}00]. variant [RDN^{+}03]. variants [vdFKK^{+}02, IHK^{+}00]. Various [SMS^{+}01b]. vasa [KPB^{+}00]. vascular [CS04, CLG^{+}03, EPH^{+}02, LZC^{+}03, SKN^{+}03, Tum03i, WFI^{+}04, AML00, ECV^{+}00, HGK^{+}01]. vascularization [KPL^{+}02, BMG^{+}01]. Vasodilator-Stimulated [KSK^{+}00]. VASP [KSK^{+}00, LeB04-106, SRK^{+}03, SAG^{+}01]. Vault [KLK^{+}01]. Vault-Associated [KLK^{+}01]. Vav [GCO^{+}04, VBH^{+}02]. Vav-dependent [VBH^{+}02]. Vav1 [VBR^{+}01b]. Vav1/ [VBR^{+}01b]. Vav2 [MC01]. VCAM [BYMS^{+}02]. VCAM-1 [BYMS^{+}02]. VCIP135 [UJK^{+}02, WSWM04b]. VDAC [MH01]. VDAC-dependent [MH01]. VEGF [CBL^{+}02, Dov03-37, EGC^{+}03, GGF^{+}03, LZC^{+}03, LeB02l, LeB03-98, WCBC04, ZVPK03]. VEGF-induced [LZC^{+}03]. VEGF-mediated [WCBC04]. velocity [SCB02]. Venus [Dov03-58]. version [Dov03-32]. versus [MB03c]. Vertebrate [FJK^{+}04, KRS^{+}02, PBD^{+}02, SSP^{+}03, VSO^{+}01, WBT^{+}03, YDPK04, PMK^{+}00, Rut00, SVG^{+}00]. Vertebrates [KR01]. vertex [FJM^{+}04, WMCW03]. very [DBH^{+}01, TWS^{+}00b]. Vesicle [Tum04-30, BBP02, BET^{+}03, BJ03, CEGZ^{+}04, FCBH01, HMG03, HCCB03, LeB02-41, LeB02-64, LeB04-66, MPV^{+}01, NB03b, PL01, SCB02, SBS^{+}02b, SDL02, Tum03c, WDFNN04, XH04, BMS^{+}00a, DSSWW00, HHHJ00, MAAZ^{+}00, MTV^{+}00, NS01, NKH^{+}00, PWU00, SM00, TOM00]. vesicle-associated [MPV^{+}01]. vesicle-tethering [SBS^{+}02b]. Vesicles [BHPN04, EWSN00, LeB02-107, LeB02-106, AQHO03, GK03a, IBS^{+}02, JAS02, JCPWS01, LOS^{+}01, LeB03-68, MJS02, MMPO^{+}01, MPV^{+}01, NOM^{+}04, OMWSN02, SVT^{+}02, SMH^{+}02, TSL^{+}03, VCGB^{+}02, Wel02-64, Wel03g, YLG^{+}02, BPS^{+}00, BMC00, BSW^{+}00, CB00, KWO^{+}00, MNHR00, OAR^{+}00, PPR^{+}00, SKR^{+}00, TYS^{+}00, TCR00, WMT^{+}01]. Vesicoureteral [HDL^{+}00]. Vesicular [FR01, KWO^{+}00, MAAZ^{+}00, WMT^{+}01]. vesiculate [Wel03-51]. Vesiculotubular [EHCC^{+}00]. vessel [LeB03-98]. vessels [CPA^{+}03]. Vfl1 [SLT^{+}01]. VHS [HLB^{+}00]. VI [Cra00]. via [ALC^{+}03, ABF^{+}03, BBMS03, BSD^{+}01, BKZ^{+}03, EOB^{+}04, FLS^{+}03, FMC^{+}04, FNFL03, GQI^{+}02, HFG^{+}04, HSB00, HHS03, HP03, KFS^{+}00, KHC02, KEGDQ01, KRR^{+}01, KKK^{+}02, LMGM^{+}02, LSMS03, LM00a, MGL^{+}00, MKS^{+}02, MTV^{+}00, MTPT02, NMH^{+}04, NDS^{+}02, NLBK00, OSNG04, RDC^{+}04, RLC00, RBBA00, SMS^{+}04a, SWH^{+}02, SRSW04, SMZ^{+}03, TKB^{+}04, TEB^{+}03, WSDW^{+}00, WGF^{+}00, YJS^{+}03, Zam04]. Viability [HES00, EKdM^{+}04, LRF^{+}02a, BWN^{+}01]. Vid [BMC00]. view [Hoy01, LeB03-99, Wel02k]. Vigna [TOM01]. VII [CKF^{+}03, MCH^{+}00]. vimentin [LeB02-33, sGS01]. Vinculin [SPA^{+}04, DBB02b]. viral [TE01, ABP^{+}00]. Virus [Wel01-64, CdLvM^{+}04, CDW^{+}03, HRV^{+}01, LeB02-68, OBG^{+}03, SLB02, Wel01w, Wel01-33, Wel03-65, Wel04-62, Gau00, HWW01]. virus-1 [SLB02]. viruses [BS02]. Visualization [ASP^{+}00, BBP02, BC03, CME^{+}02, HHF^{+}00, KKR^{+}00, KLK^{+}04, MVL^{+}02, PMPH03]. Visualized [SDN^{+}00, YOK^{+}03]. Visualizing [Wel02-64]. Vitamin [PGSE^{+}01, Dov02-68]. vitro [BMY^{+}01, EGWK^{+}01, HGC02, HDL02, KK02, SVI^{+}04, SSW^{+}01a, VYW^{+}03, WCIN04, BMS^{+}00a, BMKA01, DR00b, LLH^{+}01, TCR00, WPJ^{+}00]. Vitronectin [BMG^{+}01, KH01]. Viva [LeB02-108]. Vivo [GWBW00, AMF01, CHA^{+}01, DDL^{+}04, DFJ^{+}02, DJ03, DHK^{+}04, HGC02, JKG^{+}02, KK02, KAiK^{+}02, LHK02, LL02, LWZ03, LZS^{+}03, MWAM01, MJY^{+}04, PLR03, PPK^{+}01, RMMP04, SHW01, SGC^{+}02, SK01, UJK^{+}02, XRP^{+}01, ZW04, ZZM^{+}03, BMG^{+}01, BHFL01, BGS00, CLAC00, CMMP00, NS01, PGSL00, SBZ^{+}00, SHS^{+}00, SRHV00, SKJ^{+}00, TCR00, WBM^{+}00]. Vohwinkel [SJA^{+}00]. voltage [JB01, JJM^{+}02, RPS^{+}02, SMS^{+}01b]. voltage-dependent [JJM^{+}02, RPS^{+}02]. voltage-gated [JB01]. Vpr [Wel01-65]. vps [BHW02c, WSE00]. Vps1p [HvdBP^{+}01]. Vps27 [BWK^{+}03, KSBE03]. Vps27-Hse1 [BWK^{+}03]. VPS34 [FCBH01, KNIO01, WE02]. Vps34p [TNBH01]. Vps35p [NHB00]. Vps45p [BJ03]. Vrp1p [EKT^{+}00]. Vulgaris [CdBB^{+}01].

W [Nel00, Dov03d, WAOC^{+}03]. walking [Dov02-56, Dov03-62]. wall [JJM^{+}02, NM02, TBTN01, TF00]. Waltz [Les02e]. war [Wel02m, Wel03l]. warning [Dov03q]. Warts/Lats1 [HMN^{+}00]. Wasp [BYLA^{+}01, BC04, ZH04, FSM^{+}01, HP00, LBP00, RyHHK00, TRC^{+}00, Zig00]. WASP/WAVE [BC04]. WASPs [Wel03-64]. Watching [Dov02-69, Dov03-59]. wave [Wel03-39, TSK^{+}00, BC04]. wavelength [SAWS02]. waves [Dov02-69, Tum03f, Wel02d, Wel04e, Wel04-42]. way [Del03, Dov02e, Dov02-48, Dov02-47, Dov03-45, Dov04j, LeB02-78, LeB03-61, LeB04e, Les01l, Wel01-57]. ways [Dov01-36, Dov02-29, Dov03-64, LeB03-98, Les01d]. WD [TONN02, TN00a]. weak [Dov03-63]. Weaklings [Dov03-60]. Wee1 [SEM^{+}00, YDPK04]. Wee1-Regulated [SEM^{+}00]. weight [Tum03g]. Werner [CGY01a]. West [LeB04-107]. Where [DA01, Lai03f, LeB03-102, MKT01]. whereas [BJM^{+}02]. whether [ERS^{+}04]. which [BRB^{+}01, ITF^{+}02, LeB04e, LJK^{+}01, MZT^{+}03, MWE^{+}03, HJSM00, ZEtK^{+}00]. while [DDL^{+}04, Wel04-49]. Whose [KHN00, SME^{+}00]. wide [LeB02-51]. Widening [LeB02-110]. wild [MWSL^{+}03]. wild-type [MWSL^{+}03]. wildly [Wel03-44]. will [Dov02-33, LeB02-81, RO04, Wel02-36]. wind [Wel01-36]. Winding [LeB04-108]. wing [LeB03-50]. Wingless [Wel01-66]. wings [LeB03-41]. winner [Tum02a]. wins [Wel04-68]. Wish [FSM^{+}01]. Wiskott [BYLA^{+}01, FSM^{+}01, HP00]. withdrawal [DBL^{+}02, PKR^{+}02, FCL^{+}00]. within [CSO^{+}04, DR00a, FGS^{+}02, FCBH01, MPG^{+}02, PLP02, RWH02, SPW00, SWB03, Wel02-67, WDFNN04, XRP^{+}01, ZN01]. without [CDWB01, DPO^{+}04, Dov03-43, GJB^{+}00, LeB02-92, Les01m, MBN^{+}01, Tum04h, Wel01-36, qZC01]. Wnt [Dov03-64, Wel02-68, YSC^{+}02, CGY^{+}01b, CKS^{+}04, DA01, TN00b, THK^{+}00, ADL^{+}03, BvdWD^{+}04, DMC^{+}03, iHGK^{+}02, KPL^{+}02, TJC^{+}03, WBT^{+}03, YSC^{+}21]. WNT-1 [CGY^{+}01b]. Wnt-5 [WBT^{+}03]. Wnt-5/pipetail [WBT^{+}03]. Wnt-5a [TJC^{+}03]. WNT-signaling [CKS^{+}04]. WNT/ [THK^{+}00, WBT^{+}03]. Wnts [WM03]. work [Dov02-31]. worm [LeB02-108]. Worms [Wel03-66]. wound [CAW^{+}04, FLWMG02, KHLW02, MDT^{+}01, MCG^{+}03, RDC^{+}04, Wel01-67, WC03, THE^{+}00]. Wounds [LeB04-110, MB01]. wraps [Wel02p]. write [Wel04-44]. Wrong [Dov03-65, LeB02-73].

X [HOvD^{+}00, RRL^{+}00, BLPP01, CW01, CW02, CNJ01, GJ00, PTH^{+}04, RME^{+}00, ZZM^{+}03, KSS^{+}03]. Xbp1 [SJMB04]. Xenopus [AKH00, BGFJ01, BKDH01, CGY01a, DJ03, FFY^{+}00, GTD^{+}02, HMG03, IOLA^{+}00, KTY^{+}00, KHLW02, LGP00, LYKH00, MH02, MSC^{+}03, MB01, MPB^{+}04, PMBC^{+}00, SHCM03, SEM^{+}00, SVLM02, SM03b, TMHP00, VTH^{+}02, VAHV00, WSDW^{+}00, WDMH03, WWKV00, YSN^{+}01]. Xgrip210 [ZKW^{+}00]. XIAP [PSK^{+}03, SHE^{+}02, ESH^{+}01]. XIIIb [NOS^{+}01, PLL^{+}00]. Xist [CNJ01]. XIV [GGD^{+}04]. XIX [SML^{+}04]. XMAP215 [SHCM03, vBDH03]. Xmus101 [VTH^{+}02]. Xpo1p [GDHS01]. Xpo1p-Mediated [GDHS01]. XVIII [ACE^{+}01, KLGC^{+}01]. Xylogen [Wel04-77].

Y-box [BGM03]. Yarrowia [GKN^{+}03, TCR00, TNW^{+}02]. YDR479c [VTGT^{+}03]. Year [Mel00b]. Yeast [AMG^{+}01, Dov04p, FMC^{+}04, LeB04-111, LBP00, NHB00, PSWU00, RRS03, SMWG00, WSK^{+}03, WK04, YCB04, ARQ^{+}04, BN02, BPC03, BMLU02, BVH04, Bre03, BTVB03, CMM^{+}02, CFB^{+}03, CNT03, DC02a, DPB03, Dov02-49, EWTW02, FR01, GES04, GCT^{+}04, GKN^{+}03, HS02, HJL^{+}04, HGP^{+}04, HBAF^{+}02, IKS^{+}02, JGW02, Kil03, KYW^{+}04, KKTP03, KJB^{+}02, KMS^{+}04, LeB02-62, LeB02-96, LeB04h, LJK^{+}01, LDP02, LOC03, LRF^{+}02b, LP04b, MHT^{+}04, MRC^{+}02, MYC^{+}02, MZH^{+}02, MW04, MJV^{+}03, MZR^{+}04, MMBM04, OMWSN02, POW^{+}01a, PPP^{+}01, PNSJ01, RFCD02, RMW03, SCPP02, SRW^{+}02, SHP01, TMG03, Tum04a, TBW^{+}04, UN03, VT04, WMCW03, Wel04u, WCA^{+}03, WLH^{+}04, YITe03, YHB^{+}04, dSAH02, vBDH03, AOC01, BP00, BHM^{+}00, CDTW00, CD00, DG01, DSP^{+}01, DR00a, FK01, FMP^{+}00, FSKS00, GCN00, HW00a, HLRG01, JSCR01, LSSL00, LGM^{+}01, iNFK^{+}01, NWT^{+}01, PMSB01]. Yeast [PLH^{+}01, RPTNM01, RAS^{+}00, RHM00, SCM^{+}00, SPH^{+}00, TMD^{+}01, yZCKA01]. Yersinia [Cor02]. YHR150w [VTGT^{+}03]. YidC [DK04, NSK04, vdLBK^{+}04]. Yip1p [HCCB03]. young [Wel03z]. youth [Wel02-69]. Ypk [dSAH02]. Ypt [PSWU00]. Ypt1p [SDL02, WSFN00]. Ypt32 [OMWSN02]. Ypt7 [WSE00]. YSK1 [PSD^{+}04b]. Yt521 [NHS00]. Yt521-B [NHS00].

Z [BMM^{+}01, OLB^{+}00, ZCH^{+}01]. Z-line [ZCH^{+}01]. ZBP1 [FHM^{+}03, LeB03-82]. Zebrafish [KPB^{+}00, MMBB03, MYO^{+}04, FYHH00]. Zinc [EWM^{+}04]. ZIP [KI04]. Zip1 [DR00a]. zipper [Dov02-70]. zippering [MWM^{+}02]. Zn [HCD^{+}00, UTH^{+}02]. ZNF265 [AvdWM^{+}01]. ZO [BGM03, WHS00]. ZO-1 [BGM03]. Zo-3 [WHS00]. Zombie [Sre04o]. ZONAB [BGM03]. zone [ACBG04, SML^{+}04, TRMI^{+}04]. Zones [Wel01-67]. Zonulae [FFST01]. Zyxin [HMN^{+}00, GHS^{+}03]. zyxin-dependent [GHS^{+}03].

References

Azuma:2003:SRT

[AAD03] Yoshiaki Azuma, Alexei Arnaoutov, and Mary Dasso. SUMO-2/ 3 regulates topoisomerase II in mitosis. Journal of Cell Biology, 163(3):477–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/3/477.

Andoniou:2004:NCB

[AAM^{+}04] Christopher E. Andoniou, Daniel M. Andrews, Mitali Manzur, Paola Ricciardi-Castagnoli, and Mariapia A. Degli-Esposti. A novel checkpoint in the Bcl-2–regulated apoptotic pathway revealed by murine cytomegalovirus infection of dendritic cells. Journal of Cell Biology, 166(6):827– ??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/6/827.

Alory:2000:MBR

[AB00] Christelle Alory and William E. Balch. Molecular Basis for Rab Prenylation. Journal of Cell Biology, 150(1):89–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
1/89.

Al-Bassam:2003:DCK

[ABCK^{+}03] Jawdat Al-Bassam, Yujia Cui, Dieter Klopfenstein, Bridget O. Carragher, Ronald D. Vale, and Ronald A. Milligan. Distinct conformations of the kinesin Unc104 neck regulate a monomer to dimer motor transition. Journal of Cell Biology, 163(4):743–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/4/743.

Arakawa:2003:CAE

[ABF^{+}03] Yoshiki Arakawa, Haruhiko Bito, Tomoyuki Furuyashiki, Takahiro Tsuji, Sayaka Takemoto-Kimura, Kazuhiro Kimura, Kazuhiko Nozaki, Nobuo Hashimoto, and Shuh Narumiya. Control of axon elongation via an SDF-1α /Rho/mDia pathway in cultured cerebellar granule neurons. Journal of Cell Biology, 161(2):381–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/381.

Al-Bassam:2002:MTB

[ABOS^{+}02] Jawdat Al-Bassam, Rachel S. Ozer, Daniel Safer, Shelley Halpain, and Ronald A. Milligan. MAP2 and tau bind longitudinally along the outer ridges of microtubule protofilaments. Journal of Cell Biology, 157(7):1187–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/7/1187.

Aho:2000:UNN

[ABP^{+}00] Sirpa Aho, Monique Buisson, Tiina Pajunen, Young W. Ryoo, Jean-François Giot, Henry Gruffat, Alain Sergeant, and Jouni Uitto. Ubinuclein, a Novel Nuclear Protein Interacting with Cellular and Viral Transcription Factors. Journal of Cell Biology, 148(6):1165–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/6/1165.

Anandatheerthavarada:2003:MTN

[ABRA03] Hindupur K. Anandatheerthavarada, Gopa Biswas, Marie-Anne Robin, and Narayan G. Avadhani. Mitochondrial targeting and a novel transmembrane arrest of Alzheimer’s amyloid precursor protein impairs mitochondrial function in neuronal cells. Journal of Cell Biology, 161(1):41–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/1/
41.

Adames:2000:MIC

[AC00] Neil R. Adames and John A. Cooper. Microtubule Interactions with the Cell Cortex Causing Nuclear Movements in Saccharomyces cerevisiae. Journal of Cell Biology, 149 (4):863–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/4/863.

Aguirre:2004:NEC

[ACBG04] Adan A. Aguirre, Ramesh Chittajallu, Shibeshih Belachew, and Vittorio Gallo. NG2-expressing cells in the subventricular zone are type C-like cells and contribute to interneuron generation in the postnatal hippocampus. Journal of Cell Biology, 165(4):575–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/4/575.

Ackley:2001:NED

[ACE^{+}01] Brian D. Ackley, Jennifer R. Crew, Harri Elamaa, Tania Pihlajaniemi, Calvin J. Kuo, and James M. Kramer. The Nc1/ endostatin domain of Caenorhabditis elegans Type XVIII collagen affects cell migration and axon guidance. Journal of Cell Biology, 152(6):1219–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/6/1219.

Annes:2004:IMA

[ACMR04] Justin P. Annes, Yan Chen, John S. Munger, and Daniel B. Rifkin. Integrin α_{V} β_{6}-mediated activation of latent TGF-β requires the latent TGF-β binding protein-1. Journal of Cell Biology, 165(5):723–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/5/723.

Ashby:2002:LCU

[ACP^{+}02] Michael C. Ashby, Madeleine Craske, Myoung Kyu Park, Oleg V. Gerasimenko, Robert D. Burgoyne, Ole H. Petersen, and Alexei V. Tepikin. Localized Ca^{2+} uncaging reveals polarized distribution of Ca^{2+}-sensitive Ca^{2+} release sites. Journal of Cell Biology, 158(2):283–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/2/283.

Abeliovich:2000:DAB

[ADKK00] Hagai Abeliovich, William A. Dunn, John Kim, and Daniel J. Klionsky. Dissection of Autophagosome Biogenesis into Distinct Nucleation and Expansion Steps. Journal of Cell Biology, 151(5):1025–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/5/1025.

Ai:2003:QRS

[ADL^{+}03] Xingbin Ai, Anh-Tri Do, Olga Lozynska, Marion Kusche-Gullberg, Ulf Lindahl, and Charles P. Emerson. QSulf1 remodels the 6-O sulfation states of cell surface heparan sulfate proteoglycans to promote Wnt signaling. Journal of Cell Biology, 162(2):341–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/2/341.

Avery:2000:CFS

[AEL^{+}00] Julia Avery, Darren J. Ellis, Thorsten Lang, Phillip Holroyd, Dietmar Riedel, Robert M. Henderson, J. Michael Edwardson, and Reinhard Jahn. A Cell–Free System for Regulated Exocytosis in Pc12 Cells. Journal of Cell Biology, 148(2): 317–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/2/317.

Abu-Elneel:2001:ILU

[AERD^{+}01] Kawther Abu-Elneel, Derrick R. Robinson, Mark E. Drew, Paul T. Englund, and Joseph Shlomai. Intramitochondrial Localization of Universal Minicircle Sequence–Binding Protein, a Trypanosomatid Protein That Binds Kinetoplast Minicircle Replication Origins. Journal of Cell Biology, 153 (4):725–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/4/725.

Aridor:2001:SGC

[AFB^{+}01] Meir Aridor, Kenneth N. Fish, Sergei Bannykh, Jacques Weissman, Theresa H. Roberts, Jennifer Lippincott-Schwartz, and William E. Balch. The Sar1 Gtpase Coordinates Biosynthetic Cargo Selection with Endoplasmic Reticulum Export Site Assembly. Journal of Cell Biology, 152(1):213–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
152/1/213.

Ang:2003:RGS

[AFK^{+}03] Agnes Lee Ang, Heike Fölsch, Ulla-Maija Koivisto, Marc Pypaert, and Ira Mellman. The Rab8 GTPase selectively regulates AP-1B-dependent basolateral transport in polarized Madin–Darby canine kidney cells. Journal of Cell Biology, 163(2):339–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/2/339.

Adachi:2000:NEM

[AFN00] Makoto Adachi, Makoto Fukuda, and Eisuke Nishida. Nuclear Export of Map Kinase (ERK) Involves a Map Kinase Kinase (Mek–Dependent) Active Transport Mechanism. Journal of Cell Biology, 148(5):849–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/5/849.

Allen:2001:RHS

[AFR01] Benjamin L. Allen, Mark S. Filla, and Alan C. Rapraeger. Role of heparan sulfate as a tissue-specific regulator of FGF-4 and FGF receptor recognition. Journal of Cell Biology, 155(5):845–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/5/845.

Aitchison:2003:II

[AG03] John D. Aitchison and Timothy Galitski. Inventories to insights. Journal of Cell Biology, 161(3):465–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/3/
465.

Ackerley:2000:GSA

[AGB^{+}00] Steven Ackerley, Andrew J. Grierson, Janet Brownlees, Paul Thornhill, Brian H. Anderton, P. Nigel Leigh, Christopher E. Shaw, and Christopher C. J. Miller. Glutamate Slows Axonal Transport of Neurofilaments in Transfected Neurons. Journal of Cell Biology, 150(1):165–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/1/165.

Ahmad:2001:CSR

[AH01] Kami Ahmad and Steven Henikoff. Centromeres Are Specialized Replication Domains in Heterochromatin. Journal of Cell Biology, 153(1):101–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/1/101.

Arighi:2004:RMR

[AHA^{+}04] Cecilia N. Arighi, Lisa M. Hartnell, Ruben C. Aguilar, Carol R. Haft, and Juan S. Bonifacino. Role of the mammalian retromer in sorting of the cation-independent mannose 6-phosphate receptor. Journal of Cell Biology, 165(1): 123–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/1/123.

Assaad:2001:CGK

[AHMJ01] Farhah F. Assaad, Yoann Huet, Ulrike Mayer, and Gerd Jürgens. The Cytokinesis Gene KEULE Encodes a Sec1 Protein That Binds the Syntaxin Knolle. Journal of Cell Biology, 152(3):531–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/3/531.

Almeida:2000:MSS

[AIH^{+}00] Eduardo A. C. Almeida, Duško Ilić, Qin Han, Christof R. Hauck, Fang Jin, Hisaaki Kawakatsu, David D. Schlaepfer, and Caroline H. Damsky. Matrix Survival Signaling. Journal of Cell Biology, 149(3):741–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/149/3/741.

Aplin:2001:RNT

[AJ01] Andrew E. Aplin and R. L. Juliano. Regulation of nucleocytoplasmic trafficking by cell adhesion receptors and the cytoskeleton. Journal of Cell Biology, 155(2):187–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
2/187.

Akassoglou:2000:TPA

[AKDS00] Katerina Akassoglou, Keith W. Kombrinck, Jay L. Degen, and Sidney Strickland. Tissue plasminogen activatormediated fibrinolysis protects against axonal degeneration and demyelination after sciatic nerve injury. Journal of Cell Biology, 149(5):1157–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/5/1157.

Akimov:2000:TTI

[AKFB00] Sergey S. Akimov, Dmitry Krylov, Laurie F. Fleischman, and Alexey M. Belkin. Tissue Transglutaminase Is an Integrin– Binding Adhesion Coreceptor for Fibronectin. Journal of Cell Biology, 148(4):825–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/148/4/825.

Arnal:2000:STM

[AKH00] Isabelle Arnal, Eric Karsenti, and Anthony A. Hyman. Structural Transitions at Microtubule Ends Correlate with Their Dynamic Properties in Xenopus Egg Extracts. Journal of Cell Biology, 149(4):767–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/149/4/767.

Arya:2004:MRF

[AKH^{+}04] Ranjana Arya, Vishram Kedar, Jae Ryoung Hwang, Holly McDonough, Hui-Hua Li, Joan Taylor, and Cam Patterson. Muscle ring finger protein-1 inhibits PKCϵ activation and prevents cardiomyocyte hypertrophy. Journal of Cell Biology, 167(6):1147–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/6/1147.

Adams:2000:ASL

[AKK^{+}00] Marvin E. Adams, Neal Kramarcy, Stuart P. Krall, Susana G. Rossi, Richard L. Rotundo, Robert Sealock, and Stanley C. Froehner. Absence of α-Syntrophin Leads to Structurally Aberrant Neuromuscular Synapses Deficient in Utrophin. Journal of Cell Biology, 150(6):1385–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
150/6/1385.

Akazawa:2004:NLP

[AKM^{+}04] Hiroshi Akazawa, Sumiyo Kudoh, Naoki Mochizuki, Noboru Takekoshi, Hiroyuki Takano, Toshio Nagai, and Issei Komuro. A novel LIM protein Cal promotes cardiac differentiation by association with CSX/NKX2-5. Journal of Cell Biology, 164(3):395–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/3/395.

Adams:2001:RSC

[AKT01] Josephine C. Adams, Nina Kureishy, and Amanda L. Taylor. A Role for Syndecan-1 in Coupling Fascin Spike Formation by Thrombospondin-1. Journal of Cell Biology, 152(6): 1169–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/6/1169.

Armstrong:2000:TDI

[AKW00] R. Todd Armstrong, Anna S. Kushnir, and Judith M. White. The Transmembrane Domain of Influenza Hemagglutinin Exhibits a Stringent Length Requirement to Support the Hemifusion to Fusion Transition. Journal of Cell Biology, 151(2): 425–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/2/425.

Amara:2003:AHH

[AL03] Ali Amara and Dan R. Littman. After Hrs with HIV. Journal of Cell Biology, 162(3):371–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/3/371.

Abrami:2003:ATT

[ALC^{+}03] Laurence Abrami, Shihui Liu, Pierre Cosson, Stephen H. Leppla, and F. Gisou van der Goot. Anthrax toxin triggers endocytosis of its receptor via a lipid raft–mediated clathrin-dependent process. Journal of Cell Biology, 160(3): 321–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/3/321.

Anderson:2002:CCD

[ALEH02] David E. Anderson, Ana Losada, Harold P. Erickson, and Tatsuya Hirano. Condensin and cohesin display different arm conformations with characteristic hinge angles. Journal of Cell Biology, 156(3):419–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/3/419.

Alahari:2000:NNP

[ALJ00] Suresh K. Alahari, Jung Weon Lee, and Rudy L. Juliano. Nischarin, a Novel Protein That Interacts with the Integrin α5 Subunit and Inhibits Cell Migration. Journal of Cell Biology, 151(6):1141–??, December 2000. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/6/1141.

Abrami:2004:MIA

[ALP^{+}04] Laurence Abrami, Margaret Lindsay, Robert G. Parton, Stephen H. Leppla, and F. Gisou van der Goot. Membrane insertion of anthrax protective antigen and cytoplasmic delivery of lethal factor occur at different stages of the endocytic pathway. Journal of Cell Biology, 166(5):645–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
5/645.

Arin:2001:FAM

[ALWR01] Meral J. Arin, Mary Ann Longley, Xiao-Jing Wang, and Dennis R. Roop. Focal activation of a mutant allele defines the role of stem cells in mosaic skin disorders. Journal of Cell Biology, 152(3):645–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/3/645.

Aikawa:2003:ARP

[AM03] Yoshikatsu Aikawa and Thomas F. J. Martin. ARF6 regulates a plasma membrane pool of phosphatidylinositol(4,5)bisphosphate required for regulated exocytosis. Journal of Cell Biology, 162(4):647–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/4/647.

Ahimou:2004:AFN

[AMBW04] François Ahimou, Lee-Peng Mok, Boris Bardot, and Cedric Wesley. The adhesion force of Notch with Delta and the rate of Notch signaling. Journal of Cell Biology, 167(6): 1217–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/6/1217.

Adams:2001:ERD

[AMEC01] Richard R. Adams, Helder Maiato, William C. Earnshaw, and Mar Carmena. Essential Roles of Drosophila Inner Centromere Protein (Incenp) and Aurora B in Histone H3 Phosphorylation, Metaphase Chromosome Alignment, Kinetochore Disjunction, and Chromosome Segregation. Journal of Cell Biology, 153(4):865–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/4/865.

Adams:2001:VRS

[AMF01] Marvin E. Adams, Heather A. Mueller, and Stanley C. Froehner. In vivo requirement of the α-syntrophin PDZ domain for the sarcolemmal localization of nNOS and aquaporin-4. Journal of Cell Biology, 155(1):113–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
1/113.

Adamo:2001:YCF

[AMG^{+}01] Joan E. Adamo, John J. Moskow, Amy S. Gladfelter, Domenic Viterbo, Daniel J. Lew, and Patrick J. Brennwald. Yeast Cdc42 functions at a late step in exocytosis, specifically during polarized growth of the emerging bud. Journal of Cell Biology, 155(4):581–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/4/581.

Allport:2000:MIR

[AML00] Jennifer R. Allport, William A. Muller, and Francis W. Luscinskas. Monocytes Induce Reversible Focal Changes in Vascular Endothelial Cadherin Complex during Transendothelial Migration under Flow. Journal of Cell Biology, 148(1): 203–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/1/203.

Arredondo:2002:CRN

[ANC^{+}02] Juan Arredondo, Vu Thuong Nguyen, Alexander I. Chernyavsky, Dani Bercovich, Avi Orr-Urtreger, Wolfgang Kummer, Katrin Lips, Douglas E. Vetter, and Sergei A. Grando. Central role of α7 nicotinic receptor in differentiation of the stratified squamous epithelium. Journal of Cell Biology, 159(2): 325–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/2/325.

Andrews:2002:LPM

[And02] Norma W. Andrews. Lysosomes and the plasma membrane. Journal of Cell Biology, 158(3):389–??, August 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/3/
389.

Adames:2001:SMS

[AOC01] Neil R. Adames, Jessica R. Oberle, and John A. Cooper. The Surveillance Mechanism of the Spindle Position Checkpoint in Yeast. Journal of Cell Biology, 153(1):159–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
1/159.

Akune:2002:IRS

[AOH^{+}02] Toru Akune, Naoshi Ogata, Kazuto Hoshi, Naoto Kubota, Yasuo Terauchi, Kazuyuki Tobe, Hideko Takagi, Yoshiaki Azuma, Takashi Kadowaki, Kozo Nakamura, and Hiroshi Kawaguchi. Insulin receptor substrate-2 maintains predominance of anabolic function over catabolic function of osteoblasts. Journal of Cell Biology, 159(1):147–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
1/147.

Anyanful:2004:RBP

[AOJ^{+}04] Akwasi Anyanful, Kanako Ono, Robert C. Johnsen, Hinh Ly, Victor Jensen, David L. Baillie, and Shoichiro Ono. The RNA-binding protein SUP-12 controls muscle-specific splicing of the ADF/cofilin pre-mRNA in C. elegans. Journal of Cell Biology, 167(4):639–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/4/639.

Ao:2000:CEU

[AP00] Wanyuan Ao and Dave Pilgrim. Caenorhabditis elegans Unc-45 is a component of muscle thick filaments and colocalizes with myosin heavy chain B, but not myosin heavy chain A. Journal of Cell Biology, 148(2):375–??, January 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/2/

Abal:2002:MRC

[APBC^{+}02] Miguel Abal, Matthieu Piel, Veronique Bouckson-Castaing, Mette Mogensen, Jean-Baptiste Sibarita, and Michel Bornens. Microtubule release from the centrosome in migrating cells. Journal of Cell Biology, 159(5):731–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
5/731.

Arregui:2000:NTK

[APLB00] Carlos Arregui, Purnima Pathre, Jack Lilien, and Janne Balsamo. The Nonreceptor Tyrosine Kinase Fer Mediates Cross– Talk between N–Cadherin and β1-Integrins. Journal of Cell Biology, 149(6):1263–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/6/1263.

Arnoult:2002:MRA

[APM^{+}02] Damien Arnoult, Philippe Parone, Jean-Claude Martinou, Bruno Antonsson, Jérôme Estaquier, and Jean Claude Ameisen. Mitochondrial release of apoptosis-inducing factor occurs downstream of cytochrome c release in response to several proapoptotic stimuli. Journal of Cell Biology, 159(6): 923–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/6/923.

Arthur:2004:RPC

[AQC04] William T. Arthur, Lawrence A. Quilliam, and Jonathan A. Cooper. Rap1 promotes cell spreading by localizing Rac guanine nucleotide exchange factors. Journal of Cell Biology, 167(1):111–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/1/111.

Abe:2003:LGB

[AQHO03] Mitsuhiro Abe, Hiroshi Qadota, Aiko Hirata, and Yoshikazu Ohya. Lack of GTP-bound Rho1p in secretory vesicles of Saccharomyces cerevisiae. Journal of Cell Biology, 162(1):85–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
162/1/85.

Aitchison:2000:RR

[AR00] John D. Aitchison and Michael P. Rout. The Road to Ribosomes. Journal of Cell Biology, 151(5):0–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
5/F23.

Allen:2003:STE

[AR03] Benjamin L. Allen and Alan C. Rapraeger. Spatial and temporal expression of heparan sulfate in mouse development regulates FGF and FGF receptor assembly. Journal of Cell Biology, 163(3):637–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/3/637.

Ahmed:2000:PVS

[ARK^{+}00] Sharif U. Ahmed, Enrique Rojo, Valentina Kovaleva, Sridhar Venkataraman, James E. Dombrowski, Ken Matsuoka, and Natasha V. Raikhel. The Plant Vacuolar Sorting Receptor Atelp Is Involved in Transport of Nh_{2} -Terminal Propeptide– Containing Vacuolar Proteins in Arabidopsis thaliana. Journal of Cell Biology, 149(7):1335–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/7/1335.

Abad-Rodriguez:2004:NMC

[ARLC^{+}04] Jose Abad-Rodriguez, Maria Dolores Ledesma, Katleen Craessaerts, Simona Perga, Miguel Medina, Andre Delacourte, Colin Dingwall, Bart De Strooper, and Carlos G. Dotti. Neuronal membrane cholesterol loss enhances amyloid peptide generation. Journal of Cell Biology, 167(5): 953–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/5/953.

Amack:2002:MDU

[ARM02] Jeffrey D. Amack, Shannon R. Reagan, and Mani S. Mahadevan. Mutant DMPK 3’-UTR transcripts disrupt C2C12 myogenic differentiation by compromising MyoD. Journal of Cell Biology, 159(3):419–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/3/419.

Arrizabalaga:2004:IRM

[ARMB04] Gustavo Arrizabalaga, Felix Ruiz, Silvia Moreno, and John C. Boothroyd. Ionophore-resistant mutant of Toxoplasma gondii reveals involvement of a sodium/hydrogen exchanger in calcium regulation. Journal of Cell Biology, 165 (5):653–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/5/653.

Alberghina:2004:CSN

[ARQ^{+}04] Lilia Alberghina, Riccardo L. Rossi, Lorenzo Querin, Valeria Wanke, and Marco Vanoni. A cell sizer network involving Cln3 and Far1 controls entrance into S phase in the mitotic cycle of budding yeast. Journal of Cell Biology, 167(3): 433–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/3/433.

Adams:2000:SFS

[AS00] Josephine Clare Adams and Martin Alexander Schwartz. Stimulation of Fascin Spikes by Thrombospondin-1 Is Mediated by the Gtpases Rac and Cdc42. Journal of Cell Biology, 150(4):807–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/4/807.

Aplin:2001:IMA

[ASAJ01] Andrew E. Aplin, Sheryl A. Stewart, Richard K. Assoian, and R. L. Juliano. Integrin-mediated adhesion regulates ERK nuclear translocation and phosphorylation of Elk-1. Journal of Cell Biology, 153(2):273–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/2/273.

Asakura:2002:MSS

[ASGGR02] Atsushi Asakura, Patrick Seale, Adele Girgis-Gabardo, and Michael A. Rudnicki. Myogenic specification of side population cells in skeletal muscle. Journal of Cell Biology, 159(1): 123–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/1/123.

Arteaga-Solis:2001:RLP

[ASGL^{+}01] Emilio Arteaga-Solis, Barbara Gayraud, Sui Y. Lee, Lillian Shum, Lynn Sakai, and Francesco Ramirez. Regulation of limb patterning by extracellular microfibrils. Journal of Cell Biology, 154(2):275–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/2/275.

Atorino:2003:LMA

[ASK^{+}03] Luigia Atorino, Laura Silvestri, Mirko Koppen, Laura Cassina, Andrea Ballabio, Roberto Marconi, Thomas Langer, and Giorgio Casari. Loss of m–AAA protease in mitochondria causes complex I deficiency and increased sensitivity to oxidative stress in hereditary spastic paraplegia. Journal of Cell Biology, 163(4):777–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/4/777.

Alexander:2001:SRA

[ASMW01] Caroline M. Alexander, Sushma Selvarajan, John Mudgett, and Zena Werb. Stromelysin-1 Regulates Adipogenesis during Mammary Gland Involution. Journal of Cell Biology, 152(4):693–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/4/693.

Agrawal:2000:VTM

[ASP^{+}00] Rajendra K. Agrawal, Christian M. T. Spahn, Pawel Penczek, Robert A. Grassucci, Knud H. Nierhaus, and Joachim Frank. Visualization of Trna Movements on the Escherichia coli 70s Ribosome during the Elongation Cycle. Journal of Cell Biology, 150(3):447–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/3/
447.

Alto:2002:RKA

[ASS02] Neal M. Alto, Jacquelyn Soderling, and John D. Scott. Rab32 is an A-kinase anchoring protein and participates in mitochondrial dynamics. Journal of Cell Biology, 158(4): 659–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/4/659.

Adolfsen:2004:STD

[ASYL04] Bill Adolfsen, Sudipta Saraswati, Motojiro Yoshihara, and J. Troy Littleton. Synaptotagmins are trafficked to distinct subcellular domains including the postsynaptic compartment. Journal of Cell Biology, 166(2):249–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/166/2/
249.

Ahting:2001:TPF

[ATE^{+}01] Uwe Ahting, Michel Thieffry, Harald Engelhardt, Reiner Hegerl, Walter Neupert, and Stephan Nussberger. Tom40, the Pore–Forming Component of the Protein–Conducting Tom Channel in the Outer Membrane of Mitochondria. Journal of Cell Biology, 153(6):1151–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/6/1151.

Ang:2004:REC

[ATF^{+}04] Agnes Lee Ang, Tomohiko Taguchi, Stephen Francis, Heike Fölsch, Lindsay J. Murrells, Marc Pypaert, Graham Warren, and Ira Mellman. Recycling endosomes can serve as intermediates during transport from the Golgi to the plasma membrane of MDCK cells. Journal of Cell Biology, 167(3): 531–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/3/531.

Ackerley:2003:NHC

[ATG^{+}03] Steven Ackerley, Paul Thornhill, Andrew J. Grierson, Janet Brownlees, Brian H. Anderton, P. Nigel Leigh, Christopher E. Shaw, and Christopher C. J. Miller. Neurofilament heavy chain side arm phosphorylation regulates axonal transport of neurofilaments. Journal of Cell Biology, 161 (3):489–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/3/489.

Aoudjit:2001:MAR

[AV01] Fawzi Aoudjit and Kristiina Vuori. Matrix Attachment Regulates FAS–Induced Apoptosis in Endothelial Cells. Journal of Cell Biology, 152(3):633–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/3/633.

Adams:2001:ZNS

[AvdWM^{+}01] David J. Adams, Louise van der Weyden, Akila Mayeda, Stefan Stamm, Brian J. Morris, and John E. J. Rasko. ZNF265 — a novel spliceosomal protein able to induce alternative splicing. Journal of Cell Biology, 154(1):25–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/154/1/
25.

Aoyagi:2000:GED

[AW00] Norikazu Aoyagi and David A. Wassarman. Genes Encoding Drosophila melanogaster RNA Polymerase II General Transcription Factors. Journal of Cell Biology, 150(2):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
2/F45.

Alsop:2003:MOS

[AZ03] G. Bradley Alsop and Dahong Zhang. Microtubules are the only structural constituent of the spindle apparatus required for induction of cell cleavage. Journal of Cell Biology, 162 (3):383–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/3/383.

Aharonovitz:2000:IPR

[AZB^{+}00] Orit Aharonovitz, Hans C. Zaun, Tamas Balla, John D. York, John Orlowski, and Sergio Grinstein. Intracellular Ph regulation by Na^{+} /H^{+} exchange requires phosphatidylinositol 4,5bisphosphate. Journal of Cell Biology, 150(1):213–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
1/213.

Auchincloss:2002:CTN

[AZP^{+}02] Andrea H. Auchincloss, William Zerges, Karl Perron, Jacqueline Girard-Bascou, and Jean-David Rochaix. Characterization of Tbc2, a nucleus-encoded factor specifically required for translation of the chloroplast psbC mRNA in Chlamydomonas reinhardtii. Journal of Cell Biology, 157(6):953–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
157/6/953.

Berghs:2000:SNS

[BAD^{+}00] Stanny Berghs, Diego Aggujaro, Ronald Dirkx, Elena Maksimova, Paul Stabach, Jean-Michel Hermel, Jian-Ping Zhang, William Philbrick, Vladimir Slepnev, Tatiana Ort, and Michele Solimena. β iv Spectrin, a New Spectrin Localized at Axon Initial Segments and Nodes of Ranvier in the Central and Peripheral Nervous System. Journal of Cell Biology, 151(5):985–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/5/985.

Bahadoran:2001:R

[BAM^{+}01] Philippe Bahadoran, Edith Aberdam, Frédéric Mantoux, Roser Buscà, Karine Bille, Nevin Yalman, Geneviève de Saint-Basile, Ricardo Casaroli-Marano, Jean-Paul Ortonne, and Robert Ballotti. Rab27a. Journal of Cell Biology, 152(4):843–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/4/843.

Barr:2004:GI

[Bar04] Francis A. Barr. Golgi inheritance. Journal of Cell Biology, 164(7):955–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/7/955.

Biswas:2003:MNS

[BAZA03] Gopa Biswas, Hindupur K. Anandatheerthavarada, Mone Zaidi, and Narayan G. Avadhani. Mitochondria to nucleus stress signaling. Journal of Cell Biology, 161(3):507–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
3/507.

Berke:2004:SFA

[BBBS04] Ian C. Berke, Thomas Boehmer, Günter Blobel, and Thomas U. Schwartz. Structural and functional analysis of Nup133 domains reveals modular building blocks of the nuclear pore complex. Journal of Cell Biology, 167(4):591– ??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/4/591.

Bakker:2004:FRE

[BBDK^{+}04] Walbert J. Bakker, Montserrat Blázquez-Domingo, Andrea Kolbus, Janey Besooyen, Peter Steinlein, Hartmut Beug, Paul J. Coffer, Bob Löwenberg, Marieke von Lindern, and Thamar B. van Dijk. FoxO3a regulates erythroid differentiation and induces BTG1, an activator of protein arginine methyl transferase 1. Journal of Cell Biology, 164(2): 175–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/2/175.

Braz:2002:PRH

[BBDM02] Julian C. Braz, Orlando F. Bueno, Leon J. De Windt, and Jeffery D. Molkentin. PKCα regulates the hypertrophic growth of cardiomyocytes through extracellular signal-regulated kinase1/2 (ERK1/2). Journal of Cell Biology, 156(5):905–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/5/905.

Bagnato:2003:BAI

[BBG^{+}03] Paola Bagnato, Virigina Barone, Emiliana Giacomello, Daniela Rossi, and Vincenzo Sorrentino. Binding of an ankyrin-1 isoform to obscurin suggests a molecular link between the sarcoplasmic reticulum and myofibrils in striated muscles. Journal of Cell Biology, 160(2):245–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
2/245.

Bache:2003:HRM

[BBMS03] Kristi G. Bache, Andreas Brech, Anja Mehlum, and Harald Stenmark. Hrs regulates multivesicular body formation via ESCRT recruitment to endosomes. Journal of Cell Biology, 162(3):435–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/3/435.

Barbero:2002:VRM

[BBP02] Pierre Barbero, Lenka Bittova, and Suzanne R. Pfeffer. Visualization of Rab9-mediated vesicle transport from endosomes to the trans-Golgi in living cells. Journal of Cell Biology, 156(3):511–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/3/511.

Bolino:2004:DMP

[BBP^{+}04] Alessandra Bolino, Annalisa Bolis, Stefano Carlo Previtali, Giorgia Dina, Simona Bussini, Gabriele Dati, Stefano Amadio, Ubaldo Del Carro, Dolores D. Mruk, Maria Laura Feltri, C. Yan Cheng, Angelo Quattrini, and Lawrence Wrabetz. Disruption of Mtmr2 produces CMT4B1-like neuropathy with myelin outfolding and impaired spermatogenesis. Journal of Cell Biology, 167(4):711–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/4/711.

Beauvais:2004:SER

[BBR04] DeannaLee M. Beauvais, Brandon J. Burbach, and Alan C. Rapraeger. The syndecan-1 ectodomain regulates α_{v}β_{3} integrin activity in human mammary carcinoma cells. Journal of Cell Biology, 167(1):171–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/1/171.

Brickner:2001:TSC

[BBSF01] Jason H. Brickner, Jennifer M. Blanchette, György Sipos, and Robert S. Fuller. The Tlg SNARE complex is required for TGN homotypic fusion. Journal of Cell Biology, 155(6): 969–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/6/969.

Brody:2000:DMG

[BC00] Thomas Brody and Anibal Cravchik. Drosophila melanogaster G Protein–Coupled Receptors. Journal of Cell Biology, 150(2):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/2/F83.

Byrd:2003:VCD

[BC03] Keith Byrd and Victor G. Corces. Visualization of chromatin domains created by the gypsy insulator of Drosophila. Journal of Cell Biology, 162(4):565–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/4/565.

Bompard:2004:RWW

[BC04] Guillaume Bompard and Emmanuelle Caron. Regulation of WASP/WAVE proteins. Journal of Cell Biology, 166(7):957– ??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/7/957.

Belachew:2003:PNP

[BCA^{+}03] Shibeshih Belachew, Ramesh Chittajallu, Adan A. Aguirre, Xiaoqing Yuan, Martha Kirby, Stacie Anderson, and Vittorio Gallo. Postnatal NG2 proteoglycan–expressing progenitor cells are intrinsically multipotent and generate functional neurons. Journal of Cell Biology, 161(1):169–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
1/169.

Boisvert:2002:SDM

[BCB^{+}02] François-Michel Boisvert, Jocelyn Côté, Marie-Chloé Boulanger, Patrick Cléroux, François Bachand, Chantal Autexier, and Stéphane Richard. Symmetrical dimethylarginine methylation is required for the localization of SMN in Cajal bodies and pre-mRNA splicing. Journal of Cell Biology, 159(6): 957–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/6/957.

Bednenko:2003:ICC

[BCG03] Janna Bednenko, Gino Cingolani, and Larry Gerace. Importin β contains a COOH-terminal nucleoporin binding region important for nuclear transport. Journal of Cell Biology, 162(3):391–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/3/391.

Brieher:2004:FMP

[BCM04] William M. Brieher, Margaret Coughlin, and Timothy J. Mitchison. Fascin-mediated propulsion of Listeria monocytogenes independent of frequent nucleation by the Arp2/3 complex. Journal of Cell Biology, 165(2):233–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/165/2/
233.

Borgese:2003:TTA

[BCP03] Nica Borgese, Sara Colombo, and Emanuela Pedrazzini. The tale of tail-anchored proteins. Journal of Cell Biology, 161 (6):1013–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/6/1013.

Babiychuk:2000:ACM

[BD00] Eduard B. Babiychuk and Annette Draeger. Annexins in Cell Membrane Dynamics. Journal of Cell Biology, 150(5):1113– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/5/1113.

Bazzoni:2002:KJE

[BD02] Gianfranco Bazzoni and Elisabetta Dejana. Keratinocyte junctions and the epidermal barrier. Journal of Cell Biology, 156(6):947–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/6/947.

Beningo:2001:NFA

[BDK^{+}01] Karen A. Beningo, Micah Dembo, Irina Kaverina, J. Victor Small, and Yu li Wang. Nascent Focal Adhesions Are Responsible for the Generation of Strong Propulsive Forces in Migrating Fibroblasts. Journal of Cell Biology, 153(4):881–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/4/881.

Berdeaux:2004:ARL

[BDKM04] Rebecca L. Berdeaux, Begoña Díaz, Lomi Kim, and G. Steven Martin. Active Rho is localized to podosomes induced by oncogenic Src and is required for their assembly and function. Journal of Cell Biology, 166(3):317–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
3/317.

Blokzijl:2003:CTB

[BDR^{+}03] Andries Blokzijl, Camilla Dahlqvist, Eva Reissmann, Anna Falk, Annalena Moliner, Urban Lendahl, and Carlos F. Ibáñez. Cross-talk between the Notch and TGF-β signaling pathways mediated by interaction of the Notch intracellular domain with Smad3. Journal of Cell Biology, 163(4): 723–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/4/723.

Ben-Efraim:2001:GIA

[BEG01] Iris Ben-Efraim and Larry Gerace. Gradient of Increasing Affinity of Importin β for Nucleoporins along the Pathway of Nuclear Import. Journal of Cell Biology, 152(2):411–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
152/2/411.

Bloom:2003:CSA

[BET^{+}03] Ona Bloom, Emma Evergren, Nikolay Tomilin, Ole Kjaerulff, Peter Löw, Lennart Brodin, Vincent A. Pieribone, Paul Greengard, and Oleg Shupliakov. Colocalization of synapsin and actin during synaptic vesicle recycling. Journal of Cell Biology, 161(4):737–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/4/737.

Borges:2001:AIP

[BF01] Lúcia S. Borges and Michael Ferns. Agrin-induced Phosphorylation of the Acetylcholine Receptor Regulates Cytoskeletal Anchoring and Clustering. Journal of Cell Biology, 153 (1):1–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/1/1.

Baldanzi:2002:GAG

[BFC^{+}02] Gianluca Baldanzi, Nicoletta Filigheddu, Santina Cutrupi, Filomena Catapano, Sara Bonissoni, Alberto Fubini, Daniela Malan, Germano Baj, Riccarda Granata, Fabio Broglio, Mauro Papotti, Nicola Surico, Federico Bussolino, Jorgen Isgaard, Romano Deghenghi, Fabiola Sinigaglia, Maria Prat, Giampiero Muccioli, Ezio Ghigo, and Andrea Graziani. Ghrelin and des-acyl ghrelin inhibit cell death in cardiomyocytes and endothelial cells through ERK1/2 and PI 3-kinase/AKT. Journal of Cell Biology, 159(6):1029–??, December 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/6/
1029.

Bix:2004:ECE

[BFG^{+}04] Gregory Bix, Jian Fu, Eva M. Gonzalez, Laura Macro, Amy Barker, Shelly Campbell, Mary M. Zutter, Samuel A. Santoro, Jiyeun K. Kim, Magnus Höök, Charles C. Reed, and Renato V. Iozzo. Endorepellin causes endothelial cell disassembly of actin cytoskeleton and focal adhesions through α2β1 integrin. Journal of Cell Biology, 166(1):97–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
1/97.

Bloch:2001:DCI

[BFH^{+}01] Wilhelm Bloch, Yun Fan, Ji Han, Sheng Xue, Torsten Schöneberg, Guanju Ji, Zhong J. Lu, Michael Walther, Reinhard Fässler, Jürgen Hescheler, Klaus Addicks, and Bernd K. Fleischmann. Disruption of cytoskeletal integrity impairs G_{i}- mediated signaling due to displacement of G_{i} proteins. Journal of Cell Biology, 154(4):753–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/4/753.

Brun:2004:DLT

[BFSO^{+}04] Thierry Brun, Isobel Franklin, Luc St-Onge, Anna Biason-Lauber, Eugene J. Schoenle, Claes B. Wollheim, and Benoit R. Gauthier. The diabetes-linked transcription factor PAX4 promotes β-cell proliferation and survival in rat and human islets. Journal of Cell Biology, 167(6):1123– ??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/6/1123.

Benaud:2004:AIA

[BGA^{+}04] Christelle Benaud, Benoît J. Gentil, Nicole Assard, Magalie Court, Jerome Garin, Christian Delphin, and Jacques Baudier. AHNAK interaction with the annexin 2/S100A10 complex regulates cell membrane cytoarchitecture. Journal of Cell Biology, 164(1):133–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/1/133.

Bucciarelli:2003:SAC

[BGBG03] Elisabetta Bucciarelli, Maria Grazia Giansanti, Silvia Bonaccorsi, and Maurizio Gatti. Spindle assembly and cytokinesis in the absence of chromosomes during Drosophila male meiosis. Journal of Cell Biology, 160(7):993–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/7/
993.

Blow:2001:ROX

[BGFJ01] J. Julian Blow, Peter J. Gillespie, Dennis Francis, and Dean A. Jackson. Replication Origins in Xenopus Egg Extract Are 5–15 Kilobases Apart and Are Activated in Clusters That Fire at Different Times. Journal of Cell Biology, 152 (1):15–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/1/15.

Bonne:2003:DDP

[BGH^{+}03] Stefan Bonné, Barbara Gilbert, Mechthild Hatzfeld, Xinyu Chen, Kathleen J. Green, and Frans van Roy. Defining desmosomal plakophilin-3 interactions. Journal of Cell Biology, 161(2):403–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/403.

Balda:2003:ZAB

[BGM03] Maria S. Balda, Michelle D. Garrett, and Karl Matter. The ZO-1–associated Y-box factor ZONAB regulates epithelial cell proliferation and cell density. Journal of Cell Biology, 160(3):423–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/3/423.

Bierne:2001:RCL

[BGR^{+}01] Hélène Bierne, Edith Gouin, Pascal Roux, Pico Caroni, Helen L. Yin, and Pascale Cossart. A role for cofilin and LIM kinase in Listeria-induced phagocytosis. Journal of Cell Biology, 155(1):101–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/1/101.

Brennan:2000:PLH

[BGS00] Christopher M. Brennan, Imed-Eddine Gallouzi, and Joan A. Steitz. Protein Ligands to Hur Modulate Its Interaction with Target Mrnas in Vivo. Journal of Cell Biology, 151 (1):1–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/1/1.

Bougneres:2004:CCC

[BGW^{+}04] Laurence Bougnères, Stéphane E. Girardin, Scott A. Weed, Andrei V. Karginov, Jean-Christophe Olivo-Marin, J. Thomas Parsons, Philippe J. Sansonetti, and Guy Tran Van Nhieu. Cortactin and Crk cooperate to trigger actin polymerization during Shigella invasion of epithelial cells. Journal of Cell Biology, 166(2):225–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/2/225.

Boisvert:2000:PLP

[BHBJ00] François-Michel Boisvert, Michael J. Hendzel, and David P. Bazett-Jones. Promyelocytic Leukemia (Pml) Nuclear Bodies Are Protein Structures That Do Not Accumulate RNA. Journal of Cell Biology, 148(2):283–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/2/283.

Bezakova:2001:EPR

[BHFL01] Gabriela Bezakova, Johannes P. Helm, Maura Francolini, and Terje Lømo. Effects of Purified Recombinant Neural and Muscle Agrin on Skeletal Muscle Fibers in Vivo. Journal of Cell Biology, 153(7):1441–??, June 2001. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/7/1441.

Ballestrem:2001:MFD

[BHIWH01] Christoph Ballestrem, Boris Hinz, Beat A. Imhof, and Bernhard Wehrle-Haller. Marching at the front and dragging behind. Journal of Cell Biology, 155(7):1319–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
7/1319.

Bunnell:2002:CRL

[BHK^{+}02] Stephen C. Bunnell, David I. Hong, Julia R. Kardon, Tetsuo Yamazaki, C. Jane McGlade, Valarie A. Barr, and Lawrence E. Samelson. T cell receptor ligation induces the formation of dynamically regulated signaling assemblies. Journal of Cell Biology, 158(7):1263–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/7/
1263.

Benharouga:2001:CTT

[BHKL01] Mohamed Benharouga, Martin Haardt, Norbert Kartner, and Gergely L. Lukacs. Cooh-terminal Truncations Promote Proteasome–Dependent Degradation of Mature Cystic Fibrosis Transmembrane Conductance Regulator from post-Golgi Compartments. Journal of Cell Biology, 153(5):957–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
5/957.

Black:2001:NNT

[BHL^{+}01] Ben E. Black, James M. Holaska, Lyne Lévesque, Batool Ossareh-Nazari, Carol Gwizdek, Catherine Dargemont, and Bryce M. Paschal. Nxt1 is necessary for the terminal step of Crm1-mediated nuclear export. Journal of Cell Biology, 152(1):141–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/1/141.

Bosher:2003:CEV

[BHL^{+}03] Julia M. Bosher, Bum-Soo Hahn, Renaud Legouis, Satis Sookhareea, Robby M. Weimer, Anne Gansmuller, Andrew D. Chisholm, Ann M. Rose, Jean-Louis Bessereau, and Michel Labouesse. The Caenorhabditis elegans vab-10 spectraplakin isoforms protect the epidermis against internal and external forces. Journal of Cell Biology, 161(4):757–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
4/757.

Browning:2000:TKL

[BHM^{+}00] Heidi Browning, Jacqueline Hayles, Juan Mata, Lauren Aveline, Paul Nurse, and J. Richard McIntosh. Tea2p Is a Kinesin-like Protein Required to Generate Polarized Growth in Fission Yeast. Journal of Cell Biology, 151(1):15–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
151/1/15.

Bashaw:2001:NDF

[BHNG01] Greg J. Bashaw, Hailan Hu, Catherine D. Nobes, and Corey S. Goodman. A novel Dbl family RhoGEF promotes Rho-dependent axon attraction to the central nervous system midline in Drosophila and overcomes Robo repulsion. Journal of Cell Biology, 155(7):1117–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/7/1117.

Boyd:2004:VCM

[BHPN04] Charles Boyd, Thom Hughes, Marc Pypaert, and Peter Novick. Vesicles carry most exocyst subunits to exocytic sites marked by the remaining two subunits, Sec3p and Exo70p. Journal of Cell Biology, 167(5):889–??, December 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/167/5/
889.

Bauer:2002:ERG

[BHW^{+}02a] Jörg Bauer, Andreas Hiltbrunner, Petra Weibel, Pierre-Alexandre Vidi, Mayte Alvarez-Huerta, Matthew D. Smith, Danny J. Schnell, and Felix Kessler. Essential role of the Gdomain in targeting of the protein import receptor atToc159 to the chloroplast outer membrane. Journal of Cell Biology, 159(5):845–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/5/845.

Becker:2002:CER

[BHW02b] Thalia Becker, F.-Ulrich Hartl, and Felix Wieland. CD40, an extracellular receptor for binding and uptake of Hsp70– peptide complexes. Journal of Cell Biology, 158(7):1277– ??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/7/1277.

Bishop:2002:MCV

[BHW02c] Naomi Bishop, Alistair Horman, and Philip Woodman. Mammalian class E vps proteins recognize ubiquitin and act in the removal of endosomal protein–ubiquitin conjugates. Journal of Cell Biology, 157(1):91–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/1/91.

Bisogno:2003:CFS

[BHW^{+}03] Tiziana Bisogno, Fiona Howell, Gareth Williams, Alberto Minassi, Maria Grazia Cascio, Alessia Ligresti, Isabel Matias, Aniello Schiano-Moriello, Praveen Paul, Emma-Jane Williams, Uma Gangadharan, Carl Hobbs, Vincenzo Di Marzo, and Patrick Doherty. Cloning of the first sn1-DAG lipases points to the spatial and temporal regulation of endocannabinoid signaling in the brain. Journal of Cell Biology, 163(3):463–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/3/463.

Beauchamp:2000:ECM

[BHY^{+}00] Jonathan R. Beauchamp, Louise Heslop, David S. W. Yu, Shahragim Tajbakhsh, Robert G. Kelly, Anton Wernig, Margaret E. Buckingham, Terence A. Partridge, and Peter S. Zammit. Expression of Cd34 and Myf5 Defines the Majority of Quiescent Adult Skeletal Muscle Satellite Cells. Journal of Cell Biology, 151(6):1221–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/6/1221.

Barkalow:2003:ADF

[BIC^{+}03] Kurt L. Barkalow, Joseph E. Italiano, Denise E. Chou, Yoichiro Matsuoka, Vann Bennett, and John H. Hartwig. αadducin dissociates from F-actin and spectrin during platelet activation. Journal of Cell Biology, 161(3):557–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/3/
557.

Birnbaum:2003:L

[Bir03] Morris J. Birnbaum. Lipolysis. Journal of Cell Biology, 161 (6):1011–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/6/1011.

Birmingham:2004:BEF

[Bir04a] Karen Birmingham. Bigger equals faster. Journal of Cell Biology, 164(6):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/6/792.2.

Birmingham:2004:BB

[Bir04b] Karen Birmingham. Binding to the basement. Journal of Cell Biology, 164(6):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/6/793.3.

Birmingham:2004:CS

[Bir04c] Karen Birmingham. Cilia that sense. Journal of Cell Biology, 164(6):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/6/792.1.

Birmingham:2004:KMY

[Bir04d] Karen Birmingham. Keratins make you regular. Journal of Cell Biology, 164(6):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/6/793.1.

Birmingham:2004:SDS

[Bir04e] Karen Birmingham. Stay dynamic to slow aging. Journal of Cell Biology, 164(6):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/6/793.2.

Bryant:2003:SMS

[BJ03] Nia J. Bryant and David E. James. The Sec1p/Munc18 (SM) protein, Vps45p, cycles on and off membranes during vesicle transport. Journal of Cell Biology, 161(4):691–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
4/691.

Babu:2003:REM

[BJB^{+}03] J. Ramesh Babu, Karthik B. Jeganathan, Darren J. Baker, Xiaosheng Wu, Ningling Kang-Decker, and Jan M. van Deursen. Rae1 is an essential mitotic checkpoint regulator that cooperates with Bub3 to prevent chromosome missegregation. Journal of Cell Biology, 160(3):341–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
3/341.

Beqaj:2002:HRA

[BJM^{+}02] Safedin Beqaj, Sandhya Jakkaraju, Raymond R. Mattingly, Desi Pan, and Lucia Schuger. High RhoA activity maintains the undifferentiated mesenchymal cell phenotype, whereas RhoA down-regulation by laminin-2 induces smooth muscle myogenesis. Journal of Cell Biology, 156(5):893–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/156/
5/893.

Boisvert:2001:TCC

[BKB^{+}01] François-Michel Boisvert, Michael J. Kruhlak, Alan K. Box, Michael J. Hendzel, and David P. Bazett-Jones. The Transcription Coactivator Cbp Is a Dynamic Component of the Promyelocytic Leukemia Nuclear Body. Journal of Cell Biology, 152(5):1099–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/5/1099.

Bialkowska:2000:EII

[BKD^{+}00] Katarzyna Bialkowska, Sucheta Kulkarni, Xiaoping Du, Darrel E. Goll, Takaomi C. Saido, and Joan E. B. Fox. Evidence That β3 Integrin–Induced Rac Activation Involves the Calpain–Dependent Formation of Integrin Clusters That Are Distinct from the Focal Complexes and Focal Adhesions That Form as Rac and Rhoa Become Active. Journal of Cell Biology, 151(3):685–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/3/685.

Broday:2004:LDP

[BKD^{+}04] Limor Broday, Irina Kolotuev, Christine Didier, Anindita Bhoumik, Benjamin Podbilewicz, and Ze’ev Ronai. The LIM domain protein UNC-95 is required for the assembly of muscle attachment structures and is regulated by the RING finger protein RNF-5 in C. elegans. Journal of Cell Biology, 165 (6):857–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/6/857.

Budde:2001:ROD

[BKDH01] Priya Prakash Budde, Akiko Kumagai, William G. Dunphy, and Rebecca Heald. Regulation of Op18 during Spindle Assembly in Xenopus Egg Extracts. Journal of Cell Biology, 153 (1):149–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/1/149.

Bischof:2001:RLB

[BKI^{+}01] Oliver Bischof, Sahn-Ho Kim, John Irving, Sergey Beresten, Nathan A. Ellis, and Judith Campisi. Regulation and Localization of the Bloom Syndrome Protein in Response to DNA Damage. Journal of Cell Biology, 153(2):367–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
2/367.

Brunet:2002:TNP

[BKS^{+}02] Anne Brunet, Fumihiko Kanai, Justine Stehn, Jian Xu, Dilara Sarbassova, John V. Frangioni, Sorab N. Dalal, James A. DeCaprio, Michael E. Greenberg, and Michael B. Yaffe. 14-3-3 transits to the nucleus and participates in dynamic nucleocytoplasmic transport. Journal of Cell Biology, 156(5): 817–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/5/817.

Baldock:2001:SOF

[BKZ^{+}01] Clair Baldock, Abraham J. Koster, Ulrike Ziese, Matthew J. Rock, Michael J. Sherratt, Karl E. Kadler, C. Adrian Shuttleworth, and Cay M. Kielty. The Supramolecular Organization of Fibrillin–Rich Microfibrils. Journal of Cell Biology, 152(5): 1045–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/5/1045.

Bruzzone:2003:STP

[BKZ^{+}03] Santina Bruzzone, Svenja Kunerth, Elena Zocchi, Antonio De Flora, and Andreas H. Guse. Spatio-temporal propagation of Ca^{2+} signals by cyclic ADP-ribose in 3T3 cells stimulated via purinergic P2Y receptors. Journal of Cell Biology, 163(4):837–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/4/837.

Bezakova:2001:MAM

[BL01] Gabriela Bezakova and Terje Lømo. Muscle Activity and Muscle Agrin Regulate the Organization of Cytoskeletal Proteins and Attached Acetylcholine Receptor (Achr) Aggregates in Skeletal Muscle Fibers. Journal of Cell Biology, 153 (7):1453–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/7/1453.

Bhamidipati:2000:ARF

[BLC00] Arunashree Bhamidipati, Sally A. Lewis, and Nicholas J. Cowan. Adp Ribosylation Factor-like Protein 2 (Arl2) Regulates the Interaction of Tubulin–Folding Cofactor D with Native Tubulin. Journal of Cell Biology, 149(5):1087–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
5/1087.

Bonapace:2002:NRE

[BLP^{+}02] Ian Marc Bonapace, Lucia Latella, Roberto Papait, Francesco Nicassio, Alessandra Sacco, Masahiro Muto, Marco Crescenzi, and Pier Paolo Di Fiore. Np95 is regulated by E1A during mitotic reactivation of terminally differentiated cells and is essential for S phase entry. Journal of Cell Biology, 157 (6):909–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/6/909.

Bui:2001:ANF

[BLPP01] Nguyen Truc Bui, Antonia Livolsi, Jean-François Peyron, and Jochen H. M. Prehn. Activation of Nuclear Factor κ b and bcl-x Survival Gene Expression by Nerve Growth Factor Requires Tyrosine Phosphorylation of IκBα. Journal of Cell Biology, 152(4):753–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/4/753.

Bentele:2004:MMR

[BLU^{+}04] M. Bentele, I. Lavrik, M. Ulrich, S. Stößer, D. W. Heermann, H. Kalthoff, P. H. Krammer, and R. Eils. Mathematical modeling reveals threshold mechanism in CD95-induced apoptosis. Journal of Cell Biology, 166(6):839–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
6/839.

Brownawell:2002:ENK

[BM02] Amy M. Brownawell and Ian G. Macara. Exportin-5, a novel karyopherin, mediates nuclear export of double-stranded RNA binding proteins. Journal of Cell Biology, 156(1): 53–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/1/53.

Brown:2000:HSP

[BMC00] C. Randell Brown, Jameson A. McCann, and Hui-Ling Chiang. The Heat Shock Protein Ssa2p Is Required for Import of Fructose-1,6-Bisphosphatase into Vid Vesicles. Journal of Cell Biology, 150(1):65–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/1/65.

Berrier:2000:ARR

[BMD^{+}00] Allison L. Berrier, Anthony M. Mastrangelo, Julian Downward, Mark Ginsberg, and Susan E. LaFlamme. Activated R–Ras, Rac1, Pi 3-Kinase and Pkcϵ Can Each Restore Cell Spreading Inhibited by Isolated Integrin β1 Cytoplasmic Domains. Journal of Cell Biology, 151(7):1549–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
7/1549.

Bowe:2000:SLR

[BMF00] Mark A. Bowe, Duane B. Mendis, and Justin R. Fallon. The Small Leucine–Rich Repeat Proteoglycan Biglycan Binds to α-Dystroglycan and Is Upregulated in Dystrophic Muscle. Journal of Cell Biology, 148(4):801–??, February 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/4/
801.

Bajou:2001:PAI

[BMG^{+}01] Khalid Bajou, Véronique Masson, Robert D. Gerard, Petra M. Schmitt, Valérie Albert, Michael Praus, Leif R. Lund, Thomas L. Frandsen, Nils Brunner, Keld Dano, Norbert E. Fusenig, Ulrich Weidle, Geert Carmeliet, David Loskutoff, Desiré Collen, Peter Carmeliet, Jean Michel Foidart, and Agnès Noël. The Plasminogen Activator Inhibitor PAI-1 Controls in Vivo Tumor Vascularization by Interaction with Proteases, Not Vitronectin. Journal of Cell Biology, 152(4): 777–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/4/777.

Bonnelye:2001:ONE

[BMKA01] E. Bonnelye, L. Merdad, V. Kung, and J. E. Aubin. The Orphan Nuclear Estrogen Receptor–Related Receptor α (Errα) Is Expressed Throughout Osteoblast Differentiation and Regulates Bone Formation in Vitro. Journal of Cell Biology, 153 (5):971–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/5/971.

Boeddinghaus:2002:CVR

[BMLU02] Christine Boeddinghaus, Alexey J. Merz, Rico Laage, and Christian Ungermann. A cycle of Vam7p release from and PtdIns 3-P-dependent rebinding to the yeast vacuole is required for homotypic vacuole fusion. Journal of Cell Biology, 157(1):79–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/1/79.

Bang:2001:MNK

[BMM^{+}01] Marie-Louise Bang, Ryan E. Mudry, Abigail S. McElhinny, Karoly Trombitás, Adam J. Geach, Rob Yamasaki, Hiroyuki Sorimachi, Henk Granzier, Carol C. Gregorio, and Siegfried Labeit. Myopalladin, a Novel 145-Kilodalton Sarcomeric Protein with Multiple Roles in Z–Disc and I–Band Protein Assemblies. Journal of Cell Biology, 153(2):413–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
2/413.

Barton:2002:MSE

[BMM^{+}02] Elisabeth R. Barton, Linda Morris, Antonio Musaro, Nadia Rosenthal, and H. Lee Sweeney. Muscle-specific expression of insulin-like growth factor I counters muscle decline in mdx mice. Journal of Cell Biology, 157(1):137–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/157/
1/137.

Bananis:2000:MMD

[BMS^{+}00a] Eustratios Bananis, John W. Murray, Richard J. Stockert, Peter Satir, and Allan W. Wolkoff. Microtubule and Motor–Dependent Endocytic Vesicle Sorting in Vitro. Journal of Cell Biology, 151(1):179–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/1/179.

Becker:2000:CLC

[BMS^{+}00b] Matthias Becker, Elke Martin, Jean Schneikert, Harald F. Krug, and Andrew C. B. Cato. Cytoplasmic Localization and the Choice of Ligand Determine Aggregate Formation by Androgen Receptor with Amplified Polyglutamine Stretch. Journal of Cell Biology, 149(2):255–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/2/255.

Bieberich:2003:RCD

[BMS^{+}03] Erhard Bieberich, Sarah MacKinnon, Jeane Silva, Scott Noggle, and Brian G. Condie. Regulation of cell death in mitotic neural progenitor cells by asymmetric distribution of prostate apoptosis response 4 (PAR-4) and simultaneous elevation of endogenous ceramide. Journal of Cell Biology, 162 (3):469–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/3/469.

Bousquet:2001:NTD

[BMY^{+}01] Olivier Bousquet, Linglei Ma, Soichiro Yamada, Changhong Gu, Toshihiro Idei, Kenzo Takahashi, Denis Wirtz, and Pierre A. Coulombe. The nonhelical tail domain of keratin 14 promotes filament bundling and enhances the mechanical properties of keratin intermediate filaments in vitro. Journal of Cell Biology, 155(5):747–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/5/747.

Behrens:2002:RFY

[BN02] Ralf Behrens and Paul Nurse. Roles of fission yeast tea1p in the localization of polarity factors and in organizing the microtubular cytoskeleton. Journal of Cell Biology, 157(5): 783–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/5/783.

Bonangelino:2002:OSI

[BND^{+}02] Cecilia J. Bonangelino, Johnathan J. Nau, Jason E. Duex, Mikala Brinkman, Andrew E. Wurmser, Jonathan D. Gary, Scott D. Emr, and Lois S. Weisman. Osmotic stress– induced increase of phosphatidylinositol 3,5-bisphosphate requires Vac14p, an activator of the lipid kinase Fab1p. Journal of Cell Biology, 156(6):1015–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/6/1015.

Burakov:2003:CPI

[BNSR03] Anton Burakov, Elena Nadezhdina, Boris Slepchenko, and Vladimir Rodionov. Centrosome positioning in interphase cells. Journal of Cell Biology, 162(6):963–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
6/963.

Black:2000:STR

[BP00] Michael W. Black and Hugh R. B. Pelham. A Selective Transport Route from Golgi to Late Endosomes That Requires the Yeast Gga Proteins. Journal of Cell Biology, 151(3): 587–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/3/587.

Berlin:2003:MSS

[BPC03] Ana Berlin, Anne Paoletti, and Fred Chang. Mid2p stabilizes septin rings during cytokinesis in fission yeast. Journal of Cell Biology, 160(7):1083–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/7/1083.

Bubulya:2004:HSS

[BPD^{+}04] Paula A. Bubulya, Kannanganattu V. Prasanth, Thomas J. Deerinck, Daniel Gerlich, Joel Beaudouin, Mark H. Ellisman, Jan Ellenberg, and David L. Spector. Hypophosphorylated SR splicing factors transiently localize around active nucleolar organizing regions in telophase daughter nuclei. Journal of Cell Biology, 167(1):51–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/1/51.

Barr:2001:GMP

[BPKK01] Francis A. Barr, Christian Preisinger, Robert Kopajtich, and Roman Körner. Golgi matrix proteins interact with p24 cargo receptors and aid their efficient retention in the Golgi apparatus. Journal of Cell Biology, 155(6):885–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
6/885.

Bossenmeyer-Pourie:2002:TFP

[BPKR^{+}02] Carine Bossenmeyer-Pourié, Rama Kannan, Stéphane Ribieras, Corinne Wendling, Isabelle Stoll, Lars Thim, Catherine Tomasetto, and Marie-Christine Rio. The trefoil factor 1 participates in gastrointestinal cell differentiation by delaying G1-S phase transition and reducing apoptosis. Journal of Cell Biology, 157(5):761–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/5/761.

Berg:2000:NAA

[BPMG00] Ingeborg Berg, Barry V. L. Potter, Georg W. Mayr, and Andreas H. Guse. Nicotinic acid adenine dinucleotide phosphate (Naadp^{+}) is an essential regulator of T-lymphocyte Ca^{2+}- signaling. Journal of Cell Biology, 150(3):581–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
3/581.

Benais-Pont:2003:ITJ

[BPPFM^{+}03] Gaëlle Benais-Pont, Anu Punn, Catalina Flores-Maldonado, Judith Eckert, Graça Raposo, Tom P. Fleming, Marcelino Cereijido, Maria S. Balda, and Karl Matter. Identification of a tight junction–associated guanine nucleotide exchange factor that activates Rho and regulates paracellular permeability. Journal of Cell Biology, 160(5):729–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/5/
729.

Bajno:2000:FEV

[BPS^{+}00] Lydia Bajno, Xiao-Rong Peng, Alan D. Schreiber, Hsiao-Ping Moore, William S. Trimble, and Sergio Grinstein. Focal Exocytosis of Vamp3-Containing Vesicles at Sites of Phagosome Formation. Journal of Cell Biology, 149(3):697–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
3/697.

Belgareh:2001:ECN

[BRB^{+}01] Naïma Belgareh, Gwénaël Rabut, Siau Wei Baï, Megan van Overbeek, Joël Beaudouin, Nathalie Daigle, Olga V. Zatsepina, Fabien Pasteau, Valérie Labas, Micheline Fromont-Racine, Jan Ellenberg, and Valérie Doye. An evolutionarily conserved NPC subcomplex, which redistributes in part to kinetochores in mammalian cells. Journal of Cell Biology, 154(6):1147–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/6/1147.

Bayer:2003:VMF

[BRB^{+}03] Martin J. Bayer, Christoph Reese, Susanne Bühler, Christopher Peters, and Andreas Mayer. Vacuole membrane fusion.

Journal of Cell Biology, 162(2):211–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/2/211.

Brennwald:2000:RF

[Bre00] Patrick Brennwald. Reversal of Fortune. Journal of Cell Biology, 149(1):??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/1/1.2.

Bretscher:2003:PGO

[Bre03] Anthony Bretscher. Polarized growth and organelle segregation in yeast. Journal of Cell Biology, 160(6):811–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
6/811.

Barbieri:2000:EGF

[BRG^{+}00] M. Alejandro Barbieri, Richard L. Roberts, Aysel Gumusboga, Hilary Highfield, Carmen Alvarez-Dominguez, Alan Wells, and Philip D. Stahl. Epidermal Growth Factor and Membrane Trafficking. Journal of Cell Biology, 151(3):539– ??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/3/539.

Barret:2000:MPB

[BRM^{+}00] Cécile Barret, Christian Roy, Philippe Montcourrier, Paul Mangeat, and Verena Niggli. Mutagenesis of the phosphatidylinositol 4,5-bisphosphate (Pip_{2}) binding site in the Nh_{2}-terminal domain of ezrin correlates with its altered cellular distribution. Journal of Cell Biology, 151(5):1067– ??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/5/1067.

Brown:2003:ATM

[Bro03] Anthony Brown. Axonal transport of membranous and nonmembranous cargoes. Journal of Cell Biology, 160(6): 817–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/6/817.

Bourdon:2002:SNP

[BRR^{+}02] J.-C. Bourdon, J. Renzing, P. L. Robertson, K. N. Fernandes, and D. P. Lane. Scotin, a novel p53-inducible proapoptotic protein located in the ER and the nuclear membrane. Journal of Cell Biology, 158(2):235–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/2/235.

Bony:2001:SRP

[BRS^{+}01] Claire Bony, Serge Roche, Ueno Shuichi, Takehiko Sasaki, Michael A. Crackower, Josef Penninger, Hiroyuki Mano, and Michel Pucéat. A Specific Role of Phosphatidylinositol 3– Kinase γ. Journal of Cell Biology, 152(4):717–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
4/717.

Brown:2001:PBA

[BRY^{+}01] Fraser D. Brown, Andrew L. Rozelle, Helen L. Yin, Tamás Balla, and Julie G. Donaldson. Phosphatidylinositol 4,5bisphosphate and Arf6-regulated membrane traffic. Journal of Cell Biology, 154(5):1007–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/5/1007.

Bushell:2002:HTA

[BS02] Martin Bushell and Peter Sarnow. Hijacking the translation apparatus by RNA viruses. Journal of Cell Biology, 158(3): 395–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/3/395.

Bao:2004:IVM

[BS04a] Wenjie Bao and Staffan Strömblad. Integrin αv-mediated inactivation of p53 controls a MEK1-dependent melanoma cell survival pathway in three-dimensional collagen. Journal of Cell Biology, 167(4):745–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/4/745.

Bidlingmaier:2004:RPG

[BS04b] Scott Bidlingmaier and Michael Snyder. Regulation of polarized growth initiation and termination cycles by the polarisome and Cdc42 regulators. Journal of Cell Biology, 164(2): 207–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/2/207.

Bantel:2001:TMS

[BSD^{+}01] Heike Bantel, Bhanu Sinha, Wolfram Domschke, Georg Peters, Klaus Schulze-Osthoff, and Reiner U. Jänicke. α-toxin is a mediator of Staphylococcus aureus –induced cell death and activates caspases via the intrinsic death pathway independently of death receptor signaling. Journal of Cell Biology, 155(4):637–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/4/637.

Bucior:2004:CCI

[BSEB04] Iwona Bucior, Simon Scheuring, Andreas Engel, and Max M. Burger. Carbohydrate–carbohydrate interaction provides adhesion force and specificity for cellular recognition. Journal of Cell Biology, 165(4):529–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/4/529.

Brandenberger:2001:ICN

[BSL^{+}01] Ralph Brandenberger, Andrea Schmidt, James Linton, Denan Wang, Carey Backus, Sumiko Denda, Ullrich Müller, and Louis F. Reichardt. Identification and characterization of a novel extracellular matrix protein nephronectin that is associated with integrin α8β1 in the embryonic kidney. Journal of Cell Biology, 154(2):447–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/2/447.

Breckenridge:2003:CCP

[BSMS03] David G. Breckenridge, Marina Stojanovic, Richard C. Marcellus, and Gordon C. Shore. Caspase cleavage product of BAP31 induces mitochondrial fission through endoplasmic reticulum calcium signals, enhancing cytochrome c release to the cytosol. Journal of Cell Biology, 160(7):1115–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
7/1115.

Brock:2003:RGM

[BSR^{+}03] Carsten Brock, Michael Schaefer, H. Peter Reusch, Cornelia Czupalla, Manuela Michalke, Karsten Spicher, Günter Schultz, and Bernd Nürnberg. Roles of Gβγ in membrane recruitment and activation of p110γ /p101 phosphoinositide 3-kinase γ. Journal of Cell Biology, 160(1):89–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
1/89.

Burgess:2000:AID

[BSS00] Robert W. Burgess, William C. Skarnes, and Joshua R. Sanes. Agrin Isoforms with Distinct Amino Termini. Journal of Cell Biology, 151(1):41–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/1/41.

Brugger:2000:ESS

[BSW^{+}00] Britta Brügger, Roger Sandhoff, Sabine Wegehingel, Karin Gorgas, Jörg Malsam, J. Bernd Helms, Wolf-Dieter Lehmann, Walter Nickel, and Felix T. Wieland. Evidence for Segregation of Sphingomyelin and Cholesterol during Formation of Copi–Coated Vesicles. Journal of Cell Biology, 151(3): 507–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/3/507.

Bieberich:2004:SAP

[BSW^{+}04] Erhard Bieberich, Jeane Silva, Guanghu Wang, Kannan Krishnamurthy, and Brian G. Condie. Selective apoptosis of pluripotent mouse and human stem cells by novel ceramide analogues prevents teratoma formation and enriches for neural precursors in ES cell–derived neural transplants. Journal of Cell Biology, 167(4):723–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/4/723.

Bloecher:2000:DLP

[BT00] Andrew Bloecher and Kelly Tatchell. Dynamic Localization of Protein Phosphatase Type 1 in the Mitotic Cell Cycle of Saccharomyces cerevisiae. Journal of Cell Biology, 149 (1):125–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/1/125.

Botelho:2000:LBC

[BTD^{+}00] Roberto J. Botelho, Mary Teruel, Renee Dierckman, Richard Anderson, Alan Wells, John D. York, Tobias Meyer, and Sergio Grinstein. Localized Biphasic Changes in Phosphatidylinositol-4,5-Bisphosphate at Sites of Phagocytosis. Journal of Cell Biology, 151(7):1353–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
7/1353.

Berson:2003:PCC

[BTH^{+}03] Joanne F. Berson, Alexander C. Theos, Dawn C. Harper, Danielle Tenza, Graça Raposo, and Michael S. Marks. Proprotein convertase cleavage liberates a fibrillogenic fragment of a resident glycoprotein to initiate melanosome biogenesis. Journal of Cell Biology, 161(3):521–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/3/521.

Buvelot:2003:BYI

[BTVB03] Stéphanie Buvelot, Sean Y. Tatsutani, Danielle Vermaak, and Sue Biggins. The budding yeast Ipl1/Aurora protein kinase regulates mitotic spindle disassembly. Journal of Cell Biology, 160(3):329–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/3/329.

Burke:2001:LA

[Bur01] Brian Burke. Lamins and Apoptosis. Journal of Cell Biology, 153(3):0–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/3/F5.

Blache:2004:SIC

[BvdWD^{+}04] Philippe Blache, Marc van de Wetering, Isabelle Duluc, Claire Domon, Philippe Berta, Jean-Noël Freund, Hans Clevers, and Philippe Jay. SOX9 is an intestine crypt transcription factor, is regulated by the Wnt pathway, and represses the CDX2 and MUC2 genes. Journal of Cell Biology, 166(1):37–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/1/37.

Bressan:2004:MTD

[BVH04] Debra A. Bressan, Julio Vazquez, and James E. Haber. Mating type-dependent constraints on the mobility of the left arm of yeast chromosome III. Journal of Cell Biology, 164(3): 361–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/3/361.

Bivona:2004:RRA

[BWA^{+}04] Trever G. Bivona, Heidi H. Wiener, Ian M. Ahearn, Joseph Silletti, Vi K. Chiu, and Mark R. Philips. Rap1 up-regulation and activation on plasma membrane regulates T cell adhesion. Journal of Cell Biology, 164(3):461–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/3/
461.

Bilodeau:2003:VHE

[BWK^{+}03] Patricia S. Bilodeau, Stanley C. Winistorfer, William R. Kearney, Andrew D. Robertson, and Robert C. Piper. Vps27-Hse1 and ESCRT–I complexes cooperate to increase efficiency of sorting ubiquitinated proteins at the endosome. Journal of Cell Biology, 163(2):237–??, October 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/163/2/
237.

Burkin:2001:EEI

[BWN^{+}01] Dean J. Burkin, Gregory Q. Wallace, Kimberly J. Nicol, David J. Kaufman, and Stephen J. Kaufman. Enhanced Expression of the α7β1 Integrin Reduces Muscular Dystrophy and Restores Viability in Dystrophic Mice. Journal of Cell Biology, 152(6):1207–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/6/1207.

Bao:2003:BSN

[BWRT03] Jianxin Bao, Deon Wolpowitz, Lorna W. Role, and David A. Talmage. Back signaling by the Nrg-1 intracellular domain. Journal of Cell Biology, 161(6):1133–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/6/1133.

Bercher:2001:MGR

[BWV^{+}01] Mark Bercher, Jim Wahl, Bruce E. Vogel, Charles Lu, Edward M. Hedgecock, David H. Hall, and John D. Plenefisch. mua-3, a gene required for mechanical tissue integrity in Caenorhabditis elegans, encodes a novel transmembrane protein of epithelial attachment complexes. Journal of Cell Biology, 154(2):415–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/2/415.

Burgess:2002:PSC

[BWW^{+}02] Stan Burgess, Matt Walker, Fei Wang, James R. Sellers, Howard D. White, Peter J. Knight, and John Trinick. The prepower stroke conformation of myosin V. Journal of Cell Biology, 159(6):983–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/6/983.

Bommel:2002:MMT

[BXR^{+}02] Heike Bömmel, Gang Xie, Wilfried Rossoll, Stefan Wiese, Sibylle Jablonka, Thomas Boehm, and Michael Sendtner. Missense mutation in the tubulin-specific chaperone E (Tbce) gene in the mouse mutant progressive motor neuronopathy, a model of human motoneuron disease. Journal of Cell Biology, 159(4):563–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/4/563.

Ben-Yaacov:2001:WDW

[BYLA^{+}01] Sari Ben-Yaacov, Roland Le Borgne, Irit Abramson, François Schweisguth, and Eyal D. Schejter. Wasp, the Drosophila Wiskott–Aldrich syndrome gene homologue, is required for cell fate decisions mediated by notch signaling. Journal of Cell Biology, 152(1):1–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/1/1.

Barreiro:2002:DIV

[BYMS^{+}02] Olga Barreiro, María Yáñez-Mó, Juan M. Serrador, María C. Montoya, Miguel Vicente-Manzanares, Reyes Tejedor, Heinz Furthmayr, and Francisco Sánchez-Madrid. Dynamic interaction of VCAM-1 and ICAM-1 with moesin and ezrin in a novel endothelial docking structure for adherent leukocytes. Journal of Cell Biology, 157(7):1233–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/7/1233.

Bourguignon:2000:ATI

[BZSC00] Lilly Y. W. Bourguignon, Hongbo Zhu, Lijun Shao, and Yue Wei Chen. Ankyrin–Tiam1 Interaction Promotes Rac1 Signaling and Metastatic Breast Tumor Cell Invasion and Migration. Journal of Cell Biology, 150(1):177–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/1/
177.

Chen:2000:DAB

[CA00] Po Chen and John M. Abrams. Drosophila Apoptosis and Bcl-2 Genes. Journal of Cell Biology, 148(4):625–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
148/4/625.

Cochilla:2000:DRG

[CAB00] Amanda J. Cochilla, Joseph K. Angleson, and William J. Betz. Differential Regulation of Granule–to–Granule and Granule–to–Plasma Membrane Fusion during Secretion from Rat Pituitary Lactotrophs. Journal of Cell Biology, 150(4): 839–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/4/839.

Czekay:2003:PAI

[CACL03] Ralf-Peter Czekay, Kathleen Aertgeerts, Scott A. Curriden, and David J. Loskutoff. Plasminogen activator inhibitor-1 detaches cells from extracellular matrices by inactivating integrins. Journal of Cell Biology, 160(5):781–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/5/
781.

Churin:2003:HPC

[CAGK^{+}03] Yuri Churin, Laila Al-Ghoul, Oliver Kepp, Thomas F. Meyer, Walter Birchmeier, and Michael Naumann. Helicobacter pylori CagA protein targets the c–Met receptor and enhances the motogenic response. Journal of Cell Biology, 161(2): 249–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/2/249.

Camacho:2003:MPS

[Cam03] Patricia Camacho. Malaria parasites solve the problem of a low calcium environment. Journal of Cell Biology, 161 (1):17–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/1/17.

Czubryt:2000:HPI

[CAP00] Michael P. Czubryt, J. Alejandro Austria, and Grant N. Pierce. Hydrogen Peroxide Inhibition of Nuclear Protein Import Is Mediated by the Mitogen–Activated Protein Kinase, Erk2. Journal of Cell Biology, 148(1):7–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/1/7.

Chittka:2004:PIP

[CARG^{+}04] Alexandra Chittka, Juan Carlos Arevalo, Maria Rodriguez-Guzman, Pilar Pérez, Moses V. Chao, and Michael Sendtner. The p75NTR-interacting protein SC1 inhibits cell cycle progression by transcriptional repression of cyclin E. Journal of Cell Biology, 164(7):985–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/7/985.

Chernyavsky:2004:NSP

[CAW^{+}04] Alex I. Chernyavsky, Juan Arredondo, Jürgen Wess, Evert Karlsson, and Sergei A. Grando. Novel signaling pathways mediating reciprocal control of keratinocyte migration and wound epithelialization through M_{3} and M_{4} muscarinic receptors. Journal of Cell Biology, 166(2):261–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/166/2/
261.

Cao:2000:ARR

[CB00] Xiaochun Cao and Charles Barlowe. Asymmetric Requirements for a Rab Gtpase and Snare Proteins in Fusion of Copii Vesicles with Acceptor Membranes. Journal of Cell Biology, 149(1):55–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/1/55.

Cupers:2001:DBP

[CBC^{+}01] Philippe Cupers, Mustapha Bentahir, Katleen Craessaerts, Isabelle Orlans, Hugo Vanderstichele, Paul Saftig, Bart De Strooper, and Wim Annaert. The discrepancy between presenilin subcellular localization and γ-secretase processing of amyloid precursor protein. Journal of Cell Biology, 154(4): 731–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/4/731.

Chidgey:2001:MLD

[CBG^{+}01] Martyn Chidgey, Cord Brakebusch, Erika Gustafsson, Alan Cruchley, Chris Hail, Sarah Kirk, Anita Merritt, Alison North, Chris Tselepis, Jane Hewitt, Carolyn Byrne, Reinhard Fassler, and David Garrod. Mice lacking desmocollin 1 show epidermal fragility accompanied by barrier defects and abnormal differentiation. Journal of Cell Biology, 155(5): 821–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/5/821.

Cirulli:2000:EFI

[CBK^{+}00] Vincenzo Cirulli, Gillian M. Beattie, George Klier, Mark Ellisman, Camillo Ricordi, Vito Quaranta, Francine Frasier, Jennifer K. Ishii, Alberto Hayek, and Daniel R. Salomon. Expression and function of α_{v}β_{3} and α_{v}β_{5} integrins in the developing pancreas. Journal of Cell Biology, 150(6):1445– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/6/1445.

Chung:2002:IRE

[CBL^{+}02] Jun Chung, Robin E. Bachelder, Elizabeth A. Lipscomb, Leslie M. Shaw, and Arthur M. Mercurio. Integrin (α6β4) regulation of eIF-4E activity and VEGF translation. Journal of Cell Biology, 158(1):165–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/1/165.

Clarke:2004:PAR

[CBLT04] Dominic M. Clarke, Michael C. Brown, David P. LaLonde, and Christopher E. Turner. Phosphorylation of actopaxin regulates cell spreading and migration. Journal of Cell Biology, 166(6):901–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/6/901.

Cohen:2004:MPD

[CBRBM04] David Cohen, Patrick J. Brennwald, Enrique Rodriguez-Boulan, and Anne Müsch. Mammalian PAR-1 determines epithelial lumen polarity by organizing the microtubule cytoskeleton. Journal of Cell Biology, 164(5):717–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
5/717.

Cougot:2004:CFS

[CBS04] Nicolas Cougot, Sylvie Babajko, and Bertrand Séraphin. Cytoplasmic foci are sites of mRNA decay in human cells. Journal of Cell Biology, 165(1):31–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/1/31.

Cheeseman:2001:INM

[CBW^{+}01] Iain M. Cheeseman, Christine Brew, Michael Wolyniak, Arshad Desai, Scott Anderson, Nemone Muster, John R. Yates, Tim C. Huffaker, David G. Drubin, and Georjana Barnes. Implication of a novel multiprotein Dam1p complex in outer kinetochore function. Journal of Cell Biology, 155(7):1137– ??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/7/1137.

Chan:2000:RCE

[CBZ^{+}00] Amanda Y. Chan, Maryse Bailly, Noureddine Zebda, Jeffrey E. Segall, and John S. Condeelis. Role of Cofilin in Epidermal Growth Factor–Stimulated Actin Polymerization and Lamellipod Protrusion. Journal of Cell Biology, 148(3): 531–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/3/531.

Carlton:2002:TAA

[CC02] Peter M. Carlton and W. Zacheus Cande. Telomeres act autonomously in maize to organize the meiotic bouquet from a semipolarized chromosome orientation. Journal of Cell Biology, 157(2):231–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/2/231.

Chen:2000:RYF

[CD00] Linyi Chen and Nicholas G. Davis. Recycling of the Yeast a -Factor Receptor. Journal of Cell Biology, 151(3):731–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
151/3/731.

Cheeseman:2002:SCC

[CDB02] Iain M. Cheeseman, David G. Drubin, and Georjana Barnes. Simple centromere, complex kinetochore. Journal of Cell Biology, 157(2):199–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/2/199.

Caldelari:2001:CRA

[CdBB^{+}01] Reto Caldelari, Alain de Bruin, Dominique Baumann, Maja M. Suter, Christiane Bierkamp, Vreni Balmer, and Eliane Müller. A Central Role for the Armadillo Protein Plakoglobin in the Autoimmune Disease Pemphigus Vulgaris. Journal of Cell Biology, 153(4):823–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/4/823.

Chen:2003:MMM

[CDE^{+}03] Hsiuchen Chen, Scott A. Detmer, Andrew J. Ewald, Erik E. Griffin, Scott E. Fraser, and David C. Chan. Mitofusins Mfn1 and Mfn2 coordinately regulate mitochondrial fusion and are essential for embryonic development. Journal of Cell Biology, 160(2):189–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/2/189.

Colanzi:2000:SAM

[CDEM00] Antonino Colanzi, Thomas J. Deerinck, Mark H. Ellisman, and Vivek Malhotra. A Specific Activation of the Mitogen– Activated Protein Kinase Kinase 1 (Mek1) Is Required for Golgi Fragmentation during Mitosis. Journal of Cell Biology, 149(2):331–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/2/331.

Cifuentes-Diaz:2001:DMS

[CDFT^{+}01] Carmen Cifuentes-Diaz, Tony Frugier, Francesco D. Tiziano, Emmanuelle Lacène, Natacha Roblot, Vandana Joshi, Marie Helene Moreau, and Judith Melki. Deletion of Murine SMN Exon 7 Directed to Skeletal Muscle Leads to Severe Muscular Dystrophy. Journal of Cell Biology, 152(5):1107– ??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/152/5/1107.

Cakouros:2004:ERD

[CDK04] Dimitrios Cakouros, Tasman J. Daish, and Sharad Kumar. Ecdysone receptor directly binds the promoter of the Drosophila caspase dronc, regulating its expression in specific tissues. Journal of Cell Biology, 165(5):631–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/165/
5/631.

Cambi:2004:MCT

[CdLvM^{+}04] Alessandra Cambi, Frank de Lange, Noortje M. van Maarseveen, Monique Nijhuis, Ben Joosten, Erik M. H. P. van Dijk, Bärbel I. de Bakker, Jack A. M. Fransen, Petra H. M. Bovee-Geurts, Frank N. van Leeuwen, Niek F. Van Hulst, and Carl G. Figdor. Microdomains of the C-type lectin DC– SIGN are portals for virus entry into dendritic cells. Journal of Cell Biology, 164(1):145–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/1/145.

Cakouros:2002:EIE

[CDM^{+}02] Dimitrios Cakouros, Tasman Daish, Damali Martin, Eric H. Baehrecke, and Sharad Kumar. Ecdysone-induced expression of the caspase DRONC during hormone-dependent programmed cell death in Drosophila is regulated by Broad– Complex. Journal of Cell Biology, 157(6):985–??, June 2002.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/157/6/
985.

Catlett:2000:TDR

[CDTW00] Natalie L. Catlett, Jason E. Duex, Fusheng Tang, and Lois S. Weisman. Two Distinct Regions in a Yeast Myosin–V Tail Domain Are Required for the Movement of Different Cargoes. Journal of Cell Biology, 150(3):513–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/3/
513.

Cheshenko:2003:HSV

[CDW^{+}03] Natalia Cheshenko, Brian Del Rosario, Craig Woda, Daniel Marcellino, Lisa M. Satlin, and Betsy C. Herold. Herpes simplex virus triggers activation of calcium-signaling pathways. Journal of Cell Biology, 163(2):283–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/163/2/
283.

Crouch:2001:AET

[CDWB01] Michael F. Crouch, Deborah A. Davy, Francis S. Willard, and Leise A. Berven. Activation of Endogenous Thrombin Receptors Causes Clustering and Sensitization of Epidermal Growth Factor Receptors of Swiss 3t3 Cells without Transactivation. Journal of Cell Biology, 152(2):263–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
2/263.

Carreno:2004:ADC

[CEGZ^{+}04] Sebastien Carreno, Åsa E. Engqvist-Goldstein, Claire X. Zhang, Kent L. McDonald, and David G. Drubin. Actin dynamics coupled to clathrin-coated vesicle formation at the trans-Golgi network. Journal of Cell Biology, 165(6):781–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
165/6/781.

Cheeseman:2001:MSI

[CENMR^{+}01] Iain M. Cheeseman, Maria Enquist-Newman, Thomas Müller-Reichert, David G. Drubin, and Georjana Barnes.

Mitotic spindle integrity and kinetochore function linked by the Duo1p/Dam1p complex. Journal of Cell Biology, 152(1): 197–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/1/197.

Chiroli:2003:BYP

[CFB^{+}03] Elena Chiroli, Roberta Fraschini, Alessia Beretta, Mariagrazia Tonelli, Giovanna Lucchini, and Simonetta Piatti. Budding yeast PAK kinases regulate mitotic exit by two different mechanisms. Journal of Cell Biology, 160(6):857–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/160/6/857.

Crawford:2000:ADA

[CFC^{+}00] Gregory E. Crawford, John A. Faulkner, Rachelle H. Crosbie, Kevin P. Campbell, Stanley C. Froehner, and Jeffrey S. Chamberlain. Assembly of the Dystrophin–Associated Protein Complex Does Not Require the Dystrophin Cooh– Terminal Domain. Journal of Cell Biology, 150(6):1399– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/6/1399.

Cau:2001:NPA

[CFC^{+}01] Julien Cau, Sandrine Faure, Michel Comps, Claude Delsert, and Nathalie Morin. A novel p21-activated kinase binds the actin and microtubule networks and induces microtubule stabilization. Journal of Cell Biology, 155(6):1029–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
6/1029.

Cregan:2002:AIF

[CFM^{+}02] Sean P. Cregan, Andre Fortin, Jason G. MacLaurin, Steven M. Callaghan, Francesco Cecconi, Seong-Woon Yu, Ted M. Dawson, Valina L. Dawson, David S. Park, Guido Kroemer, and Ruth S. Slack. Apoptosis-inducing factor is involved in the regulation of caspase-independent neuronal cell death. Journal of Cell Biology, 158(3):507–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
3/507.

Chappuis-Flament:2001:MCE

[CFWH^{+}01] Sophie Chappuis-Flament, Ellen Wong, Les D. Hicks, Cyril M. Kay, and Barry M. Gumbiner. Multiple cadherin extracellular repeats mediate homophilic binding and adhesion. Journal of Cell Biology, 154(1):231–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/1/231.

Chou:2000:IFM

[CG00] Ying-Hao Chou and Robert D. Goldman. Intermediate Filaments on the Move. Journal of Cell Biology, 150(3):??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
150/3/F101.

Carnahan:2003:PFP

[CG03] Robert H. Carnahan and Kathleen L. Gould. The PCH family protein, Cdc15p, recruits two F-actin nucleation pathways to coordinate cytokinetic actin ring formation in Schizosaccharomyces pombe. Journal of Cell Biology, 162(5):851– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/5/851.

Cornforth:2002:CPL

[CGBL^{+}02] Michael N. Cornforth, Karin M. Greulich-Bode, Bradford D. Loucas, Javier Arsuaga, Mariel Vázquez, Rainer K. Sachs, Martina Brückner, Michael Molls, Philip Hahnfeldt, Lynn Hlatky, and David J. Brenner. Chromosomes are predominantly located randomly with respect to each other in interphase human cells. Journal of Cell Biology, 159(2):237– ??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/2/237.

Caroppo:2004:ECA

[CGF^{+}04] Rosa Caroppo, Andrea Gerbino, Gregorio Fistetto, Matilde Colella, Lucantonio Debellis, Aldebaran M. Hofer, and Silvana Curci. Extracellular calcium acts as a “third messenger” to regulate enzyme and alkaline secretion. Journal of Cell Biology, 166(1):111–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/1/111.

Chami:2001:STP

[CGL^{+}01] Mounia Chami, Devrim Gozuacik, David Lagorce, Marisa Brini, Pierre Falson, Gérard Peaucellier, Paolo Pinton, Hervé Lecoeur, Marie-Lyse Gougeon, Marc le Maire, Rosario Rizzuto, Christian Bréchot, and Patrizia Paterlini-Bréchot. Serca1 Truncated Proteins Unable to Pump Calcium Reduce the Endoplasmic Reticulum Calcium Concentration and Induce Apoptosis. Journal of Cell Biology, 153(6):1301–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/6/1301.

Calautti:2002:FTK

[CGM^{+}02] Enzo Calautti, Maddalena Grossi, Cristina Mammucari, Yumi Aoyama, Maria Pirro, Yoshitaka Ono, Jie Li, and G. Paolo Dotto. Fyn tyrosine kinase is a downstream mediator of Rho/PRK2 function in keratinocyte cell–cell adhesion. Journal of Cell Biology, 156(1):137–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/1/
137.

Chen:2001:ERF

[CGY01a] Chin-Yi Chen, Jeanine Graham, and Hong Yan. Evidence for a Replication Function of Ffa-1, the Xenopus Orthologue of Werner Syndrome Protein. Journal of Cell Biology, 152 (5):985–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/5/985.

Chen:2001:WSI

[CGY^{+}01b] Shaoqiong Chen, Denis C. Guttridge, Zongbing You, Zhaochen Zhang, Andrew Fribley, Marty W. Mayo, Jan Kitajewski, and Cun-Yu Wang. WNT-1 signaling inhibits apoptosis by activating β-catenin/T cell factor-mediated transcription. Journal of Cell Biology, 152(1):87–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
1/87.

Chen:2001:NCI

[CH01] Danyang Chen and Sui Huang. Nucleolar Components Involved in Ribosome Biogenesis Cycle between the Nucleolus and Nucleoplasm in Interphase Cells. Journal of Cell Biology, 153(1):169–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/1/169.

Cuvier:2003:RTI

[CH03] Olivier Cuvier and Tatsuya Hirano. A role of topoisomerase II in linking DNA replication to chromosome condensation. Journal of Cell Biology, 160(5):645–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/5/645.

Chamberlain:2000:MDM

[Cha00] Jeffrey S. Chamberlain. Muscular Dystrophy Meets the Gene Chip. Journal of Cell Biology, 151(6):0–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/6/
F43.

Caplan:2001:HVP

[CHA^{+}01] Steve Caplan, Lisa M. Hartnell, Rubén C. Aguilar, Naava Naslavsky, and Juan S. Bonifacino. Human Vam6p promotes lysosome clustering and fusion in vivo. Journal of Cell Biology, 154(1):109–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/1/109.

Chen:2002:BEK

[Che02] Rey-Huei Chen. BubR1 is essential for kinetochore localization of other spindle checkpoint proteins and its phosphorylation requires Mad1. Journal of Cell Biology, 158(3): 487–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/3/487.

Cimini:2001:MKO

[CHM^{+}01] Daniela Cimini, Bonnie Howell, Paul Maddox, Alexey Khodjakov, Francesca Degrassi, and E. D. Salmon. Merotelic Kinetochore Orientation Is a Major Mechanism of Aneuploidy in Mitotic Mammalian Tissue Cells. Journal of Cell Biology, 153(3):517–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/3/517.

Chang:2004:JDP

[CHM04] Henry C. Chang, Michael Hull, and Ira Mellman. The J-domain protein Rme-8 interacts with Hsc70 to control clathrin-dependent endocytosis in Drosophila. Journal of Cell Biology, 164(7):1055–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/7/1055.

Chen:2003:PCA

[CiKBG03] Xinyu Chen, Shin ichiro Kojima, Gary G. Borisy, and Kathleen J. Green. p120 catenin associates with kinesin and facilitates the transport of cadherin–catenin complexes to intercellular junctions. Journal of Cell Biology, 163(3):547– ??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/3/547.

Chang:2002:CIC

[CJ02] Louis K. Chang and Eugene M. Johnson. Cyclosporin A inhibits caspase-independent death of NGF-deprived sympathetic neurons. Journal of Cell Biology, 157(5):771–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/157/
5/771.

Cho:2000:ERK

[CK00] Samuel Y. Cho and Richard L. Klemke. Extracellularregulated kinase activation and CAS/Crk coupling regulate cell migration and suppress apoptosis during invasion of the extracellular matrix. Journal of Cell Biology, 149(1):223–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
149/1/223.

Cho:2002:PPP

[CK02] Samuel Y. Cho and Richard L. Klemke. Purification of pseudopodia from polarized cells reveals redistribution and activation of Rac through assembly of a CAS/Crk scaffold. Journal of Cell Biology, 156(4):725–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/4/725.

Chen:2003:LNR

[CK03] Esther J. Chen and Chris A. Kaiser. LST8 negatively regulates amino acid biosynthesis as a component of the TOR pathway. Journal of Cell Biology, 161(2):333–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/2/
333.

Chakrabarti:2003:IMR

[CKF^{+}03] Sabyasachi Chakrabarti, Koichi S. Kobayashi, Richard A. Flavell, Carolyn B. Marks, Katsuya Miyake, David R. Liston, Kimberly T. Fowler, Fred S. Gorelick, and Norma W. Andrews. Impaired membrane resealing and autoimmune myositis in synaptotagmin VII–deficient mice. Journal of Cell Biology, 162(4):543–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/4/543.

Cronin:2000:RHC

[CKFH00] Stephen R. Cronin, Afif Khoury, Dana K. Ferry, and Randolph Y. Hampton. Regulation of HMG–Coa reductase degradation requires the P-type ATPase Cod1p/Spf1p. Journal of Cell Biology, 148(5):915–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/5/915.

Cho:2003:INA

[CKL^{+}03] Ssang-Goo Cho, Jin Woo Kim, Yong Hee Lee, Hyun Sub Hwang, Mi-Sung Kim, Kanghyun Ryoo, Myung Jin Kim, Kyung Tae Noh, Eun Kyung Kim, Jun-Ho Cho, Kyoung Wan Yoon, Eun-Gyung Cho, Hee-Sae Park, Sung Wook Chi, Min-Jae Lee, Sang Sun Kang, Hidenori Ichijo, and Eui-Ju Choi. Identification of a novel antiapoptotic protein that antagonizes ASK1 and CAD activities. Journal of Cell Biology, 163 (1):71–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/1/71.

Chellaiah:2000:GDB

[CKS^{+}00] Meenakshi Chellaiah, Neil Kizer, Matthew Silva, Ulises Alvarez, David Kwiatkowski, and Keith A. Hruska. Gelsolin Deficiency Blocks Podosome Assembly and Produces Increased Bone Mass and Strength. Journal of Cell Biology, 148(4): 665–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/4/665.

Chen:2002:NMR

[CKS02] Qi Chen, Hideo Kimura, and David Schubert. A novel mechanism for the regulation of amyloid precursor protein metabolism. Journal of Cell Biology, 158(1):79–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
1/79.

Ciani:2004:DCW

[CKS^{+}04] Lorenza Ciani, Olga Krylova, Matthew J. Smalley, Trevor C. Dale, and Patricia C. Salinas. A divergent canonical WNTsignaling pathway regulates microtubule dynamics. Journal of Cell Biology, 164(2):243–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/2/243.

Cremer:2003:IGD

[CKW^{+}03] Marion Cremer, Katrin Küpper, Babett Wagler, Leah Wizelman, Johann v. Hase, Yanina Weiland, Ludwika Kreja, Joachim Diebold, Michael R. Speicher, and Thomas Cremer. Inheritance of gene density–related higher order chromatin arrangements in normal and tumor cell nuclei. Journal of Cell Biology, 162(5):809–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/5/809.

Cronshaw:2002:PAM

[CKZ^{+}02] Janet M. Cronshaw, Andrew N. Krutchinsky, Wenzhu Zhang, Brian T. Chait, and Michael J. Matunis. Proteomic analysis of the mammalian nuclear pore complex. Journal of Cell Biology, 158(5):915–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/5/915.

Charpentier:2000:PSE

[CLAC00] Emmanuelle Charpentier, Robert M. Lavker, Elizabeth Acquista, and Pamela Cowin. Plakoglobin Suppresses Epithelial Proliferation and Hair Growth in Vivo. Journal of Cell Biology, 149(2):503–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/2/503.

Christensen:2002:DHD

[CLB^{+}02] Morten O. Christensen, Morten K. Larsen, Hans Ullrich Barthelmes, Robert Hock, Claus L. Andersen, Eigil Kjeldsen, Birgitta R. Knudsen, Ole Westergaard, Fritz Boege, and Christian Mielke. Dynamics of human DNA topoisomerases IIα and IIβ in living cells. Journal of Cell Biology, 157 (1):31–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/1/31.

Castoria:2003:ASD

[CLB^{+}03] Gabriella Castoria, Maria Lombardi, Maria Vittoria Barone, Antonio Bilancio, Marina Di Domenico, Daniela Bottero, Flavia Vitale, Antimo Migliaccio, and Ferdinando Auricchio. Androgen-stimulated DNA synthesis and cytoskeletal changes in fibroblasts by a nontranscriptional receptor action. Journal of Cell Biology, 161(3):547–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/3/547.

Cattelino:2003:CIC

[CLG^{+}03] Anna Cattelino, Stefan Liebner, Radiosa Gallini, Adriana Zanetti, Giovanna Balconi, Alessandro Corsi, Paolo Bianco, Hartwig Wolburg, Robert Moore, Boussadia Oreda, Rolf Kemler, and Elisabetta Dejana. The conditional inactivation of the β-catenin gene in endothelial cells causes a defective vascular pattern and increased vascular fragility. Journal of Cell Biology, 162(6):1111–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/6/1111.

Chen:2003:ABM

[CLM^{+}03] Yuzhi Chen, Wenyun Liu, Donna L. McPhie, Linda Hassinger, and Rachael L. Neve. APP-BP1 mediates APPinduced apoptosis and DNA synthesis and is increased in Alzheimer’s disease brain. Journal of Cell Biology, 163(1): 27–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/1/27.

Canty:2004:CPM

[CLM^{+}04] Elizabeth G. Canty, Yinhui Lu, Roger S. Meadows, Michael K. Shaw, David F. Holmes, and Karl E. Kadler. Coalignment of plasma membrane channels and protrusions (fibripositors) specifies the parallelism of tendon. Journal of Cell Biology, 165(4):553–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/4/553.

Curran:2002:RTT

[CLSK02] Sean P. Curran, Danielle Leuenberger, Einhard Schmidt, and Carla M. Koehler. The role of the Tim8p–Tim13p complex in a conserved import pathway for mitochondrial polytopic inner membrane proteins. Journal of Cell Biology, 158(6):1017– ??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/6/1017.

Chun:2001:TTD

[CLT^{+}01] Wanjoo Chun, Mathieu Lesort, Janusz Tucholski, Christopher A. Ross, and Gail V. W. Johnson. Tissue Transglutaminase Does Not Contribute to the Formation of Mutant Huntingtin Aggregates. Journal of Cell Biology, 153 (1):25–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/1/25.

Cao:2001:IMM

[CLWR01] Tongyu Cao, Mary Ann Longley, Xiao-Jing Wang, and Dennis R. Roop. An Inducible Mouse Model for Epidermolysis Bullosa Simplex. Journal of Cell Biology, 152(3):651– ??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/3/651.

Cline:2001:TPD

[CM01] Kenneth Cline and Hiroki Mori. Thylakoid Δ pH-dependent precursor proteins bind to a cpTatC–Hcf106 complex before Tha4-dependent transport. Journal of Cell Biology, 154(4): 719–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/4/719.

Charrasse:2002:CDC

[CMC^{+}02] Sophie Charrasse, Mayya Meriane, Franck Comunale, Anne Blangy, and Cécile Gauthier-Rouvière. N-cadherin-dependent cell–cell contact regulates Rho GTPases and βcatenin localization in mouse C2C12 myoblasts. Journal of Cell Biology, 158(5):953–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/5/953.

Claycomb:2002:VRI

[CME^{+}02] Julie M. Claycomb, David M. MacAlpine, James G. Evans, Stephen P. Bell, and Terry L. Orr-Weaver. Visualization of replication initiation and elongation in Drosophila. Journal of Cell Biology, 159(2):225–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/2/225.

Castillo:2002:YPK

[CMM^{+}02] Andrea R. Castillo, Janet B. Meehl, Garry Morgan, Amy Schutz-Geschwender, and Mark Winey. The yeast protein kinase Mps1p is required for assembly of the integral spindle pole body component Spc42p. Journal of Cell Biology, 156(3):453–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/3/453.

Creancier:2000:FGF

[CMMP00] Laurent Créancier, Dominique Morello, Pascale Mercier, and Anne-Catherine Prats. Fibroblast Growth Factor 2 Internal Ribosome Entry Site (Ires) Activity Ex Vivo and in Transgenic Mice Reveals a Stringent Tissue–Specific Regulation. Journal of Cell Biology, 150(1):275–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/1/275.

Cioffi:2002:DRI

[CMS^{+}02] Donna L. Cioffi, Timothy M. Moore, Jerry Schaack, Judy R. Creighton, Dermot M. F. Cooper, and Troy Stevens. Dominant regulation of interendothelial cell gap formation by calcium-inhibited type 6 adenylyl cyclase. Journal of Cell Biology, 157(7):1267–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/7/1267.

Carenini:2001:RMD

[CMW^{+}01] Stefano Carenini, Mathias Mäurer, Alexander Werner, Heinrich Blazyca, Klaus V. Toyka, Christoph D. Schmid, Gennadij Raivich, and Rudolf Martini. The Role of Macrophages in Demyelinating Peripheral Nervous System of Mice Heterozygously Deficient in P0. Journal of Cell Biology, 152(2): 301–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/2/301.

Charron:2000:CCP

[CNBWN00] Audra J. Charron, Sakie Nakamura, Robert Bacallao, and Angela Wandinger-Ness. Compromised Cytoarchitecture and Polarized Trafficking in Autosomal Dominant Polycystic Kidney Disease Cells. Journal of Cell Biology, 149(1):111–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
149/1/111.

Chang:2002:HRE

[CNH^{+}02] Henry C. Chang, Sherri L. Newmyer, Michael J. Hull, Melanie Ebersold, Sandra L. Schmid, and Ira Mellman. Hsc70 is required for endocytosis and clathrin function in Drosophila. Journal of Cell Biology, 159(3):477–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
159/3/477.

Cai:2002:GSP

[CNHK02] Ti Cai, Keigo Nishida, Toshio Hirano, and Paul A. Khavari. Gab1 and SHP-2 promote Ras/MAPK regulation of epidermal growth and differentiation. Journal of Cell Biology, 159(1):103–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/1/103.

Csankovszki:2001:SXR

[CNJ01] Györgyi Csankovszki, András Nagy, and Rudolf Jaenisch. Synergism of Xist RNA, DNA Methylation, and Histone Hypoacetylation in Maintaining X Chromosome Inactivation. Journal of Cell Biology, 153(4):773–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/4/773.

Chiu:2002:CCF

[CNMS02] Raymond Chiu, Leonid Novikov, Shaeri Mukherjee, and Dennis Shields. A caspase cleavage fragment of p115 induces fragmentation of the Golgi apparatus and apoptosis. Journal of Cell Biology, 159(4):637–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/4/637.

Ciliberto:2003:MMM

[CNT03] Andrea Ciliberto, Bela Novak, and John J. Tyson. Mathematical model of the morphogenesis checkpoint in budding yeast. Journal of Cell Biology, 163(6):1243–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
6/1243.

Chen:2001:LCE

[COB01] Lihsia Chen, Bryan Ong, and Vann Bennett. LAD-1, the Caenorhabditis elegans L1CAM homologue, participates in embryonic and gonadal morphogenesis and is a substrate for fibroblast growth factor receptor pathway-dependent phosphotyrosine-based signaling. Journal of Cell Biology, 154 (4):841–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/4/841.

Cooper:2001:NRI

[Coo01] Ellis Cooper. A New Role for Ion Channels in Myoblast Fusion. Journal of Cell Biology, 153(4):0–??, May 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/153/4/
F9.

Cornelis:2002:YTI

[Cor02] Guy R. Cornelis. Yersinia type III secretion. Journal of Cell Biology, 158(3):401–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/3/401.

Carrano:2001:RFB

[CP01] Andrea C. Carrano and Michele Pagano. Role of the F–Box Protein Skp2 in Adhesion–Dependent Cell Cycle Progression. Journal of Cell Biology, 153(7):1381–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/7/1381.

Cichy:2003:LC

[CP03] Joanna Cichy and Ellen Puré. The liberation of CD44. Journal of Cell Biology, 161(5):839–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/5/839.

Christian:2003:NEC

[CPA^{+}03] Sven Christian, Jan Pilch, Maria E. Akerman, Kimmo Porkka, Pirjo Laakkonen, and Erkki Ruoslahti. Nucleolin expressed at the cell surface is a marker of endothelial cells in angiogenic blood vessels. Journal of Cell Biology, 163(4): 871–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/4/871.

Calapez:2002:IMM

[CPC^{+}02a] Alexandre Calapez, Henrique M. Pereira, Angelo Calado, José Braga, José Rino, Célia Carvalho, João Paulo Tavanez, Elmar Wahle, Agostinho C. Rosa, and Maria Carmo-Fonseca. The intranuclear mobility of messenger RNA binding proteins is ATP dependent and temperature sensitive. Journal of Cell Biology, 159(5):795–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/5/795.

Chevrier:2002:RAP

[CPC^{+}02b] Véronique Chevrier, Matthieu Piel, Nora Collomb, Yasmina Saoudi, Ronald Frank, Michel Paintrand, Shuh Narumiya, Michel Bornens, and Didier Job. The Rho-associated protein kinase p160ROCK is required for centrosome positioning. Journal of Cell Biology, 157(5):807–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/5/807.

Chiarugi:2003:ROS

[CPG^{+}03] Paola Chiarugi, Giovambattista Pani, Elisa Giannoni, Letizia Taddei, Renata Colavitti, Giovanni Raugei, Mark Symons, Silvia Borrello, Tommaso Galeotti, and Giampietro Ramponi. Reactive oxygen species as essential mediators of cell adhesion. Journal of Cell Biology, 161(5):933–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/5/
933.

Carmosino:2001:HTI

[CPN^{+}01] Monica Carmosino, Giuseppe Procino, Grazia Paola Nicchia, Roberta Mannucci, Jean-Marc Verbavatz, Renèe Gobin, Maria Svelto, and Giovanna Valenti. Histamine treatment induces rearrangements of orthogonal arrays of particles (OAPs) in human AQP4-expressing gastric cells. Journal of Cell Biology, 154(6):1235–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/6/1235.

Charroux:2000:G

[CPP^{+}00] Bernard Charroux, Livio Pellizzoni, Robert A. Perkinson, Jeongsik Yong, Andrej Shevchenko, Matthias Mann, and Gideon Dreyfuss. Gemin4. Journal of Cell Biology, 148(6): 1177–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/6/1177.

Colussi:2000:DPB

[CQH^{+}00] Paul A. Colussi, Leonie M. Quinn, David C. S. Huang, Michelle Coombe, Stuart H. Read, Helena Richardson, and Sharad Kumar. Debcl, a proapoptotic Bcl-2 homologue, is a component of the Drosophila melanogaster cell death machinery. Journal of Cell Biology, 148(4):703–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
4/703.

Cramer:2000:MV

[Cra00] Louise P. Cramer. Myosin VI. Journal of Cell Biology, 150 (6):??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/6/F121.

Chaudhry:2002:GC

[CRE02] Farrukh A. Chaudhry, Richard J. Reimer, and Robert H. Edwards. The glutamine commute. Journal of Cell Biology, 157 (3):349–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/3/349.

Cronin:2002:CSP

[CRH02] Stephen R. Cronin, Rajini Rao, and Randolph Y. Hampton. Cod1p/ spf1p is a P-type ATPase involved in ER function and Ca^{2+} homeostasis. Journal of Cell Biology, 157 (6):1017–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/6/1017.

Colognato:2004:IDS

[CROfC04] Holly Colognato, Shwetha Ramachandrappa, Inger M. Olsen, and Charles ffrench Constant. Integrins direct Src family kinases to regulate distinct phases of oligodendrocyte development. Journal of Cell Biology, 167(2):365–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
2/365.

Campellone:2004:CNR

[CRP^{+}04] Kenneth G. Campellone, Susannah Rankin, Tony Pawson, Marc W. Kirschner, Donald J. Tipper, and John M. Leong. Clustering of Nck by a 12-residue Tir phosphopeptide is sufficient to trigger localized actin assembly. Journal of Cell Biology, 164(3):407–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/3/407.

Chun:2003:ILK

[CRS^{+}03] Soo Jin Chun, Matthew N. Rasband, Richard L. Sidman, Amyn A. Habib, and Timothy Vartanian. Integrin-linked kinase is required for laminin-2–induced oligodendrocyte cell spreading and CNS myelination. Journal of Cell Biology, 163(2):397–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/2/397.

Chen:2000:SBD

[CS00] Yu Chen and Sanford M. Simon. In Situ Biochemical Demonstration That P–Glycoprotein Is a Drug Efflux Pump with Broad Specificity. Journal of Cell Biology, 148(5):863–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/148/5/863.

Cha:2001:TPE

[CS01] Hyukjin Cha and Paul Shapiro. Tyrosine-phosphorylated extracellular signal-regulated kinase associates with the Golgi complex during G2/M phase of the cell cycle. Journal of Cell Biology, 153(7):1355–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/7/1355.

Conner:2002:IAA

[CS02] Sean D. Conner and Sandra L. Schmid. Identification of an adaptor-associated kinase, AAK1, as a regulator of clathrinmediated endocytosis. Journal of Cell Biology, 156(5):921– ??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/156/5/921.

Chen:2003:LCS

[CS03a] Zu-Lin Chen and Sidney Strickland. Laminin γ1 is critical for Schwann cell differentiation, axon myelination, and regeneration in the peripheral nerve. Journal of Cell Biology, 163(4): 889–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/4/889.

Conner:2003:DRA

[CS03b] Sean D. Conner and Sandra L. Schmid. Differential requirements for AP-2 in clathrin-mediated endocytosis. Journal of Cell Biology, 162(5):773–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/5/773.

Carman:2004:TCL

[CS04] Christopher V. Carman and Timothy A. Springer. A transmigratory cup in leukocyte diapedesis both through individual vascular endothelial cells and between them. Journal of Cell Biology, 167(2):377–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/2/377.

Choy:2000:RAT

[CSD00] Lisa Choy, Jeremy Skillington, and Rik Derynck. Roles of Autocrine TGF-β Receptor and Smad Signaling in Adipocyte Differentiation. Journal of Cell Biology, 149(3):667–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
3/667.

Coscoy:2001:NCH

[CSG01] Laurent Coscoy, David Jesse Sanchez, and Don Ganem. A novel class of herpesvirus-encoded membrane-bound E3 ubiquitin ligases regulates endocytosis of proteins involved in immune recognition. Journal of Cell Biology, 155(7):1265– ??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/7/1265.

Cartaud:2004:MRA

[CSG^{+}04] Annie Cartaud, Laure Strochlic, Manuel Guerra, Benoît Blanchard, Monique Lambergeon, Eric Krejci, Jean Cartaud, and Claire Legay. MuSK is required for anchoring acetylcholinesterase at the neuromuscular junction. Journal of Cell Biology, 165(4):505–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/4/505.

Cepinskas:2003:PTM

[CSIK03] Gediminas Cepinskas, Jurate Savickiene, Carmen V. Ionescu, and Peter R. Kvietys. PMN transendothelial migration decreases nuclear NFκB in IL-1β-activated endothelial cells. Journal of Cell Biology, 161(3):641–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/3/641.

Coppens:2000:TGE

[CSJ00] Isabelle Coppens, Anthony P. Sinai, and Keith A. Joiner. Toxoplasma gondii exploits host low-density lipoprotein receptor-mediated endocytosis for cholesterol acquisition. Journal of Cell Biology, 149(1):167–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/1/167.

Chang:2003:AMP

[CSJ03a] Louis K. Chang, Robert E. Schmidt, and Eugene M. Johnson. Alternating metabolic pathways in NGF-deprived sympathetic neurons affect caspase-independent death. Journal of Cell Biology, 162(2):245–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/2/245.

Clarke:2003:CMD

[CSJ^{+}03b] Murray C. H. Clarke, John Savill, David B. Jones, Brendon S. Noble, and Simon B. Brown. Compartmentalized megakaryocyte death generates functional platelets committed to caspase-independent death. Journal of Cell Biology, 160(4):577–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/4/577.

Chazaud:2003:SCA

[CSL^{+}03] Bénédicte Chazaud, Corinne Sonnet, Peggy Lafuste, Guillaume Bassez, Anne-Cécile Rimaniol, Françoise Poron, François-Jérôme Authier, Patrick A. Dreyfus, and Romain K. Gherardi. Satellite cells attract monocytes and use macrophages as a support to escape apoptosis and enhance muscle growth. Journal of Cell Biology, 163(5):1133– ??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/5/1133.

Colanzi:2003:RAM

[CSM03] Antonino Colanzi, Christine Sutterlin, and Vivek Malhotra. RAF1-activated MEK1 is found on the Golgi apparatus in late prophase and is required for Golgi complex fragmentation in mitosis. Journal of Cell Biology, 161(1):27–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
1/27.

Chan:2004:CRC

[CSM04] Raymond C. Chan, Aaron F. Severson, and Barbara J. Meyer. Condensin restructures chromosomes in preparation for meiotic divisions. Journal of Cell Biology, 167(4):613– ??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/4/613.

Chun:2004:MMD

[CSO^{+}04] Tae-Hwa Chun, Farideh Sabeh, Ichiro Ota, Hedwig Murphy, Kevin T. McDonagh, Kenn Holmbeck, Henning Birkedal-Hansen, Edward D. Allen, and Stephen J. Weiss. MT1-MMP-dependent neovessel formation within the confines of the three-dimensional extracellular matrix. Journal of Cell Biology, 167(4):757–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/4/757.

Confalonieri:2000:TPE

[CSP^{+}00] Stefano Confalonieri, Anna Elisabetta Salcini, Claudia Puri, Carlo Tacchetti, and Pier Paolo Di Fiore. Tyrosine Phosphorylation of Eps15 Is Required for Ligand–Regulated, but Not Constitutive, Endocytosis. Journal of Cell Biology, 150 (4):905–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/4/905.

Camarda:2004:PIM

[CSP^{+}04] Grazia Camarda, Francesca Siepi, Deborah Pajalunga, Camilla Bernardini, Rossella Rossi, Alessandra Montecucco, Ettore Meccia, and Marco Crescenzi. A pRb-independent mechanism preserves the postmitotic state in terminally differentiated skeletal muscle cells. Journal of Cell Biology, 167(3):417–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/417.

Conacci-Sorrell:2003:ACE

[CSSBY^{+}03] Maralice Conacci-Sorrell, Inbal Simcha, Tamar Ben-Yedidia, Janna Blechman, Pierre Savagner, and Avri Ben-Ze’ev. Autoregulation of E-cadherin expression by cadherin–cadherin interactions. Journal of Cell Biology, 163(4):847–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
163/4/847.

Chu:2004:FMC

[CTE^{+}04] Yeh-Shiu Chu, William A. Thomas, Olivier Eder, Frederic Pincet, Eric Perez, Jean Paul Thiery, and Sylvie Dufour. Force measurements in E-cadherin-mediated cell doublets reveal rapid adhesion strengthened by actin cytoskeleton remodeling through Rac and Cdc42. Journal of Cell Biology, 167(6):1183–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/6/1183.

Chen:2004:BFI

[CVZ^{+}04] Rui Chen, Ignacio Valencia, Fei Zhong, Karen S. McColl, H. Llewelyn Roderick, Martin D. Bootman, Michael J. Berridge, Stuart J. Conway, Andrew B. Holmes, Gregory A. Mignery, Patricio Velez, and Clark W. Distelhorst. Bcl-2 functionally interacts with inositol 1,4,5-trisphosphate receptors to regulate calcium release from the ER in response to inositol 1,4,5-trisphosphate. Journal of Cell Biology, 166 (2):193–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/2/193.

Chadwick:2001:NCP

[CW01] Brian P. Chadwick and Huntington F. Willard. A Novel Chromatin Protein, Distantly Related to Histone H2a, Is Largely Excluded from the Inactive X Chromosome. Journal of Cell Biology, 152(2):375–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/2/375.

Chadwick:2002:CCD

[CW02] Brian P. Chadwick and Huntington F. Willard. Cell cycledependent localization of macroH2A in chromatin of the inactive X chromosome. Journal of Cell Biology, 157(7):1113–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
157/7/1113.

Canman:2004:RFO

[CW04] Julie C. Canman and William A. Wells. Rappaport Furrows on Our Minds. Journal of Cell Biology, 166(7):943– ??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/7/943.

Chattopadhyay:2003:ICC

[CWA^{+}03] Nibedita Chattopadhyay, Zemin Wang, Leonie K. Ashman, Susann M. Brady-Kalnay, and Jordan A. Kreidberg. α3β1 integrin–CD151, a component of the cadherin–catenin complex, regulates PTPμ expression and cell–cell adhesion. Journal of Cell Biology, 163(6):1351–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/6/1351.

Chew:2002:FRE

[CWG^{+}02] Teng-Leong Chew, Wendy A. Wolf, Patricia J. Gallagher, Fumio Matsumura, and Rex L. Chisholm. A fluorescent resonant energy transfer–based biosensor reveals transient and regional myosin light chain kinase activation in lamella and cleavage furrows. Journal of Cell Biology, 156(3):543– ??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/3/543.

Caulin:2000:KDE

[CWMO00] Carlos Caulin, Carl F. Ware, Thomas M. Magin, and Robert G. Oshima. Keratin-dependent, Epithelial Resistance to Tumor Necrosis Factor–Induced Apoptosis. Journal of Cell Biology, 149(1):17–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/1/17.

Chen:2004:SSA

[CYC^{+}04] Chun Chen, Bradford A. Young, Catherine S. Coleman, Anthony E. Pegg, and Dean Sheppard. Spermidine/ spermine N^{1}-acetyltransferase specifically binds to the integrin α9 subunit cytoplasmic domain and enhances cell migration. Journal of Cell Biology, 167(1):161–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/1/161.

Chen:2000:EPM

[CZBH00] Yi-Wen Chen, Po Zhao, Rehannah Borup, and Eric P. Hoffman. Expression Profiling in the Muscular Dystrophies. Journal of Cell Biology, 151(6):1321–??, December 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/6/
1321.

DeStrooper:2001:WNW

[DA01] Bart De Strooper and Wim Annaert. Where Notch and WNT Signaling Meet. Journal of Cell Biology, 152(4):0–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
152/4/F17.

Dunbar:2000:TSD

[DAC00] Lisa A. Dunbar, Paul Aronson, and Michael J. Caplan. A Transmembrane Segment Determines the Steady–State Localization of an Ion–Transporting Adenosine Triphosphatase. Journal of Cell Biology, 148(4):769–??, February 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/4/
769.

Du:2003:RPD

[DAV^{+}03] Guangwei Du, Yelena M. Altshuller, Nicolas Vitale, Ping Huang, Sylvette Chasserot-Golaz, Andrew J. Morris, Marie-France Bader, and Michael A. Frohman. Regulation of phospholipase D1 subcellular cycling through coordination of multiple membrane association motifs. Journal of Cell Biology, 162(2):305–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/2/305.

Daniely:2000:FCB

[DB00] Yaron Daniely and James A. Borowiec. Formation of a Complex between Nucleolin and Replication Protein a after Cell Stress Prevents Initiation of DNA Replication. Journal of Cell Biology, 149(4):799–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/149/4/799.

Denker:2002:CMR

[DB02] Sheryl P. Denker and Diane L. Barber. Cell migration requires both ion translocation and cytoskeletal anchoring by the Na–H exchanger NHE1. Journal of Cell Biology, 159(6): 1087–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/6/1087.

Degani:2002:ICD

[DBB^{+}02a] Simona Degani, Fiorella Balzac, Mara Brancaccio, Simona Guazzone, Saverio Francesco Retta, Lorenzo Silengo, Alessandra Eva, and Guido Tarone. The integrin cytoplasmic domain-associated protein ICAP-1 binds and regulates Rho family GTPases during cell spreading. Journal of Cell Biology, 156(2):377–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/2/377.

DeMali:2002:RAC

[DBB02b] Kris A. DeMali, Christy A. Barlow, and Keith Burridge. Recruitment of the Arp2/3 complex to vinculin. Journal of Cell Biology, 159(5):881–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/5/881.

Daigle:2001:NPC

[DBH^{+}01] Nathalie Daigle, Joël Beaudouin, Lisa Hartnell, Gabriela Imreh, Einar Hallberg, Jennifer Lippincott-Schwartz, and Jan Ellenberg. Nuclear pore complexes form immobile networks and have a very low turnover in live mammalian cells. Journal of Cell Biology, 154(1):71–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/1/71.

Dijkers:2002:FLC

[DBL^{+}02] Pascale F. Dijkers, Kim U. Birkenkamp, Eric W.-F. Lam, N. Shaun B. Thomas, Jan-Willem J. Lammers, Leo Koenderman, and Paul J. Coffer. FKHR-l1 can act as a critical effector of cell death induced by cytokine withdrawal. Journal of Cell Biology, 156(3):531–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/3/531.

Dou:2002:PAD

[DBLG02] Yali Dou, Josephine Bowen, Yifan Liu, and Martin A. Gorovsky. Phosphorylation and an ATP-dependent process increase the dynamic exchange of H1 in chromatin. Journal of Cell Biology, 158(7):1161–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/7/1161.

Degryse:2001:HMG

[DBS^{+}01] Bernard Degryse, Tiziana Bonaldi, Paola Scaffidi, Susanne Müller, Massimo Resnati, Francesca Sanvito, Gianluigi Arrigoni, and Marco E. Bianchi. The High Mobility Group (Hmg) Boxes of the Nuclear Protein Hmg1 Induce Chemotaxis and Cytoskeleton Reorganization in Rat Smooth Muscle Cells. Journal of Cell Biology, 152(6):1197–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
6/1197.

Dujardin:2003:RCD

[DBS^{+}03] Denis L. Dujardin, Lora E. Barnhart, Stephanie A. Stehman, Edgar R. Gomes, Gregg G. Gundersen, and Richard B. Vallee. A role for cytoplasmic dynein and LIS1 in directed cell movement. Journal of Cell Biology, 163(6):1205–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
163/6/1205.

Degterev:2001:CD

[DBY01] Alexei Degterev, Michael Boyce, and Junying Yuan. The channel of death. Journal of Cell Biology, 155(5):695– ??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/5/695.

Denis:2002:ICR

[DC02a] Valérie Denis and Martha S. Cyert. Internal Ca^{2+} release in yeast is triggered by hypertonic shock and mediated by a TRP channel homologue. Journal of Cell Biology, 156(1): 29–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/1/29.

Dramsi:2002:L

[DC02b] Shaynoor Dramsi and Pascale Cossart. Listeriolysin O. Journal of Cell Biology, 156(6):943–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/6/943.

Davalos:2003:BSC

[DC03] Albert R. Davalos and Judith Campisi. Bloom syndrome cells undergo p53-dependent apoptosis and delayed assembly of BRCA1 and NBS1 repair complexes at stalled replication forks. Journal of Cell Biology, 162(7):1197–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
7/1197.

Dabovic:2002:BAL

[DCC^{+}02] Branka Dabovic, Yan Chen, Cristina Colarossi, Hiroto Obata, Laura Zambuto, Mary Ann Perle, and Daniel B. Rifkin. Bone abnormalities in latent TGF-β binding protein (Ltbp)-3–null mice indicate a role for Ltbp-3 in modulating TGF-β bioavailability. Journal of Cell Biology, 156(2): 227–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/2/227.

Dunn:2000:MCA

[DCM00] Shannon E. Dunn, Eva R. Chin, and Robin N. Michel. Matching of Calcineurin Activity to Upstream Effectors Is Critical for Skeletal Muscle Fiber Growth. Journal of Cell Biology, 151(3):663–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/3/663.

Dacquin:2004:AIB

[DDL^{+}04] Romain Dacquin, Rachel A. Davey, Catherine Laplace, Régis Levasseur, Howard A. Morris, Steven R. Goldring, Samuel Gebre-Medhin, Deborah L. Galson, Jeffrey D. Zajac, and Gérard Karsenty. Amylin inhibits bone resorption while the calcitonin receptor controls bone formation in vivo. Journal of Cell Biology, 164(4):509–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/4/509.

DeBari:2003:SMR

[DDV^{+}03] Cosimo De Bari, Francesco Dell’Accio, Frank Vandenabeele, Joris R. Vermeesch, Jean-Marc Raymackers, and Frank P. Luyten. Skeletal muscle repair by adult human mesenchymal stem cells from synovial membrane. Journal of Cell Biology, 160(6):909–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/6/909.

Dul:2001:HAP

[DDW^{+}01] Jeanne L. Dul, David P. Davis, Edward K. Williamson, Fred J. Stevens, and Yair Argon. Hsp70 and antifibrillogenic peptides promote degradation and inhibit intracellular aggregation of amyloidogenic light chains. Journal of Cell Biology, 152(4):705–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/4/705.

Drew:2001:ILD

[DE01] Mark E. Drew and Paul T. Englund. Intramitochondrial Location and Dynamics of Crithidia fasciculata Kinetoplast Minicircle Replication Intermediates. Journal of Cell Biology, 153(4):735–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/4/735.

Djian:2000:TAM

[DEG00] Philippe Djian, Karen Easley, and Howard Green. Targeted Ablation of the Murine Involucrin Gene. Journal of Cell Biology, 151(2):381–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/2/381.

Dell:2003:DPT

[Del03] Karen R. Dell. Dynactin polices two-way organelle traffic. Journal of Cell Biology, 160(3):291–??, February 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/3/
291.

denElzen:2001:CDP

[dEP01] Nicole den Elzen and Jonathon Pines. Cyclin a Is Destroyed in Prometaphase and Can Delay Chromosome Alignment and Anaphase. Journal of Cell Biology, 153(1):121–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
1/121.

Dernburg:2001:HTE

[Der01] Abby F. Dernburg. Here, There, and Everywhere. Journal of Cell Biology, 153(6):??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/6/F33.

Deugnier:2002:ECV

[DFJ^{+}02] Marie-Ange Deugnier, Marisa M. Faraldo, Bassam Janji, Patricia Rousselle, Jean Paul Thiery, and Marina A. Glukhova. EGF controls the in vivo developmental potential of a mammary epithelial cell line possessing progenitor properties. Journal of Cell Biology, 159(3):453–??, November 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/3/
453.

Du:2000:ACD

[DFYL00] Jing Du, Linyin Feng, Feng Yang, and Bai Lu. Activity- and Ca^{2+}-dependent modulation of surface expression of brainderived neurotrophic factor receptors in hippocampal neurons. Journal of Cell Biology, 150(6):1423–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
6/1423.

Du:2003:RTR

[DFZ^{+}03] Jing Du, Linyin Feng, Eugene Zaitsev, Hyun-Soo Je, Xu wen Liu, and Bai Lu. Regulation of TrkB receptor tyrosine kinase and its internalization by neuronal activity and Ca^{2+} influx. Journal of Cell Biology, 163(2):385–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/163/2/
385.

Deacon:2001:YMM

[DG01] Sean W. Deacon and Vladimir I. Gelfand. Of Yeast, Mice, and Men. Journal of Cell Biology, 152(4):??, February 2001.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/152/4/
F21.

Dellambra:2000:DPC

[DGB^{+}00] Elena Dellambra, Osvaldo Golisano, Sergio Bondanza, Emanuela Siviero, Pedro Lacal, Marta Molinari, Stefania D’Atri, and Michele De Luca. Downregulation of 14-3-3σ prevents clonal evolution and leads to immortalization of primary human keratinocytes. Journal of Cell Biology, 149 (5):1117–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/5/1117.

Dewitt:2002:CFC

[DH02] Sharon Dewitt and Maurice B. Hallett. Cytosolic free Ca^{2+} changes and calpain activation are required for β integrin– accelerated phagocytosis by human neutrophils. Journal of Cell Biology, 159(1):181–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/1/181.

Dundr:2004:VKC

[DHK^{+}04] Miroslav Dundr, Michael D. Hebert, Tatiana S. Karpova, David Stanek, Hongzi Xu, Karl B. Shpargel, U. Thomas Meier, Karla M. Neugebauer, A. Gregory Matera, and Tom Misteli. In vivo kinetics of Cajal body components. Journal of Cell Biology, 164(6):831–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/6/831.

Davis:2003:CFP

[DIR03] Michael A. Davis, Renee C. Ireton, and Albert B. Reynolds. A core function for p120-catenin in cadherin turnover. Journal of Cell Biology, 163(3):525–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/3/525.

Doyle:2003:DVR

[DJ03] Michael Doyle and Michael F. Jantsch. Distinct in vivo roles for double-stranded RNA-binding domains of the Xenopus RNA-editing enzyme ADAR1 in chromosomal targeting. Journal of Cell Biology, 161(2):309–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/2/309.

Ditchfield:2003:ABC

[DJT^{+}03] Claire Ditchfield, Victoria L. Johnson, Anthony Tighe, Rebecca Ellston, Carolyn Haworth, Trevor Johnson, Andrew Mortlock, Nicholas Keen, and Stephen S. Taylor. Aurora B couples chromosome alignment with anaphase by targeting BubR1, Mad2, and Cenp–E to kinetochores. Journal of Cell Biology, 161(2):267–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/267.

Dalbey:2004:YFM

[DK04] Ross E. Dalbey and Andreas Kuhn. YidC family members are involved in the membrane insertion, lateral integration, folding, and assembly of membrane proteins. Journal of Cell Biology, 166(6):769–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/6/769.

Dwir:2001:CAS

[DKA01] Oren Dwir, Geoffrey S. Kansas, and Ronen Alon. Cytoplasmic anchorage of L-selectin controls leukocyte capture and rolling by increasing the mechanical stability of the selectin tether. Journal of Cell Biology, 155(1):145–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
1/145.

Dunn:2004:CDR

[DKAH04] Rebecca Dunn, Deborah A. Klos, Adam S. Adler, and Linda Hicke. The C2 domain of the Rsp5 ubiquitin ligase binds membrane phosphoinositides and directs ubiquitination of endosomal cargo. Journal of Cell Biology, 165(1):135–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
165/1/135.

Deshmukh:2000:CIE

[DKJ00] Mohanish Deshmukh, Keisuke Kuida, and Eugene M. Johnson. Caspase Inhibition Extends the Commitment to Neuronal Death Beyond Cytochrome c Release to the Point of Mitochondrial Depolarization. Journal of Cell Biology, 150 (1):131–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/1/131.

DiColandrea:2000:SDE

[DKMW00] Teresa DiColandrea, Tadashi Karashima, Arto Määttä, and Fiona M. Watt. Subcellular Distribution of Envoplakin and Periplakin. Journal of Cell Biology, 151(3):573–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
3/573.

Ding:2002:MAP

[DLK^{+}02] Jianqing Ding, Jia-Jia Liu, Anthony S. Kowal, Timothy Nardine, Priyanka Bhattacharya, Arthur Lee, and Yanmin Yang. Microtubule-associated protein 1B. Journal of Cell Biology, 158(3):427–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/3/427.

Dailey:2003:NTS

[DLPB03] Lisa Dailey, Emmanuel Laplantine, Riccardo Priore, and Claudio Basilico. A network of transcriptional and signaling events is activated by FGF to induce chondrocyte growth arrest and differentiation. Journal of Cell Biology, 161(6): 1053–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/6/1053.

Dimitroff:2001:CMS

[DLR^{+}01] Charles J. Dimitroff, Jack Y. Lee, Shahin Rafii, Robert C. Fuhlbrigge, and Robert Sackstein. Cd44 Is a Major E– Selectin Ligand on Human Hematopoietic Progenitor Cells. Journal of Cell Biology, 153(6):1277–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/6/1277.

Dostie:2000:NEI

[DLS00] Josée Dostie, Flavio Lejbkowicz, and Nahum Sonenberg. Nuclear Eukaryotic Initiation Factor 4e (Eif4e) Colocalizes with Splicing Factors in Speckles. Journal of Cell Biology, 148(2): 239–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/2/239.

Deng:2002:UIU

[DLT^{+}02] Fang-Ming Deng, Feng-Xia Liang, Liyu Tu, Katheryn A. Resing, Ping Hu, Mark Supino, Chih-Chi Andrew Hu, Ge Zhou, Mingxiao Ding, Gert Kreibich, and Tung-Tien Sun. Uroplakin IIIb, a urothelial differentiation marker, dimerizes with uroplakin Ib as an early step of urothelial plaque assembly. Journal of Cell Biology, 159(4):685–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
4/685.

Dai:2000:ADM

[DLXP00] Zhengshan Dai, Xiaoyan Luo, Hongbo Xie, and H. Benjamin Peng. The actin-driven movement and formation of acetylcholine receptor clusters. Journal of Cell Biology, 150 (6):1321–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/6/1321.

Diefenbach:2002:MMI

[DLY^{+}02] Thomas J. Diefenbach, Vaughan M. Latham, Dean Yimlamai, Canwen A. Liu, Ira M. Herman, and Daniel G. Jay. Myosin 1c and myosin IIB serve opposing roles in lamellipodial dynamics of the neuronal growth cone. Journal of Cell Biology, 158(7):1207–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/7/1207.

Doms:2000:HMF

[DM00] Robert W. Doms and John P. Moore. HIV-1 Membrane Fusion. Journal of Cell Biology, 151(2):??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/2/
F9.

Dammermann:2002:ACP

[DM02] Alexander Dammermann and Andreas Merdes. Assembly of centrosomal proteins and microtubule organization depends on PCM-1. Journal of Cell Biology, 159(2):255–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
2/255.

Denning:2001:NNF

[DMA^{+}01] Daniel Denning, Brook Mykytka, Nadia P. C. Allen, Lan Huang, Al Burlingame, and Michael Rexach. The nucleoporin Nup60p functions as a Gsp1p–GTP–sensitive tether for Nup2p at the nuclear pore complex. Journal of Cell Biology, 154(5):937–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/5/937.

Dickson:2002:FBI

[DMBS02] Tracey C. Dickson, C. David Mintz, Deanna L. Benson, and Stephen R. J. Salton. Functional binding interaction identified between the axonal CAM L1 and members of the ERM family. Journal of Cell Biology, 157(7):1105–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/157/7/
1105.

DeCat:2003:PPC

[DMC^{+}03] Bart De Cat, Sin-Ya Muyldermans, Christien Coomans, Gisèle Degeest, Bernadette Vanderschueren, John Creemers, Frédéric Biemar, Bernard Peers, and Guido David. Processing by proprotein convertases is required for glypican-3 modulation of cell survival, Wnt signaling, and gastrulation movements. Journal of Cell Biology, 163(3):625–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
3/625.

Davidoff:2004:PCT

[DME^{+}04] Michail S. Davidoff, Ralf Middendorff, Grigori Enikolopov, Dieter Riethmacher, Adolf F. Holstein, and Dieter Müller. Progenitor cells of the testosterone-producing Leydig cells revealed. Journal of Cell Biology, 167(5):935–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
5/935.

DeLuca:2002:HIB

[DMH^{+}02] Jennifer G. DeLuca, Ben Moree, Jennifer M. Hickey, John V. Kilmartin, and E. D. Salmon. hNuf2 inhibition blocks stable kinetochore–microtubule attachment and induces mitotic cell death in HeLa cells. Journal of Cell Biology, 159(4): 549–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/4/549.

Dorstyn:2004:TCC

[DMLK04] Loretta Dorstyn, Kathryn Mills, Yuri Lazebnik, and Sharad Kumar. The two cytochrome c species, DC3 and DC4, are not required for caspase activation and apoptosis in Drosophila cells. Journal of Cell Biology, 167(3):405–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/167/3/
405.

deMarco:2002:MNR

[dMMBK^{+}02] María C. de Marco, Fernando Martín-Belmonte, Leonor Kremer, Juan P. Albar, Isabel Correas, Jean P. Vaerman, Mónica Marazuela, Jennifer A. Byrne, and Miguel A. Alonso. MAL2, a novel raft protein of the MAL family, is an essential component of the machinery for transcytosis in hepatoma HepG2 cells. Journal of Cell Biology, 159(1):37–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/1/
37.

Dundr:2000:DPR

[DMO00] Miroslav Dundr, Tom Misteli, and Mark O. J. Olson. The Dynamics of Postmitotic Reassembly of the Nucleolus. Journal of Cell Biology, 150(3):433–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/3/433.

Dyson:2001:SCI

[DOB^{+}01] Jennifer M. Dyson, Cindy J. O’Malley, Jelena Becanovic, Adam D. Munday, Michael C. Berndt, Imogen D. Coghill, Harshal H. Nandurkar, Lisa M. Ooms, and Christina A. Mitchell. The SH2-containing inositol polyphosphate 5phosphatase, SHIP-2, binds filamin and regulates submembraneous actin. Journal of Cell Biology, 155(6):1065–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/155/6/1065.

Draviam:2001:LHC

[DOL^{+}01] Viji Mythily Draviam, Simona Orrechia, Martin Lowe, Ruggero Pardi, and Jonathon Pines. The Localization of Human Cyclins B1 and B2 Determines Cdk1 Substrate Specificity and Neither Enzyme Requires Mek to Disassemble the Golgi Apparatus. Journal of Cell Biology, 152(5):945–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
5/945.

Dove:2001:ARF

[Dov01a] Alan W. Dove. Ankyrin recruits a fleet. Journal of Cell Biology, 155(5):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/5/690.2.

Dove:2001:ANC

[Dov01b] Alan W. Dove. Another Notch in CD44. Journal of Cell Biology, 155(5):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/5/691.3.

Dove:2001:ATG

[Dov01c] Alan W. Dove. Apical transport gets motorized. Journal of Cell Biology, 155(1):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/1/11.1.

Dove:2001:AGT

[Dov01d] Alan W. Dove. ARNO gets things moving. Journal of Cell Biology, 154(3):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/3/479.2.

Dove:2001:CN

[Dov01e] Alan W. Dove. Cataracts and NrCAM. Journal of Cell Biology, 154(1):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/1/11.1.

Dove:2001:CCA

[Dov01f] Alan W. Dove. Chk1, Chk2, is the amplifier on? Journal of Cell Biology, 154(5):??, September 2001. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/5/903.2.

Dove:2001:CL

[Dov01g] Alan W. Dove. Clumping lysosomes. Journal of Cell Biology, 154(1):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/1/11.3.

Dove:2001:CRH

[Dov01h] Alan W. Dove. Collagen reveals a hidden talent. Journal of Cell Biology, 154(5):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/5/902.1.

Dove:2001:DUI

[Dov01i] Alan W. Dove. Death by unligated integrin. Journal of Cell Biology, 155(3):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/3/323.2.

Dove:2001:DKF

[Dov01j] Alan W. Dove. Deleting keratins to find one. Journal of Cell Biology, 154(3):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/3/479.3.

Dove:2001:DM

[Dov01k] Alan W. Dove. Disgorging the MVB. Journal of Cell Biology, 155(1):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/1/11.3.

Dove:2001:FRA

[Dov01l] Alan W. Dove. Finding Rho attractive. Journal of Cell Biology, 155(7):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/7/1093.2.

Dove:2001:FT

[Dov01m] Alan W. Dove. Flight of the titins. Journal of Cell Biology, 154(5):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/5/903.1.

Dove:2001:GCG

[Dov01n] Alan W. Dove. The growth cone gets a grip. Journal of Cell Biology, 155(3):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/3/323.3.

Dove:2001:HSG

[Dov01o] Alan W. Dove. Heparan sulfate gets specific. Journal of Cell Biology, 155(5):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/5/691.1.

Dove:2001:HPP

[Dov01p] Alan W. Dove. The hidden power of platelets. Journal of Cell Biology, 154(3):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/3/479.1.

Dove:2001:HBY

[Dov01q] Alan W. Dove. How big is your organelle? Journal of Cell Biology, 155(3):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/3/323.1.

Dove:2001:HSM

[Dov01r] Alan W. Dove. How to survive a macrophage attack. Journal of Cell Biology, 154(3):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/3/478.2.

Dove:2001:HTH

[Dov01s] Alan W. Dove. How Toxoplasma hooks up. Journal of Cell Biology, 154(1):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/1/11.2.

Dove:2001:KTB

[Dov01t] Alan W. Dove. Knock twice before invading. Journal of Cell Biology, 155(1):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/1/11.2.

Dove:2001:LBC

[Dov01u] Alan W. Dove. The leg bone’s connected to the … cerebellum? Journal of Cell Biology, 155(5):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/155/5/
691.2.

Dove:2001:LLP

[Dov01v] Alan W. Dove. Long live the pore. Journal of Cell Biology, 154(1):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/1/10.1.

Dove:2001:MTC

[Dov01w] Alan W. Dove. Motoring through the checkpoint. Journal of Cell Biology, 155(7):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/7/1092.2.

Dove:2001:MCE

[Dov01x] Alan W. Dove. Muscle cells experience a nuclear buildup. Journal of Cell Biology, 155(1):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/1/10.2.

Dove:2001:PCE

[Dov01y] Alan W. Dove. A PHD in T cell evasion. Journal of Cell Biology, 155(7):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/7/1092.1.

Dove:2001:PME

[Dov01z] Alan W. Dove. Potential muscles everywhere. Journal of Cell Biology, 155(3):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/3/322.1.

Dove:2001:PCP

[Dov01-27] Alan W. Dove. Putting cell polarity on the map. Journal of Cell Biology, 154(3):??, August 2001. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/3/478.1.

Dove:2001:RLS

[Dov01-28] Alan W. Dove. Ready … set … fuse. Journal of Cell Biology, 155(5):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/5/690.1.

Dove:2001:RBR

[Dov01-29] Alan W. Dove. RhoA brings up the rear. Journal of Cell Biology, 154(1):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/1/10.3.

Dove:2001:SNM

[Dov01-30] Alan W. Dove. Saying NO to muscular dystrophy. Journal of Cell Biology, 155(1):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/1/10.1.

Dove:2001:SSF

[Dov01-31] Alan W. Dove. Self-stimulation feels different. Journal of Cell Biology, 155(7):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/7/1093.3.

Dove:2001:SS

[Dov01-32] Alan W. Dove. Shape up or ship out. Journal of Cell Biology, 155(3):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/3/322.2.

Dove:2001:SAC

[Dov01-33] Alan W. Dove. The spindle assembly checkpoint gets a complex. Journal of Cell Biology, 154(5):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/154/5/
902.2.

Dove:2001:TTQ

[Dov01-34] Alan W. Dove. Transcription in tight quarters. Journal of Cell Biology, 154(1):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/1/10.2.

Dove:2001:TBI

[Dov01-35] Alan W. Dove. Turning back an invasion. Journal of Cell Biology, 155(7):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/7/1093.1.

Dove:2001:TNW

[Dov01-36] Alan W. Dove. Two new ways to eat beans. Journal of Cell Biology, 154(5):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/5/903.3.

Dove:2002:AI

[Dov02a] Alan W. Dove. An acid invasion. Journal of Cell Biology, 156(6):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/6/939.3.

Dove:2002:AA

[Dov02b] Alan W. Dove. Activating actin. Journal of Cell Biology, 156 (4):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/4/586.2.

Dove:2002:AP

[Dov02c] Alan W. Dove. Adapting to the pit. Journal of Cell Biology, 157(2):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/2/195.2.

Dove:2002:ANA

[Dov02d] Alan W. Dove. Another name for autotaxin. Journal of Cell Biology, 158(2):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/2/192.3.

Dove:2002:AWB

[Dov02e] Alan W. Dove. Another way to break down. Journal of Cell Biology, 158(1):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/1/9.1.

Dove:2002:AGI

[Dov02f] Alan W. Dove. Arp2/ 3 gets inhibited. Journal of Cell Biology, 159(6):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/6/910.3.

Dove:2002:AD

[Dov02g] Alan W. Dove. Arresting development. Journal of Cell Biology, 157(4):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/4/545.2.

Dove:2002:BGN

[Dov02h] Alan W. Dove. BAF goes nuclear. Journal of Cell Biology, 158(3):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/3/383.2.

Dove:2002:BCC

[Dov02i] Alan W. Dove. Birth control for centrioles. Journal of Cell Biology, 158(7):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/7/1152.1.

Dove:2002:BP

[Dov02j] Alan W. Dove. Building the pore. Journal of Cell Biology, 159(2):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/2/198.3.

Dove:2002:CS

[Dov02k] Alan W. Dove. CAST into the synapse. Journal of Cell Biology, 158(3):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/3/382.1.

Dove:2002:CMD

[Dov02l] Alan W. Dove. A chaperone in motoneuron disease. Journal of Cell Biology, 159(4):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/4/537.2.

Dove:2002:CC

[Dov02m] Alan W. Dove. Chromosomal cliques. Journal of Cell Biology, 159(2):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/2/199.1.

Dove:2002:CR

[Dov02n] Alan W. Dove. CLIPs to the rescue. Journal of Cell Biology, 159(4):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/4/536.2.

Dove:2002:CTJ

[Dov02o] Alan W. Dove. Consequences of a traffic jam. Journal of Cell Biology, 157(2):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/2/194.1.

Dove:2002:CMM

[Dov02p] Alan W. Dove. Coordinating myosin with mitosis. Journal of Cell Biology, 158(1):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/1/9.2.

Dove:2002:CUS

[Dov02q] Alan W. Dove. Crumbs under the skin. Journal of Cell Biology, 158(5):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/5/829.2.

Dove:2002:DCM

[Dov02r] Alan W. Dove. Do cells have mouths? Journal of Cell Biology, 156(4):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/4/587.1.

Dove:2002:ES

[Dov02s] Alan W. Dove. An endosomal switchboard. Journal of Cell Biology, 158(7):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/7/1153.3.

Dove:2002:TFT

[Dov02t] Alan W. Dove. ER takes a faster track. Journal of Cell Biology, 159(4):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/4/536.3.

Dove:2002:FCF

[Dov02u] Alan W. Dove. Fences catch free-ranging lipids. Journal of Cell Biology, 157(6):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/6/904.1.

Dove:2002:FHS

[Dov02v] Alan W. Dove. Finding a heat-shock absorber. Journal of Cell Biology, 158(7):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/7/1153.2.

Dove:2002:FDC

[Dov02w] Alan W. Dove. Finding degradation at the checkpoint. Journal of Cell Biology, 156(2):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/2/219.1.

Dove:2002:FDS

[Dov02x] Alan W. Dove. Finding the dark side of tau. Journal of Cell Biology, 156(6):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/6/938.2.

Dove:2002:HEB

[Dov02y] Alan W. Dove. For healthy eyes and bones. Journal of Cell Biology, 157(2):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/2/195.3.

Dove:2002:GG

[Dov02z] Alan W. Dove. Gazing into the gap. Journal of Cell Biology, 159(2):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/2/199.2.

Dove:2002:GTR

[Dov02-27] Alan W. Dove. Getting trapped in a rough neighborhood. Journal of Cell Biology, 158(3):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/3/382.2.

Dove:2002:GMC

[Dov02-28] Alan W. Dove. A glue for microtubule cargos. Journal of Cell Biology, 158(2):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/2/192.2.

Dove:2002:GGB

[Dov02-29] Alan W. Dove. A GTPase that goes both ways. Journal of Cell Biology, 156(2):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/2/219.3.

Dove:2002:HSE

[Dov02-30] Alan W. Dove. Herpes spoils excitement. Journal of Cell Biology, 158(7):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/7/1153.1.

Dove:2002:HFO

[Dov02-31] Alan W. Dove. How fish oils work. Journal of Cell Biology, 157(6):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/6/905.2.

Dove:2002:HSM

[Dov02-32] Alan W. Dove. How to start a motor. Journal of Cell Biology, 159(6):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/6/911.1.

Dove:2002:IYS

[Dov02-33] Alan W. Dove. If you stretch it, they will come. Journal of Cell Biology, 156(4):??, February 2002. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/4/587.2.

Dove:2002:ITE

[Dov02-34] Alan W. Dove. Initiate then elongate. Journal of Cell Biology, 159(2):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/2/198.2.

Dove:2002:CLG

[Dov02-35] Alan W. Dove. It’s a chaperone … it’s a gatekeeper … it’s BiP. Journal of Cell Biology, 156(2):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/2/
218.1.

Dove:2002:KWL

[Dov02-36] Alan W. Dove. Knowing when to let go. Journal of Cell Biology, 157(2):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/2/195.1.

Dove:2002:LP

[Dov02-37] Alan W. Dove. Less is pore. Journal of Cell Biology, 158 (5):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/5/828.2.

Dove:2002:MDS

[Dov02-38] Alan W. Dove. Making Dictyostelium stick. Journal of Cell Biology, 159(6):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/6/910.2.

Dove:2002:MSS

[Dov02-39] Alan W. Dove. Making straps from small movements. Journal of Cell Biology, 157(6):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/6/904.2.

Dove:2002:MBR

[Dov02-40] Alan W. Dove. Meanwhile, back at the ring canal …. Journal of Cell Biology, 156(4):??, February 2002. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/4/587.3.

Dove:2002:MN

[Dov02-41] Alan W. Dove. A motor for neurofilaments. Journal of Cell Biology, 159(2):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/2/198.1.

Dove:2002:MJS

[Dov02-42] Alan W. Dove. A muscular job for spectrin. Journal of Cell Biology, 157(4):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/4/545.3.

Dove:2002:NP

[Dov02-43] Alan W. Dove. The needs of the pore. Journal of Cell Biology, 158(1):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/1/8.1.

Dove:2002:NPE

[Dov02-44] Alan W. Dove. A new path for endostatin. Journal of Cell Biology, 158(3):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/3/383.3.

Dove:2002:NTO

[Dov02-45] Alan W. Dove. New tricks for old neurons. Journal of Cell Biology, 158(2):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/2/192.1.

Dove:2002:NPG

[Dov02-46] Alan W. Dove. Nuclear proteins get around. Journal of Cell Biology, 157(4):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/4/544.2.

Dove:2002:OOWb

[Dov02-47] Alan W. Dove. Only one way to organize a movement? Journal of Cell Biology, 158(1):??, July 2002. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/1/8.2.

Dove:2002:OOWa

[Dov02-48] Alan W. Dove. Only one way to skin a mouse. Journal of Cell Biology, 156(6):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/6/939.1.

Dove:2002:PYH

[Dov02-49] Alan W. Dove. A prion in yeast Huntington’s disease. Journal of Cell Biology, 157(6):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/6/905.3.

Dove:2002:PE

[Dov02-50] Alan W. Dove. Protecting the ends. Journal of Cell Biology, 158(5):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/5/828.1.

Dove:2002:SSS

[Dov02-51] Alan W. Dove. Selecting synapse start sites. Journal of Cell Biology, 159(4):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/4/537.1.

Dove:2002:SLG

[Dov02-52] Alan W. Dove. Shining a light on growth cones. Journal of Cell Biology, 159(6):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/6/910.1.

Dove:2002:SCC

[Dov02-53] Alan W. Dove. Signaling clusters caught in the act. Journal of Cell Biology, 158(7):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/7/1152.2.

Dove:2002:SSP

[Dov02-54] Alan W. Dove. Signaling from strange places. Journal of Cell Biology, 158(2):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/2/193.1.

Dove:2002:SS

[Dov02-55] Alan W. Dove. Sorting out secretion. Journal of Cell Biology, 156(2):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/2/218.2.

Dove:2002:SOF

[Dov02-56] Alan W. Dove. The sound of one foot walking. Journal of Cell Biology, 156(4):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/4/586.1.

Dove:2002:SAL

[Dov02-57] Alan W. Dove. Suicide attempt linked to breakdown. Journal of Cell Biology, 159(4):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/4/536.1.

Dove:2002:SGM

[Dov02-58] Alan W. Dove. Survival gets more complicated. Journal of Cell Biology, 156(2):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/2/219.2.

Dove:2002:TGA

[Dov02-59] Alan W. Dove. Topoisomerase gets around. Journal of Cell Biology, 158(1):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/1/9.3.

Dove:2002:TTS

[Dov02-60] Alan W. Dove. Transcription in tight spaces. Journal of Cell Biology, 157(4):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/4/545.1.

Dove:2002:TKA

[Dov02-61] Alan W. Dove. Tropomyosin keeps actin tough. Journal of Cell Biology, 156(6):??, March 2002. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/6/939.2.

Dove:2002:TM

[Dov02-62] Alan W. Dove. Tubulin but not microtubule. Journal of Cell Biology, 156(6):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/6/938.1.

Dove:2002:TOA

[Dov02-63] Alan W. Dove. Turn-ons for axons. Journal of Cell Biology, 157(4):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/4/544.1.

Dove:2002:TCO

[Dov02-64] Alan W. Dove. Two channels, one program. Journal of Cell Biology, 158(3):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/3/383.1.

Dove:2002:UKS

[Dov02-65] Alan W. Dove. Unattached kinetochore seeks short-term relationship. Journal of Cell Biology, 158(5):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
5/829.3.

Dove:2002:UAF

[Dov02-66] Alan W. Dove. Using acid to find direction. Journal of Cell Biology, 159(6):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/6/911.2.

Dove:2002:UGN

[Dov02-67] Alan W. Dove. Utrophin gets a new look. Journal of Cell Biology, 157(2):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/2/194.2.

Dove:2002:VMM

[Dov02-68] Alan W. Dove. A vitamin makes minerals. Journal of Cell Biology, 157(6):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/6/905.1.

Dove:2002:WCW

[Dov02-69] Alan W. Dove. Watching Ca^{2+} waves break. Journal of Cell Biology, 158(2):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/2/193.2.

Dove:2002:ZDS

[Dov02-70] Alan W. Dove. A zipper designed to snag. Journal of Cell Biology, 158(5):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/5/829.1.

Dove:2003:ACS

[Dov03a] Alan W. Dove. Agrin at CNS synapses. Journal of Cell Biology, 161(5):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/5/833.1.

Dove:2003:ATT

[Dov03b] Alan W. Dove. Anthrax toxin takes charge. Journal of Cell Biology, 160(3):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/3/285.2.

Dove:2003:AGD

[Dov03c] Alan W. Dove. AP-2 gets demoted. Journal of Cell Biology, 162(5):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/5/749.3.

Dove:2003:BWB

[Dov03d] Alan W. Dove. Bcl-w buries its tail, lets cell die. Journal of Cell Biology, 162(5):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/5/749.2.

Dove:2003:BB

[Dov03e] Alan W. Dove. Breaking down barriers. Journal of Cell Biology, 161(3):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/3/455.1.

Dove:2003:CC

[Dov03f] Alan W. Dove. Cajal compartments. Journal of Cell Biology, 160(4):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/4/463.2.

Dove:2003:CKA

[Dov03g] Alan W. Dove. Catching a kinase in the act. Journal of Cell Biology, 161(5):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/5/833.2.

Dove:2003:CNB

[Dov03h] Alan W. Dove. Cells need backward motor to move forward. Journal of Cell Biology, 163(6):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/6/1185.1.

Dove:2003:CLC

[Dov03i] Alan W. Dove. A close look at cerebellar learning. Journal of Cell Biology, 163(2):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/2/199.2.

Dove:2003:CJ

[Dov03j] Alan W. Dove. Conjunctions of junctions. Journal of Cell Biology, 161(5):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/5/832.1.

Dove:2003:CFS

[Dov03k] Alan W. Dove. Controlling flagellar strokes. Journal of Cell Biology, 162(1):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/1/11.1.

Dove:2003:CSM

[Dov03l] Alan W. Dove. Coping with the stress of morphogenesis. Journal of Cell Biology, 161(4):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/4/667.3.

Dove:2003:CPL

[Dov03m] Alan W. Dove. Cup puts a lid and a handle on mRNA. Journal of Cell Biology, 163(6):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/6/1184.2.

Dove:2003:CPB

[Dov03n] Alan W. Dove. Cut protein to boost fibers. Journal of Cell Biology, 161(3):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/3/454.2.

Dove:2003:DP

[Dov03o] Alan W. Dove. Dance of the podosomes. Journal of Cell Biology, 161(4):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/4/667.1.

Dove:2003:DSN

[Dov03p] Alan W. Dove. Death and the shrinking nucleus. Journal of Cell Biology, 163(6):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/6/1185.2.

Dove:2003:EWS

[Dov03q] Alan W. Dove. An early warning system for DNA damage. Journal of Cell Biology, 163(2):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/2/199.1.

Dove:2003:FCO

[Dov03r] Alan W. Dove. Four channels, but only one program. Journal of Cell Biology, 160(4):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/4/463.3.

Dove:2003:FFF

[Dov03s] Alan W. Dove. Fusion failure for flies. Journal of Cell Biology, 160(6):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/6/802.2.

Dove:2003:FIE

[Dov03t] Alan W. Dove. A fusion inhibitor on endosomes. Journal of Cell Biology, 162(1):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/1/10.2.

Dove:2003:FPF

[Dov03u] Alan W. Dove. Fusion promoter or fusion inhibitor? Journal of Cell Biology, 161(5):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/5/833.3.

Dove:2003:GHC

[Dov03v] Alan W. Dove. Getting a handle on centrosomes. Journal of Cell Biology, 161(3):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/3/455.2.

Dove:2003:GLB

[Dov03w] Alan W. Dove. Getting lamin B1 processed and organized. Journal of Cell Biology, 162(7):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/7/1178.1.

Dove:2003:GSA

[Dov03x] Alan W. Dove. Getting specific about polarized sorting. Journal of Cell Biology, 163(2):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/2/198.2.

Dove:2003:GAS

[Dov03y] Alan W. Dove. Glycosylation’s active site — finally. Journal of Cell Biology, 161(4):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/4/667.2.

Dove:2003:HC

[Dov03z] Alan W. Dove. Hierarchical condensation. Journal of Cell Biology, 162(1):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/1/11.3.

Dove:2003:HHC

[Dov03-27] Alan W. Dove. Histones help clones forget their pasts. Journal of Cell Biology, 162(1):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/1/10.1.

Dove:2003:HBB

[Dov03-28] Alan W. Dove. How to block botox. Journal of Cell Biology, 162(7):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/7/1178.2.

Dove:2003:IGT

[Dov03-29] Alan W. Dove. An interaction that gets things started. Journal of Cell Biology, 160(2):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/2/153.3.

Dove:2003:LB

[Dov03-30] Alan W. Dove. Laminin-2 on the brain. Journal of Cell Biology, 163(2):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/2/199.3.

Dove:2003:LFA

[Dov03-31] Alan W. Dove. Linking formins to Arp2/3. Journal of Cell Biology, 162(5):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/5/748.2.

Dove:2003:LVS

[Dov03-32] Alan W. Dove. A long version of Short stop. Journal of Cell Biology, 162(7):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/7/1179.2.

Dove:2003:MMM

[Dov03-33] Alan W. Dove. Math models morphogenesis and mitosis. Journal of Cell Biology, 163(6):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/6/1184.1.

Dove:2003:MM

[Dov03-34] Alan W. Dove. A model of missegregation. Journal of Cell Biology, 160(3):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/3/284.2.

Dove:2003:MAR

[Dov03-35] Alan W. Dove. Modulating AMPA receptors makes memories special. Journal of Cell Biology, 161(4):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/4/
666.2.

Dove:2003:MF

[Dov03-36] Alan W. Dove. Multimembrane fusion. Journal of Cell Biology, 160(3):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/3/284.3.

Dove:2003:NPF

[Dov03-37] Alan W. Dove. Neural progenitors find VEGF attractive. Journal of Cell Biology, 163(6):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/6/1185.3.

Dove:2003:NPP

[Dov03-38] Alan W. Dove. Neural progenitors play favorites. Journal of Cell Biology, 162(3):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/3/361.1.

Dove:2003:NF

[Dov03-39] Alan W. Dove. The neutrophil as firefighter. Journal of Cell Biology, 161(3):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/3/455.3.

Dove:2003:NTS

[Dov03-40] Alan W. Dove. NgCAM takes a scenic route to the axon. Journal of Cell Biology, 162(7):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/7/1179.1.

Dove:2003:PGA

[Dov03-41] Alan W. Dove. Pop goes the acrosome. Journal of Cell Biology, 162(7):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/7/1179.3.

Dove:2003:PNP

[Dov03-42] Alan W. Dove. A prolific nucleator program. Journal of Cell Biology, 162(5):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/5/749.1.

Dove:2003:RKM

[Dov03-43] Alan W. Dove. Rejection and killing, without MPR. Journal of Cell Biology, 160(2):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/2/153.2.

Dove:2003:SCDb

[Dov03-44] Alan W. Dove. Schwann cells don caps, touch nodes. Journal of Cell Biology, 162(3):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/3/360.2.

Dove:2003:SWG

[Dov03-45] Alan W. Dove. A simple way to get cleavage. Journal of Cell Biology, 162(3):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/3/361.2.

Dove:2003:SFF

[Dov03-46] Alan W. Dove. Splicing fibronectin’s functions. Journal of Cell Biology, 162(1):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/1/11.2.

Dove:2003:SCB

[Dov03-47] Alan W. Dove. Stem cells to build muscles. Journal of Cell Biology, 160(6):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/6/803.2.

Dove:2003:SCP

[Dov03-48] Alan W. Dove. Sticky cells in a prickly system. Journal of Cell Biology, 160(4):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/4/462.1.

Dove:2003:SCDa

[Dov03-49] Alan W. Dove. Stressed cells dance on a bright stage. Journal of Cell Biology, 161(3):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/3/454.1.

Dove:2003:TCU

[Dov03-50] Alan W. Dove. Taking close-ups of a rafting trip. Journal of Cell Biology, 160(2):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/2/153.1.

Dove:2003:YNC

[Dov03-51] Alan W. Dove. This is your neuron on cannabinoids. Journal of Cell Biology, 160(4):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/4/463.1.

Dove:2003:TAG

[Dov03-52] Alan W. Dove. Three AMIGOs grow a brain. Journal of Cell Biology, 160(6):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/6/803.1.

Dove:2003:TBB

[Dov03-53] Alan W. Dove. TRAP builds a better translocon. Journal of Cell Biology, 160(4):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/4/462.2.

Dove:2003:TC

[Dov03-54] Alan W. Dove. Tumor cells. Journal of Cell Biology, 161(4):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/4/666.1.

Dove:2003:TCG

[Dov03-55] Alan W. Dove. Tumor cells get primitive. Journal of Cell Biology, 160(2):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/2/152.1.

Dove:2003:TMO

[Dov03-56] Alan W. Dove. Two motors, one transmission. Journal of Cell Biology, 160(3):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/3/284.1.

Dove:2003:UME

[Dov03-57] Alan W. Dove. Urge to merge erects filopodia. Journal of Cell Biology, 160(3):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/3/285.1.

Dove:2003:VDT

[Dov03-58] Alan W. Dove. Venus dye trap catches executioners. Journal of Cell Biology, 160(2):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/2/152.2.

Dove:2003:WKS

[Dov03-59] Alan W. Dove. Watching kinetochores sweat. Journal of Cell Biology, 162(3):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/3/361.3.

Dove:2003:WUR

[Dov03-60] Alan W. Dove. Weaklings unite to reorganize cell. Journal of Cell Biology, 163(2):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/2/198.1.

Dove:2003:WDS

[Dov03-61] Alan W. Dove. When death starts with a caspase. Journal of Cell Biology, 160(6):??, March 2003. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/6/802.1.

Dove:2003:WMH

[Dov03-62] Alan W. Dove. When myosin heads go walking. Journal of Cell Biology, 162(3):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/3/360.1.

Dove:2003:WCL

[Dov03-63] Alan W. Dove. When the channels leak, the muscle is weak. Journal of Cell Biology, 160(6):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/6/802.3.

Dove:2003:WSW

[Dov03-64] Alan W. Dove. Wnt separate ways, met later. Journal of Cell Biology, 162(5):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/5/748.1.

Dove:2003:WCS

[Dov03-65] Alan W. Dove. Wrong connexin signals cataracts. Journal of Cell Biology, 161(5):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/5/832.2.

Dove:2004:AOT

[Dov04a] Alan W. Dove. Actin organizer takes pathogens for a ride. Journal of Cell Biology, 164(3):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/3/333.3.

Dove:2004:AMB

[Dov04b] Alan W. Dove. Apoptotic mitochondria blow a fuse. Journal of Cell Biology, 164(4):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/4/477.3.

Dove:2004:BJ

[Dov04c] Alan W. Dove. Boning up on JunB. Journal of Cell Biology, 164(4):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/4/477.2.

Dove:2004:CJG

[Dov04d] Alan W. Dove. Conjunction of junctions goes beyond function. Journal of Cell Biology, 164(2):??, January 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/2/
163.2.

Dove:2004:HIS

[Dov04e] Alan W. Dove. How insulin signals globally and acts locally. Journal of Cell Biology, 164(2):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/2/163.3.

Dove:2004:HMB

[Dov04f] Alan W. Dove. How meals become bones. Journal of Cell Biology, 164(4):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/4/476.1.

Dove:2004:HTS

[Dov04g] Alan W. Dove. How to terminate a shmoo. Journal of Cell Biology, 164(2):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/2/162.2.

Dove:2004:KCB

[Dov04h] Alan W. Dove. Kinase cascades are binary to the bone. Journal of Cell Biology, 164(3):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/3/333.1.

Dove:2004:KCS

[Dov04i] Alan W. Dove. Krox-20 conducts the Schwann song. Journal of Cell Biology, 164(3):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/3/332.2.

Dove:2004:MTO

[Dov04j] Alan W. Dove. More than one way to skin a phagosome. Journal of Cell Biology, 164(2):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/2/162.1.

Dove:2004:NSF

[Dov04k] Alan W. Dove. New structure found in plain sight. Journal of Cell Biology, 164(3):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/3/332.1.

Dove:2004:PMS

[Dov04l] Alan W. Dove. A primordial mover for sphingolipids. Journal of Cell Biology, 164(4):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/4/476.2.

Dove:2004:R

[Dov04m] Alan W. Dove. Retraction. Journal of Cell Biology, 167(4): 785–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/4/785.

Dove:2004:SDC

[Dov04n] Alan W. Dove. SNAREs drum up calcium. Journal of Cell Biology, 164(2):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/2/163.1.

Dove:2004:SMT

[Dov04o] Alan W. Dove. Switching mating types with one arm tied. Journal of Cell Biology, 164(3):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/3/333.2.

Dove:2004:YDN

[Dov04p] Alan W. Dove. Yeast dies for a noble cause. Journal of Cell Biology, 164(4):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/4/477.1.

Dong:2003:FDA

[DPB03] Yuqing Dong, David Pruyne, and Anthony Bretscher. Formin-dependent actin assembly is regulated by distinct modes of Rho signaling in yeast. Journal of Cell Biology, 161 (6):1081–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/6/1081.

DellAngelica:2000:G

[DPM^{+}00] Esteban C. Dell’Angelica, Rosa Puertollano, Chris Mullins, Rubén C. Aguilar, José D. Vargas, Lisa M. Hartnell, and Juan S. Bonifacino. Ggas. Journal of Cell Biology, 149 (1):81–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/1/81.

Devitt:2004:PAC

[DPO^{+}04] Andrew Devitt, Kate G. Parker, Carol Anne Ogden, Ceri Oldreive, Michael F. Clay, Lynsey A. Melville, Christopher O. Bellamy, Adam Lacy-Hulbert, Sophie C. Gangloff, Sanna M. Goyert, and Christopher D. Gregory. Persistence of apoptotic cells without autoimmune disease or inflammation in CD14^{−/}− mice. Journal of Cell Biology, 167(6): 1161–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/6/1161.

Dong:2000:OYZ

[DR00a] Hengjiang Dong and G. Shirleen Roeder. Organization of the Yeast Zip1 Protein within the Central Region of the Synaptonemal Complex. Journal of Cell Biology, 148(3): 417–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/3/417.

Dreier:2000:VFE

[DR00b] Lars Dreier and Tom A. Rapoport. In Vitro Formation of the Endoplasmic Reticulum Occurs Independently of Microtubules by a Controlled Fusion Reaction. Journal of Cell Biology, 148(5):883–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/5/883.

DeBellard:2003:DFS

[DRBF03] Maria Elena De Bellard, Yi Rao, and Marianne Bronner-Fraser. Dual function of Slit2 in repulsion and enhanced migration of trunk, but not vagal, neural crest cells. Journal of Cell Biology, 162(2):269–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/2/269.

Dorstyn:2002:RCC

[DRC^{+}02] Loretta Dorstyn, Stuart Read, Dimitrios Cakouros, Jun R. Huh, Bruce A. Hay, and Sharad Kumar. The role of cytochrome c in caspase activation in Drosophila melanogaster cells. Journal of Cell Biology, 156(6):1089–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/6/
1089.

DasGupta:2002:DC

[DRF02] Ramanuj DasGupta, Horace Rhee, and Elaine Fuchs. A developmental conundrum. Journal of Cell Biology, 158(2): 331–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/2/331.

Dong:2003:SIM

[DRG^{+}03] Min Dong, David A. Richards, Michael C. Goodnough, William H. Tepp, Eric A. Johnson, and Edwin R. Chapman. Synaptotagmins I and II mediate entry of botulinum neurotoxin B into cells. Journal of Cell Biology, 162(7):1293– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/7/1293.

Diao:2003:CCM

[DRP^{+}03] Aipo Diao, Dinah Rahman, Darryl J. C. Pappin, John Lucocq, and Martin Lowe. The coiled-coil membrane protein golgin-84 is a novel rab effector required for Golgi ribbon formation. Journal of Cell Biology, 160(2):201–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
2/201.

deHart:2002:CPY

[dSAH02] Amy K. A. deHart, Joshua D. Schnell, Damian A. Allen, and Linda Hicke. The conserved Pkh–Ypk kinase cascade is required for endocytosis in yeast. Journal of Cell Biology, 156(2):241–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/2/241.

Drees:2001:PIM

[DSB^{+}01] Becky L. Drees, Bryan Sundin, Elizabeth Brazeau, Juliane P. Caviston, Guang-Chao Chen, Wei Guo, Keith G. Kozminski, Michelle W. Lau, John J. Moskow, Amy Tong, Laura R. Schenkman, Amos McKenzie, Patrick Brennwald, Mark Longtine, Erfei Bi, Clarence Chan, Peter Novick, Charles Boone, John R. Pringle, Trisha N. Davis, Stanley Fields, and David G. Drubin. A protein interaction map for cell polarity development. Journal of Cell Biology, 154(3): 549–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/3/549.

Danen:2002:FBI

[DSB^{+}02] Erik H. J. Danen, Petra Sonneveld, Cord Brakebusch, Reinhard Fässler, and Arnoud Sonnenberg. The fibronectinbinding integrins α5β1 and αvβ3 differentially modulate RhoA–GTP loading, organization of cell matrix adhesions, and fibronectin fibrillogenesis. Journal of Cell Biology, 159 (6):1071–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/6/1071.

DAvino:2004:MSL

[DSG04] Pier Paolo D’Avino, Matthew S. Savoian, and David M. Glover. Mutations in sticky lead to defective organization of the contractile ring during cytokinesis and are enhanced by Rho and suppressed by Rac. Journal of Cell Biology, 166 (1):61–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/1/61.

Davis:2000:TIS

[DSH^{+}00] Alison J. Davis, Naresh B. Sepuri, Jason Holder, Arthur E. Johnson, and Robert E. Jensen. Two Intermembrane Space Tim Complexes Interact with Different Domains of Tim23p during Its Import into Mitochondria. Journal of Cell Biology, 150(6):1271–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/6/1271.

Desnos:2003:REM

[DSH^{+}03] Claire Desnos, Jean-Sébastien Schonn, Sébastien Huet, Viet Samuel Tran, Aziz El-Amraoui, Graça Raposo, Isabelle Fanget, Catherine Chapuis, Gaël Ménasché, Geneviève de Saint Basile, Christine Petit, Sophie Cribier, Jean-Pierre Henry, and François Darchen. Rab27A and its effector MyRIP link secretory granules to F-actin and control their motion towards release sites. Journal of Cell Biology, 163(3): 559–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/3/559.

Dwir:2003:AES

[DSM^{+}03] Oren Dwir, Ariel Solomon, Shmuel Mangan, Geoffrey S. Kansas, Ulrich S. Schwarz, and Ronen Alon. Avidity enhancement of L-selectin bonds by flow. Journal of Cell Biology, 163(3):649–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/3/649.

Delbac:2001:TGM

[DSN^{+}01] Frédéric Delbac, Astrid Sänger, Eva M. Neuhaus, Rolf Stratmann, James W. Ajioka, Catherine Toursel, Angelika Herm-Götz, Stanislas Tomavo, Thierry Soldati, and Dominique Soldati. Toxoplasma gondii myosins B/C. Journal of Cell Biology, 155(4):613–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/4/613.

Dilworth:2001:NDA

[DSP^{+}01] David J. Dilworth, Adisetyantari Suprapto, Julio C. Padovan, Brian T. Chait, Richard W. Wozniak, Michael P. Rout, and John D. Aitchison. Nup2p Dynamically Associates with the Distal Regions of the Yeast Nuclear Pore Complex. Journal of Cell Biology, 153(7):1465–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/7/1465.

Dirac-Svejstrup:2000:PVT

[DSSWW00] A. Barbara Dirac-Svejstrup, James Shorter, M. Gerard Waters, and Graham Warren. Phosphorylation of the Vesicle– Tethering Protein P115 by a Casein Kinase II-like Enzyme Is Required for Golgi Reassembly from Isolated Mitotic Fragments. Journal of Cell Biology, 150(3):475–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/3/
475.

Danen:2000:DSR

[DSSY00] Erik H. J. Danen, Petra Sonneveld, Arnoud Sonnenberg, and Kenneth M. Yamada. Dual stimulation of Ras/mitogenactivated protein kinase and Rhoa by cell adhesion to fibronectin supports growth factor-stimulated cell cycle progression. Journal of Cell Biology, 151(7):1413–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
7/1413.

DeVos:2000:TNF

[DSV^{+}00] Kurt De Vos, Fedor Severin, Franky Van Herreweghe, Katia Vancompernolle, Vera Goossens, Anthony Hyman, and Johan Grooten. Tumor necrosis factor induces hyperphosphorylation of kinesin light chain and inhibits kinesin-mediated transport of mitochondria. Journal of Cell Biology, 149(6): 1207–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/6/1207.

Deacon:2003:DRB

[DSV^{+}03] Sean W. Deacon, Anna S. Serpinskaya, Patricia S. Vaughan, Monica Lopez Fanarraga, Isabelle Vernos, Kevin T. Vaughan, and Vladimir I. Gelfand. Dynactin is required for bidirectional organelle transport. Journal of Cell Biology, 160(3): 297–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/3/297.

DeRosier:2000:FAB

[DT00] David J. DeRosier and Lewis G. Tilney. F-actin Bundles Are Derivatives of Microvilli. Journal of Cell Biology, 148 (1):1–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/1/1.

Donaldson:2001:MAC

[DTO^{+}01] Mary M. Donaldson, Álvaro A. M. Tavares, Hiroyuki Ohkura, Peter Deak, and David M. Glover. Metaphase Arrest with Centromere Separation in polo Mutants of Drosophila. Journal of Cell Biology, 153(4):663–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/4/663.

Dell:2004:TSI

[DV04] Karen R. Dell and Ronald D. Vale. A tribute to Shinya Inoue and innovation in light microscopy. Journal of Cell Biology, 165(1):21–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/1/21.

Desai:2000:CDS

[DVE^{+}00] Radhika C. Desai, Bimal Vyas, Cynthia A. Earles, J. Troy Littleton, Judith A. Kowalchyck, Thomas F. J. Martin, and E. R. Chapman. The C2b Domain of Synaptotagmin Is a Ca^{2+} –Sensing Module Essential for Exocytosis. Journal of Cell Biology, 150(5):1125–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/5/1125.

deVirgilio:2004:PSD

[dVKS04] Maddalena de Virgilio, William B. Kiosses, and Sanford J. Shattil. Proximal, selective, and dynamic interactions between integrin αIIbβ3 and protein tyrosine kinases in living cells. Journal of Cell Biology, 165(3):305–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/165/3/
305.

Drummond:2002:ICT

[DW02] Sheona P. Drummond and Katherine L. Wilson. Interference with the cytoplasmic tail of gp210 disrupts “close apposition” of nuclear membranes and blocks nuclear pore dilation. Journal of Cell Biology, 158(1):53–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/1/53.

Debnath:2003:AAD

[DWB03] Jayanta Debnath, Stephanie J. Walker, and Joan S. Brugge. Akt activation disrupts mammary acinar architecture and enhances proliferation in an mTOR-dependent manner. Journal of Cell Biology, 163(2):315–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/2/315.

Dubreuil:2000:DSF

[DWD^{+}00] Ronald R. Dubreuil, Ping Wang, Steve Dahl, John Lee, and Lawrence S. B. Goldstein. Drosophila β Spectrin Functions Independently of α Spectrin to Polarize the Na,k Atpase in Epithelial Cells. Journal of Cell Biology, 149(3):647–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
3/647.

Dupuis-Williams:2002:FRT

[DWFAdL^{+}02] Pascale Dupuis-Williams, Anne Fleury-Aubusson, Nicole Garreau de Loubresse, Hélène Geoffroy, Laurence Vayssié, Angélique Galvani, Aude Espigat, and Jean Rossier. Functional role of ϵ-tubulin in the assembly of the centriolar microtubule scaffold. Journal of Cell Biology, 158(7):1183– ??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/7/1183.

Dubreuil:2003:RAP

[DWM03] Catherine I. Dubreuil, Matthew J. Winton, and Lisa McKerracher. Rho activation patterns after spinal cord injury and the role of activated Rho in apoptosis in the central nervous system. Journal of Cell Biology, 162(2):233–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
2/233.

DeFea:2000:ADE

[DZT^{+}00] K. A. DeFea, J. Zalevsky, M. S. Thoma, O. Déry, R. D. Mullins, and N. W. Bunnett. β-arrestin–dependent Endocytosis of Proteinase–Activated Receptor 2 Is Required for Intracellular Targeting of Activated Erk1/2. Journal of Cell Biology, 148(6):1267–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/6/1267.

Eils:2003:CIC

[EA03] Roland Eils and Chaitanya Athale. Computational imaging in cell biology. Journal of Cell Biology, 161(3):477–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
3/477.

Edgar:2002:SCB

[EAD^{+}02] J. M. Edgar, T. J. Anderson, P. J. Dickinson, J. A. Barrie, M. C. McCulloch, K.-A. Nave, and I. R. Griffiths. Survival of, and competition between, oligodendrocytes expressing different alleles of the Plp gene. Journal of Cell Biology, 158(4): 719–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/4/719.

Earles:2001:TCD

[EBWC01] Cynthia A. Earles, Jihong Bai, Ping Wang, and Edwin R. Chapman. The tandem C2 domains of synaptotagmin contain redundant Ca^{2} + binding sites that cooperate to engage t–SNAREs and trigger exocytosis. Journal of Cell Biology, 154(6):1117–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/6/1117.

Earnshaw:2003:PFN

[EC03] William C. Earnshaw and Mar Carmena. A perfect funeral with no corpse. Journal of Cell Biology, 160(7): 989–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/7/989.

Evans:2003:MPA

[ECK^{+}03] James G. Evans, Ivan Correia, Olga Krasavina, Nicki Watson, and Paul Matsudaira. Macrophage podosomes assemble at the leading lamella by growth and fragmentation. Journal of Cell Biology, 161(4):697–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/4/697.

Ezura:2000:DEL

[ECO^{+}00] Yoichi Ezura, Shukti Chakravarti, Åke Oldberg, Inna Chervoneva, and David E. Birk. Differential Expression of Lumican and Fibromodulin Regulate Collagen Fibrillogenesis in Developing Mouse Tendons. Journal of Cell Biology, 151(4): 779–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/4/779.

Engsig:2000:MMV

[ECV^{+}00] Michael T. Engsig, Qing-Jun Chen, Thiennu H. Vu, Anne-Cecilie Pedersen, Bente Therkidsen, Leif R. Lund, Kim Henriksen, Thomas Lenhard, Niels T. Foged, Zena Werb, and Jean-Marie Delaissé. Matrix Metalloproteinase 9 and Vascular Endothelial Growth Factor Are Essential for Osteoclast Recruitment into Developing Long Bones. Journal of Cell Biology, 151(4):879–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/4/879.

Egea:2001:NSI

[EES^{+}01] Joaquim Egea, Carme Espinet, Rosa M. Soler, Xavier Dolcet, Víctor J. Yuste, Mario Encinas, Montserrat Iglesias, Nativitat Rocamora, and Joan X. Comella. Neuronal survival induced by neurotrophins requires calmodulin. Journal of Cell Biology, 154(3):585–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/3/585.

Enciso:2003:EGI

[EGC^{+}03] Josephine M. Enciso, Dita Gratzinger, Todd D. Camenisch, Sandra Canosa, Emese Pinter, and Joseph A. Madri. Elevated glucose inhibits VEGF–A-mediated endocardial cushion formation. Journal of Cell Biology, 160(4):605–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
160/4/605.

Engler:2004:MDO

[EGS^{+}04] Adam J. Engler, Maureen A. Griffin, Shamik Sen, Carsten G. Bönnemann, H. Lee Sweeney, and Dennis E. Discher. Myotubes differentiate optimally on substrates with tissue-like stiffness. Journal of Cell Biology, 166(6):877–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
6/877.

Engqvist-Goldstein:2001:ABP

[EGWK^{+}01] Åsa E. Y. Engqvist-Goldstein, Robin A. Warren, Michael M. Kessels, James H. Keen, John Heuser, and David G. Drubin. The actin-binding protein Hip1R associates with clathrin during early stages of endocytosis and promotes clathrin assembly in vitro. Journal of Cell Biology, 154(6):1209– ??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/6/1209.

El-Husseini:2000:DPP

[EHCC^{+}00] Alaa E. El-Husseini, Sarah E. Craven, Dane M. Chetkovich, Bonnie L. Firestein, Eric Schnell, Chiye Aoki, and David S. Bredt. Dual Palmitoylation of Psd-95 Mediates Its Vesiculotubular Sorting, Postsynaptic Targeting, and Ion Channel Clustering. Journal of Cell Biology, 148(1):159–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
1/159.

Ethell:2000:SNS

[EHM^{+}00] Iryna M. Ethell, Kazuki Hagihara, Yoshiaki Miura, Fumitoshi Irie, and Yu Yamaguchi. Synbindin, a Novel Syndecan-2–Binding Protein in Neuronal Dendritic Spines. Journal of Cell Biology, 151(1):53–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/1/53.

Ehler:2001:AID

[EHZ^{+}01] Elisabeth Ehler, Robert Horowits, Christian Zuppinger, Robert L. Price, Evelyne Perriard, Martin Leu, Pico Caroni, Mark Sussman, Hans M. Eppenberger, and Jean-Claude Perriard. Alterations at the Intercalated Disk Associated with the Absence of Muscle Lim Protein. Journal of Cell Biology, 153(4):763–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/4/763.

Eisner:2003:LCN

[EKB^{+}03] Gottfried Eisner, Hans-Georg Koch, Konstanze Beck, Joseph Brunner, and Matthias Müller. Ligand crowding at a nascent signal sequence. Journal of Cell Biology, 163(1):35–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
163/1/35.

Estrada:2003:MSR

[EKC^{+}03] Paula Estrada, Jiwon Kim, Jeff Coleman, Lee Walker, Brian Dunn, Peter Takizawa, Peter Novick, and Susan Ferro-Novick. Myo4p and She3p are required for cortical ER inheritance in Saccharomyces cerevisiae. Journal of Cell Biology, 163(6):1255–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/6/1255.

Espejel:2004:ITA

[EKdM^{+}04] Silvia Espejel, Peter Klatt, Josiane Ménissier de Murcia, Juan Martín-Caballero, Juana M. Flores, Guillermo Taccioli, Gilbert de Murcia, and María A. Blasco. Impact of telomerase ablation on organismal viability, aging, and tumorigenesis in mice lacking the DNA repair proteins PARP-1, Ku86, or DNA–PKcs. Journal of Cell Biology, 167(4): 627–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/4/627.

Ehehalt:2003:APA

[EKH^{+}03] Robert Ehehalt, Patrick Keller, Christian Haass, Christoph Thiele, and Kai Simons. Amyloidogenic processing of the Alzheimer β-amyloid precursor protein depends on lipid rafts. Journal of Cell Biology, 160(1):113–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/1/
113.

Evangelista:2000:RMA

[EKT^{+}00] Marie Evangelista, Bert M. Klebl, Amy H. Y. Tong, Bradley A. Webb, Thomas Leeuw, Ekkehard Leberer, Malcolm Whiteway, David Y. Thomas, and Charles Boone. A role for myosin-I in actin assembly through interactions with Vrp1p, Bee1p, and the Arp2/3 complex. Journal of Cell Biology, 148(2):353–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/2/353.

Engelholm:2003:UEE

[ELNA^{+}03] Lars H. Engelholm, Karin List, Sarah Netzel-Arnett, Edna Cukierman, David J. Mitola, Hannah Aaronson, Lars Kjøller, Jørgen K. Larsen, Kenneth M. Yamada, Dudley K. Strickland, Kenn Holmbeck, Keld Danø, Henning Birkedal-Hansen, Niels Behrendt, and Thomas H. Bugge. uPARAP/ endo180 is essential for cellular uptake of collagen and promotes fibroblast collagen adhesion. Journal of Cell Biology, 160(7): 1009–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/7/1009.

Erb:2001:RSP

[ELO^{+}01] Laurie Erb, Jun Liu, Jonathan Ockerhausen, Qiongman Kong, Richard C. Garrad, Korey Griffin, Chris Neal, Brent Krugh, Laura I. Santiago-Pérez, Fernando A. González, Hattie D. Gresham, John T. Turner, and Gary A. Weisman. An Rgd sequence in the P2y_{2} receptor interacts with α_{V} β_{3} integrins and is required for G_{o}-mediated signal transduction. Journal of Cell Biology, 153(3):491–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/3/491.

Efimov:2000:LRN

[EM00] Vladimir P. Efimov and N. Ronald Morris. The Lis1-related Nudf protein of Aspergillus nidulans interacts with the coiledcoil domain of the Nude/Ro11 protein. Journal of Cell Biology, 150(3):681–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/3/681.

Elias:2001:DID

[EMW^{+}01] Peter M. Elias, Norihisa Matsuyoshi, Hong Wu, Chenyan Lin, Zhi Hong Wang, Barbara E. Brown, and John R. Stanley. Desmoglein Isoform Distribution Affects Stratum Corneum Structure and Function. Journal of Cell Biology, 153(2): 243–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/2/243.

Edgar:2004:OMF

[EMY^{+}04] Julia M. Edgar, Mark McLaughlin, Donald Yool, Su-Chun Zhang, Jill H. Fowler, Paul Montague, Jennifer A. Barrie, Mailis C. McCulloch, Ian D. Duncan, James Garbern, Klaus A. Nave, and Ian R. Griffiths. Oligodendroglial modulation of fast axonal transport in a mouse model of hereditary spastic paraplegia. Journal of Cell Biology, 166(1):121–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
166/1/121.

Esselens:2004:PMT

[EOB^{+}04] Cary Esselens, Viola Oorschot, Veerle Baert, Tim Raemaekers, Kurt Spittaels, Lutgarde Serneels, Hui Zheng, Paul Saftig, Bart De Strooper, Judith Klumperman, and Wim Annaert. Presenilin 1 mediates the turnover of telencephalin in hippocampal neurons via an autophagic degradative pathway. Journal of Cell Biology, 166(7):1041–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
7/1041.

Eijpe:2003:MCR

[EOJ^{+}03] Maureen Eijpe, Hildo Offenberg, Rolf Jessberger, Ekaterina Revenkova, and Christa Heyting. Meiotic cohesin REC8 marks the axial elements of rat synaptonemal complexes before cohesins SMC1β and SMC3. Journal of Cell Biology, 160 (5):657–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/5/657.

Eliceiri:2002:SMC

[EPH^{+}02] Brian P. Eliceiri, Xose S. Puente, John D. Hood, Dwayne G. Stupack, David D. Schlaepfer, Xiaozhu Z. Huang, Dean Sheppard, and David A. Cheresh. Src-mediated coupling of focal adhesion kinase to integrin α v β 5 in vascular endothelial growth factor signaling. Journal of Cell Biology, 157 (1):149–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/1/149.

Evans:2003:TAI

[EPH^{+}03] Richard D. Evans, Vivienne C. Perkins, Alistair Henry, Paul E. Stephens, Martyn K. Robinson, and Fiona M. Watt. A tumor-associated β1 integrin mutation that abrogates epithelial differentiation control. Journal of Cell Biology, 160 (4):589–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/4/589.

Erbay:2003:IIT

[EPN^{+}03] Ebru Erbay, In-Hyun Park, Paul D. Nuzzi, Christopher J. Schoenherr, and Jie Chen. IGF-II transcription in skeletal myogenesis is controlled by mTOR and nutrients. Journal of Cell Biology, 163(5):931–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/5/931.

Erickson:2000:DF

[Eri00] Harold P. Erickson. Dynamin and Ftsz. Journal of Cell Biology, 148(6):1103–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/6/1103.

Ekholm-Reed:2004:DCH

[ERMT^{+}04] Susanna Ekholm-Reed, Juan Méndez, Donato Tedesco, Anders Zetterberg, Bruce Stillman, and Steven I. Reed. Deregulation of cyclin E in human cells interferes with prereplication complex assembly. Journal of Cell Biology, 165(6):789–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
165/6/789.

Ernfors:2000:NFR

[Ern00] Patrik Ernfors. Nuclear Factor-κ b to the Rescue of Cytokine–Induced Neuronal Survival. Journal of Cell Biology, 148(2):223–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/2/223.

Ekert:2004:ACA

[ERS^{+}04] Paul G. Ekert, Stuart H. Read, John Silke, Vanessa S. Marsden, Hitto Kaufmann, Christine J. Hawkins, Robert Gerl, Sharad Kumar, and David L. Vaux. Apaf-1 and caspase-9 accelerate apoptosis, but do not determine whether factordeprived or drug-treated cells die. Journal of Cell Biology, 165(6):835–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/6/835.

Edens:2001:TCL

[ESC^{+}01] William A. Edens, Lisa Sharling, Guangjie Cheng, Raymond Shapira, Joseph M. Kinkade, Taehoon Lee, Heather A. Edens, Xuexin Tang, Cameron Sullards, Denise B. Flaherty, Guy M. Benian, and J. David Lambeth. Tyrosine crosslinking of extracellular matrix is catalyzed by Duox, a multidomain oxidase/peroxidase with homology to the phagocyte oxidase subunit gp91 phox. Journal of Cell Biology, 154 (4):879–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/4/879.

Ekert:2001:DPA

[ESH^{+}01] Paul G. Ekert, John Silke, Christine J. Hawkins, Anne M. Verhagen, and David L. Vaux. Diablo promotes apoptosis by removing Miha/Xiap from processed caspase 9. Journal of Cell Biology, 152(3):483–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/3/483.

Eger:2000:EMT

[ESS^{+}00] Andreas Eger, Andreas Stockinger, Birgit Schaffhauser, Hartmut Beug, and Roland Foisner. Epithelial mesenchymal transition by c-Fos estrogen receptor activation involves nuclear translocation of β-catenin and upregulation of βcatenin/lymphoid enhancer binding factor-1 transcriptional activity. Journal of Cell Biology, 148(1):173–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
1/173.

Emoto:2000:ERM

[EU00] Kazuo Emoto and Masato Umeda. An Essential Role for a Membrane Lipid in Cytokinesis. Journal of Cell Biology, 149 (6):1215–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/6/1215.

Eickholt:2002:IPG

[EWD02] Britta J. Eickholt, Frank S. Walsh, and Patrick Doherty. An inactive pool of GSK-3 at the leading edge of growth cones is implicated in Semaphorin 3A signaling. Journal of Cell Biology, 157(2):211–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/2/211.

Ellis:2004:ZMZ

[EWM^{+}04] Charissa D. Ellis, Fudi Wang, Colin W. MacDiarmid, Suzanne Clark, Thomas Lyons, and David J. Eide. Zinc and the Msc2 zinc transporter protein are required for endoplasmic reticulum function. Journal of Cell Biology, 166 (3):325–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/3/325.

Elkind:2000:RCT

[EWSN00] N. Barry Elkind, Christiane Walch-Solimena, and Peter J. Novick. The role of the COOH terminus of Sec2p in the transport of post-Golgi vesicles. Journal of Cell Biology, 149 (1):95–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/1/95.

Eitzen:2002:ROB

[EWTW02] Gary Eitzen, Li Wang, Naomi Thorngren, and William Wickner. Remodeling of organelle-bound actin is required for yeast vacuole fusion. Journal of Cell Biology, 158(4):669–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
158/4/669.

Freeman:2000:CCG

[FAAS00] Lita Freeman, Luis Aragon-Alcaide, and Alexander Strunnikov. The Condensin Complex Governs Chromosome Condensation and Mitotic Transmission of Rdna. Journal of Cell Biology, 149(4):811–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/4/811.

Fujita:2004:RIO

[FAF^{+}04] Takashi Fujita, Yasutaka Azuma, Ryo Fukuyama, Yuji Hattori, Carolina Yoshida, Masao Koida, Kiyokazu Ogita, and Toshihisa Komori. Runx2 induces osteoblast and chondrocyte differentiation and enhances their migration by coupling with PI3K–Akt signaling. Journal of Cell Biology, 166 (1):85–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/1/85.

Facanha:2002:ERC

[FAT^{+}02] Anna L. Okorokova Façanha, Henrik Appelgren, Mohammad Tabish, Lev Okorokov, and Karl Ekwall. The endoplasmic reticulum cation P-type ATPase Cta4p is required for control of cell shape and microtubule dynamics. Journal of Cell Biology, 157(6):1029–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/6/1029.

Fratti:2001:RPK

[FBG^{+}01] Rutilio A. Fratti, Jonathan M. Backer, Jean Gruenberg, Silvia Corvera, and Vojo Deretic. Role of phosphatidylinositol 3-kinase and Rab5 effectors in phagosomal biogenesis and mycobacterial phagosome maturation arrest. Journal of Cell Biology, 154(3):631–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/3/631.

Fons:2003:SSF

[FBH03] Ryen D. Fons, Brigitte A. Bogert, and Ramanujan S. Hegde. Substrate-specific function of the translocon-associated protein complex during translocation across the ER membrane. Journal of Cell Biology, 160(4):529–??, February 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/4/
529.

Fabrizio:2004:SMA

[FBV^{+}04] Paola Fabrizio, Luisa Battistella, Raffaello Vardavas, Cristina Gattazzo, Lee-Loung Liou, Alberto Diaspro, Janis W. Dossen, Edith Butler Gralla, and Valter D. Longo. Superoxide is a mediator of an altruistic aging program in Saccharomyces cerevisiae. Journal of Cell Biology, 166(7):1055– ??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/7/1055.

Futter:2001:HVR

[FCBH01] C. E. Futter, L. M. Collinson, J. M. Backer, and C. R. Hopkins. Human VPS34 is required for internal vesicle formation within multivesicular endosomes. Journal of Cell Biology, 155(7):1251–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/7/1251.

Flanagan:2001:FAC

[FCF^{+}01] Lisa A. Flanagan, Janet Chou, Hervé Falet, Ralph Neujahr, John H. Hartwig, and Thomas P. Stossel. Filamin A, the Arp2/3 complex, and the morphology and function of cortical actin filaments in human melanoma cells. Journal of Cell Biology, 155(4):511–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/4/511.

Frey:2000:PKC

[FCL^{+}00] Mark R. Frey, Jennifer A. Clark, Olga Leontieva, Joshua M. Uronis, Adrian R. Black, and Jennifer D. Black. Protein Kinase C Signaling Mediates a Program of Cell Cycle Withdrawal in the Intestinal Epithelium. Journal of Cell Biology, 151(4):763–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/4/763.

Fernandez-Capetillo:2003:HRM

[FCLSN03] Oscar Fernandez-Capetillo, Bodo Liebe, Harry Scherthan, and André Nussenzweig. H2AX regulates meiotic telomere clustering. Journal of Cell Biology, 163(1):15–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
1/15.

Fortin:2001:AKT

[FCM^{+}01] Andre Fortin, Sean P. Cregan, Jason G. MacLaurin, Neena Kushwaha, Emma S. Hickman, Charlie S. Thompson, Antoine Hakim, Paul R. Albert, Francesco Cecconi, Kristian Helin, David S. Park, and Ruth S. Slack. APAF1 is a key transcriptional target for p53 in the regulation of neuronal cell death. Journal of Cell Biology, 155(2):207–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
2/207.

Fomina:2000:SCP

[FFKC00] Alla F. Fomina, Christopher M. Fanger, J. Ashot Kozak, and Michael D. Cahalan. Single Channel Properties and Regulated Expression of Ca^{2+} Release–Activated Ca^{2+} (Crac) Channels in Human T Cells. Journal of Cell Biology, 150 (6):1435–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/6/1435.

Fischer:2003:PEC

[FFSF03] Robert S. Fischer, Kimberly L. Fritz-Six, and Velia M. Fowler. Pointed-end capping by tropomodulin3 negatively regulates endothelial cell motility. Journal of Cell Biology, 161(2):371–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/371.

Furuse:2001:CZO

[FFST01] Mikio Furuse, Kyoko Furuse, Hiroyuki Sasaki, and Shoichiro Tsukita. Conversion of Zonulae Occludentes from Tight to Leaky Strand Type by Introducing Claudin-2 into Madin– Darby Canine Kidney I Cells. Journal of Cell Biology, 153 (2):263–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/2/263.

Farr:2000:IAG

[FFY^{+}00] Gist H. Farr, Denise M. Ferkey, Cynthia Yost, Sarah B. Pierce, Carole Weaver, and David Kimelman. Interaction among Gsk-3, Gbp, Axin, and APC in Xenopus Axis Specification. Journal of Cell Biology, 148(4):691–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
4/691.

Fievet:2004:PBP

[FGR^{+}04] Bruno T. Fievet, Alexis Gautreau, Christian Roy, Laurence Del Maestro, Paul Mangeat, Daniel Louvard, and Monique Arpin. Phosphoinositide binding and phosphorylation act sequentially in the activation mechanism of ezrin. Journal of Cell Biology, 164(5):653–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/5/653.

Frosst:2002:TLW

[FGS^{+}02] Phyllis Frosst, Tinglu Guan, Cecilia Subauste, Klaus Hahn, and Larry Gerace. Tpr is localized within the nuclear basket of the pore complex and has a role in nuclear protein export. Journal of Cell Biology, 156(4):617–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/156/
4/617.

Fukuda:2003:CIR

[FGSW03] Tomohiko Fukuda, Lida Guo, Xiaohua Shi, and Chuanyue Wu. CH-ILKBP regulates cell survival by facilitating the membrane translocation of protein kinase B/Akt. Journal of Cell Biology, 160(7):1001–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/7/1001.

Furuse:2002:CBT

[FHF^{+}02] Mikio Furuse, Masaki Hata, Kyoko Furuse, Yoko Yoshida, Akinori Haratake, Yoshinobu Sugitani, Tetsuo Noda, Akiharu Kubo, and Shoichiro Tsukita. Claudin-based tight junctions are crucial for the mammalian epidermal barrier. Journal of Cell Biology, 156(6):1099–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/6/1099.

Feijoo:2001:AMC

[FHJW^{+}01] Carmen Feijoo, Clare Hall-Jackson, Rong Wu, David Jenkins, Jane Leitch, David M. Gilbert, and Carl Smythe. Activation of mammalian Chk1 during DNA replication arrest. Journal of Cell Biology, 154(5):913–??, September 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/154/5/

Feil:2003:ILC

[FHL^{+}03] Robert Feil, Jana Hartmann, Chongde Luo, Wiebke Wolfsgruber, Karl Schilling, Susanne Feil, Jaroslaw J. Barski, Michael Meyer, Arthur Konnerth, Chris I. De Zeeuw, and Franz Hofmann. Impairment of LTD and cerebellar learning by Purkinje cell–specific ablation of cGMP-dependent protein kinase I. Journal of Cell Biology, 163(2):295–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
2/295.

Farina:2003:TZK

[FHM^{+}03] Kim L. Farina, Stefan Hüttelmaier, Kiran Musunuru, Robert Darnell, and Robert H. Singer. Two ZBP1 KH domains facilitate β-actin mRNA localization, granule formation, and cytoskeletal attachment. Journal of Cell Biology, 160(1): 77–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/1/77.

Friday:2000:CAR

[FHP00] Bret B. Friday, Valerie Horsley, and Grace K. Pavlath. Calcineurin Activity Is Required for the Initiation of Skeletal Muscle Differentiation. Journal of Cell Biology, 149(3): 657–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/3/657.

Fang:2004:VDR

[FJK^{+}04] Xiang Fang, Hong Ji, Si-Wan Kim, Jae-Il Park, Travis G. Vaught, Panos Z. Anastasiadis, Malgorzata Ciesiolka, and Pierre D. McCrea. Vertebrate development requires ARVCF and p120 catenins and their interplay with RhoA and Rac. Journal of Cell Biology, 165(1):87–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/1/87.

Fratti:2004:IAS

[FJM^{+}04] Rutilio A. Fratti, Youngsoo Jun, Alexey J. Merz, Nathan Margolis, and William Wickner. Interdependent assembly of specific regulatory lipids and membrane fusion proteins into the vertex ring domain of docked vacuoles. Journal of Cell Biology, 167(6):1087–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/6/1087.

Farkasovsky:2001:CNI

[FK01] Marian Farkasovsky and Hans Küntzel. Cortical Num1p Interacts with the Dynein Intermediate Chain Pac11p and Cytoplasmic Microtubules in Budding Yeast. Journal of Cell Biology, 152(2):251–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/2/251.

Flucher:2000:TTS

[FKG00] Bernhard E. Flucher, Nicole Kasielke, and Manfred Grabner. The Triad Targeting Signal of the Skeletal Muscle Calcium Channel Is Localized in the Cooh Terminus of the α_{1S} Subunit. Journal of Cell Biology, 151(2):467–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
2/467.

Fukumoto:2004:ACA

[FKH^{+}04] Satoshi Fukumoto, Takayoshi Kiba, Bradford Hall, Noriyuki Iehara, Takashi Nakamura, Glenn Longenecker, Paul H. Krebsbach, Antonio Nanci, Ashok B. Kulkarni, and Yoshihiko Yamada. Ameloblastin is a cell adhesion molecule required for maintaining the differentiation state of ameloblasts. Journal of Cell Biology, 167(5):973–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
167/5/973.

Fujimoto:2001:CTL

[FKI^{+}01] Toyoshi Fujimoto, Hiroshi Kogo, Kimiko Ishiguro, Kumi Tauchi, and Ryuji Nomura. Caveolin-2 Is Targeted to Lipid Droplets, a New “Membrane Domain” in the Cell. Journal of Cell Biology, 152(5):1079–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/5/1079.

Furstenthal:2001:CUC

[FKSJ01] Laura Furstenthal, Brett K. Kaiser, Craig Swanson, and Peter K. Jackson. Cyclin E Uses Cdc6 as a Chromatin– Associated Receptor Required for DNA Replication. Journal of Cell Biology, 152(6):1267–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/6/1267.

Faleiro:2000:CDN

[FL00] Lavina Faleiro and Yuri Lazebnik. Caspases Disrupt the Nuclear–Cytoplasmic Barrier. Journal of Cell Biology, 151 (5):951–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/5/951.

Fischer-Lougheed:2001:HMF

[FLLE^{+}01] Jacqueline Fischer-Lougheed, Jian-Hui Liu, Estelle Espinos, David Mordasini, Charles R. Bader, Dominique Belin, and Laurent Bernheim. Human Myoblast Fusion Requires Expression of Functional Inward Rectifier Kir2.1 Channels. Journal of Cell Biology, 153(4):677–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/4/677.

Foletta:2003:DSB

[FLS^{+}03] Victoria C. Foletta, Mei Ann Lim, Juliana Soosairajah, April P. Kelly, Edouard G. Stanley, Mark Shannon, Wei He, Supratik Das, Joan Massagué, and Ora Bernard. Direct signaling by the BMP type II receptor via the cytoskeletal regulator LIMK1. Journal of Cell Biology, 162(6):1089– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/6/1089.

Fu:2004:SHR

[FLS^{+}04] Ming Fu, Vincent Chi Hang Lui, Mai Har Sham, Vassilis Pachnis, and Paul Kwong Hang Tam. Sonic hedgehog regulates the proliferation, differentiation, and migration of enteric neural crest cells in gut. Journal of Cell Biology, 166 (5):673–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/5/673.

Feugate:2002:CCC

[FLWMG02] Jo Ellen Feugate, QiJing Li, Lina Wong, and Manuela Martins-Green. The cxc chemokine cCAF stimulates differentiation of fibroblasts into myofibroblasts and accelerates wound closure. Journal of Cell Biology, 156(1):161–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
156/1/161.

Frey:2000:SUR

[FLX^{+}00] Dunja Frey, Thorsten Laux, Lan Xu, Corinna Schneider, and Pico Caroni. Shared and Unique Roles of Cap23 and Gap43 in Actin Regulation, Neurite Outgrowth, and Anatomical Plasticity. Journal of Cell Biology, 149(7):1443–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
7/1443.

Frey:2001:RMI

[FM01] Mark R. Frey and A. Gregory Matera. RNA-mediated interaction of Cajal bodies and U2 snRNA genes. Journal of Cell Biology, 154(3):499–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/3/499.

Frederick:2004:YMG

[FMC^{+}04] Rebecca L. Frederick, J. Michael McCaffery, Kyle W. Cunningham, Koji Okamoto, and Janet M. Shaw. Yeast Miro GTPase, Gem1p, regulates mitochondrial morphology via a novel pathway. Journal of Cell Biology, 167(1):87–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
167/1/87.

Fache:2004:EED

[FMF^{+}04] Marie-Pierre Fache, Anissa Moussif, Fanny Fernandes, Pierre Giraud, Juan José Garrido, and Bénédicte Dargent. Endocytotic elimination and domain-selective tethering constitute a potential mechanism of protein segregation at the axonal initial segment. Journal of Cell Biology, 166(4):571–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
4/571.

Fang:2004:PFP

[FMJG04] Yi Fang, James C. Morrell, Jacob M. Jones, and Stephen J. Gould. PEX3 functions as a PEX19 docking factor in the import of class I peroxisomal membrane proteins. Journal of Cell Biology, 164(6):863–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/6/863.

Funamoto:2001:RPK

[FMMF01] Satoru Funamoto, Kristina Milan, Ruedi Meili, and Richard A. Firtel. Role of Phosphatidylinositol 3’ Kinase and a Downstream Pleckstrin Homology Domain–Containing Protein in Controlling Chemotaxis in Dictyostelium. Journal of Cell Biology, 153(4):795–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/4/795.

Fath:2000:AYR

[FMP^{+}00] Stephan Fath, Philipp Milkereit, Alexandre V. Podtelejnikov, Nicolas Bischler, Patrick Schultz, Mirko Bier, Matthias Mann, and Herbert Tschochner. Association of Yeast RNA Polymerase I with a Nucleolar Substructure Active in Rrna Synthesis and Processing. Journal of Cell Biology, 149(3): 575–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/3/575.

Fu:2003:CRC

[FNFL03] Xinrong Fu, Christine Ng, Daorong Feng, and Chun Liang. Cdc48p is required for the cell cycle commitment point at Start via degradation of the G1-CDK inhibitor Far1p. Journal of Cell Biology, 163(1):21–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/1/21.

Fournier:2000:SEE

[FNK^{+}00] Alyson E. Fournier, Fumio Nakamura, Susumu Kawamoto, Yoshio Goshima, Robert G. Kalb, and Stephen M. Strittmatter. Semaphorin3a Enhances Endocytosis at Sites of Receptor–F–Actin Colocalization during Growth Cone Collapse. Journal of Cell Biology, 149(2):411–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/149/2/

Fuchsova:2002:NDH

[FNKH02] Beata Fuchsová, Petr Novák, Jarmila Kafková, and Pavel Hozák. Nuclear DNA helicase II is recruited to IFN-α-activated transcription sites at PML nuclear bodies. Journal of Cell Biology, 158(3):463–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/3/463.

Faccio:2003:DCO

[FNZ^{+}03] Roberta Faccio, Deborah V. Novack, Alberta Zallone, F. Patrick Ross, and Steven L. Teitelbaum. Dynamic changes in the osteoclast cytoskeleton in response to growth factors and cell attachment are controlled by β3 integrin. Journal of Cell Biology, 162(3):499–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/3/499.

Franchitto:2002:BSP

[FP02] Annapaola Franchitto and Pietro Pichierri. Bloom’s syndrome protein is required for correct relocalization of RAD50/MRE11/NBS1 complex after replication fork arrest. Journal of Cell Biology, 157(1):19–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/1/19.

Folsch:2003:AAC

[FPM^{+}03] Heike Fölsch, Marc Pypaert, Sandra Maday, Laurence Pelletier, and Ira Mellman. The AP-1A and AP-1B clathrin adaptor complexes define biochemically and functionally distinct membrane domains. Journal of Cell Biology, 163(2): 351–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/2/351.

Feltri:2002:CDI

[FPP^{+}02] M. Laura Feltri, Diana Graus Porta, Stefano C. Previtali, Alessandro Nodari, Barbara Migliavacca, Arianna Cassetti, Amanda Littlewood-Evans, Louis F. Reichardt, Albee Messing, Angelo Quattrini, Ulrich Mueller, and Lawrence Wrabetz. Conditional disruption of β1 integrin in Schwann cells impedes interactions with axons. Journal of Cell Biology, 156(1):199–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/1/199.

Folsch:2001:DFA

[FPSM01] Heike Fölsch, Marc Pypaert, Peter Schu, and Ira Mellman. Distribution and Function of Ap-1 Clathrin Adaptor Complexes in Polarized Epithelial Cells. Journal of Cell Biology, 152(3):595–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/3/595.

Funato:2001:VNT

[FR01] Kouichi Funato and Howard Riezman. Vesicular and nonvesicular transport of ceramide from ER to the Golgi apparatus in yeast. Journal of Cell Biology, 155(6):949–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
155/6/949.

Freeze:2002:SS

[Fre02] Hudson H. Freeze. Sweet solution. Journal of Cell Biology, 158(4):615–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/4/615.

Fritz:2001:CMO

[FRK^{+}01] Stefan Fritz, Doron Rapaport, Elisabeth Klanner, Walter Neupert, and Benedikt Westermann. Connection of the Mitochondrial Outer and Inner Membranes by Fzo1 Is Critical for Organellar Fusion. Journal of Cell Biology, 152(4): 683–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/4/683.

Fujiwara:2002:PUH

[FRM^{+}02] Takahiro Fujiwara, Ken Ritchie, Hideji Murakoshi, Ken Jacobson, and Akihiro Kusumi. Phospholipids undergo hop diffusion in compartmentalized cell membrane. Journal of Cell Biology, 157(6):1071–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/6/1071.

Fedorov:2001:ORI

[FRO01] Yuri V. Fedorov, R. Scott Rosenthal, and Bradley B. Olwin. Oncogenic Ras–Induced Proliferation Requires Autocrine Fibroblast Growth Factor 2 Signaling in Skeletal Muscle Cells. Journal of Cell Biology, 152(6):1301–??, March 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/152/6/
1301.

Farsad:2001:GHC

[FRT^{+}01] Khashayar Farsad, Niels Ringstad, Kohji Takei, Scott R. Floyd, Kristin Rose, and Pietro De Camilli. Generation of high curvature membranes mediated by direct endophilin bilayer interactions. Journal of Cell Biology, 155(2):193– ??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/2/193.

Fu:2003:TIF

[FS03] Lianwu Fu and Elizabeth Sztul. Traffic-independent function of the Sar1p/COPII machinery in proteasomal sorting of the cystic fibrosis transmembrane conductance regulator. Journal of Cell Biology, 160(2):157–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/2/157.

Fortini:2000:SHD

[FSBH00] Mark E. Fortini, Marian P. Skupski, Mark S. Boguski, and Iswar K. Hariharan. A Survey of Human Disease Gene Counterparts in the Drosophila Genome. Journal of Cell Biology, 150(2):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/2/F23.

Fritz-Six:2003:AMA

[FSCF^{+}03] Kimberly L. Fritz-Six, Patrick R. Cox, Robert S. Fischer, Bisong Xu, Carol C. Gregorio, Huda Y. Zoghbi, and Velia M. Fowler. Aberrant myofibril assembly in tropomodulin1 null mice leads to aborted heart development and embryonic lethality. Journal of Cell Biology, 163(5):1033–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
5/1033.

Fish:2000:EDF

[FSD00] Kenneth N. Fish, Sandra L. Schmid, and Hanna Damke. Evidence That Dynamin-2 Functions as a Signal–Transducing Gtpase. Journal of Cell Biology, 150(1):145–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/1/
145.

Faivre-Sarrailh:2000:GIA

[FSGDN^{+}00] Catherine Faivre-Sarrailh, France Gauthier, Natalia Denisenko-Nehrbass, André Le Bivic, Geneviève Rougon, and Jean-Antoine Girault. The glycosylphosphatidyl inositol-anchored adhesion molecule F3/contactin is required for surface transport of paranodin/contactin-associated protein (Caspr). Journal of Cell Biology, 149(2):491–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/2/491.

Fitzli:2000:DIA

[FSK^{+}00] Dora Fitzli, Esther T. Stoeckli, Stefan Kunz, Kingsley Siribour, Christoph Rader, Beat Kunz, Serguei V. Kozlov, Andrea Buchstaller, Robert P. Lane, Daniel M. Suter, William J. Dreyer, and Peter Sonderegger. A Direct Interaction of Axonin-1 with Ngcam–Related Cell Adhesion Molecule (Nrcam) Results in Guidance, but Not Growth of Commissural Axons. Journal of Cell Biology, 149(4):951–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
149/4/951.

Fukuhara:2004:ACT

[FSK^{+}04] Tatsuro Fukuhara, Kazuya Shimizu, Tomomi Kawakatsu, Taihei Fukuyama, Yukiko Minami, Tomoyuki Honda, Takashi Hoshino, Tomohiro Yamada, Hisakazu Ogita, Masato Okada, and Yoshimi Takai. Activation of Cdc42 by trans interactions of the cell adhesion molecules nectins through c–Src and Cdc42-GEF FRG. Journal of Cell Biology, 166(3):393–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/3/393.

Fleig:2000:FYR

[FSKS00] Ursula Fleig, Sandra S. Salus, Inga Karig, and Shelley Sazer. The Fission Yeast Ran Gtpase Is Required for Microtubule Integrity. Journal of Cell Biology, 151(5):1101–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
5/1101.

Fukuoka:2001:NNW

[FSM^{+}01] Maiko Fukuoka, Shiro Suetsugu, Hiroaki Miki, Kiyoko Fukami, Takeshi Endo, and Tadaomi Takenawa. A novel neural Wiskott–Aldrich syndrome protein (N–Wasp) binding protein, wish, induces Arp2/3 complex activation independent of Cdc42. Journal of Cell Biology, 152(3):471–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
152/3/471.

Fey:2002:SNA

[FSTC02] Petra Fey, Stephen Stephens, Margaret A. Titus, and Rex L. Chisholm. SadA, a novel adhesion receptor in Dictyostelium. Journal of Cell Biology, 159(6):1109–??, December 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/6/
1109.

Fekkes:2000:GOM

[FSY00] Peter Fekkes, Kelly A. Shepard, and Michael P. Yaffe. Gag3p, an Outer Membrane Protein Required for Fission of Mitochondrial Tubules. Journal of Cell Biology, 151(2):333– ??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/2/333.

Franzen:2001:MAP

[FTD^{+}01] Rachelle Franzen, Sandra L. Tanner, Suzanne M. Dashiell, Catherine A. Rottkamp, Jeffrey A. Hammer, and Richard H. Quarles. Microtubule-associated protein 1B. Journal of Cell Biology, 155(6):893–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/6/893.

Feierbach:2004:RFC

[FVC04] Becket Feierbach, Fulvia Verde, and Fred Chang. Regulation of a formin complex by the microtubule plus end protein tea1p. Journal of Cell Biology, 165(5):697–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/165/
5/697.

Foghsgaard:2001:CBA

[FWM^{+}01] Lasse Foghsgaard, Dorte Wissing, Daniel Mauch, Ulrik Lademann, Lone Bastholm, Marianne Boes, Folmer Elling, Marcel Leist, and Marja Jäättelä. Cathepsin B Acts as a Dominant Execution Protease in Tumor Cell Apoptosis Induced by Tumor Necrosis Factor. Journal of Cell Biology, 153(5):999–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/5/999.

Fitzgerald:2000:SSE

[FWP^{+}00] Marilyn L. Fitzgerald, Zihua Wang, Pyong Woo Park, Gillian Murphy, and Merton Bernfield. Shedding of Syndecan-1 and -4 ectodomains is regulated by multiple signaling pathways and mediated by a Timp-3–sensitive metalloproteinase. Journal of Cell Biology, 148(4):811–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/4/811.

Fu:2001:RGD

[FWY01] Ying Fu, Guang Wu, and Zhenbiao Yang. ROP Gtpase– Dependent Dynamics of Tip–Localized F–Actin Controls Tip Growth in Pollen Tubes. Journal of Cell Biology, 152(5): 1019–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/5/1019.

Fletcher:2000:DCP

[FXPT00] Graham C. Fletcher, Luzheng Xue, Shareta K. Passingham, and Aviva M. Tolkovsky. Death Commitment Point Is Advanced by Axotomy in Sympathetic Neurons. Journal of Cell Biology, 150(4):741–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/4/741.

Fujii:2000:APA

[FYHH00] Ritsuko Fujii, Susumu Yamashita, Masahiko Hibi, and Toshio Hirano. Asymmetric p38 Activation in Zebrafish. Journal of Cell Biology, 150(6):1335–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/6/1335.

Fukami:2003:PCR

[FYI^{+}03] Kiyoko Fukami, Manabu Yoshida, Takafumi Inoue, Manabu Kurokawa, Rafael A. Fissore, Nobuaki Yoshida, Katsuhiko Mikoshiba, and Tadaomi Takenawa. Phospholipase Cδ4 is required for Ca^{2+} mobilization essential for acrosome reaction in sperm. Journal of Cell Biology, 161(1):79–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
1/79.

Gillespie:2001:MN

[GAB^{+}01] Peter G. Gillespie, Joseph P. Albanesi, Martin Bähler, William M. Bement, Jonathan S. Berg, David R. Burgess, Beth Burnside, Richard E. Cheney, David P. Corey, Evelyne Coudrier, Primal de Lanerolle, John A. Hammer, Tama Hasson, Jeffrey R. Holt, A. J. Hudspeth, Mitsuo Ikebe, John Kendrick-Jones, Edward D. Korn, Rong Li, John A. Mercer, Ronald A. Milligan, Mark S. Mooseker, E. Michael Ostap, Christine Petit, Thomas D. Pollard, James R. Sellers, Thierry Soldati, and Margaret A. Titus. Myosin-i nomenclature. Journal of Cell Biology, 155(5):703–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
5/703.

Grady:2003:TPN

[GAC^{+}03] R. Mark Grady, Mohammed Akaaboune, Alexander L. Cohen, Margaret M. Maimone, Jeff W. Lichtman, and Joshua R. Sanes. Tyrosine-phosphorylated and nonphosphorylated isoforms of α-dystrobrevin. Journal of Cell Biology, 160(5): 741–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/5/741.

Galan:2002:CBM

[Gal02] Jorge E. Galán. The cell biology of microbial infections. Journal of Cell Biology, 158(3):387–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/3/387.

Gao:2003:GAM

[GAT^{+}03] Xiang-Dong Gao, Stefan Albert, Serguei E. Tcheperegine, Christopher G. Burd, Dieter Gallwitz, and Erfei Bi. The GAP activity of Msb3p and Msb4p for the Rab GTPase Sec4p is required for efficient exocytosis and actin organization. Journal of Cell Biology, 162(4):635–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/162/4/
635.

Gaudin:2000:RVI

[Gau00] Yves Gaudin. Rabies Virus–Induced Membrane Fusion Pathway. Journal of Cell Biology, 150(3):601–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/3/
601.

Gerhart:2000:DDL

[GBD^{+}00] Jacquelyn Gerhart, Michael Baytion, Steven DeLuca, Robert Getts, Christian Lopez, Robert Niewenhuis, Thor Nilsen, Scott Olex, Harold Weintraub, and Mindy George-Weinstein. DNA Dendrimers Localize Myod mRNA in Presomitic Tissues of the Chick Embryo. Journal of Cell Biology, 149 (4):825–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/4/825.

Gramolini:2001:DRU

[GBJ01] Anthony O. Gramolini, Guy Bélanger, and Bernard J. Jasmin. Distinct regions in the 3’ untranslated region are responsible for targeting and stabilizing utrophin transcripts in skeletal muscle cells. Journal of Cell Biology, 154(6):1173– ??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/6/1173.

Gnessi:2000:LCL

[GBM^{+}00] Lucio Gnessi, Sabrina Basciani, Stefania Mariani, Mario Arizzi, Giovanni Spera, Chiayeng Wang, Cecilia Bondjers, Linda Karlsson, and Christer Betsholtz. Leydig Cell Loss and Spermatogenic Arrest in Platelet–Derived Growth Factor (Pdgf)–a–Deficient Mice. Journal of Cell Biology, 149 (5):1019–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/5/1019.

Guenatri:2004:MCP

[GBMA04] Mounia Guenatri, Delphine Bailly, Christèle Maison, and Geneviève Almouzni. Mouse centric and pericentric satellite repeats form distinct functional heterochromatin. Journal of Cell Biology, 166(4):493–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/4/493.

Gerhart:2001:MPM

[GBN^{+}01] Jacquelyn Gerhart, Brian Bast, Christine Neely, Stephanie Iem, Paula Amegbe, Robert Niewenhuis, Steven Miklasz, Pei Feng Cheng, and Mindy George-Weinstein. MyoDpositive myoblasts are present in mature fetal organs lacking skeletal muscle. Journal of Cell Biology, 155(3):381–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
155/3/381.

Gerbi:2003:USN

[GBOL03] Susan A. Gerbi, Anton V. Borovjagin, Federico E. Odreman, and Thilo Sascha Lange. U4 snRNA nucleolar localization requires the NHPX/15.5-kD protein binding site but not Sm protein or U6 snRNA association. Journal of Cell Biology, 162(5):821–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/5/821.

Grote:2000:GSD

[GBON00] Eric Grote, Misuzu Baba, Yoshinori Ohsumi, and Peter J. Novick. Geranylgeranylated Snares Are Dominant Inhibitors of Membrane Fusion. Journal of Cell Biology, 151(2):453– ??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/2/453.

Gu:2003:CRS

[GBY^{+}03] Haihua Gu, Roberto J. Botelho, Min Yu, Sergio Grinstein, and Benjamin G. Neel. Critical role for scaffolding adapter Gab2 in FcγR-mediated phagocytosis. Journal of Cell Biology, 161(6):1151–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/6/1151.

Gladfelter:2002:SRA

[GBZ^{+}02] Amy S. Gladfelter, Indrani Bose, Trevin R. Zyla, Elaine S. G. Bardes, and Daniel J. Lew. Septin ring assembly involves cycles of GTP loading and hydrolysis by Cdc42p. Journal of Cell Biology, 156(2):315–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/2/315.

Ganem:2004:KKK

[GC04] Neil J. Ganem and Duane A. Compton. The KinI kinesin Kif2a is required for bipolar spindle assembly through a functional relationship with MCAK. Journal of Cell Biology, 166 (4):473–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/4/473.

Gao:2001:HAT

[GCG^{+}01] Xiangming Gao, Tanya Chandra, Michel-Olivier Gratton, Isabelle Quélo, Josée Prud’homme, Stefano Stifani, and René St-Arnaud. HES6 acts as a transcriptional repressor in myoblasts and can induce the myogenic differentiation program. Journal of Cell Biology, 154(6):1161–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/154/6/
1161.

Gilley:2003:FTF

[GCH03] Jonathan Gilley, Paul J. Coffer, and Jonathan Ham. FOXO transcription factors directly activate bim gene expression and promote apoptosis in sympathetic neurons. Journal of Cell Biology, 162(4):613–??, August 2003. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/4/613.

Gassmann:2004:B

[GCH^{+}04] Reto Gassmann, Ana Carvalho, Alexander J. Henzing, Sandrine Ruchaud, Damien F. Hudson, Reiko Honda, Erich A. Nigg, Dietlind L. Gerloff, and William C. Earnshaw. Borealin. Journal of Cell Biology, 166(2):179–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/166/2/
179.

Grote:2000:OFE

[GCN00] Eric Grote, Chavela M. Carr, and Peter J. Novick. Ordering the Final Events in Yeast Exocytosis. Journal of Cell Biology, 151(2):439–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/2/439.

Gakidis:2004:VGR

[GCO^{+}04] M. Angelica Martinez Gakidis, Xavier Cullere, Timothy Olson, Julie L. Wilsbacher, Bin Zhang, Sheri L. Moores, Klaus Ley, Wojciech Swat, Tanya Mayadas, and Joan S. Brugge. Vav GEFs are required for β_{2} integrin-dependent functions of neutrophils. Journal of Cell Biology, 166(2):273–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
2/273.

Guild:2003:LCA

[GCR^{+}03] Gregory M. Guild, Patricia S. Connelly, Linda Ruggiero, Kelly A. Vranich, and Lewis G. Tilney. Long continuous actin bundles in Drosophila bristles are constructed by overlapping short filaments. Journal of Cell Biology, 162(6):1069– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/6/1069.

Geneste:2002:LKD

[GCT02] Olivier Geneste, John W. Copeland, and Richard Treisman. LIM kinase and Diaphanous cooperate to regulate serum response factor and actin dynamics. Journal of Cell Biology, 157(5):831–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/5/831.

Gourlay:2004:RAC

[GCT^{+}04] Campbell W. Gourlay, Lindsay N. Carpp, Paul Timpson, Steven J. Winder, and Kathryn R. Ayscough. A role for the actin cytoskeleton in cell death and aging in yeast. Journal of Cell Biology, 164(6):803–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/6/803.

Grosshans:2001:BSR

[GDHS01] Helge Grosshans, Karina Deinert, Ed Hurt, and George Simos. Biogenesis of the signal recognition particle (Srp) involves import of Srp proteins into the nucleolus, assembly with the Srp–RNA, and Xpo1p-mediated export. Journal of Cell Biology, 153(4):745–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/4/745.

Garin:2001:PP

[GDK^{+}01] Jérome Garin, Roberto Diez, Sylvie Kieffer, Jean-François Dermine, Sophie Duclos, Etienne Gagnon, Remy Sadoul, Christiane Rondeau, and Michel Desjardins. The Phagosome Proteome. Journal of Cell Biology, 152(1):165–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
1/165.

Gaillard:2001:FRS

[GDRS01] Anne Roush Gaillard, Dennis R. Diener, Joel L. Rosenbaum, and Winfield S. Sale. Flagellar Radial Spoke Protein 3 Is an a–Kinase Anchoring Protein (Akap). Journal of Cell Biology, 153(2):443–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/2/443.

Gillett:2004:SCP

[GES04] Emily S. Gillett, Christopher W. Espelin, and Peter K. Sorger. Spindle checkpoint proteins and chromosome– microtubule attachment in budding yeast. Journal of Cell Biology, 164(4):535–??, February 2004. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/4/535.

Genova:2003:NGK

[GF03] Jennifer L. Genova and Richard G. Fehon. Neuroglian, Gliotactin, and the Na^{+} /K^{+} ATPase are essential for septate junction function in Drosophila. Journal of Cell Biology, 161 (5):979–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/5/979.

Grevengoed:2003:BDT

[GFGP03] Elizabeth E. Grevengoed, Donald T. Fox, Julie Gates, and Mark Peifer. Balancing different types of actin polymerization at distinct sites. Journal of Cell Biology, 163(6):1267– ??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/6/1267.

Goel:2004:C

[GFM^{+}04a] Hira Lal Goel, Mara Fornaro, Loredana Moro, Natalia Teider, Johng S. Rhim, Michael King, and Lucia R. Languino. Correction. Journal of Cell Biology, 167(3):565–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
167/3/565.

Goel:2004:SMT

[GFM^{+}04b] Hira Lal Goel, Mara Fornaro, Loredana Moro, Natalia Teider, Johng S. Rhim, Michael King, and Lucia R. Languino. Selective modulation of type 1 insulin-like growth factor receptor signaling and functions by β_{1} integrins. Journal of Cell Biology, 166(3):407–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/3/407.

Goldstein:2000:FTD

[GG00] Lawrence S. B. Goldstein and Shermali Gunawardena. Flying through the Drosophila Cytoskeletal Genome. Journal of Cell Biology, 150(2):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/2/F63.

Giet:2001:DAB

[GG01] Régis Giet and David M. Glover. Drosophila Aurora B Kinase Is Required for Histone H3 Phosphorylation and Condensin Recruitment during Chromosome Condensation and to Organize the Central Spindle during Cytokinesis. Journal of Cell Biology, 152(4):669–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/4/669.

Gottardi:2004:DMF

[GG04] Cara J. Gottardi and Barry M. Gumbiner. Distinct molecular forms of β-catenin are targeted to adhesive or transcriptional complexes. Journal of Cell Biology, 167(2):339–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
2/339.

Gentle:2004:OFP

[GGB^{+}04] Ian Gentle, Kipros Gabriel, Peter Beech, Ross Waller, and Trevor Lithgow. The Omp85 family of proteins is essential for outer membrane biogenesis in mitochondria and bacteria. Journal of Cell Biology, 164(1):19–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/1/
19.

Gaskins:2004:IMR

[GGD^{+}04] Elizabeth Gaskins, Stacey Gilk, Nicolette DeVore, Tara Mann, Gary Ward, and Con Beckers. Identification of the membrane receptor of a class XIV myosin in Toxoplasma gondii. Journal of Cell Biology, 165(3):383–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/165/3/
383.

Gerhardt:2003:VGA

[GGF^{+}03] Holger Gerhardt, Matthew Golding, Marcus Fruttiger, Christiana Ruhrberg, Andrea Lundkvist, Alexandra Abramsson, Michael Jeltsch, Christopher Mitchell, Kari Alitalo, David Shima, and Christer Betsholtz. VEGF guides angiogenic sprouting utilizing endothelial tip cell filopodia. Journal of Cell Biology, 161(6):1163–??, June 2003. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/6/1163.

Ghosh:2003:MGA

[GGGK03] Pradipta Ghosh, Janice Griffith, Hans J. Geuze, and Stuart Kornfeld. Mammalian GGAs act together to sort mannose 6-phosphate receptors. Journal of Cell Biology, 163(4): 755–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/4/755.

Gruneberg:2002:CPR

[GGGN02] Ulrike Gruneberg, Michael Glotzer, Anton Gartner, and Erich A. Nigg. The CeCDC-14 phosphatase is required for cytokinesis in the Caenorhabditis elegans embryo. Journal of Cell Biology, 158(5):901–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/5/901.

Grossmann:2004:RPH

[GGH^{+}04] Katja S. Grossmann, Christine Grund, Joerg Huelsken, Martin Behrend, Bettina Erdmann, Werner W. Franke, and Walter Birchmeier. Requirement of plakophilin 2 for heart morphogenesis and cardiac junction formation. Journal of Cell Biology, 167(1):149–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/1/149.

Goncharova:2004:TMA

[GGNK04] Elena Goncharova, Dmitry Goncharov, Daniel Noonan, and Vera P. Krymskaya. TSC2 modulates actin cytoskeleton and focal adhesion through TSC1-binding domain and the Rac1 GTPase. Journal of Cell Biology, 167(6):1171–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
6/1171.

Glomski:2002:LMH

[GGT^{+}02] Ian J. Glomski, Margaret M. Gedde, Albert W. Tsang, Joel A. Swanson, and Daniel A. Portnoy. The Listeria monocytogenes hemolysin has an acidic pH optimum to compartmentalize activity and prevent damage to infected host cells.

Journal of Cell Biology, 156(6):1029–??, March 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/6/
1029.

Gill:2000:SCE

[GH00] R. Montgomery Gill and Paul A. Hamel. Subcellular Compartmentalization of E2f Family Members Is Required for Maintenance of the Postmitotic State in Terminally Differentiated Muscle. Journal of Cell Biology, 148(6):1187–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/148/6/1187.

Gordon:2001:CMM

[GHC01] Michael B. Gordon, Louisa Howard, and Duane A. Compton. Chromosome Movement in Mitosis Requires Microtubule Anchorage at Spindle Poles. Journal of Cell Biology, 152(3): 425–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/3/425.

Goldfinger:2003:SRI

[GHK^{+}03] Lawrence E. Goldfinger, Jaewon Han, William B. Kiosses, Alan K. Howe, and Mark H. Ginsberg. Spatial restriction of α4 integrin phosphorylation regulates lamellipodial stability and α4β1-dependent cell migration. Journal of Cell Biology, 162(4):731–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/4/731.

Glogauer:2000:TPT

[GHS00] M. Glogauer, J. Hartwig, and T. Stossel. Two Pathways through Cdc42 Couple the N -Formyl Receptor to Actin Nucleation in Permeabilized Human Neutrophils. Journal of Cell Biology, 150(4):785–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/4/785.

Garvalov:2003:CST

[GHS^{+}03] Boyan K. Garvalov, Theresa E. Higgins, James D. Sutherland, Markus Zettl, Niki Scaplehorn, Thomas Köcher, Eugenia Piddini, Gareth Griffiths, and Michael Way. The conformational state of Tes regulates its zyxin-dependent recruitment to focal adhesions. Journal of Cell Biology, 161 (1):33–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/1/33.

Gatfield:2002:RAA

[GI02] David Gatfield and Elisa Izaurralde. REF1/ aly and the additional exon junction complex proteins are dispensable for nuclear mRNA export. Journal of Cell Biology, 159(4): 579–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/4/579.

Gilbert:2001:NPL

[Gil01] David M. Gilbert. Nuclear Position Leaves Its Mark on Replication Timing. Journal of Cell Biology, 152(2):0–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
2/F11.

Gress:2000:GPC

[GJ00] Catherine J. Gress and Olena Jacenko. Growth Plate Compressions and Altered Hematopoiesis in Collagen X Null Mice. Journal of Cell Biology, 149(4):983–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/149/4/
983.

Gonzalez:2000:GAA

[GJB^{+}00] María Victoria González, Benilde Jiménez, María T. Berciano, José Manuel González-Sancho, Carme Caelles, Miguel Lafarga, and Alberto Muñoz. Glucocorticoids antagonize Ap-1 by inhibiting the activation/phosphorylation of Jnk without affecting its subcellular distribution. Journal of Cell Biology, 150(5):1199–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/5/1199.

Giannone:2003:TCF

[GJS^{+}03a] Grégory Giannone, Guoying Jiang, Deborah H. Sutton, David R. Critchley, and Michael P. Sheetz. Talin1 is critical for force-dependent reinforcement of initial integrin– cytoskeleton bonds but not tyrosine kinase activation. Journal of Cell Biology, 163(2):409–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/2/409.

Gromley:2003:NHP

[GJS^{+}03b] Adam Gromley, Agata Jurczyk, James Sillibourne, Ensar Halilovic, Mette Mogensen, Irina Groisman, Maureen Blomberg, and Stephen Doxsey. A novel human protein of the maternal centriole is required for the final stages of cytokinesis and entry into S phase. Journal of Cell Biology, 161 (3):535–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/3/535.

Ghosh:2003:ABS

[GK03a] Pradipta Ghosh and Stuart Kornfeld. AP-1 binding to sorting signals and release from clathrin-coated vesicles is regulated by phosphorylation. Journal of Cell Biology, 160(5): 699–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/5/699.

Green:2003:CIC

[GK03b] Rebecca A. Green and Kenneth B. Kaplan. Chromosome instability in colorectal tumor cells is associated with defects in microtubule plus-end attachments caused by a dominant mutation in APC. Journal of Cell Biology, 163(5): 949–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/5/949.

Gaetz:2004:DDR

[GK04] Jedidiah Gaetz and Tarun M. Kapoor. Dynein/ dynactin regulate metaphase spindle length by targeting depolymerizing activities to spindle poles. Journal of Cell Biology, 166 (4):465–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/4/465.

Geley:2001:APC

[GKG^{+}01] Stephan Geley, Edgar Kramer, Christian Gieffers, Julian Gannon, Jan-Michael Peters, and Tim Hunt. Anaphasepromoting complex/cyclosome-dependent proteolysis of human cyclin a starts at the beginning of mitosis and is not subject to the spindle assembly checkpoint. Journal of Cell Biology, 153(1):137–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/1/137.

Govind:2001:CFC

[GKM^{+}01] Sheila Govind, Robert Kozma, Clinton Monfries, Louis Lim, and Sohail Ahmed. Cdc42hs Facilitates Cytoskeletal Reorganization and Neurite Outgrowth by Localizing the 58-Kd Insulin Receptor Substrate to Filamentous Actin. Journal of Cell Biology, 152(3):579–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/3/579.

Guo:2003:PDY

[GKN^{+}03] Tong Guo, Yuriy Y. Kit, Jean-Marc Nicaud, Marie-Therese Le Dall, S. Kelly Sears, Hojatollah Vali, Honey Chan, Richard A. Rachubinski, and Vladimir I. Titorenko. Peroxisome division in the yeast Yarrowia lipolytica is regulated by a signal from inside the peroxisome. Journal of Cell Biology, 162(7):1255–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/7/1255.

Gayraud:2000:NIA

[GKSR00] Barbara Gayraud, Douglas R. Keene, Lynn Y. Sakai, and Francesco Ramirez. New Insights into the Assembly of Extracellular Microfibrils from the Analysis of the Fibrillin 1 Mutation in the Tight skin Mouse. Journal of Cell Biology, 150(3):667–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/3/667.

Goshima:2003:HCC

[GKYY03] Gohta Goshima, Tomomi Kiyomitsu, Kinya Yoda, and Mitsuhiro Yanagida. Human centromere chromatin protein hMis12, essential for equal segregation, is independent of CENP–A loading pathway. Journal of Cell Biology, 160(1): 25–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/1/25.

Gautreau:2000:MEE

[GLA00] Alexis Gautreau, Daniel Louvard, and Monique Arpin. Morphogenic Effects of Ezrin Require a Phosphorylation–Induced Transition from Oligomers to Monomers at the Plasma Membrane. Journal of Cell Biology, 150(1):193–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/1/
193.

Glaser:2001:ADS

[Gla01a] Vicki Glaser. Antagonistic drivers share BuMPy pathway. Journal of Cell Biology, 154(6):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/6/1100.2.

Glaser:2001:ABS

[Gla01b] Vicki Glaser. Arp bent out of shape. Journal of Cell Biology, 154(6):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/6/1101.1.

Glaser:2001:DSI

[Gla01c] Vicki Glaser. Desmosomes sort it out. Journal of Cell Biology, 154(6):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/6/1101.2.

Glaser:2001:TB

[Gla01d] Vicki Glaser. Tagged From birth. Journal of Cell Biology, 154 (6):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/6/1101.3.

Glaser:2001:TC

[Gla01e] Vicki Glaser. Tales from the crypt. Journal of Cell Biology, 154(6):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/6/1100.1.

Gilbert:2001:SEK

[GLDM01] Stéphane Gilbert, Anne Loranger, Nathalie Daigle, and Normand Marceau. Simple epithelium keratins 8 and 18 provide resistance to Fas-mediated apoptosis. The protection occurs through a receptor-targeting modulation. Journal of Cell Biology, 154(4):763–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/4/763.

Grevengoed:2001:AKR

[GLJP01] Elizabeth E. Grevengoed, Joseph J. Loureiro, Traci L. Jesse, and Mark Peifer. Abelson kinase regulates epithelial morphogenesis in Drosophila. Journal of Cell Biology, 155(7): 1185–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/7/1185.

Ghatan:2000:PMK

[GLK^{+}00] Saadi Ghatan, Stephen Larner, Yoshito Kinoshita, Michal Hetman, Leena Patel, Zhengui Xia, Richard J. Youle, and Richard S. Morrison. p38 Map Kinase Mediates Bax Translocation in Nitric Oxide–Induced Apoptosis in Neurons. Journal of Cell Biology, 150(2):335–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/2/335.

Glotzer:2004:CFP

[Glo04] Michael Glotzer. Cleavage furrow positioning. Journal of Cell Biology, 164(3):347–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/3/347.

Garcia:2003:NME

[GLS^{+}03] Michael L. Garcia, Christian S. Lobsiger, Sameer B. Shah, Tom J. Deerinck, John Crum, Darren Young, Christopher M. Ward, Thomas O. Crawford, Takahiro Gotow, Yasuo Uchiyama, Mark H. Ellisman, Nigel A. Calcutt, and Don W. Cleveland. NF-m is an essential target for the myelin-directed “outside-in” signaling cascade that mediates radial axonal growth. Journal of Cell Biology, 163(5):1011–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
5/1011.

Genoud:2002:NCO

[GLSG^{+}02] Stéphane Genoud, Corinna Lappe-Siefke, Sandra Goebbels, Freddy Radtke, Michel Aguet, Steven S. Scherer, Ueli Suter, Klaus-Armin Nave, and Ned Mantei. Notch1 control of oligodendrocyte differentiation in the spinal cord. Journal of Cell Biology, 158(4):709–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/4/709.

Gunawardane:2000:CRD

[GMC^{+}00] Ruwanthi N. Gunawardane, Ona C. Martin, Kan Cao, Lijun Zhang, Kimberly Dej, Akihiro Iwamatsu, and Yixian Zheng. Characterization and Reconstitution of Drosophila γ-Tubulin Ring Complex Subunits. Journal of Cell Biology, 151(7): 1513–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/7/1513.

Giet:2002:DAK

[GMD^{+}02] Régis Giet, Doris McLean, Simon Descamps, Michael J. Lee, Jordan W. Raff, Claude Prigent, and David M. Glover. Drosophila Aurora A kinase is required to localize D–TACC to centrosomes and to regulate astral microtubules. Journal of Cell Biology, 156(3):437–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/3/437.

Gomez-Mouton:2004:DRR

[GMLM^{+}04] Concepción Gómez-Moutón, Rosa Ana Lacalle, Emilia Mira, Sonia Jiménez-Baranda, Domingo F. Barber, Ana C. Carrera, Carlos Martínez-A., and Santos Mañes. Dynamic redistribution of raft domains as an organizing platform for signaling during cell chemotaxis. Journal of Cell Biology, 164 (5):759–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/5/759.

Gilmore:2000:IMS

[GMRS00] Andrew P. Gilmore, Anthony D. Metcalfe, Lewis H. Romer, and Charles H. Streuli. Integrin-mediated survival signals regulate the apoptotic function of Bax through its conformation and subcellular localization. Journal of Cell Biology, 149 (2):431–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/2/431.

Galvez:2002:ERM

[GMRYM^{+}02] Beatriz G. Gálvez, Salomón Matías-Román, María Yáñez-Mó, Francisco Sánchez-Madrid, and Alicia G. Arroyo. ECM regulates MT1-MMP localization with β1 or αvβ3 integrins at distinct cell compartments modulating its internalization and activity on human endothelial cells. Journal of Cell Biology, 159(3):509–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/3/509.

Gerasimenko:2003:NMC

[GMY^{+}03] Julia V. Gerasimenko, Yoshio Maruyama, Kojiro Yano, Nick J. Dolman, Alexei V. Tepikin, Ole H. Petersen, and Oleg V. Gerasimenko. NAADP mobilizes Ca^{2+} from a thapsigargin-sensitive store in the nuclear envelope by activating ryanodine receptors. Journal of Cell Biology, 163(2): 271–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/2/271.

Giordano:2000:RTA

[GMZ^{+}00] Roberto Giordano, Anna Rosa Magnano, Germana Zaccagnini, Carmine Pittoggi, Nicola Moscufo, Rodolfo Lorenzini, and Corrado Spadafora. Reverse Transcriptase Activity in Mature Spermatozoa of Mouse. Journal of Cell Biology, 148(6):1107–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/6/1107.

Gleizes:2001:ULR

[GNDLS^{+}01] Pierre-Emmanuel Gleizes, Jacqueline Noaillac-Depeyre, Isabelle Léger-Silvestre, Frédéric Teulières, Jean-Yves Dauxois, Denys Pommet, Marie-Claude Azum-Gelade, and Nicole Gas. Ultrastructural localization of rRNA shows defective nuclear export of preribosomes in mutants of the Nup82p complex. Journal of Cell Biology, 155(6):923–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
6/923.

Gruneberg:2004:RAB

[GNH^{+}04] Ulrike Gruneberg, Rüdiger Neef, Reiko Honda, Erich A. Nigg, and Francis A. Barr. Relocation of Aurora B from centromeres to the central spindle at the metaphase to anaphase transition requires MKlp2. Journal of Cell Biology, 166 (2):167–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/2/167.

Gerhart:2004:ECE

[GNS^{+}04] Jacquelyn Gerhart, Christine Neely, Benjamin Stewart, Jordanna Perlman, David Beckmann, Margaretha Wallon, Karen Knudsen, and Mindy George-Weinstein. Epiblast cells that express MyoD recruit pluripotent cells to the skeletal muscle lineage. Journal of Cell Biology, 164(5): 739–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/5/739.

Galkin:2001:ADF

[GOL^{+}01] Vitold E. Galkin, Albina Orlova, Natalya Lukoyanova, Willy Wriggers, and Edward H. Egelman. Actin Depolymerizing Factor Stabilizes an Existing State of F–Actin and Can Change the Tilt of F–Actin Subunits. Journal of Cell Biology, 153(1):75–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/1/75.

Galkin:2002:UAB

[GOV^{+}02] Vitold E. Galkin, Albina Orlova, Margaret S. VanLoock, Inna N. Rybakova, James M. Ervasti, and Edward H. Egelman. The utrophin actin-binding domain binds F-actin in two different modes. Journal of Cell Biology, 157(2):243–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
157/2/243.

Galkin:2003:ACU

[GOV^{+}03] Vitold E. Galkin, Albina Orlova, Margaret S. VanLoock, Alexander Shvetsov, Emil Reisler, and Edward H. Egelman. ADF/ cofilin use an intrinsic mode of F-actin instability to disrupt actin filaments. Journal of Cell Biology, 163(5):1057– ??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/5/1057.

Gagnoux-Palacios:2001:SAL

[GPAS^{+}01] Laurent Gagnoux-Palacios, Maryline Allegra, Flavia Spirito, Olivier Pommeret, Christine Romero, Jean paul Ortonne, and Guerrino Meneguzzi. The Short Arm of the Laminin γ2 Chain Plays a Pivotal Role in the Incorporation of Laminin 5 into the Extracellular Matrix and in Cell Adhesion. Journal of Cell Biology, 153(4):835–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/4/835.

Gagnoux-Palacios:2003:CIS

[GPDvH^{+}03] Laurent Gagnoux-Palacios, Michael Dans, Wouter van’t Hof, Agnese Mariotti, Angela Pepe, Guerrino Meneguzzi, Marilyn D. Resh, and Filippo G. Giancotti. Compartmentalization of integrin α6β4 signaling in lipid rafts. Journal of Cell Biology, 162(7):1189–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/7/1189.

Gillmor:2002:GRC

[GPL^{+}02] C. Stewart Gillmor, Patricia Poindexter, Justin Lorieau, Monica M. Palcic, and Chris Somerville. α-glucosidase I is required for cellulose biosynthesis and morphogenesis in Arabidopsis. Journal of Cell Biology, 156(6):1003–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/156/
6/1003.

Gupta:2002:ANF

[GPLS02] Sudhiranjan Gupta, Nicole H. Purcell, Anning Lin, and Subha Sen. Activation of nuclear factor-κB is necessary for myotrophin-induced cardiac hypertrophy. Journal of Cell Biology, 159(6):1019–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/6/1019.

Gu:2002:PNP

[GPZ^{+}02] Wei Gu, Feng Pan, Honglai Zhang, Gary J. Bassell, and Robert H. Singer. A predominantly nuclear protein affecting cytoplasmic localization of β-actin mRNA in fibroblasts and neurons. Journal of Cell Biology, 156(1):41–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/156/
1/41.

Gugasyan:2002:DRP

[GQI^{+}02] Raffi Gugasyan, Cathy Quilici, Stacey T. T. I., Dianne Grail, Anne M. Verhagen, Andrew Roberts, Toshio Kitamura, Ashley R. Dunn, and Peter Lock. Dok-related protein negatively regulates T cell development via its RasGTPase-activating protein and Nck docking sites. Journal of Cell Biology, 158 (1):115–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/1/115.

Goode:2001:AAC

[GRBD01] Bruce L. Goode, Avital A. Rodal, Georjana Barnes, and David G. Drubin. Activation of the Arp2/3 complex by the actin filament binding protein Abp1p. Journal of Cell Biology, 153(3):627–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/3/627.

Gingras:2002:APO

[GRCF02] Jacinthe Gingras, Siamak Rassadi, Ellis Cooper, and Michael Ferns. Agrin plays an organizing role in the formation of sympathetic synapses. Journal of Cell Biology, 158(6):1109– ??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/6/1109.

Green:2000:ASH

[Gre00] Douglas R. Green. Apoptosis and Sphingomyelin Hydrolysis. Journal of Cell Biology, 150(1):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/1/F5.

Griffiths:2003:EDS

[Gri03] Gillian M. Griffiths. Endocytosing the death sentence. Journal of Cell Biology, 160(2):155–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/2/155.

Gupta-Rossi:2004:C

[GRSL^{+}04a] Neetu Gupta-Rossi, Emmanuelle Six, Odile LeBail, Frédérique Logeat, Patricia Chastagner, Annie Olry, Alain Israël, and Christel Brou. Correction. Journal of Cell Biology, 167(3): 563–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/3/563.

Gupta-Rossi:2004:MED

[GRSL^{+}04b] Neetu Gupta-Rossi, Emmanuelle Six, Odile LeBail, Frédérique Logeat, Patricia Chastagner, Annie Olry, Alain Israël, and Christel Brou. Monoubiquitination and endocytosis direct γ-secretase cleavage of activated Notch receptor. Journal of Cell Biology, 166(1):73–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/1/73.

Grosshans:2002:MR

[GS02] Helge Grosshans and Frank J. Slack. Micro-RNAs. Journal of Cell Biology, 156(1):17–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/1/17.

Gardner:2000:ERD

[GSB^{+}00] Richard G. Gardner, Gwendolyn M. Swarbrick, Nathan W. Bays, Stephen R. Cronin, Sharon Wilhovsky, Linda Seelig, Christine Kim, and Randolph Y. Hampton. Endoplasmic Reticulum Degradation Requires Lumen to Cytosol Signaling. Journal of Cell Biology, 151(1):69–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/1/
69.

Grimm:2001:NPF

[GSB^{+}01] Jan Grimm, Martin Sachs, Stefan Britsch, Silvana Di Cesare, Thomas Schwarz-Romond, Kari Alitalo, and Walter Birchmeier. Novel p62dok family members, dok-4 and dok-5, are substrates of the c–Ret receptor tyrosine kinase and mediate neuronal differentiation. Journal of Cell Biology, 154 (2):345–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/2/345.

Gil:2003:ABM

[GSB^{+}03] Orlando D. Gil, Takeshi Sakurai, Ann E. Bradley, Marc Y. Fink, Melanie R. Cassella, James A. Kuo, and Dan P. Felsenfeld. Ankyrin binding mediates L1CAM interactions with static components of the cytoskeleton and inhibits retrograde movement of L1CAM on the cell surface. Journal of Cell Biology, 162(4):719–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/4/719.

Groot:2004:KNP

[GSN^{+}04] Karen R. Groot, Lisa M. Sevilla, Kazunori Nishi, Teresa Di-Colandrea, and Fiona M. Watt. Kazrin, a novel periplakininteracting protein associated with desmosomes and the keratinocyte plasma membrane. Journal of Cell Biology, 166(5): 653–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/5/653.

Gadal:2002:RAP

[GSP^{+}02a] Olivier Gadal, Daniela Strauss, Elisabeth Petfalski, Pierre-Emmanuel Gleizes, Nicole Gas, David Tollervey, and Ed Hurt. Rlp7p is associated with 60S preribosomes, restricted to the granular component of the nucleolus, and required for pre-rRNA processing. Journal of Cell Biology, 157 (6):941–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/6/941.

Gollan:2002:RCA

[GSP^{+}02b] Leora Gollan, Helena Sabanay, Sebastian Poliak, Erik O. Berglund, Barbara Ranscht, and Elior Peles. Retention of a cell adhesion complex at the paranodal junction requires the cytoplasmic region of Caspr. Journal of Cell Biology, 157 (7):1247–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/7/1247.

Gillooly:2001:PP

[GSS01] David J. Gillooly, Anne Simonsen, and Harald Stenmark. Phosphoinositides and phagocytosis. Journal of Cell Biology, 155(1):15–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/1/15.

Gollan:2003:CRP

[GSSP03] Leora Gollan, Daniela Salomon, James L. Salzer, and Elior Peles. Caspr regulates the processing of contactin and inhibits its binding to neurofascin. Journal of Cell Biology, 163(6):1213–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/6/1213.

Golden:2000:MAM

[GSW^{+}00] Andy Golden, Penny L. Sadler, Matthew R. Wallenfang, Jill M. Schumacher, Danielle R. Hamill, Gayle Bates, Bruce Bowerman, Geraldine Seydoux, and Diane C. Shakes. Metaphase to Anaphase (mat) transition–defective mutants in Caenorhabditis elegans. Journal of Cell Biology, 151(7): 1469–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/7/1469.

Gillingham:2004:GAG

[GTBM04] Alison K. Gillingham, Amy Hin Yan Tong, Charles Boone, and Sean Munro. The GTPase Arf1p and the ER to Golgi cargo receptor Erv14p cooperate to recruit the golgin Rud3p to the cis-Golgi. Journal of Cell Biology, 167(2):281–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
167/2/281.

Gross:2002:IRM

[GTD^{+}02] Steven P. Gross, M. Carolina Tuma, Sean W. Deacon, Anna S. Serpinskaya, Amy R. Reilein, and Vladimir I. Gelfand. Interactions and regulation of molecular motors in Xenopus melanophores. Journal of Cell Biology, 156(5): 855–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/5/855.

Gatti:2001:HTO

[GTM^{+}01] Giuliana Gatti, Sara Trifari, Nasrin Mesaeli, J. M. Robert Parker, Marek Michalak, and Jacopo Meldolesi. Head-to-tail oligomerization of calsequestrin. Journal of Cell Biology, 154 (3):525–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/3/525.

Gazarini:2003:CSL

[GTPG03] Marcos L. Gazarini, Andrew P. Thomas, Tullio Pozzan, and Célia R. S. Garcia. Calcium signaling in a low calcium environment. Journal of Cell Biology, 161(1):103–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
1/103.

Greenwood:2000:RFA

[GTPMU00] Jeffrey A. Greenwood, Anne B. Theibert, Glenn D. Prestwich, and Joanne E. Murphy-Ullrich. Restructuring of Focal Adhesion Plaques by Pi 3-Kinase. Journal of Cell Biology, 150(3):627–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/3/627.

Gilberger:2003:CDP

[GTR^{+}03] Tim-Wolf Gilberger, Jennifer K. Thompson, Michael B. Reed, Robert T. Good, and Alan F. Cowman. The cytoplasmic domain of the Plasmodium falciparum ligand EBA-175 is essential for invasion but not protein trafficking. Journal of Cell Biology, 162(2):317–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/2/317.

Gumbiner:2000:RCA

[Gum00] Barry M. Gumbiner. Regulation of Cadherin Adhesive Activity. Journal of Cell Biology, 148(3):399–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/3/
399.

Guiet:2000:CRD

[GV00] Chantal Guiet and Pasquale Vito. Caspase Recruitment Domain (Card–Dependent) Cytoplasmic Filaments Mediate Bcl10-Induced Nf-κ b Activation. Journal of Cell Biology, 148(6):1131–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/6/1131.

Goshima:2003:RMB

[GV03] Gohta Goshima and Ronald D. Vale. The roles of microtubule-based motor proteins in mitosis. Journal of Cell Biology, 162(6):1003–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/6/1003.

Gruss:2004:MSA

[GV04] Oliver J. Gruss and Isabelle Vernos. The mechanism of spindle assembly. Journal of Cell Biology, 166(7):949–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
166/7/949.

Garratt:2000:DRE

[GVT^{+}00] Alistair N. Garratt, Octavian Voiculescu, Piotr Topilko, Patrick Charnay, and Carmen Birchmeier. A Dual Role of erbB2 in Myelination and in Expansion of the Schwann Cell Precursor Pool. Journal of Cell Biology, 148(5):1035– ??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/148/5/1035.

Gross:2000:DMC

[GWBW00] Steven P. Gross, Michael A. Welte, Steven M. Block, and Eric F. Wieschaus. Dynein-mediated cargo transport in vivo. Journal of Cell Biology, 148(5):945–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/5/945.

Gross:2002:COP

[GWBW02] Steven P. Gross, Michael A. Welte, Steven M. Block, and Eric F. Wieschaus. Coordination of opposite-polarity microtubule motors. Journal of Cell Biology, 156(4):715–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
156/4/715.

Gottardi:2001:CSC

[GWG01] Cara J. Gottardi, Ellen Wong, and Barry M. Gumbiner. Ecadherin Suppresses Cellular Transformation by Inhibiting β-Catenin Signaling in an Adhesion–Independent Manner. Journal of Cell Biology, 153(5):1049–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/5/1049.

Gatto:2003:LEA

[GWL03] Cheryl L. Gatto, Barbara J. Walker, and Stephen Lambert. Local ERM activation and dynamic growth cones at Schwann cell tips implicated in efficient formation of nodes of Ranvier. Journal of Cell Biology, 162(3):489–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
3/489.

Gallo:2002:ATR

[GYL02] Gianluca Gallo, Hal F. Yee, and Paul C. Letourneau. Actin turnover is required to prevent axon retraction driven by endogenous actomyosin contractility. Journal of Cell Biology, 158(7):1219–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/7/1219.

Galbraith:2002:RBF

[GYS02] Catherine G. Galbraith, Kenneth M. Yamada, and Michael P. Sheetz. The relationship between force and focal complex development. Journal of Cell Biology, 159(4):695–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
4/695.

Gallo:2002:R

[GZA^{+}02] Rita Gallo, Francesca Zazzeroni, Edoardo Alesse, Claudia Mincione, Ugo Borello, Pasquale Buanne, Roberta D’Eugenio, Andrew R. Mackay, Beatrice Argenti, Roberto Gradini, Matteo A. Russo, Marella Maroder, Giulio Cossu, Luigi Frati, Isabella Screpanti, and Alberto Gulino. Ren. Journal of Cell Biology, 158(4):731–??, August 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/4/

Gineitis:2000:HST

[GZY^{+}00] Arunas A. Gineitis, Irina A. Zalenskaya, Peter M. Yau, E. Morton Bradbury, and Andrei O. Zalensky. Human Sperm Telomere–Binding Complex Involves Histone H2b and Secures Telomere Membrane Attachment. Journal of Cell Biology, 151(7):1591–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/7/1591.

Hsieh:2004:IIM

[HAK^{+}04] Jenny Hsieh, James B. Aimone, Brian K. Kaspar, Tomoko Kuwabara, Kinichi Nakashima, and Fred H. Gage. IGFi instructs multipotent adult neural progenitor cells to become oligodendrocytes. Journal of Cell Biology, 164(1): 111–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/1/111.

Hjalt:2001:PRP

[HAM01] Tord A. Hjalt, Brad A. Amendt, and Jeffrey C. Murray. Pitx2 Regulates Procollagen Lysyl Hydroxylase (Plod) Gene Expression. Journal of Cell Biology, 152(3):545–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
3/545.

Hotary:2000:RCI

[HAP^{+}00] Kevin Hotary, Edward Allen, Antonello Punturieri, Ikuo Yana, and Stephen J. Weiss. Regulation of Cell Invasion and Morphogenesis in a Three–Dimensional Type I Collagen Matrix by Membrane–Type Matrix Metalloproteinases 1, 2, and 3. Journal of Cell Biology, 149(6):1309–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
6/1309.

Hyde:2002:ISP

[HBAF^{+}02] Maureen Hyde, Laura Block-Alper, Jahaira Felix, Paul Webster, and David I. Meyer. Induction of secretory pathway components in yeast is associated with increased stability of their mRNA. Journal of Cell Biology, 156(6):993–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/156/
6/993.

Hirsch:2002:DRM

[HBB^{+}02] Emilio Hirsch, Laura Barberis, Mara Brancaccio, Ornella Azzolino, Dazhong Xu, John M. Kyriakis, Lorenzo Silengo, Filippo G. Giancotti, Guido Tarone, Reinhard Fässler, and Fiorella Altruda. Defective Rac-mediated proliferation and survival after targeted mutation of the β_{1} integrin cytodomain. Journal of Cell Biology, 157(3):481–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/157/
3/481.

Huisman:2004:DCB

[HBB^{+}04] Stephen M. Huisman, Olivia A. M. Bales, Marie Bertrand, Monique F. M. A. Smeets, Steven I. Reed, and Marisa Segal. Differential contribution of Bud6p and Kar9p to microtubule capture and spindle orientation in S. cerevisiae. Journal of Cell Biology, 167(2):231–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/2/231.

Holmbeck:2003:MMD

[HBC^{+}03] Kenn Holmbeck, Paolo Bianco, Kali Chrysovergis, Susan Yamada, and Henning Birkedal-Hansen. MT1-MMP-dependent, apoptotic remodeling of unmineralized cartilage. Journal of Cell Biology, 163(3):661–??, November 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/163/3/
661.

Humphries:2002:DRA

[HBD^{+}02] Christine L. Humphries, Heath I. Balcer, Jessica L. D’Agostino, Barbara Winsor, David G. Drubin, Georjana Barnes, Brenda J. Andrews, and Bruce L. Goode. Direct regulation of Arp2/3 complex activity and function by the actin binding protein coronin. Journal of Cell Biology, 159(6): 993–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/6/993.

Hunyady:2002:DPK

[HBG^{+}02] László Hunyady, Albert J. Baukal, Zsuzsanna Gáborik, Jesus A. Olivares-Reyes, Márta Bor, Márta Szaszák, Robert Lodge, Kevin J. Catt, and Tamas Balla. Differential PI 3kinase dependence of early and late phases of recycling of the internalized AT_{1} angiotensin receptor. Journal of Cell Biology, 157(7):1211–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/7/1211.

Herlan:2004:ATM

[HBH^{+}04] Mark Herlan, Carsten Bornhövd, Kai Hell, Walter Neupert, and Andreas S. Reichert. Alternative topogenesis of Mgm1 and mitochondrial morphology depend on ATP and a functional import motor. Journal of Cell Biology, 165(2):167–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
165/2/167.

Hari:2002:LSR

[HBK^{+}02] Lisette Hari, Véronique Brault, Maurice Kléber, Hye-Youn Lee, Fabian Ille, Rainer Leimeroth, Christian Paratore, Ueli Suter, Rolf Kemler, and Lukas Sommer. Lineage-specific requirements of β-catenin in neural crest development. Journal of Cell Biology, 159(5):867–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/5/867.

Holaska:2001:CRN

[HBL^{+}01] James M. Holaska, Ben E. Black, Dona C. Love, John A. Hanover, John Leszyk, and Bryce M. Paschal. Calreticulin Is a Receptor for Nuclear Export. Journal of Cell Biology, 152(1):127–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/1/127.

Huang:2004:PPP

[HBSJ04] Cai Huang, Christoph H. Borchers, Michael D. Schaller, and Ken Jacobson. Phosphorylation of paxillin by p38MAPK is involved in the neurite extension of PC-12 cells. Journal of Cell Biology, 164(4):593–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/4/593.

Hintermann:2001:IRA

[HBSQ01] Edith Hintermann, Martin Bilban, Andrew Sharabi, and Vito Quaranta. Inhibitory Role of α6β4-Associated Erbb-2 and Phosphoinositide 3-Kinase in Keratinocyte Haptotactic Migration Dependent on α3β1 Integrin. Journal of Cell Biology, 153(3):465–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/3/465.

Hiltbrunner:2001:TAC

[HBV^{+}01] Andreas Hiltbrunner, Jörg Bauer, Pierre-Alexandre Vidi, Sibylle Infanger, Petra Weibel, Morten Hohwy, and Felix Kessler. Targeting of an abundant cytosolic form of the protein import receptor at Toc159 to the outer chloroplast membrane. Journal of Cell Biology, 154(2):309–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/154/2/
309.

Hudson:2002:SDA

[HC02] Andrew M. Hudson and Lynn Cooley. A subset of dynamic actin rearrangements in Drosophila requires the Arp2/3 complex. Journal of Cell Biology, 156(4):677–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/4/
677.

Hayes:2002:TRI

[HCC02a] Susan Hayes, Anil Chawla, and Silvia Corvera. TGFβ receptor internalization into EEA1-enriched early endosomes. Journal of Cell Biology, 158(7):1239–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/7/
1239.

Haynes:2002:HIQ

[HCC02b] Cole M. Haynes, Sabrina Caldwell, and Antony A. Cooper. An HRD /DER-independent ER quality control mechanism involves Rsp5p-dependent ubiquitination and ER–Golgi transport. Journal of Cell Biology, 158(1):91–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/1/

Heidtman:2003:RYC

[HCCB03] Matthew Heidtman, Catherine Z. Chen, Ruth N. Collins, and Charles Barlowe. A role for Yip1p in COPII vesicle biogenesis. Journal of Cell Biology, 163(1):57–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
1/57.

Hyman:2000:EUN

[HCD^{+}00] Joel Hyman, Hong Chen, Pier Paolo Di Fiore, Pietro De Camilli, and Axel T. Brunger. Epsin 1 Undergoes Nucleocytosolic Shuttling and Its Eps15 Interactor Nh_{2} -Terminal Homology (Enth) Domain, Structurally Similar to Armadillo and Heat Repeats, Interacts with the Transcription Factor Promyelocytic Leukemia Zn^{2} + Finger Protein (Plzf). Journal of Cell Biology, 149(3):537–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/3/537.

Haq:2000:GSK

[HCK^{+}00] Syed Haq, Gabriel Choukroun, Zhao Bin Kang, Hardeep Ranu, Takashi Matsui, Anthony Rosenzweig, Jeffrey D. Molkentin, Alessandro Alessandrini, James Woodgett, Roger Hajjar, Ashour Michael, and Thomas Force. Glycogen Synthase Kinase-3β Is a Negative Regulator of Cardiomyocyte Hypertrophy. Journal of Cell Biology, 151(1):117–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
151/1/117.

Hauf:2003:SMH

[HCL^{+}03] Silke Hauf, Richard W. Cole, Sabrina LaTerra, Christine Zimmer, Gisela Schnapp, Rainer Walter, Armin Heckel, Jacques van Meel, Conly L. Rieder, and Jan-Michael Peters. The small molecule Hesperadin reveals a role for Aurora B in correcting kinetochore–microtubule attachment and in maintaining the spindle assembly checkpoint. Journal of Cell Biology, 161(2):281–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/281.

Heil-Chapdelaine:2000:CPN

[HCOC00] Richard A. Heil-Chapdelaine, Jessica R. Oberle, and John A. Cooper. The Cortical Protein Num1p Is Essential for Dynein–Dependent Interactions of Microtubules with the Cortex. Journal of Cell Biology, 151(6):1337–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
6/1337.

Hume:2001:RRP

[HCR^{+}01] Alistair N. Hume, Lucy M. Collinson, Andrzej Rapak, Anita Q. Gomes, Colin R. Hopkins, and Miguel C. Seabra. Rab27a Regulates the Peripheral Distribution of Melanosomes in Melanocytes. Journal of Cell Biology, 152 (4):795–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/4/795.

Haugh:2000:SSF

[HCTM00] Jason M. Haugh, Franca Codazzi, Mary Teruel, and Tobias Meyer. Spatial Sensing in Fibroblasts Mediated by 3’ Phosphoinositides. Journal of Cell Biology, 151(6):1269– ??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/6/1269.

Hagting:2002:HSP

[HdEV^{+}02] Anja Hagting, Nicole den Elzen, Hartmut C. Vodermaier, Irene C. Waizenegger, Jan-Michael Peters, and Jonathon Pines. Human securin proteolysis is controlled by the spindle checkpoint and reveals when the APC/C switches from activation by Cdc20 to Cdh1. Journal of Cell Biology, 157 (7):1125–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/7/1125.

Hutchings:2002:TCL

[HDH02] Nathan R. Hutchings, John E. Donelson, and Kent L. Hill. Trypanin is a cytoskeletal linker protein and is required for cell motility in African trypanosomes. Journal of Cell Biology, 156(5):867–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/5/867.

Hammarlund:2000:MSD

[HDJ00] Marc Hammarlund, Warren S. Davis, and Erik M. Jorgensen. Mutations in β-spectrin disrupt axon outgrowth and sarcomere structure. Journal of Cell Biology, 149(4):931–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
4/931.

Hu:2000:AUI

[HDL^{+}00] Ping Hu, Fang-Ming Deng, Feng-Xia Liang, Chuan-Min Hu, Anna B. Auerbach, Ellen Shapiro, Xue-Ru Wu, Bechara Kachar, and Tung-Tien Sun. Ablation of Uroplakin III Gene Results in Small Urothelial Plaques, Urothelial Leakage, and Vesicoureteral Reflux. Journal of Cell Biology, 151(5):961– ??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/5/961.

Huber:2002:IBD

[HDL02] Jochen Huber, Achim Dickmanns, and Reinhard Lührmann. The importin-β binding domain of snurportin1 is responsible for the Ran- and energy-independent nuclear import of spliceosomal U snRNPs in vitro. Journal of Cell Biology, 156(3):467–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/3/467.

Hoffmann:2001:PPI

[HdO^{+}01] Peter R. Hoffmann, Aimee M. deCathelineau, Carol Anne Ogden, Yann Leverrier, Donna L. Bratton, David L. Daleke, Anne J. Ridley, Valerie A. Fadok, and Peter M. Henson. Phosphatidylserine (PS) induces PS receptor–mediated macropinocytosis and promotes clearance of apoptotic cells. Journal of Cell Biology, 155(4):649–??, November 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/155/4/
649.

Haller:2001:FPE

[HDP^{+}01] Thomas Haller, Paul Dietl, Kristian Pfaller, Manfred Frick, Norbert Mair, Markus Paulmichl, Michael W. Hess, Johannes Fürst, and Karl Maly. Fusion pore expansion is a slow, discontinuous, and Ca^{2+}-dependent process regulating secretion from alveolar type II cells. Journal of Cell Biology, 155(2): 279–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/2/279.

Held:2003:PLS

[Hel03] Michael J. Held. Proposed legislation supports an untested publishing model. Journal of Cell Biology, 162(2):171–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
162/2/171.

Hemler:2001:STF

[Hem01] Martin E. Hemler. Specific tetraspanin functions. Journal of Cell Biology, 155(7):1103–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/7/1103.

Hoffmann:2001:AER

[HEN^{+}01] Isabelle Hoffmann, Emmanuel Eugène, Xavier Nassif, Pierre-Olivier Couraud, and Sandrine Bourdoulous. Activation of ErbB2 receptor tyrosine kinase supports invasion of endothelial cells by Neisseria meningitidis. Journal of Cell Biology, 155(1):133–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/1/133.

Hakeda:2000:RKG

[HES00] Satoko Hakeda, Sachiko Endo, and Kaoru Saigo. Requirements of kettin, a giant muscle protein highly conserved in overall structure in evolution, for normal muscle function, viability, and flight activity of Drosophila. Journal of Cell Biology, 148(1):101–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/1/101.

Hong:2001:MNT

[HEW^{+}01] Leexan Hong, Tricia Elbl, James Ward, Clara Franzini-Armstrong, Krystyna K. Rybicka, Beth K. Gatewood, David L. Baillie, and Elizabeth A. Bucher. MUP-4 is a novel transmembrane protein with functions in epithelial cell adhesion in Caenorhabditis elegans. Journal of Cell Biology, 154 (2):403–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/2/403.

Hagiwara:2001:SMA

[HF01] Hiroki Hagiwara and Justin R. Fallon. Shaping Membrane Architecture. Journal of Cell Biology, 153(7):0–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
7/F39.

Hu:2003:DCA

[HF03] Yue Hu and Mark E. Fortini. Different cofactor activities in γ-secretase assembly. Journal of Cell Biology, 161(4):685–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
161/4/685.

Harrington:2004:TTS

[HFG^{+}04] Laura S. Harrington, Greg M. Findlay, Alex Gray, Tatiana Tolkacheva, Simon Wigfield, Heike Rebholz, Jill Barnett, Nick R. Leslie, Susan Cheng, Peter R. Shepherd, Ivan Gout, C. Peter Downes, and Richard F. Lamb. The TSC1-2 tumor suppressor controls insulin–PI3K signaling via regulation of IRS proteins. Journal of Cell Biology, 166(2):213–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
2/213.

Hood:2003:DVI

[HFK^{+}03] John D. Hood, Ricardo Frausto, William B. Kiosses, Martin A. Schwartz, and David A. Cheresh. Differential αv integrin–mediated Ras–ERK signaling during two pathways of angiogenesis. Journal of Cell Biology, 162(5):933–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/162/5/933.

Horsley:2001:RGM

[HFM^{+}01] Valerie Horsley, Bret B. Friday, Sarah Matteson, Kristy Miller Kegley, Jonathan Gephart, and Grace K. Pavlath. Regulation of the Growth of Multinucleated Muscle Cells by an Nfatc2-Dependent Pathway. Journal of Cell Biology, 153 (2):329–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/2/329.

Horne:2003:RMK

[HG03] Melinda M. Horne and Thomas M. Guadagno. A requirement for MAP kinase in the assembly and maintenance of the mitotic spindle. Journal of Cell Biology, 161(6):1021–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
6/1021.

Heumann:2000:TAR

[HGB^{+}00] Rolf Heumann, Christoph Goemans, Daniela Bartsch, Kurt Lingenhöhl, Peter C. Waldmeier, Bastian Hengerer, Peter R. Allegrini, Karl Schellander, Erwin F. Wagner, Thomas Arendt, Rigobert H. Kamdem, Kirstin Obst-Pernberg, Frank Narz, Petra Wahle, and Hartmut Berns. Transgenic Activation of Ras in Neurons Promotes Hypertrophy and Protects from Lesion–Induced Degeneration. Journal of Cell Biology, 151(7):1537–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/7/1537.

Hong-Geller:2000:CRS

[HGC00] Elizabeth Hong-Geller and Richard A. Cerione. Cdc42 and Rac stimulate exocytosis of secretory granules by activating the Ip_{3} /calcium pathway in Rbl-2h3 mast cells. Journal of Cell Biology, 148(3):481–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/3/481.

Hinz:2002:NTP

[HGC02] Boris Hinz, Giulio Gabbiani, and Christine Chaponnier. The NH_{2}-terminal peptide of α-smooth muscle actin inhibits force generation by the myofibroblast in vitro and in vivo. Journal of Cell Biology, 157(4):657–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/4/657.

Hellstrom:2001:LPL

[HGK^{+}01] Mats Hellström, Holger Gerhardt, Mattias Kalén, Xuri Li, Ulf Eriksson, Hartwig Wolburg, and Christer Betsholtz. Lack of Pericytes Leads to Endothelial Hyperplasia and Abnormal Vascular Morphogenesis. Journal of Cell Biology, 153 (3):543–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/3/543.

Henkel:2000:IMP

[HGP^{+}00] Jennifer R. Henkel, Gregory A. Gibson, Paul A. Poland, Mark A. Ellis, Rebecca P. Hughey, and Ora A. Weisz. Influenza M2 proton channel activity selectively inhibits trans-Golgi network release of apical membrane and secreted proteins in polarized madin–darby canine kidney cells. Journal of Cell Biology, 148(3):495–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/3/495.

Huckaba:2004:LCI

[HGP^{+}04] Thomas M. Huckaba, Anna Card Gay, Luiz Fernando Pantalena, Hyeong-Cheol Yang, and Liza A. Pon. Live cell imaging of the assembly, disassembly, and actin cable-dependent movement of endosomes and actin patches in the budding yeast, Saccharomyces cerevisiae. Journal of Cell Biology, 167(3):519–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/519.

Heese:2001:FCK

[HGS^{+}01] Maren Heese, Xavier Gansel, Liliane Sticher, Peter Wick, Markus Grebe, Fabienne Granier, and Gerd Jürgens. Functional characterization of the KNOLLE-interacting t– SNARE AtSNAP33 and its role in plant cytokinesis. Journal of Cell Biology, 155(2):239–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/2/239.

Howell:2000:VMD

[HHF^{+}00] B. J. Howell, D. B. Hoffman, G. Fang, A. W. Murray, and E. D. Salmon. Visualization of Mad2 Dynamics at Kinetochores, along Spindle Fibers, and at Spindle Poles in Living Cells. Journal of Cell Biology, 150(6):1233–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
6/1233.

Harris:2000:MSD

[HHHJ00] Todd W. Harris, Erika Hartwieg, H. Robert Horvitz, and Erik M. Jorgensen. Mutations in Synaptojanin Disrupt Synaptic Vesicle Recycling. Journal of Cell Biology, 150(3): 589–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/3/589.

Henry:2004:UPM

[HHJS04] Rebecca M. Henry, Adam D. Hoppe, Nikhil Joshi, and Joel A. Swanson. The uniformity of phagosome maturation in macrophages. Journal of Cell Biology, 164(2):185–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
164/2/185.

He:2004:GDT

[HHM^{+}04] Cynthia Y. He, Helen H. Ho, Joerg Malsam, Cecile Chalouni, Christopher M. West, Elisabetta Ullu, Derek Toomre, and Graham Warren. Golgi duplication in Trypanosoma brucei. Journal of Cell Biology, 165(3):313–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/3/313.

Howard:2002:SST

[HHOP02] James P. Howard, Jenna L. Hutton, John M. Olson, and Gregory S. Payne. Sla1p serves as the targeting signal recognition factor for NPFX_{(1,2)} D-mediated endocytosis. Journal of Cell Biology, 157(2):315–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/2/315.

Hoover:2003:CTD

[HHS03] Cameron L. Hoover, Lutz G. W. Hilgenberg, and Martin A. Smith. The COOH-terminal domain of agrin signals via a synaptic receptor in central nervous system neurons. Journal of Cell Biology, 161(5):923–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/5/923.

Hatzfeld:2000:FPD

[HHSV00] Mechthild Hatzfeld, Christof Haffner, Katrin Schulze, and Ute Vinzens. The Function of Plakophilin 1 in Desmosome Assembly and Actin Filament Organization. Journal of Cell Biology, 149(1):209–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/1/209.

Hanyu:2001:DSE

[HIE^{+}01] Aki Hanyu, Yasuhiro Ishidou, Takanori Ebisawa, Tomomasa Shimanuki, Takeshi Imamura, and Kohei Miyazono. The N domain of Smad7 is essential for specific inhibition of transforming growth factor-β signaling. Journal of Cell Biology, 155(6):1017–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/6/1017.

Huttelmaier:2001:RDC

[HIG^{+}01] Stefan Hüttelmaier, Susanne Illenberger, Irina Grosheva, Manfred Rüdiger, Robert H. Singer, and Brigitte M. Jockusch. Raver1, a dual compartment protein, is a ligand for PTB/hnRNPI and microfilament attachment proteins. Journal of Cell Biology, 155(5):775–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/5/775.

Henkemeyer:2003:MER

[HIN^{+}03] Mark Henkemeyer, Olga S. Itkis, Michelle Ngo, Peter W. Hickmott, and Iryna M. Ethell. Multiple EphB receptor tyrosine kinases shape dendritic spines in the hippocampus. Journal of Cell Biology, 163(6):1313–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/6/1313.

Horiki:2004:SSO

[HIO^{+}04] Mitsuru Horiki, Takeshi Imamura, Mina Okamoto, Makoto Hayashi, Junko Murai, Akira Myoui, Takahiro Ochi, Kohei Miyazono, Hideki Yoshikawa, and Noriyuki Tsumaki. Smad6/ smurf1 overexpression in cartilage delays chondrocyte hypertrophy and causes dwarfism with osteopenia. Journal of Cell Biology, 165(3):433–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/3/433.

Hori:2002:TCS

[HIT^{+}02] Osamu Hori, Fusae Ichinoda, Takashi Tamatani, Atsushi Yamaguchi, Naoya Sato, Kentaro Ozawa, Yasuko Kitao, Mayuki Miyazaki, Heather P. Harding, David Ron, Masaya Tohyama, David M. Stern, and Satoshi Ogawa. Transmission of cell stress from endoplasmic reticulum to mitochondria. Journal of Cell Biology, 157(7):1151–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/7/1151.

Haigh:2002:NRB

[HJ02] Nora G. Haigh and Arthur E. Johnson. A new role for BiP. Journal of Cell Biology, 156(2):261–??, January 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/2/
261.

Herker:2004:CAL

[HJL^{+}04] Eva Herker, Helmut Jungwirth, Katharina A. Lehmann, Corinna Maldener, Kai-Uwe Fröhlich, Silke Wissing, Sabrina Büttner, Markus Fehr, Stephan Sigrist, and Frank Madeo. Chronological aging leads to apoptosis in yeast. Journal of Cell Biology, 164(4):501–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/4/501.

Hardwick:2000:MEN

[HJSM00] Kevin G. Hardwick, Raymond C. Johnston, Dana L. Smith, and Andrew W. Murray. MAD3 Encodes a Novel Component of the Spindle Checkpoint Which Interacts with Bub3p, Cdc20p, and Mad2p. Journal of Cell Biology, 148 (5):871–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/5/871.

Harder:2000:SAR

[HK00] Thomas Harder and Marina Kuhn. Selective Accumulation of Raft–Associated Membrane Protein Lat in T Cell Receptor Signaling Assemblies. Journal of Cell Biology, 151(2): 199–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/2/199.

Hwang:2003:SOS

[HKBH03] Eric Hwang, Justine Kusch, Yves Barral, and Tim C. Huffaker. Spindle orientation in Saccharomyces cerevisiae depends on the transport of microtubule ends along polarized actin cables. Journal of Cell Biology, 161(3):483–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
3/483.

Hitomi:2004:ICE

[HKE^{+}04] Junichi Hitomi, Taiichi Katayama, Yutaka Eguchi, Takashi Kudo, Manabu Taniguchi, Yoshihisa Koyama, Takayuki Manabe, Satoru Yamagishi, Yoshio Bando, Kazunori Imaizumi, Yoshihide Tsujimoto, and Masaya Tohyama. Involvement of caspase-4 in endoplasmic reticulum stressinduced apoptosis and Aβ-induced cell death. Journal of Cell Biology, 165(3):347–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/3/347.

Hannak:2001:AKR

[HKHO01] Eva Hannak, Matthew Kirkham, Anthony A. Hyman, and Karen Oegema. Aurora-A kinase is required for centrosome maturation in Caenorhabditis elegans. Journal of Cell Biology, 155(7):1109–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/7/1109.

Ho:2000:NCD

[HKJ00] Jennifer Hei-Ngam Ho, George Kallstrom, and Arlen W. Johnson. Nmd3p Is a Crm1p–Dependent Adapter Protein for Nuclear Export of the Large Ribosomal Subunit. Journal of Cell Biology, 151(5):1057–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/5/1057.

Holzl:2000:RCD

[HKK^{+}00] Harald Hölzl, Barbara Kapelari, Josef Kellermann, Erika Seemüller, Máté Sümegi, Andor Udvardy, Ohad Medalia, Joseph Sperling, Shirley A. Müller, Andreas Engel, and Wolfgang Baumeister. The Regulatory Complex of Drosophila melanogaster 26s Proteasomes. Journal of Cell Biology, 150 (1):119–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/1/119.

Hyytiainen:2003:LTB

[HKO03] Marko Hyytiäinen and Jorma Keski-Oja. Latent TGF-β binding protein LTBP-2 decreases fibroblast adhesion to fibronectin. Journal of Cell Biology, 163(6):1363–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
163/6/1363.

Helwani:2004:CNC

[HKP^{+}04] Falak M. Helwani, Eva M. Kovacs, Andrew D. Paterson, Suzie Verma, Radiya G. Ali, Alan S. Fanning, Scott A. Weed, and Alpha S. Yap. Cortactin is necessary for E-cadherin-mediated contact formation and actin reorganization. Journal of Cell Biology, 164(6):899–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/6/899.

Hirst:2000:FPA

[HLB^{+}00] Jennifer Hirst, Winnie W. Y. Lui, Nicholas A. Bright, Nicholas Totty, Matthew N. J. Seaman, and Margaret S. Robinson. A family of proteins with γ-adaptin and VHS domains that facilitate trafficking between the trans-Golgi network and the vacuole/lysosome. Journal of Cell Biology, 149(1):67–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/1/67.

Hausser:2002:SRL

[HLB^{+}02] Angelika Hausser, Gisela Link, Linda Bamberg, Annett Burzlaff, Sylke Lutz, Klaus Pfizenmaier, and Franz-Josef Johannes. Structural requirements for localization and activation of protein kinase C μ (PKCμ) at the Golgi compartment. Journal of Cell Biology, 156(1):65–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/156/
1/65.

Helliwell:2001:CUL

[HLK01] Stephen B. Helliwell, Sascha Losko, and Chris A. Kaiser. Components of a Ubiquitin Ligase Complex Specify Polyubiquitination and Intracellular Trafficking of the General Amino Acid Permease. Journal of Cell Biology, 153(4): 649–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/4/649.

Heun:2001:PDO

[HLRG01] Patrick Heun, Thierry Laroche, M. K. Raghuraman, and Susan M. Gasser. The Positioning and Dynamics of Origins of Replication in the Budding Yeast Nucleus. Journal of Cell Biology, 152(2):385–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/2/385.

Harvey:2000:CDC

[HLU00] Kevin J. Harvey, Dunja Lukovic, and David S. Ucker. Caspase-dependent Cdk Activity Is a Requisite Effector of Apoptotic Death Events. Journal of Cell Biology, 148(1): 59–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/1/59.

Huang:2000:RIL

[HLZW00] Yao Huang, Ji Li, Yongjun Zhang, and Chuanyue Wu. The Roles of Integrin–Linked Kinase in the Regulation of Myogenic Differentiation. Journal of Cell Biology, 150(4): 861–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/4/861.

Hetzer:2000:ADR

[HM00] Martin Hetzer and Iain W. Mattaj. An ATP–dependent, ran– independent mechanism for nuclear import of the U1a and U2b” spliceosome proteins. Journal of Cell Biology, 148(2): 293–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/2/293.

Howe:2001:HRK

[HMAM01] Mary Howe, Kent L. McDonald, Donna G. Albertson, and Barbara J. Meyer. Him-10 Is Required for Kinetochore Structure and Function on Caenorhabditis elegans Holocentric Chromosomes. Journal of Cell Biology, 153(6):1227–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
6/1227.

Howell:2001:CDD

[HMC^{+}01] B. J. Howell, B. F. McEwen, J. C. Canman, D. B. Hoffman, E. M. Farrar, C. L. Rieder, and E. D. Salmon. Cytoplasmic dynein/dynactin drives kinetochore protein transport to the spindle poles and has a role in mitotic spindle checkpoint inactivation. Journal of Cell Biology, 155(7):1159–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
155/7/1159.

Handwerger:2003:SSD

[HMG03] Korie E. Handwerger, Christine Murphy, and Joseph G. Gall. Steady-state dynamics of Cajal body components in the Xenopus germinal vesicle. Journal of Cell Biology, 160(4): 495–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/4/495.

Hsia:2003:DRC

[HMH^{+}03] Datsun A. Hsia, Satyajit K. Mitra, Christof R. Hauck, Daniel N. Streblow, Jay A. Nelson, Dusko Ilic, Shuang Huang, Erguang Li, Glen R. Nemerow, Jay Leng, Kathryn S. R. Spencer, David A. Cheresh, and David D. Schlaepfer. Differential regulation of cell motility and invasion by FAK. Journal of Cell Biology, 160(5):753–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/5/753.

Hirota:2000:ZRA

[HMN^{+}00] Toru Hirota, Tetsuro Morisaki, Yasuyuki Nishiyama, Tomotoshi Marumoto, Kenji Tada, Toshihiro Hara, Norio Masuko, Masaki Inagaki, Katsuyoshi Hatakeyama, and Hideyuki Saya. Zyxin, a regulator of actin filament assembly, targets the mitotic apparatus by interacting with H-Warts/Lats1 tumor suppressor. Journal of Cell Biology, 149(5):1073–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
149/5/1073.

Hillmer:2001:VSP

[HMRH01] Stefan Hillmer, Ali Movafeghi, David G. Robinson, and Giselbert Hinz. Vacuolar Storage Proteins Are Sorted in the Cis– Cisternae of the Pea Cotyledon Golgi Apparatus. Journal of Cell Biology, 152(1):41–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/1/41.

Helfand:2002:RCD

[HMVG02] Brian T. Helfand, Atsushi Mikami, Richard B. Vallee, and Robert D. Goldman. A requirement for cytoplasmic dynein and dynactin in intermediate filament network assembly and organization. Journal of Cell Biology, 157(5):795–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/157/
5/795.

Hoshino:2003:SGR

[HN03] Mitsunobu Hoshino and Shun Nakamura. Small GTPase Rin induces neurite outgrowth through Rac/Cdc42 and calmodulin in PC12 cells. Journal of Cell Biology, 163(5):1067– ??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/5/1067.

Hribal:2003:RIL

[HNK^{+}03] Marta L. Hribal, Jun Nakae, Tadahiro Kitamura, John R. Shutter, and Domenico Accili. Regulation of insulin-like growth factor-dependent myoblast differentiation by Foxo forkhead transcription factors. Journal of Cell Biology, 162 (4):535–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/4/535.

Hannak:2002:KDA

[HOK^{+}02] Eva Hannak, Karen Oegema, Matthew Kirkham, Pierre Gönczy, Bianca Habermann, and Anthony A. Hyman.

The kinetically dominant assembly pathway for centrosomal asters in Caenorhabditis elegans is γ-tubulin dependent. Journal of Cell Biology, 157(4):591–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/4/591.

Hollenbeck:2001:KD

[Hol01] P. J. Hollenbeck. Kinesin Delivers. Journal of Cell Biology, 152(5):0–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/5/F25.

Hausmann:2000:PAA

[HOvD^{+}00] George Hausmann, Lorraine A. O’Reilly, Rosemary van Driel, Jennifer G. Beaumont, Andreas Strasser, Jerry M. Adams, and David C. S. Huang. Pro-apoptotic Apoptosis Protease–Activating Factor 1 (Apaf-1) Has a Cytoplasmic Localization Distinct from Bcl-2 or Bcl–X_{L}. Journal of Cell Biology, 149(3):623–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/3/623.

Hoyt:2001:NVS

[Hoy01] M. Andrew Hoyt. A new view of the spindle checkpoint. Journal of Cell Biology, 154(5):909–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/5/909.

Higgs:2000:ACP

[HP00] Henry N. Higgs and Thomas D. Pollard. Activation by Cdc42 and Pip_{2} of Wiskott–Aldrich syndrome protein (Wasp) stimulates actin nucleation by Arp2/3 complex. Journal of Cell Biology, 150(6):1311–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/6/1311.

Horsley:2002:N

[HP02] Valerie Horsley and Grace K. Pavlath. Nfat. Journal of Cell Biology, 156(5):771–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/5/771.

Horsley:2003:PFS

[HP03] Valerie Horsley and Grace K. Pavlath. Prostaglandin F_{2α} stimulates growth of skeletal muscle cells via an NFATC2dependent pathway. Journal of Cell Biology, 161(1):111–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
161/1/111.

Harris:2004:AJD

[HP04] Tony J. C. Harris and Mark Peifer. Adherens junctiondependent and -independent steps in the establishment of epithelial cell polarity in Drosophila. Journal of Cell Biology, 167(1):135–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/1/135.

Huang:2001:AAF

[HPE^{+}01] James N. Huang, Iha Park, Eric Ellingson, Laurie E. Littlepage, and David Pellman. Activity of the APC^{Cdh1} form of the anaphase-promoting complex persists until S phase and prevents the premature expression of Cdc20p. Journal of Cell Biology, 154(1):85–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/1/85.

Hernandez:2003:SPC

[HPFG03] Lorraine D. Hernandez, Marc Pypaert, Richard A. Flavell, and Jorge E. Galán. A Salmonella protein causes macrophage cell death by inducing autophagy. Journal of Cell Biology, 163(5):1123–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/5/1123.

Huen:2002:IFM

[HPG^{+}02] Arthur C. Huen, Jung K. Park, Lisa M. Godsel, Xuejun Chen, Leslie J. Bannon, Evangeline V. Amargo, Tracie Y. Hudson, Anne K. Mongiu, Irene M. Leigh, David P. Kelsell, Barry M. Gumbiner, and Kathleen J. Green. Intermediate filament–membrane attachments function synergistically with actin-dependent contacts to regulate intercellular adhesive strength. Journal of Cell Biology, 159(6):1005– ??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/6/1005.

Hazan:2000:EEC

[HPQ^{+}00] Rachel B. Hazan, Greg R. Phillips, Rui Fang Qiao, Larry Norton, and Stuart A. Aaronson. Exogenous Expression of N–Cadherin in Breast Cancer Cells Induces Cell Migration, Invasion, and Metastasis. Journal of Cell Biology, 148(4): 779–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/4/779.

Huh:2004:RRP

[HPS^{+}04] Michael S. Huh, Maura H. Parker, Anthony Scimè, Robin Parks, and Michael A. Rudnicki. Rb is required for progression through myogenic differentiation but not maintenance of terminal differentiation. Journal of Cell Biology, 166(6): 865–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/6/865.

Hirai:2001:EMM

[HRB^{+}01] Yohei Hirai, Derek Radisky, Rosanne Boudreau, Marina Simian, Mary E. Stevens, Yumiko Oka, Kyoko Takebe, Shinichiro Niwa, and Mina J. Bissell. Epimorphin mediates mammary luminal morphogenesis through control of C/ EBPβ. Journal of Cell Biology, 153(4):785–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/153/4/
785.

Hofmann:2001:CRN

[HRE^{+}01] Wilma Hofmann, Beate Reichart, Andrea Ewald, Eleonora Müller, Iris Schmitt, Roland H. Stauber, Friedrich Lottspeich, Brigitte M. Jockusch, Ulrich Scheer, Joachim Hauber, and Marie-Christine Dabauvalle. Cofactor requirements for nuclear export of rev response element (Rre)–and constitutive transport element (Cte)–containing retroviral RNAs. Journal of Cell Biology, 152(5):895–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/5/895.

Hu:2002:NPT

[HRM02] Ke Hu, David S. Roos, and John M. Murray. A novel polymer of tubulin forms the conoid of Toxoplasma gondii. Journal of Cell Biology, 156(6):1039–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/6/1039.

Hollinshead:2001:VVU

[HRV^{+}01] Michael Hollinshead, Gaener Rodger, Henriette Van Eijl, Mansun Law, Ruth Hollinshead, David J. T. Vaux, and Geoffrey L. Smith. Vaccinia virus utilizes microtubules for movement to the cell surface. Journal of Cell Biology, 154 (2):389–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/2/389.

Harsay:2002:SYV

[HS02] Edina Harsay and Randy Schekman. A subset of yeast vacuolar protein sorting mutants is blocked in one branch of the exocytic pathway. Journal of Cell Biology, 156(2):271–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
156/2/271.

Hofken:2004:NRM

[HS04] Thomas Höfken and Elmar Schiebel. Novel regulation of mitotic exit by the Cdc42 effectors Gic1 and Gic2. Journal of Cell Biology, 164(2):219–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/2/219.

Homolya:2000:CCC

[HSB00] László Homolya, Thomas H. Steinberg, and Richard C. Boucher. Cell to Cell Communication in Response to Mechanical Stress via Bilateral Release of ATP and Utp in Polarized Epithelia. Journal of Cell Biology, 150(6):1349– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/6/1349.

Hsu:2001:IMG

[HSC01] Wei Hsu, Reena Shakya, and Frank Costantini. Impaired mammary gland and lymphoid development caused by inducible expression of Axin in transgenic mice. Journal of Cell Biology, 155(6):1055–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/6/1055.

Hartman:2000:PEC

[HSKG00] Theresa Hartman, Kristen Stead, Douglas Koshland, and Vincent Guacci. Pds5p Is an Essential Chromosomal Protein Required for Both Sister Chromatid Cohesion and Condensation in Saccharomyces cerevisiae. Journal of Cell Biology, 151(3):613–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/3/613.

Herskovits:2002:AEM

[HSMB02] Anat A. Herskovits, Eyal Shimoni, Abraham Minsky, and Eitan Bibi. Accumulation of endoplasmic membranes and novel membrane-bound ribosome–signal recognition particle receptor complexes in Escherichia coli. Journal of Cell Biology, 159(3):403–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/3/403.

Hobbs:2001:MMO

[HSMJ01] Alyson E. Aiken Hobbs, Maithreyan Srinivasan, J. Michael McCaffery, and Robert E. Jensen. Mmm1p, a mitochondrial outer membrane protein, is connected to mitochondrial DNA (Mtdna) nucleoids and required for Mtdna stability. Journal of Cell Biology, 152(2):401–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/2/401.

Hornemann:2000:ISI

[HSW00] Thorsten Hornemann, Martin Stolz, and Theo Wallimann. Isoenzyme-specific Interaction of Muscle–Type Creatine Kinase with the Sarcomeric M–Line Is Mediated by Nh_{2} - Terminal Lysine Charge–Clamps. Journal of Cell Biology, 149(6):1225–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/6/1225.

Horikawa:2001:RJD

[HT01] Kazuki Horikawa and Masatoshi Takeichi. Requirement of the juxtamembrane domain of the cadherin cytoplasmic tail for morphogenetic cell rearrangement during myotome development. Journal of Cell Biology, 155(7):1297–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
7/1297.

Halet:2004:CPR

[HTPC04] Guillaume Halet, Richard Tunwell, Scott J. Parkinson, and John Carroll. Conventional PKCs regulate the temporal pattern of Ca^{2+} oscillations at fertilization in mouse eggs. Journal of Cell Biology, 164(7):1033–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/7/1033.

Heit:2002:ISH

[HTRK02] Bryan Heit, Samantha Tavener, Eko Raharjo, and Paul Kubes. An intracellular signaling hierarchy determines direction of migration in opposing chemotactic gradients. Journal of Cell Biology, 159(1):91–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/1/91.

Horowitz:2002:FGF

[HTS02] Arie Horowitz, Eugene Tkachenko, and Michael Simons. Fibroblast growth factor–specific modulation of cellular response by syndecan-4. Journal of Cell Biology, 157(4): 715–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/4/715.

Harada:2002:MRD

[HTT^{+}02] Akihiro Harada, Junlin Teng, Yosuke Takei, Keiko Oguchi, and Nobutaka Hirokawa. MAP2 is required for dendrite elongation, PKA anchoring in dendrites, and proper PKA signal transduction. Journal of Cell Biology, 158(3):541–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
3/541.

Huang:2002:BEF

[Hua02] Sui Huang. Building an efficient factory. Journal of Cell Biology, 157(5):739–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/5/739.

Hillier:2003:AOP

[HV03] Brian J. Hillier and Victor D. Vacquier. Amassin, an olfactomedin protein, mediates the massive intercellular adhesion of sea urchin coelomocytes. Journal of Cell Biology, 160(4): 597–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/4/597.

Huelsken:2000:RCA

[HVB^{+}00] Joerg Huelsken, Regina Vogel, Volker Brinkmann, Bettina Erdmann, Carmen Birchmeier, and Walter Birchmeier. Requirement for β-Catenin in Anterior–Posterior Axis Formation in Mice. Journal of Cell Biology, 148(3):567–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
148/3/567.

Hoepfner:2001:RVA

[HvdBP^{+}01] Dominic Hoepfner, Marlene van den Berg, Peter Philippsen, Henk F. Tabak, and Ewald H. Hettema. A role for Vps1p, actin, and the Myo2p motor in peroxisome abundance and inheritance in Saccharomyces cerevisiae. Journal of Cell Biology, 155(6):979–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/6/979.

Herrmann:2003:MTD

[HvdHG^{+}03] Thomas Herrmann, Frank van der Hoeven, Hermann-Josef Gröne, Adrian Francis Stewart, Lutz Langbein, Iris Kaiser, Gerhard Liebisch, Isabella Gosch, Florian Buchkremer, Wolfgang Drobnik, Gerd Schmitz, and Wolfgang Stremmel. Mice with targeted disruption of the fatty acid transport protein 4 (Fatp 4, Slc27a4) gene show features of lethal restrictive dermopathy. Journal of Cell Biology, 161(6):1105–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
6/1105.

Hill:2001:RDB

[HvdKDS01] Elaine Hill, Jeroen van der Kaay, C. Peter Downes, and Elizabeth Smythe. The Role of Dynamin and Its Binding Partners in Coated Pit Invagination and Scission. Journal of Cell Biology, 152(2):309–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/2/309.

Heremans:2002:REN

[HViV^{+}02] Yves Heremans, Mark Van De Casteele, Peter in’t Veld, Gerard Gradwohl, Palle Serup, Ole Madsen, Daniel Pipeleers, and Harry Heimberg. Recapitulation of embryonic neuroendocrine differentiation in adult human pancreatic duct cells expressing neurogenin 3. Journal of Cell Biology, 159(2): 303–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/2/303.

Homburg:2000:FSI

[HVM^{+}00] S. Homburg, L. Visochek, N. Moran, F. Dantzer, E. Priel, E. Asculai, D. Schwartz, V. Rotter, N. Dekel, and M. Cohen-Armon. A Fast Signal–Induced Activation of Poly(Adp– Ribose) Polymerase. Journal of Cell Biology, 150(2):293–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
150/2/293.

Hikita:2000:ITD

[HVT^{+}00] Chinami Hikita, Soundarapandian Vijayakumar, Jiro Takito, Hediyet Erdjument-Bromage, Paul Tempst, and Qais Al-Awqati. Induction of Terminal Differentiation in Epithelial Cells Requires Polymerization of Hensin by Galectin 3. Journal of Cell Biology, 151(6):1235–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/6/1235.

Heiman:2000:PPR

[HW00a] Maxwell G. Heiman and Peter Walter. Prm1p, a Pheromone– Regulated Multispanning Membrane Protein, Facilitates Plasma Membrane Fusion during Yeast Mating. Journal of Cell Biology, 151(3):719–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/3/719.

Hunter:2000:REN

[HW00b] Christine Hunter and Eric Wieschaus. Regulated Expression of nullo Is Required for the Formation of Distinct Apical and Basal Adherens Junctions in the Drosophila Blastoderm. Journal of Cell Biology, 150(2):391–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/2/391.

Hashemolhosseini:2004:INT

[HW04] Said Hashemolhosseini and Michael Wegner. Impacts of a new transcription factor family. Journal of Cell Biology, 166(6):765–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/6/765.

Huang:2001:CSS

[HWBD^{+}01] Mingdong Huang, Jacques T. Weissman, Sophie Béraud-Dufour, Peng Luan, Chenqian Wang, Wei Chen, Meir Aridor, Ian A. Wilson, and William E. Balch. Crystal structure of Sar1-GDP at 1.7 å resolution and the role of the NH_{2} terminus in ER export. Journal of Cell Biology, 155(6): 937–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/6/937.

Haddad:2001:DGE

[HWHH01] Elias K. Haddad, Xufeng Wu, John A. Hammer, and Pierre A. Henkart. Defective Granule Exocytosis in Rab27a– Deficient Lymphocytes from Ashen Mice. Journal of Cell Biology, 152(4):835–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/4/835.

Heath:2001:ARS

[HWW01] Colin M. Heath, Miriam Windsor, and Thomas Wileman. Aggresomes Resemble Sites Specialized for Virus Assembly. Journal of Cell Biology, 153(3):449–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/3/449.

Hosaka:2002:SDS

[HYMS^{+}02] Yukio Hosaka, Toshifumi Yokota, Yuko Miyagoe-Suzuki, Katsutoshi Yuasa, Michihiro Imamura, Ryoichi Matsuda, Takaaki Ikemoto, Shuhei Kameya, and Shin’ichi Takeda. α1-syntrophin–deficient skeletal muscle exhibits hypertrophy and aberrant formation of neuromuscular junctions during regeneration. Journal of Cell Biology, 158(6):1097–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
158/6/1097.

Hynes:2000:ECA

[HZ00] Richard O. Hynes and Qi Zhao. The Evolution of Cell Adhesion. Journal of Cell Biology, 150(2):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/2/F89.

Hatch:2001:MLC

[HZS^{+}01] Victoria Hatch, Gang Zhi, Lula Smith, James T. Stull, Roger Craig, and William Lehman. Myosin light chain kinase binding to a unique site on F-actin revealed by threedimensional image reconstruction. Journal of Cell Biology, 154(3):611–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/3/611.

Iba:2000:CRD

[IAG^{+}00] Kousuke Iba, Reidar Albrechtsen, Brent Gilpin, Camilla Fröhlich, Frosty Loechel, Anna Zolkiewska, Kazuhiro Ishiguro, Tetsuhito Kojima, Wei Liu, J. Kevin Langford, Ralph D. Sanderson, Cord Brakebusch, Reinhard Fässler, and Ulla M. Wewer. The Cysteine–Rich Domain of Human Adam 12 Supports Cell Adhesion through Syndecans and Triggers Signaling Events That Lead to β1 Integrin–Dependent Cell Spreading. Journal of Cell Biology, 149(5):1143–??, May 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/149/5/
1143.

Iomini:2001:PPC

[IBKSP01] Carlo Iomini, Veronica Babaev-Khaimov, Massimo Sassaroli, and Gianni Piperno. Protein Particles in Chlamydomonas Flagella Undergo a Transport Cycle Consisting of Four Phases. Journal of Cell Biology, 153(1):13–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/153/1/

Imanishi:2004:NTP

[IBP^{+}04] Yoshikazu Imanishi, Matthew L. Batten, David W. Piston, Wolfgang Baehr, and Krzysztof Palczewski. Noninvasive twophoton imaging reveals retinyl ester storage structures in the eye. Journal of Cell Biology, 164(3):373–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
3/373.

Inbal:2002:DKD

[IBS^{+}02] Boaz Inbal, Shani Bialik, Ilana Sabanay, Gidi Shani, and Adi Kimchi. DAP kinase and DRP-1 mediate membrane blebbing and the formation of autophagic vesicles during programmed cell death. Journal of Cell Biology, 157(3):455–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
157/3/455.

Itoh:2004:BDM

[ICK^{+}04] Kyoko Itoh, Ling Cheng, Yoshimasa Kamei, Shinji Fushiki, Hiroyuki Kamiguchi, Paul Gutwein, Alexander Stoeck, Bernd Arnold, Peter Altevogt, and Vance Lemmon. Brain development in mice lacking L1–L1 homophilic adhesion. Journal of Cell Biology, 165(1):145–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/1/145.

Ishikawa:2003:IOS

[ICN^{+}03] Kuniko Ishikawa, Natalie L. Catlett, Jennifer L. Novak, Fusheng Tang, Johnathan J. Nau, and Lois S. Weisman. Identification of an organelle-specific myosin V receptor. Journal of Cell Biology, 160(6):887–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/6/887.

Inoue:2000:ONM

[IdCAS^{+}00] Yoshihiro H. Inoue, Maria do Carmo Avides, Michina Shiraki, Peter Deak, Masamitsu Yamaguchi, Yoshio Nishimoto, Akio Matsukage, and David M. Glover. Orbit, a Novel Microtubule–Associated Protein Essential for Mitosis in Drosophila melanogaster. Journal of Cell Biology, 149 (1):153–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/1/153.

Ireton:2002:NRP

[IDvH^{+}02] Reneé C. Ireton, Michael A. Davis, Jolanda van Hengel, Deborah J. Mariner, Kirk Barnes, Molly A. Thoreson, Panos Z. Anastasiadis, Linsey Matrisian, Linda M. Bundy, Linda Sealy, Barbara Gilbert, Frans van Roy, and Albert B. Reynolds. A novel role for p120 catenin in E-cadherin function. Journal of Cell Biology, 159(3):465–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/3/
465.

Imbert:2001:NCI

[IEJ^{+}01] Alexandra Imbert, Rachel Eelkema, Sara Jordan, Helen Feiner, and Pamela Cowin. Δn89β-catenin Induces Precocious Development, Differentiation, and Neoplasia in Mammary Gland. Journal of Cell Biology, 153(3):555–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
3/555.

Innocenti:2003:PKA

[IFP^{+}03] Metello Innocenti, Emanuela Frittoli, Isabella Ponzanelli, John R. Falck, Saskia M. Brachmann, Pier Paolo Di Fiore, and Giorgio Scita. Phosphoinositide 3-kinase activates Rac by entering in a complex with Eps8, Abi1, and Sos-1. Journal of Cell Biology, 160(1):17–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/1/17.

Imanishi:2004:R

[IGP04] Yoshikazu Imanishi, Volker Gerke, and Krzysztof Palczewski. Retinosomes. Journal of Cell Biology, 166(4):447–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
4/447.

Hanai:2002:EPI

[iHGK^{+}02] Jun ichi Hanai, Joachim Gloy, S. Ananth Karumanchi, Sujata Kale, Jian Tang, Guang Hu, Barden Chan, Ramani Ramchandran, Vivek Jha, Vikas P. Sukhatme, and Sergei Sokol.

Endostatin is a potential inhibitor of Wnt signaling. Journal of Cell Biology, 158(3):529–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/3/529.

Iwamoto:2000:TFE

[IHK^{+}00] Masahiro Iwamoto, Yoshinobu Higuchi, Eiki Koyama, Motomi Enomoto-Iwamoto, Kojiro Kurisu, Helena Yeh, William R. Abrams, Joel Rosenbloom, and Maurizio Pacifici. Transcription Factor Erg Variants and Functional Diversification of Chondrocytes during Limb Long Bone Development. Journal of Cell Biology, 150(1):27–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/1/
27.

Im:2001:IMT

[IHN^{+}01] Dong-Soon Im, Christopher E. Heise, Tuan Nguyen, Brian F. O’Dowd, and Kevin R. Lynch. Identification of a Molecular Target of Psychosine and Its Role in Globoid Cell Formation. Journal of Cell Biology, 153(2):429–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/2/429.

Inada:2001:KAT

[IIN^{+}01] Hiroyasu Inada, Ichiro Izawa, Miwako Nishizawa, Eriko Fujita, Tohru Kiyono, Toshitada Takahashi, Takashi Momoi, and Masaki Inagaki. Keratin attenuates tumor necrosis factor–induced cytotoxicity through association with TRADD. Journal of Cell Biology, 155(3):415–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
3/415.

Iwata:2003:NMM

[IKA^{+}03] Yuko Iwata, Yuki Katanosaka, Yuji Arai, Kazuo Komamura, Kunio Miyatake, and Munekazu Shigekawa. A novel mechanism of myocyte degeneration involving the Ca^{2+}-permeable growth factor–regulated channel. Journal of Cell Biology, 161 (5):957–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/5/957.

Kitajiri:2004:RDC

[iKFH^{+}04] Shin ichiro Kitajiri, Kanehisa Fukumoto, Masaki Hata, Hiroyuki Sasaki, Tatsuya Katsuno, Takayuki Nakagawa, Juichi Ito, Shoichiro Tsukita, and Sachiko Tsukita. Radixin deficiency causes deafness associated with progressive degeneration of cochlear stereocilia. Journal of Cell Biology, 166(4): 559–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/4/559.

Ito:2001:DTJ

[IKS^{+}01] Koichi Ito, Shinji Komazaki, Kazushige Sasamoto, Morikatsu Yoshida, Miyuki Nishi, Kenji Kitamura, and Hiroshi Takeshima. Deficiency of triad junction and contraction in mutant skeletal muscle lacking junctophilin type 1. Journal of Cell Biology, 154(5):1059–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/5/1059.

Iouk:2002:YNP

[IKS^{+}02] Tatiana Iouk, Oliver Kerscher, Robert J. Scott, Munira A. Basrai, and Richard W. Wozniak. The yeast nuclear pore complex functionally interacts with components of the spindle assembly checkpoint. Journal of Cell Biology, 159(5): 807–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/5/807.

Nishikawa:2001:MCY

[iNFK^{+}01] Shuh ichi Nishikawa, Sheara W. Fewell, Yoshihito Kato, Jeffrey L. Brodsky, and Toshiya Endo. Molecular Chaperones in the Yeast Endoplasmic Reticulum Maintain the Solubility of Proteins for Retrotranslocation and Degradation. Journal of Cell Biology, 153(5):1061–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/5/1061.

Izumi:2004:DFG

[IOIF^{+}04] Yasushi Izumi, Nao Ohta, Asako Itoh-Furuya, Naoyuki Fuse, and Fumio Matsuzaki. Differential functions of G protein and Baz–aPKC signaling pathways in Drosophila neuroblast asymmetric division. Journal of Cell Biology, 164(5):

729–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/5/729.

Ishizuya-Oka:2000:RMM

[IOLA^{+}00] Atsuko Ishizuya-Oka, Qing Li, Tosikazu Amano, Sashko Damjanovski, Shuichi Ueda, and Yun-Bo Shi. Requirement for Matrix Metalloproteinase Stromelysin-3 in Cell Migration and Apoptosis during Tissue Remodeling in Xenopus laevis. Journal of Cell Biology, 150(5):1177–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
5/1177.

Ohzeki:2002:CBB

[iONOM02] Jun ichirou Ohzeki, Megumi Nakano, Teruaki Okada, and Hiroshi Masumoto. CENP-b box is required for de novo centromere chromatin assembly on human alphoid DNA. Journal of Cell Biology, 159(5):765–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/5/765.

Izumi:2004:ECR

[ISB^{+}04] Genkichi Izumi, Toshiaki Sakisaka, Takeshi Baba, Shintaro Tanaka, Koji Morimoto, and Yoshimi Takai. Endocytosis of E-cadherin regulated by Rac and Cdc42 small G proteins through IQGAP1 and actin filaments. Journal of Cell Biology, 166(2):237–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/2/237.

Itoh:2001:JAM

[ISF^{+}01] Masahiko Itoh, Hiroyuki Sasaki, Mikio Furuse, Harunobu Ozaki, Toru Kita, and Shoichiro Tsukita. Junctional adhesion molecule (JAM) binds to PAR-3. Journal of Cell Biology, 154(3):491–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/3/491.

Inoue:2003:IMO

[ISID^{+}03] Osamu Inoue, Katsue Suzuki-Inoue, William L. Dean, Jon Frampton, and Steve P. Watson. Integrin α_{2}β_{1} mediates outside-in regulation of platelet spreading on collagen through activation of Src kinases and PLCγ2. Journal of Cell Biology, 160(5):769–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/5/769.

Inoue:2004:MOM

[ISS^{+}04] Yoshihiro H. Inoue, Matthew S. Savoian, Takao Suzuki, Endre Máthé, Masa-Toshi Yamamoto, and David M. Glover. Mutations in orbit/mast reveal that the central spindle is comprised of two microtubule populations, those that initiate cleavage and those that propagate furrow ingression. Journal of Cell Biology, 166(1):49–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/1/49.

Innocenti:2002:MTW

[ITF^{+}02] Metello Innocenti, Pierluigi Tenca, Emanuela Frittoli, Mario Faretta, Arianna Tocchetti, Pier Paolo Di Fiore, and Giorgio Scita. Mechanisms through which Sos-1 coordinates the activation of Ras and Rac. Journal of Cell Biology, 156(1): 125–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/1/125.

Iomini:2004:PCH

[ITM^{+}04] Carlo Iomini, Karla Tejada, Wenjun Mo, Heikki Vaananen, and Gianni Piperno. Primary cilia of human endothelial cells disassemble under laminar shear stress. Journal of Cell Biology, 164(6):811–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/6/811.

Iioka:2004:ERM

[IUK04] Hidekazu Iioka, Naoto Ueno, and Noriyuki Kinoshita. Essential role of MARCKS in cortical actin dynamics during gastrulation movements. Journal of Cell Biology, 164(2): 169–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/2/169.

Impagnatiello:2001:MSM

[IWG^{+}01] Maria-Antonietta Impagnatiello, Stefan Weitzer, Grainne Gannon, Amelia Compagni, Matt Cotten, and Gerhard Christofori. Mammalian Sprouty-1 and -2 are membraneanchored phosphoprotein inhibitors of growth factor signaling in endothelial cells. Journal of Cell Biology, 152(5): 1087–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/5/1087.

Ishikawa:2004:TNP

[IYT^{+}04] Daigo Ishikawa, Hayashi Yamamoto, Yasushi Tamura, Kaori Moritoh, and Toshiya Endo. Two novel proteins in the mitochondrial outer membrane mediate β-barrel protein assembly. Journal of Cell Biology, 166(5):621–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/166/5/
621.

Jaiswal:2002:MPL

[JAS02] Jyoti K. Jaiswal, Norma W. Andrews, and Sanford M. Simon. Membrane proximal lysosomes are the major vesicles responsible for calcium-dependent exocytosis in nonsecretory cells. Journal of Cell Biology, 159(4):625–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/4/
625.

Jay:2001:SAR

[Jay01] Daniel G. Jay. A Src-astic response to mounting tension. Journal of Cell Biology, 155(3):327–??, October 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/155/3/
327.

Jenkins:2001:AGC

[JB01] Scott M. Jenkins and Vann Bennett. Ankyrin-g coordinates assembly of the spectrin-based membrane skeleton, voltage-gated sodium channels, and L1 CAMs at Purkinje neuron initial segments. Journal of Cell Biology, 155(5): 739–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/5/739.

Jimenez:2001:RCI

[JBA^{+}01] Maria J. G. Jiménez, Milagros Balbín, Jesús Alvarez, Toshihisa Komori, Paolo Bianco, Kenn Holmbeck, Henning Birkedal-Hansen, José M. López, and Carlos López-Otín. A regulatory cascade involving retinoic acid, Cbfa1, and matrix metalloproteinases is coupled to the development of a process of perichondrial invasion and osteogenic differentiation during bone formation. Journal of Cell Biology, 155(7): 1333–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/7/1333.

Jady:2004:HTR

[JBK04] Beáta E. Jády, Edouard Bertrand, and Tamás Kiss. Human telomerase RNA and box H/ACA scaRNAs share a common Cajal body–specific localization signal. Journal of Cell Biology, 164(5):647–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/5/647.

Jackson:2000:AAG

[JBN^{+}00] Trevor R. Jackson, Fraser D. Brown, Zhongzhen Nie, Koichi Miura, Letizia Foroni, Jianlan Sun, Victor W. Hsu, Julie G. Donaldson, and Paul A. Randazzo. Acaps Are Arf6 Gtpase– Activating Proteins That Function in the Cell Periphery. Journal of Cell Biology, 151(3):627–??, October 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/3/
627.

Johnson:2001:ERS

[JC01] Jill L. Johnson and Elizabeth A. Craig. An Essential Role for the Substrate–Binding Region of Hsp40s in Saccharomyces cerevisiae. Journal of Cell Biology, 152(4):851–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
4/851.

Jokitalo:2001:GCV

[JCPWS01] Eija Jokitalo, Noemi Cabrera-Poch, Graham Warren, and David T. Shima. Golgi clusters and vesicles mediate mitotic inheritance independently of the endoplasmic reticulum. Journal of Cell Biology, 154(2):317–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/2/317.

Jacobson:2001:DCN

[JCR^{+}01] Christian Jacobson, Patrice D. Côté, Susana G. Rossi, Richard L. Rotundo, and Salvatore Carbonetto. The Dystroglycan Complex Is Necessary for Stabilization of Acetylcholine Receptor Clusters at Neuromuscular Junctions and Formation of the Synaptic Basement Membrane. Journal of Cell Biology, 152(3):435–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/3/435.

Janson:2003:DIM

[JdDD03] Marcel E. Janson, Mathilde E. de Dood, and Marileen Dogterom. Dynamic instability of microtubules is regulated by force. Journal of Cell Biology, 161(6):1029–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/6/
1029.

Jegalian:2001:BCN

[Jeg01a] Karin Jegalian. Bringing channels into the node. Journal of Cell Biology, 154(2):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/2/253.3.

Jegalian:2001:DLB

[Jeg01b] Karin Jegalian. A dystroglycan ligand in the brain. Journal of Cell Biology, 154(2):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/2/252.1.

Jegalian:2001:MPD

[Jeg01c] Karin Jegalian. Matrix patterning of digits. Journal of Cell Biology, 154(2):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/2/253.1.

Jegalian:2001:CCB

[Jeg01d] Karin Jegalian. Of coilin and Cajal bodies. Journal of Cell Biology, 154(2):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/2/252.2.

Jegalian:2001:TTE

[Jeg01e] Karin Jegalian. Tenascin turns on the EGFR. Journal of Cell Biology, 154(2):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/2/253.2.

Jackson:2003:CPI

[JFS^{+}03] Antony P. Jackson, Alexander Flett, Carl Smythe, Lindsay Hufton, Frank R. Wettey, and Elizabeth Smythe. Clathrin promotes incorporation of cargo into coated pits by activation of the AP2 adaptor μ2 kinase. Journal of Cell Biology, 163(2):231–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/2/231.

Jurczyk:2004:PFC

[JGR^{+}04] Agata Jurczyk, Adam Gromley, Sambra Redick, Jovenal San Agustin, George Witman, Gregory J. Pazour, Dorien J. M. Peters, and Stephen Doxsey. Pericentrin forms a complex with intraflagellar transport proteins and polycystin-2 and is required for primary cilia assembly. Journal of Cell Biology, 166(5):637–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/5/637.

Jaspersen:2002:MNC

[JGW02] Sue L. Jaspersen, Thomas H. Giddings, and Mark Winey. Mps3p is a novel component of the yeast spindle pole body that interacts with the yeast centrin homologue Cdc31p. Journal of Cell Biology, 159(6):945–??, December 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/6/
945.

Jaeschke:2002:TSC

[JHS^{+}02] Anja Jaeschke, Joerg Hartkamp, Masao Saitoh, Wendy Roworth, Takahiro Nobukuni, Angela Hodges, Julian Sampson, George Thomas, and Richard Lamb. Tuberous sclerosis complex tumor suppressor–mediated S6 kinase inhibition by phosphatidylinositide-3-OH kinase is mTOR independent. Journal of Cell Biology, 159(2):217–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
2/217.

Jospin:2002:TVD

[JJM^{+}02] Maëlle Jospin, Vincent Jacquemond, Marie-Christine Mariol, Laurent Ségalat, and Bruno Allard. The L-type voltagedependent Ca^{2+} channel EGL-19 controls body wall muscle function in Caenorhabditis elegans. Journal of Cell Biology, 159(2):337–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/2/337.

Jarousse:2001:ABS

[JK01] Nadine Jarousse and Regis B. Kelly. The AP2 binding site of synaptotagmin 1 is not an internalization signal but a regulator of endocytosis. Journal of Cell Biology, 154(4): 857–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/4/857.

Janvier:2003:RDB

[JKB^{+}03] Katy Janvier, Yukio Kato, Markus Boehm, Jeremy R. Rose, José A. Martina, Bong-Yoon Kim, Sundararajan Venkatesan, and Juan S. Bonifacino. Recognition of dileucine-based sorting signals from HIV-1 Nef and LIMP–II by the AP-1 γ–σ1 and AP-3 δ–σ3 hemicomplexes. Journal of Cell Biology, 163(6):1281–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/6/1281.

Jolly:2002:VBA

[JKG^{+}02] Caroline Jolly, Lara Konecny, Deborah L. Grady, Yulia A. Kutskova, José J. Cotto, Richard I. Morimoto, and Claire Vourc’h. In vivo binding of active heat shock transcription factor 1 to human chromosome 9 heterochromatin during stress. Journal of Cell Biology, 156(5):775–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/156/
5/775.

Jaquemar:2003:KPP

[JKW^{+}03] Daniel Jaquemar, Sergey Kupriyanov, Miriam Wankell, Jacqueline Avis, Kurt Benirschke, Hélène Baribault, and Robert G. Oshima. Keratin 8 protection of placental barrier function. Journal of Cell Biology, 161(4):749–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
4/749.

Jordan:2003:RSM

[JLJD03] Theresa Jordan, Jinyuan Li, Hongbin Jiang, and Joseph X. DiMario. Repression of slow myosin heavy chain 2 gene expression in fast skeletal muscle fibers by muscarinic acetylcholine receptor and Gαq signaling. Journal of Cell Biology, 162(5):843–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/5/843.

Janda:2002:RTC

[JLK^{+}02] Elzbieta Janda, Kerstin Lehmann, Iris Killisch, Martin Jechlinger, Michaela Herzig, Julian Downward, Hartmut Beug, and Stefan Grünert. Ras and TGFβ cooperatively regulate epithelial cell plasticity and metastasis. Journal of Cell Biology, 156(2):299–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/2/299.

Johns:2001:RSG

[JLS^{+}01] Laura M. Johns, Edwin S. Levitan, Eric A. Shelden, Ronald W. Holz, and Daniel Axelrod. Restriction of Secretory Granule Motion near the Plasma Membrane of Chromaffin Cells. Journal of Cell Biology, 153(1):177–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
1/177.

Jouhet:2004:PDI

[JMB^{+}04] Juliette Jouhet, Eric Maréchal, Barbara Baldan, Richard Bligny, Jacques Joyard, and Maryse A. Block. Phosphate deprivation induces transfer of DGDG galactolipid from chloroplast to mitochondria. Journal of Cell Biology, 167(5):863– ??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/5/863.

Jones:2001:MDT

[JMG01] Jacob M. Jones, James C. Morrell, and Stephen J. Gould. Multiple Distinct Targeting Signals in Integral Peroxisomal Membrane Proteins. Journal of Cell Biology, 153(6):1141–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/6/1141.

Jolly:2004:SIT

[JMG^{+}04a] Caroline Jolly, Alexandra Metz, Jérôme Govin, Marc Vigneron, Bryan M. Turner, Saadi Khochbin, and Claire Vourc’h. Stress-induced transcription of satellite III repeats. Journal of Cell Biology, 164(1):25–??, January 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/1/
25.

Jones:2004:PPC

[JMG04b] Jacob M. Jones, James C. Morrell, and Stephen J. Gould. PEX19 is a predominantly cytosolic chaperone and import receptor for class 1 peroxisomal membrane proteins. Journal of Cell Biology, 164(1):57–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/1/57.

Ji:2002:PCR

[JOF^{+}02] Guangju Ji, Christopher D. O’Brien, Morris Feldman, Yefim Manevich, Poay Lim, Jing Sun, Steven M. Albelda, and Michael I. Kotlikoff. PECAM-1 (CD31) regulates a hydrogen peroxide–activated nonselective cation channel in endothelial cells. Journal of Cell Biology, 157(1):173–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/157/1/
173.

Jousse:2003:ICT

[JON^{+}03] Céline Jousse, Seiichi Oyadomari, Isabel Novoa, Phoebe Lu, Yuhong Zhang, Heather P. Harding, and David Ron. Inhibition of a constitutive translation initiation factor 2α phosphatase, CReP, promotes survival of stressed cells. Journal of Cell Biology, 163(4):767–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/4/767.

Jantsch-Plunger:2000:C

[JPGR^{+}00] Verena Jantsch-Plunger, Pierre Gönczy, Alper Romano, Heinke Schnabel, Danielle Hamill, Ralf Schnabel, Anthony A. Hyman, and Michael Glotzer. Cyk-4. Journal of Cell Biology, 149(7):1391–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/7/1391.

Jiang:2001:PSV

[JPH^{+}01] Liwen Jiang, Thomas E. Phillips, Christopher A. Hamm, Yolanda M. Drozdowicz, Philip A. Rea, Masayoshi Maeshima, Sally W. Rogers, and John C. Rogers. The protein storage vacuole. Journal of Cell Biology, 155(6):991–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
6/991.

Jimenez:2000:RPR

[JPM^{+}00a] Concepción Jiménez, Rosario Armas Portela, Mario Mellado, Jose Miguel Rodríguez-Frade, John Collard, Antonio Serrano, Carlos Martínez-A, Jesus Avila, and Ana C. Carrera. Role of the Pi3k Regulatory Subunit in the Control of Actin Organization and Cell Migration. Journal of Cell Biology, 151(2):249–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/2/249.

Johnson:2000:TCR

[JPM^{+}00b] Carol Johnson, Dragan Primorac, Monique McKinstry, John McNeil, David Rowe, and Jeanne Bentley Lawrence. Tracking Col1a1 RNA in Osteogenesis Imperfecta. Journal of Cell Biology, 150(3):417–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/3/417.

Jiang:2000:BPS

[JPRR00] Liwen Jiang, Thomas E. Phillips, Sally W. Rogers, and John C. Rogers. Biogenesis of the Protein Storage Vacuole Crystalloid. Journal of Cell Biology, 150(4):755–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
4/755.

Joiner:2002:STE

[JR02] Keith A. Joiner and David S. Roos. Secretory traffic in the eukaryotic parasite Toxoplasma gondii. Journal of Cell Biology, 157(4):557–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/4/557.

Jacquet:2003:ONS

[JRL^{+}03] Michel Jacquet, Georges Renault, Sylvie Lallet, Jan De Mey, and Albert Goldbeter. Oscillatory nucleocytoplasmic shuttling of the general stress response transcriptional activators Msn2 and Msn4 in Saccharomyces cerevisiae. Journal of Cell Biology, 161(3):497–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/3/497.

Jung:2001:DCP

[JRW^{+}01] Goeh Jung, Kirsten Remmert, Xufeng Wu, Joanne M. Volosky, and John A. Hammer. The Dictyostelium carmil protein links capping protein and the Arp2/3 complex to type i myosins through their Sh3 domains. Journal of Cell Biology, 153(7):1479–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/7/1479.

Jensen:2001:NRB

[JSCR01] Sanne Jensen, Marisa Segal, Duncan J. Clarke, and Steven I. Reed. A Novel Role of the Budding Yeast Separin Esp1 in Anaphase Spindle Elongation. Journal of Cell Biology, 152 (1):27–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/1/27.

Joseph:2002:STR

[JTK^{+}02] Jomon Joseph, Shyh-Han Tan, Tatiana S. Karpova, James G. McNally, and Mary Dasso. SUMO-1 targets RanGAP1 to kinetochores and mitotic spindles. Journal of Cell Biology, 156(4):595–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/4/595.

Janes:2004:SVV

[JW04] Sam M. Janes and Fiona M. Watt. Switch from αvβ5 to αvβ6 integrin expression protects squamous cell carcinomas from anoikis. Journal of Cell Biology, 166(3):419–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
3/419.

Jin:2000:JCK

[JWJJ00] Ye Jin, Yanming Wang, Jørgen Johansen, and Kristen M. Johansen. Jil-1, a Chromosomal Kinase Implicated in Regulation of Chromatin Structure, Associates with the Male Specific Lethal (Msl) Dosage Compensation Complex. Journal of Cell Biology, 149(5):1005–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/149/5/1005.

Jiang:2002:ARP

[JZ02] Zhe Jiang and Eldad Zacksenhaus. Activation of retinoblastoma protein in mammary gland leads to ductal growth suppression, precocious differentiation, and adenocarcinoma. Journal of Cell Biology, 156(1):185–??, January 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/1/
185.

Karbowski:2004:QMD

[KAC^{+}04] Mariusz Karbowski, Damien Arnoult, Hsiuchen Chen, David C. Chan, Carolyn L. Smith, and Richard J. Youle. Quantitation of mitochondrial dynamics by photolabeling of individual organelles shows that mitochondrial fusion is blocked during the Bax activation phase of apoptosis. Journal of Cell Biology, 164(4):493–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/4/493.

Komarova:2002:CLP

[KAiK^{+}02] Yulia A. Komarova, Anna S. Akhmanova, Shin ichiro Kojima, Niels Galjart, and Gary G. Borisy. Cytoplasmic linker proteins promote microtubule rescue in vivo. Journal of Cell Biology, 159(4):589–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/4/589.

Krylyshkina:2003:NTM

[KAK^{+}03] Olga Krylyshkina, Kurt I. Anderson, Irina Kaverina, Irene Upmann, Dietmar J. Manstein, J. Victor Small, and Derek K. Toomre. Nanometer targeting of microtubules to focal adhesions. Journal of Cell Biology, 161(5):853–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/5/
853.

Kelly:2003:ANP

[KB03] Jeffery W. Kelly and William E. Balch. Amyloid as a natural product. Journal of Cell Biology, 161(3):461–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
3/461.

Kozlov:2004:PCF

[KB04] Michael M. Kozlov and Alexander D. Bershadsky. Processive capping by formin suggests a force-driven mechanism of actin polymerization. Journal of Cell Biology, 167(6): 1011–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/6/1011.

Kendirgi:2003:ERH

[KBG^{+}03] Frederic Kendirgi, Dianne M. Barry, Eric R. Griffis, Maureen A. Powers, and Susan R. Wente. An essential role for hGle1 nucleocytoplasmic shuttling in mRNA export. Journal of Cell Biology, 160(7):1029–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/7/1029.

Klemba:2004:C

[KBGG04a] Michael Klemba, Wandy Beatty, Ilya Gluzman, and Daniel E. Goldberg. Correction. Journal of Cell Biology, 164(4):625–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/4/625.

Klemba:2004:TPI

[KBGG04b] Michael Klemba, Wandy Beatty, Ilya Gluzman, and Daniel E. Goldberg. Trafficking of plasmepsin II to the food vacuole of the malaria parasite Plasmodium falciparum. Journal of Cell Biology, 164(1):47–??, January 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/1/
47.

Kamsteeg:2003:RPD

[KBK^{+}03] Erik-Jan Kamsteeg, Daniel G. Bichet, Irene B. M. Konings, Hubert Nivet, Michelle Lonergan, Marie-Françoise Arthus, Carel H. van Os, and Peter M. T. Deen. Reversed polarized delivery of an aquaporin-2 mutant causes dominant nephrogenic diabetes insipidus. Journal of Cell Biology, 163(5): 1099–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/5/1099.

Kallio:2002:RMI

[KBWG02] Marko J. Kallio, Victoria A. Beardmore, Jasminder Weinstein, and Gary J. Gorbsky. Rapid microtubule-independent dynamics of Cdc20 at kinetochores and centrosomes in mammalian cells. Journal of Cell Biology, 158(5):841–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
158/5/841.

Kimura:2001:KCH

[KC01] Hiroshi Kimura and Peter R. Cook. Kinetics of Core Histones in Living Human Cells. Journal of Cell Biology, 153 (7):1341–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/7/1341.

Kapoor:2002:SMG

[KC02] Tarun M. Kapoor and Duane A. Compton. Searching for the middle ground. Journal of Cell Biology, 157(4):551–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/157/
4/551.

Khodjakov:2003:MEC

[KCG^{+}03] Alexey Khodjakov, Lily Copenagle, Michael B. Gordon, Duane A. Compton, and Tarun M. Kapoor. Minus-end capture of preformed kinetochore fibers contributes to spindle morphogenesis. Journal of Cell Biology, 160(5):671–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
5/671.

Kedersha:2000:DST

[KCL^{+}00] Nancy Kedersha, Michael R. Cho, Wei Li, Patrick W. Yacono, Samantha Chen, Natalie Gilks, David E. Golan, and Paul Anderson. Dynamic Shuttling of Tia-1 Accompanies the Recruitment of mRNA to Mammalian Stress Granules. Journal of Cell Biology, 151(6):1257–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/6/1257.

Kim:2001:NRS

[KCL^{+}01] Jin Woo Kim, Tong-Shin Chang, Ji Eun Lee, Sung-Ho Huh, Seung Woo Yeon, Wan Seok Yang, Cheol O. Joe, Inhee Mook-Jung, Rudolph E. Tanzi, Tae-Wan Kim, and Eui-Ju Choi. Negative regulation of the Sapk/Jnk signaling pathway by presenilin 1. Journal of Cell Biology, 153(3):457–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
3/457.

Kang:2002:CPC

[KCWF02] Dongmin Kang, James Chen, Jim Wong, and Guowei Fang. The checkpoint protein Chfr is a ligase that ubiquitinates Plk1 and inhibits Cdc2 at the G2 to M transition. Journal of Cell Biology, 156(2):249–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/2/249.

Kim:2004:PAC

[KCY^{+}04] Minsoo Kim, Christopher V. Carman, Wei Yang, Azucena Salas, and Timothy A. Springer. The primacy of affinity over clustering in regulation of adhesiveness of the integrin α_{L} β_{2}. Journal of Cell Biology, 167(6):1241–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
6/1241.

Kong:2004:RUP

[KDH^{+}04] Xiang-Tian Kong, Fang-Ming Deng, Ping Hu, Feng-Xia Liang, Ge Zhou, Anna B. Auerbach, Nancy Genieser, Peter K. Nelson, Edith S. Robbins, Ellen Shapiro, Bechara Kachar, and Tung-Tien Sun. Roles of uroplakins in plaque formation, umbrella cell enlargement, and urinary tract diseases. Journal of Cell Biology, 167(6):1195–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
6/1195.

Koch:2000:LLD

[KdVS^{+}00] Peter J. Koch, Pierre A. de Viragh, Elisabeth Scharer, Donnie Bundman, Mary Ann Longley, Jackie Bickenbach, Yasuhiro Kawachi, Yasushi Suga, Zhijian Zhou, Marcel Huber, Daniel Hohl, Tonja Kartasova, Michal Jarnik, Alasdair C. Steven, and Dennis R. Roop. Lessons from Loricrin–Deficient Mice. Journal of Cell Biology, 151(2):389–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/2/
389.

Kessels:2001:MAS

[KEGDQ01] Michael M. Kessels, Åsa E. Y. Engqvist-Goldstein, David G. Drubin, and Britta Qualmann. Mammalian Abp1, a Signal– Responsive F–Actin–Binding Protein, Links the Actin Cytoskeleton to Endocytosis via the Gtpase Dynamin. Journal of Cell Biology, 153(2):351–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/2/351.

Kanaani:2002:CTD

[KEHAM^{+}02] Jamil Kanaani, Alaa El-Din El-Husseini, Andrea Aguilera-Moreno, Julia M. Diacovo, David S. Bredt, and Steinunn Baekkeskov. A combination of three distinct trafficking signals mediates axonal targeting and presynaptic clustering of GAD65. Journal of Cell Biology, 158(7):1229–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
7/1229.

Kellum:2003:HRD

[Kel03] Rebecca Kellum. Is HP1 an RNA detector that functions both in repression and activation? Journal of Cell Biology, 161(4):671–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/4/671.

Kaufman:2000:GYC

[KF00] Charles K. Kaufman and Elaine Fuchs. It’s Got You Covered. Journal of Cell Biology, 149(5):999–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/5/999.

Ku:2004:RAD

[KFO04] Nam-On Ku, Haian Fu, and M. Bishr Omary. Raf-1 activation disrupts its binding to keratins during cell stress. Journal of Cell Biology, 166(4):479–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/4/479.

Kowalczyk:2004:CRC

[KFR^{+}04] Anna Kowalczyk, Robert K. Filipkowski, Marcin Rylski, Grzegorz M. Wilczynski, Filip A. Konopacki, Jacek Jaworski, Maria A. Ciemerych, Piotr Sicinski, and Leszek Kaczmarek. The critical role of cyclin D2 in adult neurogenesis. Journal of Cell Biology, 167(2):209–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/2/209.

Kamada:2000:TMI

[KFS^{+}00] Yoshiaki Kamada, Tomoko Funakoshi, Takahiro Shintani, Kazuya Nagano, Mariko Ohsumi, and Yoshinori Ohsumi. Tor-mediated induction of autophagy via an Apg1 protein kinase complex. Journal of Cell Biology, 150(6):1507–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/150/6/1507.

Koshikawa:2000:RCS

[KGC^{+}00] Naohiko Koshikawa, Gianluigi Giannelli, Vincenzo Cirulli, Kaoru Miyazaki, and Vito Quaranta. Role of Cell Surface Metalloprotease Mt1-Mmp in Epithelial Cell Migration over Laminin-5. Journal of Cell Biology, 148(3):615–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
3/615.

Kiehart:2000:MFC

[KGE^{+}00] Daniel P. Kiehart, Catherine G. Galbraith, Kevin A. Edwards, Wayne L. Rickoll, and Ruth A. Montague. Multiple Forces Contribute to Cell Sheet Morphogenesis for Dorsal Closure in Drosophila. Journal of Cell Biology, 149(2): 471–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/2/471.

Kuriyama:2002:CMK

[KGT^{+}02] Ryoko Kuriyama, Charles Gustus, Yasuhiko Terada, Yumi Uetake, and Jurgita Matuliene. CHO1, a mammalian kinesinlike protein, interacts with F-actin and is involved in the terminal phase of cytokinesis. Journal of Cell Biology, 156 (5):783–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/5/783.

Kulkarni:2000:RPR

[KGvdG^{+}00] Sarang V. Kulkarni, Gerald Gish, Peter van der Geer, Mark Henkemeyer, and Tony Pawson. Role of P120 Ras–Gap in Directed Cell Movement. Journal of Cell Biology, 149(2): 457–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/2/457.

Kjoller:2001:RMC

[KH01] Lars Kjøller and Alan Hall. Rac Mediates Cytoskeletal Rearrangements and Increased Cell Motility Induced by Urokinase–Type Plasminogen Activator Receptor Binding to Vitronectin. Journal of Cell Biology, 152(6):1145–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
6/1145.

Kaether:2004:LBS

[KH04] Christoph Kaether and Christian Haass. A lipid boundary separates APP and secretases and limits amyloid β-peptide generation. Journal of Cell Biology, 167(5):809–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
5/809.

Kenner:2004:MLJ

[KHB^{+}04] Lukas Kenner, Astrid Hoebertz, F. Timo Beil, Niamh Keon, Florian Karreth, Robert Eferl, Harald Scheuch, Agnieszka Szremska, Michael Amling, Marina Schorpp-Kistner, Peter Angel, and Erwin F. Wagner. Mice lacking JunB are osteopenic due to cell-autonomous osteoblast and osteoclast defects. Journal of Cell Biology, 164(4):613–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
4/613.

Kelso:2002:DKR

[KHC02] Reed J. Kelso, Andrew M. Hudson, and Lynn Cooley. Drosophila Kelch regulates actin organization via Src64dependent tyrosine phosphorylation. Journal of Cell Biology, 156(4):703–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/4/703.

Klingbeil:2001:TPI

[KHH^{+}01] Candice K. Klingbeil, Christof R. Hauck, Datsun A. Hsia, K. C. Jones, Shannon R. Reider, and David D. Schlaepfer. Targeting Pyk2 to β1–Integrin–Containing Focal Contacts Rescues Fibronectin–Stimulated Signaling and Haptotactic Motility Defects of Focal Adhesion Kinase–Null Cells. Journal of Cell Biology, 152(1):97–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/1/97.

Kim:2001:MRA

[KHK01] John Kim, Wei-Pang Huang, and Daniel J. Klionsky. Membrane Recruitment of Aut7p in the Autophagy and Cytoplasm to Vacuole Targeting Pathways Requires Aut1p, Aut2p, and the Autophagy Conjugation Complex. Journal of Cell Biology, 152(1):51–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/1/51.

Kofron:2002:PRM

[KHLW02] Matthew Kofron, Janet Heasman, Stephanie A. Lang, and Christopher C. Wylie. Plakoglobin is required for maintenance of the cortical actin skeleton in early Xenopus embryos and for cdc42-mediated wound healing. Journal of Cell Biology, 158(4):695–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/4/695.

King:2000:DTK

[KHN00] Jennifer M. King, Tom S. Hays, and R. Bruce Nicklas. Dynein Is a Transient Kinetochore Component Whose Binding Is Regulated by Microtubule Attachment, Not Tension. Journal of Cell Biology, 151(4):739–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/4/739.

Kamsteeg:2000:SLA

[KHvOD00] E. J. Kamsteeg, I. Heijnen, C. H. van Os, and P. M. T. Deen. The Subcellular Localization of an Aquaporin-2 Tetramer Depends on the Stoichiometry of Phosphorylated and Nonphosphorylated Monomers. Journal of Cell Biology, 151(4): 919–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/4/919.

Komatsu:2004:ZKR

[KI04] Satoshi Komatsu and Mitsuo Ikebe. ZIP kinase is responsible for the phosphorylation of myosin II and necessary for cell motility in mammalian fibroblasts. Journal of Cell Biology, 165(2):243–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/2/243.

Kajita:2001:MTM

[KIC^{+}01] Masahiro Kajita, Yoshifumi Itoh, Tadashige Chiba, Hidetoshi Mori, Akiko Okada, Hiroaki Kinoh, and Motoharu Seiki. Membrane-type 1 Matrix Metalloproteinase Cleaves Cd44 and Promotes Cell Migration. Journal of Cell Biology, 153 (5):893–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/5/893.

Kanaji:2000:CSD

[KIK^{+}00a] Sachiko Kanaji, Jun Iwahashi, Yuichiro Kida, Masao Sakaguchi, and Katsuyoshi Mihara. Characterization of the Signal That Directs Tom20 to the Mitochondrial Outer Membrane. Journal of Cell Biology, 151(2):277–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/2/

Kim:2000:CER

[KIK^{+}00b] Jae-Beom Kim, Shahidul Islam, Young J. Kim, Ryan S. Prudoff, Kristin M. Sass, Margaret J. Wheelock, and Keith R. Johnson. N-cadherin Extracellular Repeat 4 Mediates Epithelial to Mesenchymal Transition and Increased Motility. Journal of Cell Biology, 151(6):1193–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
6/1193.

Kilmartin:2003:SCC

[Kil03] John V. Kilmartin. Sfi1p has conserved centrin-binding sites and an essential function in budding yeast spindle pole body duplication. Journal of Cell Biology, 162(7):1211– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/7/1211.

Kirisako:2000:RMR

[KIO^{+}00] Takayoshi Kirisako, Yoshinobu Ichimura, Hisashi Okada, Yukiko Kabeya, Noboru Mizushima, Tamotsu Yoshimori, Mariko Ohsumi, Toshifumi Takao, Takeshi Noda, and Yoshinori Ohsumi. The reversible modification regulates the membrane-binding state of Apg8/Aut7 essential for autophagy and the cytoplasm to vacuole targeting pathway. Journal of Cell Biology, 151(2):263–??, October 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/2/
263.

Kruse:2002:RDL

[KJB^{+}02] Claudia Kruse, Andreas Jaedicke, Joël Beaudouin, Florian Böhl, Dunja Ferring, Thomas Güttler, Jan Ellenberg, and Ralf-Peter Jansen. Ribonucleoprotein-dependent localization of the yeast class V myosin Myo4p. Journal of Cell Biology, 159(6):971–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/6/971.

Kwon:2003:RPH

[KJK^{+}03] Yong-Kook Kwon, Hyeung-Jin Jang, Sutapa Kole, Hua-Jun He, and Michel Bernier. Role of the pleckstrin homology domain of PLCγ1 in its interaction with the insulin receptor. Journal of Cell Biology, 163(2):375–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/163/2/
375.

Karbowski:2004:EBR

[KJY04] Mariusz Karbowski, Seon-Yong Jeong, and Richard J. Youle. Endophilin B1 is required for the maintenance of mitochondrial morphology. Journal of Cell Biology, 166(7):1027– ??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/7/1027.

Kim:2002:SNS

[KK02] Kyoungtae Kim and Thomas C. S. Keller. Smitin, a novel smooth muscle titin-like protein, interacts with myosin filaments in vivo and in vitro. Journal of Cell Biology, 156(1): 101–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/1/101.

Katoh:2001:RKM

[KKA^{+}01] Kazuo Katoh, Yumiko Kano, Mutsuki Amano, Hirofumi Onishi, Kozo Kaibuchi, and Keigi Fujiwara. Rho-kinase– mediated contraction of isolated stress fibers. Journal of Cell Biology, 153(3):569–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/3/569.

Ko:2000:NLH

[KKK^{+}00] Young-Gyu Ko, Young-Sun Kang, Eun-Kyoung Kim, Sang Gyu Park, and Sunghoon Kim. Nucleolar Localization of Human Methionyl–Trna Synthetase and Its Role in Ribosomal RNA Synthesis. Journal of Cell Biology, 149 (3):567–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/3/567.

Krylyshkina:2002:MSA

[KKK^{+}02] Olga Krylyshkina, Irina Kaverina, Wolfgang Kranewitter, Walter Steffen, Maria C. Alonso, Robert A. Cross, and J. Victor Small. Modulation of substrate adhesion dynamics via microtubule targeting requires kinesin-1. Journal of Cell Biology, 156(2):349–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/2/349.

Klopfenstein:2001:SSL

[KKL^{+}01] Dieter R. Klopfenstein, Judith Klumperman, Ariel Lustig, Richard A. Kammerer, Viola Oorschot, and Hans-Peter Hauri. Subdomain-specific Localization of Climp-63 (P63) in the Endoplasmic Reticulum Is Mediated by Its Luminal α-Helical Segment. Journal of Cell Biology, 153(6):1287–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/6/1287.

Kaufman:2003:FMC

[KKPB03] Brett A. Kaufman, Jill E. Kolesar, Philip S. Perlman, and Ronald A. Butow. A function for the mitochondrial chaperonin Hsp60 in the structure and transmission of mitochondrial DNA nucleoids in Saccharomyces cerevisiae. Journal of Cell Biology, 163(3):457–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/3/457.

Kiessling:2000:VCL

[KKR^{+}00] Justine Kiessling, Sven Kruse, Stefan A. Rensing, Klaus Harter, Eva L. Decker, and Ralf Reski. Visualization of a Cytoskeleton-like Ftsz Network in Chloroplasts. Journal of Cell Biology, 151(4):945–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/4/945.

Kim:2001:CGF

[KKS^{+}01] John Kim, Yoshiaki Kamada, Per E. Stromhaug, Ju Guan, Ann Hefner-Gravink, Misuzu Baba, Sidney V. Scott, Yoshinori Ohsumi, William A. Dunn, and Daniel J. Klionsky. Cvt9/ gsa9 Functions in Sequestering Selective Cytosolic Cargo Destined for the Vacuole. Journal of Cell Biology, 153 (2):381–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/2/381.

Krementsov:2004:MV

[KKT04] Dimitry N. Krementsov, Elena B. Krementsova, and Kathleen M. Trybus. Myosin V. Journal of Cell Biology, 164 (6):877–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/6/877.

Kovar:2003:FYC

[KKTP03] David R. Kovar, Jeffrey R. Kuhn, Andrea L. Tichy, and Thomas D. Pollard. The fission yeast cytokinesis formin Cdc12p is a barbed end actin filament capping protein gated by profilin. Journal of Cell Biology, 161(5):875–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
5/875.

Krauss:2003:ASC

[KKW^{+}03] Michael Krauss, Masahiro Kinuta, Markus R. Wenk, Pietro De Camilli, Kohji Takei, and Volker Haucke. ARF6 stimulates clathrin/AP-2 recruitment to synaptic membranes by activating phosphatidylinositol phosphate kinase type Iγ. Journal of Cell Biology, 162(1):113–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/1/113.

Korhonen:2004:UPS

[KL04] Laura Korhonen and Dan Lindholm. The ubiquitin proteasome system in synaptic and axonal degeneration. Journal of Cell Biology, 165(1):27–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/1/27.

Kalai:2003:REP

[KLD^{+}03] Michael Kalai, Mohamed Lamkanfi, Geertrui Denecker, Michael Boogmans, Saskia Lippens, Ann Meeus, Wim Declercq, and Peter Vandenabeele. Regulation of the expression and processing of caspase-12. Journal of Cell Biology, 162(3): 457–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/3/457.

Kaether:2002:PAT

[KLE^{+}02] Christoph Kaether, Sven Lammich, Dieter Edbauer, Michaela Ertl, Jens Rietdorf, Anja Capell, Harald Steiner, and Christian Haass. Presenilin-1 affects trafficking and processing of βAPP and is targeted in a complex with nicastrin to the plasma membrane. Journal of Cell Biology, 158(3): 551–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/3/551.

Kruhlak:2000:RMA

[KLF^{+}00] Michael J. Kruhlak, Melody A. Lever, Wolfgang Fischle, Eric Verdin, David P. Bazett-Jones, and Michael J. Hendzel. Reduced Mobility of the Alternate Splicing Factor (Asf) through the Nucleoplasm and Steady State Speckle Compartments. Journal of Cell Biology, 150(1):41–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/1/
41.

Kodama:2004:CCT

[KLF04] Atsuko Kodama, Terry Lechler, and Elaine Fuchs. Coordinating cytoskeletal tracks to polarize cellular movements. Journal of Cell Biology, 167(2):203–??, October 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/167/2/
203.

Karbowski:2002:STA

[KLG^{+}02] Mariusz Karbowski, Yang-Ja Lee, Brigitte Gaume, Seon-Yong Jeong, Stephan Frank, Amotz Nechushtan, Ansgar Santel, Margaret Fuller, Carolyn L. Smith, and Richard J. Youle. Spatial and temporal association of Bax with mitochondrial fission sites, Drp1, and Mfn2 during apoptosis. Journal of Cell Biology, 159(6):931–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/6/931.

Kuo:2001:ODR

[KLGC^{+}01] Calvin J. Kuo, Kenneth R. LaMontagne, Guillermo Garcia-Cardeña, Brian D. Ackley, Daniel Kalman, Susan Park, Rolf Christofferson, Junne Kamihara, Yuan-Hua Ding, Kin-Ming Lo, Stephen Gillies, Judah Folkman, Richard C. Mulligan, and Kashi Javaherian. Oligomerization-dependent regulation of motility and morphogenesis by the collagen XVIII Nc1/endostatin domain. Journal of Cell Biology, 152(6): 1233–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/6/1233.

Kickhoefer:2001:TVA

[KLK^{+}01] Valerie A. Kickhoefer, Yie Liu, Lawrence B. Kong, Bryan E. Snow, Phoebe L. Stewart, Lea Harrington, and Leonard H. Rome. The telomerase/vault-associated protein Tep1 is required for vault RNA stability and its association with the vault particle. Journal of Cell Biology, 152(1):157–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
152/1/157.

Kireeva:2004:VEC

[KLK^{+}04] Natashe Kireeva, Margot Lakonishok, Igor Kireev, Tatsuya Hirano, and Andrew S. Belmont. Visualization of early chromosome condensation. Journal of Cell Biology, 166(6):775– ??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/6/775.

Kim:2003:CHA

[KLT^{+}03] Jin-Moon Kim, Honglin Liu, Mayuko Tazaki, Masao Nagata, and Fugaku Aoki. Changes in histone acetylation during mouse oocyte meiosis. Journal of Cell Biology, 162(1):37–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
162/1/37.

Kohler:2004:RRP

[KLZ04] Katja Köhler, Daniel Louvard, and Ahmed Zahraoui. Rab13 regulates PKA signaling during tight junction assembly. Journal of Cell Biology, 165(2):175–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/2/175.

Koch:2000:DTI

[KM00] Hans-Georg Koch and Matthias Müller. Dissecting the Translocase and Integrase Functions of the Escherichia coli Secyeg Translocon. Journal of Cell Biology, 150(3):689–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/150/3/689.

Kapoor:2001:ESB

[KM01] Tarun M. Kapoor and Timothy J. Mitchison. Eg5 is static in bipolar spindles relative to tubulin. Journal of Cell Biology, 154(6):1125–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/6/1125.

Keilhack:2001:NRR

[KMB^{+}01] Heike Keilhack, Marit Müller, Sylvia-Annette Böhmer, Carsten Frank, K. Michael Weidner, Walter Birchmeier, Tanja Ligensa, Alexander Berndt, Hartwig Kosmehl, Bernd Günther, Thomas Müller, Carmen Birchmeier, and Frank D. Böhmer. Negative Regulation of Ros Receptor Tyrosine Kinase Signaling. Journal of Cell Biology, 152(2):325–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
152/2/325.

Kapoor:2000:PSA

[KMCM00] Tarun M. Kapoor, Thomas U. Mayer, Margaret L. Coughlin, and Timothy J. Mitchison. Probing Spindle Assembly Mechanisms with Monastrol, a Small Molecule Inhibitor of the Mitotic Kinesin, Eg5. Journal of Cell Biology, 150(5): 975–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/5/975.

Kuznetsov:2001:CSS

[KMG^{+}01] Sergei A. Kuznetsov, Mahesh H. Mankani, Stan Gronthos, Kazuhito Satomura, Paolo Bianco, and Pamela Gehron Robey. Circulating Skeletal Stem Cells. Journal of Cell Biology, 153(5):1133–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/5/1133.

Kao:2003:HDI

[KMG^{+}03] Gary D. Kao, W. Gillies McKenna, Matthew G. Guenther, Ruth J. Muschel, Mitchell A. Lazar, and Tim J. Yen. Histone deacetylase 4 interacts with 53BP1 to mediate the DNA damage response. Journal of Cell Biology, 160(7):1017–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
7/1017.

Koch:2000:NMN

[KMH^{+}00] Manuel Koch, Julie R. Murrell, Dale D. Hunter, Pamela F. Olson, William Jin, Douglas R. Keene, William J. Brunken, and Robert E. Burgeson. A Novel Member of the Netrin Family, β-Netrin, Shares Homology with the β Chain of Laminin. Journal of Cell Biology, 151(2):221–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
2/221.

Kanzaki:2004:APK

[KMH^{+}04] Makoto Kanzaki, Silvia Mora, Joseph B. Hwang, Alan R. Saltiel, and Jeffrey E. Pessin. Atypical protein kinase C (PKCζ /λ) is a convergent downstream target of the insulinstimulated phosphatidylinositol 3-kinase and TC10 signaling pathways. Journal of Cell Biology, 164(2):279–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
2/279.

Kawasaki:2001:DAK

[KMiM^{+}01] Akira Kawasaki, Itaru Matsumura, Jun ichiro Miyagawa, Sachiko Ezoe, Hirokazu Tanaka, Yasuhiko Terada, Masaaki Tatsuka, Takashi Machii, Hiroshi Miyazaki, Yusuke Furukawa, and Yuzuru Kanakura. Downregulation of an Aim-1 Kinase Couples with Megakaryocytic Polyploidization of Human Hematopoietic Cells. Journal of Cell Biology, 152(2): 275–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/2/275.

Katsumi:2002:ECT

[KMK^{+}02] Akira Katsumi, Julie Milanini, William B. Kiosses, Miguel A. del Pozo, Roland Kaunas, Shu Chien, Klaus M. Hahn, and Martin Alexander Schwartz. Effects of cell tension on the small GTPase Rac. Journal of Cell Biology, 158(1):153–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
158/1/153.

Kummer:2004:NIF

[KMLS04] Terrance T. Kummer, Thomas Misgeld, Jeff W. Lichtman, and Joshua R. Sanes. Nerve-independent formation of a topologically complex postsynaptic apparatus. Journal of Cell Biology, 164(7):1077–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/7/1077.

Koberna:2002:RGF

[KMP^{+}02a] Karel Koberna, Jan Malínský, Artem Pliss, Martin Mašata, Jaromíra Večeřová, Markéta Fialová, Jan Bednár, and Ivan Raška. Ribosomal genes in focus. Journal of Cell Biology, 157 (5):743–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/5/743.

Kumaran:2002:LCS

[KMP02b] R. Ileng Kumaran, Bhattiprolu Muralikrishna, and Veena K. Parnaik. Lamin A/C speckles mediate spatial organization of splicing factor compartments and RNA polymerase II transcription. Journal of Cell Biology, 159(5):783–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
5/783.

Krylova:2000:DRM

[KMS00] Olga Krylova, Marcus J. Messenger, and Patricia C. Salinas. Dishevelled-1 Regulates Microtubule Stability. Journal of Cell Biology, 151(1):83–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/1/83.

Kusari:2004:CPI

[KMS^{+}04] Anasua B. Kusari, Douglas M. Molina, Walid Sabbagh, Chang S. Lau, and Lee Bardwell. A conserved protein interaction network involving the yeast MAP kinases Fus3 and Kss1. Journal of Cell Biology, 164(2):267–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/2/

Kasahara:2004:TLT

[KNI^{+}04] Kousuke Kasahara, Yuji Nakayama, Kikuko Ikeda, Yuka Fukushima, Daisuke Matsuda, Shinya Horimoto, and Naoto Yamaguchi. Trafficking of Lyn through the Golgi caveolin involves the charged residues on αE and αI helices in the kinase domain. Journal of Cell Biology, 165(5):641–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/165/
5/641.

Kihara:2001:TDV

[KNIO01] Akio Kihara, Takeshi Noda, Naotada Ishihara, and Yoshinori Ohsumi. Two Distinct Vps34 Phosphatidylinositol 3–Kinase Complexes Function in Autophagy and Carboxypeptidase Y Sorting in Saccharomyces cerevisiae. Journal of Cell Biology, 152(3):519–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/3/519.

Kulke:2001:KMS

[KNK^{+}01] Michael Kulke, Ciprian Neagoe, Bernhard Kolmerer, Ave Minajeva, Horst Hinssen, Belinda Bullard, and Wolfgang A. Linke. Kettin, a major source of myofibrillar stiffness in Drosophila indirect flight muscle. Journal of Cell Biology, 154(5):1045–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/5/1045.

Kanazawa:2000:ICL

[KNM^{+}00] Takayuki Kanazawa, Sachiko Nakamura, Michiko Momoi, Toshiyuki Yamaji, Hiromu Takematsu, Hajime Yano, Hisataka Sabe, Akitsugu Yamamoto, Toshisuke Kawasaki, and Yasunori Kozutsumi. Inhibition of Cytokinesis by a Lipid Metabolite, Psychosine. Journal of Cell Biology, 149(4): 943–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/4/943.

Kenworthy:2004:DPR

[KNR^{+}04] Anne K. Kenworthy, Benjamin J. Nichols, Catha L. Remmert, Glenn M. Hendrix, Mukesh Kumar, Joshua Zimmerberg, and Jennifer Lippincott-Schwartz. Dynamics of putative raft-associated proteins at the cell surface. Journal of Cell Biology, 165(5):735–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/5/735.

Ku:2000:KTU

[KO00] Nam-On Ku and M. Bishr Omary. Keratins Turn over by Ubiquitination in a Phosphorylation–Modulated Fashion. Journal of Cell Biology, 149(3):547–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/3/547.

Kimata:2004:RBA

[KOS^{+}04] Yukio Kimata, Daisuke Oikawa, Yusuke Shimizu, Yuki Ishiwata-Kimata, and Kenji Kohno. A role for BiP as an adjustor for the endoplasmic reticulum stress-sensing protein Ire1. Journal of Cell Biology, 167(3):445–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
3/445.

Kobielak:2003:DBF

[KPA^{+}03] Krzysztof Kobielak, H. Amalia Pasolli, Laura Alonso, Lisa Polak, and Elaine Fuchs. Defining BMP functions in the hair follicle by conditional ablation of BMP receptor IA. Journal of Cell Biology, 163(3):609–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/3/609.

Knaut:2000:ZVR

[KPB^{+}00] Holger Knaut, Francisco Pelegri, Kerstin Bohmann, Heinz Schwarz, and Christiane Nüsslein-Volhard. Zebrafish vasa RNA but Not Its Protein Is a Component of the Germ Plasm and Segregates Asymmetrically before Germline Specification. Journal of Cell Biology, 149(4):875–??, May 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/149/4/
875.

Kuja-Panula:2003:ATP

[KPKY^{+}03] Juha Kuja-Panula, Marjaana Kiiltomäki, Takashi Yamashiro, Ari Rouhiainen, and Heikki Rauvala. AMIGO, a transmembrane protein implicated in axon tract development, defines a novel protein family with leucine-rich repeats.

Journal of Cell Biology, 160(6):963–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/6/963.

Kato:2002:CID

[KPL^{+}02] Masaki Kato, Millan S. Patel, Regis Levasseur, Ivan Lobov, Benny H.-J. Chang, Donald A. Glass, Christine Hartmann, Lan Li, Tae-Ho Hwang, Cory F. Brayton, Richard A. Lang, Gerard Karsenty, and Lawrence Chan. Cbfa1-independent decrease in osteoblast proliferation, osteopenia, and persistent embryonic eye vascularization in mice deficient in Lrp5, a Wnt coreceptor. Journal of Cell Biology, 157(2):303–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
157/2/303.

Khodjakov:2001:CEF

[KR01] Alexey Khodjakov and Conly L. Rieder. Centrosomes Enhance the Fidelity of Cytokinesis in Vertebrates and Are Required for Cell Cycle Progression. Journal of Cell Biology, 153(1):237–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/1/237.

Kondylis:2003:NRD

[KR03] Vangelis Kondylis and Catherine Rabouille. A novel role for dp115 in the organization of tER sites in Drosophila. Journal of Cell Biology, 162(2):185–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/2/185.

Kramer:2000:UDL

[Krä00] Helmut Krämer. The Ups and Downs of Life in an Epithelium. Journal of Cell Biology, 151(4):0–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
4/F15.

Koivunen:2001:IIM

[KRA^{+}01] Erkki Koivunen, Tanja-Maria Ranta, Arto Annila, Seija Taube, Asko Uppala, Marjukka Jokinen, Gijsbert van Willigen, Eveliina Ihanus, and Carl G. Gahmberg. Inhibition of β_{2} integrin-mediated leukocyte cell adhesion by leucine–leucine– glycine motif-containing peptides. Journal of Cell Biology, 153(5):905–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/5/905.

Kad:2003:MHM

[KRF^{+}03] Neil M. Kad, Arthur S. Rovner, Patricia M. Fagnant, Peteranne B. Joel, Guy G. Kennedy, Joseph B. Patlak, David M. Warshaw, and Kathleen M. Trybus. A mutant heterodimeric myosin with one inactive head generates maximal displacement. Journal of Cell Biology, 162(3):481–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/162/3/
481.

Kohler:2003:DDL

[KRMB03] Danny Köhler, Christine Ruff, Edgar Meyhöfer, and Martin Bähler. Different degrees of lever arm rotation control myosin step size. Journal of Cell Biology, 161(2):237–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
161/2/237.

Krimmer:2001:BPO

[KRR^{+}01] Thomas Krimmer, Doron Rapaport, Michael T. Ryan, Chris Meisinger, C. Kenneth Kassenbrock, Elizabeth Blachly-Dyson, Michael Forte, Michael G. Douglas, Walter Neupert, Frank E. Nargang, and Nikolaus Pfanner. Biogenesis of Porin of the Outer Mitochondrial Membrane Involves an Import Pathway via Receptors and the General Import Pore of the Tom Complex. Journal of Cell Biology, 152(2):289–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
152/2/289.

Kleijmeer:2001:RMB

[KRS^{+}01] Monique Kleijmeer, Georg Ramm, Danita Schuurhuis, Janice Griffith, Maria Rescigno, Paola Ricciardi-Castagnoli, Alexander Y. Rudensky, Ferry Ossendorp, Cornelis J. M. Melief, Willem Stoorvogel, and Hans J. Geuze. Reorganization of multivesicular bodies regulates MHC class II antigen presentation by dendritic cells. Journal of Cell Biology, 155 (1):53–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/1/53.

Khodjakov:2002:NFC

[KRS^{+}02] Alexey Khodjakov, Conly L. Rieder, Greenfield Sluder, Grisel Cassels, Ody Sibon, and Chuo-Lung Wang. De novo formation of centrosomes in vertebrate cells arrested during S phase. Journal of Cell Biology, 158(7):1171–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
7/1171.

Kamimura:2004:RNS

[KRU^{+}04] Keisuke Kamimura, John M. Rhodes, Ryu Ueda, Melissa McNeely, Deepak Shukla, Koji Kimata, Patricia G. Spear, Nicholas W. Shworak, and Hiroshi Nakato. Regulation of Notch signaling by Drosophila heparan sulfate 3- O sulfotransferase. Journal of Cell Biology, 166(7):1069–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
166/7/1069.

Kuznetsov:2004:IOH

[KRZ^{+}04] Sergei A. Kuznetsov, Mara Riminucci, Navid Ziran, Takeo W. Tsutsui, Alessandro Corsi, Laura Calvi, Henry M. Kronenberg, Ernestina Schipani, Pamela Gehron Robey, and Paolo Bianco. The interplay of osteogenesis and hematopoiesis. Journal of Cell Biology, 167(6):1113–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/167/6/1113.

Komada:2002:SRS

[KS02] Masayuki Komada and Philippe Soriano. βiv-spectrin regulates sodium channel clustering through ankyrin–G at axon initial segments and nodes of Ranvier. Journal of Cell Biology, 156(2):337–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/2/337.

Katzmann:2003:VRE

[KSBE03] David J. Katzmann, Christopher J. Stefan, Markus Babst, and Scott D. Emr. Vps27 recruits ESCRT machinery to endosomes during MVB sorting. Journal of Cell Biology, 162 (3):413–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/3/413.

Kimura:2002:TCR

[KSC02] Hiroshi Kimura, Kimihiko Sugaya, and Peter R. Cook. The transcription cycle of RNA polymerase II in living cells. Journal of Cell Biology, 159(5):777–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/5/777.

Kadrmas:2004:IEP

[KSC^{+}04] Julie L. Kadrmas, Mark A. Smith, Kathleen A. Clark, Stephen M. Pronovost, Nemone Muster, John R. Yates, and Mary C. Beckerle. The integrin effector PINCH regulates JNK activity and epithelial migration in concert with Ras suppressor 1. Journal of Cell Biology, 167(6):1019– ??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/6/1019.

Khater:2004:MMS

[KSD04] Emad I. Khater, Robert E. Sinden, and Johannes T. Dessens. A malaria membrane skeletal protein is essential for normal morphogenesis, motility, and infectivity of sporozoites. Journal of Cell Biology, 167(3):425–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/3/425.

Kida:2000:MTT

[KSF^{+}00] Yuichiro Kida, Masao Sakaguchi, Mitsunori Fukuda, Katsuhiko Mikoshiba, and Katsuyoshi Mihara. Membrane Topogenesis of a Type I Signal–Anchor Protein, Mouse Synaptotagmin Ii, on the Endoplasmic Reticulum. Journal of Cell Biology, 150(4):719–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/4/719.

Krause:2000:FBP

[KSK^{+}00] Matthias Krause, Antonio S. Sechi, Marlies Konradt, David Monner, Frank B. Gertler, and Jürgen Wehland. Fynbinding protein (Fyb)/SLP-76-associated protein (SLAP), Ena/vasodilator-stimulated phosphoprotein (VASP) proteins and the Arp2/3 complex link T cell receptor (Tcr) signaling to the actin cytoskeleton. Journal of Cell Biology, 149 (1):181–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/1/181.

Kornitzer:2001:AEP

[KSK01] Daniel Kornitzer, Rakefet Sharf, and Tamar Kleinberger. Adenovirus E4orf4 protein induces PP2A-dependent growth arrest in Saccharomyces cerevisiae and interacts with the anaphase-promoting complex/cyclosome. Journal of Cell Biology, 154(2):331–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/2/331.

Kufer:2002:HTR

[KSK^{+}02] Thomas A. Kufer, Herman H. W. Silljé, Roman Körner, Oliver J. Gruss, Patrick Meraldi, and Erich A. Nigg. Human TPX2 is required for targeting Aurora–A kinase to the spindle. Journal of Cell Biology, 158(4):617–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/4/
617.

Kunz:2001:MAI

[KSM^{+}01] Stefan Kunz, Noemí Sevilla, Dorian B. McGavern, Kevin P. Campbell, and Michael B. A. Oldstone. Molecular analysis of the interaction of LCMV with its cellular receptor αdystroglycan. Journal of Cell Biology, 155(2):301–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
155/2/301.

Kabir:2001:PKC

[KSN^{+}01] Nurul Kabir, Andrew W. Schaefer, Arash Nakhost, Wayne S. Sossin, and Paul Forscher. Protein Kinase C Activation Promotes Microtubule Advance in Neuronal Growth Cones by Increasing Average Microtubule Growth Lifetimes. Journal of Cell Biology, 152(5):1033–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/5/1033.

Kee:2004:SNT

[KSNS^{+}04] Anthony J. Kee, Galina Schevzov, Visalini Nair-Shalliker, C. Stephen Robinson, Bernadette Vrhovski, Majid Ghoddusi, Min Ru Qiu, Jim J.-C. Lin, Ron Weinberger, Peter W. Gunning, and Edna C. Hardeman. Sorting of a nonmuscle tropomyosin to a novel cytoskeletal compartment in skeletal muscle results in muscular dystrophy. Journal of Cell Biology, 166(5):685–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/5/685.

Kreuz:2004:NAF

[KSR^{+}04] Sebastian Kreuz, Daniela Siegmund, Jost-Julian Rumpf, Dierk Samel, Martin Leverkus, Ottmar Janssen, Georg Häcker, Oliver Dittrich-Breiholz, Michael Kracht, Peter Scheurich, and Harald Wajant. NFκB activation by Fas is mediated through FADD, caspase-8, and RIP and is inhibited by FLIP. Journal of Cell Biology, 166(3):369–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
3/369.

Kaufmann:2003:CSD

[KSS^{+}03] Thomas Kaufmann, Sarah Schlipf, Javier Sanz, Karin Neubert, Reuven Stein, and Christoph Borner. Characterization of the signal that directs Bcl-x_{L}, but not Bcl-2, to the mitochondrial outer membrane. Journal of Cell Biology, 160 (1):53–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/1/53.

Kline-Smith:2000:MO

[KSW00] Susan L. Kline-Smith and Claire E. Walczak. Microtubules in orbit. Journal of Cell Biology, 149(1):5–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/149/1/5.

Kano:2000:MCK

[KTY^{+}00] Fumi Kano, Katsuya Takenaka, Akitsugu Yamamoto, Kuniaki Nagayama, Eisuke Nishida, and Masayuki Murata. Mek and Cdc2 Kinase Are Sequentially Required for Golgi Disassembly in Mdck Cells by the Mitotic Xenopus Extracts.

Journal of Cell Biology, 149(2):357–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/2/357.

Kurzchalia:2003:ATR

[Kur03] Teymuras Kurzchalia. Anthrax toxin rafts into cells. Journal of Cell Biology, 160(3):295–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/3/295.

Kayali:2003:SCD

[KVC^{+}03] Ayse G. Kayali, Kurt Van Gunst, Iain L. Campbell, Aleksandr Stotland, Marcie Kritzik, Guoxun Liu, Malin Flodström-Tullberg, You-Qing Zhang, and Nora Sarvetnick. The stromal cell–derived factor-1α/CXCR4 ligand–receptor axis is critical for progenitor survival and migration in the pancreas. Journal of Cell Biology, 163(4):859–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
4/859.

Kotelevets:2001:LPA

[KvHB^{+}01] Larissa Kotelevets, Jolanda van Hengel, Erik Bruyneel, Marc Mareel, Frans van Roy, and Eric Chastre. The lipid phosphatase activity of PTEN is critical for stabilizing intercellular junctions and reverting invasiveness. Journal of Cell Biology, 155(7):1129–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/7/1129.

Kikkawa:2003:MCO

[KVM03] Yamato Kikkawa, Ismo Virtanen, and Jeffrey H. Miner. Mesangial cells organize the glomerular capillaries by adhering to the G domain of laminin α5 in the glomerular basement membrane. Journal of Cell Biology, 161(1):187–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
1/187.

Kang:2000:DRC

[KWH^{+}00] Shin-Jung Kang, Suyue Wang, Hideaki Hara, Erin P. Peterson, Shobu Namura, Sepideh Amin-Hanjani, Zhihong Huang, Anu Srinivasan, Kevin J. Tomaselli, Nancy A. Thornberry, Michael A. Moskowitz, and Junying Yuan. Dual Role of Caspase-11 in Mediating Activation of Caspase-1 and Caspase-3 under Pathological Conditions. Journal of Cell Biology, 149(3):613–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/3/613.

Krantz:2000:PSR

[KWO^{+}00] David E. Krantz, Clarissa Waites, Viola Oorschot, Yongjian Liu, Rachel I. Wilson, Philip K. Tan, Judith Klumperman, and Robert H. Edwards. A Phosphorylation Site Regulates Sorting of the Vesicular Acetylcholine Transporter to Dense Core Vesicles. Journal of Cell Biology, 149(2):379–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
2/379.

Komeili:2000:MMC

[KWOP00] Arash Komeili, Karen P. Wedaman, Erin K. O’Shea, and Ted Powers. Mechanism of Metabolic Control. Journal of Cell Biology, 151(4):863–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/4/863.

Kim:2002:BDN

[KWS^{+}02] Sun H. Kim, Seok J. Won, Seonghyang Sohn, Hyuk J. Kwon, Jee Y. Lee, Jong H. Park, and Byoung J. Gwag. Brain-derived neurotrophic factor can act as a pronecrotic factor through transcriptional and translational activation of NADPH oxidase. Journal of Cell Biology, 159(5):821– ??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/5/821.

Koli:2004:DLF

[KWSK^{+}04] Katri Koli, Frank Wempe, Anja Sterner-Kock, Anna Kantola, Martina Komor, Wolf-K. Hofmann, Harald von Melchner, and Jorma Keski-Oja. Disruption of LTBP-4 function reduces TGF-β activation and enhances BMP-4 signaling in the lung. Journal of Cell Biology, 167(1):123–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
1/123.

Kazama:2000:OKR

[KY00] Hirotaka Kazama and Shin Yonehara. Oncogenic K–Ras and basic fibroblast growth factor prevent FAS-mediated apoptosis in fibroblasts through activation of mitogen-activated protein kinase. Journal of Cell Biology, 148(3):557–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
3/557.

Kempiak:2003:LSE

[KYBS03] Stephan J. Kempiak, Shu-Chin Yip, Jonathan M. Backer, and Jeffrey E. Segall. Local signaling by the EGF receptor. Journal of Cell Biology, 162(5):781–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/162/5/
781.

Karakesisoglou:2000:EPI

[KYF00] Iakowos Karakesisoglou, Yanmin Yang, and Elaine Fuchs. An Epidermal Plakin That Integrates Actin and Microtubule Networks at Cellular Junctions. Journal of Cell Biology, 149 (1):195–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/1/195.

Kress:2004:NRC

[KYM04] Tracy L. Kress, Young J. Yoon, and Kimberly L. Mowry. Nuclear RNP complex assembly initiates cytoplasmic RNA localization. Journal of Cell Biology, 165(2):203–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/165/
2/203.

Kazarov:2002:EST

[KYS^{+}02] Alexander R. Kazarov, Xiuwei Yang, Christopher S. Stipp, Bantoo Sehgal, and Martin E. Hemler. An extracellular site on tetraspanin CD151 determines α3 and α6 integrindependent cellular morphology. Journal of Cell Biology, 158(7):1299–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/7/1299.

Kim:2004:CPB

[KYW^{+}04] Kyoungtae Kim, Atsuko Yamashita, Martin A. Wear, Yuichiro Maéda, and John A. Cooper. Capping protein binding to actin in yeast. Journal of Cell Biology, 164(4): 567–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/4/567.

Kang:2004:NNP

[KYZ^{+}04] Jong-Sun Kang, Min-Jeong Yi, Wei Zhang, Jessica L. Feinleib, Francesca Cole, and Robert S. Krauss. Netrins and neogenin promote myotube formation. Journal of Cell Biology, 167(3):493–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/493.

Lindholm:2004:CD

[LA04] Dan Lindholm and Urmas Arumäe. Cell differentiation. Journal of Cell Biology, 167(2):193–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/2/193.

Leissring:2000:CCE

[LAF^{+}00] Malcolm A. Leissring, Yama Akbari, Christopher M. Fanger, Michael D. Cahalan, Mark P. Mattson, and Frank M. LaFerla. Capacitative Calcium Entry Deficits and Elevated Luminal Calcium Content in Mutant Presenilin-1 Knockin Mice. Journal of Cell Biology, 149(4):793–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/149/4/
793.

Lai:2003:ACE

[Lai03a] Eric C. Lai. APP causes an energy crisis. Journal of Cell Biology, 161(1):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/1/12.3.

Lai:2003:LL

[Lai03b] Eric C. Lai. Lamas in loops. Journal of Cell Biology, 161(1):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/1/13.2.

Lai:2003:LLC

[Lai03c] Eric C. Lai. Life in a low calcium home. Journal of Cell Biology, 161(1):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/1/12.2.

Lai:2003:LRM

[Lai03d] Eric C. Lai. Lipid rafts make for slippery platforms. Journal of Cell Biology, 162(3):365–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/3/365.

Lai:2003:RPC

[Lai03e] Eric C. Lai. RB pockets the cell cycle. Journal of Cell Biology, 161(1):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/1/12.1.

Lai:2003:WNN

[Lai03f] Eric C. Lai. Where new neurons come from. Journal of Cell Biology, 161(1):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/1/13.1.

Levraud:2003:DCD

[LAL^{+}03] Jean-Pierre Levraud, Myriam Adam, Marie-Françoise Luciani, Chantal de Chastellier, Richard L. Blanton, and Pierre Golstein. Dictyostelium cell death. Journal of Cell Biology, 160(7):1105–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/7/1105.

Lasko:2000:DMG

[Las00] Paul Lasko. The Drosophila melanogaster Genome. Journal of Cell Biology, 150(2):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/2/F51.

Lasko:2003:CLO

[Las03] Paul Lasko. Cup-ling oskar RNA localization and translational control. Journal of Cell Biology, 163(6):1189–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/163/6/1189.

Lemichez:2003:HHQ

[LB03] Emmanuel Lemichez and Patrice Boquet. To be helped or not helped, that is the question. Journal of Cell Biology, 160 (7):991–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/7/991.

Lee:2002:CPT

[LBH^{+}02] Shaun W. Lee, Robert A. Bonnah, Dustin L. Higashi, John P. Atkinson, Sharon L. Milgram, and Magdalene So. CD46 is phosphorylated at tyrosine 354 upon infection of epithelial cells by Neisseria gonorrhoeae. Journal of Cell Biology, 156 (6):951–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/6/951.

Lee:2000:FYM

[LBP00] Wei-Lih Lee, Magdalena Bezanilla, and Thomas D. Pollard. Fission yeast myosin-I, Myo1p, stimulates actin assembly by Arp2/3 complex and shares functions with Wasp. Journal of Cell Biology, 151(4):789–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/4/789.

Longva:2002:UPA

[LBS^{+}02] Karianne E. Longva, Frøydis D. Blystad, Espen Stang, Astrid M. Larsen, Lene E. Johannessen, and Inger H. Madshus. Ubiquitination and proteasomal activity is required for transport of the EGF receptor to inner membranes of multivesicular bodies. Journal of Cell Biology, 156(5):843–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/156/5/843.

Lopez-Boado:2000:BEI

[LBWH^{+}00] Yolanda S. López-Boado, Carole L. Wilson, Lora V. Hooper, Jeffrey I. Gordon, Scott J. Hultgren, and William C. Parks. Bacterial Exposure Induces and Activates Matrilysin in Mucosal Epithelial Cells. Journal of Cell Biology, 148(6): 1305–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/6/1305.

Levesque:2001:CKN

[LC01] Aime A. Levesque and Duane A. Compton. The chromokinesin Kid is necessary for chromosome arm orientation and oscillation, but not congression, on mitotic spindles. Journal of Cell Biology, 154(6):1135–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/6/1135.

Li:2004:CDR

[LC04] Yun Li and Patricia Camacho. Ca^{2+}-dependent redox modulation of SERCA 2b by ERp57. Journal of Cell Biology, 164 (1):35–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/1/35.

Liebner:2004:CRE

[LCG^{+}04] Stefan Liebner, Anna Cattelino, Radiosa Gallini, Noemi Rudini, Monica Iurlaro, Stefano Piccolo, and Elisabetta Dejana. β-catenin is required for endothelial-mesenchymal transformation during heart cushion development in the mouse. Journal of Cell Biology, 166(3):359–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/3/359.

Lavoie:2000:AEO

[LCGR00] Josée N. Lavoie, Claudia Champagne, Marie-Claude Gingras, and Amélie Robert. Adenovirus E4 Open Reading Frame 4–Induced Apoptosis Involves Dysregulation of Src Family Kinases. Journal of Cell Biology, 150(5):1037–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
5/1037.

Li:2001:RTP

[LCI^{+}01] Feng Li, Jianhua Chen, Masako Izumi, Mark C. Butler, Susan M. Keezer, and David M. Gilbert. The replication timing program of the Chinese hamster β-globin locus is established coincident with its repositioning near peripheral heterochromatin in early G1 phase. Journal of Cell Biology, 154 (2):283–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/2/283.

Lane:2000:SDN

[LCM00] Cynthia M. Lane, Ian Cushman, and Mary Shannon Moore. Selective Disruption of Nuclear Import by a Functional Mutant Nuclear Transport Carrier. Journal of Cell Biology, 151(2):321–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/2/321.

Lambert:2002:DLI

[LCM02] Mireille Lambert, Daniel Choquet, and René-Marc Mège. Dynamics of ligand-induced, Rac1-dependent anchoring of cadherins to the actin cytoskeleton. Journal of Cell Biology, 157(3):469–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/3/469.

Liu:2001:ADN

[LCRS01] Yewei Liu, Zoltán Cseresnyés, William R. Randall, and Martin F. Schneider. Activity-dependent nuclear translocation and intranuclear distribution of NFATc in adult skeletal muscle fibers. Journal of Cell Biology, 155(1):27–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
1/27.

Li:2001:FGF

[LCS^{+}01] Xiaofeng Li, Yali Chen, Susanne Schéele, Esther Arman, Rebecca Haffner-Krausz, Peter Ekblom, and Peter Lonai. Fibroblast Growth Factor Signaling and Basement Membrane Assembly Are Connected during Epithelial Morphogenesis of the Embryoid Body. Journal of Cell Biology, 153(4):811–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/4/811.

Li:2002:ITI

[LCS^{+}02] Zhuqing Li, Maria J. Calzada, John M. Sipes, Jo Anne Cashel, Henry C. Krutzsch, Douglas S. Annis, Deane F. Mosher, and David D. Roberts. Interactions of thrombospondins with α4β1 integrin and CD47 differentially modulate T cell behavior. Journal of Cell Biology, 157(3):509–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
157/3/509.

Li:2003:SDS

[LCSG03] Feng Li, Jianhua Chen, Eduardo Solessio, and David M. Gilbert. Spatial distribution and specification of mammalian replication origins during G1 phase. Journal of Cell Biology, 161(2):257–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/257.

Li:2004:PAM

[LCT^{+}04] Jingxia Li, Haobin Chen, Moon-Shong Tang, Xianglin Shi, Shantu Amin, Dhimant Desai, Max Costa, and Chuanshu Huang. PI-3K and Akt are mediators of AP-1 induction by 5-MCDE in mouse epidermal Cl41 cells. Journal of Cell Biology, 165(1):77–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/1/77.

Lawrence:2004:SKN

[LDC^{+}04] Carolyn J. Lawrence, R. Kelly Dawe, Karen R. Christie, Don W. Cleveland, Scott C. Dawson, Sharyn A. Endow, Lawrence S. B. Goldstein, Holly V. Goodson, Nobutaka Hirokawa, Jonathon Howard, Russell L. Malmberg, J. Richard McIntosh, Harukata Miki, Timothy J. Mitchison, Yasushi Okada, Anireddy S. N. Reddy, William M. Saxton, Manfred Schliwa, Jonathan M. Scholey, Ronald D. Vale, Claire E. Walczak, and Linda Wordeman. A standardized kinesin nomenclature. Journal of Cell Biology, 167(1):19–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/167/1/
19.

Lin:2001:PRB

[LdCK^{+}01] Haijiang Lin, Pedro de Carvalho, David Kho, Chin-Yin Tai, Philippe Pierre, Gerald R. Fink, and David Pellman. Polyploids require Bik1 for kinetochore–microtubule attachment. Journal of Cell Biology, 155(7):1173–??, December 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/155/7/

Ling:2003:TPT

[LDI^{+}03] Kun Ling, Renee L. Doughman, Vidhya V. Iyer, Ari J. Firestone, Shawn F. Bairstow, Deane F. Mosher, Michael D. Schaller, and Richard A. Anderson. Tyrosine phosphorylation of type Iγ phosphatidylinositol phosphate kinase by Src regulates an integrin–talin switch. Journal of Cell Biology, 163(6):1339–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/6/1339.

Liu:2003:BER

[LDK^{+}03] Jia-Jia Liu, Jianqing Ding, Anthony S. Kowal, Timothy Nardine, Elizabeth Allen, Jean-Dominique Delcroix, Chengbiao Wu, William Mobley, Elaine Fuchs, and Yanmin Yang. BPAG1n4 is essential for retrograde axonal transport in sensory neurons. Journal of Cell Biology, 163(2):223–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
163/2/223.

Lee:2002:YMS

[LDP02] Margaret S. Lee, Karen A. D’Amour, and Jackie Papkoff. A yeast model system for functional analysis of β-catenin signaling. Journal of Cell Biology, 158(6):1067–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
6/1067.

Lechardeur:2000:DNL

[LDS^{+}00] Delphine Lechardeur, Luke Drzymala, Manu Sharma, Danuta Zylka, Robert Kinach, Joanna Pacia, Christopher Hicks, Nawaid Usmani, Johanna M. Rommens, and Gergely L. Lukacs. Determinants of the nuclear localization of the heterodimeric DNA fragmentation factor (Icad/Cad). Journal of Cell Biology, 150(2):321–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/2/321.

Lindsten:2002:MUF

[LdVV^{+}02] Kristina Lindsten, Femke M. S. de Vrij, Lisette G. G. C. Verhoef, David F. Fischer, Fred W. van Leeuwen, Elly M. Hol, Maria G. Masucci, and Nico P. Dantuma. Mutant ubiquitin found in neurodegenerative disorders is a ubiquitin fusion degradation substrate that blocks proteasomal degradation. Journal of Cell Biology, 157(3):417–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/3/417.

LeBrasseur:2002:APU

[LeB02a] Nicole LeBrasseur. ABU to perform under stress. Journal of Cell Biology, 158(4):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/4/605.1.

LeBrasseur:2002:AC

[LeB02b] Nicole LeBrasseur. Actin to completion. Journal of Cell Biology, 156(5):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/5/766.3.

LeBrasseur:2002:ACC

[LeB02c] Nicole LeBrasseur. Active cells choose life. Journal of Cell Biology, 158(3):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/3/384.2.

LeBrasseur:2002:ADK

[LeB02d] Nicole LeBrasseur. An amidation a day keeps apoptosis away. Journal of Cell Biology, 159(2):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/2/201.2.

LeBrasseur:2002:AL

[LeB02e] Nicole LeBrasseur. Anticancer lubrication. Journal of Cell Biology, 156(6):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/6/940.1.

LeBrasseur:2002:ACM

[LeB02f] Nicole LeBrasseur. Arrested cells are Mad as Hec. Journal of Cell Biology, 159(2):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/2/201.3.

LeBrasseur:2002:APF

[LeB02g] Nicole LeBrasseur. Arteries play follow the leader. Journal of Cell Biology, 158(1):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/1/10.2.

LeBrasseur:2002:BTM

[LeB02h] Nicole LeBrasseur. Bacteria tear down microtubules. Journal of Cell Biology, 158(1):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/1/10.1.

LeBrasseur:2002:BPR

[LeB02i] Nicole LeBrasseur. Big proteins Ran into import difficulties. Journal of Cell Biology, 159(1):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/1/12.2.

LeBrasseur:2002:BCI

[LeB02j] Nicole LeBrasseur. Blocking caspase-independent death. Journal of Cell Biology, 157(5):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/5/734.2.

LeBrasseur:2002:BCE

[LeB02k] Nicole LeBrasseur. Blood cells expand their numbers. Journal of Cell Biology, 157(3):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/3/346.2.

LeBrasseur:2002:BRV

[LeB02l] Nicole LeBrasseur. Branching out requires VEGF. Journal of Cell Biology, 159(2):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/2/201.1.

LeBrasseur:2002:BDB

[LeB02m] Nicole LeBrasseur. The brightest dots in the bunch. Journal of Cell Biology, 159(6):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/6/913.1.

LeBrasseur:2002:BAT

[LeB02n] Nicole LeBrasseur. The bud is all in the timing. Journal of Cell Biology, 156(5):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/5/766.1.

LeBrasseur:2002:BAR

[LeB02o] Nicole LeBrasseur. Build actin at the right place, right time. Journal of Cell Biology, 159(5):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/5/726.1.

LeBrasseur:2002:BLM

[LeB02p] Nicole LeBrasseur. Bundle loss makes turning possible. Journal of Cell Biology, 157(5):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/5/734.1.

LeBrasseur:2002:BC

[LeB02q] Nicole LeBrasseur. Bundle up for cytokinesis. Journal of Cell Biology, 157(7):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/7/1100.2.

LeBrasseur:2002:CIA

[LeB02r] Nicole LeBrasseur. Calcium increases the appetite. Journal of Cell Biology, 159(1):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/1/12.1.

LeBrasseur:2002:CSA

[LeB02s] Nicole LeBrasseur. cGMP steers axons home. Journal of Cell Biology, 159(3):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/3/388.1.

LeBrasseur:2002:CS

[LeB02t] Nicole LeBrasseur. Choose a side. Journal of Cell Biology, 159(2):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/2/200.1.

LeBrasseur:2002:CT

[LeB02u] Nicole LeBrasseur. Clash of the titin. Journal of Cell Biology, 157(1):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/1/10.2.

LeBrasseur:2002:CM

[LeB02v] Nicole LeBrasseur. Collagen melt-down. Journal of Cell Biology, 156(4):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/4/589.2.

LeBrasseur:2002:CUS

[LeB02w] Nicole LeBrasseur. Coming undone at the seams. Journal of Cell Biology, 156(6):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/6/940.2.

LeBrasseur:2002:CD

[LeB02x] Nicole LeBrasseur. Connexin the dots. Journal of Cell Biology, 157(4):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/4/547.2.

LeBrasseur:2002:CR

[LeB02y] Nicole LeBrasseur. Coordination of replication. Journal of Cell Biology, 158(3):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/3/384.3.

LeBrasseur:2002:CHS

[LeB02z] Nicole LeBrasseur. COPs are held up by Sec16p. Journal of Cell Biology, 158(6):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/6/990.2.

LeBrasseur:2002:CTC

[LeB02-27] Nicole LeBrasseur. Crumbs turns the corner. Journal of Cell Biology, 156(6):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/6/941.1.

LeBrasseur:2002:DCS

[LeB02-28] Nicole LeBrasseur. Decoding Ca^{2+} signals through cAMP. Journal of Cell Biology, 158(3):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/3/385.2.

LeBrasseur:2002:DSD

[LeB02-29] Nicole LeBrasseur. DEDD spells death to caspase substrates. Journal of Cell Biology, 158(6):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/6/991.1.

LeBrasseur:2002:DCC

[LeB02-30] Nicole LeBrasseur. Dendritic cells can be self sustaining. Journal of Cell Biology, 159(4):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/4/539.3.

LeBrasseur:2002:DI

[LeB02-31] Nicole LeBrasseur. Desensitizing interactions. Journal of Cell Biology, 157(6):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/6/907.2.

LeBrasseur:2002:DPA

[LeB02-32] Nicole LeBrasseur. A dual purpose antideath agent. Journal of Cell Biology, 156(6):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/6/941.2.

LeBrasseur:2002:DTV

[LeB02-33] Nicole LeBrasseur. Dynein takes its vimentin. Journal of Cell Biology, 157(5):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/5/735.2.

LeBrasseur:2002:CSR

[LeB02-34] Nicole LeBrasseur. E-cadherin in a stabilizing relationship. Journal of Cell Biology, 159(3):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/3/389.1.

LeBrasseur:2002:ECM

[LeB02-35] Nicole LeBrasseur. ECM communicates with MMP. Journal of Cell Biology, 159(3):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/3/388.2.

LeBrasseur:2002:EA

[LeB02-36] Nicole LeBrasseur. Enabling actin. Journal of Cell Biology, 156(5):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/5/767.2.

LeBrasseur:2002:CE

[LeB02-37] Nicole LeBrasseur. The ER contributes to engulfment. Journal of Cell Biology, 158(3):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/3/385.1.

LeBrasseur:2002:ENH

[LeB02-38] Nicole LeBrasseur. Everyone needs a home. Journal of Cell Biology, 159(1):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/1/13.1.

LeBrasseur:2002:ENM

[LeB02-39] Nicole LeBrasseur. Exchange is necessary for migration. Journal of Cell Biology, 157(3):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/3/347.1.

LeBrasseur:2002:MC

[LeB02-40] Nicole LeBrasseur. From migration to circulation. Journal of Cell Biology, 157(2):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/2/197.2.

LeBrasseur:2002:GCV

[LeB02-41] Nicole LeBrasseur. A GAP in COPI vesicle formation is filled. Journal of Cell Biology, 159(1):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/1/13.3.

LeBrasseur:2002:GGK

[LeB02-42] Nicole LeBrasseur. Getting a GRIP on kinesin movement. Journal of Cell Biology, 157(4):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/4/546.1.

LeBrasseur:2002:GSS

[LeB02-43] Nicole LeBrasseur. Getting sick from Sic. Journal of Cell Biology, 157(6):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/6/906.1.

LeBrasseur:2002:GB

[LeB02-44] Nicole LeBrasseur. Gnawing at the bone. Journal of Cell Biology, 158(6):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/6/991.2.

LeBrasseur:2002:HTC

[LeB02-45] Nicole LeBrasseur. Hiding telomerase from chromatin. Journal of Cell Biology, 158(7):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/7/1154.1.

LeBrasseur:2002:HTE

[LeB02-46] Nicole LeBrasseur. How tumors escape oblivion. Journal of Cell Biology, 157(6):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/6/906.3.

LeBrasseur:2002:KCP

[LeB02-47] Nicole LeBrasseur. Keeping the cord in place. Journal of Cell Biology, 156(4):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/4/588.3.

LeBrasseur:2002:KKE

[LeB02-48] Nicole LeBrasseur. A kinase kick for endocytosis. Journal of Cell Biology, 156(5):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/5/767.1.

LeBrasseur:2002:LSA

[LeB02-49] Nicole LeBrasseur. L1 sticks around. Journal of Cell Biology, 157(7):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/7/1101.2.

LeBrasseur:2002:LP

[LeB02-50] Nicole LeBrasseur. Learning by palmitoylation. Journal of Cell Biology, 157(2):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/2/196.1.

LeBrasseur:2002:LTC

[LeB02-51] Nicole LeBrasseur. Lipids tell channel to open wide. Journal of Cell Biology, 158(5):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/5/831.2.

LeBrasseur:2002:LPL

[LeB02-52] Nicole LeBrasseur. Little PIP, little PIP, let me in. Journal of Cell Biology, 159(2):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/2/200.2.

LeBrasseur:2002:LPC

[LeB02-53] Nicole LeBrasseur. Lon plays the chaperone. Journal of Cell Biology, 157(7):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/7/1100.3.

LeBrasseur:2002:MNP

[LeB02-54] Nicole LeBrasseur. Mads at the nuclear pore. Journal of Cell Biology, 159(5):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/5/726.2.

LeBrasseur:2002:MMB

[LeB02-55] Nicole LeBrasseur. Mapping MAP binding. Journal of Cell Biology, 157(7):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/7/1101.1.

LeBrasseur:2002:MGE

[LeB02-56] Nicole LeBrasseur. Microtubules get early release. Journal of Cell Biology, 159(5):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/5/727.3.

LeBrasseur:2002:MGS

[LeB02-57] Nicole LeBrasseur. Microtubules get swept up in the excitement of mitosis. Journal of Cell Biology, 158(6):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
158/6/990.1.

LeBrasseur:2002:MST

[LeB02-58] Nicole LeBrasseur. A molecule with a sense of timing. Journal of Cell Biology, 158(3):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/3/384.1.

LeBrasseur:2002:NTS

[LeB02-59] Nicole LeBrasseur. Neurons are translationally smart. Journal of Cell Biology, 158(5):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/5/831.1.

LeBrasseur:2002:NGT

[LeB02-60] Nicole LeBrasseur. Neurons go toward the light. Journal of Cell Biology, 159(6):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/6/912.3.

LeBrasseur:2002:NCR

[LeB02-61] Nicole LeBrasseur. Neutrophils choose the right direction. Journal of Cell Biology, 159(1):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/1/13.2.

LeBrasseur:2002:NMJ

[LeB02-62] Nicole LeBrasseur. A new member joins the yeast filament party. Journal of Cell Biology, 157(3):??, April 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/157/3/
347.3.

LeBrasseur:2002:NTC

[LeB02-63] Nicole LeBrasseur. Nuclear travel costs energy. Journal of Cell Biology, 159(5):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/5/727.1.

LeBrasseur:2002:PSV

[LeB02-64] Nicole LeBrasseur. p115 SNAREs a vesicle. Journal of Cell Biology, 157(1):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/1/11.2.

LeBrasseur:2002:PSP

[LeB02-65] Nicole LeBrasseur. PECAM-1 senses the pressure. Journal of Cell Biology, 158(4):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/4/604.3.

LeBrasseur:2002:PP

[LeB02-66] Nicole LeBrasseur. Peroxide perturbations. Journal of Cell Biology, 157(1):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/1/10.1.

LeBrasseur:2002:P

[LeB02-67] Nicole LeBrasseur. PilT. Journal of Cell Biology, 159(6): ??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/6/912.2.

LeBrasseur:2002:PHV

[LeB02-68] Nicole LeBrasseur. Pinosomes help a virus escape. Journal of Cell Biology, 158(6):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/6/990.3.

LeBrasseur:2002:PFB

[LeB02-69] Nicole LeBrasseur. Playing FtsZ before sporulation. Journal of Cell Biology, 157(4):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/4/546.2.

LeBrasseur:2002:PHM

[LeB02-70] Nicole LeBrasseur. Poking a hole in mitochondria. Journal of Cell Biology, 159(4):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/4/539.1.

LeBrasseur:2002:PAT

[LeB02-71] Nicole LeBrasseur. Polyamino acids take amyloid form. Journal of Cell Biology, 159(4):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/4/538.2.

LeBrasseur:2002:PMK

[LeB02-72] Nicole LeBrasseur. Pressured into making a kidney. Journal of Cell Biology, 157(2):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/2/197.1.

LeBrasseur:2002:PRT

[LeB02-73] Nicole LeBrasseur. Promiscuous Ras turns on the wrong partner. Journal of Cell Biology, 158(7):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/7/
1154.3.

LeBrasseur:2002:PSC

[LeB02-74] Nicole LeBrasseur. Prostate stem cells identified. Journal of Cell Biology, 157(7):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/7/1100.1.

LeBrasseur:2002:PCM

[LeB02-75] Nicole LeBrasseur. Proteases commit microbicide. Journal of Cell Biology, 157(2):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/2/196.2.

LeBrasseur:2002:PCR

[LeB02-76] Nicole LeBrasseur. Putting a cap on reinitiation. Journal of Cell Biology, 159(4):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/4/539.2.

LeBrasseur:2002:QA

[LeB02-77] Nicole LeBrasseur. Quantity assurance. Journal of Cell Biology, 158(4):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/4/604.2.

LeBrasseur:2002:RMT

[LeB02-78] Nicole LeBrasseur. Rabs make their way to mitochondria. Journal of Cell Biology, 158(4):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/4/604.1.

LeBrasseur:2002:RI

[LeB02-79] Nicole LeBrasseur. Ras on the inside. Journal of Cell Biology, 157(3):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/3/346.1.

LeBrasseur:2002:RP

[LeB02-80] Nicole LeBrasseur. Ribosomes pile up. Journal of Cell Biology, 159(3):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/3/389.3.

LeBrasseur:2002:RRW

[LeB02-81] Nicole LeBrasseur. ROCK and Rho will never Dia. Journal of Cell Biology, 157(5):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/5/734.3.

LeBrasseur:2002:RSC

[LeB02-82] Nicole LeBrasseur. ROCK-steady centriole migration. Journal of Cell Biology, 157(5):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/5/735.1.

LeBrasseur:2002:SCS

[LeB02-83] Nicole LeBrasseur. Securin correct separation. Journal of Cell Biology, 158(1):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/1/11.2.

LeBrasseur:2002:SM

[LeB02-84] Nicole LeBrasseur. Seeing in motion. Journal of Cell Biology, 158(5):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/5/830.1.

LeBrasseur:2002:STC

[LeB02-85] Nicole LeBrasseur. Seeing through the cilium. Journal of Cell Biology, 157(1):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/1/10.3.

LeBrasseur:2002:SFC

[LeB02-86] Nicole LeBrasseur. Selectin flexible cells. Journal of Cell Biology, 158(4):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/4/605.2.

LeBrasseur:2002:SJA

[LeB02-87] Nicole LeBrasseur. Semaphorins — not just for axons. Journal of Cell Biology, 158(7):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/7/1154.2.

LeBrasseur:2002:SSC

[LeB02-88] Nicole LeBrasseur. Sequences specify centromeric chromatin. Journal of Cell Biology, 159(5):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/5/727.2.

LeBrasseur:2002:SSSa

[LeB02-89] Nicole LeBrasseur. Sex, Sox, and splicing. Journal of Cell Biology, 156(4):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/4/588.1.

LeBrasseur:2002:SL

[LeB02-90] Nicole LeBrasseur. Shocking localization. Journal of Cell Biology, 156(5):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/5/766.2.

LeBrasseur:2002:SSM

[LeB02-91] Nicole LeBrasseur. Slow and steady moves the cell. Journal of Cell Biology, 157(6):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/6/907.1.

LeBrasseur:2002:SIC

[LeB02-92] Nicole LeBrasseur. Sorting it out, without clathrin. Journal of Cell Biology, 157(4):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/4/546.3.

LeBrasseur:2002:SGN

[LeB02-93] Nicole LeBrasseur. Stable globin needs a partner. Journal of Cell Biology, 158(1):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/1/10.3.

LeBrasseur:2002:STH

[LeB02-94] Nicole LeBrasseur. Stand up and take hold. Journal of Cell Biology, 158(7):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/7/1155.2.

LeBrasseur:2002:SMN

[LeB02-95] Nicole LeBrasseur. Structural Mad-ness. Journal of Cell Biology, 156(4):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/4/588.2.

LeBrasseur:2002:SY

[LeB02-96] Nicole LeBrasseur. Suicidal yeast. Journal of Cell Biology, 157(3):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/3/347.2.

LeBrasseur:2002:SPR

[LeB02-97] Nicole LeBrasseur. Swallowing pain relief. Journal of Cell Biology, 156(4):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/4/589.1.

LeBrasseur:2002:SSSb

[LeB02-98] Nicole LeBrasseur. SynCAM and sidekick synchronize synapse synthesis. Journal of Cell Biology, 158(7):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/158/7/1155.1.

LeBrasseur:2002:TSA

[LeB02-99] Nicole LeBrasseur. Tension seals axons fate. Journal of Cell Biology, 159(3):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/3/389.2.

LeBrasseur:2002:TC

[LeB02-100] Nicole LeBrasseur. A time for cytokinesis. Journal of Cell Biology, 157(6):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/6/906.2.

LeBrasseur:2002:TIM

[LeB02-101] Nicole LeBrasseur. Topo II on the move? Journal of Cell Biology, 157(1):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/1/11.1.

LeBrasseur:2002:TJS

[LeB02-102] Nicole LeBrasseur. Traffic jam suppresses immunity. Journal of Cell Biology, 157(2):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/2/196.3.

LeBrasseur:2002:TB

[LeB02-103] Nicole LeBrasseur. Traveling backward. Journal of Cell Biology, 158(5):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/5/830.2.

LeBrasseur:2002:T

[LeB02-104] Nicole LeBrasseur. A trip to the ER. Journal of Cell Biology, 156(6):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/6/940.3.

LeBrasseur:2002:URF

[LeB02-105] Nicole LeBrasseur. Ubiquitin receptor found. Journal of Cell Biology, 158(1):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/1/11.1.

LeBrasseur:2002:VSS

[LeB02-106] Nicole LeBrasseur. Vesicles slide on a synaptic body. Journal of Cell Biology, 159(6):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/6/912.1.

LeBrasseur:2002:VPE

[LeB02-107] Nicole LeBrasseur. Vesicles that promise enlargement. Journal of Cell Biology, 159(6):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/6/913.2.

LeBrasseur:2002:VW

[LeB02-108] Nicole LeBrasseur. Viva the worm. Journal of Cell Biology, 159(4):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/4/538.1.

LeBrasseur:2002:WLS

[LeB02-109] Nicole LeBrasseur. When LDH says “open,” channels listen. Journal of Cell Biology, 158(5):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/5/830.3.

LeBrasseur:2002:WTC

[LeB02-110] Nicole LeBrasseur. Widening the target for cancer treatment. Journal of Cell Biology, 157(4):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/4/547.1.

LeBrasseur:2003:ATC

[LeB03a] Nicole LeBrasseur. The APC–tumor connection. Journal of Cell Biology, 163(5):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/5/926.2.

LeBrasseur:2003:ALM

[LeB03b] Nicole LeBrasseur. Apoptosis leaves its mark on chromatin. Journal of Cell Biology, 161(5):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/5/835.1.

LeBrasseur:2003:ABP

[LeB03c] Nicole LeBrasseur. Aurora B puts chromosomes in their place. Journal of Cell Biology, 161(2):??, April 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/2/
218.1.

LeBrasseur:2003:AAE

[LeB03d] Nicole LeBrasseur. Auxin auxin everywhere. Journal of Cell Biology, 163(6):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/6/1186.2.

LeBrasseur:2003:BTC

[LeB03e] Nicole LeBrasseur. Binding takes cooperation. Journal of Cell Biology, 163(2):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/2/201.3.

LeBrasseur:2003:BHS

[LeB03f] Nicole LeBrasseur. BMP in a hairy situation. Journal of Cell Biology, 163(3):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/3/432.1.

LeBrasseur:2003:BCC

[LeB03g] Nicole LeBrasseur. BubR1 checks in with CENP–E. Journal of Cell Biology, 162(4):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/4/531.3.

LeBrasseur:2003:BBN

[LeB03h] Nicole LeBrasseur. Building and breaking the NPC. Journal of Cell Biology, 160(7):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/7/980.1.

LeBrasseur:2003:BBT

[LeB03i] Nicole LeBrasseur. A building block of two organelles. Journal of Cell Biology, 162(2):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/2/167.1.

LeBrasseur:2003:CSC

[LeB03j] Nicole LeBrasseur. Calcium for strong clotting. Journal of Cell Biology, 160(7):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/7/980.2.

LeBrasseur:2003:CMF

[LeB03k] Nicole LeBrasseur. Calcium makes a fast getaway. Journal of Cell Biology, 163(6):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/6/1187.2.

LeBrasseur:2003:CCE

[LeB03l] Nicole LeBrasseur. Caspases chew on electron transport. Journal of Cell Biology, 160(1):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/1/6.1.

LeBrasseur:2003:CUP

[LeB03m] Nicole LeBrasseur. Catastrophe under pressure. Journal of Cell Biology, 161(6):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/6/1006.2.

LeBrasseur:2003:CLS

[LeB03n] Nicole LeBrasseur. Centrosomes lead stem cell orientation. Journal of Cell Biology, 162(7):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/7/1181.1.

LeBrasseur:2003:CRO

[LeB03o] Nicole LeBrasseur. Choosing replication origins. Journal of Cell Biology, 161(2):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/219.2.

LeBrasseur:2003:CT

[LeB03p] Nicole LeBrasseur. Choreographing telomeres. Journal of Cell Biology, 163(1):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/1/11.3.

LeBrasseur:2003:CLG

[LeB03q] Nicole LeBrasseur. The complex life of a GEF. Journal of Cell Biology, 160(1):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/1/7.1.

LeBrasseur:2003:CBK

[LeB03r] Nicole LeBrasseur. Cyclin B knows its place. Journal of Cell Biology, 161(5):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/5/835.3.

LeBrasseur:2003:CPP

[LeB03s] Nicole LeBrasseur. A cytokine that packs a punch. Journal of Cell Biology, 161(5):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/5/835.2.

LeBrasseur:2003:DUD

[LeB03t] Nicole LeBrasseur. Deacetylase undoes DNA damage. Journal of Cell Biology, 160(7):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/7/981.1.

LeBrasseur:2003:DT

[LeB03u] Nicole LeBrasseur. Desmosomes in a tangle. Journal of Cell Biology, 163(2):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/2/200.1.

LeBrasseur:2003:DCSb

[LeB03v] Nicole LeBrasseur. Desulfation on the cell surface. Journal of Cell Biology, 162(2):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/2/166.2.

LeBrasseur:2003:DFD

[LeB03w] Nicole LeBrasseur. DNA-free division. Journal of Cell Biology, 160(7):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/7/981.2.

LeBrasseur:2003:CIC

[LeB03x] Nicole LeBrasseur. E-cadherin immobilizes cells from the inside. Journal of Cell Biology, 161(6):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/6/1007.1.

LeBrasseur:2003:ENF

[LeB03y] Nicole LeBrasseur. Eggs never forget. Journal of Cell Biology, 163(6):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/6/1187.1.

LeBrasseur:2003:EMT

[LeB03z] Nicole LeBrasseur. Endosomes make tracks. Journal of Cell Biology, 160(6):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/6/805.2.

LeBrasseur:2003:ESB

[LeB03-27] Nicole LeBrasseur. Evolution steps back. Journal of Cell Biology, 162(7):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/7/1180.1.

LeBrasseur:2003:FPF

[LeB03-28] Nicole LeBrasseur. Fewer phosphates but fatter axons. Journal of Cell Biology, 163(5):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/5/926.1.

LeBrasseur:2003:FB

[LeB03-29] Nicole LeBrasseur. Fibronectin for branching. Journal of Cell Biology, 162(1):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/1/12.1.

LeBrasseur:2003:FA

[LeB03-30] Nicole LeBrasseur. Finding the axon. Journal of Cell Biology, 162(6):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/6/959.3.

LeBrasseur:2003:FFF

[LeB03-31] Nicole LeBrasseur. Flies fight famine. Journal of Cell Biology, 163(2):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/2/200.2.

LeBrasseur:2003:FML

[LeB03-32] Nicole LeBrasseur. Fold me or leave me. Journal of Cell Biology, 163(4):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/4/691.2.

LeBrasseur:2003:FCS

[LeB03-33] Nicole LeBrasseur. Forcing cells to stick around. Journal of Cell Biology, 163(3):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/3/433.2.

LeBrasseur:2003:FIS

[LeB03-34] Nicole LeBrasseur. FRET-ing over scaffolds. Journal of Cell Biology, 160(1):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/1/7.2.

LeBrasseur:2003:FFK

[LeB03-35] Nicole LeBrasseur. Fusion fuels killing. Journal of Cell Biology, 163(5):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/5/927.2.

LeBrasseur:2003:FGG

[LeB03-36] Nicole LeBrasseur. Fusion gets in the groove. Journal of Cell Biology, 163(4):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/4/691.3.

LeBrasseur:2003:GPS

[LeB03-37] Nicole LeBrasseur. G-proteins for spindle geometry. Journal of Cell Biology, 162(4):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/4/531.2.

LeBrasseur:2003:GBB

[LeB03-38] Nicole LeBrasseur. The glue that builds bristles. Journal of Cell Biology, 162(6):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/6/959.1.

LeBrasseur:2003:GUD

[LeB03-39] Nicole LeBrasseur. Golgi united are also divided. Journal of Cell Biology, 161(6):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/6/1007.2.

LeBrasseur:2003:GSU

[LeB03-40] Nicole LeBrasseur. GPR-1/ 2 support unequal division. Journal of Cell Biology, 161(5):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/5/834.2.

LeBrasseur:2003:GWL

[LeB03-41] Nicole LeBrasseur. Grow wings for a limited time only. Journal of Cell Biology, 162(7):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/7/1181.2.

LeBrasseur:2003:HPM

[LeB03-42] Nicole LeBrasseur. H. pylori mobilizes cells. Journal of Cell Biology, 161(2):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/219.1.

LeBrasseur:2003:HPR

[LeB03-43] Nicole LeBrasseur. The healing power of Ron. Journal of Cell Biology, 162(5):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/5/751.1.

LeBrasseur:2003:HSB

[LeB03-44] Nicole LeBrasseur. Heparin sustains the brain. Journal of Cell Biology, 163(1):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/1/10.1.

LeBrasseur:2003:HCC

[LeB03-45] Nicole LeBrasseur. How cholesterol clogs the arteries. Journal of Cell Biology, 162(5):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/5/750.1.

LeBrasseur:2003:HSD

[LeB03-46] Nicole LeBrasseur. Hox signals death for neuronal precursors. Journal of Cell Biology, 160(4):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/4/464.1.

LeBrasseur:2003:ICT

[LeB03-47] Nicole LeBrasseur. Immune cells take in unknown enemies. Journal of Cell Biology, 163(4):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/4/691.1.

LeBrasseur:2003:ISD

[LeB03-48] Nicole LeBrasseur. Inducing Sprouty degradation. Journal of Cell Biology, 160(6):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/6/804.1.

LeBrasseur:2003:IUU

[LeB03-49] Nicole LeBrasseur. Integrin is on uPAR to unleash cells. Journal of Cell Biology, 163(1):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/1/11.2.

LeBrasseur:2003:KWC

[LeB03-50] Nicole LeBrasseur. Keeping wing cells in line. Journal of Cell Biology, 160(4):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/4/465.1.

LeBrasseur:2003:KSC

[LeB03-51] Nicole LeBrasseur. Keratin for supple cells. Journal of Cell Biology, 162(7):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/7/1181.3.

LeBrasseur:2003:KAP

[LeB03-52] Nicole LeBrasseur. Kicking out antiviral proteins. Journal of Cell Biology, 162(2):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/2/167.3.

LeBrasseur:2003:KK

[LeB03-53] Nicole LeBrasseur. The kinase knows. Journal of Cell Biology, 163(2):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/2/201.1.

LeBrasseur:2003:KNA

[LeB03-54] Nicole LeBrasseur. Knocked down a Notch with age. Journal of Cell Biology, 163(6):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/6/1187.3.

LeBrasseur:2003:LRK

[LeB03-55] Nicole LeBrasseur. Lonely receptors kill cells. Journal of Cell Biology, 162(5):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/5/750.2.

LeBrasseur:2003:LSA

[LeB03-56] Nicole LeBrasseur. Losing self-awareness. Journal of Cell Biology, 162(5):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/5/751.3.

LeBrasseur:2003:MEM

[LeB03-57] Nicole LeBrasseur. A minus-end motor sees the plus side. Journal of Cell Biology, 160(4):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/4/465.2.

LeBrasseur:2003:MSC

[LeB03-58] Nicole LeBrasseur. Mitochondria sensitive to ER Ca^{2+} stores. Journal of Cell Biology, 160(7):??, March 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/7/
981.3.

LeBrasseur:2003:MOD

[LeB03-59] Nicole LeBrasseur. More order in DNA. Journal of Cell Biology, 162(6):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/6/959.2.

LeBrasseur:2003:NCT

[LeB03-60] Nicole LeBrasseur. Neural cells turn and run. Journal of Cell Biology, 162(2):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/2/167.2.

LeBrasseur:2003:NOW

[LeB03-61] Nicole LeBrasseur. Neurons die one way or another. Journal of Cell Biology, 163(5):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/5/927.1.

LeBrasseur:2003:NTB

[LeB03-62] Nicole LeBrasseur. NRG-1 talks back to neurons. Journal of Cell Biology, 161(6):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/6/1007.3.

LeBrasseur:2003:NCU

[LeB03-63] Nicole LeBrasseur. Nucleoli come undone. Journal of Cell Biology, 160(6):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/6/804.2.

LeBrasseur:2003:NGP

[LeB03-64] Nicole LeBrasseur. Nucleosomes get peeled and rewrapped. Journal of Cell Biology, 160(6):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/6/805.1.

LeBrasseur:2003:NCO

[LeB03-65] Nicole LeBrasseur. Nucleotides choose origins. Journal of Cell Biology, 162(5):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/5/751.2.

LeBrasseur:2003:OSK

[LeB03-66] Nicole LeBrasseur. The other side of kinetochore fibers. Journal of Cell Biology, 160(5):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/5/624.1.

LeBrasseur:2003:PSN

[LeB03-67] Nicole LeBrasseur. PDE5 says NO to cGMP. Journal of Cell Biology, 160(5):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/5/625.1.

LeBrasseur:2003:PCV

[LeB03-68] Nicole LeBrasseur. Phosphorylation cycles on vesicles. Journal of Cell Biology, 160(5):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/5/624.2.

LeBrasseur:2003:PI

[LeB03-69] Nicole LeBrasseur. Polyamines in inflammation. Journal of Cell Biology, 162(2):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/2/166.1.

LeBrasseur:2003:PPC

[LeB03-70] Nicole LeBrasseur. Prions prefer caveosomes. Journal of Cell Biology, 162(4):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/4/530.1.

LeBrasseur:2003:PPP

[LeB03-71] Nicole LeBrasseur. Promoting passage in plasmodesmata. Journal of Cell Biology, 160(4):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/4/464.3.

LeBrasseur:2003:PIU

[LeB03-72] Nicole LeBrasseur. A protease inhibitor unsticks cells. Journal of Cell Biology, 160(5):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/5/625.2.

LeBrasseur:2003:QCT

[LeB03-73] Nicole LeBrasseur. Quick cartilage transformation. Journal of Cell Biology, 163(3):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/3/433.3.

LeBrasseur:2003:REA

[LeB03-74] Nicole LeBrasseur. Rafts and the evil amyloids. Journal of Cell Biology, 160(1):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/1/6.2.

LeBrasseur:2003:RSG

[LeB03-75] Nicole LeBrasseur. Ran sticks a GEF to chromatin. Journal of Cell Biology, 160(5):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/5/625.3.

LeBrasseur:2003:RRa

[LeB03-76] Nicole LeBrasseur. Rec’d and repaired. Journal of Cell Biology, 160(4):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/4/464.2.

LeBrasseur:2003:RMQ

[LeB03-77] Nicole LeBrasseur. Receptors make a quick exit. Journal of Cell Biology, 163(4):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/4/690.2.

LeBrasseur:2003:RE

[LeB03-78] Nicole LeBrasseur. Recycling endosomes. Journal of Cell Biology, 163(1):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/1/10.2.

LeBrasseur:2003:RN

[LeB03-79] Nicole LeBrasseur. Remodeling the nucleus. Journal of Cell Biology, 162(1):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/1/13.1.

LeBrasseur:2003:RKB

[LeB03-80] Nicole LeBrasseur. Riding Kar9 into the bud. Journal of Cell Biology, 160(6):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/6/805.3.

LeBrasseur:2003:RAC

[LeB03-81] Nicole LeBrasseur. Ring around the chromatin. Journal of Cell Biology, 161(1):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/1/15.1.

LeBrasseur:2003:RTZ

[LeB03-82] Nicole LeBrasseur. RNA travels with ZBP1. Journal of Cell Biology, 160(1):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/1/7.3.

LeBrasseur:2003:RRb

[LeB03-83] Nicole LeBrasseur. ROCK and Rho in 3D. Journal of Cell Biology, 163(3):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/3/432.2.

LeBrasseur:2003:SNI

[LeB03-84] Nicole LeBrasseur. Say NO to insulin. Journal of Cell Biology, 161(2):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/219.3.

LeBrasseur:2003:SD

[LeB03-85] Nicole LeBrasseur. Searching DNA in 3D. Journal of Cell Biology, 161(1):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/1/15.2.

LeBrasseur:2003:SP

[LeB03-86] Nicole LeBrasseur. Shaping up plants. Journal of Cell Biology, 161(1):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/1/14.1.

LeBrasseur:2003:SMS

[LeB03-87] Nicole LeBrasseur. Spare the Mbl, spoil the rod? Journal of Cell Biology, 162(1):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/1/12.2.

LeBrasseur:2003:SGA

[LeB03-88] Nicole LeBrasseur. The spindle gets the Axs. Journal of Cell Biology, 161(1):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/1/14.2.

LeBrasseur:2003:SGL

[LeB03-89] Nicole LeBrasseur. Stop and go L1CAM. Journal of Cell Biology, 162(4):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/4/530.2.

LeBrasseur:2003:STE

[LeB03-90] Nicole LeBrasseur. Synapse too exclusive for CD148. Journal of Cell Biology, 162(4):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/4/531.1.

LeBrasseur:2003:TA

[LeB03-91] Nicole LeBrasseur. Tearing down actin. Journal of Cell Biology, 163(5):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/5/927.3.

LeBrasseur:2003:TSa

[LeB03-92] Nicole LeBrasseur. Tearing down the spindle. Journal of Cell Biology, 163(4):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/4/690.1.

LeBrasseur:2003:GSL

[LeB03-93] Nicole LeBrasseur. To grow or to shrink …. Journal of Cell Biology, 161(2):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/218.2.

LeBrasseur:2003:KBC

[LeB03-94] Nicole LeBrasseur. To kill a brain cell. Journal of Cell Biology, 163(1):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/1/11.1.

LeBrasseur:2003:TSb

[LeB03-95] Nicole LeBrasseur. Transcribed in synchronicity. Journal of Cell Biology, 163(6):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/6/1186.1.

LeBrasseur:2003:TKI

[LeB03-96] Nicole LeBrasseur. Transcription keeps it simple. Journal of Cell Biology, 163(2):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/2/201.2.

LeBrasseur:2003:T

[LeB03-97] Nicole LeBrasseur. Translocon. Journal of Cell Biology, 162 (7):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/7/1180.2.

LeBrasseur:2003:TWV

[LeB03-98] Nicole LeBrasseur. Two ways VEGF–A builds a vessel. Journal of Cell Biology, 161(6):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/6/1006.1.

LeBrasseur:2003:VAM

[LeB03-99] Nicole LeBrasseur. A view of antibody maturation. Journal of Cell Biology, 161(5):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/5/834.1.

LeBrasseur:2003:WCD

[LeB03-100] Nicole LeBrasseur. We can deplete it, we have the technology. Journal of Cell Biology, 162(6):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/6/958.1.

LeBrasseur:2003:WNP

[LeB03-101] Nicole LeBrasseur. When nuclear proteins go mitotic. Journal of Cell Biology, 162(6):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/6/958.2.

LeBrasseur:2003:WMM

[LeB03-102] Nicole LeBrasseur. Where mtDNA is made. Journal of Cell Biology, 163(3):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/3/433.1.

LeBrasseurr:2003:DBO

[LeB03-103] Nicole LeBrasseurr. Dying of a bad overhang. Journal of Cell Biology, 161(1):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/1/15.3.

LeBrasseur:2004:A

[LeB04a] Nicole LeBrasseur. Ameloblastin. Journal of Cell Biology, 167(5):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/5/804.1.

LeBrasseur:2004:APC

[LeB04b] Nicole LeBrasseur. Another passenger on chromosomes. Journal of Cell Biology, 166(2):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/2/152.2.

LeBrasseur:2004:AC

[LeB04c] Nicole LeBrasseur. Antagonizing T cells. Journal of Cell Biology, 166(4):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/4/442.2.

LeBrasseur:2004:AMP

[LeB04d] Nicole LeBrasseur. Antigens make a pit stop. Journal of Cell Biology, 164(5):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/5/633.3.

LeBrasseur:2004:WWL

[LeB04e] Nicole LeBrasseur. Any which way but loose. Journal of Cell Biology, 164(6):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/6/795.1.

LeBrasseur:2004:AAT

[LeB04f] Nicole LeBrasseur. Arp2/ 3 for axon turning. Journal of Cell Biology, 166(3):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/3/307.1.

LeBrasseur:2004:ACF

[LeB04g] Nicole LeBrasseur. Arpless comets are fascin-ating. Journal of Cell Biology, 165(2):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/2/162.2.

LeBrasseur:2004:AYC

[LeB04h] Nicole LeBrasseur. Assembling yeast complexes. Journal of Cell Biology, 165(1):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/1/17.1.

LeBrasseur:2004:AH

[LeB04i] Nicole LeBrasseur. Awakening heterochromatin. Journal of Cell Biology, 164(1):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/1/7.1.

LeBrasseur:2004:BTG

[LeB04j] Nicole LeBrasseur. BARS takes Golgi apart. Journal of Cell Biology, 166(3):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/3/306.1.

LeBrasseur:2004:BSS

[LeB04k] Nicole LeBrasseur. Best stiffness for striation. Journal of Cell Biology, 166(6):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/6/760.1.

LeBrasseur:2004:BTB

[LeB04l] Nicole LeBrasseur. Breathe, then breed. Journal of Cell Biology, 164(4):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/4/478.2.

LeBrasseur:2004:CPN

[LeB04m] Nicole LeBrasseur. Calcium pumps need a chaperone. Journal of Cell Biology, 164(1):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/1/7.3.

LeBrasseur:2004:CBL

[LeB04n] Nicole LeBrasseur. Cells have a bubbly look. Journal of Cell Biology, 167(2):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/2/190.2.

LeBrasseur:2004:CSB

[LeB04o] Nicole LeBrasseur. Cells step back in time. Journal of Cell Biology, 165(5):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/5/605.3.

LeBrasseur:2004:CFH

[LeB04p] Nicole LeBrasseur. A chaperone feels the heat. Journal of Cell Biology, 167(2):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/2/190.1.

LeBrasseur:2004:CSS

[LeB04q] Nicole LeBrasseur. Checking in on spindle size. Journal of Cell Biology, 167(6):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/6/995.3.

LeBrasseur:2004:CHA

[LeB04r] Nicole LeBrasseur. Cholesterol hastens Alzheimer’s. Journal of Cell Biology, 165(1):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/1/16.2.

LeBrasseur:2004:CPA

[LeB04s] Nicole LeBrasseur. Cholesterol prevents Alzheimer’s. Journal of Cell Biology, 167(5):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/5/805.1.

LeBrasseur:2004:CS

[LeB04t] Nicole LeBrasseur. Collagen for swallowing. Journal of Cell Biology, 166(4):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/4/443.3.

LeBrasseur:2004:CF

[LeB04u] Nicole LeBrasseur. Condensation by folding. Journal of Cell Biology, 166(6):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/6/761.3.

LeBrasseur:2004:C

[LeB04v] Nicole LeBrasseur. Correction. Journal of Cell Biology, 166(6):925–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/6/925.

LeBrasseur:2004:DCS

[LeB04w] Nicole LeBrasseur. Death comes in small packages. Journal of Cell Biology, 165(3):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/3/297.2.

LeBrasseur:2004:DDT

[LeB04x] Nicole LeBrasseur. Death DNase tied down. Journal of Cell Biology, 167(5):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/5/804.3.

LeBrasseur:2004:DTH

[LeB04y] Nicole LeBrasseur. Death takes a holiday. Journal of Cell Biology, 166(3):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/3/307.2.

LeBrasseur:2004:DRA

[LeB04z] Nicole LeBrasseur. Division at right angles. Journal of Cell Biology, 166(5):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/5/609.2.

LeBrasseur:2004:DFF

[LeB04-27] Nicole LeBrasseur. Dynamin in fission and fusion. Journal of Cell Biology, 167(6):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/6/994.1.

LeBrasseur:2004:DIH

[LeB04-28] Nicole LeBrasseur. A dynamite intercellular highway. Journal of Cell Biology, 164(6):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/6/795.2.

LeBrasseur:2004:CEE

[LeB04-29] Nicole LeBrasseur. E-cadherin evades endocytosis. Journal of Cell Biology, 166(2):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/2/153.3.

LeBrasseur:2004:ERA

[LeB04-30] Nicole LeBrasseur. Endosomes are ready and APPL to signal. Journal of Cell Biology, 164(6):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/6/794.1.

LeBrasseur:2004:ES

[LeB04-31] Nicole LeBrasseur. Escorting STAT5A. Journal of Cell Biology, 167(3):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/394.1.

LeBrasseur:2004:EI

[LeB04-32] Nicole LeBrasseur. Expaaaaanded inclusions. Journal of Cell Biology, 167(3):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/394.2.

LeBrasseur:2004:ECF

[LeB04-33] Nicole LeBrasseur. Export in cystic fibrosis. Journal of Cell Biology, 167(1):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/1/12.2.

LeBrasseur:2004:FBa

[LeB04-34] Nicole LeBrasseur. Fat and bone. Journal of Cell Biology, 165(3):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/3/297.3.

LeBrasseur:2004:FSM

[LeB04-35] Nicole LeBrasseur. Feudal system in metabolic flux. Journal of Cell Biology, 164(6):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/6/794.2.

LeBrasseur:2004:FFP

[LeB04-36] Nicole LeBrasseur. Fighting foreigners prevents selfdestruction. Journal of Cell Biology, 165(3):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/165/3/
296.2.

LeBrasseur:2004:FLF

[LeB04-37] Nicole LeBrasseur. Fingers lay fibers. Journal of Cell Biology, 165(4):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/4/454.1.

LeBrasseur:2004:FHR

[LeB04-38] Nicole LeBrasseur. A force for holocentrics to reckon with. Journal of Cell Biology, 166(7):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/7/938.2.

LeBrasseur:2004:FBb

[LeB04-39] Nicole LeBrasseur. Forced to bond. Journal of Cell Biology, 166(6):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/6/761.2.

LeBrasseur:2004:GHD

[LeB04-40] Nicole LeBrasseur. Geminin halts DNA synthesis. Journal of Cell Biology, 165(4):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/4/455.1.

LeBrasseur:2004:GIC

[LeB04-41] Nicole LeBrasseur. Getting to the inner circle. Journal of Cell Biology, 166(5):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/5/609.3.

LeBrasseur:2004:GMM

[LeB04-42] Nicole LeBrasseur. GTPase modes modeled. Journal of Cell Biology, 167(4):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/4/581.2.

LeBrasseur:2004:HSC

[LeB04-43] Nicole LeBrasseur. Heart-shaped clones. Journal of Cell Biology, 164(1):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/1/6.1.

LeBrasseur:2004:HGS

[LeB04-44] Nicole LeBrasseur. Hitched genes still independent. Journal of Cell Biology, 166(6):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/6/761.1.

LeBrasseur:2004:IA

[LeB04-45] Nicole LeBrasseur. Immune to autoimmunity. Journal of Cell Biology, 165(5):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/5/604.2.

LeBrasseur:2004:IM

[LeB04-46] Nicole LeBrasseur. Inflexibility in motion. Journal of Cell Biology, 164(4):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/4/479.1.

LeBrasseur:2004:IBH

[LeB04-47] Nicole LeBrasseur. Insulin breaks hearts. Journal of Cell Biology, 167(6):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/6/995.1.

LeBrasseur:2004:IKT

[LeB04-48] Nicole LeBrasseur. Insulin keeps time. Journal of Cell Biology, 167(2):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/2/191.2.

LeBrasseur:2004:ITC

[LeB04-49] Nicole LeBrasseur. Integrins talk to N-cadherin. Journal of Cell Biology, 166(2):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/2/153.2.

LeBrasseur:2004:ICT

[LeB04-50] Nicole LeBrasseur. Intimacy is a cellular turn-on. Journal of Cell Biology, 165(3):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/3/296.1.

LeBrasseur:2004:KCM

[LeB04-51] Nicole LeBrasseur. Kif2a cures monopolar disorder. Journal of Cell Biology, 166(4):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/4/443.1.

LeBrasseur:2004:KSG

[LeB04-52] Nicole LeBrasseur. Kinases speak from the Golgi. Journal of Cell Biology, 164(7):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/7/950.2.

LeBrasseur:2004:LML

[LeB04-53] Nicole LeBrasseur. Lethargic mitochondria are left out. Journal of Cell Biology, 165(2):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/2/162.1.

LeBrasseur:2004:LT

[LeB04-54] Nicole LeBrasseur. Loopy for telomeres. Journal of Cell Biology, 166(2):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/2/153.1.

LeBrasseur:2004:LLP

[LeB04-55] Nicole LeBrasseur. Lumens laid out by PAR-1. Journal of Cell Biology, 164(5):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/5/632.2.

LeBrasseur:2004:MPT

[LeB04-56] Nicole LeBrasseur. Malaria protease takes an indirect route. Journal of Cell Biology, 164(1):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/1/7.2.

LeBrasseur:2004:MCM

[LeB04-57] Nicole LeBrasseur. MARK clears MAP-muddled traffic. Journal of Cell Biology, 167(1):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/1/13.1.

LeBrasseur:2004:MD

[LeB04-58] Nicole LeBrasseur. Mass dedifferentiation. Journal of Cell Biology, 165(1):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/1/17.2.

LeBrasseur:2004:MFS

[LeB04-59] Nicole LeBrasseur. Membrane fusion SNAREs an inhibitor. Journal of Cell Biology, 164(1):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/1/6.2.

LeBrasseur:2004:MR

[LeB04-60] Nicole LeBrasseur. Migration on rafts. Journal of Cell Biology, 164(5):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/5/632.1.

LeBrasseur:2004:MCD

[LeB04-61] Nicole LeBrasseur. A model of cell death. Journal of Cell Biology, 166(6):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/6/760.2.

LeBrasseur:2004:MLC

[LeB04-62] Nicole LeBrasseur. Muscle loves company. Journal of Cell Biology, 164(5):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/5/633.2.

LeBrasseur:2004:NRS

[LeB04-63] Nicole LeBrasseur. Nanofibers have the right stuff. Journal of Cell Biology, 164(4):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/4/479.3.

LeBrasseur:2004:NCA

[LeB04-64] Nicole LeBrasseur. Narcissistic CLIP-170 also attracts dynactin. Journal of Cell Biology, 166(7):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/166/7/
939.2.

LeBrasseur:2004:NMS

[LeB04-65] Nicole LeBrasseur. Nerveless muscles shape up. Journal of Cell Biology, 164(7):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/7/950.1.

LeBrasseur:2004:NVS

[LeB04-66] Nicole LeBrasseur. Next vesicle SNAP’d up. Journal of Cell Biology, 165(2):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/2/163.2.

LeBrasseur:2004:NCF

[LeB04-67] Nicole LeBrasseur. NG2^{+} cells feed the brain. Journal of Cell Biology, 165(4):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/4/455.3.

LeBrasseur:2004:NEA

[LeB04-68] Nicole LeBrasseur. No endocytosis in the axonal entry. Journal of Cell Biology, 166(4):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/4/442.1.

LeBrasseur:2004:NMD

[LeB04-69] Nicole LeBrasseur. No more division. Journal of Cell Biology, 164(7):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/7/951.1.

LeBrasseur:2004:NSM

[LeB04-70] Nicole LeBrasseur. The nucleus shuns MAPK. Journal of Cell Biology, 164(5):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/5/633.1.

LeBrasseur:2004:OCP

[LeB04-71] Nicole LeBrasseur. Opening closed pores. Journal of Cell Biology, 167(6):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/6/994.2.

LeBrasseur:2004:PPb

[LeB04-72] Nicole LeBrasseur. Palindromic protection. Journal of Cell Biology, 167(4):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/4/580.1.

LeBrasseur:2004:PRC

[LeB04-73] Nicole LeBrasseur. Patches ride cables. Journal of Cell Biology, 167(3):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/394.3.

LeBrasseur:2004:PPP

[LeB04-74] Nicole LeBrasseur. Ping pong in the pore. Journal of Cell Biology, 166(5):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/5/609.1.

LeBrasseur:2004:PPT

[LeB04-75] Nicole LeBrasseur. PKCs pick through Ca^{2+}. Journal of Cell Biology, 165(2):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/2/163.1.

LeBrasseur:2004:PPA

[LeB04-76] Nicole LeBrasseur. Polarity precedes adhesion. Journal of Cell Biology, 167(1):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/1/13.2.

LeBrasseur:2004:PSA

[LeB04-77] Nicole LeBrasseur. Pollen spares all but self. Journal of Cell Biology, 165(5):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/5/605.2.

LeBrasseur:2004:PTC

[LeB04-78] Nicole LeBrasseur. The pore–transcription connection. Journal of Cell Biology, 165(5):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/5/604.1.

LeBrasseur:2004:PPc

[LeB04-79] Nicole LeBrasseur. Profilin for processivity. Journal of Cell Biology, 167(4):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/4/581.3.

LeBrasseur:2004:PSI

[LeB04-80] Nicole LeBrasseur. Protein splicing for immunity. Journal of Cell Biology, 164(4):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/4/478.1.

LeBrasseur:2004:QRT

[LeB04-81] Nicole LeBrasseur. Quick results for thyroid hormone. Journal of Cell Biology, 167(5):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/5/805.2.

LeBrasseur:2004:RK

[LeB04-82] Nicole LeBrasseur. A Rho for the kinetochore. Journal of Cell Biology, 165(3):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/3/297.1.

LeBrasseur:2004:RI

[LeB04-83] Nicole LeBrasseur. RNA de-inhibition. Journal of Cell Biology, 165(1):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/1/16.1.

LeBrasseur:2004:RML

[LeB04-84] Nicole LeBrasseur. Running Mad-ly into mitosis. Journal of Cell Biology, 166(3):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/3/307.3.

LeBrasseur:2004:SBB

[LeB04-85] Nicole LeBrasseur. Scaffolds from blood to brain. Journal of Cell Biology, 166(7):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/7/939.1.

LeBrasseur:2004:SLb

[LeB04-86] Nicole LeBrasseur. Shh is out on a limb. Journal of Cell Biology, 166(5):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/5/608.2.

LeBrasseur:2004:SLa

[LeB04-87] Nicole LeBrasseur. Slingshot to the lamellipod. Journal of Cell Biology, 165(4):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/4/455.2.

LeBrasseur:2004:SHC

[LeB04-88] Nicole LeBrasseur. SNX1 hugs the curves. Journal of Cell Biology, 167(4):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/4/581.1.

LeBrasseur:2004:SNH

[LeB04-89] Nicole LeBrasseur. SOS needs help getting active. Journal of Cell Biology, 167(4):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/4/580.2.

LeBrasseur:2004:SR

[LeB04-90] Nicole LeBrasseur. Speed from recycling. Journal of Cell Biology, 167(3):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/395.1.

LeBrasseur:2004:STN

[LeB04-91] Nicole LeBrasseur. Splicing takes a NAP. Journal of Cell Biology, 167(1):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/1/12.1.

LeBrasseur:2004:SS

[LeB04-92] Nicole LeBrasseur. Sticky sugar. Journal of Cell Biology, 165(4):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/4/454.2.

LeBrasseur:2004:ST

[LeB04-93] Nicole LeBrasseur. Swelling with tumors. Journal of Cell Biology, 166(5):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/5/608.1.

LeBrasseur:2004:SSQ

[LeB04-94] Nicole LeBrasseur. A synaptic scaffold quiets T cells. Journal of Cell Biology, 166(2):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/2/152.1.

LeBrasseur:2004:TPE

[LeB04-95] Nicole LeBrasseur. Talin for presynaptic endocytosis. Journal of Cell Biology, 167(1):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/1/12.3.

LeBrasseur:2004:TTR

[LeB04-96] Nicole LeBrasseur. Teeny tiny rafts. Journal of Cell Biology, 164(6):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/6/795.3.

LeBrasseur:2004:TSP

[LeB04-97] Nicole LeBrasseur. Telomeric spindle proteins. Journal of Cell Biology, 167(5):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/5/804.2.

LeBrasseur:2004:TKA

[LeB04-98] Nicole LeBrasseur. Time keepers abound. Journal of Cell Biology, 167(6):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/6/995.2.

LeBrasseur:2004:TT

[LeB04-99] Nicole LeBrasseur. Tracking transcripts. Journal of Cell Biology, 165(2):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/2/163.3.

LeBrasseur:2004:TAT

[LeB04-100] Nicole LeBrasseur. Two arms of TGFβ signaling. Journal of Cell Biology, 164(7):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/7/951.3.

LeBrasseur:2004:TFM

[LeB04-101] Nicole LeBrasseur. The two faces of Mad2. Journal of Cell Biology, 165(1):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/1/16.3.

LeBrasseur:2004:UCG

[LeB04-102] Nicole LeBrasseur. Ubiquitin cycles for Golgi fusion. Journal of Cell Biology, 164(7):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/7/951.2.

LeBrasseur:2004:USb

[LeB04-103] Nicole LeBrasseur. Ubiquitin stripping. Journal of Cell Biology, 166(4):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/4/443.2.

LeBrasseur:2004:USa

[LeB04-104] Nicole LeBrasseur. Uncertainty in structures. Journal of Cell Biology, 165(5):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/5/605.1.

LeBrasseur:2004:UIC

[LeB04-105] Nicole LeBrasseur. Uncooperative immune cells. Journal of Cell Biology, 167(2):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/2/190.3.

LeBrasseur:2004:VCS

[LeB04-106] Nicole LeBrasseur. VASP comet sightings. Journal of Cell Biology, 167(2):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/2/191.1.

LeBrasseur:2004:WME

[LeB04-107] Nicole LeBrasseur. West Meets East. Journal of Cell Biology, 166(3):308–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/3/308.

LeBrasseur:2004:WUA

[LeB04-108] Nicole LeBrasseur. Winding and unwinding ATP synthase. Journal of Cell Biology, 164(4):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/4/479.2.

LeBrasseur:2004:TDB

[LeB04-109] Nicole LeBrasseur. With their dying breath. Journal of Cell Biology, 166(3):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/3/306.2.

LeBrasseur:2004:WQB

[LeB04-110] Nicole LeBrasseur. Wounds quiet basal cells. Journal of Cell Biology, 167(3):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/395.2.

LeBrasseur:2004:YAD

[LeB04-111] Nicole LeBrasseur. Yeast apoptosis debate continues. Journal of Cell Biology, 166(7):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/7/938.1.

Leslie:2001:DM

[Les01a] Mitch Leslie. Death by mistake. Journal of Cell Biology, 155 (2):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/2/179.3.

Leslie:2001:FB

[Les01b] Mitch Leslie. A fatal break up. Journal of Cell Biology, 155 (4):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/4/496.1.

Leslie:2001:FL

[Les01c] Mitch Leslie. Freed by light. Journal of Cell Biology, 155 (6):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/6/870.1.

Leslie:2001:GSW

[Les01d] Mitch Leslie. Going separate ways. Journal of Cell Biology, 155(6):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/6/870.2.

Leslie:2001:GT

[Les01e] Mitch Leslie. Growing together. Journal of Cell Biology, 155 (6):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/6/871.1.

Leslie:2001:NT

[Les01f] Mitch Leslie. In the nick of time. Journal of Cell Biology, 155 (4):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/4/496.2.

Leslie:2001:KD

[Les01g] Mitch Leslie. Kiss of death. Journal of Cell Biology, 155 (6):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/6/871.3.

Leslie:2001:LS

[Les01h] Mitch Leslie. Leaf starter. Journal of Cell Biology, 155 (2):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/2/179.2.

Leslie:2001:LE

[Les01i] Mitch Leslie. A lupus explanation. Journal of Cell Biology, 155(6):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/6/871.2.

Leslie:2001:MTF

[Les01j] Mitch Leslie. A mobile transcription factor. Journal of Cell Biology, 155(2):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/2/178.2.

Leslie:2001:SCD

[Les01k] Mitch Leslie. Sizing up cell division. Journal of Cell Biology, 155(2):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/2/179.1.

Leslie:2001:SW

[Les01l] Mitch Leslie. Slide this way. Journal of Cell Biology, 155 (4):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/4/497.3.

Leslie:2001:SAA

[Les01m] Mitch Leslie. Stayin’ alive, without antigen. Journal of Cell Biology, 155(2):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/2/178.1.

Leslie:2001:SC

[Les01n] Mitch Leslie. Staying connected. Journal of Cell Biology, 155 (4):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/4/497.1.

Leslie:2001:BB

[Les01o] Mitch Leslie. To branch or not to branch. Journal of Cell Biology, 155(4):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/4/497.2.

Leslie:2002:GSM

[Les02a] Mitch Leslie. The good side of a maligned protein. Journal of Cell Biology, 156(1):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/1/11.2.

Leslie:2002:GRR

[Les02b] Mitch Leslie. Greater risk from radiation? Journal of Cell Biology, 156(1):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/1/10.1.

Leslie:2002:GYO

[Les02c] Mitch Leslie. Grow-your-own synapses. Journal of Cell Biology, 156(1):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/1/11.1.

Leslie:2002:NG

[Les02d] Mitch Leslie. Numbers game. Journal of Cell Biology, 156 (1):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/1/10.2.

Leslie:2002:WC

[Les02e] Mitch Leslie. Waltz of the chromosomes. Journal of Cell Biology, 156(1):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/1/11.3.

Lelouard:2004:DCA

[LFM^{+}04] Hugues Lelouard, Vincent Ferrand, Didier Marguet, Jacek Bania, Voahirana Camosseto, Alexandre David, Evelina Gatti, and Philippe Pierre. Dendritic cell aggresome-like induced structures are dedicated areas for ubiquitination and storage of newly synthesized defective proteins. Journal of Cell Biology, 164(5):667–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/5/667.

Laux:2000:GMC

[LFT^{+}00] Thorsten Laux, Kiyoko Fukami, Marcus Thelen, Tamara Golub, Dunja Frey, and Pico Caroni. Gap43, Marcks, and Cap23 Modulate Pi(4,5)p_{2} at Plasmalemmal Rafts, and Regulate Cell Cortex Actin Dynamics through a Common Mechanism. Journal of Cell Biology, 149(7):1455–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/149/7/
1455.

Lyman:2001:NPC

[LG01] Susan K. Lyman and Larry Gerace. Nuclear pore complexes: dynamics in unexpected places. Journal of Cell Biology, 154 (1):17–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/1/17.

Li:2002:PPP

[LG02a] Xiaoling Li and Stephen J. Gould. PEX11 promotes peroxisome division independently of peroxisome metabolism.

Journal of Cell Biology, 156(4):643–??, February 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/4/
643.

Liddington:2002:IAT

[LG02b] R. C. Liddington and M. H. Ginsberg. Integrin activation takes shape. Journal of Cell Biology, 158(5):833–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
158/5/833.

Lyman:2002:ICS

[LGB^{+}02] Susan K. Lyman, Tinglu Guan, Janna Bednenko, Harald Wodrich, and Larry Gerace. Influence of cargo size on Ran and energy requirements for nuclear protein import. Journal of Cell Biology, 159(1):55–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/1/55.

Lacas-Gervais:2004:SSN

[LGGS^{+}04] Sandra Lacas-Gervais, Jun Guo, Nicola Strenzke, Eric Scarfone, Melanie Kolpe, Monika Jahkel, Pietro De Camilli, Tobias Moser, Matthew N. Rasband, and Michele Solimena. βIVσ1 spectrin stabilizes the nodes of Ranvier and axon initial segments. Journal of Cell Biology, 166(7):983–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/166/7/983.

Laloraya:2000:CAC

[LGK00] Shikha Laloraya, Vincent Guacci, and Douglas Koshland. Chromosomal Addresses of the Cohesin Component Mcd1p. Journal of Cell Biology, 151(5):1047–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/5/
1047.

Long:2001:ENR

[LGM^{+}01] Roy M. Long, Wei Gu, Xiuhua Meng, Graydon Gonsalvez, Robert H. Singer, and Pascal Chartrand. An Exclusively Nuclear RNA–Binding Protein Affects Asymmetric Localization of ASH1 mRNA and Ash1p in Yeast. Journal of Cell Biology, 153(2):307–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/2/307.

Leung:2004:QKA

[LGM^{+}04] Anthony Kar Lun Leung, Daniel Gerlich, Gail Miller, Carol Lyon, Yun Wah Lam, David Lleres, Nathalie Daigle, Joost Zomerdijk, Jan Ellenberg, and Angus I. Lamond. Quantitative kinetic analysis of nucleolar breakdown and reassembly during mitosis in live human cells. Journal of Cell Biology, 166(6):787–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/6/787.

Landesman:2000:GJC

[LGP00] Yosef Landesman, Daniel A. Goodenough, and David L. Paul. Gap Junctional Communication in the Early Xenopus Embryo. Journal of Cell Biology, 150(4):929–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
4/929.

LeStunff:2002:SPP

[LGRP^{+}02] Hervé Le Stunff, Ismael Galve-Roperh, Courtney Peterson, Sheldon Milstien, and Sarah Spiegel. Sphingosine-1-phosphate phosphohydrolase in regulation of sphingolipid metabolism and apoptosis. Journal of Cell Biology, 158 (6):1039–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/6/1039.

Lowe:2000:MPC

[LGW00] Martin Lowe, Nicholas K. Gonatas, and Graham Warren. The mitotic phosphorylation cycle of the cis-Golgi matrix protein Gm130. Journal of Cell Biology, 149(2):341–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
2/341.

Li:2003:USI

[LH03] Shyr-Jiann Li and Mark Hochstrasser. The Ulp1 SUMO isopeptidase. Journal of Cell Biology, 160(7):1069–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
7/1069.

Li:2002:MAR

[LHC^{+}02] Shaohua Li, David Harrison, Salvatore Carbonetto, Reinhard Fässler, Neil Smyth, David Edgar, and Peter D. Yurchenco. Matrix assembly, regulation, and survival functions of laminin and its receptors in embryonic stem cell differentiation. Journal of Cell Biology, 157(7):1279–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/157/
7/1279.

Lavoie:2002:VDC

[LHK02] Brigitte D. Lavoie, Eileen Hogan, and Douglas Koshland. In vivo dissection of the chromosome condensation machinery. Journal of Cell Biology, 156(5):805–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/5/805.

Lu:2004:TRA

[LHR04] Phoebe D. Lu, Heather P. Harding, and David Ron. Translation reinitiation at alternative open reading frames regulates gene expression in an integrated stress response. Journal of Cell Biology, 167(1):27–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/1/27.

Lorson:2000:LNC

[LHvdH00] Monique A. Lorson, H. Robert Horvitz, and Sander van den Heuvel. LIN-5 Is a Novel Component of the Spindle Apparatus Required for Chromosome Segregation and Cleavage Plane Specification in Caenorhabditis elegans. Journal of Cell Biology, 148(1):73–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/148/1/73.

Lindsay:2001:RBP

[LHW^{+}01] Mark E. Lindsay, James M. Holaska, Katie Welch, Bryce M. Paschal, and Ian G. Macara. Ran-binding protein 3 is a cofactor for Crm1-mediated nuclear protein export. Journal of Cell Biology, 153(7):1391–??, June 2001. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/7/1391.

Littleton:2000:GAM

[Lit00] J. Troy Littleton. A Genomic Analysis of Membrane Trafficking and Neurotransmitter Release in Drosophila. Journal of Cell Biology, 150(2):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/2/F77.

Lechler:2001:TTM

[LJK^{+}01] Terry Lechler, Gudrun A. Jonsdottir, Saskia K. Klee, David Pellman, and Rong Li. A two-tiered mechanism by which Cdc42 controls the localization and activation of an Arp2/ 3-activating motor complex in yeast. Journal of Cell Biology, 155(2):261–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/2/261.

Lee:2004:RPE

[LKH^{+}04] Jong-Hwan Lee, Tomoya Katakai, Takahiro Hara, Hiroyuki Gonda, Manabu Sugai, and Akira Shimizu. Roles of p–ERM and Rho–ROCK signaling in lymphocyte polarity and uropod formation. Journal of Cell Biology, 167(2):327–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
167/2/327.

Lee:2003:APR

[LKL^{+}03] Ming-Sum Lee, Shih-Chu Kao, Cynthia A. Lemere, Weiming Xia, Huang-Chun Tseng, Ying Zhou, Rachael Neve, Michael K. Ahlijanian, and Li-Huei Tsai. APP processing is regulated by cytoplasmic phosphorylation. Journal of Cell Biology, 163(1):83–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/1/83.

Li:2004:ABL

[LKLD04] Wenhui Li, Soo-Mi Kim, Joon Lee, and William G. Dunphy. Absence of BLM leads to accumulation of chromosomal DNA breaks during both unperturbed and disrupted S phases. Journal of Cell Biology, 165(6):801–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/6/801.

Lansbergen:2004:CCC

[LKM^{+}04] Gideon Lansbergen, Yulia Komarova, Mauro Modesti, Claire Wyman, Casper C. Hoogenraad, Holly V. Goodson, Régis P. Lemaitre, David N. Drechsel, Erik van Munster, Theodorus W. J. Gadella, Frank Grosveld, Niels Galjart, Gary G. Borisy, and Anna Akhmanova. Conformational changes in CLIP-170 regulate its binding to microtubules and dynactin localization. Journal of Cell Biology, 166(7):1003–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
7/1003.

Leung:2002:VAN

[LL02] Anthony K. L. Leung and Angus I. Lamond. In vivo analysis of NHPX reveals a novel nucleolar localization pathway involving a transient accumulation in splicing speckles. Journal of Cell Biology, 157(4):615–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/4/615.

LeBrasseur:2003:DCSa

[LL03a] Nicole LeBrasseur and Nicole LeBrasseur. Dying cells say come-hither. Journal of Cell Biology, 162(1):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/162/1/
13.3.

LeBrasseur:2003:GUP

[LL03b] Nicole LeBrasseur and Nicole LeBrasseur. Glutamate in unusual places. Journal of Cell Biology, 162(1):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/162/1/
13.2.

Lacazette:2003:NPM

[LLGB03] Eric Lacazette, Sophie Le Calvez, Nadesan Gajendran, and Hans Rudolf Brenner. A novel pathway for MuSK to induce key genes in neuromuscular synapse formation. Journal of Cell Biology, 161(4):727–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/4/727.

Li:2000:CRC

[LLH^{+}00] Hedong Li, Tin-Chung Leung, Stanley Hoffman, Janne Balsamo, and Jack Lilien. Coordinate Regulation of Cadherin and Integrin Function by the Chondroitin Sulfate Proteoglycan Neurocan. Journal of Cell Biology, 149(6):1275–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
6/1275.

Lehnert:2001:ASH

[LLH^{+}01] Lasse Lehnert, Markus M. Lerch, Yohei Hirai, Marie-Luise Kruse, Wolff Schmiegel, and Holger Kalthoff. Autocrine Stimulation of Human Pancreatic Duct-like Development by Soluble Isoforms of Epimorphin in Vitro. Journal of Cell Biology, 152(5):911–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/5/911.

Lane:2002:CMC

[LLP^{+}02] Jon D. Lane, John Lucocq, James Pryde, Francis A. Barr, Philip G. Woodman, Victoria J. Allan, and Martin Lowe. Caspase-mediated cleavage of the stacking protein GRASP65 is required for Golgi fragmentation during apoptosis. Journal of Cell Biology, 156(3):495–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/3/495.

Li:2000:ORD

[LM00a] Danxi Li and Randall J. Mrsny. Oncogenic Raf-1 Disrupts Epithelial Tight Junctions via Downregulation of Occludin. Journal of Cell Biology, 148(4):791–??, February 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/4/
791.

Li:2000:TPP

[LM00b] Hong-Sheng Li and Craig Montell. TRP and the PDZ Protein, Inad, Form the Core Complex Required for Retention of the Signalplex in Drosophila Photoreceptor Cells. Journal of Cell Biology, 150(6):1411–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/6/1411.

Le:2001:NRF

[LM01] Anh-Chi N. Le and Linda S. Musil. A novel role for FGF and extracellular signal–regulated kinase in gap junction– mediated intercellular communication in the lens. Journal of Cell Biology, 154(1):197–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/1/197.

Liu:2003:DPI

[LMDS^{+}03] Hongjun Liu, Michelle Mardahl-Dumesnil, Sean T. Sweeney, Cahir J. O’Kane, and Sanford I. Bernstein. Drosophila paramyosin is important for myoblast fusion and essential for myofibril formation. Journal of Cell Biology, 160(6): 899–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/6/899.

Liu:2004:R

[LMDS^{+}04] Hongjun Liu, Michelle Mardahl-Dumesnil, Sean T. Sweeney, Cahir J. O’Kane, and Sanford I. Bernstein. Retraction. Journal of Cell Biology, 167(4):783–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/4/783.

Labbe:2004:FPM

[LMG04] Jean-Claude Labbé, Erin K. McCarthy, and Bob Goldstein. The forces that position a mitotic spindle asymmetrically are tethered until after the time of spindle assembly. Journal of Cell Biology, 167(2):245–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/2/245.

Lacalle:2002:SSP

[LMGM^{+}02] Rosa Ana Lacalle, Emilia Mira, Concepción Gómez-Moutón, Sonia Jiménez-Baranda, Carlos Martínez-A., and Santos Mañes. Specific SHP-2 partitioning in raft domains triggers integrin-mediated signaling via Rho activation. Journal of Cell Biology, 157(2):277–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/2/277.

Lang:2002:SNP

[LMHJ02] Thorsten Lang, Martin Margittai, Helmut Hölzler, and Reinhard Jahn. SNAREs in native plasma membranes are active and readily form core complexes with endogenous and exogenous SNAREs. Journal of Cell Biology, 158(4):751–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
158/4/751.

Lusk:2002:KNP

[LMM^{+}02] C. Patrick Lusk, Taras Makhnevych, Marcello Marelli, John D. Aitchison, and Richard W. Wozniak. Karyopherins in nuclear pore biogenesis. Journal of Cell Biology, 159(2): 267–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/2/267.

Larson:2003:DMG

[LMVW03] Daniel R. Larson, Yu May Ma, Volker M. Vogt, and Watt W. Webb. Direct measurement of Gag–Gag interaction during retrovirus assembly with FRET and fluorescence correlation spectroscopy. Journal of Cell Biology, 162(7):1233– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/7/1233.

Lu:2000:MIE

[LMW^{+}00] Q. L. Lu, G. E. Morris, S. D. Wilton, T. Ly, O. V. Artem’yeva, P. Strong, and T. A. Partridge. Massive Idiosyncratic Exon Skipping Corrects the Nonsense Mutation in Dystrophic Mouse Muscle and Produces Functional Revertant Fibers by Clonal Expansion. Journal of Cell Biology, 148 (5):985–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/5/985.

Luhtala:2004:BCD

[LO04] Natalie Luhtala and Greg Odorizzi. Bro1 coordinates deubiquitination in the multivesicular body pathway by recruiting Doa4 to endosomes. Journal of Cell Biology, 166(5): 717–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/5/717.

Lee:2003:RLP

[LOC03] Wei-Lih Lee, Jessica R. Oberle, and John A. Cooper. The role of the lissencephaly protein Pac1 during nuclear migration in budding yeast. Journal of Cell Biology, 160(3): 355–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/3/355.

Lanoix:2001:SGR

[LOS^{+}01] Joel Lanoix, Joke Ouwendijk, Annika Stark, Edith Szafer, Dan Cassel, Kurt Dejgaard, Matthias Weiss, and Tommy Nilsson. Sorting of Golgi resident proteins into different subpopulations of COPI vesicles. Journal of Cell Biology, 155(7): 1199–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/7/1199.

Lindon:2004:OPA

[LP04a] Catherine Lindon and Jonathon Pines. Ordered proteolysis in anaphase inactivates Plk1 to contribute to proper mitotic exit in human cells. Journal of Cell Biology, 164(2): 233–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/2/233.

Lord:2004:UPR

[LP04b] Matthew Lord and Thomas D. Pollard. UCS protein Rng3p activates actin filament gliding by fission yeast myosin–II. Journal of Cell Biology, 167(2):315–??, October 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/167/2/
315.

Lin:2004:RCM

[LPL^{+}04] Yi Hsing Lin, Zee-Yong Park, Dayin Lin, Anar A. Brahmbhatt, Marie-Christine Rio, John R. Yates, and Richard L. Klemke. Regulation of cell migration and survival by focal adhesion targeting of Lasp-1. Journal of Cell Biology, 165 (3):421–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/3/421.

Li:2002:CRC

[LPPT^{+}02] Jian Li, Michel Pucéat, Carmen Perez-Terzic, Annabelle Mery, Kimitoshi Nakamura, Marek Michalak, Karl-Heinz Krause, and Marisa E. Jaconi. Calreticulin reveals a critical Ca^{2+} checkpoint in cardiac myofibrillogenesis. Journal of Cell Biology, 158(1):103–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/1/103.

Luo:2003:ADK

[LPT03] Bai Luo, Stephen M. Prescott, and Matthew K. Topham. Association of diacylglycerol kinase ζ with protein kinase C α. Journal of Cell Biology, 160(6):929–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/6/929.

Lee:2000:CIM

[LQPC^{+}00] Joon Yung Lee, Zhuqing Qu-Petersen, Baohong Cao, Shigemi Kimura, Ron Jankowski, James Cummins, Arvydas Usas, Charley Gates, Paul Robbins, Anton Wernig, and Johnny Huard. Clonal Isolation of Muscle–Derived Cells Capable of Enhancing Muscle Regeneration and Bone Healing. Journal of Cell Biology, 150(5):1085–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/5/1085.

Lawlor:2000:CCM

[LR00] Margaret A. Lawlor and Peter Rotwein. Coordinate Control of Muscle Cell Survival by Distinct Insulin-like Growth Factor Activated Signaling Pathways. Journal of Cell Biology, 151(6):1131–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/6/1131.

Lenz:2002:CCC

[LRA^{+}02] Johannes C. Lenz, H. Peter Reusch, Nadine Albrecht, Günter Schultz, and Michael Schaefer. Ca^{2+}-controlled competitive diacylglycerol binding of protein kinase C isoenzymes in living cells. Journal of Cell Biology, 159(2):291–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
2/291.

Lohez:2003:AMF

[LRB^{+}03] Olivier D. Lohez, Caroline Reynaud, Franck Borel, Paul R. Andreassen, and Robert L. Margolis. Arrest of mammalian fibroblasts in G1 in response to actin inhibition is dependent on retinoblastoma pocket proteins but not on p53. Journal of Cell Biology, 161(1):67–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/1/67.

Lamoureux:2002:MTC

[LRBH02] Phillip Lamoureux, Gordon Ruthel, Robert E. Buxbaum, and Steven R. Heidemann. Mechanical tension can specify axonal fate in hippocampal neurons. Journal of Cell Biology, 159(3):499–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/3/499.

Lenart:2003:NEB

[LRD^{+}03] Péter Lénárt, Gwénaël Rabut, Nathalie Daigle, Arthur R. Hand, Mark Terasaki, and Jan Ellenberg. Nuclear envelope breakdown in starfish oocytes proceeds by partial NPC disassembly followed by a rapidly spreading fenestration of nuclear membranes. Journal of Cell Biology, 160(7):1055–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
7/1055.

Lee:2002:CPK

[LRF^{+}02a] Paul Lee, Jie Rao, Albert Fliss, Emy Yang, Stephen Garrett, and Avrom J. Caplan. The Cdc37 protein kinase–binding domain is sufficient for protein kinase activity and cell viability. Journal of Cell Biology, 159(6):1051–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
6/1051.

Li:2002:AOB

[LRF^{+}02b] Xinmin Li, Marcos P. Rivas, Min Fang, Jennifer Marchena, Bharat Mehrotra, Anu Chaudhary, Li Feng, Glenn D. Prestwich, and Vytas A. Bankaitis. Analysis of oxysterol binding protein homologue Kes1p function in regulation of Sec14pdependent protein transport from the yeast Golgi complex.

Journal of Cell Biology, 157(1):63–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/1/63.

Laplantine:2002:FST

[LRS^{+}02] Emmanuel Laplantine, Ferdinand Rossi, Malika Sahni, Claudio Basilico, and David Cobrinik. FGF signaling targets the pRb-related p107 and p130 proteins to induce chondrocyte growth arrest. Journal of Cell Biology, 158(4):741–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
4/741.

Leonhardt:2000:DDR

[LRW^{+}00] Heinrich Leonhardt, Hans-Peter Rahn, Peter Weinzierl, Anje Sporbert, Thomas Cremer, Daniele Zink, and M. Cristina Cardoso. Dynamics of DNA Replication Factories in Living Cells. Journal of Cell Biology, 149(2):271–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/149/2/
271.

Liu:2004:PPC

[LRWB04] Hong Liu, Derek C. Radisky, Fei Wang, and Mina J. Bissell. Polarity and proliferation are controlled by distinct signaling pathways downstream of PI3-kinase in breast epithelial tumor cells. Journal of Cell Biology, 164(4):603–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
4/603.

Lander:2000:EFP

[LS00] Arthur D. Lander and Scott B. Selleck. The Elusive Functions of Proteoglycans. Journal of Cell Biology, 148(2): 227–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/2/227.

Lalli:2002:ART

[LS02] Giovanna Lalli and Giampietro Schiavo. Analysis of retrograde transport in motor neurons reveals common endocytic carriers for tetanus toxin and neurotrophin receptor p75^{NTR}. Journal of Cell Biology, 156(2):233–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/156/
2/233.

LeNegrate:2000:SPL

[LSA^{+}00] Gaëlle Le’Negrate, Eric Selva, Patrick Auberger, Bernard Rossi, and Paul Hofman. Sustained Polymorphonuclear Leukocyte Transmigration Induces Apoptosis in T84 Intestinal Epithelial Cells. Journal of Cell Biology, 150(6):1479– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/6/1479.

Lee:2001:PRC

[LSK01] Ethan Lee, Adrian Salic, and Marc W. Kirschner. Physiological regulation of β-catenin stability by Tcf3 and CK1ϵ. Journal of Cell Biology, 154(5):983–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/5/983.

Lopez-Soler:2001:RNL

[LSMS^{+}01] Reynold I. Lopez-Soler, Robert D. Moir, Timothy P. Spann, Reimer Stick, and Robert D. Goldman. A role for nuclear lamins in nuclear envelope assembly. Journal of Cell Biology, 154(1):61–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/1/61.

Leshchynska:2003:NCA

[LSMS03] Iryna Leshchyns’ka, Vladimir Sytnyk, Jon S. Morrow, and Melitta Schachner. Neural cell adhesion molecule (NCAM) association with PKCβ_{2} via βI spectrin is implicated in NCAM-mediated neurite outgrowth. Journal of Cell Biology, 161(3):625–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/3/625.

Lee:2002:DRD

[LSS^{+}02] Justine C. Lee, Olaf Schickling, Alexander H. Stegh, Robert G. Oshima, David Dinsdale, Gerald M. Cohen, and Marcus E. Peter. DEDD regulates degradation of intermediate filaments during apoptosis. Journal of Cell Biology, 158(6):1051–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/6/1051.

Lechler:2000:DIY

[LSSL00] Terry Lechler, Anna Shevchenko, Andrej Shevchenko, and Rong Li. Direct Involvement of Yeast Type I Myosins in Cdc42-Dependent Actin Polymerization. Journal of Cell Biology, 148(2):363–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/2/363.

List:2003:LPP

[LSW^{+}03] Karin List, Roman Szabo, Philip W. Wertz, Julie Segre, Christian C. Haudenschild, Soo-Youl Kim, and Thomas H. Bugge. Loss of proteolytically processed filaggrin caused by epidermal deletion of Matriptase/MT-SP1. Journal of Cell Biology, 163(4):901–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/4/901.

Lampe:2000:PCS

[LTB^{+}00] Paul D. Lampe, Erica M. TenBroek, Janis M. Burt, Wendy E. Kurata, Ross G. Johnson, and Alan F. Lau. Phosphorylation of Connexin43 on Serine368 by Protein Kinase C Regulates Gap Junctional Communication. Journal of Cell Biology, 149 (7):1503–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/7/1503.

Lindemann:2001:APM

[LTD^{+}01] Stephan Lindemann, Neal D. Tolley, Dan A. Dixon, Thomas M. McIntyre, Stephen M. Prescott, Guy A. Zimmerman, and Andrew S. Weyrich. Activated platelets mediate inflammatory signaling by regulated interleukin 1β synthesis. Journal of Cell Biology, 154(3):485–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/154/3/
485.

Lefebvre:2002:MSS

[LTD^{+}02] Christophe Lefebvre, M. Emilie Terret, Alexandre Djiane, Pascale Rassinier, Bernard Maro, and Marie-Hélène Verlhac. Meiotic spindle stability depends on MAPK-interacting and spindle-stabilizing protein (MISS), a new MAPK substrate. Journal of Cell Biology, 157(4):603–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/4/603.

Liu:2001:OCO

[LTF^{+}01] Wenguang Liu, Satoru Toyosawa, Tatsuya Furuichi, Naoko Kanatani, Carolina Yoshida, Yang Liu, Miki Himeno, Satoru Narai, Akira Yamaguchi, and Toshihisa Komori. Overexpression of Cbfa1 in osteoblasts inhibits osteoblast maturation and causes osteopenia with multiple fractures. Journal of Cell Biology, 155(1):157–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/1/157.

Larsen:2002:RPF

[LUB^{+}02] Elaine C. Larsen, Takehiko Ueyama, Pamela M. Brannock, Yasuhito Shirai, Naoaki Saito, Christer Larsson, Daniel Loegering, Peter B. Weber, and Michelle R. Lennartz. A role for PKC-ϵ in FcγR-mediated phagocytosis by RAW 264.7 cells. Journal of Cell Biology, 159(6):939–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/6/
939.

Luo:2004:IFA

[LVD^{+}04] Jianying Luo, Elizabeth A. Vallen, Christopher Dravis, Serguei E. Tcheperegine, Becky Drees, and Erfei Bi. Identification and functional analysis of the essential and regulatory light chains of the only type II myosin Myo1p in Saccharomyces cerevisiae. Journal of Cell Biology, 165(6):843–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
165/6/843.

Lane:2001:ACC

[LVWA01] Jon D. Lane, Maïlys A. S. Vergnolle, Philip G. Woodman, and Victoria J. Allan. Apoptotic Cleavage of Cytoplasmic Dynein Intermediate Chain and P150 ^{Glued} Stops Dynein– Dependent Membrane Motility. Journal of Cell Biology, 153 (7):1415–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/7/1415.

Lecuit:2000:PIN

[LW00] Thomas Lecuit and Eric Wieschaus. Polarized Insertion of New Membrane from a Cytoplasmic Reservoir during Cleavage of the Drosophila Embryo. Journal of Cell Biology, 150 (4):849–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/4/849.

Lin:2003:TPC

[LW03] Joseph Lin and Arthur Weiss. The tyrosine phosphatase CD148 is excluded from the immunologic synapse and downregulates prolonged T cell signaling. Journal of Cell Biology, 162(4):673–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/4/673.

Lin:2001:VSP

[LWCKL01] Rui Lin, Bonnie J. Warn-Cramer, Wendy E. Kurata, and Alan F. Lau. v-Src phosphorylation of connexin 43 on Tyr247 and Tyr265 disrupts gap junctional communication. Journal of Cell Biology, 154(4):815–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/4/815.

Laukaitis:2001:DDI

[LWDH01] Christina M. Laukaitis, Donna J. Webb, Karen Donais, and Alan F. Horwitz. Differential Dynamics of α5 Integrin, Paxillin, and α-Actinin during Formation and Disassembly of Adhesions in Migrating Cells. Journal of Cell Biology, 153 (7):1427–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/7/1427.

Lecanda:2000:CDC

[LWS^{+}00] Fernando Lecanda, Pamela M. Warlow, Sharmin Sheikh, Federico Furlan, Thomas H. Steinberg, and Roberto Civitelli. Connexin43 Deficiency Causes Delayed Ossification, Craniofacial Abnormalities, and Osteoblast Dysfunction. Journal of Cell Biology, 151(4):931–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/4/931.

Li:2003:MCR

[LWZ03] Hoi Yeung Li, Denis Wirtz, and Yixian Zheng. A mechanism of coupling RCC1 mobility to RanGTP production on the chromatin in vivo. Journal of Cell Biology, 160(5): 635–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/5/635.

Lechardeur:2004:CND

[LXL04a] Delphine Lechardeur, Ming Xu, and Gergely L. Lukacs. Contrasting nuclear dynamics of the caspase-activated DNase (CAD) in dividing and apoptotic cells. Journal of Cell Biology, 167(5):851–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/5/851.

Lee:2004:API

[LXL^{+}04b] Jiunn-Tay Lee, Jan Xu, Jin-Moo Lee, Grace Ku, Xianlin Han, Ding-I Yang, Shawei Chen, and Chung Y. Hsu. Amyloidβ peptide induces oligodendrocyte death by activating the neutral sphingomyelinase–ceramide pathway. Journal of Cell Biology, 164(1):123–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/1/123.

Losada:2000:ICS

[LYKH00] Ana Losada, Tomoki Yokochi, Ryuji Kobayashi, and Tatsuya Hirano. Identification and characterization of Sa/Scc3p subunits in the Xenopus and human cohesin complexes. Journal of Cell Biology, 150(3):405–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/3/405.

Liang:2004:NFM

[LYL^{+}04] Yun Liang, Wei Yu, Yan Li, Zhenye Yang, Xiumin Yan, Qiongping Huang, and Xueliang Zhu. Nudel functions in membrane traffic mainly through association with Lis1 and cytoplasmic dynein. Journal of Cell Biology, 164(4):557– ??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/4/557.

Lord:2000:CCP

[LYMC00] Matthew Lord, Melody C. Yang, Michelle Mischke, and John Chant. Cell Cycle Programs of Gene Expression Control Morphogenetic Protein Localization. Journal of Cell Biology, 151(7):1501–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/7/1501.

Lampugnani:2003:CIV

[LZC^{+}03] Maria Grazia Lampugnani, Adriana Zanetti, Monica Corada, Takamune Takahashi, Giovanna Balconi, Ferruccio Breviario, Fabrizio Orsenigo, Anna Cattelino, Rolf Kemler, Thomas O. Daniel, and Elisabetta Dejana. Contact inhibition of VEGFinduced proliferation requires vascular endothelial cadherin, β-catenin, and the phosphatase DEP-1/CD148. Journal of Cell Biology, 161(4):793–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/4/793.

Leung:2001:BL

[LZPL01] Conrad L. Leung, Min Zheng, Susan M. Prater, and Ronald K. H. Liem. The BPAG1 locus. Journal of Cell Biology, 154 (4):691–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/4/691.

Loomis:2003:ECL

[LZS^{+}03] Patricia A. Loomis, Lili Zheng, Gabriella Sekerková, Benjarat Changyaleket, Enrico Mugnaini, and James R. Bartles. Espin cross-links cause the elongation of microvillus-type parallel actin bundles in vivo. Journal of Cell Biology, 163(5): 1045–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/5/1045.

Machado:2000:DT

[MA00] Cristina Machado and Deborah J. Andrew. D-titin. Journal of Cell Biology, 151(3):639–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/3/639.

Martinez-Arca:2000:RTN

[MAAZ^{+}00] Sonia Martinez-Arca, Philipp Alberts, Ahmed Zahraoui, Daniel Louvard, and Thierry Galli. Role of Tetanus Neurotoxin Insensitive Vesicle–Associated Membrane Protein (Ti– Vamp) in Vesicular Transport Mediating Neurite Outgrowth. Journal of Cell Biology, 149(4):889–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/4/889.

Machesky:2000:TTM

[Mac00] Laura M. Machesky. The Tails of Two Myosins. Journal of Cell Biology, 148(2):219–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/2/219.

Molenaar:2004:PRR

[MAG^{+}04] Chris Molenaar, Abadir Abdulle, Aarti Gena, Hans J. Tanke, and Roeland W. Dirks. Poly(a)^{+} RNAs roam the cell nucleus and pass through speckle domains in transcriptionally active and inactive cells. Journal of Cell Biology, 165(2):191–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
165/2/191.

Manning:2004:BAS

[Man04a] Brendan D. Manning. Balancing Akt with S6K. Journal of Cell Biology, 167(3):399–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/3/399.

Manning:2004:C

[Man04b] Brendan D. Manning. Correction. Journal of Cell Biology, 167(6):1255–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/6/1255.

Maraia:2001:PTN

[Mar01] Richard J. Maraia. La Protein and the Trafficking of Nascent RNA Polymerase III Transcripts. Journal of Cell Biology, 153(4):??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/4/F13.

Mandato:2001:CPC

[MB01] Craig A. Mandato and William M. Bement. Contraction and polymerization cooperate to assemble and close actomyosin rings around Xenopus oocyte wounds. Journal of Cell Biology, 154(4):785–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/4/785.

Maddox:2003:RRC

[MB03a] Amy Shaub Maddox and Keith Burridge. RhoA is required for cortical retraction and rigidity during mitotic cell rounding. Journal of Cell Biology, 160(2):255–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/2/
255.

Matter:2003:HB

[MB03b] Karl Matter and Maria S. Balda. Holey barrier. Journal of Cell Biology, 161(3):459–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/3/459.

Mythreye:2003:DKP

[MB03c] Karthikeyan Mythreye and Kerry S. Bloom. Differential kinetochore protein requirements for establishment versus propagation of centromere activity in Saccharomyces cerevisiae. Journal of Cell Biology, 160(6):833–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/6/833.

Martina:2001:S

[MBAB01] José A. Martina, Cecilia J. Bonangelino, Rubén C. Aguilar, and Juan S. Bonifacino. Stonin 2. Journal of Cell Biology, 153(5):1111–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/5/1111.

Meier:2002:ATM

[MBH^{+}02] Oliver Meier, Karin Boucke, Silvija Vig Hammer, Stephan Keller, Robert P. Stidwill, Silvio Hemmi, and Urs F. Greber. Adenovirus triggers macropinocytosis and endosomal leakage together with its clathrin-mediated uptake. Journal of Cell Biology, 158(6):1119–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/6/1119.

Martin:2003:TCS

[MBLCE03] Doris Martin, Marianne Brown-Luedi, and Ruth Chiquet-Ehrismann. Tenascin-C signaling through induction of 14-3-3 tau. Journal of Cell Biology, 160(2):171–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
2/171.

Melixetian:2004:LGI

[MBM^{+}04] Marina Melixetian, Andrea Ballabeni, Laura Masiero, Patrizia Gasparini, Raffaella Zamponi, Jiri Bartek, Jiri Lukas, and Kristian Helin. Loss of Geminin induces rereplication in the presence of functional p53. Journal of Cell Biology, 165 (4):473–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/4/473.

Martin-Belmonte:2003:MRC

[MBMMA^{+}03] Fernando Martín-Belmonte, José A. Martínez-Menárguez, Juan F. Aranda, José Ballesta, María C. de Marco, and Miguel A. Alonso. MAL regulates clathrin-mediated endocytosis at the apical surface of Madin–Darby canine kidney cells. Journal of Cell Biology, 163(1):155–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
1/155.

Mironov:2001:SCP

[MBN^{+}01] Alexander A. Mironov, Galina V. Beznoussenko, Paolo Nicoziani, Oliviano Martella, Alvar Trucco, Hee-Seok Kweon, Daniele Di Giandomenico, Roman S. Polishchuk, Aurora Fusella, Pietro Lupetti, Eric G. Berger, Willie J. C. Geerts, Abraham J. Koster, Koert N. J. Burger, and Alberto Luini. Small cargo proteins and large aggregates can traverse the Golgi by a common mechanism without leaving the lumen of cisternae. Journal of Cell Biology, 155(7):1225–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
7/1225.

Marshall:2001:RAP

[MBS^{+}01] John G. Marshall, James W. Booth, Vuk Stambolic, Tak Mak, Tamas Balla, Alan D. Schreiber, Tobias Meyer, and Sergio Grinstein. Restricted accumulation of phosphatidylinositol 3-kinase products in a plasmalemmal subdomain during Fcγ receptor-mediated phagocytosis. Journal of Cell Biology, 153(7):1369–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/7/1369.

Mansukhani:2000:SFG

[MBSB00] Alka Mansukhani, Paola Bellosta, Malika Sahni, and Claudio Basilico. Signaling by Fibroblast Growth Factors (Fgf) and Fibroblast Growth Factor Receptor 2 (Fgfr2)–Activating Mutations Blocks Mineralization and Induces Apoptosis in Osteoblasts. Journal of Cell Biology, 149(6):1297–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
6/1297.

Motley:2003:CME

[MBSR03] Alison Motley, Nicholas A. Bright, Matthew N. J. Seaman, and Margaret S. Robinson. Clathrin-mediated endocytosis in AP-2–depleted cells. Journal of Cell Biology, 162(5):909– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/5/909.

Marignani:2001:VRC

[MC01] Paola A. Marignani and Christopher L. Carpenter. Vav2 is required for cell spreading. Journal of Cell Biology, 154 (1):177–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/1/177.

Mori:2002:TAS

[MC02] Hiroki Mori and Kenneth Cline. A twin arginine signal peptide and the pH gradient trigger reversible assembly of the thylakoid Δ pH/Tat translocase. Journal of Cell Biology, 157 (2):205–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/2/205.

Myster:2003:DPP

[MCA^{+}03] Steven H. Myster, Robert Cavallo, Charles T. Anderson, Donald T. Fox, and Mark Peifer. Drosophila p120catenin plays a supporting role in cell adhesion but is not an essential adherens junction component. Journal of Cell Biology, 160(3):433–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/3/433.

Moorthy:2000:CEG

[MCB00a] Suraj Moorthy, Lihsia Chen, and Vann Bennett. Caenorhabditis elegans β-G spectrin is dispensable for establishment of epithelial polarity, but essential for muscular and neuronal function. Journal of Cell Biology, 149(4):915–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
4/915.

Myers:2000:HBP

[MCB00b] Connie Myers, Aubri Charboneau, and Nancy Boudreau. Homeobox B3 Promotes Capillary Morphogenesis and Angiogenesis. Journal of Cell Biology, 148(2):343–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
2/343.

Martinez-Campos:2004:DPL

[MCBB^{+}04] Maruxa Martinez-Campos, Renata Basto, James Baker, Maurice Kernan, and Jordan W. Raff. The Drosophila pericentrin-like protein is essential for cilia/flagella function, but appears to be dispensable for mitosis. Journal of Cell Biology, 165(5):673–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/5/673.

McFadden:2000:SC

[McF00] Geoffrey I. McFadden. Skeletons in the Closet. Journal of Cell Biology, 151(4):??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/4/F19.

Mu:2002:IVM

[MCF^{+}02] Dezhi Mu, Stephanie Cambier, Lars Fjellbirkeland, Jody L. Baron, John S. Munger, Hisaaki Kawakatsu, Dean Sheppard, V. Courtney Broaddus, and Stephen L. Nishimura. The integrin αvβ8 mediates epithelial homeostasis through MT1-MMP-dependent activation of TGF-β1. Journal of Cell Biology, 157(3):493–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/3/493.

Mal:2000:PRP

[MCG^{+}00] Asoke Mal, Debasis Chattopadhyay, Mrinal K. Ghosh, Randy Y. C. Poon, Tony Hunter, and Marian L. Harter. P21 and Retinoblastoma Protein Control the Absence of DNA Replication in Terminally Differentiated Muscle Cells. Journal of Cell Biology, 149(2):281–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/149/2/281.

Muro:2003:RSF

[MCG^{+}03] Andrés F. Muro, Anil K. Chauhan, Srecko Gajovic, Alessandra Iaconcig, Fabiola Porro, Giorgio Stanta, and Francisco E. Baralle. Regulated splicing of the fibronectin EDA exon is essential for proper skin wound healing and normal lifespan. Journal of Cell Biology, 162(1):149–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/1/149.

Martinez:2000:SVR

[MCH^{+}00] Iñigo Martinez, Sabyasachi Chakrabarti, Turid Hellevik, Jennifer Morehead, Kimberly Fowler, and Norma W. Andrews. Synaptotagmin VII Regulates Ca^{2+} -Dependent Exocytosis of Lysosomes in Fibroblasts. Journal of Cell Biology, 148(6): 1141–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/6/1141.

McIntosh:2001:EMC

[McI01] J. Richard McIntosh. Electron Microscopy of Cells. Journal of Cell Biology, 153(6):0–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/6/F25.

McCullough:2004:AEA

[MCU04] John McCullough, Michael J. Clague, and Sylvie Urbé. AMSH is an endosome-associated ubiquitin isopeptidase. Journal of Cell Biology, 166(4):487–??, August 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/166/4/
487.

Mardahl-Dumesnil:2001:TFE

[MDF01] Michelle Mardahl-Dumesnil and Velia M. Fowler. Thin filaments elongate from their pointed ends during myofibril assembly in Drosophila indirect flight muscle. Journal of Cell Biology, 155(6):1043–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/6/1043.

McCallum:2000:ICT

[MDJF00] Christine D. McCallum, Hung Do, Arthur E. Johnson, and Judith Frydman. The Interaction of the Chaperonin Tailless Complex Polypeptide 1 (Tcp1) Ring Complex (Tric) with Ribosome–Bound Nascent Chains Examined Using Photo– Cross–Linking. Journal of Cell Biology, 149(3):591–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
3/591.

Marcus:2002:MAG

[MDP02] Jill Marcus, Jeffrey L. Dupree, and Brian Popko. Myelinassociated glycoprotein and myelin galactolipids stabilize developing axo-glial interactions. Journal of Cell Biology, 156 (3):567–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/3/567.

Marsh:2003:SNR

[MDQ^{+}03] H. Nicholas Marsh, Catherine I. Dubreuil, Celia Quevedo, Anna Lee, Marta Majdan, Gregory S. Walsh, Sharon Hausdorff, Farid Arab Said, Olga Zoueva, Maya Kozlowski, Katherine Siminovitch, Benjamin G. Neel, Freda D. Miller, and David R. Kaplan. SHP-1 negatively regulates neuronal survival by functioning as a TrkA phosphatase. Journal of Cell Biology, 163(5):999–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/5/999.

Michalik:2001:ISW

[MDT^{+}01] Liliane Michalik, Béatrice Desvergne, Nguan Soon Tan, Sharmila Basu-Modak, Pascal Escher, Jennifer Rieusset, Jeffrey M. Peters, Gürkan Kaya, Frank J. Gonzalez, Jozsef Zakany, Daniel Metzger, Pierre Chambon, Denis Duboule, and Walter Wahli. Impaired skin wound healing in peroxisome proliferator–activated receptor (PPAR) α and PPAR β mutant mice. Journal of Cell Biology, 154(4):799–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/154/
4/799.

Morgan:2004:RTP

[MDW^{+}04] Jennifer R. Morgan, Gilbert Di Paolo, Hauke Werner, Valentina A. Shchedrina, Marc Pypaert, Vincent A. Pieribone, and Pietro De Camilli. A role for talin in presynaptic function. Journal of Cell Biology, 167(1):43–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
1/43.

Murray:2000:RPC

[ME00] Patricia Murray and David Edgar. Regulation of Programmed Cell Death by Basement Membranes in Embryonic Development. Journal of Cell Biology, 150(5):1215– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/5/1215.

Martins:2003:HLM

[MEFC03] Sandra Martins, Sissel Eikvar, Kazuhiro Furukawa, and Philippe Collas. HA95 and LAP2β mediate a novel chromatin–nuclear envelope interaction implicated in initiation of DNA replication. Journal of Cell Biology, 160(2): 177–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/2/177.

Meilhac:2004:OCC

[MEK^{+}04] Sigolène M. Meilhac, Milan Esner, Michel Kerszberg, Julie E. Moss, and Margaret E. Buckingham. Oriented clonal cell growth in the developing mouse myocardium underlies cardiac morphogenesis. Journal of Cell Biology, 164(1):97– ??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/1/97.

Mellman:2000:GCC

[Mel00a] Ira Mellman. Genomics Comes to Cell Biology. Journal of Cell Biology, 150(2):1–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/2/F21.

Mellman:2000:NYL

[Mel00b] Ira Mellman. A New Year’s letter from The Editor. Journal of Cell Biology, 148(1):??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/148/1/NP.2.

Mellman:2000:QVP

[Mel00c] Ira Mellman. Quo vadis: Polarized membrane recycling in motility and phagocytosis. Journal of Cell Biology, 149 (3):529–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/3/529.

Mellman:2001:CBJ

[Mel01] Ira Mellman. Cell biology’s journal gets a new look. Journal of Cell Biology, 154(1):9–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/1/9.

Mellman:2004:AES

[Mel04a] Ira Mellman. Another evolutionary step for the JCB. Journal of Cell Biology, 167(1):17–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/1/17.

Mellman:2004:PRA

[Mel04b] Ira Mellman. Providing realistic access. Journal of Cell Biology, 165(1):19–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/1/19.

Marsden:2004:BRA

[MEV^{+}04] Vanessa S. Marsden, Paul G. Ekert, Mark Van Delft, David L. Vaux, Jerry M. Adams, and Andreas Strasser. Bcl-2-regulated apoptosis and cytochrome c release can occur independently of both caspase-2 and caspase-9. Journal of Cell Biology, 165(6):775–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/6/775.

Mahmoud:2001:MHC

[MF01a] Sahar F. Mahmoud and Clare Fewtrell. Microdomains of High Calcium Are Not Required for Exocytosis in Rbl-2h3 Mucosal Mast Cells. Journal of Cell Biology, 153(2):339–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/2/339.

Meletis:2001:BCG

[MF01b] Konstantinos Meletis and Jonas Frisén. Have the bloody cells gone to our heads? Journal of Cell Biology, 155(5): 699–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/5/699.

Mullershausen:2003:DAP

[MFF^{+}03] Florian Mullershausen, Andreas Friebe, Robert Feil, W. Joseph Thompson, Franz Hofmann, and Doris Koesling. Direct activation of PDE5 by cGMP. Journal of Cell Biology, 160 (5):719–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/5/719.

Mechtersheimer:2001:ESL

[MGAL^{+}01] Sabine Mechtersheimer, Paul Gutwein, Nancy Agmon-Levin, Alexander Stoeck, Matthias Oleszewski, Svenja Riedle, Rolf Postina, Falk Fahrenholz, Mina Fogel, Vance Lemmon, and Peter Altevogt. Ectodomain shedding of L1 adhesion molecule promotes cell migration by autocrine binding to integrins. Journal of Cell Biology, 155(4):661–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
4/661.

Mahnke:2000:DCR

[MGL^{+}00] Karsten Mahnke, Ming Guo, Sena Lee, Homero Sepulveda, Suzy L. Swain, Michel Nussenzweig, and Ralph M. Steinman. The Dendritic Cell Receptor for Endocytosis, Dec-205, Can Recycle and Enhance Antigen Presentation via Major Histocompatibility Complex Class II–Positive Lysosomal Compartments. Journal of Cell Biology, 151(3):673–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
3/673.

Miller:2000:NFO

[MGLPM00] Brent J. Miller, Elisabeth Georges-Labouesse, Paul Primakoff, and Diana G. Myles. Normal Fertilization Occurs with Eggs Lacking the Integrin α6β1 and Is Cd9-Dependent. Journal of Cell Biology, 149(6):1289–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/6/1289.

Moore:2003:FIC

[MGMH03] Misty Moore, Robyn L. Goforth, Hiroki Mori, and Ralph Henry. Functional interaction of chloroplast SRP/FtsY with the ALB3 translocase in thylakoids. Journal of Cell Biology, 162(7):1245–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/7/1245.

Molinari:2002:SAM

[MGP^{+}02a] Maurizio Molinari, Carmela Galli, Verena Piccaluga, Michel Pieren, and Paolo Paganetti. Sequential assistance of molecular chaperones and transient formation of covalent complexes during protein degradation from the ER. Journal of Cell Biology, 158(2):247–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/2/247.

Morgan:2002:MCP

[MGP^{+}02b] Jennifer E. Morgan, Jacqueline G. Gross, Charles N. Pagel, Jonathan R. Beauchamp, Ariberto Fassati, Adrian J. Thrasher, James P. Di Santo, Ivan B. Fisher, Xu Shiwen, David J. Abraham, and Terence A. Partridge. Myogenic cell proliferation and generation of a reversible tumorigenic phenotype are triggered by preirradiation of the recipient site. Journal of Cell Biology, 157(4):693–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/4/693.

Mishra-Gorur:2002:RDP

[MGRPVAT02] Ketu Mishra-Gorur, Matthew D. Rand, Beatriz Perez-Villamil, and Spyros Artavanis-Tsakonas. Down-regulation of Delta by proteolytic processing. Journal of Cell Biology, 159(2):313–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/2/313.

Madesh:2001:VDP

[MH01] Muniswamy Madesh and György Hajnóczky. VDACdependent permeabilization of the outer mitochondrial membrane by superoxide induces rapid and massive cytochrome c release. Journal of Cell Biology, 155(6):1003–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
6/1003.

Machaca:2002:IMP

[MH02] Khaled Machaca and Shirley Haun. Induction of maturationpromoting factor during Xenopus oocyte maturation uncouples Ca^{2+} store depletion from store-operated Ca^{2+} entry. Journal of Cell Biology, 156(1):75–??, January 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/1/
75.

Middleton:2000:CIN

[MHE^{+}00] Gayle Middleton, Makoto Hamanoue, Yasushi Enokido, Sean Wyatt, Diane Pennica, Ellis Jaffray, Ronald T. Hay, and Alun M. Davies. Cytokine-induced Nuclear Factor Kappa B Activation Promotes the Survival of Developing Neurons. Journal of Cell Biology, 148(2):325–??, January 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/2/

Maske:2003:CTI

[MHH^{+}03] Christopher P. Maske, Michael S. Hollinshead, Niall C. Higbee, Martin O. Bergo, Stephen G. Young, and David J. Vaux. A carboxyl-terminal interaction of lamin B1 is dependent on the CAAX endoprotease Rce1 and carboxymethylation. Journal of Cell Biology, 162(7):1223–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
7/1223.

Monzen:2001:STT

[MHK^{+}01] Koshiro Monzen, Yukio Hiroi, Sumiyo Kudoh, Hiroshi Akazawa, Toru Oka, Eiki Takimoto, Doubun Hayashi, Toru Hosoda, Masahiro Kawabata, Kohei Miyazono, Shunsuke Ishii, Yoshio Yazaki, Ryozo Nagai, and Issei Komuro. Smads, Tak1, and Their Common Target Atf-2 Play a Critical Role in Cardiomyocyte Differentiation. Journal of Cell Biology, 153 (4):687–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/4/687.

Markevich:2004:SSB

[MHK04] Nick I. Markevich, Jan B. Hoek, and Boris N. Kholodenko. Signaling switches and bistability arising from multisite phosphorylation in protein kinase cascades. Journal of Cell Biology, 164(3):353–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/3/353.

Murata-Hori:2002:BMA

[MHlW02] Maki Murata-Hori and Yu li Wang. Both midzone and astral microtubules are involved in the delivery of cytokinesis signals. Journal of Cell Biology, 159(1):45–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
1/45.

Moores:2003:RKK

[MHNSM03] Carolyn A. Moores, Mohammad Hekmat-Nejad, Roman Sakowicz, and Ronald A. Milligan. Regulation of KinI kinesin ATPase activity by binding to the microtubule lattice. Journal of Cell Biology, 163(5):963–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/5/963.

Merdes:2000:FSP

[MHS^{+}00] Andreas Merdes, Rebecca Heald, Kumiko Samejima, William C. Earnshaw, and Don W. Cleveland. Formation of spindle poles by dynein/dynactin-dependent transport of Numa. Journal of Cell Biology, 149(4):851–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/149/4/
851.

Micheva:2001:RPP

[MHS01] Kristina D. Micheva, Ronald W. Holz, and Stephen J. Smith. Regulation of presynaptic phosphatidylinositol 4,5biphosphate by neuronal activity. Journal of Cell Biology, 154(2):355–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/2/355.

Malathi:2004:MPS

[MHT^{+}04] Krishnamurthy Malathi, Katsumi Higaki, Arthur H. Tinkelenberg, Dina A. Balderes, Dorca Almanzar-Paramio, Lisa J. Wilcox, Naz Erdeniz, Francis Redican, Mahajabeen Padamsee, Ying Liu, Sohail Khan, Frederick Alcantara, Eugene D. Carstea, Jill A. Morris, and Stephen L. Sturley. Mutagenesis of the putative sterol-sensing domain of yeast Niemann Pick C–related protein reveals a primordial role in subcellular sphingolipid distribution. Journal of Cell Biology, 164(4):547–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/4/547.

Meixner:2000:MRA

[MHW^{+}00] Arabella Meixner, Silke Haverkamp, Heinz Wässle, Susanne Führer, Johann Thalhammer, Nina Kropf, Reginald E. Bittner, Hans Lassmann, Gerhard Wiche, and Friedrich Propst. Map1b Is Required for Axon Guidance and Is Involved in the Development of the Central and Peripheral Nervous System. Journal of Cell Biology, 151(6):1169–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
6/1169.

Miller:2002:EAC

[Mil02] Kathryn G. Miller. Extending the Arp2/3 complex and its regulation beyond the leading edge. Journal of Cell Biology, 156(4):591–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/4/591.

Miner:2001:MS

[Min01] Jeffrey H. Miner. Mystery solved. Journal of Cell Biology, 154(2):257–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/2/257.

Misteli:2001:CSO

[Mis01] Tom Misteli. The concept of self-organization in cellular architecture. Journal of Cell Biology, 155(2):181–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
2/181.

Mitchison:2001:PCO

[Mit01] T. J. Mitchison. Psychosine, Cytokinesis, and Orphan Receptors. Journal of Cell Biology, 153(2):0–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/153/2/
F1.

Mosammaparast:2001:NIH

[MJG^{+}01] Nima Mosammaparast, Kelley R. Jackson, Yurong Guo, Cynthia J. Brame, Jeffrey Shabanowitz, Donald F. Hunt, and Lucy F. Pemberton. Nuclear Import of Histone H2a and H2b Is Mediated by a Network of Karyopherins. Journal of Cell Biology, 153(2):251–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/2/251.

Mandon:2003:DRR

[MJG03] Elisabet C. Mandon, Ying Jiang, and Reid Gilmore. Dual recognition of the ribosome and the signal recognition particle by the SRP receptor during protein targeting to the endoplasmic reticulum. Journal of Cell Biology, 162(4): 575–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/4/575.

Malkus:2002:CSS

[MJS02] Per Malkus, Feng Jiang, and Randy Schekman. Concentrative sorting of secretory cargo proteins into COPII-coated vesicles. Journal of Cell Biology, 159(6):915–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
6/915.

Messerschmitt:2003:IMP

[MJV^{+}03] Marlies Messerschmitt, Stefan Jakobs, Frank Vogel, Stefan Fritz, Kai Stefan Dimmer, Walter Neupert, and Benedikt Westermann. The inner membrane protein Mdm33 controls mitochondrial morphology in yeast. Journal of Cell Biology, 160(4):553–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/4/553.

Murray:2004:PKC

[MJY^{+}04] Nicole R. Murray, Lee Jamieson, Wangsheng Yu, Jie Zhang, Yesim Gökmen-Polar, Deborah Sier, Panos Anastasiadis, Zoran Gatalica, E. Aubrey Thompson, and Alan P. Fields. Protein kinase Cι is required for Ras transformation and colon carcinogenesis in vivo. Journal of Cell Biology, 164 (6):797–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/6/797.

Middendorp:2000:RCC

[MKA^{+}00] Sandrine Middendorp, Thomas Küntziger, Yann Abraham, Simon Holmes, Nicole Bordes, Michel Paintrand, Anne Paoletti, and Michel Bornens. A Role for Centrin 3 in Centrosome Reproduction. Journal of Cell Biology, 148(3):405– ??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/3/405.

Mariotti:2001:ERS

[MKD^{+}01] Agnese Mariotti, Paul A. Kedeshian, Michael Dans, Anna Maria Curatola, Laurent Gagnoux-Palacios, and Filippo G. Giancotti. EGF-r signaling through Fyn kinase disrupts the function of integrin α6β4 at hemidesmosomes. Journal of Cell Biology, 155(3):447–??, October 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/155/3/
447.

Mollinari:2002:PMB

[MKJ^{+}02] Cristiana Mollinari, Jean-Philippe Kleman, Wei Jiang, Guy Schoehn, Tony Hunter, and Robert L. Margolis. PRC1 is a microtubule binding and bundling protein essential to maintain the mitotic spindle midzone. Journal of Cell Biology, 157 (7):1175–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/7/1175.

McNiven:2000:RIB

[MKK^{+}00a] Mark A. McNiven, Leung Kim, Eugene W. Krueger, James D. Orth, Hong Cao, and Tai Wai Wong. Regulated Interactions between Dynamin and the Actin–Binding Protein Cortactin Modulate Cell Shape. Journal of Cell Biology, 151(1):187–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/1/187.

Miyazaki:2000:RRE

[MKK^{+}00b] Tsuyoshi Miyazaki, Hideki Katagiri, Yumi Kanegae, Hiroshi Takayanagi, Yasuhiro Sawada, Aiichiro Yamamoto, Mattew P. Pando, Tomoichiro Asano, Inder M. Verma, Hiromi Oda, Kozo Nakamura, and Sakae Tanaka. Reciprocal Role of ERK and Nf-κ b Pathways in Survival and Activation of Osteoclasts. Journal of Cell Biology, 148(2):333–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
2/333.

MacDonald:2004:NNA

[MKM04] James I. S. MacDonald, Chris J. Kubu, and Susan O. Meakin. Nesca, a novel adapter, translocates to the nuclear envelope and regulates neurotrophin-induced neurite outgrowth. Journal of Cell Biology, 164(6):851–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/6/851.

More:2001:TAN

[MKR01] Margret I. Moré, Frank-P. Kirsch, and Fritz G. Rathjen. Targeted ablation of NrCAM or ankyrin–B results in disorganized lens fibers leading to cataract formation. Journal of Cell Biology, 154(1):187–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/1/187.

McElhinny:2002:MSR

[MKS^{+}02] Abigail S. McElhinny, Kazumi Kakinuma, Hiroyuki Sorimachi, Siegfried Labeit, and Carol C. Gregorio. Musclespecific RING finger-1 interacts with titin to regulate sarcomeric M-line and thick filament structure and may have nuclear functions via its interaction with glucocorticoid modulatory element binding protein-1. Journal of Cell Biology, 157 (1):125–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/1/125.

Mimori-Kiyosue:2000:APC

[MKST00] Yuko Mimori-Kiyosue, Nobuyuki Shiina, and Shoichiro Tsukita. Adenomatous Polyposis Coli (APC) Protein Moves along Microtubules and Concentrates at Their Growing Ends in Epithelial Cells. Journal of Cell Biology, 148(3):505– ??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/3/505.

Mimori-Kiyosue:2001:WAG

[MKT01] Yuko Mimori-Kiyosue and Shoichiro Tsukita. Where is APC going? Journal of Cell Biology, 154(6):1105–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/154/
6/1105.

Moore:2001:MVE

[MKTW01] Jeffrey R. Moore, Elena B. Krementsova, Kathleen M. Trybus, and David M. Warshaw. Myosin V exhibits a high duty cycle and large unitary displacement. Journal of Cell Biology, 155(4):625–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/4/625.

McHugh:2004:I

[MKY^{+}04] Brian McHugh, Sue A. Krause, Bin Yu, Anne-Marie Deans, Sarah Heasman, Paul McLaughlin, and Margarete M. S. Heck. Invadolysin. Journal of Cell Biology, 167(4):673– ??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/4/673.

Montana:2004:CHI

[ML04] Enrico S. Montana and J. Troy Littleton. Characterization of a hypercontraction-induced myopathy in Drosophila caused by mutations in Mhc. Journal of Cell Biology, 164(7): 1045–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/7/1045.

Marelli:2001:LBS

[MLC^{+}01] Marcello Marelli, C. Patrick Lusk, Honey Chan, John D. Aitchison, and Richard W. Wozniak. A Link between the Synthesis of Nucleoporins and the Biogenesis of the Nuclear Envelope. Journal of Cell Biology, 153(4):709–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/153/4/
709.

Matsuura:2004:BNP

[MLKH04] Kumi Matsuura, Paul A. Lefebvre, Ritsu Kamiya, and Masafumi Hirono. Bld10p, a novel protein essential for basal body assembly in Chlamydomonas. Journal of Cell Biology, 165 (5):663–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/5/663.

Machleidt:2000:MDC

[MLLA00] Thomas Machleidt, Wei-Ping Li, Pingsheng Liu, and Richard G. W. Anderson. Multiple Domains in Caveolin-1 Control Its Intracellular Traffic. Journal of Cell Biology, 148(1): 17–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/1/17.

Muraoka:2001:CDK

[MLS^{+}01] Rebecca S. Muraoka, Anne E. G. Lenferink, Jean Simpson, Dana M. Brantley, L. Renee Roebuck, F. Michael Yakes, and Carlos L. Arteaga. Cyclin-dependent Kinase Inhibitor P27^{Kip1} Is Required for Mouse Mammary Gland Morphogenesis and Function. Journal of Cell Biology, 153(5):917–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/5/917.

Markovic:2001:SAR

[MLZ^{+}01] Ingrid Markovic, Eugenia Leikina, Mikhail Zhukovsky, Joshua Zimmerberg, and Leonid V. Chernomordik. Synchronized activation and refolding of influenza hemagglutinin in multimeric fusion machines. Journal of Cell Biology, 155(5): 833–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/5/833.

Mushegian:2001:EPI

[MM01] Arcady Mushegian and Ruslan Medzhitov. Evolutionary perspective on innate immune recognition. Journal of Cell Biology, 155(5):705–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/5/705.

Mellman:2003:CCB

[MM03] Ira Mellman and Tom Misteli. Computational cell biology. Journal of Cell Biology, 161(3):463–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/3/463.

McWhorter:2003:KSM

[MMBB03] Michelle L. McWhorter, Umrao R. Monani, Arthur H. M. Burghes, and Christine E. Beattie. Knockdown of the survival motor neuron (Smn) protein in zebrafish causes defects in motor axon outgrowth and pathfinding. Journal of Cell Biology, 162(5):919–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/5/919.

Motegi:2004:MIR

[MMBM04] Fumio Motegi, Mithilesh Mishra, Mohan K. Balasubramanian, and Issei Mabuchi. Myosin-II reorganization during mitosis is controlled temporally by its dephosphorylation and spatially by Mid1 in fission yeast. Journal of Cell Biology, 165 (5):685–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/5/685.

Morrison:2000:PKP

[MMC00] Deborah K. Morrison, Monica S. Murakami, and Vaughn Cleghon. Protein Kinases and Phosphatases in the Drosophila Genome. Journal of Cell Biology, 150(2):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
2/F57.

MacFarlane:2000:ACC

[MMDC00] Marion MacFarlane, Wendy Merrison, David Dinsdale, and Gerald M. Cohen. Active Caspases and Cleaved Cytokeratins Are Sequestered into Cytoplasmic Inclusions in Trail– Induced Apoptosis. Journal of Cell Biology, 148(6):1239–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/148/6/1239.

Miles:2001:EEG

[MMFS01] Suzanne Miles, Heather McManus, Kimberly E. Forsten, and Brian Storrie. Evidence that the entire Golgi apparatus cycles in interphase HeLa cells. Journal of Cell Biology, 155(4): 543–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/4/543.

Meriin:2001:IAP

[MMG^{+}01] Anatoli B. Meriin, Katsuhide Mabuchi, Vladimir L. Gabai, Julia A. Yaglom, Alex Kazantsev, and Michael Y. Sherman. Intracellular Aggregation of Polypeptides with Expanded Polyglutamine Domain Is Stimulated by Stress–Activated Kinase Mekk1. Journal of Cell Biology, 153(4):851–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
4/851.

Murtagh:2004:OME

[MMG^{+}04] Janice Murtagh, Emmett McArdle, Emer Gilligan, Laura Thornton, Fiona Furlong, and Finian Martin. Organization of mammary epithelial cells into 3D acinar structures requires glucocorticoid and JNK signaling. Journal of Cell Biology, 166(1):133–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/1/133.

Melikyan:2000:ETH

[MMH^{+}00] Grigory B. Melikyan, Ruben M. Markosyan, Hila Hemmati, Mary K. Delmedico, Dennis M. Lambert, and Fredric S. Cohen. Evidence That the Transition of HIV-1 Gp41 into a Six– Helix Bundle, Not the Bundle Configuration, Induces Membrane Fusion. Journal of Cell Biology, 151(2):413–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
151/2/413.

Matheos:2004:PIP

[MMM^{+}04] Dina Matheos, Metodi Metodiev, Eric Muller, David Stone, and Mark D. Rose. Pheromone-induced polarization is dependent on the Fus3p MAPK acting through the formin Bni1p. Journal of Cell Biology, 165(1):99–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/165/1/
99.

Martinez-Menarguez:2001:PGV

[MMPO^{+}01] José A. Martínez-Menárguez, Rytis Prekeris, Viola M. J. Oorschot, Richard Scheller, Jan W. Slot, Hans J. Geuze, and Judith Klumperman. Peri-Golgi vesicles contain retrograde but not anterograde proteins consistent with the cisternal progression model of intra-Golgi transport. Journal of Cell Biology, 155(7):1213–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/7/1213.

Mancini:2000:CLG

[MMR^{+}00] Marie Mancini, Carolyn E. Machamer, Sophie Roy, Donald W. Nicholson, Nancy A. Thornberry, Livia A. Casciola-Rosen, and Antony Rosen. Caspase-2 Is Localized at the Golgi Complex and Cleaves Golgin-160 during Apoptosis. Journal of Cell Biology, 149(3):603–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/3/603.

Mozdy:2000:DGM

[MMS00] A. D. Mozdy, J. M. McCaffery, and J. M. Shaw. Dnm1p GTPase-mediated mitochondrial fission is a multi-step process requiring the novel integral membrane component Fis1p. Journal of Cell Biology, 151(2):367–??, October 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/2/
367.

Milner:2000:DCL

[MMWC00] Derek J. Milner, Manolis Mavroidis, Noah Weisleder, and Yassemi Capetanaki. Desmin Cytoskeleton Linked to Muscle Mitochondrial Distribution and Respiratory Function. Journal of Cell Biology, 150(6):1283–??, September 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/6/
1283.

Meeusen:2003:ETM

[MN03] Shelly Meeusen and Jodi Nunnari. Evidence for a two membrane–spanning autonomous mitochondrial DNA replisome. Journal of Cell Biology, 163(3):503–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
3/503.

Mitchell:2004:BPD

[MN04] David R. Mitchell and Masako Nakatsugawa. Bend propagation drives central pair rotation in Chlamydomonas reinhardtii flagella. Journal of Cell Biology, 166(5):709–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
166/5/709.

Miao:2003:EKA

[MNC^{+}03] Hui Miao, Christian H. Nickel, Lloyd G. Cantley, Leslie A. Bruggeman, Laura N. Bennardo, and Bingcheng Wang. EphA kinase activation regulates HGF-induced epithelial branching morphogenesis. Journal of Cell Biology, 162(7): 1281–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/7/1281.

Muniz:2000:ECR

[MNHR00] Manuel Muñiz, Claude Nuoffer, Hans-Peter Hauri, and Howard Riezman. The Emp24 Complex Recruits a Specific Cargo Molecule into Endoplasmic Reticulum–Derived Vesicles. Journal of Cell Biology, 148(5):925–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/5/
925.

Mizoguchi:2002:N

[MNK^{+}02] Akira Mizoguchi, Hiroyuki Nakanishi, Kazushi Kimura, Kaho Matsubara, Kumi Ozaki-Kuroda, Tatsuo Katata, Tomoyuki Honda, Yoshimoto Kiyohara, Kyun Heo, Mikito Higashi, Tomonari Tsutsumi, Satomi Sonoda, Chizuka Ide, and Yoshimi Takai. Nectin. Journal of Cell Biology, 156(3): 555–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/3/555.

Miyazaki:2003:RCC

[MNT^{+}03] Tsuyoshi Miyazaki, Lynn Neff, Sakae Tanaka, William C. Horne, and Roland Baron. Regulation of cytochrome c oxidase activity by c–Src in osteoclasts. Journal of Cell Biology, 160(5):709–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/5/709.

Matsubara:2001:ECC

[MO01] Shyuichiro Matsubara and Masayuki Ozawa. Expression of α-catenin in α-catenin–deficient cells increases resistance to sphingosine-induced apoptosis. Journal of Cell Biology, 154 (3):573–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/3/573.

Martinez:2003:DRN

[MOMK03] Maria Cruz Martinez, Tomoyo Ochiishi, Michael Majewski, and Kenneth S. Kosik. Dual regulation of neuronal morphogenesis by a δ-catenin–cortactin complex and Rho. Journal of Cell Biology, 162(1):99–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/1/99.

Moolenaar:2002:LL

[Moo02] Wouter H. Moolenaar. Lysophospholipids in the limelight. Journal of Cell Biology, 158(2):197–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/2/197.

Morris:2000:NM

[Mor00] N. Ronald Morris. Nuclear Migration. Journal of Cell Biology, 148(6):1097–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/6/1097.

Miralles:2000:ETR

[MÖS^{+}00] Francesc Miralles, Lars-Göran Öfverstedt, Nafiseh Sabri, Youssef Aissouni, Ulf Hellman, Ulf Skoglund, and Neus Visa. Electron Tomography Reveals Posttranscriptional Binding of Pre–Mrnps to Specific Fibers in the Nucleoplasm. Journal of Cell Biology, 148(2):271–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/2/271.

Matsusaka:2004:CAP

[MP04] Takahiro Matsusaka and Jonathon Pines. Chfr acts with the p38 stress kinases to block entry to mitosis in mammalian cells. Journal of Cell Biology, 166(4):507–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
4/507.

Mostafavi-Pour:2003:ISS

[MPAP^{+}03] Zohreh Mostafavi-Pour, Janet A. Askari, Scott J. Parkinson, Peter J. Parker, Tony T. C. Ng, and Martin J. Humphries. Integrin-specific signaling pathways controlling focal adhesion formation and cell migration. Journal of Cell Biology, 161(1):155–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/1/155.

Mahy:2002:GDT

[MPB02] Nicola L. Mahy, Paul E. Perry, and Wendy A. Bickmore. Gene density and transcription influence the localization of chromatin outside of chromosome territories detectable by FISH. Journal of Cell Biology, 159(5):753–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
5/753.

Miyamoto:2004:KED

[MPB^{+}04] David T. Miyamoto, Zachary E. Perlman, Kendra S. Burbank, Aaron C. Groen, and Timothy J. Mitchison. The kinesin Eg5 drives poleward microtubule flux in Xenopus laevis egg extract spindles. Journal of Cell Biology, 167(5): 813–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/5/813.

Morsomme:2003:VSR

[MPBR03] Pierre Morsomme, Cristina Prescianotto-Baschong, and Howard Riezman. The ER v–SNAREs are required for GPIanchored protein sorting from other secretory proteins upon exit from the ER. Journal of Cell Biology, 162(3):403–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
162/3/403.

Mahy:2002:SOA

[MPG^{+}02] Nicola L. Mahy, Paul E. Perry, Susan Gilchrist, Richard A. Baldock, and Wendy A. Bickmore. Spatial organization of active and inactive genes and noncoding DNA within chromosome territories. Journal of Cell Biology, 157(4):579–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/157/
4/579.

Monani:2003:TCA

[MPG^{+}03] Umrao R. Monani, Matthew T. Pastore, Tatiana O. Gavrilina, Sibylle Jablonka, Thanh T. Le, Catia Andreassi, Jennifer M. DiCocco, Christian Lorson, Elliot J. Androphy, Michael Sendtner, Michael Podell, and Arthur H. M. Burghes. A transgene carrying an A2G missense mutation in the SMN gene modulates phenotypic severity in mice with severe (type I) spinal muscular atrophy. Journal of Cell Biology, 160(1):41–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/1/41.

Matsuda:2001:RTF

[MPR^{+}01] Miho Matsuda, Hugh F. Paterson, Rosie Rodriguez, Amanda C. Fensome, Moira V. Ellis, Karl Swann, and Matilda Katan. Real Time Fluorescence Imaging of Plcγ Translocation and Its Interaction with the Epidermal Growth Factor Receptor. Journal of Cell Biology, 153(3):599–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/3/599.

Mudry:2003:ITT

[MPR^{+}03] Ryan E. Mudry, Cynthia N. Perry, Meredith Richards, Velia M. Fowler, and Carol C. Gregorio. The interaction of tropomodulin with tropomyosin stabilizes thin filaments in cardiac myocytes. Journal of Cell Biology, 162(6):1057– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/6/1057.

Mah:2000:IUN

[MPSM00a] Alex L. Mah, George Perry, Mark A. Smith, and Mervyn J. Monteiro. Identification of Ubiquilin, a Novel Presenilin Interactor That Increases Presenilin Protein Accumulation. Journal of Cell Biology, 151(4):847–??, November 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/4/
847.

Martinez-Perez:2000:PIC

[MPSM00b] Enrique Martinez-Perez, Peter J. Shaw, and Graham Moore. Polyploidy Induces Centromere Association. Journal of Cell Biology, 148(2):233–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/148/2/233.

Mitchell:2001:DCS

[MPV^{+}01] Kathryn J. Mitchell, Paolo Pinton, Aniko Varadi, Carlo Tacchetti, Edward K. Ainscow, Tullio Pozzan, Rosario Rizzuto, and Guy A. Rutter. Dense core secretory vesicles revealed as a dynamic Ca^{2+} store in neuroendocrine cells with a vesicleassociated membrane protein aequorin chimaera. Journal of Cell Biology, 155(1):41–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/1/41.

Mills:2003:E

[MPV^{+}03] Ian G. Mills, Gerrit J. K. Praefcke, Yvonne Vallis, Brian J. Peter, Lene E. Olesen, Jennifer L. Gallop, P. Jonathan G. Butler, Philip R. Evans, and Harvey T. McMahon. EpsinR. Journal of Cell Biology, 160(2):213–??, January 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/2/
213.

Marshall:2001:ITB

[MR01a] Wallace F. Marshall and Joel L. Rosenbaum. Intraflagellar transport balances continuous turnover of outer doublet microtubules. Journal of Cell Biology, 155(3):405–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
3/405.

Moore:2001:HCC

[MR01b] Landon L. Moore and Mark B. Roth. Hcp-4, a Cenp–C-like Protein in Caenorhabditis elegans, Is Required for Resolution of Sister Centromeres. Journal of Cell Biology, 153(6): 1199–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/6/1199.

Mellman:2002:RFI

[MR02] Ira Mellman and Anne Ridley. Regulation and Functional Insights in Cellular Polarity Madrid, Spain June 3–5, 2002. Journal of Cell Biology, 158(1):12–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/1/12.

May:2002:PBC

[MRC^{+}02] Karen M. May, Nicola Reynolds, C. Fiona Cullen, Mitsuhiro Yanagida, and Hiroyuki Ohkura. Polo boxes and Cut23 (Apc8) mediate an interaction between polo kinase and the anaphase-promoting complex for fission yeast mitosis. Journal of Cell Biology, 156(1):23–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/1/23.

Marx:2001:PDR

[MRH^{+}01] Steven O. Marx, Steven Reiken, Yuji Hisamatsu, Marta Gaburjakova, Jana Gaburjakova, Yi-Ming Yang, Nora Rosemblit, and Andrew R. Marks. Phosphorylationdependent Regulation of Ryanodine Receptors. Journal of Cell Biology, 153(4):699–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/4/699.

Maiato:2004:KDF

[MRK04] Helder Maiato, Conly L. Rieder, and Alexey Khodjakov. Kinetochore-driven formation of kinetochore fibers contributes to spindle assembly during animal mitosis. Journal of Cell Biology, 167(5):831–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/5/831.

McGarvey:2000:AML

[MRM^{+}00] Tim McGarvey, Emanuel Rosonina, Susan McCracken, Qiyu Li, Ramy Arnaout, Edwin Mientjes, Jeffrey A. Nickerson, Don Awrey, Jack Greenblatt, Gerard Grosveld, and Benjamin J. Blencowe. The Acute Myeloid Leukemia–Associated Protein, Dek, Forms a Splicing–Dependent Interaction with Exon–Product Complexes. Journal of Cell Biology, 150(2): 309–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/2/309.

Mann:2004:CIS

[MRM^{+}04] Kris M. Mann, Jenna Lynn Ray, Edward S. Moon, Kristin M. Sass, and Mark R. Benson. Calcineurin initiates smooth muscle differentiation in neural crest stem cells. Journal of Cell Biology, 165(4):483–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/4/483.

Mullershausen:2001:RNO

[MRT^{+}01] Florian Mullershausen, Michael Russwurm, W. Joseph Thompson, Li Liu, Doris Koesling, and Andreas Friebe. Rapid nitric oxide–induced desensitization of the cGMP response is caused by increased activity of phosphodiesterase type 5 paralleled by phosphorylation of the enzyme. Journal of Cell Biology, 155(2):271–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/2/271.

Mount:2000:PMR

[MS00] Stephen M. Mount and Helen K. Salz. Pre-messenger RNA Processing Factors in the Drosophila Genome. Journal of Cell Biology, 150(2):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/2/F37.

Maul:2003:ERA

[MSA^{+}03] Raymond S. Maul, Yuhong Song, Kurt J. Amann, Sachi C. Gerbin, Thomas D. Pollard, and David D. Chang. EPLIN regulates actin dynamics by cross-linking and stabilizing filaments. Journal of Cell Biology, 160(3):399–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
3/399.

Maddox:2003:DOM

[MSC^{+}03] Paul Maddox, Aaron Straight, Peg Coughlin, Timothy J. Mitchison, and Edward D. Salmon. Direct observation of microtubule dynamics at kinetochores in Xenopus extract spindles. Journal of Cell Biology, 162(3):377–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
3/377.

Mouneimne:2004:PCC

[MSD^{+}04] Ghassan Mouneimne, Lilian Soon, Vera DesMarais, Mazen Sidani, Xiaoyan Song, Shu-Chin Yip, Mousumi Ghosh, Robert Eddy, Jonathan M. Backer, and John Condeelis. Phospholipase C and cofilin are required for carcinoma cell directionality in response to EGF stimulation. Journal of Cell Biology, 166(5):697–??, August 2004. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/5/697.

Muller:2000:AT

[MSF^{+}00] Oliver Müller, Tanja Sattler, Matthias Flötenmeyer, Heinz Schwarz, Helmut Plattner, and Andreas Mayer. Autophagic Tubes. Journal of Cell Biology, 151(3):519–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/3/
519.

Mayer:2000:SSP

[MSFH00] Wolfgang Mayer, Avril Smith, Reinald Fundele, and Thomas Haaf. Spatial Separation of Parental Genomes in Preimplantation Mouse Embryos. Journal of Cell Biology, 148(4): 629–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/4/629.

Mahal:2002:CIS

[MSGS02] Lara K. Mahal, Sonia M. Sequeira, Jodi M. Gureasko, and Thomas H. Söllner. Calcium-independent stimulation of membrane fusion and SNAREpin formation by synaptotagmin I. Journal of Cell Biology, 158(2):273–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/2/
273.

Mukherjee:2000:LSR

[MSH^{+}00] Konark Mukherjee, Shadab A. Siddiqi, Shehla Hashim, Manoj Raje, Sandip K. Basu, and Amitabha Mukhopadhyay. Live Salmonella Recruits N –Ethylmaleimide–Sensitive Fusion Protein on Phagosomal Membrane and Promotes Fusion with Early Endosome. Journal of Cell Biology, 148(4): 741–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/4/741.

Moir:2000:DNL

[MSHG00] Robert D. Moir, Timothy P. Spann, Harald Herrmann, and Robert D. Goldman. Disruption of Nuclear Lamin Organization Blocks the Elongation Phase of DNA Replication. Journal of Cell Biology, 149(6):1179–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/6/1179.

McKean:2003:FIE

[MSI^{+}03] David M. McKean, Lila Sisbarro, Dusko Ilic, Nihal Kaplan-Alburquerque, Raphael Nemenoff, Mary Weiser-Evans, Michael J. Kern, and Peter Lloyd Jones. FAK induces expression of Prx1 to promote tenascin–C-dependent fibroblast migration. Journal of Cell Biology, 161(2):393–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
2/393.

Martin:2002:LOP

[MSJ^{+}02] Bernd Martin, Richard Schneider, Stefanie Janetzky, Zoe Waibler, Petra Pandur, Michael Kühl, Jürgen Behrens, Klaus von der Mark, Anna Starzinski-Powitz, and Viktor Wixler. The LIM-only protein FHL2 interacts with β-catenin and promotes differentiation of mouse myoblasts. Journal of Cell Biology, 159(1):113–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/1/113.

Marelli:2004:QMS

[MSJ^{+}04] Marcello Marelli, Jennifer J. Smith, Sunhee Jung, Eugene Yi, Alexey I. Nesvizhskii, Rowan H. Christmas, Ramsey A. Saleem, Yuen Yi C. Tam, Andrei Fagarasanu, David R. Goodlett, Ruedi Aebersold, Richard A. Rachubinski, and John D. Aitchison. Quantitative mass spectrometry reveals a role for the GTPase Rho1p in actin organization on the peroxisome membrane. Journal of Cell Biology, 167(6):1099– ??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/6/1099.

Maiato:2002:MOR

[MSL^{+}02] Helder Maiato, Paula Sampaio, Catarina L. Lemos, John Findlay, Mar Carmena, William C. Earnshaw, and Claudio E. Sunkel. MAST/ orbit has a role in microtubule– kinetochore attachment and is essential for chromosome alignment and maintenance of spindle bipolarity. Journal of Cell Biology, 157(5):749–??, May 2002. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/5/749.

Michaelson:2001:DLR

[MSM^{+}01] David Michaelson, Joseph Silletti, Gretchen Murphy, Peter D’Eustachio, Mark Rush, and Mark R. Philips. Differential Localization of Rho Gtpases in Live Cells. Journal of Cell Biology, 152(1):111–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/1/111.

Marg:2004:NSN

[MSM^{+}04a] Andreas Marg, Ying Shan, Thomas Meyer, Torsten Meissner, Martin Brandenburg, and Uwe Vinkemeier. Nucleocytoplasmic shuttling by nucleoporins Nup153 and Nup214 and CRM1-dependent nuclear export control the subcellular distribution of latent Stat1. Journal of Cell Biology, 165 (6):823–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/6/823.

Mazumdar:2004:HCK

[MSM04b] Manjari Mazumdar, Suma Sundareshan, and Tom Misteli. Human chromokinesin KIF4A functions in chromosome condensation and segregation. Journal of Cell Biology, 166(5): 613–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/5/613.

Murshed:2004:EMM

[MSMK04] Monzur Murshed, Thorsten Schinke, Marc D. McKee, and Gerard Karsenty. Extracellular matrix mineralization is regulated locally; different roles of two gla-containing proteins. Journal of Cell Biology, 165(5):625–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/5/625.

Muller:2002:SSD

[MSN^{+}02] Joyce M. M. Müller, James Shorter, Richard Newman, Katrin Deinhardt, Yuval Sagiv, Zvulun Elazar, Graham Warren, and David T. Shima. Sequential SNARE disassembly and GATE-16–GOS-28 complex assembly mediated by distinct NSF activities drives Golgi membrane fusion. Journal of Cell Biology, 157(7):1161–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/7/1161.

Matsumura:2000:NDP

[MSO^{+}00] Hirotaka Matsumura, Yusuke Shimizu, Yoshiyuki Ohsawa, Atsuo Kawahara, Yasuo Uchiyama, and Shigekazu Nagata. Necrotic Death Pathway in FAS Receptor Signaling. Journal of Cell Biology, 151(6):1247–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/6/1247.

Mannick:2001:NMC

[MSP^{+}01] Joan B. Mannick, Christopher Schonhoff, Natalia Papeta, Pedram Ghafourifar, Marten Szibor, Kezhong Fang, and Benjamin Gaston. S-nitrosylation of mitochondrial caspases. Journal of Cell Biology, 154(6):1111–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/154/6/
1111.

Mikhailov:2004:TIH

[MSR04] Alexei Mikhailov, Mio Shinohara, and Conly L. Rieder. Topoisomerase II and histone deacetylase inhibitors delay the G2/M transition by triggering the p38 MAPK checkpoint pathway. Journal of Cell Biology, 166(4):517–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
4/517.

Miyoshi:2001:STA

[MSS^{+}01] Keiko Miyoshi, Jonathan M. Shillingford, Gilbert H. Smith, Sandra L. Grimm, Kay-Uwe Wagner, Takami Oka, Jeffrey M. Rosen, Gertraud W. Robinson, and Lothar Hennighausen. Signal transducer and activator of transcription (Stat) 5 controls the proliferation and differentiation of mammary alveolar epithelium. Journal of Cell Biology, 155(4):531– ??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/4/531.

Mayor:2000:CPC

[MST^{+}00] Thibault Mayor, York-Dieter Stierhof, Kayoko Tanaka, Andrew M. Fry, and Erich A. Nigg. The Centrosomal Protein C–Nap1 Is Required for Cell Cycle–Regulated Centrosome Cohesion. Journal of Cell Biology, 151(4):837–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
4/837.

Marzia:2000:DCS

[MSV^{+}00] Marilena Marzia, Natalie A. Sims, Susanne Voit, Silvia Migliaccio, Anna Taranta, Silvia Bernardini, Tullio Faraggiana, Toshiyuki Yoneda, Gregory R. Mundy, Brendan F. Boyce, Roland Baron, and Anna Teti. Decreased C–Src Expression Enhances Osteoblast Differentiation and Bone Formation. Journal of Cell Biology, 151(2):311–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/2/
311.

Miyamoto:2004:CSI

[MSY^{+}04] Yoichi Miyamoto, Takuya Saiwaki, Junichi Yamashita, Yoshinari Yasuda, Ippei Kotera, Satoshi Shibata, Masaki Shigeta, Yasushi Hiraoka, Tokuko Haraguchi, and Yoshihiro Yoneda. Cellular stresses induce the nuclear accumulation of importin α and cause a conventional nuclear import block. Journal of Cell Biology, 165(5):617–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/5/617.

Matsumoto:2002:PMK

[MTB^{+}02] Taro Matsumoto, Ingela Turesson, Majlis Book, Pär Gerwins, and Lena Claesson-Welsh. p38 MAP kinase negatively regulates endothelial cell survival, proliferation, and differentiation in FGF-2–stimulated angiogenesis. Journal of Cell Biology, 156(1):149–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/1/149.

Mallard:2002:ERE

[MTG^{+}02] Frédéric Mallard, Bor Luen Tang, Thierry Galli, Danièle Tenza, Agnès Saint-Pol, Xu Yue, Claude Antony, Wanjin Hong, Bruno Goud, and Ludger Johannes. Early/ recycling endosomes–to–TGN transport involves two SNARE complexes and a Rab6 isoform. Journal of Cell Biology, 156(4): 653–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/4/653.

Mogensen:2002:APC

[MTM^{+}02] Mette M. Mogensen, John B. Tucker, John B. Mackie, Alan R. Prescott, and Inke S. Näthke. The adenomatous polyposis coli protein unambiguously localizes to microtubule plus ends and is involved in establishing parallel arrays of microtubule bundles in highly polarized epithelial cells. Journal of Cell Biology, 157(6):1041–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/6/1041.

McCroskery:2003:MNR

[MTM^{+}03] Seumas McCroskery, Mark Thomas, Linda Maxwell, Mridula Sharma, and Ravi Kambadur. Myostatin negatively regulates satellite cell activation and self-renewal. Journal of Cell Biology, 162(6):1135–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/6/1135.

Morgan:2002:NCS

[MTPT02] Michael Morgan, Jacqueline Thorburn, Pier Paolo Pandolfi, and Andrew Thorburn. Nuclear and cytoplasmic shuttling of TRADD induces apoptosis via different mechanisms. Journal of Cell Biology, 157(6):975–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/6/975.

Mandelkow:2004:MPK

[MTT^{+}04] Eva-Maria Mandelkow, Edda Thies, Bernhard Trinczek, Jacek Biernat, and Eckard Mandelkow. MARK/ PAR1 kinase is a regulator of microtubule-dependent transport in axons. Journal of Cell Biology, 167(1):99–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/167/1/

Minshall:2000:ECS

[MTV^{+}00] Richard D. Minshall, Chinnaswamy Tiruppathi, Stephen M. Vogel, Walter D. Niles, Annette Gilchrist, Heidi E. Hamm, and Asrar B. Malik. Endothelial Cell–Surface Gp60 Activates Vesicle Formation and Trafficking via G_{i} -Coupled Src Kinase Signaling Pathway. Journal of Cell Biology, 150(5):1057– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/5/1057.

Ma:2002:RRA

[MTW^{+}02] Zhong Ma, Keena S. Thomas, Donna J. Webb, Radim Moravec, Ana Maria Salicioni, Wendy M. Mars, and Steven L. Gonias. Regulation of Rac1 activation by the low density lipoprotein receptor–related protein. Journal of Cell Biology, 159(6):1061–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/6/1061.

Meriane:2004:PDF

[MTW^{+}04] Mayya Meriane, Joseph Tcherkezian, Christine A. Webber, Eric I. Danek, Ibtissem Triki, Sarah McFarlane, Evelyne Bloch-Gallego, and Nathalie Lamarche-Vane. Phosphorylation of DCC by Fyn mediates Netrin-1 signaling in growth cone guidance. Journal of Cell Biology, 167(4):687– ??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/4/687.

McDonald:2002:VIB

[MVL^{+}02] David McDonald, Marie A. Vodicka, Ginger Lucero, Tatyana M. Svitkina, Gary G. Borisy, Michael Emerman, and Thomas J. Hope. Visualization of the intracellular behavior of HIV in living cells. Journal of Cell Biology, 159(3): 441–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/3/441.

Merz:2004:TSI

[MW04] Alexey J. Merz and William T. Wickner. Trans-SNARE interactions elicit Ca^{2+} efflux from the yeast vacuole lumen.

Journal of Cell Biology, 164(2):195–??, January 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/2/
195.

Majdan:2001:TMD

[MWAM01] Marta Majdan, Gregory S. Walsh, Raquel Aloyz, and Freda D. Miller. TrkA mediates developmental sympathetic neuron survival in vivo by silencing an ongoing p75NTRmediated death signal. Journal of Cell Biology, 155(7):1275– ??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/7/1275.

Murray:2002:PKC

[MWC^{+}02] Nicole R. Murray, Capella Weems, Lu Chen, Jessica Leon, Wangsheng Yu, Laurie A. Davidson, Lee Jamieson, Robert S. Chapkin, E. Aubrey Thompson, and Alan P. Fields. Protein kinase C βII and TGFβRII in ω-3 fatty acid–mediated inhibition of colon carcinogenesis. Journal of Cell Biology, 157 (6):915–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/6/915.

Metkar:2003:GBA

[MWE^{+}03] Sunil S. Metkar, Baikun Wang, Michelle L. Ebbs, Jin H. Kim, Yong J. Lee, Srikumar M. Raja, and Christopher J. Froelich. Granzyme B activates procaspase-3 which signals a mitochondrial amplification loop for maximal apoptosis. Journal of Cell Biology, 160(6):875–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/6/875.

Marangi:2002:LRP

[MWF02] P. Angelo Marangi, Simon T. Wieland, and Christian Fuhrer. Laminin-1 redistributes postsynaptic proteins and requires rapsyn, tyrosine phosphorylation, and Src and Fyn to stably cluster acetylcholine receptors. Journal of Cell Biology, 157 (5):883–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/5/883.

Muller:2001:LSC

[MWHM01] Waltraud G. Müller, Dawn Walker, Gordon L. Hager, and James G. McNally. Large-scale chromatin decondensation and recondensation regulated by transcription from a natural promoter. Journal of Cell Biology, 154(1):33–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/154/
1/33.

Medina:2002:CIM

[MWK^{+}02] Emmanuelle Médina, Janice Williams, Elizabeth Klipfell, Daniela Zarnescu, Graham Thomas, and André Le Bivic. Crumbs interacts with moesin and β_{Heavy}-spectrin in the apical membrane skeleton of Drosophila. Journal of Cell Biology, 158(5):941–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/5/941.

Maheshwari:2001:AEG

[MWL01] Gargi Maheshwari, H. Steven Wiley, and Douglas A. Lauffenburger. Autocrine epidermal growth factor signaling stimulates directionally persistent mammary epithelial cell migration. Journal of Cell Biology, 155(7):1123–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
7/1123.

Melia:2002:RMF

[MWM^{+}02] Thomas J. Melia, Thomas Weber, James A. McNew, Lillian E. Fisher, Robert J. Johnston, Frank Parlati, Lara K. Mahal, Thomas H. Söllner, and James E. Rothman. Regulation of membrane fusion by the membrane-proximal coil of the t–SNARE during zippering of SNAREpins. Journal of Cell Biology, 158(5):929–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/5/929.

Miller:2004:ARG

[MWMK04] Ann L. Miller, Yinxiang Wang, Mark S. Mooseker, and Anthony J. Koleske. The Abl-related gene (Arg) requires its F-actin–microtubule cross-linking activity to regulate lamellipodial dynamics during fibroblast adhesion. Journal of Cell Biology, 165(3):407–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/3/407.

Matsuura:2004:CFS

[MWN^{+}04a] Katsuhisa Matsuura, Hiroshi Wada, Toshio Nagai, Yoshihiro Iijima, Tohru Minamino, Masanori Sano, Hiroshi Akazawa, Jeffery D. Molkentin, Hiroshi Kasanuki, and Issei Komuro. Cardiomyocytes fuse with surrounding noncardiomyocytes and reenter the cell cycle. Journal of Cell Biology, 167(2): 351–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/2/351.

Matsuura:2004:C

[MWN^{+}04b] Katsuhisa Matsuura, Hiroshi Wada, Toshio Nagai, Yoshihiro Iijima, Tohru Minamino, Masanori Sano, Hiroshi Akazawa, Jeffery D. Molkentin, Hiroshi Kasanuki, and Issei Komuro. Correction. Journal of Cell Biology, 167(5):985–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
167/5/985.

McNew:2000:CE

[MWP^{+}00] James A. McNew, Thomas Weber, Francesco Parlati, Robert J. Johnston, Thomas J. Melia, Thomas H. Söllner, and James E. Rothman. Close Is Not Enough. Journal of Cell Biology, 150(1):105–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/1/105.

Martinez-Wittinghan:2003:DCR

[MWSL^{+}03] Francisco J. Martinez-Wittinghan, Caterina Sellitto, Leping Li, Xiaohua Gong, Peter R. Brink, Richard T. Mathias, and Thomas W. White. Dominant cataracts result from incongruous mixing of wild-type lens connexins. Journal of Cell Biology, 161(5):969–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/5/969.

Mehta:2002:MPP

[MYC^{+}02] Shwetal Mehta, Xian Mei Yang, Clarence S. Chan, Melanie J. Dobson, Makkuni Jayaram, and Soundarapandian Velmurugan. The 2 micron plasmid purloins the yeast cohesin complex. Journal of Cell Biology, 158(4):625–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/4/
625.

Mizushima:2001:DAF

[MYH^{+}01] Noboru Mizushima, Akitsugu Yamamoto, Masahiko Hatano, Yoshinori Kobayashi, Yukiko Kabeya, Kuninori Suzuki, Takeshi Tokuhisa, Yoshinori Ohsumi, and Tamotsu Yoshimori. Dissection of Autophagosome Formation Using Apg5-Deficient Mouse Embryonic Stem Cells. Journal of Cell Biology, 152(4):657–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/4/657.

Moir:2000:NLB

[MYKG00] Robert D. Moir, Miri Yoon, Satya Khuon, and Robert D. Goldman. Nuclear Lamins a and B1. Journal of Cell Biology, 151(6):1155–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/6/1155.

Miyagi:2004:SNA

[MYO^{+}04] Chiemi Miyagi, Susumu Yamashita, Yusuke Ohba, Hisayoshi Yoshizaki, Michiyuki Matsuda, and Toshio Hirano. STAT3 noncell-autonomously controls planar cell polarity during zebrafish convergence and extension. Journal of Cell Biology, 166(7):975–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/7/975.

McCloskey:2000:PFR

[MZ00] Michael A. McCloskey and Lei Zhang. Potentiation of Fcϵ receptor I-activated Ca^{2+} current (I_{CRAC}) by cholera toxin: Possible mediation by ADP ribosylation factor. Journal of Cell Biology, 148(1):137–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/1/137.

Meriin:2002:HTY

[MZH^{+}02] Anatoli B. Meriin, Xiaoqian Zhang, Xiangwei He, Gary P. Newnam, Yury O. Chernoff, and Michael Y. Sherman. Huntingtin toxicity in yeast model depends on polyglutamine aggregation mediated by a prion-like protein Rnq1. Journal of Cell Biology, 157(6):997–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/6/997.

Mitra:2004:NPP

[MZR^{+}04] Prasenjit Mitra, Yingjie Zhang, Lucia E. Rameh, Maria P. Ivshina, Dannel McCollum, John J. Nunnari, Gregory M. Hendricks, Monica L. Kerr, Seth J. Field, Lewis C. Cantley, and Alonzo H. Ross. A novel phosphatidylinositol(3,4,5)P_{3} pathway in fission yeast. Journal of Cell Biology, 166(2): 205–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/2/205.

Manes:2003:IER

[MZT^{+}03] Thomas Manes, Duo-Qi Zheng, Simona Tognin, Amy S. Woodard, Pier Carlo Marchisio, and Lucia R. Languino. α_{v}β_{3} integrin expression up-regulates cdc2, which modulates cell migration. Journal of Cell Biology, 161(4):817–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
4/817.

Nern:2000:NSC

[NA00] Aljoscha Nern and Robert A. Arkowitz. Nucleocytoplasmic Shuttling of the Cdc42p Exchange Factor Cdc24p. Journal of Cell Biology, 148(6):1115–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/148/6/1115.

NDiaye:2003:URP

[NB03a] Elsa-Noah N’Diaye and Eric J. Brown. The ubiquitin-related protein PLIC-1 regulates heterotrimeric G protein function through association with Gβγ. Journal of Cell Biology, 163(5):1157–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/5/1157.

Nguyen:2003:CMS

[NB03b] Chan Nguyen and James A. Bibb. Cdk5 and the mystery of synaptic vesicle endocytosis. Journal of Cell Biology, 163(4):

697–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/4/697.

Nakamura:2000:CER

[NBWB^{+}00] Kimitoshi Nakamura, Ella Bossy-Wetzel, Kimberly Burns, Marc P. Fadel, Mira Lozyk, Ing Swie Goping, Michal Opas, R. Chris Bleackley, Douglas R. Green, and Marek Michalak. Changes in Endoplasmic Reticulum Luminal Environment Affect Cell Sensitivity to Apoptosis. Journal of Cell Biology, 150(4):731–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/4/731.

Niedergang:2003:ARF

[NCGD^{+}03] Florence Niedergang, Emma Colucci-Guyon, Thierry Dubois, Graça Raposo, and Philippe Chavrier. ADP ribosylation factor 6 is activated and controls membrane delivery during phagocytosis in macrophages. Journal of Cell Biology, 161 (6):1143–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/6/1143.

Nunbhakdi-Craig:2002:PPA

[NCMO^{+}02] Viyada Nunbhakdi-Craig, Thomas Machleidt, Egon Ogris, Dennis Bellotto, Charles L. White, and Estelle Sontag. Protein phosphatase 2A associates with and regulates atypical PKC and the epithelial tight junction complex. Journal of Cell Biology, 158(5):967–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/5/967.

Nielsen:2000:RNR

[NCUJ^{+}00] Erik Nielsen, Savvas Christoforidis, Sandrine Uttenweiler-Joseph, Marta Miaczynska, Frederique Dewitte, Matthias Wilm, Bernard Hoflack, and Marino Zerial. Rabenosyn-5, a Novel Rab5 Effector, Is Complexed with Hvps45 and Recruited to Endosomes through a Fyve Finger Domain. Journal of Cell Biology, 151(3):601–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/3/601.

Nicole:2003:ISC

[NDM^{+}03] Sophie Nicole, Benedicte Desforges, Gaelle Millet, Jeanne Lesbordes, Carmen Cifuentes-Diaz, Dora Vertes, My Linh Cao, Fabienne De Backer, Laetitia Languille, Natacha Roblot, Vandana Joshi, Jean-Marie Gillis, and Judith Melki. Intact satellite cells lead to remarkable protection against Smn gene defect in differentiated skeletal muscle. Journal of Cell Biology, 161(3):571–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/3/571.

Niethammer:2002:CNL

[NDS^{+}02] Philipp Niethammer, Markus Delling, Vladimir Sytnyk, Alexander Dityatev, Kiyoko Fukami, and Melitta Schachner. Cosignaling of NCAM via lipid rafts and the FGF receptor is required for neuritogenesis. Journal of Cell Biology, 157 (3):521–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/3/521.

Nedelec:2002:CSR

[Néd02] François Nédélec. Computer simulations reveal motor properties generating stable antiparallel microtubule interactions. Journal of Cell Biology, 158(6):1005–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/6/
1005.

Nelson:2000:WHI

[Nel00] W. James Nelson. W(h)ither the Golgi during Mitosis? Journal of Cell Biology, 149(2):243–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/2/243.

Nogales:2001:MM

[NG01] Eva Nogales and Nikolaus Grigorieff. Molecular Machines. Journal of Cell Biology, 152(1):0–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/1/F1.

Niessen:2002:CMC

[NG02] Carien M. Niessen and Barry M. Gumbiner. Cadherinmediated cell sorting not determined by binding or adhesion specificity. Journal of Cell Biology, 156(2):389–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/156/
2/389.

Nesbitt:2003:ICC

[NGK^{+}03] Warwick S. Nesbitt, Simon Giuliano, Suhasini Kulkarni, Sacha M. Dopheide, Ian S. Harper, and Shaun P. Jackson. Intercellular calcium communication regulates platelet aggregation and thrombus growth. Journal of Cell Biology, 160(7): 1151–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/7/1151.

Niv:2002:AKR

[NGKH02] Hagit Niv, Orit Gutman, Yoel Kloog, and Yoav I. Henis. Activated K–Ras and H–Ras display different interactions with saturable nonraft sites at the surface of live cells. Journal of Cell Biology, 157(5):865–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/5/865.

Newman:2001:CEM

[NGS^{+}01] Bill Newman, Laure I. Gigout, Laure Sudre, Michael E. Grant, and Gillian A. Wallis. Coordinated expression of matrix Gla protein is required during endochondral ossification for chondrocyte survival. Journal of Cell Biology, 154(3): 659–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/3/659.

Nakata:2003:MPD

[NH03a] Takao Nakata and Nobutaka Hirokawa. Microtubules provide directional cues for polarized axonal transport through interaction with kinesin motor head. Journal of Cell Biology, 162(6):1045–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/6/1045.

Nawshad:2003:TSA

[NH03b] Ali Nawshad and Elizabeth D. Hay. TGFβ_{3} signaling activates transcription of the LEF1 gene to induce epithelial mesenchymal transformation during mouse palate development. Journal of Cell Biology, 163(6):1291–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
6/1291.

Nothwehr:2000:SYM

[NHB00] Steven F. Nothwehr, Seon-Ah Ha, and Paul Bruinsma. Sorting of yeast membrane proteins into an endosome-to-Golgi pathway involves direct interaction of their cytosolic domains with Vps35p. Journal of Cell Biology, 151(2):297–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
151/2/297.

Nitta:2003:SSL

[NHG^{+}03] Takehiro Nitta, Masaki Hata, Shimpei Gotoh, Yoshiteru Seo, Hiroyuki Sasaki, Nobuo Hashimoto, Mikio Furuse, and Shoichiro Tsukita. Size-selective loosening of the blood-brain barrier in claudin-5–deficient mice. Journal of Cell Biology, 161(3):653–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/3/653.

Nagai:2000:ELM

[NHI^{+}00] Naoko Nagai, Masanori Hosokawa, Shigeyoshi Itohara, Eijiro Adachi, Takatoshi Matsushita, Nobuko Hosokawa, and Kazuhiro Nagata. Embryonic Lethality of Molecular Chaperone Hsp47 Knockout Mice Is Associated with Defects in Collagen Biosynthesis. Journal of Cell Biology, 150(6):1499– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/6/1499.

Nayler:2000:RPY

[NHS00] Oliver Nayler, Annette M. Hartmann, and Stefan Stamm. The ER Repeat Protein Yt521-B Localizes to a Novel Subnuclear Compartment. Journal of Cell Biology, 150(5):949– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/5/949.

Nakai:2002:MNG

[NK02] Yoko Nakai and Hiroyuki Kamiguchi. Migration of nerve growth cones requires detergent-resistant membranes in a spatially defined and substrate-dependent manner. Journal of Cell Biology, 159(6):1097–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/6/1097.

Noda:2000:ACI

[NKH^{+}00] Takeshi Noda, John Kim, Wei-Pang Huang, Misuzu Baba, Chikara Tokunaga, Yoshinori Ohsumi, and Daniel J. Klionsky. Apg9p/ cvt7p Is an Integral Membrane Protein Required for Transport Vesicle Formation in the Cvt and Autophagy Pathways. Journal of Cell Biology, 148(3):465–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
3/465.

Nilsson:2003:PLN

[NKM^{+}03] IngMarie Nilsson, Daniel J. Kelleher, Yiwei Miao, Yuanlong Shao, Gert Kreibich, Reid Gilmore, Gunnar von Heijne, and Arthur E. Johnson. Photocross-linking of nascent chains to the STT3 subunit of the oligosaccharyltransferase complex. Journal of Cell Biology, 161(4):715–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/4/715.

Nichols:2001:RCL

[NKP^{+}01] Benjamin J. Nichols, Anne K. Kenworthy, Roman S. Polishchuk, Robert Lodge, Theresa H. Roberts, Koret Hirschberg, Robert D. Phair, and Jennifer Lippincott-Schwartz. Rapid Cycling of Lipid Raft Markers between the Cell Surface and Golgi Complex. Journal of Cell Biology, 153 (3):529–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/3/529.

Nabi:2003:CRD

[NL03] Ivan R. Nabi and Phuong U. Le. Caveolae/ raft-dependent endocytosis. Journal of Cell Biology, 161(4):673–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
4/673.

Noren:2000:PCR

[NLBK00] Nicole K. Noren, Betty P. Liu, Keith Burridge, and Bertolt Kreft. P120 Catenin Regulates the Actin Cytoskeleton via Rho Family Gtpases. Journal of Cell Biology, 150(3): 567–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/3/567.

Nakamura:2001:CIM

[NLRD01] Ichiro Nakamura, Lorraine Lipfert, Gideon A. Rodan, and Le T. Duong. Convergence of α_{v}β_{3} integrin- and macrophage colony stimulating factor-mediated signals on phospholipase Cγ in prefusion osteoclasts. Journal of Cell Biology, 152(2): 361–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/2/361.

Nemergut:2000:NIR

[NM00] Michael E. Nemergut and Ian G. Macara. Nuclear Import of the Ran Exchange Factor, Rcc1, Is Mediated by at Least Two Distinct Mechanisms. Journal of Cell Biology, 149(4):835–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
149/4/835.

Norman:2002:SEM

[NM02] Kenneth R. Norman and Donald G. Moerman. α spectrin is essential for morphogenesis and body wall muscle formation in Caenorhabditis elegans. Journal of Cell Biology, 157 (4):665–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/4/665.

Nasrallah:2004:PTE

[NMG04] Ilya M. Nasrallah, Jeremy C. Minarcik, and Jeffrey A. Golden. A polyalanine tract expansion in Arx forms intranuclear inclusions and results in increased cell death. Journal of Cell Biology, 167(3):411–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/3/411.

Nagano:2004:CMI

[NMH^{+}04] Osamu Nagano, Daizo Murakami, Dieter Hartmann, Bart de Strooper, Paul Saftig, Takeshi Iwatsubo, Motowo Nakajima, Masanori Shinohara, and Hideyuki Saya. Cell–matrix interaction via CD44 is independently regulated by different metalloproteinases activated in response to extracellular Ca^{2+} influx and PKC activation. Journal of Cell Biology, 165 (6):893–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/6/893.

Nanba:2003:PRC

[NMHH03] Daisuke Nanba, Akiko Mammoto, Koji Hashimoto, and Shigeki Higashiyama. Proteolytic release of the carboxyterminal fragment of proHB–EGF causes nuclear export of PLZF. Journal of Cell Biology, 163(3):489–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
3/489.

Noble:2000:PCT

[Nob00] Mark Noble. Precursor Cell Transitions in Oligodendrocyte Development. Journal of Cell Biology, 148(5):839–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
5/839.

Nakatsu:2004:DFG

[NOM^{+}04] Fubito Nakatsu, Motohiro Okada, Fumiaki Mori, Noriko Kumazawa, Hiroto Iwasa, Gang Zhu, Yasufumi Kasagi, Haruyuki Kamiya, Akihiro Harada, Kazuhiro Nishimura, Arata Takeuchi, Taisuke Miyazaki, Masahiko Watanabe, Shigeki Yuasa, Toshiya Manabe, Koichi Wakabayashi, Sunao Kaneko, Takashi Saito, and Hiroshi Ohno. Defective function of GABA-containing synaptic vesicles in mice lacking the AP-3B clathrin adaptor. Journal of Cell Biology, 167(2): 293–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/2/293.

Nagata-Ohashi:2004:PNI

[NOOG^{+}04] Kyoko Nagata-Ohashi, Yusaku Ohta, Kazumichi Goto, Shuhei Chiba, Reiko Mori, Michiru Nishita, Kazumasa Ohashi, Kazuyoshi Kousaka, Akihiro Iwamatsu, Ryusuke Niwa, Tadashi Uemura, and Kensaku Mizuno. A pathway of neuregulin-induced activation of cofilin-phosphatase Slingshot and cofilin in lamellipodia. Journal of Cell Biology, 165 (4):465–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/4/465.

Noda:2001:KMM

[NOS^{+}01] Yasuko Noda, Yasushi Okada, Nobuhito Saito, Mitsutoshi Setou, Ying Xu, Zheizeng Zhang, and Nobutaka Hirokawa. KIFC3, a microtubule minus end–directed motor for the apical transport of annexin XIIIb–associated Triton-insoluble membranes. Journal of Cell Biology, 155(1):77–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
1/77.

Nair:2004:PTP

[NPL04] Devi M. Nair, P. Edward Purdue, and Paul B. Lazarow. Pex7p translocates in and out of peroxisomes in Saccharomyces cerevisiae. Journal of Cell Biology, 167(4):599– ??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/4/599.

Neef:2003:PMK

[NPS^{+}03] Rüdiger Neef, Christian Preisinger, Josephine Sutcliffe, Robert Kopajtich, Erich A. Nigg, Thomas U. Mayer, and Francis A. Barr. Phosphorylation of mitotic kinesin-like protein 2 by polo-like kinase 1 is required for cytokinesis. Journal of Cell Biology, 162(5):863–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/5/863.

Neuhaus:2000:MIM

[NS00] Eva M. Neuhaus and Thierry Soldati. A Myosin I Is Involved in Membrane Recycling from Early Endosomes. Journal of Cell Biology, 150(5):1013–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/5/1013.

Newmyer:2001:DIH

[NS01] Sherri L. Newmyer and Sandra L. Schmid. Dominantinterfering Hsc70 Mutants Disrupt Multiple Stages of the Clathrin–Coated Vesicle Cycle in Vivo. Journal of Cell Biology, 152(3):607–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/3/607.

Nagamori:2004:RYF

[NSK04] Shushi Nagamori, Irina N. Smirnova, and H. Ronald Kaback. Role of YidC in folding of polytopic membrane proteins. Journal of Cell Biology, 165(1):53–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/1/53.

Nechushtan:2001:BBC

[NSL^{+}01] Amotz Nechushtan, Carolyn L. Smith, Itschak Lamensdorf, Soo-Han Yoon, and Richard J. Youle. Bax and Bak Coalesce into Novel Mitochondria–Associated Clusters during Apoptosis. Journal of Cell Biology, 153(6):1265–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
6/1265.

Nikonov:2002:ATC

[NSLSK02] Andrei V. Nikonov, Erik Snapp, Jennifer Lippincott-Schwartz, and Gert Kreibich. Active translocon complexes labeled with GFP–Dad1 diffuse slowly as large polysome arrays in the endoplasmic reticulum. Journal of Cell Biology, 158(3):497–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/3/497.

Ng:2000:UPR

[NSW00] Davis T. W. Ng, Eric D. Spear, and Peter Walter. The Unfolded Protein Response Regulates Multiple Aspects of Secretory and Membrane Protein Biogenesis and Endoplasmic Reticulum Quality Control. Journal of Cell Biology, 150 (1):77–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/1/77.

Nikolopoulos:2000:ANF

[NT00] Sotiris N. Nikolopoulos and Christopher E. Turner. Actopaxin, a New Focal Adhesion Protein That Binds Paxillin Ld Motifs and Actin and Regulates Cell Adhesion. Journal of Cell Biology, 151(7):1435–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/7/1435.

Nikitina:2004:CCL

[NW04] Tatiana Nikitina and Christopher L. Woodcock. Closed chromatin loops at the ends of chromosomes. Journal of Cell Biology, 166(2):161–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/2/161.

Norgaard:2001:FDY

[NWT^{+}01] Per Nørgaard, Vibeke Westphal, Christine Tachibana, Lene Alsøe, Bjørn Holst, and Jakob R. Winther. Functional Differences in Yeast Protein Disulfide Isomerases. Journal of Cell Biology, 152(3):553–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/3/553.

Nakagawa:2000:CTB

[NY00] Toshiyuki Nakagawa and Junying Yuan. Cross-talk between Two Cysteine Protease Families. Journal of Cell Biology, 150 (4):887–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/4/887.

Nishimura:2003:LDN

[NYT^{+}03] Kazunari Nishimura, Fumie Yoshihara, Takuro Tojima, Noriko Ooashi, Woohyun Yoon, Katsuhiko Mikoshiba, Vann Bennett, and Hiroyuki Kamiguchi. L1-dependent neuritogenesis involves ankyrin_{B} that mediates L1-CAM coupling with retrograde actin flow. Journal of Cell Biology, 163(5): 1077–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/5/1077.

Nakamura:2001:FSC

[NZA^{+}01] Kimitoshi Nakamura, Anna Zuppini, Serge Arnaudeau, Jeffery Lynch, Irfan Ahsan, Ryoko Krause, Sylvia Papp, Humbert De Smedt, Jan B. Parys, Werner Müller-Esterl, Daniel P. Lew, Karl-Heinz Krause, Nicolas Demaurex, Michal Opas, and Marek Michalak. Functional specialization of calreticulin domains. Journal of Cell Biology, 154(5):961– ??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/5/961.

Novoa:2001:FIU

[NZHR01] Isabel Novoa, Huiqing Zeng, Heather P. Harding, and David Ron. Feedback inhibition of the unfolded protein response by GADD34-mediated dephosphorylation of eIF2α. Journal of Cell Biology, 153(5):1011–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/5/1011.

Orci:2000:EGR

[OAR^{+}00] Lelio Orci, Mylène Amherdt, Mariella Ravazzola, Alain Perrelet, and James E. Rothman. Exclusion of Golgi Residents from Transport Vesicles Budding from Golgi Cisternae in Intact Cells. Journal of Cell Biology, 150(6):1263–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
150/6/1263.

Ohtani:2003:EBV

[OBG^{+}03] Naoko Ohtani, Paul Brennan, Stefan Gaubatz, Elaine Sanij, Paul Hertzog, Ernst Wolvetang, Jacques Ghysdael, Martin Rowe, and Eiji Hara. Epstein–Barr virus LMP1 blocks p16^{INK4a}-RB pathway by promoting nuclear export of E2F4/ 5. Journal of Cell Biology, 162(2):173–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/2/173.

Otte:2001:EE

[OBH^{+}01] Stefan Otte, William J. Belden, Matthew Heidtman, Jay Liu, Ole N. Jensen, and Charles Barlowe. Erv41p and Erv46p. Journal of Cell Biology, 152(3):503–??, February 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/152/3/
503.

Oegema:2001:FAK

[ODR^{+}01] Karen Oegema, Arshad Desai, Sonja Rybina, Matthew Kirkham, and Anthony A. Hyman. Functional Analysis of Kinetochore Assembly in Caenorhabditis elegans. Journal of Cell Biology, 153(6):1209–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/6/1209.

Obergfell:2002:CIC

[OEM^{+}02] Achim Obergfell, Koji Eto, Attila Mocsai, Charito Buensuceso, Sheri L. Moores, Joan S. Brugge, Clifford A. Lowell, and Sanford J. Shattil. Coordinate interactions of Csk, Src, and Syk kinases with αIIbβ3 initiate integrin signaling to the cytoskeleton. Journal of Cell Biology, 157(2):265–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/157/
2/265.

Ohta:2002:CCN

[OER^{+}02] Toshiro Ohta, Russell Essner, Jung-Hwa Ryu, Robert E. Palazzo, Yumi Uetake, and Ryoko Kuriyama. Characterization of Cep135, a novel coiled-coil centrosomal protein involved in microtubule organization in mammalian cells. Journal of Cell Biology, 156(1):87–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/1/87.

Oliveria:2003:IKA

[OGD03] Seth F. Oliveria, Lisa L. Gomez, and Mark L. Dell’Acqua. Imaging kinase–AKAP79–phosphatase scaffold complexes at the plasma membrane in living cells using FRET microscopy. Journal of Cell Biology, 160(1):101–??, January 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/1/
101.

Osman:2002:ILS

[OKC02] Mahasin A. Osman, James B. Konopka, and Richard A. Cerione. Iqg1p links spatial and secretion landmarks to polarity and cytokinesis. Journal of Cell Biology, 159(4):601– ??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/4/601.

Okamoto:2001:PRC

[OKM^{+}01] Isamu Okamoto, Yoshiaki Kawano, Daizo Murakami, Takashi Sasayama, Norie Araki, Toru Miki, Albert J. Wong, and Hideyuki Saya. Proteolytic release of CD44 intracellular domain and its role in the CD44 signaling pathway. Journal of Cell Biology, 155(5):755–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/5/755.

Oliferenko:2000:HAB

[OKSH00] Snezhana Oliferenko, Irina Kaverina, J. Victor Small, and Lukas A. Huber. Hyaluronic Acid (Ha) Binding to Cd44 Activates Rac1 and Induces Lamellipodia Outgrowth. Journal of Cell Biology, 148(6):1159–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/6/1159.

Ogg:2002:CBC

[OL02] Stephen C. Ogg and Angus I. Lamond. Cajal bodies and coilin — moving towards function. Journal of Cell Biology, 159(1):17–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/1/17.

Ojima:2000:DFZ

[OLB^{+}00] K. Ojima, Z. X. Lin, M.-L. Bang, S. Holtzer, R. Matsuda, S. Labeit, H. L. Sweeney, and H. Holtzer. Distinct Families of Z–Line Targeting Modules in the Cooh–Terminal Region of Nebulin. Journal of Cell Biology, 150(3):553–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
3/553.

Olazabal:2001:ACN

[OM01] Isabel M. Olazabal and Laura M. Machesky. Abp1p and cortactin, new “hand-holds” for actin. Journal of Cell Biology, 154(4):679–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/4/679.

Orike:2001:RPK

[OMB^{+}01] Nina Orike, Gayle Middleton, Emma Borthwick, Vladimir Buchman, Timothy Cowen, and Alun M. Davies. Role of PI 3-kinase, Akt and Bcl-2–related proteins in sustaining the survival of neurotrophic factor–independent adult sympathetic neurons. Journal of Cell Biology, 154(5):995– ??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/5/995.

Osawa:2002:ERP

[OMiKF02] Masaki Osawa, Michitaka Masuda, Ken ichi Kusano, and Keigi Fujiwara. Evidence for a role of platelet endothelial cell adhesion molecule-1 in endothelial cell mechanosignal transduction. Journal of Cell Biology, 158(4):773–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
4/773.

OToole:2003:MDM

[OMM^{+}03] Eileen T. O’Toole, Kent L. McDonald, Jana Mäntler, J. Richard McIntosh, Anthony A. Hyman, and Thomas Müller-Reichert. Morphologically distinct microtubule ends in the mitotic centrosome of Caenorhabditis elegans. Journal of Cell Biology, 163(3):451–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/3/451.

Ortiz:2002:YRS

[OMWSN02] Darinel Ortiz, Martina Medkova, Christiane Walch-Solimena, and Peter Novick. Ypt32 recruits the Sec4p guanine nucleotide exchange factor, Sec2p, to secretory vesicles; evidence for a Rab cascade in yeast. Journal of Cell Biology, 157 (6):1005–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/6/1005.

OConnor:2000:RFL

[ONM00] Kathleen L. O’Connor, Bao-Kim Nguyen, and Arthur M. Mercurio. Rhoa Function in Lamellae Formation and Migration Is Regulated by the α6β4 Integrin and Camp Metabolism. Journal of Cell Biology, 148(2):253–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
2/253.

Ono:2001:CEU

[Ono01] Shoichiro Ono. The Caenorhabditis elegans unc-78 Gene Encodes a Homologue of Actin–Interacting Protein 1 Required for Organized Assembly of Muscle Actin Filaments. Journal of Cell Biology, 152(6):1313–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/6/1313.

Okagaki:2000:ASM

[ONS^{+}00] Tsuyoshi Okagaki, Akio Nakamura, Tomohiko Suzuki, Kazuhiro Ohmi, and Kazuhiro Kohama. Assembly of Smooth Muscle Myosin by the 38k Protein, a Homologue of a Subunit of Pre-mRNA Splicing Factor-2. Journal of Cell Biology, 148(4):653–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/4/653.

Ono:2002:TIA

[OO02] Shoichiro Ono and Kanako Ono. Tropomyosin inhibits ADF/ cofilin-dependent actin filament dynamics. Journal of Cell Biology, 156(6):1065–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/6/1065.

Orr:2003:LDL

[OPP^{+}03] Anthony Wayne Orr, Claudio E. Pedraza, Manuel Antonio Pallero, Carrie A. Elzie, Silvia Goicoechea, Dudley K. Strickland, and Joanne E. Murphy-Ullrich. Low density lipoprotein receptor–related protein is a calreticulin coreceptor that signals focal adhesion disassembly. Journal of Cell Biology, 161 (6):1179–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/6/1179.

Ostermeyer:2001:ACE

[OPZ^{+}01] Anne G. Ostermeyer, James M. Paci, Youchun Zeng, Douglas M. Lublin, Sean Munro, and Deborah A. Brown. Accumulation of Caveolin in the Endoplasmic Reticulum Redirects the Protein to Lipid Storage Droplets. Journal of Cell Biology, 152(5):1071–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/5/1071.

Ostermeyer:2004:RHD

[ORZ^{+}04] Anne G. Ostermeyer, Lynne T. Ramcharan, Youchun Zeng, Douglas M. Lublin, and Deborah A. Brown. Role of the hydrophobic domain in targeting caveolin-1 to lipid droplets. Journal of Cell Biology, 164(1):69–??, January 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/1/
69.

Oehlmann:2004:RCE

[OSB04] Maren Oehlmann, Alan J. Score, and J. Julian Blow. The role of Cdc6 in ensuring complete genome licensing and S phase checkpoint activation. Journal of Cell Biology, 165 (2):181–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/2/181.

Oegema:2000:FAH

[OSMF00] Karen Oegema, Matthew S. Savoian, Timothy J. Mitchison, and Christine M. Field. Functional Analysis of a Human Homologue of the Drosophila Actin Binding Protein Anillin Suggests a Role in Cytokinesis. Journal of Cell Biology, 150 (3):539–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/3/539.

Ochoa:2000:FLB

[OSN^{+}00] Gian-Carlo Ochoa, Vladimir I. Slepnev, Lynn Neff, Niels Ringstad, Kohji Takei, Laurie Daniell, Warren Kim, Hong Cao, Mark McNiven, Roland Baron, and Pietro De Camilli. A Functional Link between Dynamin and the Actin Cytoskeleton at Podosomes. Journal of Cell Biology, 150(2): 377–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/2/377.

Ohba:2004:ETD

[OSNG04] Tomoyuki Ohba, Eric C. Schirmer, Takeharu Nishimoto, and Larry Gerace. Energy- and temperature-dependent transport of integral proteins to the inner nuclear membrane via the nuclear pore. Journal of Cell Biology, 167(6):1051–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
167/6/1051.

Offterdinger:2002:CE

[OSWG02] Martin Offterdinger, Christian Schöfer, Klara Weipoltshammer, and Thomas W. Grunt. c-erbB-3. Journal of Cell Biology, 157(6):929–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/6/929.

OCallaghan:2003:DCS

[OTB03] Dermott W. O’Callaghan, Alexei V. Tepikin, and Robert D. Burgoyne. Dynamics and calcium sensitivity of the Ca^{2+} /myristoyl switch protein hippocalcin in living cells. Journal of Cell Biology, 163(4):715–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/4/715.

Ohtsuka:2002:C

[OTRI^{+}02] Toshihisa Ohtsuka, Etsuko Takao-Rikitsu, Eiji Inoue, Marie Inoue, Masakazu Takeuchi, Kaho Matsubara, Maki Deguchi-Tawarada, Keiko Satoh, Koji Morimoto, Hiroyuki Nakanishi, and Yoshimi Takai. Cast. Journal of Cell Biology, 158(3): 577–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/3/577.

Ostergaard:2004:MDB

[ØTW04] Henrik Østergaard, Christine Tachibana, and Jakob R. Winther. Monitoring disulfide bond formation in the eukaryotic cytosol. Journal of Cell Biology, 166(3):337–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
3/337.

Okamura:2004:CAR

[OTY^{+}04] Ko Okamura, Hidekazu Tanaka, Yoshiki Yagita, Yoshinaga Saeki, Akihiko Taguchi, Yasushi Hiraoka, Ling-Hui Zeng, David R. Colman, and Naomasa Miki. Cadherin activity is required for activity-induced spine remodeling. Journal of Cell Biology, 167(5):961–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/5/961.

Ovechkina:2002:KLI

[OWW02] Yulia Ovechkina, Michael Wagenbach, and Linda Wordeman. K-loop insertion restores microtubule depolymerizing activity of a “neckless” MCAK mutant. Journal of Cell Biology, 159(4):557–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/4/557.

Portnoy:2002:CBL

[PAG02] Daniel A. Portnoy, Victoria Auerbuch, and Ian J. Glomski. The cell biology of Listeria monocytogenes infection. Journal of Cell Biology, 158(3):409–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/3/409.

Peterson:2004:CMS

[PAR^{+}04] Richard S. Peterson, Roma A. Andhare, Kathleen T. Rousche, Warren Knudson, Weihua Wang, Jami B. Grossfield, Raymond O. Thomas, Robert E. Hollingsworth, and Cheryl B. Knudson. CD44 modulates Smad1 activation in the BMP-7 signaling pathway. Journal of Cell Biology, 166(7):1081–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/7/1081.

Pardo:2004:APS

[PBB^{+}04] Julián Pardo, Alberto Bosque, Reina Brehm, Reinhard Wallich, Javier Naval, Arno Müllbacher, Alberto Anel, and Markus M. Simon. Apoptotic pathways are selectively activated by granzyme A and/or granzyme B in CTL-mediated target cell lysis. Journal of Cell Biology, 167(3):457–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/167/3/457.

Pazour:2002:ITP

[PBD^{+}02] Gregory J. Pazour, Sheila A. Baker, James A. Deane, Douglas G. Cole, Bethany L. Dickert, Joel L. Rosenbaum, George B. Witman, and Joseph C. Besharse. The intraflagellar transport protein, IFT88, is essential for vertebrate photoreceptor assembly and maintenance. Journal of Cell Biology, 157(1):103–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/1/103.

Parkinson:2004:KIJ

[PBD^{+}04] David B. Parkinson, Ambily Bhaskaran, Anna Droggiti, Sarah Dickinson, Maurizio D’Antonio, Rhona Mirsky, and Kristjan R. Jessen. Krox-20 inhibits Jun–NH_{2}-terminal kinase/c–Jun to control Schwann cell proliferation and death. Journal of Cell Biology, 164(3):385–??, February 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/3/
385.

Petit:2000:PTR

[PBL^{+}00] Valérie Petit, Brigitte Boyer, Delphine Lentz, Christopher E. Turner, Jean Paul Thiery, and Ana M. Vallés. Phosphorylation of Tyrosine Residues 31 and 118 on Paxillin Regulates Cell Migration through an Association with Crk in Nbt–II Cells. Journal of Cell Biology, 148(5):957–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
5/957.

Paumet:2001:SEP

[PBP^{+}01] Fabienne Paumet, Britta Brügger, Francesco Parlati, James A. McNew, Thomas H. Söllner, and James E. Rothman. A t–SNARE of the endocytic pathway must be activated for fusion. Journal of Cell Biology, 155(6):961–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/155/6/961.

Prudovsky:2002:ITC

[PBT^{+}02] Igor Prudovsky, Cinzia Bagala, Francesca Tarantini, Anna Mandinova, Raffaella Soldi, Stephen Bellum, and Thomas Maciag. The intracellular translocation of the components of the fibroblast growth factor 1 release complex precedes their assembly prior to export. Journal of Cell Biology, 158 (2):201–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/2/201.

Polizotto:2001:CDN

[PC01] Renée S. Polizotto and Martha S. Cyert. Calcineurindependent nuclear import of the transcription factor Crz1p requires Nmd5p. Journal of Cell Biology, 154(5):951–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/154/5/951.

Pettitt:2003:CEP

[PCB^{+}03] Jonathan Pettitt, Elisabeth A. Cox, Ian D. Broadbent, Aileen Flett, and Jeff Hardin. The Caenorhabditis elegans p120 catenin homologue, JAC-1, modulates cadherin–catenin function during epidermal morphogenesis. Journal of Cell Biology, 162(1):15–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/1/15.

Porrello:2000:PRM

[PCC^{+}00] Alessandro Porrello, Maria Antonietta Cerone, Sabrina Coen, Aymone Gurtner, Giulia Fontemaggi, Letizia Cimino, Giulia Piaggio, Ada Sacchi, and Silvia Soddu. P53 Regulates Myogenesis by Triggering the Differentiation Activity of Prb. Journal of Cell Biology, 151(6):1295–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
6/1295.

Pankov:2000:IDM

[PCK^{+}00] Roumen Pankov, Edna Cukierman, Ben-Zion Katz, Kazue Matsumoto, Diane C. Lin, Shin Lin, Cornelia Hahn, and Kenneth M. Yamada. Integrin Dynamics and Matrix Assembly. Journal of Cell Biology, 148(5):1075–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/5/
1075.

Pellizzoni:2001:FIB

[PCR^{+}01] Livio Pellizzoni, Bernard Charroux, Juri Rappsilber, Matthias Mann, and Gideon Dreyfuss. A Functional Interaction between the Survival Motor Neuron Complex and RNA Polymerase II. Journal of Cell Biology, 152(1):75–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
1/75.

Putcha:2000:IAS

[PDJ00] Girish V. Putcha, Mohanish Deshmukh, and Eugene M. Johnson. Inhibition of Apoptotic Signaling Cascades Causes Loss of Trophic Factor Dependence during Neuronal Maturation. Journal of Cell Biology, 149(5):1011–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/149/5/
1011.

Parisi:2003:NDC

[PDL^{+}03] Silvia Parisi, Daniela D’Andrea, Carmine T. Lago, Eileen D. Adamson, M. Graziella Persico, and Gabriella Minchiotti. Nodal-dependent Cripto signaling promotes cardiomyogenesis and redirects the neural fate of embryonic stem cells. Journal of Cell Biology, 163(2):303–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/2/303.

Papp:2003:AER

[PDMO03] S. Papp, E. Dziak, M. Michalak, and M. Opas. Is all of the endoplasmic reticulum created equal? The effects of the heterogeneous distribution of endoplasmic reticulum Ca^{2+}-handling proteins. Journal of Cell Biology, 160(4):475–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
4/475.

Prigent:2003:ACP

[PDR^{+}03] Magali Prigent, Thierry Dubois, Graça Raposo, Valérie Derrien, Danièle Tenza, Carine Rossé, Jacques Camonis, and Philippe Chavrier. ARF6 controls post-endocytic recycling through its downstream exocyst complex effector. Journal of Cell Biology, 163(5):1111–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/5/1111.

Porter:2002:HS

[PDT^{+}02] Lisa A. Porter, Ryan W. Dellinger, John A. Tynan, Elizabeth A. Barnes, Monica Kong, Jean-Luc Lenormand, and Daniel J. Donoghue. Human Speedy. Journal of Cell Biology, 157(3):357–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/3/357.

Pazour:2000:CIM

[PDV^{+}00] Gregory J. Pazour, Bethany L. Dickert, Yvonne Vucica, E. Scott Seeley, Joel L. Rosenbaum, George B. Witman, and Douglas G. Cole. Chlamydomonas IFT 88 and Its Mouse Homologue, Polycystic Kidney Disease Gene Tg 737, Are Required for Assembly of Cilia and Flagella. Journal of Cell Biology, 151(3):709–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/3/709.

Plath:2000:NFT

[PDW^{+}00] Thomas Plath, Katharina Detjen, Martina Welzel, Zofia von Marschall, Derek Murphy, Michael Schirner, Bertram Wiedenmann, and Stefan Rosewicz. A Novel Function for the Tumor Suppressor p16^{INK4a}. Journal of Cell Biology, 150(6):1467–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/6/1467.

Pederson:2004:CTP

[Ped04] Thoru Pederson. Can telomerase be put in its place? Journal of Cell Biology, 164(5):637–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/5/637.

Pelham:2001:TTG

[Pel01] Hugh R. B. Pelham. Traffic through the Golgi apparatus. Journal of Cell Biology, 155(7):1099–??, December 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/155/7/
1099.

Piacentini:2003:HPH

[PFB^{+}03] Lucia Piacentini, Laura Fanti, Maria Berloco, Barbara Perrini, and Sergio Pimpinelli. Heterochromatin protein 1 (HP1) is associated with induced gene expression in Drosophila euchromatin. Journal of Cell Biology, 161(4):707–??, May 2003.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/4/
707.

Pfeffer:2001:CGC

[Pfe01] Suzanne R. Pfeffer. Constructing a Golgi complex. Journal of Cell Biology, 155(6):873–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/6/873.

Pinton:2000:RLI

[PFM^{+}00] Paolo Pinton, Davide Ferrari, Paulo Magalhães, Klaus Schulze-Osthoff, Francesco Di Virgilio, Tullio Pozzan, and Rosario Rizzuto. Reduced Loading of Intracellular Ca^{2+} Stores and Downregulation of Capacitative Ca^{2+} Influx in Bcl-2–Overexpressing Cells. Journal of Cell Biology, 148(5): 857–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/5/857.

Petiot:2003:PSR

[PFSG03] A. Petiot, J. Fauré, H. Stenmark, and J. Gruenberg. PI3P signaling regulates receptor sorting but not transport in the endosomal pathway. Journal of Cell Biology, 162(6):971– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/6/971.

Pan:2000:GHG

[PFW^{+}00] Julie Y. Pan, William E. Fieles, Anne M. White, Mark M. Egerton, and David S. Silberstein. Ges, a human GTPase of the Rad/Gem/Kir family, promotes endothelial cell sprouting and cytoskeleton reorganization. Journal of Cell Biology, 149 (5):1107–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/5/1107.

Peng:2002:SPB

[PG02] Renwang Peng and Dieter Gallwitz. Sly1 protein bound to Golgi syntaxin Sed5p allows assembly and contributes to specificity of SNARE fusion complexes. Journal of Cell Biology, 157(4):645–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/4/645.

Palmer:2001:SMS

[PGS^{+}01] Steve Palmer, Nicola Groves, Aaron Schindeler, Thomas Yeoh, Christine Biben, Cheng-Chun Wang, Duncan B. Sparrow, Louise Barnett, Nancy A. Jenkins, Neal G. Copeland, Frank Koentgen, Tim Mohun, and Richard P. Harvey. The Small Muscle–Specific Protein Csl Modifies Cell Shape and Promotes Myocyte Fusion in an Insulin-like Growth Factor 1– Dependent Manner. Journal of Cell Biology, 153(5):985–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/5/985.

Palmer:2001:VDP

[PGSE^{+}01] Héctor G. Pálmer, José Manuel González-Sancho, Jesús Espada, María T. Berciano, Isabel Puig, Josep Baulida, Miguel Quintanilla, Amparo Cano, Antonio García de Herreros, Miguel Lafarga, and Alberto Muñoz. Vitamin D_{3} promotes the differentiation of colon carcinoma cells by the induction of E-cadherin and the inhibition of β-catenin signaling. Journal of Cell Biology, 154(2):369–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/2/369.

Platani:2000:VAC

[PGSL00] Melpomeni Platani, Ilya Goldberg, Jason R. Swedlow, and Angus I. Lamond. In Vivo Analysis of Cajal Body Movement, Separation, and Joining in Live Human Cells. Journal of Cell Biology, 151(7):1561–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/7/1561.

Prinz:2000:MAS

[PGV^{+}00] William A. Prinz, Lara Grzyb, Marten Veenhuis, Jason A. Kahana, Pamela A. Silver, and Tom A. Rapoport. Mutants Affecting the Structure of the Cortical Endoplasmic Reticulum in Saccharomyces cerevisiae. Journal of Cell Biology, 150 (3):461–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/3/461.

Putcha:2002:IEP

[PHM^{+}02] Girish V. Putcha, Charles A. Harris, Krista L. Moulder, Rachael M. Easton, Craig B. Thompson, and Eugene M. Johnson. Intrinsic and extrinsic pathway signaling during neuronal apoptosis. Journal of Cell Biology, 157(3):441–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
157/3/441.

Pornillos:2003:HGM

[PHS^{+}03] Owen Pornillos, Daniel S. Higginson, Kirsten M. Stray, Robert D. Fisher, Jennifer E. Garrus, Marielle Payne, Gong-Ping He, Hubert E. Wang, Scott G. Morham, and Wesley I. Sundquist. HIV Gag mimics the Tsg101-recruiting activity of the human Hrs protein. Journal of Cell Biology, 162(3): 425–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/3/425.

Pepperkok:2000:IDM

[PHWK^{+}00] Rainer Pepperkok, Agnes Hotz-Wagenblatt, Norbert König, Andreas Girod, Dirk Bossemeyer, and Volker Kinzel. Intracellular Distribution of Mammalian Protein Kinase a Catalytic Subunit Altered by Conserved Asn2 Deamidation. Journal of Cell Biology, 148(4):715–??, February 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/4/
715.

He:2000:ICC

[pHYpXL00] Xiang ping He, Feng Yang, Zuo ping Xie, and Bai Lu. Intracellular Ca^{2+} and Ca^{2+} /calmodulin–dependent kinase II mediate acute potentiation of neurotransmitter release by neurotrophin-3. Journal of Cell Biology, 149(4):783–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
4/783.

Pralle:2000:SCR

[PKF^{+}00] A. Pralle, P. Keller, E.-L. Florin, K. Simons, and J. K. H. Hörber. Sphingolipid–cholesterol Rafts Diffuse as Small Entities in the Plasma Membrane of Mammalian Cells. Journal of Cell Biology, 148(5):997–??, March 2000. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/148/5/997.

Palmada:2002:CJE

[PKR^{+}02] M. Palmada, S. Kanwal, N. J. Rutkoski, C. Gustafson-Brown, R. S. Johnson, R. Wisdom, and B. D. Carter. c-jun is essential for sympathetic neuronal death induced by NGF withdrawal but not by p75 activation. Journal of Cell Biology, 158(3):453–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/3/453.

Puthenveedu:2001:EGS

[PL01] Manojkumar A. Puthenveedu and Adam D. Linstedt. Evidence that Golgi structure depends on a p115 activity that is independent of the vesicle tether components giantin and GM130. Journal of Cell Biology, 155(2):227–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
2/227.

Prasad:2002:PKR

[PLC^{+}02] Sathyamangla V. Naga Prasad, Stéphane A. Laporte, Dean Chamberlain, Marc G. Caron, Larry Barak, and Howard A. Rockman. Phosphoinositide 3-kinase regulates β2-adrenergic receptor endocytosis by AP-2 recruitment to the receptor/βarrestin complex. Journal of Cell Biology, 158(3):563–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
158/3/563.

Preuss:2001:MNC

[PLH^{+}01] Marc Preuss, Klaus Leonhard, Kai Hell, Rosemary A. Stuart, Walter Neupert, and Johannes M. Herrmann. Mba1, a Novel Component of the Mitochondrial Protein Export Machinery of the Yeast Saccharomyces cerevisiae. Journal of Cell Biology, 153(5):1085–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/5/1085.

Plant:2000:AMT

[PLL^{+}00] Pamela J. Plant, Frank Lafont, Sandra Lecat, Paul Verkade, Kai Simons, and Daniela Rotin. Apical Membrane Targeting of Nedd4 Is Mediated by an Association of Its C2 Domain with Annexin Xiiib. Journal of Cell Biology, 149(7):1473–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
149/7/1473.

Pol:2001:CDN

[PLL^{+}01] Albert Pol, Robert Luetterforst, Margaret Lindsay, Sanna Heino, Elina Ikonen, and Robert G. Parton. A Caveolin Dominant Negative Mutant Associates with Lipid Bodies and Induces Intracellular Cholesterol Imbalance. Journal of Cell Biology, 152(5):1057–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/5/1057.

Politz:2002:SRP

[PLP02] Joan C. Politz, Laura B. Lewandowski, and Thoru Pederson. Signal recognition particle RNA localization within the nucleolus differs from the classical sites of ribosome synthesis. Journal of Cell Biology, 159(3):411–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/3/
411.

Pacquelet:2003:BSP

[PLR03] Anne Pacquelet, Li Lin, and Pernille Rørth. Binding site for p120/δ-catenin is not required for Drosophila E-cadherin function in vivo. Journal of Cell Biology, 160(3):313–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/160/3/313.

Pinton:2004:LTM

[PLW^{+}04] Paolo Pinton, Sara Leo, Mariusz R. Wieckowski, Giulietta Di Benedetto, and Rosario Rizzuto. Long-term modulation of mitochondrial Ca^{2+} signals by protein kinase C isozymes. Journal of Cell Biology, 165(2):223–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/2/223.

Pryor:2000:RIC

[PMB^{+}00] Paul R. Pryor, Barbara M. Mullock, Nicholas A. Bright, Sally R. Gray, and J. Paul Luzio. The Role of Intraorganellar Ca^{2+} In Late Endosome–Lysosome Heterotypic Fusion and in the Reformation of Lysosomes from Hybrid Organelles. Journal of Cell Biology, 149(5):1053–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/5/1053.

Perez-Mongiovi:2000:NMA

[PMBC^{+}00] Daniel Pérez-Mongiovi, Clare Beckhelling, Patrick Chang, Christopher C. Ford, and Evelyn Houliston. Nuclei and microtubule asters stimulate maturation/M phase promoting factor (Mpf) activation in Xenopus eggs and egg cytoplasmic extracts. Journal of Cell Biology, 150(5):963–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
5/963.

Pereira:2002:RBB

[PMGS02] Gislene Pereira, Claire Manson, Joan Grindlay, and Elmar Schiebel. Regulation of the Bfa1p–Bub2p complex at spindle pole bodies by the cell cycle phosphatase Cdc14p. Journal of Cell Biology, 157(3):367–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/3/367.

Piel:2000:RCM

[PMK^{+}00] Matthieu Piel, Pablo Meyer, Alexey Khodjakov, Conly L. Rieder, and Michel Bornens. The Respective Contributions of the Mother and Daughter Centrioles to Centrosome Activity and Behavior in Vertebrate Cells. Journal of Cell Biology, 149 (2):317–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/2/317.

Pelchen-Matthews:2003:IHA

[PMKM03] Annegret Pelchen-Matthews, Beatrice Kramer, and Mark Marsh. Infectious HIV-1 assembles in late endosomes in primary macrophages. Journal of Cell Biology, 162(3):443–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/162/3/443.

Park:2001:EGR

[PMKV01] Song-Kyu Park, Robert Miller, Ian Krane, and Timothy Vartanian. The erbB2 gene is required for the development of terminally differentiated spinal cord oligodendrocytes. Journal of Cell Biology, 154(6):1245–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/6/1245.

Peters:2003:TPP

[PMP^{+}03] Peter J. Peters, Alexander Mironov, David Peretz, Elly van Donselaar, Estelle Leclerc, Susanne Erpel, Stephen J. DeArmond, Dennis R. Burton, R. Anthony Williamson, Martin Vey, and Stanley B. Prusiner. Trafficking of prion proteins through a caveolae-mediated endosomal pathway. Journal of Cell Biology, 162(4):703–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/4/703.

Prior:2003:DVR

[PMPH03] Ian A. Prior, Cornelia Muncke, Robert G. Parton, and John F. Hancock. Direct visualization of Ras proteins in spatially distinct cell surface microdomains. Journal of Cell Biology, 160(2):165–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/2/165.

Pearson:2001:BYC

[PMSB01] Chad G. Pearson, Paul S. Maddox, E. D. Salmon, and Kerry Bloom. Budding Yeast Chromosome Structure and Dynamics during Mitosis. Journal of Cell Biology, 152(6): 1255–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/6/1255.

Poliak:2002:DCA

[PMU^{+}02] Sebastian Poliak, Sean Matlis, Christoph Ullmer, Steven S. Scherer, and Elior Peles. Distinct claudins and associated PDZ proteins form different autotypic tight junctions in myelinating Schwann cells. Journal of Cell Biology, 159(2): 361–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/2/361.

Paliulis:2000:RCN

[PN00] Leocadia V. Paliulis and R. Bruce Nicklas. The Reduction of Chromosome Number in Meiosis Is Determined by Properties Built into the Chromosomes. Journal of Cell Biology, 150(6):1223–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/6/1223.

Poon:2001:GAA

[PNSJ01] Pak Phi Poon, Steven F. Nothwehr, Richard A. Singer, and Gerald C. Johnston. The Gcs1 and Age2 ArfGAP proteins provide overlapping essential function for transport from the yeast trans-Golgi network. Journal of Cell Biology, 155(7): 1239–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/7/1239.

Parast:2000:CPN

[PO00] Mana M. Parast and Carol A. Otey. Characterization of Palladin, a Novel Protein Localized to Stress Fibers and Cell Adhesions. Journal of Cell Biology, 150(3):643–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
3/643.

Peden:2004:LAA

[POH^{+}04] Andrew A. Peden, Viola Oorschot, Boris A. Hesser, Cary D. Austin, Richard H. Scheller, and Judith Klumperman. Localization of the AP-3 adaptor complex defines a novel endosomal exit site for lysosomal membrane proteins. Journal of Cell Biology, 164(7):1065–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/7/1065.

Palmgren:2001:IPA

[POW^{+}01a] Sandra Palmgren, Pauli J. Ojala, Martin A. Wear, John A. Cooper, and Pekka Lappalainen. Interactions with PIP_{2}, ADP-actin monomers, and capping protein regulate the activity and localization of yeast twinfilin. Journal of Cell Biology, 155(2):251–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/2/251.

Powledge:2001:CCR

[Pow01b] Tabitha M. Powledge. Cell cycling with Rac. Journal of Cell Biology, 154(4):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/4/677.1.

Powledge:2001:CJ

[Pow01c] Tabitha M. Powledge. Christmas in July. Journal of Cell Biology, 154(2):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/2/255.1.

Powledge:2001:FSR

[Pow01d] Tabitha M. Powledge. Fail-safe replication. Journal of Cell Biology, 154(2):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/2/254.1.

Powledge:2001:FML

[Pow01e] Tabitha M. Powledge. The (formerly) missing link. Journal of Cell Biology, 154(4):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/4/677.2.

Powledge:2001:GG

[Pow01f] Tabitha M. Powledge. Garbage in, garbage out. Journal of Cell Biology, 154(4):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/4/676.1.

Powledge:2001:KLD

[Pow01g] Tabitha M. Powledge. Keeping the lid on destruction. Journal of Cell Biology, 154(2):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/2/255.2.

Powledge:2001:RI

[Pow01h] Tabitha M. Powledge. Running interference. Journal of Cell Biology, 154(4):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/4/676.2.

Powledge:2001:SLP

[Pow01i] Tabitha M. Powledge. Shedding light on protein folding. Journal of Cell Biology, 154(4):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/4/677.3.

Powledge:2001:TL

[Pow01j] Tabitha M. Powledge. Tensegrity lives. Journal of Cell Biology, 154(2):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/2/254.2.

Powell:2004:ASG

[Pow04] Kendall Powell. All systems go. Journal of Cell Biology, 165 (3):299–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/3/299.

Pederson:2000:NFR

[PP00] Thoru Pederson and Joan C. Politz. The Nucleolus and the Four Ribonucleoproteins of Translation. Journal of Cell Biology, 148(6):1091–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/6/1091.

Popsueva:2003:GPT

[PPA^{+}03] Anna Popsueva, Dmitry Poteryaev, Elena Arighi, Xiaojuan Meng, Alexandre Angers-Loustau, David Kaplan, Mart Saarma, and Hannu Sariola. GDNF promotes tubulogenesis of GFRα1-expressing MDCK cells by Src-mediated phosphorylation of Met receptor tyrosine kinase. Journal of Cell Biology, 161(1):119–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/1/119.

Perez:2002:CNT

[PPGN^{+}02] Franck Perez, Karin Pernet-Gallay, Clément Nizak, Holly V. Goodson, Thomas E. Kreis, and Bruno Goud. CLIPR-59, a new trans-Golgi/TGN cytoplasmic linker protein belonging to the CLIP-170 family. Journal of Cell Biology, 156(4): 631–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/4/631.

Priller:2001:NCP

[PPK^{+}01] Josef Priller, Derek A. Persons, Francisco F. Klett, Gerd Kempermann, Georg W. Kreutzberg, and Ulrich Dirnagl. Neogenesis of cerebellar Purkinje neurons from gene-marked bone marrow cells in vivo. Journal of Cell Biology, 155(5): 733–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/5/733.

Polishchuk:2000:CLE

[PPM^{+}00] Roman S. Polishchuk, Elena V. Polishchuk, Pierfrancesco Marra, Saverio Alberti, Roberto Buccione, Alberto Luini, and Alexander A. Mironov. Correlative Light–Electron Microscopy Reveals the Tubular–Saccular Ultrastructure of Carriers Operating between Golgi Apparatus and Plasma Membrane. Journal of Cell Biology, 148(1):45–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
1/45.

Pavlov:2001:NHC

[PPP^{+}01] Evgeny V. Pavlov, Muriel Priault, Dawn Pietkiewicz, Emily H.-Y. Cheng, Bruno Antonsson, Stephen Manon, Stanley J. Korsmeyer, Carmen A. Mannella, and Kathleen W. Kinnally. A novel, high conductance channel of mitochondria linked to apoptosis in mammalian cells and Bax expression in yeast. Journal of Cell Biology, 155(5):725–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
5/725.

Peretti:2000:EIK

[PPR^{+}00] Diego Peretti, Leticia Peris, Silvana Rosso, Santiago Quiroga, and Alfredo Cáceres. Evidence for the Involvement of Kif4 in the Anterograde Transport of L1-Containing Vesicles. Journal of Cell Biology, 149(1):141–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/1/141.

Plafker:2004:UCH

[PPWM04] Scott M. Plafker, Kendra S. Plafker, Allan M. Weissman, and Ian G. Macara. Ubiquitin charging of human class III ubiquitin-conjugating enzymes triggers their nuclear import. Journal of Cell Biology, 167(4):649–??, November 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/167/4/
649.

Previtali:2000:ETP

[PQF^{+}00] Stefano C. Previtali, Angelo Quattrini, Marina Fasolini, Maria Carla Panzeri, Antonello Villa, Marie T. Filbin, Wenhui Li, Shing-Yan Chiu, Albee Messing, Lawrence Wrabetz, and M. Laura Feltri. Epitope-tagged P_{0} Glycoprotein Causes Charcot–Marie–Tooth-like Neuropathy in Transgenic Mice. Journal of Cell Biology, 151(5):1035–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/5/
1035.

Plath:2000:SRC

[PR00] Kathrin Plath and Tom A. Rapoport. Spontaneous Release of Cytosolic Proteins from Posttranslational Substrates before Their Transport into the Endoplasmic Reticulum. Journal of Cell Biology, 151(1):167–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/1/167.

Paul:2002:DMH

[PR02] Angelika C. Paul and Nadia Rosenthal. Different modes of hypertrophy in skeletal muscle fibers. Journal of Cell Biology, 156(4):751–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/4/751.

Pidoux:2003:S

[PRA03] Alison L. Pidoux, William Richardson, and Robin C. Allshire. Sim4. Journal of Cell Biology, 161(2):295–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
2/295.

Pin:2001:BTF

[PRJK01] Christopher L. Pin, J. Michael Rukstalis, Charis Johnson, and Stephen F. Konieczny. The bHLH transcription factor Mist1 is required to maintain exocrine pancreas cell organization and acinar cell identity. Journal of Cell Biology, 155(4): 519–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/4/519.

Peden:2002:AFA

[PRLR02] Andrew A. Peden, Rachel E. Rudge, Winnie W. Y. Lui, and Margaret S. Robinson. Assembly and function of AP-3 complexes in cells expressing mutant subunits. Journal of Cell Biology, 156(2):327–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/2/327.

Prockop:2003:FPP

[Pro03] Darwin J. Prockop. Further proof of the plasticity of adult stem cells and their role in tissue repair. Journal of Cell Biology, 160(6):807–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/6/807.

Powers:2004:LKP

[PRS^{+}04] James Powers, Debra J. Rose, Adam Saunders, Steven Dunkelbarger, Susan Strome, and William M. Saxton. Loss of KLP-19 polar ejection force causes misorientation and missegregation of holocentric chromosomes. Journal of Cell Biology, 166(7):991–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/7/991.

Porter:2000:AAM

[PS00] Mary E. Porter and Winfield S. Sale. The 9 + 2 Axoneme Anchors Multiple Inner Arm Dyneins and a Network of Kinases and Phosphatases That Control Motility. Journal of Cell Biology, 151(5):??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/5/F37.

Pepper:2003:LE

[PS03] Michael S. Pepper and Mihaela Skobe. Lymphatic endothelium. Journal of Cell Biology, 163(2):209–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/163/2/
209.

Prokopenko:2000:UGK

[PSB00] Sergei N. Prokopenko, Robert Saint, and Hugo J. Bellen. Untying the Gordian Knot of Cytokinesis. Journal of Cell Biology, 148(5):843–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/5/843.

Palumbo:2004:EHS

[PSD^{+}04a] Roberta Palumbo, Maurilio Sampaolesi, Francesco De Marchis, Rossana Tonlorenzi, Sara Colombetti, Anna Mondino, Giulio Cossu, and Marco E. Bianchi. Extracellular HMGB1, a signal of tissue damage, induces mesoangioblast migration and proliferation. Journal of Cell Biology, 164(3): 441–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/3/441.

Preisinger:2004:YAG

[PSD^{+}04b] Christian Preisinger, Benjamin Short, Veerle De Corte, Erik Bruyneel, Alexander Haas, Robert Kopajtich, Jan Gettemans, and Francis A. Barr. YSK1 is activated by the Golgi matrix protein GM130 and plays a role in cell migration through its substrate 14-3-3ζ. Journal of Cell Biology, 164 (7):1009–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/7/1009.

Poliak:2003:JCS

[PSE^{+}03] Sebastian Poliak, Daniela Salomon, Hadas Elhanany, Helena Sabanay, Brent Kiernan, Larysa Pevny, Colin L. Stewart, Xiaorong Xu, Shing-Yan Chiu, Peter Shrager, Andrew J. W. Furley, and Elior Peles. Juxtaparanodal clustering of Shakerlike K^{+} channels in myelinated axons depends on Caspr2 and TAG-1. Journal of Cell Biology, 162(6):1149–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
6/1149.

Potts:2003:CFE

[PSK^{+}03] Patrick Ryan Potts, Shweta Singh, Malia Knezek, Craig B. Thompson, and Mohanish Deshmukh. Critical function of endogenous XIAP in regulating caspase activation during sympathetic neuronal apoptosis. Journal of Cell Biology, 163(4): 789–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/4/789.

Paasinen-Sohns:2000:CJA

[PSKK^{+}00] Aino Paasinen-Sohns, Mari Kielosto, Essi Kääriäinen, Terho Eloranta, Aire Laine, Olli A. Jänne, Michael J. Birrer, and Erkki Hölttä. C-jun Activation–Dependent Tumorigenic Transformation Induced Paradoxically by Overexpression or Block of S -Adenosylmethionine Decarboxylase. Journal of Cell Biology, 151(4):801–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/4/801.

Paladino:2004:POM

[PSP^{+}04] Simona Paladino, Daniela Sarnataro, Rudolf Pillich, Simona Tivodar, Lucio Nitsch, and Chiara Zurzolo. Protein oligomerization modulates raft partitioning and apical sorting of GPIanchored proteins. Journal of Cell Biology, 167(4):699– ??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/4/699.

Pece:2004:LNR

[PSS^{+}04] Salvatore Pece, Michela Serresi, Elisa Santolini, Maria Capra, Esther Hulleman, Viviana Galimberti, Stefano Zurrida, Patrick Maisonneuve, Giuseppe Viale, and Pier Paolo Di Fiore. Loss of negative regulation by Numb over Notch is relevant to human breast carcinogenesis. Journal of Cell Biology, 167(2):215–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/2/215.

Pot:2002:NES

[PSW^{+}02] Caroline Pot, Marjo Simonen, Oliver Weinmann, Lisa Schnell, Franziska Christ, Sascha Stoeckle, Philipp Berger, Thomas Rülicke, Ueli Suter, and Martin E. Schwab. Nogo-A expressed in Schwann cells impairs axonal regeneration after peripheral nerve injury. Journal of Cell Biology, 159(1): 29–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/1/29.

Price:2000:DSY

[PSWU00] Albert Price, Darren Seals, William Wickner, and Christian Ungermann. The docking stage of yeast vacuole fusion requires the transfer of proteins from a cis-snare complex to a Rab/Ypt protein. Journal of Cell Biology, 148(6): 1231–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/6/1231.

Plath:2004:DRA

[PTH^{+}04] Kathrin Plath, Dale Talbot, Karien M. Hamer, Arie P. Otte, Thomas P. Yang, Rudolf Jaenisch, and Barbara Panning. Developmentally regulated alterations in Polycomb repressive complex 1 proteins on the inactive X chromosome. Journal of Cell Biology, 167(6):1025–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/6/1025.

Persad:2001:TSP

[PTM^{+}01] Sujata Persad, Armelle A. Troussard, Timothy R. McPhee, David J. Mulholland, and Shoukat Dedhar. Tumor suppressor Pten inhibits nuclear accumulation of β-catenin and T cell/lymphoid enhancer factor 1-mediated transcriptional activation. Journal of Cell Biology, 153(6):1161–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/153/6/
1161.

Peterziel:2002:TIG

[PUK02] H. Peterziel, K. Unsicker, and K. Krieglstein. TGFβ induces GDNF responsiveness in neurons by recruitment of GFRα1 to the plasma membrane. Journal of Cell Biology, 159(1):

157–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/1/157.

Poperechnaya:2000:LAM

[PVjL^{+}00] Angela Poperechnaya, Olga Varlamova, Pei ju Lin, James T. Stull, and Anne R. Bresnick. Localization and Activity of Myosin Light Chain Kinase Isoforms during the Cell Cycle. Journal of Cell Biology, 151(3):697–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/3/
697.

Pedrazzini:2000:MRC

[PVL^{+}00] Emanuela Pedrazzini, Antonello Villa, Renato Longhi, Alessandra Bulbarelli, and Nica Borgese. Mechanism of Residence of Cytochrome B(5), a Tail–Anchored Protein, in the Endoplasmic Reticulum. Journal of Cell Biology, 148(5): 899–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/5/899.

Poon:2001:IMN

[PWC^{+}01] Betty Y. Poon, Christopher A. Ward, Conan B. Cooper, Wayne R. Giles, Alan R. Burns, and Paul Kubes. α_{4} -Integrin Mediates Neutrophil–Induced Free Radical Injury to Cardiac Myocytes. Journal of Cell Biology, 152(5):857–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
5/857.

Puri:2001:CDI

[PWS^{+}01] Vishwajeet Puri, Rikio Watanabe, Raman Deep Singh, Michel Dominguez, Jennifer C. Brown, Christine L. Wheatley, David L. Marks, and Richard E. Pagano. Clathrindependent and -independent internalization of plasma membrane sphingolipids initiates two Golgi targeting pathways. Journal of Cell Biology, 154(3):535–??, August 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/154/3/

Price:2000:PNV

[PWU00] Albert Price, William Wickner, and Christian Ungermann. Proteins Needed for Vesicle Budding from the Golgi Complex Are Also Required for the Docking Step of Homotypic Vacuole Fusion. Journal of Cell Biology, 148(6):1223–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
6/1223.

Padron:2003:PPK

[PWY^{+}03] David Padrón, Ying Jie Wang, Masaya Yamamoto, Helen Yin, and Michael G. Roth. Phosphatidylinositol phosphate 5-kinase Iβ recruits AP-2 to the plasma membrane and regulates rates of constitutive endocytosis. Journal of Cell Biology, 162(4):693–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/4/693.

Peifer:2003:TC

[PY03] Mark Peifer and Alpha S. Yap. Traffic control. Journal of Cell Biology, 163(3):437–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/3/437.

Percipalle:2001:ABH

[PZP^{+}01] Piergiorgio Percipalle, Jian Zhao, Brian Pope, Alan Weeds, Uno Lindberg, and Bertil Daneholt. Actin Bound to the Heterogeneous Nuclear Ribonucleoprotein Hrp36 Is Associated with Balbiani Ring mRNA from the Gene to Polysomes. Journal of Cell Biology, 153(1):229–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/1/229.

Qin:2004:C

[QDG^{+}04a] Hongmin Qin, Dennis R. Diener, Stefan Geimer, Douglas G. Cole, and Joel L. Rosenbaum. Correction. Journal of Cell Biology, 164(6):943–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/6/943.

Qin:2004:ITI

[QDG^{+}04b] Hongmin Qin, Dennis R. Diener, Stefan Geimer, Douglas G. Cole, and Joel L. Rosenbaum. Intraflagellar transport (IFT) cargo. Journal of Cell Biology, 164(2):255–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/2/
255.

Qualmann:2000:SIP

[QK00] Britta Qualmann and Regis B. Kelly. Syndapin isoforms participate in receptor-mediated endocytosis and actin organization. Journal of Cell Biology, 148(5):1047–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
5/1047.

Qualmann:2000:MLB

[QKK00] Britta Qualmann, Michael M. Kessels, and Regis B. Kelly. Molecular Links between Endocytosis and the Actin Cytoskeleton. Journal of Cell Biology, 150(5):0–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
5/F111.

Qu-Petersen:2002:INP

[QPDJ^{+}02] Zhuqing Qu-Petersen, Bridget Deasy, Ron Jankowski, Makato Ikezawa, James Cummins, Ryan Pruchnic, John Mytinger, Baohong Cao, Charley Gates, Anton Wernig, and Johnny Huard. Identification of a novel population of muscle stem cells in mice. Journal of Cell Biology, 157(5):851–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
157/5/851.

Quist:2000:PRG

[QRLL00] Arjan Pieter Quist, Seung Keun Rhee, Hai Lin, and Ratneshwar Lal. Physiological Role of Gap–Junctional Hemichannels. Journal of Cell Biology, 148(5):1063–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/5/
1063.

Quintyne:2002:DCC

[QS02] Nicholas J. Quintyne and Trina A. Schroer. Distinct cell cycle-dependent roles for dynactin and dynein at centrosomes. Journal of Cell Biology, 159(2):245–??, October 2002.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/2/
245.

Quaranta:2000:CMT

[Qua00] Vito Quaranta. Cell Migration through Extracellular Matrix. Journal of Cell Biology, 149(6):1167–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/6/1167.

Zhou:2001:GCC

[qZC01] Feng quan Zhou and Christopher S. Cohan. Growth Cone Collapse through Coincident Loss of Actin Bundles and Leading Edge Actin without Actin Depolymerization. Journal of Cell Biology, 153(5):1071–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/5/1071.

Rein:2002:AGM

[RAD^{+}02] Ulrike Rein, Uwe Andag, Rainer Duden, Hans Dieter Schmitt, and Anne Spang. ARF-GAP-mediated interaction between the ER–Golgi v–SNAREs and the COPI coat. Journal of Cell Biology, 157(3):395–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/3/395.

Rapraeger:2000:SRR

[Rap00] Alan C. Rapraeger. Syndecan-regulated Receptor Signaling. Journal of Cell Biology, 149(5):995–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/5/995.

Rout:2000:YNP

[RAS^{+}00] Michael P. Rout, John D. Aitchison, Adisetyantari Suprapto, Kelly Hjertaas, Yingming Zhao, and Brian T. Chait. The Yeast Nuclear Pore Complex. Journal of Cell Biology, 148(4): 635–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/4/635.

Rolls:2003:DAR

[RAS^{+}03] Melissa M. Rolls, Roger Albertson, Hsin-Pei Shih, Cheng-Yu Lee, and Chris Q. Doe. Drosophila aPKC regulates cell polarity and cell proliferation in neuroblasts and epithelia. Journal of Cell Biology, 163(5):1089–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/5/1089.

Reczek:2001:IET

[RB01] David Reczek and Anthony Bretscher. Identification of Epi64, a Tbc/Rabgap domain–containing microvillar protein that binds to the first PDZ domain of Ebp50 and E3karp. Journal of Cell Biology, 153(1):191–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/1/191.

Roper:2003:MEI

[RB03] Katja Röper and Nicholas H. Brown. Maintaining epithelial integrity. Journal of Cell Biology, 162(7):1305–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
7/1305.

Ruf:2000:SCE

[RBB00] Stephanie Ruf, Klaus Biehler, and Ralph Bock. A Small Chloroplast–Encoded Protein as a Novel Architectural Component of the Light–Harvesting Antenna. Journal of Cell Biology, 149(2):369–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/2/369.

Roll:2000:PPI

[RBBA00] Richard L. Roll, Eve Marie Bauman, Joel S. Bennett, and Charles S. Abrams. Phosphorylated Pleckstrin Induces Cell Spreading via an Integrin–Dependent Pathway. Journal of Cell Biology, 150(6):1461–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/6/1461.

Robertson:2001:ADN

[RBD^{+}01] Janice Robertson, Jean-Martin Beaulieu, Mohammad M. Doroudchi, Heather D. Durham, Jean-Pierre Julien, and Walter E. Mushynski. Apoptotic death of neurons exhibiting peripherin aggregates is mediated by the proinflammatory cytokine tumor necrosis factor-α. Journal of Cell Biology, 155(2):217–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/2/217.

Read:2002:NAI

[RBE^{+}02] Stuart H. Read, Belinda C. Baliga, Paul G. Ekert, David L. Vaux, and Sharad Kumar. A novel Apaf-1–independent putative caspase-2 activation complex. Journal of Cell Biology, 159(5):739–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/5/739.

Raghavan:2000:CAI

[RBM^{+}00] Srikala Raghavan, Christoph Bauer, Gina Mundschau, Qingqin Li, and Elaine Fuchs. Conditional Ablation of β1 Integrin in Skin. Journal of Cell Biology, 150(5):1149– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/5/1149.

Reedy:2000:FET

[RBV00] Mary C. Reedy, Belinda Bullard, and Jim O. Vigoreaux. Flightin Is Essential for Thick Filament Assembly and Sarcomere Stability in Drosophila Flight Muscles. Journal of Cell Biology, 151(7):1483–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/7/1483.

Rogers:2002:AKA

[RBW^{+}02] Eric Rogers, John D. Bishop, James A. Waddle, Jill M. Schumacher, and Rueyling Lin. The aurora kinase AIR-2 functions in the release of chromosome cohesion in Caenorhabditis elegans meiosis. Journal of Cell Biology, 157(2):219–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/157/
2/219.

Rogers:2000:KRP

[RCL^{+}00] Gregory C. Rogers, Kitty K. Chui, Edwin W. Lee, Karen P. Wedaman, David J. Sharp, Gina Holland, Robert L. Morris, and Jonathan M. Scholey. A Kinesin–Related Protein, Krp_{180}, Positions Prometaphase Spindle Poles during Early Sea Urchin Embryonic Cell Division. Journal of Cell Biology, 150(3):499–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/3/499.

Ricotta:2002:PAS

[RCS^{+}02] Doris Ricotta, Sean D. Conner, Sandra L. Schmid, Kurt von Figura, and Stefan Höning. Phosphorylation of the AP2 μ subunit by AAK1 mediates high affinity binding to membrane protein sorting signals. Journal of Cell Biology, 156 (5):791–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/5/791.

Rao:2003:NMM

[RCY^{+}03] Mala V. Rao, Jabbar Campbell, Aidong Yuan, Asok Kumar, Takahiro Gotow, Yasuo Uchiyama, and Ralph A. Nixon. The neurofilament middle molecular mass subunit carboxylterminal tail domains is essential for the radial growth and cytoskeletal architecture of axons but not for regulating neurofilament transport rate. Journal of Cell Biology, 163(5): 1021–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/5/1021.

Richards:2004:PKC

[RDC^{+}04] Theresa S. Richards, Clarence A. Dunn, William G. Carter, Marcia L. Usui, John E. Olerud, and Paul D. Lampe. Protein kinase C spatially and temporally regulates gap junctional communication during human wound repair via phosphorylation of connexin43 on serine368. Journal of Cell Biology, 167(3):555–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/555.

Ren:2001:RCG

[RDH^{+}01] Xiu-Rong Ren, Quan-Sheng Du, Yang-Zhong Huang, Shi-Zhou Ao, Lin Mei, and Wen-Cheng Xiong. Regulation of Cdc42 Gtpase by proline-rich tyrosine kinase 2 interacting with Psgap, a novel pleckstrin homology and Src homology 3 domain containing rhogap protein. Journal of Cell Biology, 152(5):971–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/5/971.

Robertson:2003:NPS

[RDN^{+}03] Janice Robertson, Mohammad M. Doroudchi, Minh Dang Nguyen, Heather D. Durham, Michael J. Strong, Gerry Shaw, Jean-Pierre Julien, and Walter E. Mushynski. A neurotoxic peripherin splice variant in a mouse model of ALS. Journal of Cell Biology, 160(6):939–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/6/939.

Ryten:2002:ARD

[RDNB02] Mina Ryten, Philip M. Dunn, Joseph T. Neary, and Geoffrey Burnstock. ATP regulates the differentiation of mammalian skeletal muscle by activation of a P2X_{5} receptor on satellite cells. Journal of Cell Biology, 158(2):345–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
2/345.

Rehm:2003:RTS

[RDP03] Markus Rehm, Heiko Düßmann, and Jochen H. M. Prehn. Real-time single cell analysis of Smac/DIABLO release during apoptosis. Journal of Cell Biology, 162(6):1031–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/162/6/1031.

Rutledge:2002:CCS

[RDS02] Eric Rutledge, Jerod Denton, and Kevin Strange. Cell cycle– and swelling-induced activation of a Caenorhabditis elegans ClC channel is mediated by CeGLC-7α /β phosphatases. Journal of Cell Biology, 158(3):435–??, August 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/3/
435.

Rangarajan:2003:CAI

[REK^{+}03] Savithri Rangarajan, Jorrit M. Enserink, H. Bea Kuiperij, Johan de Rooij, Leo S. Price, Frank Schwede, and Johannes L. Bos. Cyclic AMP induces integrin-mediated cell adhesion through Epac and Rap1 upon stimulation of the β_{2}-adrenergic receptor. Journal of Cell Biology, 160(4):487– ??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/4/487.

Rao:2002:MVB

[REM^{+}02] Mala V. Rao, Linda J. Engle, Panaiyur S. Mohan, Aidong Yuan, Dike Qiu, Anne Cataldo, Linda Hassinger, Stephen Jacobsen, Virginia M-Y. Lee, Athena Andreadis, Jean-Pierre Julien, Paul C. Bridgman, and Ralph A. Nixon. Myosin Va binding to neurofilaments is essential for correct myosin Va distribution and transport and neurofilament density. Journal of Cell Biology, 159(2):279–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/2/279.

Roth:2002:YDC

[RFCD02] Amy F. Roth, Ying Feng, Linyi Chen, and Nicholas G. Davis. The yeast DHHC cysteine-rich domain protein Akr1p is a palmitoyl transferase. Journal of Cell Biology, 159(1): 23–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/1/23.

Rovner:2002:CTI

[RFLT02] Arthur S. Rovner, Patricia M. Fagnant, Susan Lowey, and Kathleen M. Trybus. The carboxyl-terminal isoforms of smooth muscle myosin heavy chain determine thick filament assembly properties. Journal of Cell Biology, 156(1): 113–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/1/113.

Rouach:2000:ADN

[RGG00] Nathalie Rouach, Jacques Glowinski, and Christian Giaume. Activity-dependent Neuronal Control of Gap–Junctional Communication in Astrocytes. Journal of Cell Biology, 149 (7):1513–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/7/1513.

Ricci:2003:CML

[RGG03] Jean-Ehrland Ricci, Roberta A. Gottlieb, and Douglas R. Green. Caspase-mediated loss of mitochondrial function and generation of reactive oxygen species during apoptosis. Journal of Cell Biology, 160(1):65–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/1/65.

Reuveni:2000:RRI

[RGGL00] Hadas Reuveni, Tamar Geiger, Benjamin Geiger, and Alexander Levitzki. Reversal of the Ras-induced transformed phenotype by Hr12, a novel Ras farnesylation inhibitor, is mediated by the Mek/ERK pathway. Journal of Cell Biology, 151(6):1179–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/6/1179.

Romano:2003:C

[RGK^{+}03] Alper Romano, Annika Guse, Ivica Krascenicova, Heinke Schnabel, Ralf Schnabel, and Michael Glotzer. Csc-1. Journal of Cell Biology, 161(2):229–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/2/229.

Rao:2002:GRM

[RGM^{+}02] Mala V. Rao, Michael L. Garcia, Yukio Miyazaki, Takahiro Gotow, Aidong Yuan, Salvatore Mattina, Chris M. Ward, Nigel A. Calcutt, Yasuo Uchiyama, Ralph A. Nixon, and Don W. Cleveland. Gene replacement in mice reveals that the heavily phosphorylated tail of neurofilament heavy subunit does not affect axonal caliber or the transit of cargoes in slow axonal transport. Journal of Cell Biology, 158(4): 681–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/4/681.

Reese:2000:DDK

[RH00] Eric L. Reese and Leah T. Haimo. Dynein, dynactin, and kinesin II’s interaction with microtubules is regulated during bidirectional organelle transport. Journal of Cell Biology, 151(1):155–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/1/155.

Ron:2004:MBU

[RH04] David Ron and Randolph Y. Hampton. Membrane biogenesis and the unfolded protein response. Journal of Cell Biology, 167(1):23–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/1/23.

Rudner:2000:CAE

[RHM00] Adam D. Rudner, Kevin G. Hardwick, and Andrew W. Murray. Cdc28 Activates Exit from Mitosis in Budding Yeast. Journal of Cell Biology, 149(7):1361–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/7/1361.

Ridley:2000:RG

[Rid00] Anne Ridley. Rho Gtpases. Journal of Cell Biology, 150 (4):0–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/4/F107.

Rappold:2001:TSP

[RIDC01] Irene Rappold, Kuniyoshi Iwabuchi, Takayasu Date, and Junjie Chen. Tumor Suppressor P53 Binding Protein 1 (53bp1) Is Involved in DNA Damage–Signaling Pathways. Journal of Cell Biology, 153(3):613–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/3/613.

Rizzuto:2003:CMM

[Riz03] Rosario Rizzuto. Calcium mobilization from mitochondria in synaptic transmitter release. Journal of Cell Biology, 163(3): 441–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/3/441.

Rossoll:2003:SSM

[RJA^{+}03] Wilfried Rossoll, Sibylle Jablonka, Catia Andreassi, Ann-Kathrin Kröning, Kathrin Karle, Umrao R. Monani, and Michael Sendtner. Smn, the spinal muscular atrophy– determining gene product, modulates axon growth and localization of β-actin mRNA in growth cones of motoneurons.

Journal of Cell Biology, 163(4):801–??, November 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/163/4/
801.

Raff:2002:RFC

[RJyH02] Jordan W. Raff, Kim Jeffers, and Jun yong Huang. The roles of Fzy/Cdc20 and Fzr/Cdh1 in regulating the destruction of cyclin B in space and time. Journal of Cell Biology, 157 (7):1139–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/7/1139.

Runker:2004:PMM

[RKF^{+}04] Annette E. Rünker, Igor Kobsar, Torsten Fink, Gabriele Loers, Thomas Tilling, Peggy Putthoff, Carsten Wessig, Rudolf Martini, and Melitta Schachner. Pathology of a mouse mutation in peripheral myelin protein P0 is characteristic of a severe and early onset form of human Charcot–Marie–Tooth type 1B disorder. Journal of Cell Biology, 165(4):565–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
165/4/565.

Russell:2003:DFP

[RKJL03] Charles J. Russell, Karen L. Kantor, Theodore S. Jardetzky, and Robert A. Lamb. A dual-functional paramyxovirus F protein regulatory switch segment. Journal of Cell Biology, 163(2):363–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/2/363.

Rashkova:2002:GPT

[RKKP02] Svetlana Rashkova, Sarah E. Karam, Rebecca Kellum, and Mary-Lou Pardue. Gag proteins of the two Drosophila telomeric retrotransposons are targeted to chromosome ends. Journal of Cell Biology, 159(3):397–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/3/397.

Reimertz:2003:GED

[RKR^{+}03] Claus Reimertz, Donat Kögel, Abdelhaq Rami, Thomas Chittenden, and Jochen H. M. Prehn. Gene expression during ER stress–induced apoptosis in neurons. Journal of Cell Biology, 162(4):587–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/4/587.

Rice:2003:SMH

[RL03] Nancy A. Rice and Leslie A. Leinwand. Skeletal myosin heavy chain function in cultured lung myofibroblasts. Journal of Cell Biology, 163(1):119–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/1/119.

Roderick:2000:CPC

[RLC00] H. Llewelyn Roderick, James D. Lechleiter, and Patricia Camacho. Cytosolic Phosphorylation of Calnexin Controls Intracellular Ca^{2+} Oscillations via an Interaction with Serca2b. Journal of Cell Biology, 149(6):1235–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/6/1235.

Rivera:2002:BIT

[RLTC^{+}02] Claudio Rivera, Hong Li, Judith Thomas-Crusells, Hannele Lahtinen, Tero Viitanen, Avtandil Nanobashvili, Zaal Kokaia, Matti S. Airaksinen, Juha Voipio, Kai Kaila, and Mart Saarma. BDNF-induced TrkB activation downregulates the K^{+} -Cl− cotransporter KCC2 and impairs neuronal Cl− extrusion. Journal of Cell Biology, 159(5):747– ??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/5/747.

Reiken:2003:PPA

[RLZ^{+}03] Steven Reiken, Alain Lacampagne, Hua Zhou, Aftab Kherani, Stephan E. Lehnart, Chris Ward, Fannie Huang, Marta Gaburjakova, Jana Gaburjakova, Nora Rosemblit, Michelle S. Warren, Kun lun He, Geng hua Yi, Jie Wang, Daniel Burkhoff, Guy Vassort, and Andrew R. Marks. PKA phosphorylation activates the calcium release channel (ryanodine receptor) in skeletal muscle. Journal of Cell Biology, 160(6):919–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/6/919.

Rudner:2000:PCA

[RM00] Adam D. Rudner and Andrew W. Murray. Phosphorylation by Cdc28 Activates the Cdc20-Dependent Activity of the Anaphase–Promoting Complex. Journal of Cell Biology, 149 (7):1377–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/7/1377.

Reilly:2001:IMN

[RM01] John F. Reilly and Pamela A. Maher. Importin β–mediated nuclear import of fibroblast growth factor receptor. Journal of Cell Biology, 152(6):1307–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/6/1307.

Robert:2002:DCD

[RMC^{+}02] Amélie Robert, Marie-Joëlle Miron, Claudia Champagne, Marie-Claude Gingras, Philip E. Branton, and Josée N. Lavoie. Distinct cell death pathways triggered by the adenovirus early region 4 ORF 4 protein. Journal of Cell Biology, 158(3):519–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/3/519.

Rasmussen:2000:DRH

[RME^{+}00] Theodore P. Rasmussen, Mary-Ann Mastrangelo, Amir Eden, John R. Pehrson, and Rudolf Jaenisch. Dynamic Relocalization of Histone Macroh2a1 from Centrosomes to Inactive X Chromosomes during X Inactivation. Journal of Cell Biology, 150(5):1189–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/5/1189.

Rogalski:2000:UGC

[RMG^{+}00] Teresa M. Rogalski, Gregory P. Mullen, Mary M. Gilbert, Benjamin D. Williams, and Donald G. Moerman. The UNC-112 Gene in Caenorhabditis elegans Encodes a Novel Component of Cell–Matrix Adhesion Structures Required for Integrin Localization in the Muscle Cell Membrane. Journal of Cell Biology, 150(1):253–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/1/253.

Reinecke:2000:ECB

[RMHM00] Hans Reinecke, Glen H. MacDonald, Stephen D. Hauschka, and Charles E. Murry. Electromechanical Coupling between Skeletal and Cardiac Muscle. Journal of Cell Biology, 149 (3):731–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/3/731.

Roh:2002:MPP

[RML^{+}02] Michael H. Roh, Olga Makarova, Chia-Jen Liu, Shin, Seonok Lee, Stephanie Laurinec, Meera Goyal, Roger Wiggins, and Ben Margolis. The Maguk protein, Pals1, functions as an adapter, linking mammalian homologues of Crumbs and Discs Lost. Journal of Cell Biology, 157(1):161–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/157/
1/161.

Roberts:2001:IAP

[RMMC01] Darren L. Roberts, Wendy Merrison, Marion MacFarlane, and Gerald M. Cohen. The Inhibitor of Apoptosis Protein– Binding Domain of Smac Is Not Essential for Its Proapoptotic Activity. Journal of Cell Biology, 153(1):221–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
1/221.

Rudolf:2004:VMC

[RMMP04] Rüdiger Rudolf, Marco Mongillo, Paulo J. Magalhães, and Tullio Pozzan. In vivo monitoring of Ca^{2+} uptake into mitochondria of mouse skeletal muscle during contraction. Journal of Cell Biology, 166(4):527–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/4/527.

Russell:2002:MCM

[RMR02] David G. Russell, Henry C. Mwandumba, and Elizabeth E. Rhoades. Mycobacterium and the coat of many lipids. Journal of Cell Biology, 158(3):421–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/3/421.

Ryan:2003:RGC

[RMW03] Kathryn J. Ryan, J. Michael McCaffery, and Susan R. Wente. The Ran GTPase cycle is required for yeast nuclear pore complex assembly. Journal of Cell Biology, 160 (7):1041–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/7/1041.

Rossner:2004:JWL

[RO04] Mike Rossner and Rob O’Donnell. The JCB will let your data shine in RGB. Journal of Cell Biology, 164(1):11–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
1/11.

Rossner:2002:FM

[Ros02] Mike Rossner. Figure manipulation. Journal of Cell Biology, 158(7):1151–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/7/1151.

Rossner:2003:J

[Ros03] Mike Rossner. The JCB 2003. Journal of Cell Biology, 161 (5):837–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/5/837.

Rizzo:2003:RCG

[RP03a] Mark A. Rizzo and David W. Piston. Regulation of β cell glucokinase by S-nitrosylation and association with nitric oxide synthase. Journal of Cell Biology, 161(2):243–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
2/243.

Rupp:2003:SDR

[RP03b] Gerald Rupp and Mary E. Porter. A subunit of the dynein regulatory complex in Chlamydomonas is a homologue of a growth arrest–specific gene product. Journal of Cell Biology, 162(1):47–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/1/47.

Rybakova:2000:DCF

[RPE00] Inna N. Rybakova, Jitandrakumar R. Patel, and James M. Ervasti. The Dystrophin Complex Forms a Mechanically Strong Link between the Sarcolemma and Costameric Actin. Journal of Cell Biology, 150(5):1209–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/5/
1209.

Rajan:2003:BSA

[RPNM03] Prithi Rajan, David M. Panchision, Laura F. Newell, and Ronald D. G. McKay. BMPs signal alternately through a SMAD or FRAP–STAT pathway to regulate fate choice in CNS stem cells. Journal of Cell Biology, 161(5):911–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
5/911.

Rapizzi:2002:REV

[RPS^{+}02] Elena Rapizzi, Paolo Pinton, György Szabadkai, Mariusz R. Wieckowski, Grégoire Vandecasteele, Geoff Baird, Richard A. Tuft, Kevin E. Fogarty, and Rosario Rizzuto. Recombinant expression of the voltage-dependent anion channel enhances the transfer of Ca^{2+} microdomains to mitochondria. Journal of Cell Biology, 159(4):613–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/4/613.

Reck-Peterson:2001:YCV

[RPTNM01] Samara L. Reck-Peterson, Matthew J. Tyska, Peter J. Novick, and Mark S. Mooseker. The Yeast Class V Myosins, Myo2p and Myo4p, Are Nonprocessive Actin–Based Motors. Journal of Cell Biology, 153(5):1121–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/5/1121.

Rasband:2002:CNP

[RPZ^{+}02] Matthew N. Rasband, Eunice W. Park, Dongkai Zhen, Margaret I. Arbuckle, Sebastian Poliak, Elior Peles, Seth G. N. Grant, and James S. Trimmer. Clustering of neuronal potassium channels is independent of their interaction with PSD-95. Journal of Cell Biology, 159(4):663–??, November 2002.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/4/
663.

Rossanese:2001:RAC

[RRB^{+}01] Olivia W. Rossanese, Catherine A. Reinke, Brooke J. Bevis, Adam T. Hammond, Irina B. Sears, James O’Connor, and Benjamin S. Glick. A Role for Actin, Cdc1p, and Myo2p in the Inheritance of Late Golgi Elements in Saccharomyces cerevisiae. Journal of Cell Biology, 153(1):47–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/153/1/
47.

Raemaekers:2003:NNM

[RRB^{+}03] Tim Raemaekers, Katharina Ribbeck, Joël Beaudouin, Wim Annaert, Mark Van Camp, Ingrid Stockmans, Nico Smets, Roger Bouillon, Jan Ellenberg, and Geert Carmeliet. NuSAP, a novel microtubule-associated protein involved in mitotic spindle organization. Journal of Cell Biology, 162(6):1017– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/6/1017.

Roy:2001:LPC

[RRJ^{+}01] Partha Roy, Zenon Rajfur, David Jones, Gerard Marriott, Leslie Loew, and Ken Jacobson. Local Photorelease of Caged Thymosin β4 in Locomoting Keratocytes Causes Cell Turning. Journal of Cell Biology, 153(5):1035–??, May 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/153/5/
1035.

Reedquist:2000:SGR

[RRK^{+}00] Kris A. Reedquist, Ewan Ross, Elianne A. Koop, Rob M. F. Wolthuis, Fried J. T. Zwartkruis, Yvette van Kooyk, Mike Salmon, Christopher D. Buckley, and Johannes L. Bos. The Small Gtpase, Rap1, Mediates Cd31-Induced Integrin Adhesion. Journal of Cell Biology, 148(6):1151–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/6/

Rosen:2000:ARP

[RRL^{+}00] Kirill Rosen, Janusz Rak, Thomas Leung, Nicholas M. Dean, Robert S. Kerbel, and Jorge Filmus. Activated ras Prevents Downregulation of Bcl–X_{L} Triggered by Detachment from the Extracellular Matrix. Journal of Cell Biology, 149 (2):447–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/2/447.

Richard:2000:LCP

[RRM^{+}00] Isabelle Richard, Carinne Roudaut, Sylvie Marchand, Stephen Baghdiguian, Muriel Herasse, Daniel Stockholm, Yasuko Ono, Laurence Suel, Nathalie Bourg, Hiroyuki Sorimachi, Gérard Lefranc, Michel Fardeau, Alain Sébille, and Jacques S. Beckmann. Loss of Calpain 3 proteolytic activity leads to muscular dystrophy and to apoptosis-associated Iκ bα/nuclear factor κ b pathway perturbation in mice. Journal of Cell Biology, 151(7):1583–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/7/1583.

Riggs:2003:ACR

[RRM^{+}03] Blake Riggs, Wendy Rothwell, Sarah Mische, Gilles R. X. Hickson, Johanne Matheson, Thomas S. Hays, Gwyn W. Gould, and William Sullivan. Actin cytoskeleton remodeling during early Drosophila furrow formation requires recycling endosomal components Nuclear-fallout and Rab11. Journal of Cell Biology, 163(1):143–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/1/143.

Reiser:2003:YOS

[RRS03] Vladimír Reiser, Desmond C. Raitt, and Haruo Saito. Yeast osmosensor Sln1 and plant cytokinin receptor Cre1 respond to changes in turgor pressure. Journal of Cell Biology, 161 (6):1035–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/6/1035.

Rogers:2002:DEI

[RRSV02] Stephen L. Rogers, Gregory C. Rogers, David J. Sharp, and Ronald D. Vale. Drosophila EB1 is important for proper assembly, dynamics, and positioning of the mitotic spindle. Journal of Cell Biology, 158(5):873–??, September 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/5/
873.

Raucher:2000:CSL

[RS00a] Drazen Raucher and Michael P. Sheetz. Cell Spreading and Lamellipodial Extension Rate Is Regulated by Membrane Tension. Journal of Cell Biology, 148(1):127–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
1/127.

Roberts:2000:ALA

[RS00b] Thomas M. Roberts and Murray Stewart. Acting like Actin. Journal of Cell Biology, 149(1):7–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/1/7.

Robinson:2000:DGL

[RS00c] Douglas N. Robinson and James A. Spudich. Dynacortin, a Genetic Link between Equatorial Contractility and Global Shape Control Discovered by Library Complementation of a Dictyostelium discoideum Cytokinesis Mutant. Journal of Cell Biology, 150(4):823–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/4/823.

Redick:2000:DFC

[RSBE00] Sambra D. Redick, Daniel L. Settles, Gina Briscoe, and Harold P. Erickson. Defining fibronectin’s cell adhesion synergy site by site-directed mutagenesis. Journal of Cell Biology, 149(2):521–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/2/521.

Rzadzinska:2004:AMT

[RSD^{+}04] Agnieszka K. Rzadzinska, Mark E. Schneider, Caroline Davies, Gavin P. Riordan, and Bechara Kachar. An actin molecular treadmill and myosins maintain stereocilia functional architecture and self-renewal. Journal of Cell Biology, 164(6):887–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/6/887.

Raden:2000:RCS

[RSG00] David Raden, Weiqun Song, and Reid Gilmore. Role of the Cytoplasmic Segments of Sec61α in the Ribosome–Binding and Translocation–Promoting Activities of the Sec61 Complex. Journal of Cell Biology, 150(1):53–??, July 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/1/
53.

Reinacher-Schick:2001:AML

[RSG01] Anke Reinacher-Schick and Barry M. Gumbiner. Apical Membrane Localization of the Adenomatous Polyposis Coli Tumor Suppressor Protein and Subcellular Distribution of the β-Catenin Destruction Complex in Polarized Epithelial Cells. Journal of Cell Biology, 152(3):491–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/152/3/
491.

Royou:2002:CRN

[RSK02] Anne Royou, William Sullivan, and Roger Karess. Cortical recruitment of nonmuscle myosin II in early syncytial Drosophila embryos. Journal of Cell Biology, 158(1):127–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
158/1/127.

Roy:2002:RLT

[RT02] Craig R. Roy and Lewis G. Tilney. The road less traveled. Journal of Cell Biology, 158(3):415–??, August 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/3/
415.

Rusan:2002:RMA

[RTFW02] Nasser M. Rusan, U. Serdar Tulu, Carey Fagerstrom, and Patricia Wadsworth. Reorganization of the microtubule array in prophase/prometaphase requires cytoplasmic dyneindependent microtubule transport. Journal of Cell Biology, 158(6):997–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/6/997.

Raposo:2001:DPS

[RTM^{+}01] Graça Raposo, Danielle Tenza, Diane M. Murphy, Joanne F. Berson, and Michael S. Marks. Distinct protein sorting and localization to premelanosomes, melanosomes, and lysosomes in pigmented melanocytic cells. Journal of Cell Biology, 152(4):809–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/4/809.

Rutishauser:2000:DRM

[Rut00] Urs Rutishauser. Defining a Role and Mechanism for Igcam Function in Vertebrate Axon Guidance. Journal of Cell Biology, 149(4):757–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/4/757.

Reiss:2001:ICE

[RVB^{+}01] Matthias Reiss, Nicola Viebig, Susan Brecht, Marie-Noelle Fourmaux, Martine Soete, Manlio Di Cristina, Jean François Dubremetz, and Dominique Soldati. Identification and characterization of an escorter for two secretory adhesins in Toxoplasma gondii. Journal of Cell Biology, 152(3):563–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
152/3/563.

Ratcliffe:2001:SCS

[RWCC01] Charlotte F. Ratcliffe, Ruth E. Westenbroek, Rory Curtis, and William A. Catterall. Sodium channel β1 and β3 subunits associate with neurofascin through their extracellular immunoglobulin-like domain. Journal of Cell Biology, 154 (2):427–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/2/427.

Rohde:2002:PBB

[RWH02] Gundula Rohde, Dirk Wenzel, and Volker Haucke. A phosphatidylinositol (4,5)-bisphosphate binding site within μ2adaptin regulates clathrin-mediated endocytosis. Journal of Cell Biology, 158(2):209–??, July 2002. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/2/209.

Roeth:2004:HND

[RWK^{+}04] Jeremiah F. Roeth, Maya Williams, Matthew R. Kasper, Tracey M. Filzen, and Kathleen L. Collins. HIV-1 Nef disrupts MHC–I trafficking by recruiting AP-1 to the MHC– I cytoplasmic tail. Journal of Cell Biology, 167(5):903– ??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/5/903.

Rogers:2003:MRA

[RWSV03] Stephen L. Rogers, Ursula Wiedemann, Nico Stuurman, and Ronald D. Vale. Molecular requirements for actin-based lamella formation in Drosophila S2 cells. Journal of Cell Biology, 162(6):1079–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/6/1079.

Roth:2003:DND

[RXS^{+}03] Peggy Roth, Nikos Xylourgidis, Nafiseh Sabri, Anne Uv, Maarten Fornerod, and Christos Samakovlis. The Drosophila nucleoporin DNup88 localizes DNup214 and CRM1 on the nuclear envelope and attenuates NES-mediated nuclear export. Journal of Cell Biology, 163(4):701–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/163/4/
701.

Rossner:2004:WPT

[RY04] Mike Rossner and Kenneth M. Yamada. What’s in a picture? The temptation of image manipulation. Journal of Cell Biology, 166(1):11–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/1/11.

Rohatgi:2000:MWA

[RyHHK00] Rajat Rohatgi, Hsin yi Henry Ho, and Marc W. Kirschner. Mechanism of N–Wasp Activation by Cdc42 and Phosphatidylinositol 4,5-Bisphosphate. Journal of Cell Biology, 150(6):1299–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/6/1299.

Riveline:2001:FCM

[RZB^{+}01] Daniel Riveline, Eli Zamir, Nathalie Q. Balaban, Ulrich S. Schwarz, Toshimasa Ishizaki, Shuh Narumiya, Zvi Kam, Benjamin Geiger, and Alexander D. Bershadsky. Focal Contacts as Mechanosensors. Journal of Cell Biology, 153(6):1175–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/6/1175.

Ratts:2003:CED

[RZB^{+}03] Ryan Ratts, Huiyan Zeng, Eric A. Berg, Clare Blue, Mark E. McComb, Cathy E. Costello, Johanna C. vanderSpek, and John R. Murphy. The cytosolic entry of diphtheria toxin catalytic domain requires a host cell cytosolic translocation factor complex. Journal of Cell Biology, 160(7):1139–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
7/1139.

Skoble:2001:PRV

[SAG^{+}01] Justin Skoble, Victoria Auerbuch, Erin D. Goley, Matthew D. Welch, and Daniel A. Portnoy. Pivotal role of VASP in Arp2/3 complex–mediated actin nucleation, actin branchformation, and Listeria monocytogenes motility. Journal of Cell Biology, 155(1):89–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/1/89.

Stead:2003:PRM

[SAH^{+}03] Kristen Stead, Cristina Aguilar, Theresa Hartman, Melissa Drexel, Pamela Meluh, and Vincent Guacci. Pds5p regulates the maintenance of sister chromatid cohesion and is sumoylated to promote the dissolution of cohesion. Journal of Cell Biology, 163(4):729–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/4/729.

Su:2003:PCC

[SAP03] Ling Su, Joyce M. Agati, and Sarah J. Parsons. p190RhoGAP is cell cycle regulated and affects cytokinesis. Journal of Cell Biology, 163(3):571–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/3/571.

Skerrett:2002:IAA

[SAS^{+}02] I. M. Skerrett, J. Aronowitz, J. H. Shin, G. Cymes, E. Kasperek, F. L. Cao, and B. J. Nicholson. Identification of amino acid residues lining the pore of a gap junction channel. Journal of Cell Biology, 159(2):349–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/159/
2/349.

Shao:2002:PPC

[SATA^{+}02] Yufang Shao, Wendy Akmentin, Juan Jose Toledo-Aral, Julie Rosenbaum, Gregorio Valdez, John B. Cabot, Brian S. Hilbush, and Simon Halegoua. Pincher, a pinocytic chaperone for nerve growth factor/TrkA signaling endosomes. Journal of Cell Biology, 157(4):679–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/4/679.

Salmon:2002:DWF

[SAWS02] Wendy C. Salmon, Michael C. Adams, and Clare M. Waterman-Storer. Dual-wavelength fluorescent speckle microscopy reveals coupling of microtubule and actin movements in migrating cells. Journal of Cell Biology, 158(1): 31–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/1/31.

Somersan:2001:TT

[SB01] Selin Somersan and Nina Bhardwaj. Tethering and tickling. Journal of Cell Biology, 155(4):501–??, November 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/155/4/
501.

Severson:2003:MPP

[SB03] Aaron F. Severson and Bruce Bowerman. Myosin and the PAR proteins polarize microfilament-dependent forces that shape and position mitotic spindles in Caenorhabditis elegans. Journal of Cell Biology, 161(1):21–??, April 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/1/
21.

Sekelsky:2000:DRD

[SBB00] Jeff J. Sekelsky, Michael H. Brodsky, and Kenneth C. Burtis. DNA Repair in Drosophila. Journal of Cell Biology, 150(2):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/2/F31.

Sharp-Baker:2001:SCP

[SBC01] Hilary Sharp-Baker and Rey-Huei Chen. Spindle Checkpoint Protein Bub1 Is Required for Kinetochore Localization of Mad1, Mad2, Bub3, and Cenp–E, Independently of Its Kinase Activity. Journal of Cell Biology, 153(6):1239–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
6/1239.

Svitkina:2003:MFI

[SBC^{+}03] Tatyana M. Svitkina, Elena A. Bulanova, Oleg Y. Chaga, Danijela M. Vignjevic, Shin ichiro Kojima, Jury M. Vasiliev, and Gary G. Borisy. Mechanism of filopodia initiation by reorganization of a dendritic network. Journal of Cell Biology, 160(3):409–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/3/409.

Seveau:2004:RLR

[SBG^{+}04] Stéphanie Seveau, Hélène Bierne, Stéphanie Giroux, Marie-Christine Prévost, and Pascale Cossart. Role of lipid rafts in E-cadherin- and HGF-R/Met-mediated entry of Listeria monocytogenes into host cells. Journal of Cell Biology, 166 (5):743–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/5/743.

Strasser:2000:BMM

[SBH00] Katja Sträßer, Jochen Baßler, and Ed Hurt. Binding of the Mex67p/Mtr2p heterodimer to Fxfg, Glfg, and Fg repeat nucleoporins is essential for nuclear mRNA export. Journal of Cell Biology, 150(4):695–??, August 2000. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/4/695.

Smith:2000:EIC

[SBI^{+}00] Robert M. Smith, Boris Baibakov, Yoshihide Ikebuchi, Benjamin H. White, Nevin A. Lambert, Leonard K. Kaczmarek, and Steven S. Vogel. Exocytotic Insertion of Calcium Channels Constrains Compensatory Endocytosis to Sites of Exocytosis. Journal of Cell Biology, 148(4):755–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
4/755.

Stinchcombe:2001:RRR

[SBM^{+}01] Jane C. Stinchcombe, Duarte C. Barral, Emilie H. Mules, Sarah Booth, Alistair N. Hume, Laura M. Machesky, Miguel C. Seabra, and Gillian M. Griffiths. Rab27a Is Required for Regulated Secretion in Cytotoxic T Lymphocytes. Journal of Cell Biology, 152(4):825–??, February 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/152/4/
825.

Salinas:2004:SRN

[SBMB^{+}04] Sara Salinas, Anne Briançon-Marjollet, Guillaume Bossis, Marie-Aude Lopez, Marc Piechaczyk, Isabelle Jariel-Encontre, Anne Debant, and Robert A. Hipskind. SUMOylation regulates nucleo-cytoplasmic shuttling of Elk-1. Journal of Cell Biology, 165(6):767–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/6/767.

Sacco-Bubulya:2002:DIG

[SBS02a] Paula Sacco-Bubulya and David L. Spector. Disassembly of interchromatin granule clusters alters the coordination of transcription and pre-mRNA splicing. Journal of Cell Biology, 156(3):425–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/3/425.

Shorter:2002:STG

[SBS^{+}02b] James Shorter, Matthew B. Beard, Joachim Seemann, A. Barbara Dirac-Svejstrup, and Graham Warren. Sequential tethering of Golgins and catalysis of SNAREpin assembly by the vesicle-tethering protein p115. Journal of Cell Biology, 157(1):45–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/1/45.

Sakisaka:2004:RST

[SBT^{+}04] Toshiaki Sakisaka, Takeshi Baba, Shintaro Tanaka, Genkichi Izumi, Masato Yasumi, and Yoshimi Takai. Regulation of SNAREs by tomosyn and ROCK. Journal of Cell Biology, 166(1):17–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/1/17.

Sachs:2000:ERG

[SBZ^{+}00] Martin Sachs, Henning Brohmann, Dietmar Zechner, Thomas Müller, Jörg Hülsken, Ingrid Walther, Ute Schaeper, Carmen Birchmeier, and Walter Birchmeier. Essential Role of Gab1 for Signaling by the C–Met Receptor in Vivo. Journal of Cell Biology, 150(6):1375–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/6/1375.

Santy:2001:AAA

[SC01a] Lorraine C. Santy and James E. Casanova. Activation of ARF6 by ARNO stimulates epithelial cell migration through downstream activation of both Rac1 and phospholipase D. Journal of Cell Biology, 154(3):599–??, August 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/154/3/
599.

Steen:2001:MBT

[SC01b] Rikke L. Steen and Philippe Collas. Mistargeting of B–Type Lamins at the End of Mitosis. Journal of Cell Biology, 153 (3):621–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/3/621.

Strachan:2004:CNC

[SC04] Lauren R. Strachan and Maureen L. Condic. Cranial neural crest recycle surface integrins in a substratum-dependent manner to promote rapid motility. Journal of Cell Biology, 167(3):545–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/545.

Schott:2002:SVT

[SCB02] Daniel H. Schott, Ruth N. Collins, and Anthony Bretscher. Secretory vesicle transport velocity in living cells depends on the myosin–V lever arm length. Journal of Cell Biology, 156 (1):35–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/1/35.

Skillington:2002:BMP

[SCD02] Jeremy Skillington, Lisa Choy, and Rik Derynck. Bone morphogenetic protein and retinoic acid signaling cooperate to induce osteoblast differentiation of preadipocytes. Journal of Cell Biology, 159(1):135–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/1/135.

Schaller:2000:U

[Sch00] Michael D. Schaller. Unc112. Journal of Cell Biology, 150(1):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/1/F9.

Scherer:2002:M

[Sch02] Steven S. Scherer. Myelination. Journal of Cell Biology, 156 (1):13–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/1/13.

Schaller:2004:FP

[Sch04] Michael D. Schaller. FAK and paxillin. Journal of Cell Biology, 166(2):157–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/2/157.

Solomon:2004:DTR

[SCK04] David A. Solomon, M. Cristina Cardoso, and Erik S. Knudsen. Dynamic targeting of the replication machinery to sites of DNA damage. Journal of Cell Biology, 166(4):455–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
166/4/455.

Strochlic:2001:M

[SCL^{+}01] Laure Strochlic, Annie Cartaud, Valérie Labas, Werner Hoch, Jean Rossier, and Jean Cartaud. Magi-1c. Journal of Cell Biology, 153(5):1127–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/5/1127.

Steen:2000:KAP

[SCLC00] Rikke Lise Steen, Fabien Cubizolles, Katherine Le Guellec, and Philippe Collas. A kinase-anchoring protein (Akap95) recruits human chromosome-associated protein (Hcap–D2/ Eg7) for chromosome condensation in mitotic extract. Journal of Cell Biology, 149(3):531–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/3/531.

Segal:2000:CSA

[SCM^{+}00] Marisa Segal, Duncan J. Clarke, Paul Maddox, E. D. Salmon, Kerry Bloom, and Steven I. Reed. Coordinated Spindle Assembly and Orientation Requires Clb5p–Dependent Kinase in Budding Yeast. Journal of Cell Biology, 148(3): 441–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/3/441.

Schultz:2000:PBP

[SCMH00] Linda B. Schultz, Nabil H. Chehab, Asra Malikzay, and Thanos D. Halazonetis. P53 Binding Protein 1 (53bp1) Is an Early Participant in the Cellular Response to DNA Double–Strand Breaks. Journal of Cell Biology, 151(7):1381– ??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/7/1381.

Siebrasse:2002:RNP

[SCP02] Jan Peter Siebrasse, Elias Coutavas, and Reiner Peters. Reconstitution of nuclear protein export in isolated nuclear envelopes. Journal of Cell Biology, 158(5):849–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
5/849.

Schenkman:2002:RCC

[SCPP02] Laura R. Schenkman, Carlo Caruso, Nicolas Pagé, and John R. Pringle. The role of cell cycle–regulated expression in the localization of spatial landmark proteins in yeast. Journal of Cell Biology, 156(5):829–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/5/829.

Striepen:2000:PTG

[SCS^{+}00] Boris Striepen, Michael J. Crawford, Michael K. Shaw, Lewis G. Tilney, Frank Seeber, and David S. Roos. The Plastid of Toxoplasma gondii Is Divided by Association with the Centrosomes. Journal of Cell Biology, 151(7):1423– ??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/7/1423.

Sasaki:2004:LRS

[SCTF04] Atsuo T. Sasaki, Cheryl Chun, Kosuke Takeda, and Richard A. Firtel. Localized Ras signaling at the leading edge regulates PI3K, cell polarity, and directional cell movement. Journal of Cell Biology, 167(3):505–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
3/505.

Servant:2000:DRP

[SCTM00] Marc J. Servant, Philippe Coulombe, Benjamin Turgeon, and Sylvain Meloche. Differential Regulation of P27^{Kip1} Expression by Mitogenic and Hypertrophic Factors. Journal of Cell Biology, 148(3):543–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/148/3/543.

Sudakin:2001:CIA

[SCY01] Valery Sudakin, Gordon K. T. Chan, and Tim J. Yen. Checkpoint inhibition of the APC/C in HeLa cells is mediated by a complex of BUBR1, BUB3, CDC20, and MAD2. Journal of Cell Biology, 154(5):925–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/5/925.

Schmidtke:2000:DOM

[SD00] David W. Schmidtke and Scott L. Diamond. Direct Observation of Membrane Tethers Formed during Neutrophil Attachment to Platelets or P–Selectin under Physiological Flow. Journal of Cell Biology, 149(3):719–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/3/719.

Sprong:2001:GRS

[SDC^{+}01] Hein Sprong, Sophie Degroote, Tijs Claessens, Judith van Drunen, Viola Oorschot, Ben H. C. Westerink, Yoshio Hirabayashi, Judith Klumperman, Peter van der Sluijs, and Gerrit van Meer. Glycosphingolipids are required for sorting melanosomal proteins in the Golgi complex. Journal of Cell Biology, 155(3):369–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/3/369.

Sumen:2004:CRA

[SDD04] Cenk Sumen, Michael L. Dustin, and Mark M. Davis. T cell receptor antagonism interferes with MHC clustering and integrin patterning during immunological synapse formation. Journal of Cell Biology, 166(4):579–??, August 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/166/4/
579.

Screaton:2000:SDB

[SDDS00] Robert A. Screaton, Luisa DeMarte, Petr Dráber, and Clifford P. Stanners. The Specificity for the Differentiation Blocking Activity of Carcinoembryonic Antigen Resides in Its Glycophosphatidyl–Inositol Anchor. Journal of Cell Biology, 150(3):613–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/3/613.

Slack-Davis:2003:PPM

[SDEZ^{+}03] Jill K. Slack-Davis, Scott T. Eblen, Maja Zecevic, Scott A. Boerner, Adel Tarcsafalvi, H. Bruce Diaz, Mark S. Marshall, Michael J. Weber, J. Thomas Parsons, and Andrew D. Catling. PAK1 phosphorylation of MEK1 regulates fibronectin-stimulated MAPK activation. Journal of Cell Biology, 162(2):281–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/2/281.

Suvorova:2002:SSC

[SDL02] Elena S. Suvorova, Rainer Duden, and Vladimir V. Lupashin. The Sec34/Sec35p complex, a Ypt1p effector required for retrograde intra-Golgi trafficking, interacts with Golgi SNAREs and COPI vesicle coat proteins. Journal of Cell Biology, 157 (4):631–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/4/631.

Shin:2003:HTG

[SDL^{+}03] Dong Min Shin, Marlin Dehoff, Xiang Luo, Shin Hyeok Kang, Jiangchen Tu, Surendra K. Nayak, Elliott M. Ross, Paul F. Worley, and Shmuel Muallem. Homer 2 tunes G protein– coupled receptors stimulus intensity by regulating RGS proteins and PLCβ GAP activities. Journal of Cell Biology, 162 (2):293–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/2/293.

Sotillos:2004:DDP

[SDMC^{+}04] Sol Sotillos, María Teresa Díaz-Meco, Eva Caminero, Jorge Moscat, and Sonsoles Campuzano. DaPKC-dependent phosphorylation of Crumbs is required for epithelial cell polarity in Drosophila. Journal of Cell Biology, 166(4):549–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
166/4/549.

Smits:2004:SSN

[SDML04] Patrick Smits, Peter Dy, Srijeet Mitra, and Véronique Lefebvre. Sox5 and Sox6 are needed to develop and maintain source, columnar, and hypertrophic chondrocytes in the cartilage growth plate. Journal of Cell Biology, 164(5):747–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/164/5/747.

Sonnichsen:2000:DMD

[SDN^{+}00] Birte Sönnichsen, Stefano De Renzis, Erik Nielsen, Jens Rietdorf, and Marino Zerial. Distinct membrane domains on endosomes in the recycling pathway visualized by multicolor imaging of Rab4, Rab5, and Rab11. Journal of Cell Biology, 149(4):901–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/4/901.

Sever:2000:DGC

[SDS00] Sanja Sever, Hanna Damke, and Sandra L. Schmid. Dynamin: GTP controls the formation of constricted coated pits, the rate limiting step in clathrin-mediated endocytosis. Journal of Cell Biology, 150(5):1137–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/5/
1137.

Sole:2004:DRA

[SDS^{+}04] Carme Sole, Xavier Dolcet, Miguel F. Segura, Humberto Gutierrez, Maria-Teresa Diaz-Meco, Raffaella Gozzelino, Daniel Sanchis, Jose R. Bayascas, Carme Gallego, Jorge Moscat, Alun M. Davies, and Joan X. Comella. The death receptor antagonist FAIM promotes neurite outgrowth by a mechanism that depends on ERK and NF-κB signaling. Journal of Cell Biology, 167(3):479–??, November 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/167/3/
479.

Seaman:2004:CSE

[Sea04] Matthew N. J. Seaman. Cargo-selective endosomal sorting for retrieval to the Golgi requires retromer. Journal of Cell Biology, 165(1):111–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/1/111.

Spencer:2000:RMD

[SEI^{+}00] Jeffrey A. Spencer, Susan Eliazer, Robert L. Ilaria, James A. Richardson, and Eric N. Olson. Regulation of Microtubule Dynamics and Myogenic Differentiation by Murf, a Striated Muscle Ring–Finger Protein. Journal of Cell Biology, 150 (4):771–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/4/771.

Smith:2000:WRA

[SEM^{+}00] Jesse J. Smith, Erica K. Evans, Monica Murakami, Mary B. Moyer, M. Arthur Moseley, George Vande Woude, and Sally Kornbluth. Wee1-regulated apoptosis mediated by the Crk adaptor protein in Xenopus egg extracts. Journal of Cell Biology, 151(7):1391–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/7/1391.

Sixt:2001:ECL

[SEP^{+}01] Michael Sixt, Britta Engelhardt, Friederike Pausch, Rupert Hallmann, Olaf Wendler, and Lydia M. Sorokin. Endothelial Cell Laminin Isoforms, Laminins 8 and 10, Play Decisive Roles in T Cell Recruitment across the Blood–Brain Barrier in Experimental Autoimmune Encephalomyelitis. Journal of Cell Biology, 153(5):933–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/5/933.

Salina:2003:NIN

[SERB03] Davide Salina, Paul Enarson, J. B. Rattner, and Brian Burke. Nup358 integrates nuclear envelope breakdown with kinetochore assembly. Journal of Cell Biology, 162(6):991– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/6/991.

Stockinger:2001:CRC

[SEW^{+}01] Andreas Stockinger, Andreas Eger, Julia Wolf, Hartmut Beug, and Roland Foisner. E-cadherin regulates cell growth by modulating proliferation-dependent β-catenin transcriptional activity. Journal of Cell Biology, 154(6):1185–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/154/6/1185.

Suter:2001:TGC

[SF01] Daniel M. Suter and Paul Forscher. Transmission of growth cone traction force through apCAM–cytoskeletal linkages is regulated by Src family tyrosine kinase activity. Journal of Cell Biology, 155(3):427–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/3/427.

Sasaki:2004:L

[SFH04] Takako Sasaki, Reinhard Fässler, and Erhard Hohenester. Laminin. Journal of Cell Biology, 164(7):959–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
7/959.

Shi:2000:TCA

[SFSD00] Wen Shi, Huizhou Fan, Lillian Shum, and Rik Derynck. The Tetraspanin Cd9 Associates with Transmembrane TGFα and Regulates TGF-α-Induced Egf Receptor Activation and Cell Proliferation. Journal of Cell Biology, 148(3):591– ??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/3/591.

Sisson:2000:LLN

[SFV^{+}00] John C. Sisson, Christine Field, Richard Ventura, Anne Royou, and William Sullivan. Lava Lamp, a Novel Peripheral Golgi Protein, Is Required for Drosophila melanogaster Cellularization. Journal of Cell Biology, 151(4):905–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
151/4/905.

Straley:2000:RTI

[SG00] Kimberly S. Straley and Samuel A. Green. Rapid Transport of Internalized P–Selectin to Late Endosomes and the Tgn. Journal of Cell Biology, 151(1):107–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/1/
107.

Schmoranzer:2000:ICE

[SGAS00] Jan Schmoranzer, Mark Goulian, Dan Axelrod, and Sanford M. Simon. Imaging Constitutive Exocytosis with Total Internal Reflection Fluorescence Microscopy. Journal of Cell Biology, 149(1):23–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/1/23.

Smith:2002:CRD

[SGC^{+}02] Katherine M. Smith, Alban Gaultier, Helene Cousin, Dominique Alfandari, Judith M. White, and Douglas W. DeSimone. The cysteine-rich domain regulates ADAM protease function in vivo. Journal of Cell Biology, 159(5):893– ??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/5/893.

Samper:2001:NTL

[SGdM^{+}01] Enrique Samper, Fermín A. Goytisolo, Josiane Ménissier de Murcia, Eva González-Suárez, Juan C. Cigudosa, Gilbert de Murcia, and María A. Blasco. Normal telomere length and chromosomal end capping in poly(ADP-ribose) polymerase– deficient mice and primary cells despite increased chromosomal instability. Journal of Cell Biology, 154(1):49–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/154/
1/49.

Schaeper:2000:CGC

[SGF^{+}00] Ute Schaeper, Niels H. Gehring, Klaus P. Fuchs, Martin Sachs, Bettina Kempkes, and Walter Birchmeier. Coupling of Gab1 to C–Met, Grb2, and Shp2 Mediates Biological Responses. Journal of Cell Biology, 149(7):1419–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/149/7/
1419.

Schirmer:2001:ILR

[SGG01] Eric C. Schirmer, Tinglu Guan, and Larry Gerace. Involvement of the Lamin Rod Domain in Heterotypic Lamin Interactions Important for Nuclear Organization. Journal of Cell Biology, 153(3):479–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/3/479.

Smith:2001:T

[SGHL01] Laurie G. Smith, Suzanne M. Gerttula, Shengcheng Han, and Joshua Levy. Tangled1. Journal of Cell Biology, 152(1):

231–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/1/231.

Santini:2002:GPC

[SGK02a] Francesca Santini, Ibragim Gaidarov, and James H. Keen. G protein–coupled receptor/arrestin3 modulation of the endocytic machinery. Journal of Cell Biology, 156(4):665– ??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/4/665.

Schweizer:2002:CGA

[SGK^{+}02b] Ulrich Schweizer, Jennifer Gunnersen, Christoph Karch, Stefan Wiese, Bettina Holtmann, Kiyoshi Takeda, Shizuo Akira, and Michael Sendtner. Conditional gene ablation of Stat3 reveals differential signaling requirements for survival of motoneurons during development and after nerve injury in the adult. Journal of Cell Biology, 156(2):287–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/2/
287.

Sterk:2000:TMC

[SGO^{+}00] Lotus M. Th. Sterk, Cecile A. W. Geuijen, Lauran C. J. M. Oomen, Jero Calafat, Hans Janssen, and Arnoud Sonnenberg. The Tetraspan Molecule Cd151, a Novel Constituent of Hemidesmosomes, Associates with the Integrin α6β4 and May Regulate the Spatial Organization of Hemidesmosomes. Journal of Cell Biology, 149(4):969–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/4/969.

Spencer:2000:ENI

[SGPL^{+}00] Kathryn S. R. Spencer, Diana Graus-Porta, Jie Leng, Nancy E. Hynes, and Richard L. Klemke. ErbB2 Is Necessary for Induction of Carcinoma Cell Invasion by Erbb Family Receptor Tyrosine Kinases. Journal of Cell Biology, 148(2): 385–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/2/385.

Gao:2001:NIG

[sGS01] Ya sheng Gao and Elizabeth Sztul. A Novel Interaction of the Golgi Complex with the Vimentin Intermediate Filament Cytoskeleton. Journal of Cell Biology, 152(5):877–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
5/877.

Spann:2002:ANL

[SGW^{+}02] Timothy P. Spann, Anne E. Goldman, Chen Wang, Sui Huang, and Robert D. Goldman. Alteration of nuclear lamin organization inhibits RNA polymerase II-dependent transcription. Journal of Cell Biology, 156(4):603–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/156/
4/603.

Savino:2001:NAR

[SGYD^{+}01] Tulia Maria Savino, Jeannine Gébrane-Younès, Jan De Mey, Jean-Baptiste Sibarita, and Danièle Hernandez-Verdun. Nucleolar assembly of the Rrna processing machinery in living cells. Journal of Cell Biology, 153(5):1097–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/153/5/
1097.

Sawhney:2002:SLM

[SH02] Ravi K. Sawhney and Jonathon Howard. Slow local movements of collagen fibers by fibroblasts drive the rapid global self-organization of collagen gels. Journal of Cell Biology, 157 (6):1083–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/6/1083.

Shmelzer:2003:UTC

[SHA^{+}03] Zeev Shmelzer, Nurit Haddad, Ester Admon, Itai Pessach, Thomas L. Leto, Zahit Eitan-Hazan, Michal Hershfinkel, and Rachel Levy. Unique targeting of cytosolic phospholipase A_{2} to plasma membranes mediated by the NADPH oxidase in phagocytes. Journal of Cell Biology, 162(4):683–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
4/683.

Schenk:2003:BER

[SHB^{+}03] Susann Schenk, Edith Hintermann, Martin Bilban, Naohiko Koshikawa, Carlo Hojilla, Rama Khokha, and Vito Quaranta. Binding to EGF receptor of a laminin-5 EGF-like fragment liberated during MMP-dependent mammary gland involution. Journal of Cell Biology, 161(1):197–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/1/
197.

Shirasu-Hiza:2003:IXM

[SHCM03] Mimi Shirasu-Hiza, Peg Coughlin, and Tim Mitchison. Identification of XMAP215 as a microtubule-destabilizing factor in Xenopus egg extract by biochemical purification. Journal of Cell Biology, 161(2):349–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/2/349.

Silke:2002:AAA

[SHE^{+}02] John Silke, Christine J. Hawkins, Paul G. Ekert, Joanne Chew, Catherine L. Day, Miha Pakusch, Anne M. Verhagen, and David L. Vaux. The anti-apoptotic activity of XIAP is retained upon mutation of both the caspase 3– and caspase 9–interacting sites. Journal of Cell Biology, 157(1):115–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
157/1/115.

Snapp:2003:FSC

[SHF^{+}03] Erik L. Snapp, Ramanujan S. Hegde, Maura Francolini, Francesca Lombardo, Sara Colombo, Emanuela Pedrazzini, Nica Borgese, and Jennifer Lippincott-Schwartz. Formation of stacked ER cisternae by low affinity protein interactions. Journal of Cell Biology, 163(2):257–??, October 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/163/2/
257.

Severin:2001:SPM

[SHHH01] Fedor Severin, Bianca Habermann, Tim Huffaker, and Tony Hyman. Stu2 Promotes Mitotic Spindle Elongation in Anaphase. Journal of Cell Biology, 153(2):435–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
2/435.

Smith:2002:TAP

[SHKS02] Matthew D. Smith, Andreas Hiltbrunner, Felix Kessler, and Danny J. Schnell. The targeting of the atToc159 preprotein receptor to the chloroplast outer membrane is mediated by its GTPase domain and is regulated by GTP. Journal of Cell Biology, 159(5):833–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/5/833.

Schwoebel:2002:MIR

[SHM02] Eric D. Schwoebel, Thai H. Ho, and Mary Shannon Moore. The mechanism of inhibition of Ran-dependent nuclear transport by cellular ATP depletion. Journal of Cell Biology, 157 (6):963–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/6/963.

Sanjay:2001:CAP

[SHN^{+}01] Archana Sanjay, Adam Houghton, Lynn Neff, Emilia DiDomenico, Chantal Bardelay, Evelyne Antoine, Joan Levy, James Gailit, David Bowtell, William C. Horne, and Roland Baron. Cbl associates with Pyk2 and Src to regulate Src kinase activity, α_{v}β_{3} integrin-mediated signaling, cell adhesion, and osteoclast motility. Journal of Cell Biology, 152(1): 181–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/1/181.

Severin:2001:CSE

[SHP01] Fedor Severin, Anthony A. Hyman, and Simonetta Piatti. Correct spindle elongation at the metaphase/anaphase transition is an APC-dependent event in budding yeast. Journal of Cell Biology, 155(5):711–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/5/711.

Sengbusch:2002:DFI

[SHPY02] Jennifer K. Sengbusch, Wei He, Karen A. Pinco, and Joy T. Yang. Dual functions of α4β1 integrin in epicardial development. Journal of Cell Biology, 157(5):873–??, May 2002.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/157/5/
873.

Sanchez:2000:LCN

[SHS^{+}00] Ivelisse Sánchez, Linda Hassinger, Ram K. Sihag, Don W. Cleveland, Panaiyur Mohan, and Ralph A. Nixon. Local Control of Neurofilament Accumulation during Radial Growth of Myelinating Axons in Vivo. Journal of Cell Biology, 151(5): 1013–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/5/1013.

Sirri:2002:CDK

[SHVR02] Valentina Sirri, Danièle Hernandez-Verdun, and Pascal Roussel. Cyclin-dependent kinases govern formation and maintenance of the nucleolus. Journal of Cell Biology, 156(6): 969–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/6/969.

Sander:2001:MIV

[SHW01] Andreas Sander, Boris A. Hesser, and Veit Witzemann. MuSK induces in vivo acetylcholine receptor clusters in a ligand-independent manner. Journal of Cell Biology, 155(7): 1287–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/7/1287.

Schmitz:2000:CTE

[SHWH00] Anton Schmitz, Helga Herrgen, Alexandra Winkeler, and Volker Herzog. Cholera Toxin Is Exported from Microsomes by the Sec61p Complex. Journal of Cell Biology, 148(6): 1203–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/6/1203.

Sabo:2001:AAP

[SIBG01] Shasta L. Sabo, Annat F. Ikin, Joseph D. Buxbaum, and Paul Greengard. The Alzheimer Amyloid Precursor Protein (APP) and Fe65, an APP–Binding Protein, Regulate Cell Movement. Journal of Cell Biology, 153(7):1403–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
7/1403.

Silverstein:2002:FTR

[Sil02] Roy L. Silverstein. The face of TSR revealed. Journal of Cell Biology, 159(2):203–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/2/203.

Sayan:2001:NNM

[SISO01] Berna S. Sayan, Gulayse Ince, A. Emre Sayan, and Mehmet Ozturk. NAPO as a novel marker for apoptosis. Journal of Cell Biology, 155(5):719–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/5/719.

Sesaki:2001:UEO

[SJ01a] Hiromi Sesaki and Robert E. Jensen. UGO1 Encodes an Outer Membrane Protein Required for Mitochondrial Fusion. Journal of Cell Biology, 152(6):1123–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/152/6/
1123.

Sinai:2001:TGP

[SJ01b] Anthony P. Sinai and Keith A. Joiner. The Toxoplasma gondii protein ROP2 mediates host organelle association with the parasitophorous vacuole membrane. Journal of Cell Biology, 154(1):95–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/1/95.

Suga:2000:TME

[SJA^{+}00] Yasushi Suga, Michal Jarnik, Paul S. Attar, Mary A. Longley, Donnie Bundman, Alasdair C. Steven, Peter J. Koch, and Dennis R. Roop. Transgenic Mice Expressing a Mutant Form of Loricrin Reveal the Molecular Basis of the Skin Diseases, Vohwinkel Syndrome and Progressive Symmetric Erythrokeratoderma. Journal of Cell Biology, 151(2):401–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
2/401.

Shopland:2003:CMS

[SJB^{+}03] Lindsay S. Shopland, Carol V. Johnson, Meg Byron, John McNeil, and Jeanne B. Lawrence. Clustering of multiple specific genes and gene-rich R-bands around SC-35 domains. Journal of Cell Biology, 162(6):981–??, September 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/162/6/
981.

Sanjuan:2001:RDK

[SJIM01] Miguel Angel Sanjuán, David R. Jones, Manuel Izquierdo, and Isabel Mérida. Role of Diacylglycerol Kinase α in the Attenuation of Receptor Signaling. Journal of Cell Biology, 153 (1):207–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/1/207.

Sriburi:2004:X

[SJMB04] Rungtawan Sriburi, Suzanne Jackowski, Kazutoshi Mori, and Joseph W. Brewer. Xbp1. Journal of Cell Biology, 167(1): 35–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/1/35.

Sacksteder:2000:PBM

[SJS^{+}00] Katherine A. Sacksteder, Jacob M. Jones, Sarah T. South, Xiaoling Li, Yifei Liu, and Stephen J. Gould. Pex19 Binds Multiple Peroxisomal Membrane Proteins, Is Predominantly Cytoplasmic, and Is Required for Peroxisome Membrane Synthesis. Journal of Cell Biology, 148(5):931–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/5/
931.

Saelim:2004:NMI

[SJW^{+}04] Nuttawut Saelim, Linu M. John, Jun Wu, Jeong Soon Park, Yidong Bai, Patricia Camacho, and James D. Lechleiter. Nontranscriptional modulation of intracellular Ca^{2+} signaling by ligand stimulated thyroid hormone receptor. Journal of Cell Biology, 167(5):915–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/5/915.

Sullivan:2001:CID

[SK01] Beth Sullivan and Gary Karpen. Centromere identity in Drosophila is not determined in vivo by replication timing. Journal of Cell Biology, 154(4):683–??, August 2001. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/154/4/
683.

Setoguchi:2004:NEO

[SK04] Takao Setoguchi and Toru Kondo. Nuclear export of OLIG2 in neural stem cells is essential for ciliary neurotrophic factor– induced astrocyte differentiation. Journal of Cell Biology, 166(7):963–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/7/963.

Kang:2001:FCD

[sKCK^{+}01] Jung seog Kang, Iain M. Cheeseman, George Kallstrom, Soundarapandian Velmurugan, Georjana Barnes, and Clarence S. M. Chan. Functional cooperation of Dam1, Ipl1, and the inner centromere protein (INCENP)–related protein Sli15 during chromosome segregation. Journal of Cell Biology, 155(5):763–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/5/763.

Snyder:2000:PPI

[SKCS00] William B. Snyder, Antonius Koller, A. Jobu Choy, and Suresh Subramani. The Peroxin Pex19p Interacts with Multiple, Integral Membrane Proteins at the Peroxisomal Membrane. Journal of Cell Biology, 149(6):1171–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/149/6/
1171.

Soriano:2001:PNR

[SKF^{+}01] Salvador Soriano, David E. Kang, Maofu Fu, Richard Pestell, Nathalie Chevallier, Hui Zheng, and Edward H. Koo. Presenilin 1 negatively regulates β-catenin/T cell factor/lymphoid enhancer factor-1 signaling independently of β-amyloid precursor protein and notch processing. Journal of Cell Biology, 152(4):785–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/4/785.

Schaefer:2002:FAA

[SKF02] Andrew W. Schaefer, Nurul Kabir, and Paul Forscher. Filopodia and actin arcs guide the assembly and transport of two populations of microtubules with unique dynamic parameters in neuronal growth cones. Journal of Cell Biology, 158(1):139–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/1/139.

Sekiya-Kawasaki:2003:DPC

[SKGC^{+}03] Mariko Sekiya-Kawasaki, Aaron Chris Groen, M. Jamie T. V. Cope, Marko Kaksonen, Hadiya A. Watson, Chao Zhang, Kevan M. Shokat, Beverly Wendland, Kent L. McDonald, J. Michael McCaffery, and David G. Drubin. Dynamic phosphoregulation of the cortical actin cytoskeleton and endocytic machinery revealed by real-time chemical genetic analysis. Journal of Cell Biology, 162(5):765–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/162/5/
765.

Stipp:2003:ERI

[SKH03] Christopher S. Stipp, Tatiana V. Kolesnikova, and Martin E. Hemler. EWI-2 regulates α3β1 integrin-dependent cell functions on laminin-5. Journal of Cell Biology, 163(5):1167– ??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/5/1167.

Spencer:2000:MTR

[SKJ^{+}00] Charlotte A. Spencer, Michael J. Kruhlak, Heather L. Jenkins, Xuejun Sun, and David P. Bazett-Jones. Mitotic Transcription Repression in Vivo in the Absence of Nucleosomal Chromatin Condensation. Journal of Cell Biology, 150 (1):13–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/1/13.

Schaefer:2002:LEC

[SKK^{+}02] Andrew W. Schaefer, Yoshimasa Kamei, Hiroyuki Kamiguchi, Eric V. Wong, Iris Rapoport, Tomas Kirchhausen, Carol M. Beach, Gary Landreth, Sandra K. Lemmon, and Vance Lemmon. L1 endocytosis is controlled by a phosphorylationdephosphorylation cycle stimulated by outside-in signaling by L1. Journal of Cell Biology, 157(7):1223–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/157/7/
1223.

Swaminathan:2004:RPO

[SKK^{+}04] Sowmya Swaminathan, Florian Kiendl, Roman Körner, Raffaella Lupetti, Ludger Hengst, and Frauke Melchior. Ran-GAP1*SUMO1 is phosphorylated at the onset of mitosis and remains associated with RanBP2 upon NPC disassembly. Journal of Cell Biology, 164(7):965–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/7/965.

Simons:2002:OMP

[SKM^{+}02] Mikael Simons, Eva-Maria Krämer, Paolo Macchi, Silvia Rathke-Hartlieb, Jacqueline Trotter, Klaus-Armin Nave, and Jörg B. Schulz. Overexpression of the myelin proteolipid protein leads to accumulation of cholesterol and proteolipid protein in endosomes/lysosomes. Journal of Cell Biology, 157 (2):327–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/2/327.

Sriram:2003:DOC

[SKM03] V. Sriram, K. S. Krishnan, and Satyajit Mayor. deep-orange and carnation define distinct stages in late endosomal biogenesis in Drosophila melanogaster. Journal of Cell Biology, 161 (3):593–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/3/593.

Shimonaka:2003:RTC

[SKN^{+}03] Mika Shimonaka, Koko Katagiri, Toshinori Nakayama, Naoya Fujita, Takashi Tsuruo, Osamu Yoshie, and Tatsuo Kinashi. Rap1 translates chemokine signals to integrin activation, cell polarization, and motility across vascular endothelium under flow. Journal of Cell Biology, 161(2):417–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
2/417.

Swiercz:2004:PBR

[SKO04] Jakub M. Swiercz, Rohini Kuner, and Stefan Offermanns. Plexin–B1/ RhoGEF-mediated RhoA activation involves the receptor tyrosine kinase ErbB-2. Journal of Cell Biology, 165 (6):869–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/6/869.

Shimoni:2000:LSC

[SKR^{+}00] Yuval Shimoni, Tatsuo Kurihara, Mariella Ravazzola, Mylène Amherdt, Lelio Orci, and Randy Schekman. Lst1p and Sec24p Cooperate in Sorting of the Plasma Membrane Atpase into Copii Vesicles in Saccharomyces cerevisiae. Journal of Cell Biology, 151(5):973–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/5/973.

Schinzel:2004:CCB

[SKS^{+}04] Anna Schinzel, Thomas Kaufmann, Martin Schuler, Jorge Martinalbo, David Grubb, and Christoph Borner. Conformational control of Bax localization and apoptotic activity by Pro168. Journal of Cell Biology, 164(7):1021–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
7/1021.

Simons:2000:AMA

[SKT^{+}00] Mikael Simons, Eva-Maria Krämer, Christoph Thiele, Wilhelm Stoffel, and Jacqueline Trotter. Assembly of Myelin by Association of Proteolipid Protein with Cholesteroland Galactosylceramide–Rich Membrane Domains. Journal of Cell Biology, 151(1):143–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/1/143.

Shvartsman:2003:DAI

[SKT^{+}03] Dmitry E. Shvartsman, Mariana Kotler, Renee D. Tall, Michael G. Roth, and Yoav I. Henis. Differently anchored influenza hemagglutinin mutants display distinct interaction dynamics with mutual rafts. Journal of Cell Biology, 163(4): 879–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/4/879.

Shopland:2000:SCG

[SL00] Lindsay S. Shopland and Jeanne B. Lawrence. Seeking Common Ground in Nuclear Complexity. Journal of Cell Biology, 150(1):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/1/F1.

Song:2001:NFS

[SL01] Sukgil Song and Kyung S. Lee. A Novel Function of Saccharomyces cerevisiae CDC5 in Cytokinesis. Journal of Cell Biology, 152(3):451–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/152/3/451.

Sakurai:2001:OFC

[SLB^{+}01] Takeshi Sakurai, Marc Lustig, Joanne Babiarz, Andrew J. W. Furley, Steven Tait, Peter J. Brophy, Stephen A. Brown, Lucia Y. Brown, Carol A. Mason, and Martin Grumet. Overlapping functions of the cell adhesion molecules Nr–CAM and L1 in cerebellar granule cell development. Journal of Cell Biology, 154(6):1259–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/6/1259.

Storey:2002:SIS

[SLB02] Nina Storey, David Latchman, and Stuart Bevan. Selective internalization of sodium channels in rat dorsal root ganglion neurons infected with herpes simplex virus-1. Journal of Cell Biology, 158(7):1251–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/7/1251.

Sytnyk:2002:NCA

[SLD^{+}02] Vladimir Sytnyk, Iryna Leshchyns’ka, Markus Delling, Galina Dityateva, Alexander Dityatev, and Melitta Schachner. Neural cell adhesion molecule promotes accumulation of TGN organelles at sites of neuron-to-neuron contacts. Journal of Cell Biology, 159(4):649–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/4/649.

Shang:2002:TTC

[SLG02] Yuhua Shang, Bing Li, and Martin A. Gorovsky. Tetrahymena thermophila contains a conventional γ-tubulin that is differentially required for the maintenance of different microtubule-organizing centers. Journal of Cell Biology, 158 (7):1195–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/7/1195.

Schuyler:2003:MFA

[SLP03] Scott C. Schuyler, Jenny Y. Liu, and David Pellman. The molecular function of Ase1p. Journal of Cell Biology, 160(4): 517–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/4/517.

Sandor:2003:IEI

[SLR^{+}03] Frantisek Sandor, Eicke Latz, Fabio Re, Leisa Mandell, Galina Repik, Douglas T. Golenbock, Terje Espevik, Evelyn A. Kurt-Jones, and Robert W. Finberg. Importance of extra- and intracellular domains of TLR1 and TLR2 in NFκB signaling. Journal of Cell Biology, 162(6):1099–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
162/6/1099.

Siniossoglou:2000:SAN

[SLSR^{+}00] Symeon Siniossoglou, Malik Lutzmann, Helena Santos-Rosa, Kevin Leonard, Shirley Mueller, Ueli Aebi, and Ed Hurt. Structure and Assembly of the Nup84p Complex. Journal of Cell Biology, 149(1):41–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/149/1/41.

Silflow:2001:VPC

[SLT^{+}01] Carolyn D. Silflow, Matthew LaVoie, Lai-Wa Tam, Susan Tousey, Mark Sanders, Wei chien Wu, Mark Borodovsky, and Paul A. Lefebvre. The Vfl1 Protein in Chlamydomonas Localizes in a Rotationally Asymmetric Pattern at the Distal Ends of the Basal Bodies. Journal of Cell Biology, 153 (1):63–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/1/63.

Sattler:2000:CFR

[SM00] Tanja Sattler and Andreas Mayer. Cell-free Reconstitution of Microautophagic Vacuole Invagination and Vesicle Formation. Journal of Cell Biology, 151(3):529–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/3/
529.

Setiadi:2003:SDD

[SM03a] Hendra Setiadi and Rodger P. McEver. Signal-dependent distribution of cell surface P-selectin in clathrin-coated pits affects leukocyte rolling under flow. Journal of Cell Biology, 163(6):1385–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/6/1385.

Stokes:2003:DDI

[SM03b] Matthew P. Stokes and W. Matthew Michael. DNA damageinduced replication arrest in Xenopus egg extracts. Journal of Cell Biology, 163(2):245–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/2/245.

Steinman:2004:JJ

[SM04a] Ralph M. Steinman and Ira Mellman. The JEM and the JCB. Journal of Cell Biology, 166(5):611–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/166/5/
611.

Sun:2004:CNR

[SM04b] Lu Sun and Khaled Machaca. Ca_{cyt}^{2+} negatively regulates the initiation of oocyte maturation. Journal of Cell Biology, 165 (1):63–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/1/63.

Smith:2002:TPI

[SMC^{+}02] Jennifer J. Smith, Marcello Marelli, Rowan H. Christmas, Franco J. Vizeacoumar, David J. Dilworth, Trey Ideker, Timothy Galitski, Krassen Dimitrov, Richard A. Rachubinski, and John D. Aitchison. Transcriptome profiling to identify genes involved in peroxisome assembly and function. Journal of Cell Biology, 158(2):259–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/2/259.

Scholl:2000:DPR

[SME^{+}00] Florence A. Scholl, Patricia McLoughlin, Elisabeth Ehler, Carla de Giovanni, and Beat W. Schäfer. Dral Is a P53-Responsive Gene Whose Four and a Half Lim Domain Protein Product Induces Apoptosis. Journal of Cell Biology, 151(3):495–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/3/495.

Supek:2002:SPA

[SMH^{+}02] Frantisek Supek, David T. Madden, Susan Hamamoto, Lelio Orci, and Randy Schekman. Sec16p potentiates the action of COPII proteins to bud transport vesicles. Journal of Cell Biology, 158(6):1029–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/6/1029.

Sakaguchi:2000:RBC

[SMI^{+}00] Masakiyo Sakaguchi, Masahiro Miyazaki, Yusuke Inoue, Toshiya Tsuji, Hirosuke Kouchi, Toshio Tanaka, Hidenori Yamada, and Masayoshi Namba. Relationship between Contact Inhibition and Intranuclear S100c of Normal Human Fibroblasts. Journal of Cell Biology, 149(6):1193–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
6/1193.

Sumiyoshi:2004:EMP

[SML^{+}04] Hideaki Sumiyoshi, Niv Mor, Sui Y. Lee, Stephen Doty, Scott Henderson, Shizuko Tanaka, Hidekatsu Yoshioka, Satish Rattan, and Francesco Ramirez. Esophageal muscle physiology and morphogenesis require assembly of a collagen XIX–rich basement membrane zone. Journal of Cell Biology, 166(4): 591–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/4/591.

Skoufias:2000:HSK

[SMLM00] Dimitrios A. Skoufias, Cristiana Mollinari, Françoise B. Lacroix, and Robert L. Margolis. Human Survivin Is a Kinetochore–Associated Passenger Protein. Journal of Cell Biology, 151(7):1575–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/7/1575.

Smith:2002:CCS

[SMR^{+}02] Peter L. Smith, Jay T. Myers, Clare E. Rogers, Lan Zhou, Bronia Petryniak, Daniel J. Becker, Jonathon W. Homeister, and John B. Lowe. Conditional control of selectin ligand expression and global fucosylation events in mice with a targeted mutation at the FX locus. Journal of Cell Biology, 158 (4):801–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/4/801.

Shelby:2000:CAK

[SMS00] Richard D. Shelby, Karine Monier, and Kevin F. Sullivan. Chromatin Assembly at Kinetochores Is Uncoupled from DNA Replication. Journal of Cell Biology, 151(5):1113– ??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/5/1113.

Seewaldt:2001:SPF

[SMS^{+}01a] Victoria L. Seewaldt, Krzysztof Mrózek, Randy Sigle, Eric C. Dietze, Kevin Heine, David M. Hockenbery, Katherine B. Hobbs, and L. Elizabeth Caldwell. Suppression of p53 function in normal human mammary epithelial cells increases sensitivity to extracellular matrix–induced apoptosis. Journal of Cell Biology, 155(3):471–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/3/471.

Shimizu:2001:ERV

[SMS^{+}01b] Shigeomi Shimizu, Yosuke Matsuoka, Yasuo Shinohara, Yoshihiro Yoneda, and Yoshihide Tsujimoto. Essential Role of Voltage–Dependent Anion Channel in Various Forms of Apoptosis in Mammalian Cells. Journal of Cell Biology, 152(2):237–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/2/237.

Sakaguchi:2004:PMT

[SMS^{+}04a] Masakiyo Sakaguchi, Masahiro Miyazaki, Hiroyuki Sonegawa, Mariko Kashiwagi, Motoi Ohba, Toshio Kuroki, Masayoshi Namba, and Nam ho Huh. PKCα mediates TGFβ-induced growth inhibition of human keratinocytes via phosphorylation of S100C/A11. Journal of Cell Biology, 164 (7):979–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/7/979.

Siegel:2004:S

[SMS^{+}04b] Richard M. Siegel, Jagan R. Muppidi, Malabika Sarker, Adrian Lobito, Melinda Jen, David Martin, Stephen E. Straus, and Michael J. Lenardo. Spots. Journal of Cell Biology, 167(4):735–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/4/735.

Shimizu:2000:DCA

[SMSF00] Masayuki Shimizu, Yasunori Murakami, Fumikazu Suto, and Hajime Fujisawa. Determination of Cell Adhesion Sites of Neuropilin-1. Journal of Cell Biology, 148(6):1283–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
6/1283.

Silletti:2000:PSD

[SMSM00] Steve Silletti, Fang Mei, Dean Sheppard, and Anthony M. P. Montgomery. Plasmin-sensitive Dibasic Sequences in the Third Fibronectin-like Domain of L1–Cell Adhesion Molecule (Cam) Facilitate Homomultimerization and Concomitant Integrin Recruitment. Journal of Cell Biology, 149(7):1485–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
149/7/1485.

Sakaguchi:2003:SAK

[SMT^{+}03] Masakiyo Sakaguchi, Masahiro Miyazaki, Mikiro Takaishi, Yoshihiko Sakaguchi, Eiichi Makino, Noriyuki Kataoka, Hidenori Yamada, Masayoshi Namba, and Nam ho Huh. S100C/ a11 is a key mediator of Ca^{2+}-induced growth inhibition of human epidermal keratinocytes. Journal of Cell Biology, 163(4):825–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/4/825.

Steen:2000:RPP

[SMTC00] Rikke L. Steen, Sandra B. Martins, Kjetil Taskén, and Philippe Collas. Recruitment of Protein Phosphatase 1 to the Nuclear Envelope by a–Kinase Anchoring Protein Akap149 Is a Prerequisite for Nuclear Lamina Assembly. Journal of Cell Biology, 150(6):1251–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/6/1251.

Shin:2003:SEE

[SMW^{+}03] Jennifer H. Shin, L. Mahadevan, Guillermina S. Waller, Knut Langsetmo, and Paul Matsudaira. Stored elastic energy powers the 60-μ m extension of the Limulus polyphemus sperm actin bundle. Journal of Cell Biology, 162(7):1183– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/7/1183.

Shulga:2000:YNI

[SMWG00] Nataliya Shulga, Nima Mosammaparast, Richard Wozniak, and David S. Goldfarb. Yeast Nucleoporins Involved in Passive Nuclear Envelope Permeability. Journal of Cell Biology, 149(5):1027–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/5/1027.

Silva:2003:RRR

[SMZ^{+}03] Jorge Santos Da Silva, Miguel Medina, Cecilia Zuliani, Alessia Di Nardo, Walter Witke, and Carlos G. Dotti. RhoA/ ROCK regulation of neuritogenesis via profilin IIa-mediated control of actin stability. Journal of Cell Biology, 162(7): 1267–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/7/1267.

Stanek:2004:DSA

[SN04] David Staněk and Karla M. Neugebauer. Detection of snRNP assembly intermediates in Cajal bodies by fluorescence resonance energy transfer. Journal of Cell Biology, 166(7):1015– ??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/7/1015.

Schwer:2002:HSI

[SNF^{+}02] Björn Schwer, Brian J. North, Roy A. Frye, Melanie Ott, and Eric Verdin. The human silent information regulator (Sir)2 homologue hSIRT3 is a mitochondrial nicotinamide adenine dinucleotide-dependent deacetylase. Journal of Cell Biology, 158(4):647–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/4/647.

Sancho:2000:TKP

[SNL^{+}00] David Sancho, Marta Nieto, Manuel Llano, José L. Rodríguez-Fernández, Reyes Tejedor, Shalom Avraham, Carlos Cabañas, Miguel López-Botet, and Francisco Sánchez-Madrid. The tyrosine kinase Pyk-2/Raftk regulates natural killer (Nk) cell cytotoxic response, and is translocated and activated upon specific target cell recognition and killing. Journal of Cell Biology, 149(6):1249–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/6/1249.

Suzuki:2004:BIR

[SNS^{+}04] Shingo Suzuki, Tadahiro Numakawa, Kazuhiro Shimazu, Hisatsugu Koshimizu, Tomoko Hara, Hiroshi Hatanaka, Lin Mei, Bai Lu, and Masami Kojima. BDNF-induced recruitment of TrkB receptor into neuronal lipid rafts. Journal of Cell Biology, 167(6):1205–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/6/1205.

Sodeik:2002:UMH

[Sod02] Beate Sodeik. Unchain my heart, baby let me go — the entry and intracellular transport of HIV. Journal of Cell Biology, 159(3):393–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/3/393.

Sabeh:2004:TCT

[SOH^{+}04] Farideh Sabeh, Ichiro Ota, Kenn Holmbeck, Henning Birkedal-Hansen, Paul Soloway, Milagros Balbin, Carlos Lopez-Otin, Steven Shapiro, Masaki Inada, Stephen Krane, Edward Allen, Duane Chung, and Stephen J. Weiss. Tumor cell traffic through the extracellular matrix is controlled by the membrane-anchored collagenase MT1-MMP. Journal of Cell Biology, 167(4):769–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/4/769.

Smyth:2002:TD

[SP02] Susan S. Smyth and Cam Patterson. Tiny dancers. Journal of Cell Biology, 158(1):17–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/1/17.

Sekar:2003:FRE

[SP03] Rajesh Babu Sekar and Ammasi Periasamy. Fluorescence resonance energy transfer (FRET) microscopy imaging of live cell protein localizations. Journal of Cell Biology, 160(5): 629–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/5/629.

Subauste:2004:VMP

[SPA^{+}04] M. Cecilia Subauste, Olivier Pertz, Eileen D. Adamson, Christopher E. Turner, Sachiko Junger, and Klaus M. Hahn. Vinculin modulation of paxillin–FAK interactions regulates ERK to control survival and motility. Journal of Cell Biology, 165(3):371–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/3/371.

Stupack:2001:AAC

[SPB^{+}01] Dwayne G. Stupack, Xose S. Puente, Souphaphone Boutsaboualoy, Chris M. Storgard, and David A. Cheresh. Apoptosis of adherent cells by recruitment of caspase-8 to unligated integrins. Journal of Cell Biology, 155(3):459–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/155/3/
459.

Schratt:2002:SRF

[SPB^{+}02] Gerhard Schratt, Ulrike Philippar, Jürgen Berger, Heinz Schwarz, Olaf Heidenreich, and Alfred Nordheim. Serum response factor is crucial for actin cytoskeletal organization and focal adhesion assembly in embryonic stem cells. Journal of Cell Biology, 156(4):737–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/4/737.

Sells:2000:TSD

[SPC00] Mary Ann Sells, Amanda Pfaff, and Jonathan Chernoff. Temporal and Spatial Distribution of Activated Pak1 in Fibroblasts. Journal of Cell Biology, 151(7):1449–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
7/1449.

Suzuki:2000:PYG

[SPH^{+}00] Tadashi Suzuki, Hangil Park, Nancy M. Hollingsworth, Rolf Sternglanz, and William J. Lennarz. PNG1, a Yeast Gene Encoding a Highly Conserved Peptide: N -Glycanase. Journal of Cell Biology, 149(5):1039–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/149/5/1039.

Short:2001:GRE

[SPK^{+}01] Benjamin Short, Christian Preisinger, Roman Körner, Robert Kopajtich, Olwyn Byron, and Francis A. Barr. A GRASP55-rab2 effector complex linking Golgi structure to membrane traffic. Journal of Cell Biology, 155(6):877–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/155/6/877.

Signoret:2000:ERH

[SPMM^{+}00] Nathalie Signoret, Annegret Pelchen-Matthews, Matthias Mack, Amanda E. I. Proudfoot, and Mark Marsh. Endocytosis and Recycling of the HIV Coreceptor Ccr5. Journal of Cell Biology, 151(6):1281–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/6/1281.

Sharma:2004:MDC

[SPN^{+}04] Manu Sharma, Francesca Pampinella, Csilla Nemes, Mohamed Benharouga, Jeffrey So, Kai Du, Kristi G. Bache, Blake Papsin, Noa Zerangue, Harald Stenmark, and Gergely L. Lukacs. Misfolding diverts CFTR from recycling to degradation. Journal of Cell Biology, 164(6):923–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/164/6/923.

Sheff:2002:TRR

[SPO^{+}02] David Sheff, Laurence Pelletier, Christopher B. O’Connell, Graham Warren, and Ira Mellman. Transferrin receptor recycling in the absence of perinuclear recycling endosomes. Journal of Cell Biology, 156(5):797–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/5/797.

Santolini:2000:NEP

[SPS^{+}00] Elisa Santolini, Claudia Puri, Anna Elisabetta Salcini, Maria Cristina Gagliani, Pier Giuseppe Pelicci, Carlo Tacchetti, and Pier Paolo Di Fiore. Numb Is an Endocytic Protein. Journal of Cell Biology, 151(6):1345–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
6/1345.

Skoble:2000:TRW

[SPW00] Justin Skoble, Daniel A. Portnoy, and Matthew D. Welch. Three regions within ActA promote Arp2/3 complexmediated actin nucleation and Listeria monocytogenes motility. Journal of Cell Biology, 150(3):527–??, August 2000.

CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/3/
527.

Soulet:2003:PPC

[SR03] Denis Soulet and Serge Rivest. Polyamines play a critical role in the control of the innate immune response in the mouse central nervous system. Journal of Cell Biology, 162 (2):257–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/2/257.

Snapp:2004:OEQ

[SRB^{+}04] Erik L. Snapp, Gretchen A. Reinhart, Brigitte A. Bogert, Jennifer Lippincott-Schwartz, and Ramanujan S. Hegde. The organization of engaged and quiescent translocons in the endoplasmic reticulum of mammalian cells. Journal of Cell Biology, 164(7):997–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/7/997.

Sechler:2001:NFB

[SRC^{+}01] Jan L. Sechler, Hongwei Rao, Anne Marie Cumiskey, Irbert Vega-Colón, Michael S. Smith, Takatoshi Murata, and Jean E. Schwarzbauer. A novel fibronectin binding site required for fibronectin fibril growth during matrix assembly. Journal of Cell Biology, 154(5):1081–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/154/5/
1081.

Spessotto:2002:HCI

[SRD^{+}02] Paola Spessotto, Francesca Maria Rossi, Massimo Degan, Raffaele Di Francia, Roberto Perris, Alfonso Colombatti, and Valter Gattei. Hyaluronan-CD44 interaction hampers migration of osteoclast-like cells by down-regulating MMP-9. Journal of Cell Biology, 158(6):1133–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/6/
1133.

Sreenivasan:2004:AGH

[Sre04a] Aparna Sreenivasan. Actin gets Hip at the TGN. Journal of Cell Biology, 165(6):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/6/754.2.

Sreenivasan:2004:AAA

[Sre04b] Aparna Sreenivasan. Anti-angiogenesis is anti-actin. Journal of Cell Biology, 166(1):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/1/6.2.

Sreenivasan:2004:BC

[Sre04c] Aparna Sreenivasan. Building centrioles. Journal of Cell Biology, 165(5):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/5/603.3.

Sreenivasan:2004:CSO

[Sre04d] Aparna Sreenivasan. Calcium signals from outside. Journal of Cell Biology, 166(1):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/1/7.1.

Sreenivasan:2004:EE

[Sre04e] Aparna Sreenivasan. An exit exchanger. Journal of Cell Biology, 165(5):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/5/603.1.

Sreenivasan:2004:FR

[Sre04f] Aparna Sreenivasan. Fast rafting. Journal of Cell Biology, 165(5):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/5/603.2.

Sreenivasan:2004:FF

[Sre04g] Aparna Sreenivasan. Fusion at the front. Journal of Cell Biology, 166(1):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/1/6.1.

Sreenivasan:2004:MMa

[Sre04h] Aparna Sreenivasan. Mineralized mice. Journal of Cell Biology, 165(5):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/5/602.2.

Sreenivasan:2004:MG

[Sre04i] Aparna Sreenivasan. Mixed guidance. Journal of Cell Biology, 165(6):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/6/755.2.

Sreenivasan:2004:MMb

[Sre04j] Aparna Sreenivasan. Myelin movement. Journal of Cell Biology, 166(1):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/1/7.2.

Sreenivasan:2004:OIF

[Sre04k] Aparna Sreenivasan. Orbit(ing) the furrow. Journal of Cell Biology, 166(1):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/1/7.3.

Sreenivasan:2004:PC

[Sre04l] Aparna Sreenivasan. Pulling on a cytokine. Journal of Cell Biology, 165(5):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/5/602.1.

Sreenivasan:2004:SD

[Sre04m] Aparna Sreenivasan. Stuck in the divide. Journal of Cell Biology, 165(6):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/6/755.1.

Sreenivasan:2004:TMD

[Sre04n] Aparna Sreenivasan. Tumors make do with less matrix. Journal of Cell Biology, 165(6):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/6/755.3.

Sreenivasan:2004:ZC

[Sre04o] Aparna Sreenivasan. Zombie cells. Journal of Cell Biology, 165(6):??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/6/754.1.

Schuetz:2004:NSP

[SRG^{+}04] Gunnar Schuetz, Marta Rosário, Jan Grimm, Tobias M. Boeckers, Eckart D. Gundelfinger, and Walter Birchmeier. The neuronal scaffold protein Shank3 mediates signaling and biological function of the receptor tyrosine kinase Ret in epithelial cells. Journal of Cell Biology, 167(5):945–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
167/5/945.

Sirri:2000:VRM

[SRHV00] Valentina Sirri, Pascal Roussel, and Danièle Hernandez-Verdun. In Vivo Release of Mitotic Silencing of Ribosomal Gene Transcription Does Not Give Rise to Precursor Ribosomal RNA Processing. Journal of Cell Biology, 148(2): 259–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/2/259.

Samarin:2003:HVE

[SRK^{+}03] Stanislav Samarin, Stéphane Romero, Christine Kocks, Dominique Didry, Dominique Pantaloni, and Marie-France Carlier. How VASP enhances actin-based motility. Journal of Cell Biology, 163(1):131–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/1/131.

Stanek:2003:TUU

[SRKN03] David Stanĕk, Stephen D. Rader, Mirko Klingauf, and Karla M. Neugebauer. Targeting of U4/U6 small nuclear RNP assembly factor SART3/p110 to Cajal bodies. Journal of Cell Biology, 160(4):505–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/4/505.

Sherman:2000:CEN

[SRKR00] Larry S. Sherman, Tilat A. Rizvi, Saikumar Karyala, and Nancy Ratner. Cd44 Enhances Neuregulin Signaling by Schwann Cells. Journal of Cell Biology, 150(5):1071–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/150/5/1071.

Sengupta:2004:FIB

[SRL^{+}04] Sagar Sengupta, Ana I. Robles, Steven P. Linke, Natasha I. Sinogeeva, Ran Zhang, Remy Pedeux, Irene M. Ward, Arkady Celeste, André Nussenzweig, Junjie Chen, Thanos D. Halazonetis, and Curtis C. Harris. Functional interaction between BLM helicase and 53BP1 in a Chk1-mediated pathway during S-phase arrest. Journal of Cell Biology, 166(6):801– ??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/6/801.

Scholey:2001:MMM

[SRS01] Jonathan M. Scholey, Gregory C. Rogers, and David J. Sharp. Mitosis, microtubules, and the matrix. Journal of Cell Biology, 154(2):261–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/2/261.

See:2004:CDR

[SRSW04] Violaine Sée, Nina K. M. Rajala, David G. Spiller, and Michael R. H. White. Calcium-dependent regulation of the cell cycle via a novel MAPK-NF-κB pathway in Swiss 3T3 cells. Journal of Cell Biology, 166(5):661–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/166/5/
661.

Sciorra:2002:DRP

[SRW^{+}02] Vicki A. Sciorra, Simon A. Rudge, Jiyao Wang, Stuart McLaughlin, JoAnne Engebrecht, and Andrew J. Morris. Dual role for phosphoinositides in regulation of yeast and mammalian phospholipase D enzymes. Journal of Cell Biology, 159(6):1039–??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/6/1039.

Smith:2004:AST

[SRW^{+}04] Matthew D. Smith, Caleb M. Rounds, Fei Wang, Kunhua Chen, Meshack Afitlhile, and Danny J. Schnell. atToc159 is a selective transit peptide receptor for the import of nucleusencoded chloroplast proteins. Journal of Cell Biology, 165 (3):323–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/3/323.

Sohrt:2000:TNC

[SS00] Karen Sohrt and Jürgen Soll. Toc64, a New Component of the Protein Translocon of Chloroplasts. Journal of Cell Biology, 148(6):1213–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/6/1213.

Sanger:2001:FPB

[SS01] Joseph W. Sanger and Jean M. Sanger. Fishing out proteins that bind to titin. Journal of Cell Biology, 154(1):21–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/154/
1/21.

Sawada:2002:FTT

[SS02] Yasuhiro Sawada and Michael P. Sheetz. Force transduction by Triton cytoskeletons. Journal of Cell Biology, 156(4): 609–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/4/609.

Sandqvist:2004:NSG

[SS04] Anton Sandqvist and Lea Sistonen. Nuclear stress granules. Journal of Cell Biology, 164(1):15–??, January 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/1/
15.

Schapiro:2000:GAP

[SSA^{+}00] Florencia Schapiro, Jason Sparkowski, Alex Adduci, Frank Suprynowicz, Richard Schlegel, and Sergio Grinstein. Golgi Alkalinization by the Papillomavirus E5 Oncoprotein. Journal of Cell Biology, 148(2):305–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/2/305.

Steiner:2002:MRC

[SSG^{+}02] Pascal Steiner, J.-C. Floyd Sarria, Liliane Glauser, Sarah Magnin, Stefan Catsicas, and Harald Hirling. Modulation of receptor cycling by neuron-enriched endosomal protein of 21 kD. Journal of Cell Biology, 157(7):1197–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/157/7/
1197.

Strick:2001:CCB

[SSGLS01] Reiner Strick, Pamela L. Strissel, Konstantin Gavrilov, and Riccardo Levi-Setti. Cation–chromatin binding as shown by ion microscopy is essential for the structural integrity of chromosomes. Journal of Cell Biology, 155(6):899–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
6/899.

Shi:2004:GPR

[SSH^{+}04] Weibin Shi, Chuanxi Sun, Bin He, Wencheng Xiong, Xingming Shi, Dachun Yao, and Xu Cao. GADD34–PP1c recruited by Smad7 dephosphorylates TGFβ type I receptor. Journal of Cell Biology, 164(2):291–??, January 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/2/
291.

Schleiff:2003:CTO

[SSK^{+}03] Enrico Schleiff, Jürgen Soll, Michael Küchler, Werner Kühlbrandt, and Roswitha Harrer. Characterization of the translocon of the outer envelope of chloroplasts. Journal of Cell Biology, 160(4):541–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/4/541.

South:2000:ICC

[SSL^{+}00] Sarah T. South, Katherine A. Sacksteder, Xiaoling Li, Yifei Liu, and Stephen J. Gould. Inhibitors of COPI and COPII do not block PEX3-mediated peroxisome synthesis. Journal of Cell Biology, 149(7):1345–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/149/7/1345.

Scott:2004:ACD

[SSM^{+}04] Clare L. Scott, Martin Schuler, Vanessa S. Marsden, Alex Egle, Marc Pellegrini, Dobrila Nesic, Steve Gerondakis, Stephen L. Nutt, Douglas R. Green, and Andreas Strasser. Apaf-1 and caspase-9 do not act as tumor suppressors in mycinduced lymphomagenesis or mouse embryo fibroblast transformation. Journal of Cell Biology, 164(1):89–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
1/89.

Sato:2001:RRR

[SSN01] Ken Sato, Miyuki Sato, and Akihiko Nakano. Rer1p, a Retrieval Receptor for Endoplasmic Reticulum Membrane Proteins, Is Dynamically Localized to the Golgi Apparatus by Coatomer. Journal of Cell Biology, 152(5):935–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
5/935.

Shigeta:2003:CRE

[SSO^{+}03] Masaki Shigeta, Noriko Sanzen, Masayuki Ozawa, Jianguo Gu, Hitoshi Hasegawa, and Kiyotoshi Sekiguchi. CD151 regulates epithelial cell–cell adhesion through PKC- and Cdc42dependent actin cytoskeletal reorganization. Journal of Cell Biology, 163(1):165–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/1/165.

Sussman:2001:PKB

[SSOS01] Joshua Sussman, David Stokoe, Natalya Ossina, and Emma Shtivelman. Protein kinase B phosphorylates AHNAK and regulates its subcellular localization. Journal of Cell Biology, 154(5):1019–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/5/1019.

Sheldahl:2003:DAC

[SSP^{+}03] Laird C. Sheldahl, Diane C. Slusarski, Petra Pandur, Jeffrey R. Miller, Michael Kühl, and Randall T. Moon. Dishevelled activates Ca^{2+} flux, PKC, and CamKII in vertebrate embryos. Journal of Cell Biology, 161(4):769–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
4/769.

Smith:2004:RRM

[SSRX04] Elizabeth R. Smith, Jennifer L. Smedberg, Malgorzata E. Rula, and Xiang-Xi Xu. Regulation of Ras–MAPK pathway mitogenic activity by restricting nuclear entry of activated MAPK in endoderm differentiation of embryonic carcinoma and stem cells. Journal of Cell Biology, 164(5):689–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
5/689.

Sugita:2001:SCN

[SST^{+}01] Shuzo Sugita, Fumiaki Saito, Jiong Tang, Jakob Satz, Kevin Campbell, and Thomas C. Südhof. A stoichiometric complex of neurexins and dystroglycan in brain. Journal of Cell Biology, 154(2):435–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/2/435.

Stepanek:2001:CCN

[SSW^{+}01a] Laurie Stepanek, Qi Lun Sun, Jun Wang, Cong Wang, and John L. Bixby. CRYP-2/ cPTPRO is a neurite inhibitory repulsive guidance cue for retinal neurons in vitro. Journal of Cell Biology, 154(4):867–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/4/867.

Suhr:2001:ISP

[SSW^{+}01b] Steven T. Suhr, Marie-Claude Senut, Julian P. Whitelegge, Kym F. Faull, Denise B. Cuizon, and Fred H. Gage. Identities of Sequestered Proteins in Aggregates from Cells with Induced Polyglutamine Expression. Journal of Cell Biology, 153(2):283–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/2/283.

Shinzawa:2003:PAR

[ST03] Koei Shinzawa and Yoshihide Tsujimoto. PLA_{2} activity is required for nuclear shrinkage in caspase-independent cell death. Journal of Cell Biology, 163(6):1219–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
6/1219.

Scita:2001:ERE

[STA^{+}01] Giorgio Scita, Pierluigi Tenca, Liliana B. Areces, Arianna Tocchetti, Emanuela Frittoli, Giuseppina Giardina, Isabella Ponzanelli, Patrizia Sini, Metello Innocenti, and Pier Paolo Di Fiore. An effector region in Eps8 is responsible for the activation of the Rac-specific GEF activity of Sos-1 and for the proper localization of the Rac-based actin–polymerizing machine. Journal of Cell Biology, 154(5):1031–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/154/
5/1031.

Schulte:2003:GNM

[STA03] Joost Schulte, Ulrich Tepass, and Vanessa J. Auld. Gliotactin, a novel marker of tricellular junctions, is necessary for septate junction development in Drosophila. Journal of Cell Biology, 161(5):991–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/5/991.

Steinert:2000:CRE

[Ste00] Peter M. Steinert. The Complexity and Redundancy of Epithelial Barrier Function. Journal of Cell Biology, 151 (2):??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/2/F5.

Speransky:2001:PFN

[STE^{+}01] Vladislav V. Speransky, Kimberly L. Taylor, Herman K. Edskes, Reed B. Wickner, and Alasdair C. Steven. Prion Filament Networks in [Ure3] Cells of Saccharomyces cerevisiae. Journal of Cell Biology, 153(6):1327–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/6/1327.

Swindle:2001:EGF

[STJ^{+}01] C. Scott Swindle, Kien T. Tran, Terry D. Johnson, Pallab Banerjee, Anne M. Mayes, Linda Griffith, and Alan Wells. Epidermal growth factor (EGF)-like repeats of human tenascin–C as ligands for EGF receptor. Journal of Cell Biology, 154(2):459–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/2/459.

Segura-Totten:2002:BAF

[STKCW02] Miriam Segura-Totten, Amy K. Kowalski, Robert Craigie, and Katherine L. Wilson. Barrier-to-autointegration factor. Journal of Cell Biology, 158(3):475–??, August 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/3/
475.

Shan:2000:FCH

[STP^{+}00] Wei-Song Shan, Hidekazu Tanaka, Greg R. Phillips, Kirsten Arndt, Mika Yoshida, David R. Colman, and Lawrence Shapiro. Functional Cis–Heterodimers of N- and R– Cadherins. Journal of Cell Biology, 148(3):579–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
3/579.

Stukenberg:2004:TPA

[Stu04] P. Todd Stukenberg. Triggering p53 after cytokinesis failure. Journal of Cell Biology, 165(5):607–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/5/607.

Subramani:2002:HFR

[Sub02] Suresh Subramani. Hitchhiking fads en route to peroxisomes. Journal of Cell Biology, 156(3):415–??, February 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/3/
415.

Schuh:2001:PCD

[SUT^{+}01] Kai Schuh, Stjepan Uldrijan, Myriam Telkamp, Nicola Röthlein, and Ludwig Neyses. The plasmamembrane calmodulin-dependent calcium pump. Journal of Cell Biology, 155(2):201–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/2/201.

Sonawane:2003:DCR

[SV03] N. D. Sonawane and A. S. Verkman. Determinants of [Cl−] in recycling and late endosomes and Golgi complex measured using fluorescent ligands. Journal of Cell Biology, 160(7): 1129–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/7/1129.

Sumara:2000:CVC

[SVG^{+}00] Izabela Sumara, Elisabeth Vorlaufer, Christian Gieffers, Beate H. Peters, and Jan-Michael Peters. Characterization of Vertebrate Cohesin Complexes and Their Regulation in Prophase. Journal of Cell Biology, 151(4):749–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
4/749.

Scott:2004:SDS

[SVI^{+}04] Brenton L. Scott, Jeffrey S. Van Komen, Hassan Irshad, Song Liu, Kirilee A. Wilson, and James A. McNew. Sec1p directly stimulates SNARE-mediated membrane fusion in vitro. Journal of Cell Biology, 167(1):75–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/1/75.

Stokes:2002:DRR

[SVLM02] Matthew P. Stokes, Ruth Van Hatten, Howard D. Lindsay, and W. Matthew Michael. DNA replication is required for the checkpoint response to damaged DNA in Xenopus egg extracts. Journal of Cell Biology, 158(5):863–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
5/863.

Stamer:2002:TBT

[SVT^{+}02] K. Stamer, R. Vogel, E. Thies, E. Mandelkow, and E.-M. Mandelkow. Tau blocks traffic of organelles, neurofilaments, and APP vesicles in neurons and enhances oxidative stress. Journal of Cell Biology, 156(6):1051–??, March 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/6/

Strukov:2003:ECR

[SWB03] Yuri G. Strukov, Yan Wang, and Andrew S. Belmont. Engineered chromosome regions with altered sequence composition demonstrate hierarchical large-scale folding within metaphase chromosomes. Journal of Cell Biology, 162(1): 23–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/1/23.

Schaeren-Wiemers:2004:RAP

[SWBE^{+}04] Nicole Schaeren-Wiemers, Annick Bonnet, Michael Erb, Beat Erne, Udo Bartsch, Frances Kern, Ned Mantei, Diane Sherman, and Ueli Suter. The raft-associated protein MAL is required for maintenance of proper axon–glia interactions in the central nervous system. Journal of Cell Biology, 166(5): 731–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/5/731.

Sturge:2003:GAU

[SWE^{+}03] Justin Sturge, Dirk Wienke, Lucy East, Gareth E. Jones, and Clare M. Isacke. GPI-anchored uPAR requires Endo180 for rapid directional sensing during chemotaxis. Journal of Cell Biology, 162(5):789–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/5/789.

Srinivasan:2003:RCP

[SWG^{+}03] Supriya Srinivasan, Fei Wang, Suzana Glavas, Alexander Ott, Fred Hofmann, Klaus Aktories, Daniel Kalman, and Henry R. Bourne. Rac and Cdc42 play distinct roles in regulating PI(3,4,5)P_{3} and polarity during neutrophil chemotaxis. Journal of Cell Biology, 160(3):375–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/3/
375.

Schmidt:2002:CMS

[SWH^{+}02] Hannes Schmidt, Matthias Werner, Paul A. Heppenstall, Mechthild Henning, Margret I. Moré, Susanne Kühbandner, Gary R. Lewin, Franz Hofmann, Robert Feil, and Fritz G. Rathjen. cGMP-mediated signaling via cGKIα is required for the guidance and connectivity of sensory axons. Journal of Cell Biology, 159(3):489–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/3/489.

Snaar:2000:MAF

[SWJ^{+}00] Sabine Snaar, Karien Wiesmeijer, Aart G. Jochemsen, Hans J. Tanke, and Roeland W. Dirks. Mutational Analysis of Fibrillarin and Its Mobility in Living Human Cells. Journal of Cell Biology, 151(3):653–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/3/653.

Sahin:2004:DRA

[SWK^{+}04] Umut Sahin, Gisela Weskamp, Kristine Kelly, Hong-Ming Zhou, Shigeki Higashiyama, Jacques Peschon, Dieter Hartmann, Paul Saftig, and Carl P. Blobel. Distinct roles for ADAM10 and ADAM17 in ectodomain shedding of six EGFR ligands. Journal of Cell Biology, 164(5):769–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/5/
769.

Suzuki:2002:CIC

[SwZK^{+}02] Atsushi Suzuki, Yun wen Zheng, Shin Kaneko, Masafumi Onodera, Katashi Fukao, Hiromitsu Nakauchi, and Hideki Taniguchi. Clonal identification and characterization of selfrenewing pluripotent stem cells in the developing liver. Journal of Cell Biology, 156(1):173–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/1/173.

Sztalryd:2003:PET

[SXD^{+}03] Carole Sztalryd, Guoheng Xu, Heidi Dorward, John T. Tansey, Juan A. Contreras, Alan R. Kimmel, and Constantine Londos. Perilipin A is essential for the translocation of hormone-sensitive lipase during lipolytic activation. Journal of Cell Biology, 161(6):1093–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/6/1093.

Suzuki:2001:APK

[SYH^{+}01] Atsushi Suzuki, Tomoyuki Yamanaka, Tomonori Hirose, Naoyuki Manabe, Keiko Mizuno, Miki Shimizu, Kazunori Akimoto, Yasushi Izumi, Tetsuo Ohnishi, and Shigeo Ohno. Atypical Protein Kinase C Is Involved in the Evolutionarily Conserved Par Protein Complex and Plays a Critical Role in Establishing Epithelia–Specific Junctional Structures. Journal of Cell Biology, 152(6):1183–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/6/1183.

Sun:2003:HRE

[SYVB03] Wei Sun, Qing Yan, Thomas A. Vida, and Andrew J. Bean. Hrs regulates early endosome fusion by inhibiting formation of an endosomal SNARE complex. Journal of Cell Biology, 162(1):125–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/1/125.

Scholefield:2003:HSR

[SYW^{+}03] Zoe Scholefield, Edwin A. Yates, Gareth Wayne, Augustin Amour, William McDowell, and Jeremy E. Turnbull. Heparan sulfate regulates amyloid precursor protein processing by BACE1, the Alzheimer’s β-secretase. Journal of Cell Biology, 163(1):97–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/1/97.

Schubert:2003:RTF

[SYY^{+}03] William Schubert, Xiao Yong Yang, Teddy T. C. Yang, Stephen M. Factor, Michael P. Lisanti, Jeffery D. Molkentin, Mercedes Rincón, and Chi-Wing Chow. Requirement of transcription factor NFAT in developing atrial myocardium. Journal of Cell Biology, 161(5):861–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/5/861.

Subramaniam:2004:C

[SZvBuH^{+}04a] Srinivasa Subramaniam, Ute Zirrgiebel, Oliver von Bohlen und Halbach, Jens Strelau, Christine Laliberté, David R. Kaplan, and Klaus Unsicker. Correction. Journal of Cell Biology, 165 (4):593–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/4/593.

Subramaniam:2004:EAP

[SZvBuH^{+}04b] Srinivasa Subramaniam, Ute Zirrgiebel, Oliver von Bohlen und Halbach, Jens Strelau, Christine Laliberté, David R. Kaplan, and Klaus Unsicker. ERK activation promotes neuronal degeneration predominantly through plasma membrane damage and independently of caspase-3. Journal of Cell Biology, 165(3):357–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/3/357.

Shin:2000:MSC

[SZZ^{+}00] Dong Min Shin, Xiao-Song Zhao, Weizhong Zeng, Marina Mozhayeva, and Shmuel Muallem. The mammalian Sec6/8 complex interacts with Ca^{2+} signaling complexes and regulates their activity. Journal of Cell Biology, 150(5):1101– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/5/1101.

Tamaki:2002:IME

[TAA^{+}02] Tetsuro Tamaki, Akira Akatsuka, Kiyoshi Ando, Yoshihiko Nakamura, Hideyuki Matsuzawa, Tomomitsu Hotta, Roland R. Roy, and V. Reggie Edgerton. Identification of myogenic-endothelial progenitor cells in the interstitial spaces of skeletal muscle. Journal of Cell Biology, 157(4):571–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
157/4/571.

Takito:2004:CCC

[TAA04] Jiro Takito and Qais Al-Awqati. Conversion of ES cells to columnar epithelia by hensin and to squamous epithelia by laminin. Journal of Cell Biology, 166(7):1093–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
7/1093.

Thoreson:2000:SUP

[TAD^{+}00] Molly A. Thoreson, Panos Z. Anastasiadis, Juliet M. Daniel, Reneé C. Ireton, Margaret J. Wheelock, Keith R. Johnson, Diana K. Hummingbird, and Albert B. Reynolds. Selective Uncoupling of P120^{ctn} from E–Cadherin Disrupts Strong Adhesion. Journal of Cell Biology, 148(1):189–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
1/189.

Taylor:2000:KP

[TB00a] Edwin W. Taylor and Gary G. Borisy. Kinesin Processivity. Journal of Cell Biology, 151(5):??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/5/F27.

Tirnauer:2000:EPR

[TB00b] Jennifer S. Tirnauer and Barbara E. Bierer. Eb1 Proteins Regulate Microtubule Dynamics, Cell Polarity, and Chromosome Stability. Journal of Cell Biology, 149(4):761–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
4/761.

Tucker:2001:RCB

[TBJ^{+}01] Karen E. Tucker, Maria Teresa Berciano, Erica Y. Jacobs, David F. LePage, Karl B. Shpargel, Jennifer J. Rossire, Edward K. L. Chan, Miguel Lafarga, Ronald A. Conlon, and A. Gregory Matera. Residual Cajal bodies in coilin knockout mice fail to recruit Sm snRNPs and SMN, the spinal muscular atrophy gene product. Journal of Cell Biology, 154 (2):293–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/2/293.

Topper:2001:ELC

[TBRG01] Leana M. Topper, Holger Bastians, Joan V. Ruderman, and Gary J. Gorbsky. Elevating the level of Cdc34/Ubc3 ubiquitin-conjugating enzyme in mitosis inhibits association of CENP-E with kinetochores and blocks the metaphase alignment of chromosomes. Journal of Cell Biology, 154(4): 707–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/4/707.

Tachikawa:2001:GGP

[TBTN01] Hiroyuki Tachikawa, Andrew Bloecher, Kelly Tatchell, and Aaron M. Neiman. A Gip1p–Glc7p phosphatase complex regulates septin organization and spore wall formation. Journal of Cell Biology, 155(5):797–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/5/797.

Tuzon:2004:FYH

[TBW^{+}04] Creighton T. Tuzon, Britta Borgstrom, Dietmar Weilguny, Richard Egel, Julia Promisel Cooper, and Olaf Nielsen. The fission yeast heterochromatin protein Rik1 is required for telomere clustering during meiosis. Journal of Cell Biology, 165(6):759–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/6/759.

Thompson:2000:FFM

[TCH^{+}00] Terri G. Thompson, Yiu-Mo Chan, Andrew A. Hack, Melissa Brosius, Michael Rajala, Hart G. W. Lidov, Elizabeth M. McNally, Simon Watkins, and Louis M. Kunkel. Filamin 2 (FLN2): a muscle-specific sarcoglycan interacting protein. Journal of Cell Biology, 148(1):115–??, January 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/1/
115.

Tavormina:2002:REM

[TCH^{+}02] Penny A. Tavormina, Marie-George Côme, Joanna R. Hudson, Yin-Yuan Mo, William T. Beck, and Gary J. Gorbsky. Rapid exchange of mammalian topoisomerase IIα at kinetochores and chromosome arms in mitosis. Journal of Cell Biology, 158(1):23–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/1/23.

Tomita:2003:FSD

[TCK^{+}03] Susumu Tomita, Lu Chen, Yoshimi Kawasaki, Ronald S. Petralia, Robert J. Wenthold, Roger A. Nicoll, and David S. Bredt. Functional studies and distribution define a family of transmembrane AMPA receptor regulatory proteins. Journal of Cell Biology, 161(4):805–??, May 2003. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/4/805.

Torrente:2003:IPP

[TCP^{+}03] Yvan Torrente, Geoffrey Camirand, Federica Pisati, Marzia Belicchi, Barbara Rossi, Fabio Colombo, Mosthapha El Fahime, Nicolas J. Caron, Andrew C. Issekutz, Gabriela Constantin, Jacques P. Tremblay, and Nereo Bresolin. Identification of a putative pathway for the muscle homing of stem cells in a muscular dystrophy model. Journal of Cell Biology, 162(3):511–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/3/511.

Titorenko:2000:FSP

[TCR00] Vladimir I. Titorenko, Honey Chan, and Richard A. Rachubinski. Fusion of Small Peroxisomal Vesicles in Vitro Reconstructs an Early Step in the in Vivo Multistep Peroxisome Assembly Pathway of Yarrowia lipolytica. Journal of Cell Biology, 148(1):29–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/1/29.

Tai:2001:CDR

[TCS01] Andrew W. Tai, Jen-Zen Chuang, and Ching-Hwa Sung. Cytoplasmic Dynein Regulation by Subunit Heterogeneity and Its Role in Apical Transport. Journal of Cell Biology, 153 (7):1499–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/7/1499.

Tilney:2000:RAF

[TCV^{+}00] Lewis G. Tilney, Patricia S. Connelly, Kelly A. Vranich, Michael K. Shaw, and Gregory M. Guild. Regulation of actin filament cross-linking and bundle shape in Drosophila bristles. Journal of Cell Biology, 148(1):87–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/1/
87.

Tourriere:2003:RAE

[TCZ^{+}03] Hélène Tourrière, Karim Chebli, Latifa Zekri, Brice Courselaud, Jean Marie Blanchard, Edouard Bertrand, and Jamal Tazi. The RasGAP-associated endoribonuclease G3BP assembles stress granules. Journal of Cell Biology, 160(6): 823–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/6/823.

Tai:2002:RDD

[TDFV02] Chin-Yin Tai, Denis L. Dujardin, Nicole E. Faulkner, and Richard B. Vallee. Role of dynein, dynactin, and CLIP-170 interactions in LIS1 kinetochore function. Journal of Cell Biology, 156(6):959–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/6/959.

Tan:2002:CST

[TDhL^{+}02] Kemin Tan, Mark Duquette, Jin huan Liu, Yicheng Dong, Rongguang Zhang, Andrzej Joachimiak, Jack Lawler, and Jia huai Wang. Crystal structure of the TSP-1 type 1 repeats. Journal of Cell Biology, 159(2):373–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/2/
373.

Tam:2003:DFA

[TDL03] Lai-Wa Tam, William L. Dentler, and Paul A. Lefebvre. Defective flagellar assembly and length regulation in LF3 null mutants in Chlamydomonas. Journal of Cell Biology, 163(3): 597–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/3/597.

Tomishima:2001:CHP

[TE01] M. J. Tomishima and L. W. Enquist. A conserved αherpesvirus protein necessary for axonal localization of viral membrane proteins. Journal of Cell Biology, 154(4): 741–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/4/741.

Tucker:2003:ISE

[TEB^{+}03] Ward C. Tucker, J. Michael Edwardson, Jihong Bai, Hyun-Jung Kim, Thomas F. J. Martin, and Edwin R. Chapman. Identification of synaptotagmin effectors via acute inhibition of secretion from cracked PC12 cells. Journal of Cell Biology, 162(2):199–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/2/199.

Tunquist:2003:SCP

[TEC^{+}03] Brian J. Tunquist, Patrick A. Eyers, Lin G. Chen, Andrea L. Lewellyn, and James L. Maller. Spindle checkpoint proteins Mad1 and Mad2 are required for cytostatic factor– mediated metaphase arrest. Journal of Cell Biology, 163(6): 1231–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/6/1231.

Tsukita:2000:PW

[TF00] Shoichiro Tsukita and Mikio Furuse. Pores in the Wall. Journal of Cell Biology, 149(1):13–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/1/13.

Treves:2004:JKE

[TFAM^{+}04] Susan Treves, Clara Franzini-Armstrong, Luca Moccagatta, Christophe Arnoult, Cristiano Grasso, Adam Schrum, Sylvie Ducreux, Michael X. Zhu, Katsuhiko Mikoshiba, Thierry Girard, Sophia Smida-Rezgui, Michel Ronjat, and Francesco Zorzato. Junctate is a key element in calcium entry induced by activation of InsP_{3} receptors and/or calcium store depletion. Journal of Cell Biology, 166(4):537–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
4/537.

Tchernychev:2003:PME

[TFF03] Boris Tchernychev, Bruce Furie, and Barbara C. Furie. Peritoneal macrophages express both P-selectin and PSGL-1. Journal of Cell Biology, 163(5):1145–??, December 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/163/5/
1145.

Tomas:2004:ARM

[TFM04] Alejandra Tomas, Clare Futter, and Stephen E. Moss. Annexin 11 is required for midbody formation and completion of the terminal phase of cytokinesis. Journal of Cell Biology, 165 (6):813–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/6/813.

Traka:2003:ATC

[TGD^{+}03] Maria Traka, Laurence Goutebroze, Natalia Denisenko, Maria Bessa, Artemisia Nifli, Sophia Havaki, Yoichiro Iwakura, Fumihiko Fukamauchi, Kazutada Watanabe, Betty Soliven, Jean-Antoine Girault, and Domna Karagogeos. Association of TAG-1 with Caspr2 is essential for the molecular organization of juxtaparanodal regions of myelinated fibers. Journal of Cell Biology, 162(6):1161–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
6/1161.

Thome:2001:EHP

[TGM^{+}01] Margot Thome, Olivier Gaide, Olivier Micheau, Fabio Martinon, David Bonnet, Montserrat Gonzalez, and Jürg Tschopp. Equine Herpesvirus Protein E10 Induces Membrane Recruitment and Phosphorylation of Its Cellular Homologue, Bcl-10. Journal of Cell Biology, 152(5):1115–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/152/5/
1115.

Tait:2000:OCA

[TGMC^{+}00] Steven Tait, Frank Gunn-Moore, J. Martin Collinson, Jeffery Huang, Catherine Lubetzki, Liliana Pedraza, Diane L. Sherman, David R. Colman, and Peter J. Brophy. An Oligodendrocyte Cell Adhesion Molecule at the Site of Assembly of the Paranodal Axo–Glial Junction. Journal of Cell Biology, 150(3):657–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/3/657.

Trombetta:2000:CRG

[TH00] E. Sergio Trombetta and Ari Helenius. Conformational Requirements for Glycoprotein Reglucosylation in the Endoplasmic Reticulum. Journal of Cell Biology, 148(6):1123–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/148/6/1123.

Tokumaru:2000:ESE

[THE^{+}00] Sho Tokumaru, Shigeki Higashiyama, Takeshi Endo, Takatoshi Nakagawa, Jun ichiro Miyagawa, Katsumi Yamamori, Yasushi Hanakawa, Hiroshi Ohmoto, Kohichiro Yoshino, Yuji Shirakata, Yuji Matsuzawa, Koji Hashimoto, and Naoyuki Taniguchi. Ectodomain Shedding of Epidermal Growth Factor Receptor Ligands Is Required for Keratinocyte Migration in Cutaneous Wound Healing. Journal of Cell Biology, 151(2):209–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/2/209.

Tsurumi:2004:SAC

[THG^{+}04] Chizuko Tsurumi, Steffen Hoffmann, Stephan Geley, Ralph Graeser, and Zbigniew Polanski. The spindle assembly checkpoint is not essential for CSF arrest of mouse oocytes. Journal of Cell Biology, 167(6):1037–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/6/1037.

Toyofuku:2000:WFS

[THK^{+}00] Toshihiko Toyofuku, Zhang Hong, Tsunehiko Kuzuya, Michihiko Tada, and Masatsugu Hori. WNT/ frizzled-2 Signaling Induces Aggregation and Adhesion among Cardiac Myocytes by Increased Cadherin-β-Catenin Complex. Journal of Cell Biology, 150(1):225–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/1/225.

Takahashi:2004:SEI

[THO^{+}04] Noriko Takahashi, Hiroyasu Hatakeyama, Haruo Okado, Akiko Miwa, Takuya Kishimoto, Tatsuya Kojima, Teruo Abe, and Haruo Kasai. Sequential exocytosis of insulin granules is associated with redistribution of SNAP25. Journal of Cell Biology, 165(2):255–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/2/255.

Tu:2001:NFA

[THZ^{+}01] Yizeng Tu, Yao Huang, Yongjun Zhang, Yun Hua, and Chuanyue Wu. A New Focal Adhesion Protein That Interacts with Integrin–Linked Kinase and Regulates Cell Adhesion and Spreading. Journal of Cell Biology, 153(3):585–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
3/585.

Tsuji:2002:RMA

[TIO^{+}02] Takahiro Tsuji, Toshimasa Ishizaki, Muneo Okamoto, Chiharu Higashida, Kazuhiro Kimura, Tomoyuki Furuyashiki, Yoshiki Arakawa, Raymond B. Birge, Tetsuya Nakamoto, Hisamaru Hirai, and Shuh Narumiya. ROCK and mDia1 antagonize in Rho-dependent Rac activation in Swiss 3T3 fibroblasts. Journal of Cell Biology, 157(5):819–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/157/5/
819.

Taya:2001:DII

[TIS^{+}01] Shinichiro Taya, Naoyuki Inagaki, Hiroaki Sengiku, Hiroshi Makino, Akihiro Iwamatsu, Itaru Urakawa, Kenji Nagao, Shiro Kataoka, and Kozo Kaibuchi. Direct interaction of insulin-like growth factor-1 receptor with leukemia-associated RhoGEF. Journal of Cell Biology, 155(5):809–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
5/809.

Topol:2003:WIC

[TJC^{+}03] Lilia Topol, Xueyuan Jiang, Hosoon Choi, Lisa Garrett-Beal, Peter J. Carolan, and Yingzi Yang. Wnt-5a inhibits the canonical Wnt pathway by promoting GSK-3–independent β-catenin degradation. Journal of Cell Biology, 162(5):899– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/5/899.

Thomas:2004:TOD

[TJS^{+}04] David M. Thomas, Sandra A. Johnson, Natalie A. Sims, Melanie K. Trivett, John L. Slavin, Brian P. Rubin, Paul Waring, Grant A. McArthur, Carl R. Walkley, Andrew J. Holloway, Dileepa Diyagama, Jonathon E. Grim, Bruce E. Clurman, David D. L. Bowtell, Jong-Seo Lee, Gabriel M. Gutierrez, Denise M. Piscopo, Shannon A. Carty, and Philip W. Hinds. Terminal osteoblast differentiation, mediated by runx2 and p27 ^{KIP1}, is disrupted in osteosarcoma.

Journal of Cell Biology, 167(5):925–??, December 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/167/5/
925.

Trapp:2000:AGS

[TK00] Bruce D. Trapp and Grahame J. Kidd. Axo-glial Septate Junctions. Journal of Cell Biology, 150(3):1–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/3/
F97.

Toivola:2004:KMC

[TKB^{+}04] Diana M. Toivola, Selvi Krishnan, Henry J. Binder, Satish K. Singh, and M. Bishr Omary. Keratins modulate colonocyte electrolyte transport via protein mistargeting. Journal of Cell Biology, 164(6):911–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/6/911.

Tsou:2003:PDG

[TKHR03] Meng-Fu Bryan Tsou, Wei Ku, Adam Hayashi, and Lesilee S. Rose. PAR-dependent and geometry-dependent mechanisms of spindle positioning. Journal of Cell Biology, 160(6): 845–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/6/845.

Tsujimura:2002:PLM

[TKS^{+}02] Akira Tsujimura, Yasuhiro Koikawa, Sarah Salm, Tetsuya Takao, Sandra Coetzee, David Moscatelli, Ellen Shapiro, Herbert Lepor, Tung-Tien Sun, and E. Lynette Wilson. Proximal location of mouse prostate epithelial stem cells. Journal of Cell Biology, 157(7):1257–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/7/1257.

Troyanovsky:2001:A

[TLM^{+}01] Boris Troyanovsky, Tetyana Levchenko, Göran Månsson, Olga Matvijenko, and Lars Holmgren. Angiomotin. Journal of Cell Biology, 152(6):1247–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/6/1247.

TenBroek:2001:SCU

[TLS^{+}01] Erica M. TenBroek, Paul D. Lampe, Joell L. Solan, James K. Reynhout, and Ross G. Johnson. Ser364 of connexin43 and the upregulation of gap junction assembly by cAMP. Journal of Cell Biology, 155(7):1307–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/7/1307.

Takemaru:2000:TCC

[TM00] Ken-Ichi Takemaru and Randall T. Moon. The Transcriptional Coactivator Cbp Interacts with β-Catenin to Activate Gene Expression. Journal of Cell Biology, 149(2):249–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
149/2/249.

Tyska:2004:RML

[TM04] Matthew J. Tyska and Mark S. Mooseker. A role for myosin-1A in the localization of a brush border disaccharidase. Journal of Cell Biology, 165(3):395–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/3/395.

Trajkovski:2004:NTI

[TMA^{+}04] Mirko Trajkovski, Hassan Mziaut, Anke Altkrüger, Joke Ouwendijk, Klaus-Peter Knoch, Stefan Müller, and Michele Solimena. Nuclear translocation of an ICA512 cytosolic fragment couples granule exocytosis and insulin expression in β-cells. Journal of Cell Biology, 167(6):1063–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
6/1063.

Tran:2001:MNP

[TMD^{+}01] P. T. Tran, L. Marsh, V. Doye, S. Inoué, and F. Chang. A Mechanism for Nuclear Positioning in Fission Yeast Based on Microtubule Pushing. Journal of Cell Biology, 153(2):397–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/2/397.

Tasto:2003:AHM

[TMG03] Joseph J. Tasto, Jennifer L. Morrell, and Kathleen L. Gould. An anillin homologue, Mid2p, acts during fission yeast cytokinesis to organize the septin ring and promote cell separation. Journal of Cell Biology, 160(7):1093–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/7/
1093.

Tam:2000:IOS

[TMHP00] Beatrice M. Tam, Orson L. Moritz, Lawrence B. Hurd, and David S. Papermaster. Identification of an Outer Segment Targeting Signal in the Cooh Terminus of Rhodopsin Using Transgenic Xenopus laevis. Journal of Cell Biology, 151(7): 1369–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/7/1369.

Tiffany:2000:PSE

[TMK^{+}00] Amanda M. Tiffany, Louis N. Manganas, Eunjoon Kim, Yi-Ping Hsueh, Morgan Sheng, and James S. Trimmer. PSD-95 and SAP97 Exhibit Distinct Mechanisms for Regulating K^{+} Channel Surface Expression and Clustering. Journal of Cell Biology, 148(1):147–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/148/1/147.

Tieu:2000:MWR

[TN00a] Quinton Tieu and Jodi Nunnari. Mdv1p Is a Wd Repeat Protein That Interacts with the Dynamin–Related Gtpase, Dnm1p, to Trigger Mitochondrial Division. Journal of Cell Biology, 151(2):353–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/2/353.

Torres:2000:CRD

[TN00b] Monica A. Torres and W. James Nelson. Colocalization and Redistribution of Dishevelled and Actin during WNT– Induced Mesenchymal Morphogenesis. Journal of Cell Biology, 149(7):1433–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/7/1433.

Tomita:2001:PPI

[TNB01] Susumu Tomita, Roger A. Nicoll, and David S. Bredt. PDZ Protein Interactions Regulating Glutamate Receptor Function and Plasticity. Journal of Cell Biology, 153(5):0–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
5/F19.

Tuma:2001:VDR

[TNBH01] Pamela L. Tuma, Lydia K. Nyasae, Jonathan M. Backer, and Ann L. Hubbard. Vps34p differentially regulates endocytosis from the apical and basolateral domains in polarized hepatic cells. Journal of Cell Biology, 154(6):1197–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/154/
6/1197.

Tian:2000:IAM

[TNK^{+}00] Li Tian, Henrietta Nyman, Patrick Kilgannon, Yoshihiro Yoshihara, Kensaku Mori, Leif C. Andersson, Sami Kaukinen, Heikki Rauvala, W. Michael Gallatin, and Carl G. Gahmberg. Intercellular Adhesion Molecule-5 Induces Dendritic Outgrowth by Homophilic Adhesion. Journal of Cell Biology, 150(1):243–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/1/243.

Tachibana:2000:TCA

[TNM^{+}00] Kouichi Tachibana, Hiroyuki Nakanishi, Kenji Mandai, Kumi Ozaki, Wataru Ikeda, Yasunori Yamamoto, Akira Nagafuchi, Shoichiro Tsukita, and Yoshimi Takai. Two Cell Adhesion Molecules, Nectin and Cadherin, Interact through Their Cytoplasmic Domain–Associated Proteins. Journal of Cell Biology, 150(5):1161–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/5/1161.

Tanimura:2003:DCM

[TNM^{+}03] Natsuko Tanimura, Masakazu Nagafuku, Yasuko Minaki, Yukio Umeda, Fumie Hayashi, Junko Sakakura, Akiko Kato, Douglas R. Liddicoat, Masato Ogata, Toshiyuki Hamaoka, and Atsushi Kosugi. Dynamic changes in the mobility of LAT in aggregated lipid rafts upon T cell activation. Journal of Cell Biology, 160(1):125–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/1/125.

Takemoto:2003:STA

[TNMM03] Kiwamu Takemoto, Takeharu Nagai, Atsushi Miyawaki, and Masayuki Miura. Spatio-temporal activation of caspase revealed by indicator that is insensitive to environmental effects. Journal of Cell Biology, 160(2):235–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/2/
235.

Titorenko:2002:ACO

[TNW^{+}02] Vladimir I. Titorenko, Jean-Marc Nicaud, Huijie Wang, Honey Chan, and Richard A. Rachubinski. Acyl-CoA oxidase is imported as a heteropentameric, cofactor-containing complex into peroxisomes of Yarrowia lipolytica. Journal of Cell Biology, 156(3):481–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/3/481.

Tamai:2000:CMP

[ToIB^{+}00] Yoshitaka Tamai, Tomo o Ishikawa, Michael R. Bösl, Masahiko Mori, Masami Nozaki, Heléne Baribault, Robert G. Oshima, and Makoto M. Taketo. Cytokeratins 8 and 19 in the Mouse Placental Development. Journal of Cell Biology, 151(3):563–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/3/563.

Toyooka:2000:MTP

[TOM00] Kiminori Toyooka, Takashi Okamoto, and Takao Minamikawa. Mass Transport of Proform of a Kdel–Tailed Cysteine Proteinase (Sh–EP) to Protein Storage Vacuoles by Endoplasmic Reticulum–Derived Vesicle Is Involved in Protein Mobilization in Germinating Seeds. Journal of Cell Biology, 148(3):453–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/148/3/453.

Toyooka:2001:CCV

[TOM01] Kiminori Toyooka, Takashi Okamoto, and Takao Minamikawa. Cotyledon cells of Vigna mungo seedlings use at least two distinct autophagic machineries for degradation of starch granules and cellular components. Journal of Cell Biology, 154(5):973–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/5/973.

Tieu:2002:WRP

[TONN02] Quinton Tieu, Voytek Okreglak, Kari Naylor, and Jodi Nunnari. The WD repeat protein, Mdv1p, functions as a molecular adaptor by interacting with Dnm1p and Fis1p during mitochondrial fission. Journal of Cell Biology, 158(3): 445–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/3/445.

Tateishi:2001:NSE

[TOTC01] Keisuke Tateishi, Masao Omata, Keiji Tanaka, and Tomoki Chiba. The NEDD8 system is essential for cell cycle progression and morphogenetic pathway in mice. Journal of Cell Biology, 155(4):571–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/4/571.

Topham:2001:DKR

[TP01] Matthew K. Topham and Stephen M. Prescott. Diacylglycerol Kinase ζ Regulates Ras Activation by a Novel Mechanism. Journal of Cell Biology, 152(6):1135–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/152/6/
1135.

Terpstra:2003:RCP

[TPA^{+}03] Leonieke Terpstra, Josée Prud’homme, Alice Arabian, Shu Takeda, Gérard Karsenty, Shoukat Dedhar, and René St-Arnaud. Reduced chondrocyte proliferation and chondrodysplasia in mice lacking the integrin-linked kinase in chondrocytes. Journal of Cell Biology, 162(1):139–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/162/1/

Toman:2004:DTS

[TPW^{+}04] Rachelle E. Toman, Shawn G. Payne, Kenneth R. Watterson, Michael Maceyka, Norman H. Lee, Sheldon Milstien, John W. Bigbee, and Sarah Spiegel. Differential transactivation of sphingosine-1-phosphate receptors modulates NGFinduced neurite extension. Journal of Cell Biology, 166(3): 381–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/3/381.

Titorenko:2000:PMF

[TR00] Vladimir I. Titorenko and Richard A. Rachubinski. Peroxisomal Membrane Fusion Requires Two Aaa Family Atpases, Pex1p and Pex6p. Journal of Cell Biology, 150(4): 881–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/4/881.

Tsai:2002:UCT

[TR02] Billy Tsai and Tom A. Rapoport. Unfolded cholera toxin is transferred to the ER membrane and released from protein disulfide isomerase upon oxidation by Ero1. Journal of Cell Biology, 159(2):207–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/2/207.

Titorenko:2004:C

[TR04a] Vladimir I. Titorenko and Richard A. Rachubinski. Correction. Journal of Cell Biology, 165(1):155–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/165/1/
155.

Titorenko:2004:P

[TR04b] Vladimir I. Titorenko and Richard A. Rachubinski. The peroxisome. Journal of Cell Biology, 164(5):641–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
5/641.

Traub:2003:SI

[Tra03] Linton M. Traub. Sorting it out. Journal of Cell Biology, 163(2):203–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/2/203.

Taunton:2000:ADP

[TRC^{+}00] Jack Taunton, Brian A. Rowning, Margaret L. Coughlin, Michael Wu, Randall T. Moon, Timothy J. Mitchison, and Carolyn A. Larabell. Actin-dependent propulsion of endosomes and lysosomes by recruitment of N-Wasp. Journal of Cell Biology, 148(3):519–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/3/519.

Takao-Rikitsu:2004:PFI

[TRMI^{+}04] Etsuko Takao-Rikitsu, Sumiko Mochida, Eiji Inoue, Maki Deguchi-Tawarada, Marie Inoue, Toshihisa Ohtsuka, and Yoshimi Takai. Physical and functional interaction of the active zone proteins, CAST, RIM1, and Bassoon, in neurotransmitter release. Journal of Cell Biology, 164(2):301–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
164/2/301.

Tepper:2000:SHC

[TRW^{+}00] Annemiek D. Tepper, Paula Ruurs, Therese Wiedmer, Peter J. Sims, Jannie Borst, and Wim J. van Blitterswijk. Sphingomyelin Hydrolysis to Ceramide during the Execution Phase of Apoptosis Results from Phospholipid Scrambling and Alters Cell–Surface Morphology. Journal of Cell Biology, 150(1):155–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/1/155.

Trelles-Sticken:2000:MTP

[TSDS00] Edgar Trelles-Sticken, Michael E. Dresser, and Harry Scherthan. Meiotic Telomere Protein Ndj1p Is Required for Meiosis–Specific Telomere Distribution, Bouquet Formation and Efficient Homologue Pairing. Journal of Cell Biology, 151 (1):95–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/1/95.

Tanaka:2004:LDF

[TSH^{+}04] Teruyuki Tanaka, Finley F. Serneo, Christine Higgins, Michael J. Gambello, Anthony Wynshaw-Boris, and Joseph G. Gleeson. Lis1 and doublecortin function with dynein to mediate coupling of the nucleus to the centrosome in neuronal migration. Journal of Cell Biology, 165(5):709–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/165/5/
709.

Toomre:2000:FCM

[TSK^{+}00] Derek Toomre, Jürgen A. Steyer, Patrick Keller, Wolfhard Almers, and Kai Simons. Fusion of Constitutive Membrane Traffic with the Cell Surface Observed by Evanescent Wave Microscopy. Journal of Cell Biology, 149(1):33–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
1/33.

Tomizawa:2003:CAD

[TSL^{+}03] Kazuhito Tomizawa, Satoshi Sunada, Yun-Fei Lu, Yoshiya Oda, Masahiro Kinuta, Toshio Ohshima, Taro Saito, Fan-Yan Wei, Masayuki Matsushita, Sheng-Tian Li, Kimiko Tsutsui, Shin ichi Hisanaga, Katsuhiko Mikoshiba, Kohji Takei, and Hideki Matsui. Cophosphorylation of amphiphysin I and dynamin I by Cdk5 regulates clathrinmediated endocytosis of synaptic vesicles. Journal of Cell Biology, 163(4):813–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/4/813.

Tam:2004:SHF

[TSL04] Rose Tam, Kelly P. Smith, and Jeanne B. Lawrence. The 4q subtelomere harboring the FSHD locus is specifically anchored with peripheral heterochromatin unlike most human telomeres. Journal of Cell Biology, 167(2):269–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
2/269.

Tan:2001:RAM

[TSMS01] Shirlee Tan, Nikunj Somia, Pamela Maher, and David Schubert. Regulation of Antioxidant Metabolism by Translation Initiation Factor 2α. Journal of Cell Biology, 152(5): 997–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/5/997.

Tanentzapf:2000:ALB

[TSMT00] Guy Tanentzapf, Christian Smith, Jane McGlade, and Ulrich Tepass. Apical, Lateral, and Basal Polarization Cues Contribute to the Development of the Follicular Epithelium during Drosophila Oogenesis. Journal of Cell Biology, 151(4): 891–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/4/891.

Trapani:2003:CDM

[TST^{+}03] Joseph A. Trapani, Vivien R. Sutton, Kevin Y. T. Thia, Yu Qin Li, Christopher J. Froelich, David A. Jans, Mauro S. Sandrin, and Kylie A. Browne. A clathrin/dynaminand mannose-6-phosphate receptor–independent pathway for granzyme B–induced cell death. Journal of Cell Biology, 160(2):223–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/2/223.

Tsubouchi:2002:LSR

[TSY^{+}02] Asako Tsubouchi, Junko Sakakura, Ryohei Yagi, Yuichi Mazaki, Erik Schaefer, Hajime Yano, and Hisataka Sabe. Localized suppression of RhoA activity by Tyr31/118phosphorylated paxillin in cell adhesion and migration. Journal of Cell Biology, 159(4):673–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/4/673.

Tanoue:2004:MFC

[TT04] Takuji Tanoue and Masatoshi Takeichi. Mammalian Fat1 cadherin regulates actin dynamics and cell–cell contact. Journal of Cell Biology, 165(4):517–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/165/4/517.

Teng:2001:SEM

[TTH^{+}01] Junlin Teng, Yosuke Takei, Akihiro Harada, Takao Nakata, Jianguo Chen, and Nobutaka Hirokawa. Synergistic effects of MAP2 and MAP1B knockout in neuronal migration, dendritic outgrowth, and microtubule organization. Journal of Cell Biology, 155(1):65–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/1/65.

Takei:2000:DAE

[TTHH00] Yosuke Takei, Junlin Teng, Akihiro Harada, and Nobutaka Hirokawa. Defects in Axonal Elongation and Neuronal Migration in Mice with Disrupted tau and map1b Genes. Journal of Cell Biology, 150(5):989–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/5/989.

Takeda:2003:TCC

[TTM^{+}03] Yoshito Takeda, Isao Tachibana, Kenji Miyado, Masatoshi Kobayashi, Toru Miyazaki, Toshiki Funakoshi, Hiromi Kimura, Hiroyuki Yamane, Yoshiyuki Saito, Hiroyuki Goto, Tsutomu Yoneda, Mitsuhiro Yoshida, Toru Kumagai, Tadashi Osaki, Seiji Hayashi, Ichiro Kawase, and Eisuke Mekada. Tetraspanins CD9 and CD81 function to prevent the fusion of mononuclear phagocytes. Journal of Cell Biology, 161(5):945–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/5/945.

Torrente:2001:IIM

[TTP^{+}01] Yuan Torrente, Jacques-P Tremblay, Federica Pisati, Marzia Belicchi, Barbara Rossi, Manuela Sironi, Franco Fortunato, Mostafa El Fahime, Maria Grazia D’Angelo, Nicolas J. Caron, Gabriela Constantin, Denise Paulin, Guglielmo Scarlato, and Nereo Bresolin. Intraarterial Injection of Muscle– Derived Cd34^{+} Sca-1^{+} Stem Cells Restores Dystrophin in mdx Mice. Journal of Cell Biology, 152(2):335–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
2/335.

Tang:2000:LTC

[TTR00] Dean G. Tang, Yasuhito M. Tokumoto, and Martin C. Raff. Long-term Culture of Purified Postnatal Oligodendrocyte Precursor Cells. Journal of Cell Biology, 148(5): 971–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/5/971.

Taylor:2000:IAC

[TTS00] Kenneth A. Taylor, Dianne W. Taylor, and Fred Schachat. Isoforms of α-Actinin from Cardiac, Smooth, and Skeletal Muscle Form Polar Arrays of Actin Filaments. Journal of Cell Biology, 149(3):635–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/149/3/635.

Terada:2003:IAC

[TUK03] Yasuhiko Terada, Yumi Uetake, and Ryoko Kuriyama. Interaction of Aurora–A and centrosomin at the microtubulenucleating site in Drosophila and mammalian cells. Journal of Cell Biology, 162(5):757–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/5/757.

Tuma:2002:WLP

[Tum02a] Rabiya Tuma. And the winner is … the phosphatase! Journal of Cell Biology, 157(7):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/7/1103.2.

Tuma:2002:CBF

[Tum02b] Rabiya Tuma. Cajal body finds its function. Journal of Cell Biology, 157(7):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/7/1103.1.

Tuma:2002:GCN

[Tum02c] Rabiya Tuma. Gluing cells into the niche. Journal of Cell Biology, 157(7):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/7/1102.2.

Tuma:2002:MC

[Tum02d] Rabiya Tuma. A Molotov cocktail. Journal of Cell Biology, 157(7):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/7/1102.1.

Tuma:2002:NPC

[Tum02e] Rabiya Tuma. The nuclear pore as a chromatin boundary. Journal of Cell Biology, 157(7):??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/7/1103.3.

Tuma:2003:CTC

[Tum03a] Rabiya Tuma. Cilia take charge in L–R asymmetry. Journal of Cell Biology, 162(3):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/3/362.1.

Tuma:2003:ECC

[Tum03b] Rabiya Tuma. The elusive cell cycle–transcription link. Journal of Cell Biology, 162(3):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/3/363.2.

Tuma:2003:EOV

[Tum03c] Rabiya Tuma. Exocytosis — one vesicle at a time. Journal of Cell Biology, 162(3):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/3/363.1.

Tuma:2003:LRL

[Tum03d] Rabiya Tuma. Loss of Rb launches cell cycle. Journal of Cell Biology, 162(3):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/3/363.3.

Tuma:2003:NAD

[Tum03e] Rabiya Tuma. Network architecture determines stability. Journal of Cell Biology, 162(3):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/3/362.2.

Tuma:2003:CWM

[Tum03f] Rabiya S. Tuma. Calcium waves in membrane fusion. Journal of Cell Biology, 161(2):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/2/221.2.

Tuma:2003:CSW

[Tum03g] Rabiya S. Tuma. Conserving synaptic weight. Journal of Cell Biology, 161(2):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/220.2.

Tuma:2003:CLR

[Tum03h] Rabiya S. Tuma. Crm1 locks up replication factors. Journal of Cell Biology, 161(3):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/3/457.3.

Tuma:2003:MCV

[Tum03i] Rabiya S. Tuma. Math connects vascular network to the universe. Journal of Cell Biology, 161(3):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/3/457.1.

Tuma:2003:MMB

[Tum03j] Rabiya S. Tuma. MicroRNAs make big splash in apoptosis. Journal of Cell Biology, 161(3):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/3/456.1.

Tuma:2003:RCT

[Tum03k] Rabiya S. Tuma. Reconstituting a chloroplast transport system. Journal of Cell Biology, 161(2):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/2/221.1.

Tuma:2003:RYR

[Tum03l] Rabiya S. Tuma. ROS for your roots. Journal of Cell Biology, 161(2):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/2/220.1.

Tuma:2003:SAO

[Tum03m] Rabiya S. Tuma. Shared amyloid oligomer structure proves toxic. Journal of Cell Biology, 161(3):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/3/456.2.

Tuma:2003:TUT

[Tum03n] Rabiya S. Tuma. Toxoplasma uses tough love. Journal of Cell Biology, 161(2):??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/221.3.

Tuma:2003:TRR

[Tum03o] Rabiya S. Tuma. Turning RacGAP into a RhoGAP. Journal of Cell Biology, 161(3):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/3/457.2.

Tuma:2004:ABY

[Tum04a] Rabiya Tuma. Actin branches out in yeast. Journal of Cell Biology, 166(5):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/5/606.1.

Tuma:2004:AON

[Tum04b] Rabiya Tuma. Astrocytes orchestrate neuronal synchrony. Journal of Cell Biology, 166(7):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/7/941.3.

Tuma:2004:CLC

[Tum04c] Rabiya Tuma. Calcium lights up cell cycle. Journal of Cell Biology, 166(5):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/5/607.1.

Tuma:2004:CTS

[Tum04d] Rabiya Tuma. Centrosomes take sides. Journal of Cell Biology, 166(7):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/7/940.2.

Tuma:2004:CIB

[Tum04e] Rabiya Tuma. Clustering in integrin binding. Journal of Cell Biology, 167(6):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/6/992.3.

Tuma:2004:CDG

[Tum04f] Rabiya Tuma. Crouching dragon guards nascent proteins. Journal of Cell Biology, 166(7):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/7/941.1.

Tuma:2004:DCT

[Tum04g] Rabiya Tuma. Dendritic cells turn traitorous. Journal of Cell Biology, 166(7):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/7/940.1.

Tuma:2004:CT

[Tum04h] Rabiya Tuma. ES cells without teratomas. Journal of Cell Biology, 167(4):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/4/578.2.

Tuma:2004:FMD

[Tum04i] Rabiya Tuma. Flagellar microtubules do the twist. Journal of Cell Biology, 166(5):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/5/607.2.

Tuma:2004:IOM

[Tum04j] Rabiya Tuma. Invadolysin opens many doors. Journal of Cell Biology, 167(4):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/4/579.1.

Tuma:2004:LKM

[Tum04k] Rabiya Tuma. Less kinesin, more condensation. Journal of Cell Biology, 166(5):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/5/606.2.

Tuma:2004:MRO

[Tum04l] Rabiya Tuma. Making a raft with oligomers. Journal of Cell Biology, 167(4):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/4/579.2.

Tuma:2004:MI

[Tum04m] Rabiya Tuma. Measuring insulin. Journal of Cell Biology, 167(6):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/6/992.1.

Tuma:2004:MMC

[Tum04n] Rabiya Tuma. Mitochondria move in on calcium. Journal of Cell Biology, 167(4):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/4/578.1.

Tuma:2004:OCA

[Tum04o] Rabiya Tuma. One cyclin for adult neurogenesis. Journal of Cell Biology, 167(2):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/2/188.3.

Tuma:2004:OPT

[Tum04p] Rabiya Tuma. One protein, two pools. Journal of Cell Biology, 167(2):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/2/188.2.

Tuma:2004:PPA

[Tum04q] Rabiya Tuma. Parsing p53 activity. Journal of Cell Biology, 164(3):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/3/335.2.

Tuma:2004:PPP

[Tum04r] Rabiya Tuma. Pericentrin pivotal in primary cilia. Journal of Cell Biology, 166(5):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/5/607.3.

Tuma:2004:PTI

[Tum04s] Rabiya Tuma. PINCHing together integrins and signaling. Journal of Cell Biology, 167(6):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/6/992.2.

Tuma:2004:RGC

[Tum04t] Rabiya Tuma. Repopulating glioma cell lines. Journal of Cell Biology, 164(3):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/3/334.2.

Tuma:2004:RMD

[Tum04u] Rabiya Tuma. RNAi-mediated DNA silencing. Journal of Cell Biology, 164(3):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/3/335.3.

Tuma:2004:SMR

[Tum04v] Rabiya Tuma. A sensitive measure of repression. Journal of Cell Biology, 164(3):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/3/334.1.

Tuma:2004:SSG

[Tum04w] Rabiya Tuma. Sgo stabilizes glue and microtubules. Journal of Cell Biology, 166(7):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/7/941.2.

Tuma:2004:STM

[Tum04x] Rabiya Tuma. Slipping through the middle. Journal of Cell Biology, 167(2):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/2/189.2.

Tuma:2004:SHS

[Tum04y] Rabiya Tuma. Stick hard, signal fast. Journal of Cell Biology, 167(6):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/6/993.2.

Tuma:2004:SP

[Tum04z] Rabiya Tuma. Sticking to PIP_{2}. Journal of Cell Biology, 164 (3):??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/3/335.1.

Tuma:2004:TIN

[Tum04-27] Rabiya Tuma. Targeting the inner nuclear membrane. Journal of Cell Biology, 167(6):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/6/993.1.

Tuma:2004:TIL

[Tum04-28] Rabiya Tuma. Transport if loaded. Journal of Cell Biology, 167(4):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/4/578.3.

Tuma:2004:TPA

[Tum04-29] Rabiya Tuma. Tugging and pulling in asymmetric cell divisions. Journal of Cell Biology, 167(2):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/167/2/
189.1.

Tuma:2004:VBP

[Tum04-30] Rabiya Tuma. Vesicle biogenesis potentiation. Journal of Cell Biology, 167(2):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/2/188.1.

Thorn:2000:EPR

[TUV00] Kurt S. Thorn, Jeffrey A. Ubersax, and Ronald D. Vale. Engineering the Processive Run Length of the Kinesin Motor. Journal of Cell Biology, 151(5):1093–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/151/5/
1093.

Tomishige:2000:CKR

[TV00] Michio Tomishige and Ronald D. Vale. Controlling Kinesin by Reversible Disulfide Cross–Linking. Journal of Cell Biology, 151(5):1081–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/5/1081.

Truscott:2003:JPE

[TVF^{+}03] Kaye N. Truscott, Wolfgang Voos, Ann E. Frazier, Maria Lind, Yanfeng Li, Andreas Geissler, Jan Dudek, Hanne Müller, Albert Sickmann, Helmut E. Meyer, Chris Meisinger, Bernard Guiard, Peter Rehling, and Nikolaus Pfanner. A Jprotein is an essential subunit of the presequence translocase– associated protein import motor of mitochondria. Journal of Cell Biology, 163(4):707–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/4/707.

Tuma:2002:CMC

[TW02] Rabiya Tuma and William A. Wells. Cells and More Cells. Journal of Cell Biology, 156(2):220–??, January 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/156/2/
220.

Tu:2004:OPF

[TW04] Benjamin P. Tu and Jonathan S. Weissman. Oxidative protein folding in eukaryotes. Journal of Cell Biology, 164(3): 341–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/3/341.

Tiwari-Woodruff:2001:OCF

[TWBV^{+}01] Seema K. Tiwari-Woodruff, Alex G. Buznikov, Trung Q. Vu, Paul E. Micevych, Kendall Chen, Harley I. Kornblum, and Jeff M. Bronstein. Osp/ claudin-11 Forms a Complex with a Novel Member of the Tetraspanin Super Family and β1 Integrin and Regulates Proliferation and Migration of Oligodendrocytes. Journal of Cell Biology, 153(2):295–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
2/295.

Tashiro:2000:RAS

[TWS^{+}00a] Satoshi Tashiro, Joachim Walter, Akira Shinohara, Nanao Kamada, and Thomas Cremer. Rad51 Accumulation at Sites of DNA Damage and in Postreplicative Chromatin. Journal of Cell Biology, 150(2):283–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/150/2/283.

Tvrdik:2000:RNM

[TWS^{+}00b] Petr Tvrdik, Rolf Westerberg, Sandra Silve, Abolfazl Asadi, Andreas Jakobsson, Barbara Cannon, Gerard Loison, and Anders Jacobsson. Role of a New Mammalian Gene Family in the Biosynthesis of Very Long Chain Fatty Acids and Sphingolipids. Journal of Cell Biology, 149(3):707–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
3/707.

Totsukawa:2004:DRM

[TWS^{+}04] Go Totsukawa, Yue Wu, Yasuharu Sasaki, David J. Hartshorne, Yoshihiko Yamakita, Shigeko Yamashiro, and Fumio Matsumura. Distinct roles of MLCK and ROCK in the regulation of membrane protrusions and focal adhesion dynamics during cell migration of fibroblasts. Journal of Cell Biology, 164(3):427–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/3/427.

Tanaka:2002:CPR

[TYA^{+}02] Hiroyuki Tanaka, Toshihide Yamashita, Minoru Asada, Shuki Mizutani, Hideki Yoshikawa, and Masaya Tohyama. Cytoplasmic p21^{Cip1}/W AF 1 regulates neurite remodeling by inhibiting Rho-kinase activity. Journal of Cell Biology, 158 (2):321–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/2/321.

Takeda:2000:KSP

[TYS^{+}00] Sen Takeda, Hiroto Yamazaki, Dae-Hyun Seog, Yoshimitsu Kanai, Sumio Terada, and Nobutaka Hirokawa. Kinesin Superfamily Protein 3 (Kif3) Motor Transports Fodrin– Associating Vesicles Important for Neurite Building. Journal of Cell Biology, 148(6):1255–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/6/1255.

Totsukawa:2000:DRR

[TYY^{+}00] Go Totsukawa, Yoshihiko Yamakita, Shigeko Yamashiro, David J. Hartshorne, Yasuharu Sasaki, and Fumio Matsumura. Distinct Roles of Rock (Rho–Kinase) and Mlck in Spatial Regulation of Mlc Phosphorylation for Assembly of Stress Fibers and Focal Adhesions in 3t3 Fibroblasts. Journal of Cell Biology, 150(4):797–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/4/797.

Ueda:2000:SPK

[UAZG00] Hiroki Ueda, Smita Abbi, Chuanhai Zheng, and Jun-Lin Guan. Suppression of Pyk2 Kinase and Cellular Activities by Fip200. Journal of Cell Biology, 149(2):423–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
2/423.

Urano:2002:SPC

[UCY^{+}02] Fumihiko Urano, Marcella Calfon, Takunari Yoneda, Chi Yun, Moni Kiraly, Scott G. Clark, and David Ron. A survival pathway for Caenorhabditis elegans with a blocked unfolded protein response. Journal of Cell Biology, 158(4):639–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
158/4/639.

Umezu-Goto:2002:ALD

[UGKT^{+}02] Makiko Umezu-Goto, Yasuhiro Kishi, Akitsu Taira, Kotaro Hama, Naoshi Dohmae, Koji Takio, Takao Yamori, Gordon B. Mills, Keizo Inoue, Junken Aoki, and Hiroyuki Arai. Autotaxin has lysophospholipase D activity leading to tumor cell growth and motility by lysophosphatidic acid production. Journal of Cell Biology, 158(2):227–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/2/227.

Ullers:2003:ISR

[UHR^{+}03] Ronald S. Ullers, Edith N. G. Houben, Amanda Raine, Corinne M. ten Hagen-Jongman, Måns Ehrenberg, Joseph Brunner, Bauke Oudega, Nellie Harms, and Joen Luirink. Interplay of signal recognition particle and trigger factor at L23 near the nascent chain exit site on the Escherichia coli ribosome. Journal of Cell Biology, 161(4):679–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/4/
679.

Ueta:2001:SMC

[UIK^{+}01] Chisato Ueta, Masahiro Iwamoto, Naoko Kanatani, Carolina Yoshida, Yang Liu, Motomi Enomoto-Iwamoto, Tomoharu Ohmori, Hirayuki Enomoto, Ken Nakata, Kenji Takada, Kojiro Kurisu, and Toshihisa Komori. Skeletal Malformations Caused by Overexpression of Cbfa1 or Its Dominant Negative Form in Chondrocytes. Journal of Cell Biology, 153 (1):87–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/1/87.

Uekita:2001:CTD

[UIY^{+}01] Takamasa Uekita, Yoshifumi Itoh, Ikuo Yana, Hiroshi Ohno, and Motoharu Seiki. Cytoplasmic tail-dependent internalization of membrane-type 1 matrix metalloproteinase is important for its invasion-promoting activity. Journal of Cell Biology, 155(7):1345–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/7/1345.

Uchiyama:2002:VNE

[UJK^{+}02] Keiji Uchiyama, Eija Jokitalo, Fumi Kano, Masayuki Murata, Xiaodong Zhang, Benito Canas, Richard Newman, Catherine Rabouille, Darryl Pappin, Paul Freemont, and Hisao Kondo. VCIP135, a novel essential factor for p97/ p47-mediated membrane fusion, is required for Golgi and ER assembly in vivo. Journal of Cell Biology, 159(5):855– ??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/5/855.

Uchiyama:2003:LPP

[UJL^{+}03] Keiji Uchiyama, Eija Jokitalo, Mervi Lindman, Mark Jackman, Fumi Kano, Masayuki Murata, Xiaodong Zhang, and Hisao Kondo. The localization and phosphorylation of p47 are important for Golgi disassembly–assembly during the cell cycle. Journal of Cell Biology, 161(6):1067–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/161/6/
1067.

Uhles:2003:ISI

[UML^{+}03] Sabine Uhles, Tilo Moede, Barbara Leibiger, Per-Olof Berggren, and Ingo B. Leibiger. Isoform-specific insulin receptor signaling involves different plasma membrane domains. Journal of Cell Biology, 163(6):1327–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
6/1327.

Umebayashi:2003:ERT

[UN03] Kyohei Umebayashi and Akihiko Nakano. Ergosterol is required for targeting of tryptophan permease to the yeast plasma membrane. Journal of Cell Biology, 161(6):1117–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
161/6/1117.

Ungar:2002:CMG

[UOB^{+}02] Daniel Ungar, Toshihiko Oka, Elizabeth E. Brittle, Eliza Vasile, Vladimir V. Lupashin, Jon E. Chatterton, John E. Heuser, Monty Krieger, and M. Gerard Waters. Characterization of a mammalian Golgi-localized protein complex, COG, that is required for normal Golgi morphology and function. Journal of Cell Biology, 157(3):405–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/157/3/
405.

Uetake:2004:CCP

[US04] Yumi Uetake and Greenfield Sluder. Cell cycle progression after cleavage failure. Journal of Cell Biology, 165(5):609–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
165/5/609.

Ueno:2002:MFZ

[UTH^{+}02] Sayaka Ueno, Masako Tsukamoto, Tomoya Hirano, Kazuya Kikuchi, Maki K. Yamada, Nobuyoshi Nishiyama, Tetsuo Nagano, Norio Matsuki, and Yuji Ikegaya. Mossy fiber Zn^{2+} spillover modulates heterosynaptic N-methyl-d-aspartate receptor activity in hippocampal CA3 circuits. Journal of Cell Biology, 158(2):215–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/2/215.

Verheggen:2000:RRP

[VAHV00] Céline Verheggen, Geneviève Almouzni, and Danièle Hernandez-Verdun. The Ribosomal RNA Processing Machinery Is Recruited to the Nucleolar Domain before RNA Polymerase I during Xenopus laevis Development. Journal of Cell Biology, 149(2):293–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/2/293.

Vale:2000:AP

[Val00] Ronald D. Vale. Aaa Proteins. Journal of Cell Biology, 150(1):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/1/F13.

Vale:2003:MVM

[Val03] Ronald D. Vale. Myosin V motor proteins. Journal of Cell Biology, 163(3):445–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/3/445.

vonAhsen:2000:PMS

[vARP^{+}00] Oliver von Ahsen, Christian Renken, Guy Perkins, Ruth M. Kluck, Ella Bossy-Wetzel, and Donald D. Newmeyer. Preservation of mitochondrial structure and function after Bid- or Bax-mediated cytochrome c release. Journal of Cell Biology, 150(5):1027–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/5/1027.

vanBreugel:2003:SBY

[vBDH03] Mark van Breugel, David Drechsel, and Anthony Hyman. Stu2p, the budding yeast member of the conserved Dis1/ XMAP215 family of microtubule-associated proteins is a plus end–binding microtubule destabilizer. Journal of Cell Biology, 161(2):359–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/359.

Villalba:2002:TPC

[VBH^{+}02] Martin Villalba, Kun Bi, Junru Hu, Yoav Altman, Paul Bushway, Eric Reits, Jacques Neefjes, Gottfried Baier, Robert T. Abraham, and Amnon Altman. Translocation of PKCθ in T cells is mediated by a nonconventional, PI3-K-and Vav-dependent pathway, but does not absolutely require phospholipase C. Journal of Cell Biology, 157(2):253–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
157/2/253.

Vieira:2001:DRC

[VBR^{+}01a] Otilia V. Vieira, Roberto J. Botelho, Lucia Rameh, Saskia M. Brachmann, Tsuyoshi Matsuo, Howard W. Davidson, Alan Schreiber, Jonathan M. Backer, Lewis C. Cantley, and Sergio Grinstein. Distinct roles of class I and class III phosphatidylinositol 3-kinases in phagosome formation and maturation. Journal of Cell Biology, 155(1):19–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/155/1/
19.

Villalba:2001:VRD

[VBR^{+}01b] Martin Villalba, Kun Bi, Fernando Rodriguez, Yoshihiko Tanaka, Stephen Schoenberger, and Amnon Altman. Vav1/ rac-dependent actin cytoskeleton reorganization is required for lipid raft clustering in T cells. Journal of Cell Biology, 155(3):331–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/3/331.

Vindis:2003:ERC

[VCDHD03] Cécile Vindis, Douglas Pat Cerretti, Thomas O. Daniel, and Uyen Huynh-Do. EphB1 recruits c–Src and p52^{Shc} to activate MAPK/ERK and promote chemotaxis. Journal of Cell Biology, 162(4):661–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/4/661.

Vernooy:2000:CDR

[VCG^{+}00] Stephanie Y. Vernooy, Jeffrey Copeland, Nazli Ghaboosi, Erik E. Griffin, Soon Ji Yoo, and Bruce A. Hay. Cell Death Regulation in Drosophila. Journal of Cell Biology, 150(2):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
150/2/F69.

Vitale:2002:CRE

[VCGB^{+}02] Nicolas Vitale, Sylvette Chasserot-Golaz, Yannick Bailly, Naoko Morinaga, Michael A. Frohman, and Marie-France Bader. Calcium-regulated exocytosis of dense-core vesicles requires the activation of ADP-ribosylation factor (ARF)6 by ARF nucleotide binding site opener at the plasma membrane. Journal of Cell Biology, 159(1):79–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/1/
79.

vanderBliek:2000:MDA

[vdB00] Alexander M. van der Bliek. A Mitochondrial Division Apparatus Takes Shape. Journal of Cell Biology, 151(2):0–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
151/2/F1.

vandenBoom:2004:DDS

[vdBCH^{+}04] Vincent van den Boom, Elisabetta Citterio, Deborah Hoogstraten, Angelika Zotter, Jean-Marc Egly, Wiggert A. van Cappellen, Jan H. J. Hoeijmakers, Adriaan B. Houtsmuller, and Wim Vermeulen. DNA damage stabilizes interaction of CSB with the transcription elongation machinery. Journal of Cell Biology, 166(1):27–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/166/1/
27.

vanderFlier:2002:DSV

[vdFKK^{+}02] Arjan van der Flier, Ingrid Kuikman, Duco Kramer, Dirk Geerts, Maaike Kreft, Toshiro Takafuta, Sandor S. Shapiro, and Arnoud Sonnenberg. Different splice variants of filamin– B affect myogenesis, subcellular distribution, and determine binding to integrin β subunits. Journal of Cell Biology, 156 (2):361–??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/2/361.

vanOordt:2000:MPS

[vdHvODML^{+}00] Willemien van der Houven van Oordt, María T. Diaz-Meco, José Lozano, Adrian R. Krainer, Jorge Moscat, and Javier F. Cáceres. The Mkk_{3/6}-p38-signaling cascade alters the subcellular distribution of Hnrnp A1 and modulates alternative splicing regulation. Journal of Cell Biology, 149(2):307–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
149/2/307.

vanderLaan:2004:FFA

[vdLBK^{+}04] Martin van der Laan, Philipp Bechtluft, Stef Kol, Nico Nouwen, and Arnold J. M. Driessen. F_{1} F_{0} ATP synthase subunit c is a substrate of the novel YidC pathway for membrane protein biogenesis. Journal of Cell Biology, 165(2): 213–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/2/213.

Volkmann:2001:AMA

[VDMH01] Niels Volkmann, David DeRosier, Paul Matsudaira, and Dorit Hanein. An Atomic Model of Actin Filaments Cross– Linked by Fimbrin and Its Implications for Bundle Assembly and Function. Journal of Cell Biology, 153(5):947–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
5/947.

vanderVen:2000:INM

[vdVWS^{+}00] Peter F. M. van der Ven, Sebastian Wiesner, Paula Salmikangas, Daniel Auerbach, Mirko Himmel, Stefan Kempa, Katrin Hayeß, Dirk Pacholsky, Anu Taivainen, Rolf Schröder, Olli Carpén, and Dieter O. Fürst. Indications for a Novel Muscular Dystrophy Pathway. Journal of Cell Biology, 151(2): 235–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/2/235.

vanEeden:2001:BEP

[vEPP^{+}01] Fredericus J. M. van Eeden, Isabel M. Palacios, Mark Petronczki, Matthew J. D. Weston, and Daniel St Johnston. Barentsz is essential for the posterior localization of oskar mRNA and colocalizes with it to the posterior pole. Journal of Cell Biology, 154(3):511–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/3/511.

vanEs:2001:TNM

[vEWS^{+}01] Saskia van Es, Deborah Wessels, David R. Soll, Jane Borleis, and Peter N. Devreotes. Tortoise, a Novel Mitochondrial Protein, Is Required for Directional Responses of Dictyostelium in Chemotactic Gradients. Journal of Cell Biology, 152(3): 621–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/3/621.

Vanderluit:2004:PRN

[VFN^{+}04] Jacqueline L. Vanderluit, Kerry L. Ferguson, Vassiliki Nikoletopoulou, Maura Parker, Vladimir Ruzhynsky, Tania Alexson, Stephen M. McNamara, David S. Park, Michael Rudnicki, and Ruth S. Slack. p107 regulates neural precursor cells in the mammalian brain. Journal of Cell Biology, 166(6): 853–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/6/853.

Vilar:2003:MND

[VGL03] José M. G. Vilar, Călin C. Guet, and Stanislas Leibler. Modeling network dynamics. Journal of Cell Biology, 161 (3):471–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/3/471.

Verkade:2000:ICA

[VHLS00] Paul Verkade, Thomas Harder, Frank Lafont, and Kai Simons. Induction of Caveolae in the Apical Plasma Membrane of Madin–Darby Canine Kidney Cells. Journal of Cell Biology, 148(4):727–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/148/4/727.

Vinkemeier:2004:GMA

[Vin04] Uwe Vinkemeier. Getting the message across, STAT! Design principles of a molecular signaling circuit. Journal of Cell Biology, 167(2):197–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/2/197.

Vashist:2001:DRR

[VKB^{+}01] Shilpa Vashist, Woong Kim, William J. Belden, Eric D. Spear, Charles Barlowe, and Davis T. W. Ng. Distinct retrieval and retention mechanisms are required for the quality control of endoplasmic reticulum protein folding. Journal of Cell Biology, 155(3):355–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/3/355.

Voet:2003:TIH

[VLL^{+}03] Thierry Voet, Bodo Liebe, Charlotte Labaere, Peter Marynen, and Harry Scherthan. Telomere-independent homologue pairing and checkpoint escape of accessory ring chromosomes in male mouse meiosis. Journal of Cell Biology, 162(5):795– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/5/795.

vanMeer:2001:CCL

[vM01] Gerrit van Meer. Caveolin, Cholesterol, and Lipid Droplets? Journal of Cell Biology, 152(5):??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/5/F29.

VanSlyke:2002:DDR

[VM02] Judy K. VanSlyke and Linda S. Musil. Dislocation and degradation from the ER are regulated by cytosolic stress. Journal of Cell Biology, 157(3):381–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/3/381.

Verhey:2001:CKI

[VMD^{+}01] Kristen J. Verhey, Debra Meyer, Reneé Deehan, John Blenis, Bruce J. Schnapp, Tom A. Rapoport, and Ben Margolis. Cargo of Kinesin Identified as Jip Scaffolding Proteins and Associated Signaling Molecules. Journal of Cell Biology, 152 (5):959–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/5/959.

Vaughan:2002:RRB

[VMH^{+}02] Patricia S. Vaughan, Pedro Miura, Matthew Henderson, Belinda Byrne, and Kevin T. Vaughan. A role for regulated binding of p150 ^{Glued} to microtubule plus ends in organelle transport. Journal of Cell Biology, 158(2):305–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/2/

Valentijn:2003:STC

[VMK^{+}03] Anthony J. Valentijn, Anthony D. Metcalfe, Jane Kott, Charles H. Streuli, and Andrew P. Gilmore. Spatial and temporal changes in Bax subcellular localization during anoikis. Journal of Cell Biology, 162(4):599–??, August 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/162/4/
599.

Vitha:2001:FRF

[VMO01] Stanislav Vitha, Rosemary S. McAndrew, and Katherine W. Osteryoung. Ftsz Ring Formation at the Chloroplast Division Site in Plants. Journal of Cell Biology, 153(1):111–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/153/
1/111.

Vandebrouck:2002:ITA

[VMS^{+}02] Clarisse Vandebrouck, Dominique Martin, Monique Colson-Van Schoor, Huguette Debaix, and Philippe Gailly. Involvement of TRPC in the abnormal calcium influx observed in dystrophic (mdx) mouse skeletal muscle fibers. Journal of Cell Biology, 158(6):1089–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/6/1089.

vonMikecz:2000:CBP

[vMZM^{+}00] Anna von Mikecz, Suisheng Zhang, Marc Montminy, Eng M. Tan, and Peter Hemmerich. Creb-binding protein (Cbp/ P300) and RNA polymerase II colocalize in transcriptionally active domains in the nucleus. Journal of Cell Biology, 150(1):265–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/1/265.

Vashist:2004:MPS

[VN04] Shilpa Vashist and Davis T. W. Ng. Misfolded proteins are sorted by a sequential checkpoint mechanism of ER quality control. Journal of Cell Biology, 165(1):41–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/165/1/

Vecchi:2001:NSE

[VPP^{+}01] Manuela Vecchi, Simona Polo, Viviane Poupon, Jan-Willem van de Loo, Alexandre Benmerah, and Pier Paolo Di Fiore. Nucleocytoplasmic Shuttling of Endocytic Proteins. Journal of Cell Biology, 153(7):1511–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/153/7/1511.

vanRoermund:2000:PPP

[vRTvdB^{+}00] Carlo W. T. van Roermund, Henk F. Tabak, Marlene van den Berg, Ronald J. A. Wanders, and Ewald H. Hettema. Pex11p Plays a Primary Role in Medium–Chain Fatty Acid Oxidation, a Process That Affects Peroxisome Number and Size in Saccharomyces cerevisiae. Journal of Cell Biology, 150(3): 489–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/3/489.

Vasu:2001:NVN

[VSO^{+}01] Sanjay Vasu, Sundeep Shah, Arturo Orjalo, Minkyu Park, Wolfgang H. Fischer, and Douglass J. Forbes. Novel vertebrate nucleoporins Nup133 and Nup160 play a role in mRNA export. Journal of Cell Biology, 155(3):339–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/155/3/
339.

Versele:2004:SCF

[VT04] Matthias Versele and Jeremy Thorner. Septin collar formation in budding yeast requires GTP binding and direct phosphorylation by the PAK, Cla4. Journal of Cell Biology, 164 (5):701–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/5/701.

Vizeacoumar:2003:YYE

[VTGT^{+}03] Franco J. Vizeacoumar, Juan C. Torres-Guzman, Yuen Yi C. Tam, John D. Aitchison, and Richard A. Rachubinski. YHR150w and YDR479c encode peroxisomal integral membrane proteins involved in the regulation of peroxisome number, size, and distribution in Saccharomyces cerevisiae. Journal of Cell Biology, 161(2):321–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/2/321.

VanHatten:2002:XXP

[VTH^{+}02] Ruth A. Van Hatten, Antonin V. Tutter, Antonia H. Holway, Alyssa M. Khederian, Johannes C. Walter, and W. Matthew Michael. The Xenopus Xmus101 protein is required for the recruitment of Cdc45 to origins of DNA replication. Journal of Cell Biology, 159(4):541–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/4/541.

Varlamov:2004:S

[VVR^{+}04] Oleg Varlamov, Allen Volchuk, Vahid Rahimian, Claudia A. Doege, Fabienne Paumet, William S. Eng, Nancy Arango, Francesco Parlati, Mariella Ravazzola, Lelio Orci, Thomas H. Söllner, and James E. Rothman. i-SNAREs. Journal of Cell Biology, 164(1):79–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/1/79.

vonWichert:2003:RAT

[vWJK^{+}03] Götz von Wichert, Guoying Jiang, Ana Kostic, Kurt De Vos, Jan Sap, and Michael P. Sheetz. RPTP-α acts as a transducer of mechanical force on α_{v}/β_{3}-integrin-cytoskeleton linkages. Journal of Cell Biology, 161(1):143–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/1/143.

Velmurugan:2000:PCP

[VYC^{+}00] Soundarapandian Velmurugan, Xian-Mei Yang, Clarence S.- M. Chan, Melanie Dobson, and Makkuni Jayaram. Partitioning of the 2-μ m Circle Plasmid of Saccharomyces cerevisiae. Journal of Cell Biology, 149(3):553–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/3/553.

Vignjevic:2003:FFL

[VYW^{+}03] Danijela Vignjevic, Defne Yarar, Matthew D. Welch, John Peloquin, Tatyana Svitkina, and Gary G. Borisy. Formation of filopodia-like bundles in vitro from a dendritic network. Journal of Cell Biology, 160(6):951–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/6/951.

Violin:2003:GEF

[VZTN03] Jonathan D. Violin, Jin Zhang, Roger Y. Tsien, and Alexandra C. Newton. A genetically encoded fluorescent reporter reveals oscillatory phosphorylation by protein kinase C. Journal of Cell Biology, 161(5):899–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/5/899.

Wisco:2003:UMA

[WAC^{+}03] Dolora Wisco, Eric D. Anderson, Michael C. Chang, Caren Norden, Tatiana Boiko, Heike Fölsch, and Bettina Winckler. Uncovering multiple axonal targeting pathways in hippocampal neurons. Journal of Cell Biology, 162(7):1317– ??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/7/1317.

Wilson-Annan:2003:PBO

[WAOC^{+}03] Julie Wilson-Annan, Lorraine A. O’Reilly, Simon A. Crawford, George Hausmann, Jennifer G. Beaumont, Loes P. Parma, Lin Chen, Martin Lackmann, Trevor Lithgow, Mark G. Hinds, Catherine L. Day, Jerry M. Adams, and David C. S. Huang. Proapoptotic BH3-only proteins trigger membrane integration of prosurvival Bcl-w and neutralize its activity. Journal of Cell Biology, 162(5):877–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
5/877.

Wiederkehr:2000:FBP

[WAPB^{+}00] Andreas Wiederkehr, Sandrine Avaro, Cristina Prescianotto-Baschong, Rosine Haguenauer-Tsapis, and Howard Riezman. The F–Box Protein Rcy1p Is Involved in Endocytic Membrane Traffic and Recycling Out of an Early Endosome in Saccharomyces cerevisiae. Journal of Cell Biology, 149(2): 397–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/2/397.

Williams:2004:EHR

[WAV^{+}04] Christopher C. Williams, June G. Allison, Gregory A. Vidal, Matthew E. Burow, Barbara S. Beckman, Luis Marrero, and Frank E. Jones. The ERBB4/HER4 receptor tyrosine kinase regulates gene expression by functioning as a STAT5A nuclear chaperone. Journal of Cell Biology, 167(3): 469–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/3/469.

Wightman:2004:CAN

[WBAS04] Raymond Wightman, Steven Bates, Pat Amornrrattanapan, and Peter Sudbery. In Candida albicans, the Nim1 kinases Gin4 and Hsl1 negatively regulate pseudohypha formation and Gin4 also controls septin organization. Journal of Cell Biology, 164(4):581–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/4/581.

Wahl:2000:EAI

[WBC^{+}00] Siegfried Wahl, Holger Barth, Thomas Ciossek, Klaus Aktories, and Bernhard K. Mueller. Ephrin-A5 Induces Collapse of Growth Cones by Activating Rho and Rho Kinase. Journal of Cell Biology, 149(2):263–??, April 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/149/2/263.

Wakefield:2001:DPA

[WBG01] James G. Wakefield, Silvia Bonaccorsi, and Maurizio Gatti. The Drosophila Protein Asp Is Involved in Microtubule Organization during Spindle Formation and Cytokinesis. Journal of Cell Biology, 153(4):637–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/4/637.

Wolven:2000:VIA

[WBM^{+}00] Amy K. Wolven, Lisa D. Belmont, Nicole M. Mahoney, Steven C. Almo, and David G. Drubin. In Vivo Importance of Actin Nucleotide Exchange Catalyzed by Profilin. Journal of Cell Biology, 150(4):895–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/4/895.

Weaver:2003:CAP

[WBP^{+}03] Beth A. A. Weaver, Zahid Q. Bonday, Frances R. Putkey, Geert J. P. L. Kops, Alain D. Silk, and Don W. Cleveland. Centromere-associated protein–E is essential for the mammalian mitotic checkpoint to prevent aneuploidy due to single chromosome loss. Journal of Cell Biology, 162(4):551–??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
162/4/551.

Westfall:2003:WPF

[WBT^{+}03] Trudi A. Westfall, Ryan Brimeyer, Jen Twedt, Jean Gladon, Andrea Olberding, Makoto Furutani-Seiki, and Diane C. Slusarski. Wnt-5/pipetail functions in vertebrate axis formation as a negative regulator of Wnt/β-catenin activity. Journal of Cell Biology, 162(5):889–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/5/889.

Weston:2003:ARU

[WBU03] Andrea D. Weston, Bruce Blumberg, and T. Michael Underhill. Active repression by unliganded retinoid receptors in development. Journal of Cell Biology, 161(2):223–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
2/223.

Wittmann:2003:RLE

[WBWS03] Torsten Wittmann, Gary M. Bokoch, and Clare M. Waterman-Storer. Regulation of leading edge microtubule and actin dynamics downstream of Rac1. Journal of Cell Biology, 161(5):845–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/5/845.

Wong:2003:LKK

[WC03] Pauline Wong and Pierre A. Coulombe. Loss of keratin 6 (K6) proteins reveals a function for intermediate filaments during wound repair. Journal of Cell Biology, 163(2):327–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
163/2/327.

Westermann:2003:ABY

[WCA^{+}03] Stefan Westermann, Iain M. Cheeseman, Scott Anderson, John R. Yates, David G. Drubin, and Georjana Barnes. Architecture of the budding yeast kinetochore reveals a conserved molecular core. Journal of Cell Biology, 163(2): 215–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/2/215.

Weis:2004:EBD

[WCBC04] Sara Weis, Jianhua Cui, Leo Barnes, and David Cheresh. Endothelial barrier disruption by VEGF-mediated Src activity potentiates tumor cell extravasation and metastasis. Journal of Cell Biology, 167(2):223–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/2/223.

Wong:2000:INM

[WCGT^{+}00] Pauline Wong, Emma Colucci-Guyon, Kenzo Takahashi, Changhong Gu, Charles Babinet, and Pierre A. Coulombe. Introducing a Null Mutation in the Mouse K6α and K6β Genes Reveals Their Essential Structural Role in the Oral Mucosa. Journal of Cell Biology, 150(4):921–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/150/
4/921.

Wagner:2004:VFR

[WCIN04] Stefan Wagner, Simion Chiosea, Maria Ivshina, and Jeffrey A. Nickerson. In vitro FRAP reveals the ATP-dependent nuclear mobilization of the exon junction complex protein SRm160. Journal of Cell Biology, 164(6):843–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
6/843.

Weston:2002:RRM

[WCTU02] Andrea D. Weston, Roshantha A. S. Chandraratna, Joseph Torchia, and T. Michael Underhill. Requirement for RARmediated gene repression in skeletal progenitor differentiation. Journal of Cell Biology, 158(1):39–??, July 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/1/
39.

Wu:2001:ILK

[WD01] Chuanyue Wu and Shoukat Dedhar. Integrin-linked kinase (ILK) and its interactors. Journal of Cell Biology, 155(4): 505–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/4/505.

Wiederkehr:2004:FSW

[WDFNN04] Andreas Wiederkehr, Johan-Owen De Craene, Susan Ferro-Novick, and Peter Novick. Functional specialization within a vesicle tethering complex. Journal of Cell Biology, 167(5): 875–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/5/875.

Wrobel:2004:ACA

[WDL^{+}04] Carolyn N. Wrobel, Jayanta Debnath, Eva Lin, Sean Beausoleil, Martine F. Roussel, and Joan S. Brugge. Autocrine CSF-1R activation promotes Src-dependent disruption of mammary epithelial architecture. Journal of Cell Biology, 165(2):263–??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/2/263.

Walenta:2001:GAH

[WDLK01] Jason H. Walenta, Aaron J. Didier, Xinran Liu, and Helmut Krämer. The Golgi–Associated Hook3 Protein Is a Member of a Novel Family of Microtubule–Binding Proteins. Journal of Cell Biology, 152(5):923–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/152/5/923.

Wignall:2003:CCR

[WDMH03] Sarah M. Wignall, Renée Deehan, Thomas J. Maresca, and Rebecca Heald. The condensin complex is required for proper spindle assembly and chromosome segregation in Xenopus egg extracts. Journal of Cell Biology, 161(6):1041–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
6/1041.

Wozniak:2003:RGC

[WDS^{+}03] Michele A. Wozniak, Radhika Desai, Patricia A. Solski, Channing J. Der, and Patricia J. Keely. ROCK-generated contractility regulates breast epithelial cell differentiation in response to the physical properties of a three-dimensional collagen matrix. Journal of Cell Biology, 163(3):583–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
163/3/583.

Wood:2003:SIC

[WDW03] Chris D. Wood, Alberto Darszon, and Michael Whitaker. Speract induces calcium oscillations in the sperm tail. Journal of Cell Biology, 161(1):89–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/1/89.

Wurmser:2002:NPP

[WE02] Andrew E. Wurmser and Scott D. Emr. Novel PtdIns(3)Pbinding protein Etf1 functions as an effector of the Vps34 PtdIns 3-kinase in autophagy. Journal of Cell Biology, 158 (4):761–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/4/761.

Wells:2000:B

[Wel00] William A. Wells. In brief. Journal of Cell Biology, 148 (1):??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/1/NP.1.

Wells:2001:AO

[Wel01a] William A. Wells. Actin’s origins. Journal of Cell Biology, 155(1):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/1/13.1.

Wells:2001:ARS

[Wel01b] William A. Wells. Algal Rb sets size limits. Journal of Cell Biology, 154(1):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/1/12.1.

Wells:2001:ASB

[Wel01c] William A. Wells. All skin and brain. Journal of Cell Biology, 154(5):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/5/904.1.

Wells:2001:ATC

[Wel01d] William A. Wells. Arms together, CENs apart. Journal of Cell Biology, 155(3):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/3/325.2.

Wells:2001:BAJ

[Wel01e] William A. Wells. Battle of the actin junctions. Journal of Cell Biology, 155(4):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/4/494.2.

Wells:2001:BBb

[Wel01f] William A. Wells. Beware the Borg. Journal of Cell Biology, 155(1):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/1/13.2.

Wells:2001:BBc

[Wel01g] William A. Wells. Bind and bend. Journal of Cell Biology, 155(2):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/2/177.1.

Wells:2001:BP

[Wel01h] William A. Wells. Biotechnology on Parade. Journal of Cell Biology, 154(3):482–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/3/482.

Wells:2001:BBa

[Wel01i] William A. Wells. Bone breakage. Journal of Cell Biology, 155(1):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/1/10.3.

Wells:2001:CCC

[Wel01j] William A. Wells. Calcium-coated chromosomes. Journal of Cell Biology, 155(6):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/6/869.1.

Wells:2001:CF

[Wel01k] William A. Wells. CENP-e falls off. Journal of Cell Biology, 154(4):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/4/674.2.

Wells:2001:CSO

[Wel01l] William A. Wells. Chopping and storing in one. Journal of Cell Biology, 155(6):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/6/868.3.

Wells:2001:CM

[Wel01m] William A. Wells. Cleaving and migrating. Journal of Cell Biology, 155(4):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/4/494.1.

Wells:2001:CCa

[Wel01n] William A. Wells. Cohesin’s cohesion. Journal of Cell Biology, 154(5):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/5/905.3.

Wells:2001:CID

[Wel01o] William A. Wells. Cold-induced dwarfism. Journal of Cell Biology, 155(1):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/1/12.2.

Wells:2001:CIS

[Wel01p] William A. Wells. Compartmentalized insulin signaling. Journal of Cell Biology, 154(4):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/4/675.3.

Wells:2001:CN

[Wel01q] William A. Wells. Connect the Nerves. Journal of Cell Biology, 155(7):1096–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/7/1096.

Wells:2001:CB

[Wel01r] William A. Wells. Containing the bullet. Journal of Cell Biology, 154(1):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/1/13.1.

Wells:2001:CCb

[Wel01s] William A. Wells. Cooperation in the cerebellum. Journal of Cell Biology, 154(6):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/6/1098.3.

Wells:2001:CLC

[Wel01t] William A. Wells. Cross-linking can be good. Journal of Cell Biology, 154(4):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/4/674.1.

Wells:2001:DCG

[Wel01u] William A. Wells. Dead cells get the squeeze. Journal of Cell Biology, 155(7):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/7/1094.1.

Wells:2001:DD

[Wel01v] William A. Wells. Death and Destruction. Journal of Cell Biology, 154(1):14–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/1/14.

Wells:2001:EV

[Wel01w] William A. Wells. Enter the virus. Journal of Cell Biology, 155(7):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/7/1094.2.

Wells:2001:EC

[Wel01x] William A. Wells. Exchanging clathrin. Journal of Cell Biology, 155(2):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/2/177.3.

Wells:2001:ERS

[Wel01y] William A. Wells. An extended release for surfactant. Journal of Cell Biology, 155(2):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/2/176.1.

Wells:2001:FLM

[Wel01z] William A. Wells. Formin a link to microtubules. Journal of Cell Biology, 154(3):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/3/481.2.

Wells:2001:NPK

[Wel01-27] William A. Wells. From nuclear pore to kinetochore. Journal of Cell Biology, 154(6):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/6/1099.2.

Wells:2001:HRF

[Wel01-28] William A. Wells. Histones Rule!: The FASEB Conference on Chromatin and Transcription. July 7–12, 2001. Journal of Cell Biology, 154(5):906–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/5/906.

Wells:2001:JH

[Wel01-29] William A. Wells. Joined at the Hip. Journal of Cell Biology, 154(6):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/6/1098.1.

Wells:2001:KL

[Wel01-30] William A. Wells. Killing for love. Journal of Cell Biology, 155(3):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/3/325.3.

Wells:2001:LMG

[Wel01-31] William A. Wells. Let’s make Golgi. Journal of Cell Biology, 155(4):498–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/4/498.

Wells:2001:MRN

[Wel01-32] William A. Wells. Making room for new memories. Journal of Cell Biology, 155(7):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/7/1095.2.

Wells:2001:MVA

[Wel01-33] William A. Wells. Making virus in the axon. Journal of Cell Biology, 154(4):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/4/675.2.

Wells:2001:MM

[Wel01-34] William A. Wells. A marker for munching. Journal of Cell Biology, 155(4):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/4/495.2.

Wells:2001:MC

[Wel01-35] William A. Wells. Microtubules concentrate. Journal of Cell Biology, 154(1):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/1/13.3.

Wells:2001:MW

[Wel01-36] William A. Wells. Mitosis without wind. Journal of Cell Biology, 154(6):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/6/1099.1.

Wells:2001:MV

[Wel01-37] William A. Wells. Myosin V. Journal of Cell Biology, 155 (4):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/4/495.1.

Wells:2001:NB

[Wel01-38] William A. Wells. A nurse for bicoid. Journal of Cell Biology, 154(3):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/3/480.2.

Wells:2001:OCG

[Wel01-39] William A. Wells. Old cells gone bad. Journal of Cell Biology, 155(3):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/3/324.1.

Wells:2001:OK

[Wel01-40] William A. Wells. An organelle knockout. Journal of Cell Biology, 154(5):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/5/904.2.

Wells:2001:PU

[Wel01-41] William A. Wells. p115 untethered. Journal of Cell Biology, 155(2):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/2/177.2.

Wells:2001:PM

[Wel01-42] William A. Wells. Peroxisomes on the move. Journal of Cell Biology, 155(6):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/6/868.2.

Wells:2001:PN

[Wel01-43] William A. Wells. Pinning down NuMA. Journal of Cell Biology, 155(5):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/5/693.1.

Wells:2001:RFR

[Wel01-44] William A. Wells. Rafts to the front; rafts to the rear. Journal of Cell Biology, 154(5):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/5/905.2.

Wells:2001:RH

[Wel01-45] William A. Wells. Remaking the heart. Journal of Cell Biology, 155(1):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/1/13.3.

Wells:2001:RLP

[Wel01-46] William A. Wells. Repair long and prosper. Journal of Cell Biology, 155(7):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/7/1094.3.

Wells:2001:RL

[Wel01-47] William A. Wells. Romeo and Leishmania. Journal of Cell Biology, 155(7):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/7/1095.1.

Wells:2001:SSM

[Wel01-48] William A. Wells. Searching for a spindle matrix. Journal of Cell Biology, 154(6):1102–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/6/1102.

Wells:2001:SSd

[Wel01-49] William A. Wells. Sigma separation. Journal of Cell Biology, 155(5):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/5/693.2.

Wells:2001:SSa

[Wel01-50] William A. Wells. Squeezing in with sugar. Journal of Cell Biology, 154(5):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/5/905.1.

Wells:2001:SUM

[Wel01-51] William A. Wells. Sticky utrophin messages. Journal of Cell Biology, 154(6):??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/6/1098.2.

Wells:2001:SBD

[Wel01-52] William A. Wells. Sudden bacterial death syndrome. Journal of Cell Biology, 154(3):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/3/480.1.

Wells:2001:SLO

[Wel01-53] William A. Wells. Surviving in low oxygen. Journal of Cell Biology, 155(5):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/5/693.3.

Wells:2001:SSb

[Wel01-54] William A. Wells. Sustained specificity. Journal of Cell Biology, 155(3):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/3/324.2.

Wells:2001:SSc

[Wel01-55] William A. Wells. Swimming with sperm. Journal of Cell Biology, 155(3):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/3/325.1.

Wells:2001:TTA

[Wel01-56] William A. Wells. Taking tea with actin and microtubules. Journal of Cell Biology, 154(1):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/1/13.2.

Wells:2001:TPW

[Wel01-57] William A. Wells. Tmod points the way. Journal of Cell Biology, 155(6):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/6/868.1.

Wells:2001:TAS

[Wel01-58] William A. Wells. A TORrific ATP sensor. Journal of Cell Biology, 155(5):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/5/692.2.

Wells:2001:TCT

[Wel01-59] William A. Wells. Touching cells transform. Journal of Cell Biology, 154(3):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/3/481.3.

Wells:2001:TA

[Wel01-60] William A. Wells. Toxic apoptosis. Journal of Cell Biology, 155(4):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/4/494.3.

Wells:2001:TGL

[Wel01-61] William A. Wells. Transcription gets a licence. Journal of Cell Biology, 154(3):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/3/481.1.

Wells:2001:TN

[Wel01-62] William A. Wells. Translation in the nucleus. Journal of Cell Biology, 154(1):??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/1/12.2.

Wells:2001:UBC

[Wel01-63] William A. Wells. An unfolding breast cancer story. Journal of Cell Biology, 155(6):??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/6/869.2.

Wells:2001:VOL

[Wel01-64] William A. Wells. Virus one; laminin nil. Journal of Cell Biology, 155(2):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/2/176.2.

Wells:2001:VMH

[Wel01-65] William A. Wells. Vpr makes holes. Journal of Cell Biology, 155(5):??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/5/692.1.

Wells:2001:WA

[Wel01-66] William A. Wells. Wingless and the argosomes. Journal of Cell Biology, 155(1):??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/1/12.1.

Wells:2001:ZRW

[Wel01-67] William A. Wells. Zones and rings in wound healing. Journal of Cell Biology, 154(4):??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/4/675.1.

Wells:2002:SMF

[Wel02a] William A. Wells. 14-3-3, set me free. Journal of Cell Biology, 159(5):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/5/729.3.

Wells:2002:ATB

[Wel02b] William A. Wells. Adenovirus takes the back door. Journal of Cell Biology, 159(1):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/1/15.1.

Wells:2002:AM

[Wel02c] William A. Wells. Attack of the microvesicles. Journal of Cell Biology, 157(5):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/5/737.2.

Wells:2002:AW

[Wel02d] William A. Wells. Attack of the waves. Journal of Cell Biology, 158(2):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/2/195.2.

Wells:2002:BAF

[Wel02e] William A. Wells. A BAFfling actin function. Journal of Cell Biology, 157(1):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/1/12.1.

Wells:2002:BNE

[Wel02f] William A. Wells. Brave new eggs. Journal of Cell Biology, 158(4):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/4/607.3.

Wells:2002:BS

[Wel02g] William A. Wells. Bring in the SNAREs. Journal of Cell Biology, 157(3):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/3/344.3.

Wells:2002:BSP

[Wel02h] William A. Wells. Building a spindle pole. Journal of Cell Biology, 156(3):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/3/411.3.

Wells:2002:CJP

[Wel02i] William A. Wells. Cadherins jump into the pool. Journal of Cell Biology, 157(5):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/5/736.2.

Wells:2002:CTM

[Wel02j] William A. Wells. Cdc14p turns MEN on and off. Journal of Cell Biology, 157(3):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/3/345.1.

Wells:2002:CSV

[Wel02k] William A. Wells. Cells spin into view. Journal of Cell Biology, 159(5):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/5/728.2.

Wells:2002:CH

[Wel02l] William A. Wells. Channeling hedgehogs. Journal of Cell Biology, 158(6):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/6/993.2.

Wells:2002:CTW

[Wel02m] William A. Wells. A chemokine tug-of-war. Journal of Cell Biology, 157(1):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/1/13.1.

Wells:2002:CCa

[Wel02n] William A. Wells. Cloning confusion. Journal of Cell Biology, 156(3):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/3/412.1.

Wells:2002:CM

[Wel02o] William A. Wells. Comm with me. Journal of Cell Biology, 158(6):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/6/992.1.

Wells:2002:CWI

[Wel02p] William A. Wells. Condensin wraps it up. Journal of Cell Biology, 158(2):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/2/195.1.

Wells:2002:CCb

[Wel02q] William A. Wells. Cristae in crisis. Journal of Cell Biology, 156(5):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/5/769.2.

Wells:2002:CT

[Wel02r] William A. Wells. Crossed TRAILs. Journal of Cell Biology, 156(3):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/3/413.3.

Wells:2002:DA

[Wel02s] William A. Wells. Death by autophagy. Journal of Cell Biology, 157(3):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/3/345.2.

Wells:2002:DSM

[Wel02t] William A. Wells. Does size matter? Journal of Cell Biology, 158(7):1156–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/7/1156.

Wells:2002:EH

[Wel02u] William A. Wells. Electrical healing. Journal of Cell Biology, 159(1):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/159/1/14.1.

Wells:2002:CCc

[Wel02v] William A. Wells. An ES cell caution. Journal of Cell Biology, 157(1):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/1/13.3.

Wells:2002:ESC

[Wel02w] William A. Wells. An exit strategy for the clean-up crew. Journal of Cell Biology, 159(5):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/5/729.2.

Wells:2002:FCA

[Wel02x] William A. Wells. Flicking the chromosome–attachment switch. Journal of Cell Biology, 159(3):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/3/
390.1.

Wells:2002:FP

[Wel02y] William A. Wells. A flipping permease. Journal of Cell Biology, 157(5):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/5/737.1.

Wells:2002:GB

[Wel02z] William A. Wells. Getting biorientated. Journal of Cell Biology, 156(5):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/5/768.2.

Wells:2002:GCC

[Wel02-27] William A. Wells. The Golgi’s central COG. Journal of Cell Biology, 157(3):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/3/344.1.

Wells:2002:GGC

[Wel02-28] William A. Wells. GRASP65 gets the chop. Journal of Cell Biology, 156(3):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/3/411.2.

Wells:2002:HDI

[Wel02-29] William A. Wells. The Hippi did it. Journal of Cell Biology, 156(3):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/3/413.2.

Wells:2002:HMH

[Wel02-30] William A. Wells. How to make a helix. Journal of Cell Biology, 157(5):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/5/736.1.

Wells:2002:FYL

[Wel02-31] William A. Wells. Is that a fly in your leg? Journal of Cell Biology, 158(4):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/4/607.1.

Wells:2002:TT

[Wel02-32] William A. Wells. Is transdifferentiation in trouble? Journal of Cell Biology, 157(1):15–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/1/15.

Wells:2002:JSI

[Wel02-33] William A. Wells. A JAK/STAT invasion. Journal of Cell Biology, 156(3):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/3/413.1.

Wells:2002:LTC

[Wel02-34] William A. Wells. Let there be a channel. Journal of Cell Biology, 158(2):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/2/195.3.

Wells:2002:LML

[Wel02-35] William A. Wells. Let’s make liver. Journal of Cell Biology, 156(1):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/1/8.2.

Wells:2002:MJW

[Wel02-36] William A. Wells. Make junction, will travel. Journal of Cell Biology, 157(1):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/1/13.2.

Wells:2002:MR

[Wel02-37] William A. Wells. A mitochondrial ratchet. Journal of Cell Biology, 158(4):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/4/606.1.

Wells:2002:MM

[Wel02-38] William A. Wells. Monitoring MLCK. Journal of Cell Biology, 156(3):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/3/410.1.

Wells:2002:MDI

[Wel02-39] William A. Wells. MS as a disease of immaturity. Journal of Cell Biology, 159(1):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/1/15.2.

Wells:2002:NC

[Wel02-40] William A. Wells. Nervous chemokines. Journal of Cell Biology, 158(2):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/2/194.2.

Wells:2002:NCL

[Wel02-41] William A. Wells. The numbers come to light. Journal of Cell Biology, 158(4):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/4/606.2.

Wells:2002:OU

[Wel02-42] William A. Wells. Old and unattracted. Journal of Cell Biology, 159(1):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/1/14.2.

Wells:2002:OBF

[Wel02-43] William A. Wells. One brain fits all. Journal of Cell Biology, 156(5):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/5/768.1.

Wells:2002:PSA

[Wel02-44] William A. Wells. Plasmid segregation actin’ up. Journal of Cell Biology, 158(2):??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/2/194.1.

Wells:2002:PMA

[Wel02-45] William A. Wells. Polo meets the APC. Journal of Cell Biology, 156(1):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/1/9.1.

Wells:2002:PHH

[Wel02-46] William A. Wells. Put that hunchback over here. Journal of Cell Biology, 156(5):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/5/769.3.

Wells:2002:RRR

[Wel02-47] William A. Wells. The real role of reelin. Journal of Cell Biology, 159(1):??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/1/15.3.

Wells:2002:RRb

[Wel02-48] William A. Wells. Redox regulation. Journal of Cell Biology, 159(5):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/5/729.1.

Wells:2002:RD

[Wel02-49] William A. Wells. Repressing differentiation. Journal of Cell Biology, 159(3):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/3/390.2.

Wells:2002:RTD

[Wel02-50] William A. Wells. RNAi takes on DNA. Journal of Cell Biology, 158(6):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/6/992.2.

Wells:2002:SMM

[Wel02-51] William A. Wells. Send Me a Message: The FASEB Conference on Signal Transduction Salt Lake City, UT June 29–July 3, 2002. Journal of Cell Biology, 158(6):994–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
6/994.

Wells:2002:SH

[Wel02-52] William A. Wells. Squeezing for healing. Journal of Cell Biology, 156(1):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/1/9.3.

Wells:2002:STS

[Wel02-53] William A. Wells. Sticking together a signal. Journal of Cell Biology, 156(1):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/1/8.1.

Wells:2002:SP

[Wel02-54] William A. Wells. Sticky polarity. Journal of Cell Biology, 157(1):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/1/12.2.

Wells:2002:S

[Wel02-55] William A. Wells. Stop and die. Journal of Cell Biology, 158(4):??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/4/607.2.

Wells:2002:SC

[Wel02-56] William A. Wells. Stop That Cell. Journal of Cell Biology, 157(4):548–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/4/548.

Wells:2002:SS

[Wel02-57] William A. Wells. Switch to survival. Journal of Cell Biology, 159(3):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/3/390.3.

Wells:2002:SD

[Wel02-58] William A. Wells. Systemic destruction. Journal of Cell Biology, 156(5):??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/5/769.1.

Wells:2002:TE

[Wel02-59] William A. Wells. A titinic extension. Journal of Cell Biology, 158(6):??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/6/993.1.

Wells:2002:RRa

[Wel02-60] William A. Wells. To Rb or not to Rb. Journal of Cell Biology, 156(1):??, January 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/1/9.2.

Wells:2002:TEL

[Wel02-61] William A. Wells. The transcription–export link. Journal of Cell Biology, 157(5):??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/5/737.3.

Wells:2002:TS

[Wel02-62] William A. Wells. Transcriptional segregation. Journal of Cell Biology, 159(3):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/3/391.2.

Wells:2002:UCP

[Wel02-63] William A. Wells. An unlikely culprit in prion diseases. Journal of Cell Biology, 159(3):??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/3/391.1.

Wells:2002:VV

[Wel02-64] William A. Wells. Visualizing vesicles. Journal of Cell Biology, 156(3):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/3/410.2.

Wells:2002:WMR

[Wel02-65] William A. Wells. What makes Ran run? Journal of Cell Biology, 156(3):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/3/412.2.

Wells:2002:WMM

[Wel02-66] William A. Wells. When more is not more. Journal of Cell Biology, 159(5):??, December 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/5/728.1.

Wells:2002:DLW

[Wel02-67] William A. Wells. With Aβ, the danger lies within. Journal of Cell Biology, 156(3):??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/3/411.1.

Wells:2002:WDD

[Wel02-68] William A. Wells. Wnt defeats death. Journal of Cell Biology, 157(3):??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/3/344.2.

Wells:2002:YR

[Wel02-69] William A. Wells. The youth revolution. Journal of Cell Biology, 158(4):609–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/4/609.

Wells:2003:AR

[Wel03a] William A. Wells. Activation by reduction. Journal of Cell Biology, 162(2):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/2/169.2.

Wells:2003:AF

[Wel03b] William A. Wells. Anchored by feedback. Journal of Cell Biology, 161(6):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/6/1009.3.

Wells:2003:BA

[Wel03c] William A. Wells. Bacterial aging. Journal of Cell Biology, 160(3):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/3/286.2.

Wells:2003:BE

[Wel03d] William A. Wells. Barriers to export. Journal of Cell Biology, 163(4):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/4/689.3.

Wells:2003:BH

[Wel03e] William A. Wells. A boisterous histone. Journal of Cell Biology, 160(7):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/7/982.1.

Wells:2003:CGN

[Wel03f] William A. Wells. Calcium goes nuclear. Journal of Cell Biology, 161(4):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/4/669.1.

Wells:2003:CKV

[Wel03g] William A. Wells. Catenin keeps vesicles close. Journal of Cell Biology, 163(5):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/5/929.1.

Wells:2003:CA

[Wel03h] William A. Wells. Cautious apoptosis. Journal of Cell Biology, 163(4):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/4/688.2.

Wells:2003:CEG

[Wel03i] William A. Wells. Centromere errors get chewed out. Journal of Cell Biology, 162(6):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/6/960.2.

Wells:2003:CD

[Wel03j] William A. Wells. The checkpoint of death. Journal of Cell Biology, 162(4):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/4/533.2.

Wells:2003:CH

[Wel03k] William A. Wells. Chromosomes in the ‘hood. Journal of Cell Biology, 160(7):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/7/983.2.

Wells:2003:CWI

[Wel03l] William A. Wells. Civil war in the immune system. Journal of Cell Biology, 163(5):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/5/928.2.

Wells:2003:CRL

[Wel03m] William A. Wells. Cohesin resolves to let go. Journal of Cell Biology, 160(1):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/1/8.3.

Wells:2003:CM

[Wel03n] William A. Wells. Committing to mitosis. Journal of Cell Biology, 161(6):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/6/1009.2.

Wells:2003:DFI

[Wel03o] William A. Wells. Damage first; infect later. Journal of Cell Biology, 163(3):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/3/435.3.

Wells:2003:DSC

[Wel03p] William A. Wells. A daughter’s size is not critical. Journal of Cell Biology, 162(6):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/6/961.2.

Wells:2003:DAT

[Wel03q] William A. Wells. A disease of actin transport? Journal of Cell Biology, 163(4):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/4/689.1.

Wells:2003:DCS

[Wel03r] William A. Wells. Doing the CDC20 shuffle. Journal of Cell Biology, 162(2):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/2/169.3.

Wells:2003:EA

[Wel03s] William A. Wells. Exocytosis in action. Journal of Cell Biology, 160(5):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/5/627.3.

Wells:2003:GM

[Wel03t] William A. Wells. A graveyard for mRNA. Journal of Cell Biology, 161(4):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/4/669.3.

Wells:2003:HGF

[Wel03u] William A. Wells. Hearts grow with the flow. Journal of Cell Biology, 160(3):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/3/287.2.

Wells:2003:HBF

[Wel03v] William A. Wells. Held back by a finger. Journal of Cell Biology, 163(4):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/4/688.1.

Wells:2003:HBJ

[Wel03w] William A. Wells. Helicobacter breaks down junctions. Journal of Cell Biology, 161(6):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/6/1009.1.

Wells:2003:HD

[Wel03x] William A. Wells. Heterochromatingets dynamic. Journal of Cell Biology, 160(5):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/5/626.1.

Wells:2003:HS

[Wel03y] William A. Wells. Hide and seek. Journal of Cell Biology, 160(7):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/7/982.3.

Wells:2003:HOM

[Wel03z] William A. Wells. How old mothers can have young offspring. Journal of Cell Biology, 160(7):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/7/982.2.

Wells:2003:HRS

[Wel03-27] William A. Wells. How to renew a stem cell. Journal of Cell Biology, 163(3):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/3/435.1.

Wells:2003:IC

[Wel03-28] William A. Wells. The importance of catastrophe. Journal of Cell Biology, 160(1):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/1/8.1.

Wells:2003:BNK

[Wel03-29] William A. Wells. Is Botryllus a natural killer? Journal of Cell Biology, 160(3):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/3/286.1.

Wells:2003:KAO

[Wel03-30] William A. Wells. A kinase that activates by occupancy. Journal of Cell Biology, 163(3):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/3/434.2.

Wells:2003:LF

[Wel03-31] William A. Wells. Latching onto fibrinogen. Journal of Cell Biology, 163(3):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/3/435.2.

Wells:2003:LBN

[Wel03-32] William A. Wells. Life on a bed of needles. Journal of Cell Biology, 160(5):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/5/626.2.

Wells:2003:LSF

[Wel03-33] William A. Wells. Linking spindle to furrow. Journal of Cell Biology, 160(5):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/5/627.2.

Wells:2003:LTG

[Wel03-34] William A. Wells. A loss, or two gains? Journal of Cell Biology, 162(2):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/2/168.1.

Wells:2003:MS

[Wel03-35] William A. Wells. Making stripes. Journal of Cell Biology, 160(7):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/7/983.1.

Wells:2003:MT

[Wel03-36] William A. Wells. Managing tension. Journal of Cell Biology, 163(5):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/5/928.1.

Wells:2003:MSO

[Wel03-37] William A. Wells. Metastasizing in search of oxygen. Journal of Cell Biology, 161(4):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/4/669.2.

Wells:2003:MC

[Wel03-38] William A. Wells. Movement in Colorado. Journal of Cell Biology, 160(7):985–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/7/985.

Wells:2003:MMW

[Wel03-39] William A. Wells. My mother, the wave. Journal of Cell Biology, 162(4):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/4/533.1.

Wells:2003:NPF

[Wel03-40] William A. Wells. Nervous picket fences. Journal of Cell Biology, 162(2):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/2/168.2.

Wells:2003:NS

[Wel03-41] William A. Wells. Neurons that swell. Journal of Cell Biology, 163(1):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/1/13.2.

Wells:2003:NGT

[Wel03-42] William A. Wells. Notch gets transformed. Journal of Cell Biology, 163(4):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/4/689.2.

Wells:2003:OCE

[Wel03-43] William A. Wells. One crossover is enough. Journal of Cell Biology, 163(1):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/1/12.1.

Wells:2003:OW

[Wel03-44] William A. Wells. Oscillating wildly. Journal of Cell Biology, 160(1):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/1/9.1.

Wells:2003:PAP

[Wel03-45] William A. Wells. Passing along polarity. Journal of Cell Biology, 163(3):??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/3/434.1.

Wells:2003:PM

[Wel03-46] William A. Wells. Poised to migrate. Journal of Cell Biology, 160(3):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/3/287.3.

Wells:2003:PFB

[Wel03-47] William A. Wells. Polarity feeds back. Journal of Cell Biology, 160(5):??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/5/627.1.

Wells:2003:PAA

[Wel03-48] William A. Wells. Prions by activation not amyloid. Journal of Cell Biology, 163(1):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/1/13.1.

Wells:2003:PB

[Wel03-49] William A. Wells. Pulling backwards. Journal of Cell Biology, 162(4):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/4/532.1.

Wells:2003:PP

[Wel03-50] William A. Wells. Pulling to the protease. Journal of Cell Biology, 162(6):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/6/961.3.

Wells:2003:RRV

[Wel03-51] William A. Wells. Ready to replicate? So vesiculate! Journal of Cell Biology, 163(1):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/1/13.3.

Wells:2003:RD

[Wel03-52] William A. Wells. Replicating dominos. Journal of Cell Biology, 160(3):??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/3/287.1.

Wells:2003:RF

[Wel03-53] William A. Wells. Reprogrammed by a frog. Journal of Cell Biology, 162(4):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/4/532.2.

Wells:2003:S

[Wel03-54] William A. Wells. S is for sticky. Journal of Cell Biology, 163 (5):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/5/929.2.

Wells:2003:SCb

[Wel03-55] William A. Wells. Selectins with a catch. Journal of Cell Biology, 161(4):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/4/668.2.

Wells:2003:SLC

[Wel03-56] William A. Wells. The sex–love connection. Journal of Cell Biology, 163(1):??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/1/12.2.

Wells:2003:SMM

[Wel03-57] William A. Wells. Small minerals make strong bones. Journal of Cell Biology, 161(4):??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/4/668.1.

Wells:2003:ST

[Wel03-58] William A. Wells. Sticky transmission. Journal of Cell Biology, 162(2):??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/2/169.1.

Wells:2003:SS

[Wel03-59] William A. Wells. Survival is sweet. Journal of Cell Biology, 162(6):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/162/6/961.1.

Wells:2003:SHT

[Wel03-60] William A. Wells. Surviving heat through destruction. Journal of Cell Biology, 162(4):??, August 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/162/4/532.3.

Wells:2003:TM

[Wel03-61] William A. Wells. A trigger for myelination. Journal of Cell Biology, 160(1):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/1/9.2.

Wells:2003:TGA

[Wel03-62] William A. Wells. Tubes go apical. Journal of Cell Biology, 160(1):??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/1/8.2.

Wells:2003:TCC

[Wel03-63] William A. Wells. A tumorigenic culprit in ES cells. Journal of Cell Biology, 161(6):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/6/1008.1.

Wells:2003:WM

[Wel03-64] William A. Wells. WASPs with memory. Journal of Cell Biology, 161(6):??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/161/6/1008.2.

Wells:2003:WVE

[Wel03-65] William A. Wells. When is a virus an exosome? Journal of Cell Biology, 162(6):??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/6/960.1.

Wells:2003:WDD

[Wel03-66] William A. Wells. Worms delay in deep sleep. Journal of Cell Biology, 163(5):??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/163/5/929.3.

Wells:2004:AS

[Wel04a] William A. Wells. Actin as stairmaster. Journal of Cell Biology, 164(7):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/7/953.3.

Wells:2004:ANF

[Wel04b] William A. Wells. All in the (network) family. Journal of Cell Biology, 164(7):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/7/952.2.

Wells:2004:AM

[Wel04c] William A. Wells. An anchor for motoring. Journal of Cell Biology, 165(3):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/3/294.1.

Wells:2004:ALA

[Wel04d] William A. Wells. Arg links actin and microtubules. Journal of Cell Biology, 165(3):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/3/294.2.

Wells:2004:BWM

[Wel04e] William A. Wells. Backward waves move cells forward. Journal of Cell Biology, 164(5):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/5/634.1.

Wells:2004:BD

[Wel04f] William A. Wells. Bacteria diversify. Journal of Cell Biology, 167(5):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/5/807.1.

Wells:2004:BIT

[Wel04g] William A. Wells. Being intolerant of tumors. Journal of Cell Biology, 165(2):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/2/165.2.

Wells:2004:BCG

[Wel04h] William A. Wells. Brain construction goes straight. Journal of Cell Biology, 165(1):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/1/15.2.

Wells:2004:BR

[Wel04i] William A. Wells. Bugs to the rescue. Journal of Cell Biology, 166(4):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/4/444.1.

Wells:2004:CSa

[Wel04j] William A. Wells. A caspase for stress. Journal of Cell Biology, 165(3):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/3/295.1.

Wells:2004:CC

[Wel04k] William A. Wells. Caterpillar compaction. Journal of Cell Biology, 166(1):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/1/9.1.

Wells:2004:CR

[Wel04l] William A. Wells. Channels rush in. Journal of Cell Biology, 166(4):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/4/445.1.

Wells:2004:CFI

[Wel04m] William A. Wells. Checking folding inside and out. Journal of Cell Biology, 165(1):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/1/14.1.

Wells:2004:CUM

[Wel04n] William A. Wells. A circular uncoating mechanism. Journal of Cell Biology, 164(1):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/1/8.2.

Wells:2004:CSAa

[Wel04o] William A. Wells. Clathrin starts off alone. Journal of Cell Biology, 166(6):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/6/763.3.

Wells:2004:CGP

[Wel04p] William A. Wells. Cohesin gets pushed around. Journal of Cell Biology, 166(2):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/2/154.1.

Wells:2004:CSb

[Wel04q] William A. Wells. Corralled by septin. Journal of Cell Biology, 166(4):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/4/445.2.

Wells:2004:DL

[Wel04r] William A. Wells. Destroy and localize. Journal of Cell Biology, 167(5):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/5/806.3.

Wells:2004:DSR

[Wel04s] William A. Wells. Diabetes as a stress response. Journal of Cell Biology, 167(3):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/3/397.1.

Wells:2004:JY

[Wel04t] William A. Wells. Die — but not just yet. Journal of Cell Biology, 164(5):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/5/635.3.

Wells:2004:DYA

[Wel04u] William A. Wells. Do yeast apoptose? Journal of Cell Biology, 166(3):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/3/304.2.

Wells:2004:DDM

[Wel04v] William A. Wells. Don’t digest the messenger. Journal of Cell Biology, 165(1):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/1/14.2.

Wells:2004:DTA

[Wel04w] William A. Wells. Dormant tumors awaken. Journal of Cell Biology, 167(3):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/396.2.

Wells:2004:DMG

[Wel04x] William A. Wells. Dystrophic muscles get overexcited. Journal of Cell Biology, 166(6):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/6/763.1.

Wells:2004:ES

[Wel04y] William A. Wells. Elimination by shedding. Journal of Cell Biology, 167(5):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/5/806.1.

Wells:2004:EGS

[Wel04z] William A. Wells. Endocytosis gets squeezed. Journal of Cell Biology, 167(1):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/1/15.3.

Wells:2004:FPA

[Wel04-27] William A. Wells. Flat-out pushing by actin. Journal of Cell Biology, 166(2):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/2/154.2.

Wells:2004:FFS

[Wel04-28] William A. Wells. Flipping the fusion switch. Journal of Cell Biology, 164(5):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/5/634.2.

Wells:2004:FF

[Wel04-29] William A. Wells. Flow to the front. Journal of Cell Biology, 167(1):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/1/15.1.

Wells:2004:FD

[Wel04-30] William A. Wells. Foci of decay. Journal of Cell Biology, 165(1):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/1/15.1.

Wells:2004:FSa

[Wel04-31] William A. Wells. Formin a shmoo. Journal of Cell Biology, 165(1):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/1/15.3.

Wells:2004:FAJ

[Wel04-32] William A. Wells. Formin’ adherens junctions. Journal of Cell Biology, 164(1):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/1/9.2.

Wells:2004:FSb

[Wel04-33] William A. Wells. Freedom from substrate. Journal of Cell Biology, 166(3):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/3/305.2.

Wells:2004:GDC

[Wel04-34] William A. Wells. Get out! And don’t come back! Journal of Cell Biology, 165(2):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/2/165.3.

Wells:2004:GNP

[Wel04-35] William A. Wells. Giving neural precursors the finger. Journal of Cell Biology, 164(1):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/1/9.1.

Wells:2004:HCS

[Wel04-36] William A. Wells. How a channel springs open. Journal of Cell Biology, 166(1):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/1/8.2.

Wells:2004:HBG

[Wel04-37] William A. Wells. How to become a generalist again. Journal of Cell Biology, 165(4):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/4/457.1.

Wells:2004:HTT

[Wel04-38] William A. Wells. How transmembranes tell of their approach. Journal of Cell Biology, 164(7):??, March 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/7/
952.1.

Wells:2004:ISM

[Wel04-39] William A. Wells. Immune synapses make a choice. Journal of Cell Biology, 167(1):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/1/15.2.

Wells:2004:IMC

[Wel04-40] William A. Wells. Integrins make a choice. Journal of Cell Biology, 166(3):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/3/305.1.

Wells:2004:LG

[Wel04-41] William A. Wells. License to glue. Journal of Cell Biology, 166(4):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/4/445.3.

Wells:2004:MW

[Wel04-42] William A. Wells. Making waves. Journal of Cell Biology, 165(4):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/4/457.3.

Wells:2004:MN

[Wel04-43] William A. Wells. Man the nanoscopes. Journal of Cell Biology, 164(3):337–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/3/337.

Wells:2004:MWP

[Wel04-44] William A. Wells. Me write pretty one day. Journal of Cell Biology, 165(6):757–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/6/757.

Wells:2004:MBD

[Wel04-45] William A. Wells. A message before dying. Journal of Cell Biology, 166(1):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/1/9.2.

Wells:2004:MPL

[Wel04-46] William A. Wells. Metabolism as a production line. Journal of Cell Biology, 165(4):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/4/457.2.

Wells:2004:MI

[Wel04-47] William A. Wells. Mini inhibition. Journal of Cell Biology, 166(2):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/2/155.2.

Wells:2004:MMA

[Wel04-48] William A. Wells. Monomeric myosin anchoring. Journal of Cell Biology, 165(3):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/3/295.2.

Wells:2004:MWS

[Wel04-49] William A. Wells. Moving while sticking. Journal of Cell Biology, 166(1):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/1/8.1.

Wells:2004:MPQ

[Wel04-50] William A. Wells. A muscular path to quiescence. Journal of Cell Biology, 166(3):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/3/304.1.

Wells:2004:MCA

[Wel04-51] William A. Wells. Myc conquers all. Journal of Cell Biology, 165(2):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/2/164.2.

Wells:2004:NIA

[Wel04-52] William A. Wells. Neurons inch along. Journal of Cell Biology, 167(3):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/396.3.

Wells:2004:OMG

[Wel04-53] William A. Wells. Oxidation makes a germline. Journal of Cell Biology, 164(1):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/164/1/9.3.

Wells:2004:PCL

[Wel04-54] William A. Wells. Picking the clathrin lock. Journal of Cell Biology, 167(5):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/5/807.2.

Wells:2004:PM

[Wel04-55] William A. Wells. Polarization by migration. Journal of Cell Biology, 167(1):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/1/14.2.

Wells:2004:PP

[Wel04-56] William A. Wells. Prions in packages. Journal of Cell Biology, 166(1):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/1/9.3.

Wells:2004:POP

[Wel04-57] William A. Wells. Protein-only prions. Journal of Cell Biology, 166(4):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/4/444.2.

Wells:2004:PTP

[Wel04-58] William A. Wells. Protein translocation pulls the plug. Journal of Cell Biology, 164(1):??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/164/1/8.1.

Wells:2004:PGS

[Wel04-59] William A. Wells. Pulling out gene silencers. Journal of Cell Biology, 164(5):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/5/635.2.

Wells:2004:PSS

[Wel04-60] William A. Wells. Pulling the spindle into shape. Journal of Cell Biology, 165(4):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/4/456.2.

Wells:2004:RAP

[Wel04-61] William A. Wells. Rafting across polarized cells. Journal of Cell Biology, 165(2):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/2/164.1.

Wells:2004:RV

[Wel04-62] William A. Wells. Reprogramming by virus. Journal of Cell Biology, 166(2):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/2/155.3.

Wells:2004:RC

[Wel04-63] William A. Wells. Rewiring the cell. Journal of Cell Biology, 167(1):??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/1/14.1.

Wells:2004:RPa

[Wel04-64] William A. Wells. Rho in podosomes. Journal of Cell Biology, 166(3):??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/3/305.3.

Wells:2004:RPb

[Wel04-65] William A. Wells. Rocking on Pins. Journal of Cell Biology, 167(5):??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/5/806.2.

Wells:2004:SD

[Wel04-66] William A. Wells. Sensing detachment. Journal of Cell Biology, 165(3):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/3/295.3.

Wells:2004:STA

[Wel04-67] William A. Wells. Signaling together apart. Journal of Cell Biology, 167(3):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/3/396.1.

Wells:2004:SSW

[Wel04-68] William A. Wells. Slow but steady wins the drug race. Journal of Cell Biology, 166(6):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/6/763.2.

Wells:2004:SFE

[Wel04-69] William A. Wells. A SNARE for fast endocytosis. Journal of Cell Biology, 167(3):??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/3/397.2.

Wells:2004:SBN

[Wel04-70] William A. Wells. Snaring bugs in NETs. Journal of Cell Biology, 164(7):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/7/953.1.

Wells:2004:SP

[Wel04-71] William A. Wells. Solo polarization. Journal of Cell Biology, 164(5):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/5/635.1.

Wells:2004:SLT

[Wel04-72] William A. Wells. Structure with a light touch. Journal of Cell Biology, 166(6):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/6/762.1.

Wells:2004:SR

[Wel04-73] William A. Wells. Surviving to remember. Journal of Cell Biology, 164(7):??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/7/953.2.

Wells:2004:TT

[Wel04-74] William A. Wells. Telomeres together. Journal of Cell Biology, 165(2):??, April 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/2/165.1.

Wells:2004:TBT

[Wel04-75] William A. Wells. Tilt back to turn left. Journal of Cell Biology, 165(4):??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/4/456.1.

Wells:2004:TNL

[Wel04-76] William A. Wells. Tumors need less regulation. Journal of Cell Biology, 166(6):??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/6/762.2.

Wells:2004:XMP

[Wel04-77] William A. Wells. Xylogen makes plant channels. Journal of Cell Biology, 166(2):??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/2/155.1.

Welm:2002:IDF

[WFC^{+}02] Bryan E. Welm, Kevin W. Freeman, Mercy Chen, Alejandro Contreras, David M. Spencer, and Jeffrey M. Rosen. Inducible dimerization of FGFR1. Journal of Cell Biology, 157 (4):703–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/4/703.

Wang:2001:AMR

[WFF^{+}01] Rong Wang, Linda D. Ferrell, Saadia Faouzi, Jacquelyn J. Maher, and J. Michael Bishop. Activation of the Met Receptor by Cell Attachment Induces and Sustains Hepatocellular Carcinomas in Transgenic Mice. Journal of Cell Biology, 153 (5):1023–??, May 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/5/1023.

Wang:2004:TCI

[WFI^{+}04] Hui Wang, Weili Fu, Jae Hong Im, Zengyi Zhou, Samuel A. Santoro, Vandana Iyer, C. Mike DiPersio, Qian-Chun Yu, Vito Quaranta, Abu Al-Mehdi, and Ruth J. Muschel. Tumor cell α_{3}β_{1} integrin and vascular laminin-5 mediate pulmonary arrest and metastasis. Journal of Cell Biology, 164 (6):935–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/6/935.

Wrabetz:2000:PGO

[WFQ^{+}00] Lawrence Wrabetz, Maria Laura Feltri, Angelo Quattrini, Daniele Imperiale, Stefano Previtali, Maurizio D’Antonio, Rudolf Martini, Xinghua Yin, Bruce D. Trapp, Lei Zhou, Shing-Yan Chiu, and Albee Messing. P_{0} glycoprotein overexpression causes congenital hypomyelination of peripheral nerves. Journal of Cell Biology, 148(5):1021–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/148/5/
1021.

Wunderlich:2001:NKP

[WFT^{+}01] Winfried Wunderlich, Irene Fialka, David Teis, Arno Alpi, Andrea Pfeifer, Robert G. Parton, Friedrich Lottspeich, and Lukas A. Huber. A novel 14-kilodalton protein interacts with the mitogen-activated protein kinase scaffold Mp1 on a late endosomal/lysosomal compartment. Journal of Cell Biology, 152(4):765–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/152/4/765.

Wong:2003:AIM

[WG03] Alice S. T. Wong and Barry M. Gumbiner. Adhesionindependent mechanism for suppression of tumor cell invasion by E-cadherin. Journal of Cell Biology, 161(6):1191–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
161/6/1191.

Whitlock:2000:DRI

[WGF^{+}00] Ben B. Whitlock, Shyra Gardai, Valerie Fadok, Donna Bratton, and Peter M. Henson. Differential Roles for α_{M} β_{2} Integrin Clustering or Activation in the Control of Apoptosis via Regulation of Akt and ERK Survival Mechanisms. Journal of Cell Biology, 151(6):1305–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/6/1305.

Wang:2000:CRA

[WGP^{+}00] Hui-Jun Wang, Ginette Guay, Liviu Pogan, Remy Sauvé, and Ivan R. Nabi. Calcium Regulates the Association between Mitochondria and a Smooth Subdomain of the Endoplasmic Reticulum. Journal of Cell Biology, 150(6):1489– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/6/1489.

Waterhouse:2001:CCM

[WGvA^{+}01] Nigel J. Waterhouse, Joshua C. Goldstein, Oliver von Ahsen, Martin Schuler, Donald D. Newmeyer, and Douglas R. Green. Cytochrome C maintains mitochondrial transmembrane potential and ATP generation after outer mitochondrial membrane permeabilization during the apoptotic process. Journal of Cell Biology, 153(2):319–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/153/2/319.

Wagner:2003:MVE

[WH03] Wolfgang Wagner and John A. Hammer. Myosin V and the endoplasmic reticulum. Journal of Cell Biology, 163(6): 1193–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/6/1193.

Wilhelm:2003:CEB

[WHAH03] James E. Wilhelm, Meredith Hilton, Quinlan Amos, and William J. Henzel. Cup is an eIF4E binding protein required for both the translational repression of oskar and the recruitment of Barentsz. Journal of Cell Biology, 163(6): 1197–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/6/1197.

Webber:2004:CPO

[WHB04] Hayley A. Webber, Louisa Howard, and Sharon E. Bickel. The cohesion protein ORD is required for homologue bias during meiotic recombination. Journal of Cell Biology, 164 (6):819–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/6/819.

Wiesner:2003:BMA

[WHD^{+}03] Sebastian Wiesner, Emmanuele Helfer, Dominique Didry, Guylaine Ducouret, Françoise Lafuma, Marie-France Carlier, and Dominique Pantaloni. A biomimetic motility assay provides insight into the mechanism of actin-based motility. Journal of Cell Biology, 160(3):387–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/3/
387.

Wada:2002:CGP

[WHM^{+}02] Hiromichi Wada, Koji Hasegawa, Tatsuya Morimoto, Tsuyoshi Kakita, Tetsuhiko Yanazume, Mitsuru Abe, and Shigetake Sasayama. Calcineurin-GATA-6 pathway is involved in smooth muscle-specific transcription. Journal of Cell Biology, 156(6):983–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/156/6/983.

Widlund:2002:CIM

[WHP^{+}02] Hans R. Widlund, Martin A. Horstmann, E. Roydon Price, Junqing Cui, Stephen L. Lessnick, Min Wu, Xi He, and David E. Fisher. β-catenin-induced melanoma growth requires the downstream target Microphthalmia-associated transcription factor. Journal of Cell Biology, 158(6):1079– ??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/158/6/1079.

Wittchen:2000:EEA

[WHS00] Erika S. Wittchen, Julie Haskins, and Bruce R. Stevenson. Exogenous Expression of the Amino–Terminal Half of the Tight Junction Protein Zo-3 Perturbs Junctional Complex Assembly. Journal of Cell Biology, 151(4):825–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
4/825.

Wilcke:2000:RRC

[WJG^{+}00] Mona Wilcke, Ludger Johannes, Thierry Galli, Véronique Mayau, Bruno Goud, and Jean Salamero. Rab11 Regulates the Compartmentalization of Early Endosomes Required for Efficient Transport from Early Endosomes to the Trans-Golgi Network. Journal of Cell Biology, 151(6):1207–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
6/1207.

Wigge:2001:NCS

[WK01] Philip A. Wigge and John V. Kilmartin. The Ndc80p Complex from Saccharomyces cerevisiae Contains Conserved Centromere Components and Has a Function in Chromosome Segregation. Journal of Cell Biology, 152(2):349–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
2/349.

Wang:2002:RAS

[WK02] Wei Wang and Thorsten Kirsch. Retinoic acid stimulates annexin-mediated growth plate chondrocyte mineralization. Journal of Cell Biology, 157(6):1061–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/6/1061.

Wysocki:2004:YCD

[WK04] Robert Wysocki and Stephen J. Kron. Yeast cell death during DNA damage arrest is independent of caspase or reactive oxygen species. Journal of Cell Biology, 166(3):311–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
3/311.

Wimmer-Kleikamp:2004:RER

[WKJS^{+}04] Sabine H. Wimmer-Kleikamp, Peter W. Janes, Anthony Squire, Philippe I. H. Bastiaens, and Martin Lackmann. Recruitment of Eph receptors into signaling clusters does not require ephrin contact. Journal of Cell Biology, 164(5): 661–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/5/661.

Weed:2000:CLS

[WKS^{+}00] Scott A. Weed, Andrei V. Karginov, Dorothy A. Schafer, Alissa M. Weaver, Andrew W. Kinley, John A. Cooper, and J. Thomas Parsons. Cortactin localization to sites of actin assembly in Lamellipodia requires interactions with F-actin and the Arp2/3 complex. Journal of Cell Biology, 151(1): 29–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/1/29.

Wang:2002:DKI

[WKYC02] Won-Jing Wang, Jean-Cheng Kuo, Chung-Chen Yao, and Ruey-Hwa Chen. DAP-kinase induces apoptosis by suppressing integrin activity and disrupting matrix survival signals. Journal of Cell Biology, 159(1):169–??, October 2002. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/159/1/
169.

Weiss:2002:SCM

[WKZ^{+}02] Eric L. Weiss, Cornelia Kurischko, Chao Zhang, Kevan Shokat, David G. Drubin, and Francis C. Luca. The Saccharomyces cerevisiae Mob2p–Cbk1p kinase complex promotes polarized growth and acts with the mitotic exit network to facilitate daughter cell–specific localization of Ace2p transcription factor. Journal of Cell Biology, 158(5):885–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
158/5/885.

Wells:2003:SCa

[WL03] William A. Wells and Nicole LeBrasseur. Stories of Cells. Journal of Cell Biology, 160(4):466–??, February 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/4/
466.

Wells:2004:CP

[WL04] William A. Wells and Nicole LeBrasseur. Cells on Parade. Journal of Cell Biology, 164(4):480–??, February 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/164/4/

Wasiak:2002:E

[WLGP^{+}02] Sylwia Wasiak, Valerie Legendre-Guillemin, Rosa Puertollano, François Blondeau, Martine Girard, Elaine de Heuvel, Daniel Boismenu, Alexander W. Bell, Juan S. Bonifacino, and Peter S. McPherson. Enthoprotin. Journal of Cell Biology, 158(5):855–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/5/855.

Wissing:2004:AOR

[WLH^{+}04] Silke Wissing, Paula Ludovico, Eva Herker, Sabrina Büttner, Silvia M. Engelhardt, Thorsten Decker, Alexander Link, Astrid Proksch, Fernando Rodrigues, Manuela Corte-Real, Kai-Uwe Fröhlich, Joachim Manns, Céline Candé, Stephan J. Sigrist, Guido Kroemer, and Frank Madeo. An AIF orthologue regulates apoptosis in yeast. Journal of Cell Biology, 166(7):969–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/166/7/969.

Witke:2001:CCG

[WLKS01] Walter Witke, Wei Li, David J. Kwiatkowski, and Frederick S. Southwick. Comparisons of CapG and gelsolin-null macrophages. Journal of Cell Biology, 154(4):775–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/154/
4/775.

Woodring:2002:MFA

[WLO^{+}02] Pamela J. Woodring, E. David Litwack, Dennis D. M. O’Leary, Ginger R. Lucero, Jean Y. J. Wang, and Tony Hunter. Modulation of the F-actin cytoskeleton by c–Abl tyrosine kinase in cell spreading and neurite extension. Journal of Cell Biology, 156(5):879–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/156/5/879.

Wright:2004:DAA

[WLPD04] Kevin M. Wright, Michael W. Linhoff, Patrick Ryan Potts, and Mohanish Deshmukh. Decreased apoptosome activity with neuronal differentiation sets the threshold for strict IAP regulation of apoptosis. Journal of Cell Biology, 167(2):

303–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/2/303.

Wojcik:2001:DNM

[WLR01] Sonja M. Wojcik, Mary A. Longley, and Dennis R. Roop. Discovery of a novel murine keratin 6 (K6) isoform explains the absence of hair and nail defects in mice deficient for K6a and K6b. Journal of Cell Biology, 154(3):619–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/154/
3/619.

Wang:2004:CRP

[WLW^{+}04] Ying Jie Wang, Wen Hong Li, Jing Wang, Ke Xu, Ping Dong, Xiang Luo, and Helen L. Yin. Critical role of PIP5KIγ87 in InsP_{3}-mediated Ca^{2+} signaling. Journal of Cell Biology, 167(6):1005–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/6/1005.

Worthylake:2001:RRM

[WLWB01] Rebecca A. Worthylake, Sean Lemoine, Joanna M. Watson, and Keith Burridge. RhoA is required for monocyte tail retraction during transendothelial migration. Journal of Cell Biology, 154(1):147–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/1/147.

Weidinger:2003:WWA

[WM03] Gilbert Weidinger and Randall T. Moon. When Wnts antagonize Wnts. Journal of Cell Biology, 162(5):753–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
5/753.

Wang:2004:CDE

[WMA^{+}04a] Xiaodong Wang, Jeanne Matteson, Yu An, Bryan Moyer, Jin-San Yoo, Sergei Bannykh, Ian A. Wilson, John R. Riordan, and William E. Balch. COPII-dependent export of cystic fibrosis transmembrane conductance regulator from the ER uses a di-acidic exit code. Journal of Cell Biology, 167 (1):65–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/1/65.

Wang:2004:Cb

[WMA^{+}04b] Xiaodong Wang, Jeanne Matteson, Yu An, Bryan Moyer, Jin-San Yoo, Sergei Bannykh, Ian A. Wilson, John R. Riordan, and William E. Balch. Correction. Journal of Cell Biology, 167(3):567–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/3/567.

Wang:2003:HPA

[WMCW03] Li Wang, Alexey J. Merz, Kevin M. Collins, and William Wickner. Hierarchy of protein assembly at the vertex ring domain for yeast vacuole docking and fusion. Journal of Cell Biology, 160(3):365–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/3/365.

Wilson:2004:DOC

[WMG^{+}04] Christina A. Wilson, Diane D. Murphy, Benoit I. Giasson, Bin Zhang, John Q. Trojanowski, and Virginia M.-Y. Lee. Degradative organelles containing mislocalized α- and βsynuclein proliferate in presenilin-1 null neurons. Journal of Cell Biology, 165(3):335–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/165/3/335.

Wilhelm:2000:IRC

[WMHB^{+}00] James E. Wilhelm, Jennifer Mansfield, Nora Hom-Booher, Shengxian Wang, Christoph W. Turck, Tulle Hazelrigg, and Ronald D. Vale. Isolation of a Ribonucleoprotein Complex Involved in mRNA Localization in Drosophila Oocytes. Journal of Cell Biology, 148(3):427–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/148/3/427.

Woodring:2004:CAP

[WMJ^{+}04] Pamela J. Woodring, Jill Meisenhelder, Sam A. Johnson, Guo-Lei Zhou, Jeffrey Field, Kavita Shah, Friedhelm Bladt, Tony Pawson, Masaru Niki, Pier Paolo Pandolfi, Jean Y. J. Wang, and Tony Hunter. c-abl phosphorylates Dok1 to promote filopodia during cell spreading. Journal of Cell Biology, 165(4):493–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/4/493.

Wolf:2003:CMT

[WML^{+}03] Katarina Wolf, Irina Mazo, Harry Leung, Katharina Engelke, Ulrich H. von Andrian, Elena I. Deryugina, Alex Y. Strongin, Eva-B. Bröcker, and Peter Friedl. Compensation mechanism in tumor cell migration. Journal of Cell Biology, 160(2): 267–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/2/267.

Wenk:2000:TCS

[WMS00] Melissa B. Wenk, Kim S. Midwood, and Jean E. Schwarzbauer. Tenascin-c Suppresses Rho Activation. Journal of Cell Biology, 150(4):913–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/4/913.

Waites:2001:AMR

[WMT^{+}01] Clarissa L. Waites, Anand Mehta, Philip K. Tan, Gary Thomas, Robert H. Edwards, and David E. Krantz. An Acidic Motif Retains Vesicular Monoamine Transporter 2 on Large Dense Core Vesicles. Journal of Cell Biology, 152(6): 1159–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/6/1159.

Wood:2000:ADR

[WND^{+}00] Jonathan D. Wood, Frederick C. Nucifora, Kui Duan, Chuanyi Zhang, Jianxiang Wang, Yujin Kim, Gabriele Schilling, Nicoletta Sacchi, Johnson M. Liu, and Christopher A. Ross. Atrophin-1, the Dentato–Rubral and Pallido– Luysian atrophy gene product, interacts with Eto/Mtg8 in the nuclear matrix and represses transcription. Journal of Cell Biology, 150(5):939–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/5/939.

Watabe:2003:TRK

[WNM^{+}03] Tetsuro Watabe, Ayako Nishihara, Koichi Mishima, Jun Yamashita, Kiyoshi Shimizu, Keiji Miyazawa, Shin-Ichi Nishikawa, and Kohei Miyazono. TGF-β receptor kinase inhibitor enhances growth and integrity of embryonic stem cell–derived endothelial cells. Journal of Cell Biology, 163(6): 1303–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/6/1303.

Wordeman:2003:BNU

[Wor03] Linda Wordeman. Breathing down the neck of Unc104. Journal of Cell Biology, 163(4):693–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/4/693.

Ward:2001:MGS

[WPC^{+}01] Theresa H. Ward, Roman S. Polishchuk, Steve Caplan, Koret Hirschberg, and Jennifer Lippincott-Schwartz. Maintenance of Golgi structure and function depends on the integrity of ER export. Journal of Cell Biology, 155(4):557–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
4/557.

Walther:2002:CFN

[WPC^{+}02] Tobias C. Walther, Helen S. Pickersgill, Volker C. Cordes, Martin W. Goldberg, Terry D. Allen, Iain W. Mattaj, and Maarten Fornerod. The cytoplasmic filaments of the nuclear pore complex are dispensable for selective nuclear protein import. Journal of Cell Biology, 158(1):63–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/158/1/
63.

Wang:2000:VAS

[WPJ^{+}00] He Wang, David A. D. Parry, Leslie N. Jones, William W. Idler, Lyuben N. Marekov, and Peter M. Steinert. In Vitro Assembly and Structure of Trichocyte Keratin Intermediate Filaments. Journal of Cell Biology, 151(7):1459–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
151/7/1459.

Weber:2000:SFR

[WPM^{+}00] Thomas Weber, Francesco Parlati, James A. McNew, Robert J. Johnston, Benedikt Westermann, Thomas H.

Söllner, and James E. Rothman. Snarepins Are Functionally Resistant to Disruption by Nsf and αSNAP. Journal of Cell Biology, 149(5):1063–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/149/5/1063.

Wilson:2000:OFS

[WPO00] Bridget S. Wilson, Janet R. Pfeiffer, and Janet M. Oliver. Observing Fcϵ ri Signaling from the Inside of the Mast Cell Membrane. Journal of Cell Biology, 149(5):1131–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/149/
5/1131.

Wilson:2001:HRM

[WPS^{+}01] Bridget S. Wilson, Janet R. Pfeiffer, Zurab Surviladze, Elizabeth A. Gaudet, and Janet M. Oliver. High resolution mapping of mast cell membranes reveals primary and secondary domains of FcϵRI and LAT. Journal of Cell Biology, 154(3): 645–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/3/645.

Weston:2000:RSP

[WRCU00] Andrea D. Weston, Vicki Rosen, Roshantha A. S. Chandraratna, and T. Michael Underhill. Regulation of Skeletal Progenitor Differentiation by the Bmp and Retinoid Signaling Pathways. Journal of Cell Biology, 148(4):679–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
148/4/679.

Wodarz:2000:DAP

[WRGK00] Andreas Wodarz, Andreas Ramrath, Alexandra Grimm, and Elisabeth Knust. Drosophila Atypical Protein Kinase C Associates with Bazooka and Controls Polarity of Epithelia and Neuroblasts. Journal of Cell Biology, 150(6):1361– ??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/6/1361.

Wodarz:2004:C

[WRGK04] Andreas Wodarz, Andreas Ramrath, Alexandra Grimm, and Elisabeth Knust. Correction. Journal of Cell Biology, 165 (4):591–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/4/591.

Ward:2004:RCS

[WRSMO^{+}04] Irene M. Ward, Bernardo Reina-San-Martin, Alexandru Olaru, Kay Minn, Koji Tamada, Julie S. Lau, Marilia Cascalho, Lieping Chen, Andre Nussenzweig, Ferenc Livak, Michel C. Nussenzweig, and Junjie Chen. 53BP1 is required for class switch recombination. Journal of Cell Biology, 165 (4):459–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/4/459.

Wesley:2000:ANL

[WS00] Cedric S. Wesley and Lino Saez. Analysis of Notch Lacking the Carboxyl Terminus Identified in Drosophila Embryos. Journal of Cell Biology, 149(3):683–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/149/3/683.

Watson:2001:LRM

[WSC^{+}01] Robert T. Watson, Satoshi Shigematsu, Shian-Huey Chiang, Silvia Mora, Makoto Kanzaki, Ian G. Macara, Alan R. Saltiel, and Jeffrey E. Pessin. Lipid raft microdomain compartmentalization of TC10 is required for insulin signaling and GLUT4 translocation. Journal of Cell Biology, 154(4): 829–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/4/829.

Walter:2003:COH

[WSC^{+}03] Joachim Walter, Lothar Schermelleh, Marion Cremer, Satoshi Tashiro, and Thomas Cremer. Chromosome order in HeLa cells changes during mitosis and early G1, but is stably maintained during subsequent interphase stages. Journal of Cell Biology, 160(5):685–??, March 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/5/685.

Waterman-Storer:2000:MRA

[WSDW^{+}00] Clare Waterman-Storer, Devin Y. Duey, Kari L. Weber, John Keech, Richard E. Cheney, E. D. Salmon, and William M. Bement. Microtubules Remodel Actomyosin Networks in Xenopus Egg Extracts via Two Mechanisms of F–Actin Transport. Journal of Cell Biology, 150(2):361–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/2/361.

Wurmser:2000:NCV

[WSE00] Andrew E. Wurmser, Trey K. Sato, and Scott D. Emr. New Component of the Vacuolar Class C–Vps Complex Couples Nucleotide Exchange on the Ypt7 Gtpase to Snare– Dependent Docking and Fusion. Journal of Cell Biology, 151(3):551–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/3/551.

Weins:2001:DSD

[WSF^{+}01] Astrid Weins, Karin Schwarz, Christian Faul, Laura Barisoni, Wolfgang A. Linke, and Peter Mundel. Differentiation- and stress-dependent nuclear cytoplasmic redistribution of myopodin, a novel actin-bundling protein. Journal of Cell Biology, 155(3):393–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/3/393.

Wang:2000:TSG

[WSFN00] Wei Wang, Michael Sacher, and Susan Ferro-Novick. Trapp Stimulates Guanine Nucleotide Exchange on Ypt1p. Journal of Cell Biology, 151(2):289–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/2/289.

Wu:2004:SDC

[WSH^{+}04] Victoria M. Wu, Joost Schulte, Alexander Hirschi, Ulrich Tepass, and Greg J. Beitel. Sinuous is a Drosophila claudin required for septate junction organization and epithelial tube size control. Journal of Cell Biology, 164(2):313–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
2/313.

Wang:2003:YHV

[WSK^{+}03] Chao-Wen Wang, Per E. Stromhaug, Emily J. Kauffman, Lois S. Weisman, and Daniel J. Klionsky. Yeast homotypic vacuole fusion requires the Ccz1–Mon1 complex during the tethering/docking stage. Journal of Cell Biology, 163(5): 973–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/5/973.

Wu:2000:TPA

[WSL^{+}00] Yan Ping Wu, Chia-Jen Siao, Weiquan Lu, Tsung-Chang Sung, Michael A. Frohman, Peter Milev, Thomas H. Bugge, Jay L. Degen, Joel M. Levine, Richard U. Margolis, and Stella E. Tsirka. The tissue plasminogen activator (Tpa/ plasmin) extracellular proteolytic system regulates seizureinduced hippocampal mossy fiber outgrowth through a proteoglycan substrate. Journal of Cell Biology, 148(6):1295–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/148/6/1295.

Wu:2001:ECE

[WSL^{+}01] Songwei Wu, Jose Sangerman, Ming Li, George H. Brough, Steven R. Goodman, and Troy Stevens. Essential control of an endothelial cell I_{SOC} by the spectrin membrane skeleton. Journal of Cell Biology, 154(6):1225–??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/154/
6/1225.

Wiseman:2003:SSI

[WSL^{+}03] Bryony S. Wiseman, Mark D. Sternlicht, Leif R. Lund, Caroline M. Alexander, Joni Mott, Mina J. Bissell, Paul Soloway, Shigeyoshi Itohara, and Zena Werb. Site-specific inductive and inhibitory activities of MMP-2 and MMP-3 orchestrate mammary gland branching morphogenesis. Journal of Cell Biology, 162(6):1123–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/6/1123.

Wojciak-Stothard:2003:SSI

[WSR03] Beata Wojciak-Stothard and Anne J. Ridley. Shear stress– induced endothelial cell polarization is mediated by Rho and Rac but not Cdc42 or PI 3-kinases. Journal of Cell Biology, 161(2):429–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/161/2/429.

Wang:2004:PLN

[WSS^{+}04] Jayson Wang, Carol Shiels, Peter Sasieni, Pei Jun Wu, Suhail A. Islam, Paul S. Freemont, and Denise Sheer. Promyelocytic leukemia nuclear bodies associate with transcriptionally active genomic regions. Journal of Cell Biology, 164(4):515–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/4/515.

Wehling:2001:NOS

[WST01] Michelle Wehling, Melissa J. Spencer, and James G. Tidball. A nitric oxide synthase transgene ameliorates muscular dystrophy in mdx mice. Journal of Cell Biology, 155(1): 123–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/1/123.

Watts:2000:HRD

[WSW^{+}00] Norman R. Watts, Dan L. Sackett, Rita D. Ward, Mill W. Miller, Paul T. Wingfield, Stephen S. Stahl, and Alasdair C. Steven. HIV-1 Rev Depolymerizes Microtubules to Form Stable Bilayered Rings. Journal of Cell Biology, 150(2):349–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
150/2/349.

Wang:2004:Ca

[WSWM04a] Yanzhuang Wang, Ayano Satoh, Graham Warren, and Hemmo H. Meyer. Correction. Journal of Cell Biology, 166 (3):433–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/166/3/433.

Wang:2004:VAD

[WSWM04b] Yanzhuang Wang, Ayano Satoh, Graham Warren, and Hemmo H. Meyer. VCIP135 acts as a deubiquitinating enzyme during p97–p47-mediated reassembly of mitotic Golgi fragments. Journal of Cell Biology, 164(7):973–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
7/973.

Wedlich-Soldner:2004:RCP

[WSWSL04] Roland Wedlich-Soldner, Stephanie C. Wai, Thomas Schmidt, and Rong Li. Robust cell polarity is a dynamic state established by coupling transport and GTPase signaling. Journal of Cell Biology, 166(6):889–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/6/889.

Webb:2001:PAI

[WTG01] Donna J. Webb, Keena S. Thomas, and Steven L. Gonias. Plasminogen Activator Inhibitor 1 Functions as a Urokinase Response Modifier at the Level of Cell Signaling and Thereby Promotes Mcf-7 Cell Growth. Journal of Cell Biology, 152(4): 741–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/4/741.

Wang:2003:EFP

[WVBY^{+}03] Rui-An Wang, Ratna K. Vadlamudi, Rozita Bagheri-Yarmand, Iwan Beuvink, Nancy E. Hynes, and Rakesh Kumar. Essential functions of p21-activated kinase 1 in morphogenesis and differentiation of mammary glands. Journal of Cell Biology, 161(3):583–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/3/583.

Wahlstrom:2001:TRA

[WVY^{+}01] Gudrun Wahlström, Maria Vartiainen, Lumi Yamamoto, Pieta K. Mattila, Pekka Lappalainen, and Tapio I. Heino. Twinfilin is required for actin-dependent developmental processes in Drosophila. Journal of Cell Biology, 155(5):787– ??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/5/787.

Werner:2002:IEM

[WW02] Erica Werner and Zena Werb. Integrins engage mitochondrial function for signal transduction by a mechanism dependent on Rho GTPases. Journal of Cell Biology, 158(2):357–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
158/2/357.

Wucherpfennig:2003:RDR

[WWBGG03] Tanja Wucherpfennig, Michaela Wilsch-Bräuninger, and Marcos González-Gaitán. Role of Drosophila Rab5 during endosomal trafficking at the synapse and evoked neurotransmitter release. Journal of Cell Biology, 161(3):609–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
3/609.

Williams:2003:FRU

[WWD03] Emma-Jane Williams, Frank S. Walsh, and Patrick Doherty. The FGF receptor uses the endocannabinoid signaling system to couple to an axonal growth response. Journal of Cell Biology, 160(4):481–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/160/4/481.

Wiggins:2004:IBH

[WWD^{+}04] Amanda K. Wiggins, Guangwei Wei, Epaminondas Doxakis, Connie Wong, Amy A. Tang, Keling Zang, Esther J. Luo, Rachael L. Neve, Louis F. Reichardt, and Eric J. Huang. Interaction of Brn3a and HIPK2 mediates transcriptional repression of sensory neuron survival. Journal of Cell Biology, 167(2):257–??, October 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/167/2/257.

Wong:2000:DRG

[WWG^{+}00] Edith D. Wong, Jennifer A. Wagner, Steven W. Gorsich, J. Michael McCaffery, Janet M. Shaw, and Jodi Nunnari. The Dynamin–Related Gtpase, Mgm1p, Is an Intermembrane Space Protein Required for Maintenance of Fusion Competent Mitochondria. Journal of Cell Biology, 151(2): 341–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/2/341.

Wollert:2002:AMV

[WWGK02] Torsten Wöllert, Dieter G. Weiss, Hans-Hermann Gerdes, and Sergei A. Kuznetsov. Activation of myosin V-based motility and F-actin-dependent network formation of endoplasmic reticulum during mitosis. Journal of Cell Biology, 159(4):571–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/159/4/571.

Walker:2000:SNC

[WWJ^{+}00] Diana L. Walker, Dong Wang, Ye Jin, Uttama Rath, Yanming Wang, Jørgen Johansen, and Kristen M. Johansen. Skeletor, a Novel Chromosomal Protein That Redistributes during Mitosis Provides Evidence for the Formation of a Spindle Matrix. Journal of Cell Biology, 151(7):1401–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/151/7/1401.

Wayman:2000:CKI

[WWK^{+}00] Gary A. Wayman, Melinda J. Walters, Kathryn Kolibaba, Thomas R. Soderling, and Jan L. Christian. CAM kinase IV regulates lineage commitment and survival of erythroid progenitors in a non-cell-autonomous manner. Journal of Cell Biology, 151(4):811–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/151/4/811.

Wittmann:2000:TNX

[WWKV00] Torsten Wittmann, Matthias Wilm, Eric Karsenti, and Isabelle Vernos. Tpx2, a Novel Xenopus Map Involved in Spindle Pole Organization. Journal of Cell Biology, 149(7): 1405–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/7/1405.

Wong:2003:IDR

[WWS^{+}03] Edith D. Wong, Jennifer A. Wagner, Sidney V. Scott, Voytek Okreglak, Timothy J. Holewinske, Ann Cassidy-Stone, and Jodi Nunnari. The intramitochondrial dynamin-related GT-Pase, Mgm1p, is a component of a protein complex that mediates mitochondrial fusion. Journal of Cell Biology, 160(3):

303–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/160/3/303.

Wang:2003:RFO

[WXD^{+}03] Qiang Wang, Yi Xie, Quan-Sheng Du, Xiao-Jun Wu, Xu Feng, Lin Mei, Jay M. McDonald, and Wen-Cheng Xiong. Regulation of the formation of osteoclastic actin rings by proline-rich tyrosine kinase 2 interacting with gelsolin. Journal of Cell Biology, 160(4):565–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/160/4/565.

Wei:2000:REH

[WXQ^{+}00] Feng Wei, Zao C. Xu, Zhican Qu, Jeffrey Milbrandt, and Min Zhuo. Role of Egr1 in Hippocampal Synaptic Enhancement Induced by Tetanic Stimulation and Amputation. Journal of Cell Biology, 149(7):1325–??, June 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/149/7/1325.

Weston:2000:AIA

[WYHP00] Christi Weston, Barry Yee, Eldad Hod, and Joav Prives. Agrin-induced Acetylcholine Receptor Clustering Is Mediated by the Small Guanosine Triphosphatases Rac and Cdc42. Journal of Cell Biology, 150(1):205–??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/1/
205.

Ward:2002:GBP

[WYR^{+}02] Yvona Ward, Seow-Fong Yap, V. Ravichandran, Fumio Matsumura, Masaaki Ito, Beth Spinelli, and Kathleen Kelly. The GTP binding proteins Gem and Rad are negative regulators of the Rho–Rho kinase pathway. Journal of Cell Biology, 157 (2):291–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/2/291.

West:2001:LMP

[WZB^{+}01a] Kip A. West, Huaye Zhang, Michael C. Brown, Sotiris N. Nikolopoulos, M. C. Riedy, Alan F. Horwitz, and Christopher E. Turner. The LD4 motif of paxillin regulates cell spreading and motility through an interaction with paxillin kinase linker (PKL). Journal of Cell Biology, 154(1):161–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
154/1/161.

Wu:2001:CED

[WZB^{+}01b] Xufeng Wu, Xiaohong Zhao, Lauren Baylor, Shivani Kaushal, Evan Eisenberg, and Lois E. Greene. Clathrin exchange during clathrin-mediated endocytosis. Journal of Cell Biology, 155(2):291–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/155/2/291.

Xiao:2003:CLP

[XAB^{+}03] Kanyan Xiao, David F. Allison, Kathleen M. Buckley, Margaret D. Kottke, Peter A. Vincent, Victor Faundez, and Andrew P. Kowalczyk. Cellular levels of p120 catenin function as a set point for cadherin expression levels in microvascular endothelial cells. Journal of Cell Biology, 163(3):535– ??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/3/535.

Xu:2004:HRP

[XB04] Lifeng Xu and Elizabeth H. Blackburn. Human Rif1 protein binds aberrant telomeres and aligns along anaphase midzone microtubules. Journal of Cell Biology, 167(5):819–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
167/5/819.

Xi:2003:CRC

[XBL^{+}03] Xiaodong Xi, Richard J. Bodnar, Zhenyu Li, Stephen C.-T. Lam, and Xiaoping Du. Critical roles for the COOH-terminal NITY and RGT sequences of the integrin β_{3} cytoplasmic domain in inside-out and outside-in signaling. Journal of Cell Biology, 162(2):329–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/2/329.

Xu:2004:RCV

[XH04] Dalu Xu and Jesse C. Hay. Reconstitution of COPII vesicle fusion to generate a pre-Golgi intermediate compartment. Journal of Cell Biology, 167(6):997–??, December 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/167/6/
997.

Xia:2000:PTE

[XHG^{+}00] Lu Xia, Bing Hai, Yan Gao, Dylan Burnette, Rupal Thazhath, Jianming Duan, Marie-Helene Bré, Nicolette Levilliers, Martin A. Gorovsky, and Jacek Gaertig. Polyglycylation of tubulin is essential and affects cell motility and division in Tetrahymena thermophila. Journal of Cell Biology, 149(5):1097–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/149/5/1097.

Xiang:2003:LMP

[Xia03] Xin Xiang. LIS1 at the microtubule plus end and its role in dynein-mediated nuclear migration. Journal of Cell Biology, 160(3):289–??, February 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/160/3/289.

Xia:2004:DAF

[XJW^{+}04] Lijun Xia, Tongzhong Ju, Andrew Westmuckett, Guangyu An, Lacramioara Ivanciu, J. Michael McDaniel, Florea Lupu, Richard D. Cummings, and Rodger P. McEver. Defective angiogenesis and fatal embryonic hemorrhage in mice lacking core 1–derived O-glycans. Journal of Cell Biology, 164(3): 451–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/164/3/451.

Xu:2001:RCI

[XLGS01] Jun Xu, Dan Liu, Gordon Gill, and Zhou Songyang. Regulation of cytokine-independent survival kinase (CISK) by the Phox homology domain and phosphoinositides. Journal of Cell Biology, 154(4):699–??, August 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/4/699.

Xu:2001:MMN

[XLH^{+}01] X. Xu, W. E. I. Li, G. Y. Huang, R. Meyer, T. Chen, Y. Luo, M. P. Thomas, G. L. Radice, and C. W. Lo. Modulation of mouse neural crest cell motility by N-cadherin and connexin 43 gap junctions. Journal of Cell Biology, 154(1):217–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/154/
1/217.

Xia:2003:ANT

[XRH^{+}03] Chun-Hong Xia, Elizabeth A. Roberts, Lu-Shiun Her, Xinran Liu, David S. Williams, Don W. Cleveland, and Lawrence S. B. Goldstein. Abnormal neurofilament transport caused by targeted disruption of neuronal kinesin heavy chain KIF5A. Journal of Cell Biology, 161(1):55–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/161/1/55.

Xavier:2004:DLD

[XRI^{+}04] Ramnik Xavier, Shahrooz Rabizadeh, Kazuhiro Ishiguro, Niko Andre, J. Bernabe Ortiz, Heather Wachtel, David G. Morris, Marco Lopez-Ilasaca, Albert C. Shaw, Wojciech Swat, and Brian Seed. Discs large (Dlg1) complexes in lymphocyte activation. Journal of Cell Biology, 166(2):173–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
166/2/173.

Xu:2001:PEC

[XRP^{+}01] Jingsong Xu, Dorothy Rodriguez, Eric Petitclerc, Jenny J. Kim, Masanori Hangai, S. Moon Yuen, George E. Davis, and Peter C. Brooks. Proteolytic exposure of a cryptic site within collagen type IV is required for angiogenesis and tumor growth in vivo. Journal of Cell Biology, 154(5):1069– ??, September 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/154/5/1069.

Xu:2001:MCD

[XSK^{+}01] Wenbo Xu, Michael Shy, John Kamholz, Lisa Elferink, Gang Xu, Jack Lilien, and Janne Balsamo. Mutations in the cytoplasmic domain of P0 reveal a role for PKC-mediated phosphorylation in adhesion and myelination. Journal of Cell Biology, 155(3):439–??, October 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/155/3/439.

Xu:2002:RKM

[XTN^{+}02] Ying Xu, Sen Takeda, Takao Nakata, Yasuko Noda, Yosuke Tanaka, and Nobutaka Hirokawa. Role of KIFC3 motor protein in Golgi positioning and integration. Journal of Cell Biology, 158(2):293–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/158/2/293.

Xu:2003:KAR

[XWL03] Chang Xu, James Watras, and Leslie M. Loew. Kinetic analysis of receptor-activated phosphoinositide turnover. Journal of Cell Biology, 161(4):779–??, May 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/161/4/779.

Yoshida:2001:KKM

[YB01] Kimihisa Yoshida and Günter Blobel. The karyopherin Kap142p/Msn5p mediates nuclear import and nuclear export of different cargo proteins. Journal of Cell Biology, 152(4): 729–??, February 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/4/729.

Young:2004:YAP

[YCB04] Michael E. Young, John A. Cooper, and Paul C. Bridgman. Yeast actin patches are networks of branched actin filaments. Journal of Cell Biology, 166(5):629–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
5/629.

Yu:2003:DRG

[YCK^{+}03] Fengwei Yu, Yu Cai, Rachna Kaushik, Xiaohang Yang, and William Chia. Distinct roles of Gαi and Gβ13F subunits of the heterotrimeric G protein complex in the mediation of Drosophila neuroblast asymmetric divisions. Journal of Cell Biology, 162(4):623–??, August 2003. CODEN JCLBA3.

ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/4/623.

Yang:2001:TSS

[YCX^{+}01] Xiao Yang, Lin Chen, Xiaoling Xu, Cuiling Li, Cuifen Huang, and Chu-Xia Deng. TGF-β / smad3 Signals Repress Chondrocyte Hypertrophic Differentiation and Are Required for Maintaining Articular Cartilage. Journal of Cell Biology, 153 (1):35–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/1/35.

Yu:2000:FRM

[YD00] Hong-Guo Yu and R. Kelly Dawe. Functional Redundancy in the Maize Meiotic Kinetochore. Journal of Cell Biology, 151(1):131–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/151/1/131.

Yamada:2004:DRC

[YDPK04] Ayumi Yamada, Brad Duffy, Jennifer A. Perry, and Sally Kornbluth. DNA replication checkpoint control of Wee1 stability by vertebrate Hsl7. Journal of Cell Biology, 167(5): 841–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/5/841.

Yang:2001:LCD

[YDRS01] Pinfen Yang, Dennis R. Diener, Joel L. Rosenbaum, and Winfield S. Sale. Localization of Calmodulin and Dynein Light Chain Lc8 in Flagellar Radial Spokes. Journal of Cell Biology, 153(6):1315–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/6/1315.

Young:2001:OGS

[YEG01] Paul Young, Elisabeth Ehler, and Mathias Gautel. Obscurin, a giant sarcomeric Rho guanine nucleotide exchange factor protein involved in sarcomere assembly. Journal of Cell Biology, 154(1):123–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/154/1/123.

Yun:2000:CSP

[YF00] Chi Y. Yun and Xiang-Dong Fu. Conserved Sr Protein Kinase Functions in Nuclear Import and Its Action Is Counteracted by Arginine Methylation in Saccharomyces cerevisiae. Journal of Cell Biology, 150(4):707–??, August 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/150/4/707.

Yu:2004:GK

[YFW^{+}04] Jiangtao Yu, Shawna L. Fleming, Byron Williams, Erika V. Williams, ZeXiao Li, Patrizia Somma, Conly L. Rieder, and Michael L. Goldberg. Greatwall kinase. Journal of Cell Biology, 164(4):487–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/164/4/487.

Yeaman:2001:SCT

[YGWN01] Charles Yeaman, Kent K. Grindstaff, Jessica R. Wright, and W. James Nelson. Sec6/ 8 complexes on trans-Golgi network and plasma membrane regulate late stages of exocytosis in mammalian cells. Journal of Cell Biology, 155(4): 593–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/4/593.

Youngman:2004:MMO

[YHB^{+}04] Matthew J. Youngman, Alyson E. Aiken Hobbs, Shawn M. Burgess, Maithreyan Srinivasan, and Robert E. Jensen. Mmm2p, a mitochondrial outer membrane protein required for yeast mitochondrial shape and maintenance of mtDNA nucleoids. Journal of Cell Biology, 164(5):677–??, March 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
5/677.

Yang:2000:PEC

[YHD^{+}00] Changsong Yang, Minzhou Huang, John DeBiasio, Martin Pring, Michael Joyce, Hiroaki Miki, Tadaomi Takenawa, and Sally H. Zigmond. Profilin Enhances Cdc42-Induced Nucleation of Actin Polymerization. Journal of Cell Biology, 150(5):1001–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/150/5/1001.

Yamamoto:2001:PBI

[YHF^{+}01] Masaya Yamamoto, Donald H. Hilgemann, Siyi Feng, Haruhiko Bito, Hisamitsu Ishihara, Yoshikazu Shibasaki, and Helen L. Yin. Phosphatidylinositol 4,5-Bisphosphate Induces Actin Stress–Fiber Formation and Inhibits Membrane Ruffling in Cv1 Cells. Journal of Cell Biology, 152 (5):867–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/5/867.

Yamamoto:2000:CCA

[YHK^{+}00] Ayako Yamamoto, Yoshihiro Hashimoto, Kenjiro Kohri, Etsuro Ogata, Shige aki Kato, Kyoji Ikeda, and Makoto Nakanishi. Cyclin E as a Coactivator of the Androgen Receptor. Journal of Cell Biology, 150(4):873–??, August 2000. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/4/
873.

Yamashita:2002:PRT

[YHT02] Toshihide Yamashita, Haruhisa Higuchi, and Masaya Tohyama. The p75 receptor transduces the signal from myelinassociated glycoprotein to Rho. Journal of Cell Biology, 157 (4):565–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/157/4/565.

Ye:2001:BIL

[YHZ^{+}01] Qinong Ye, Yan-Fen Hu, Hongjun Zhong, Anne C. Nye, Andrew S. Belmont, and Rong Li. BRCA1-induced largescale chromatin unfolding and allele-specific effects of cancerpredisposing mutations. Journal of Cell Biology, 155(6):911– ??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/155/6/911.

Yamazaki:2003:MDH

[YIS^{+}03] Satoru Yamazaki, Ryo Iwamoto, Kazuko Saeki, Masanori Asakura, Seiji Takashima, Ayano Yamazaki, Rina Kimura, Hiroto Mizushima, Hiroki Moribe, Shigeki Higashiyama, Masayuki Endoh, Yasufumi Kaneda, Satoshi Takagi, Satoshi Itami, Naoki Takeda, Gen Yamada, and Eisuke Mekada. Mice with defects in HB–EGF ectodomain shedding show severe developmental abnormalities. Journal of Cell Biology, 163(3): 469–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/3/469.

Yoshida:2003:RRM

[YITe03] Satoshi Yoshida, Ryuji Ichihashi, and Akio Toh-e. Ras recruits mitotic exit regulator Lte1 to the bud cortex in budding yeast. Journal of Cell Biology, 161(5):889–??, June 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/161/
5/889.

Yamaguchi:2004:MEI

[YIW^{+}04] Yuji Yamaguchi, Satoshi Itami, Hidenori Watabe, Ken ichi Yasumoto, Zalfa A. Abdel-Malek, Tateki Kubo, François Rouzaud, Atsushi Tanemura, Kunihiko Yoshikawa, and Vincent J. Hearing. Mesenchymal-epithelial interactions in the skin. Journal of Cell Biology, 165(2):275–??, April 2004. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/165/2/
275.

Yu:2003:GDS

[YJS^{+}03] Li-Ying Yu, Eija Jokitalo, Yun-Fu Sun, Patrick Mehlen, Dan Lindholm, Mart Saarma, and Urmas Arumäe. GDNFdeprived sympathetic neurons die via a novel nonmitochondrial pathway. Journal of Cell Biology, 163(5):987–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/163/5/987.

Yap:2003:DCA

[YK03a] Alpha S. Yap and Eva M. Kovacs. Direct cadherin-activated cell signaling. Journal of Cell Biology, 160(1):11–??, January 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/160/
1/11.

Yu:2003:MCR

[YK03b] Hong-Guo Yu and Douglas E. Koshland. Meiotic condensin is required for proper chromosome compaction, SC assembly, and resolution of recombination-dependent chromosome linkages. Journal of Cell Biology, 163(5):937–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/163/
5/937.

Yang:2004:PSA

[YKT^{+}04] Xiuwei Yang, Oleg V. Kovalenko, Wei Tang, Christoph Claas, Christopher S. Stipp, and Martin E. Hemler. Palmitoylation supports assembly and function of integrin–tetraspanin complexes. Journal of Cell Biology, 167(6):1231–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
6/1231.

Yin:2000:SCM

[YKW^{+}00] X. Yin, G. J. Kidd, L. Wrabetz, M. L. Feltri, A. Messing, and B. D. Trapp. Schwann Cell Myelination Requires Timely and Precise Targeting of P_{0} Protein. Journal of Cell Biology, 148 (5):1009–??, March 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/148/5/1009.

Yang:2002:APF

[YLG^{+}02] Jia-Shu Yang, Stella Y. Lee, Minggeng Gao, Sylvain Bourgoin, Paul A. Randazzo, Richard T. Premont, and Victor W. Hsu. ARFGAP1 promotes the formation of COPI vesicles, suggesting function as a component of the coat. Journal of Cell Biology, 159(1):69–??, October 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/159/1/69.

Yago:2002:DMC

[YLQ^{+}02] Tadayuki Yago, Anne Leppänen, Haiying Qiu, Warren D. Marcus, Matthias U. Nollert, Cheng Zhu, Richard D. Cummings, and Rodger P. McEver. Distinct molecular and cellular contributions to stabilizing selectin-mediated rolling under flow. Journal of Cell Biology, 158(4):787–??, August 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/158/
4/787.

Yang:2002:RNC

[YLY^{+}02] Jun Yang, Xiaoqing Liu, Guohua Yue, Michael Adamian, Oleg Bulgakov, and Tiansen Li. Rootletin, a novel coiled-coil protein, is a structural component of the ciliary rootlet. Journal of Cell Biology, 159(3):431–??, November 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/159/3/431.

Young:2001:H

[YMH01] Jason C. Young, Ismail Moarefi, and F. Ulrich Hartl. Hsp90. Journal of Cell Biology, 154(2):267–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/154/2/267.

Yano:2004:RPS

[YMK^{+}04] Hajime Yano, Yuichi Mazaki, Kazuo Kurokawa, Steven K. Hanks, Michiyuki Matsuda, and Hisataka Sabe. Roles played by a subset of integrin signaling molecules in cadherin-based cell–cell adhesion. Journal of Cell Biology, 166(2):283–??, July 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
166/2/283.

Yucel:2000:CME

[YMM^{+}00] Jennifer K. Yucel, Janet D. Marszalek, J. Richard McIntosh, Lawrence S. B. Goldstein, Don W. Cleveland, and Alastair Valentine Philp. CENP-meta, an Essential Kinetochore Kinesin Required for the Maintenance of Metaphase Chromosome Alignment in Drosophila. Journal of Cell Biology, 150(1):??, July 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/150/1/1.1.

Ye:2003:FPU

[YMR03] Yihong Ye, Hemmo H. Meyer, and Tom A. Rapoport. Function of the p97–Ufd1–Npl4 complex in retrotranslocation from the ER to the cytosol. Journal of Cell Biology, 162 (1):71–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/162/1/71.

Yoshizaki:2003:ARF

[YOK^{+}03] Hisayoshi Yoshizaki, Yusuke Ohba, Kazuo Kurokawa, Reina E. Itoh, Takeshi Nakamura, Naoki Mochizuki, Kazuo Nagashima, and Michiyuki Matsuda. Activity of Rho-family GTPases during cell division as visualized with FRET-based probes. Journal of Cell Biology, 162(2):223–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/162/2/
223.

Yang:2003:CII

[YpHRL03] Feng Yang, Xiang ping He, James Russell, and Bai Lu. Ca^{2+} influx-independent synaptic potentiation mediated by mitochondrial Na^{+}–Ca^{2+} exchanger and protein kinase C. Journal of Cell Biology, 163(3):511–??, November 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/163/3/511.

Yang:2004:MCS

[YPN^{+}04] Jianbo Yang, Matthew A. Price, Cheryl L. Neudauer, Christopher Wilson, Soldano Ferrone, Hong Xia, Joji Iida, Melanie A. Simpson, and James B. McCarthy. Melanoma chondroitin sulfate proteoglycan enhances FAK and ERK activation by distinct mechanisms. Journal of Cell Biology, 165 (6):881–??, June 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/165/6/881.

Younger:2004:FCF

[YRC^{+}04] J. Michael Younger, Hong-Yu Ren, Liling Chen, Chun-Yang Fan, Andrea Fields, Cam Patterson, and Douglas M. Cyr. A foldable CFTRΔF508 biogenic intermediate accumulates upon inhibition of the Hsc70–CHIP E3 ubiquitin ligase. Journal of Cell Biology, 167(6):1075–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/167/6/1075.

You:2002:WSP

[YSC^{+}02] Zongbing You, Daniel Saims, Shaoqiong Chen, Zhaocheng Zhang, Denis C. Guttridge, Kun liang Guan, Ormond A. MacDougald, Anthony M. C. Brown, Gerard Evan, Jan Kitajewski, and Cun-Yu Wang. Wnt signaling promotes oncogenic transformation by inhibiting c–Myc–induced apoptosis.

Journal of Cell Biology, 157(3):429–??, April 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/157/3/429. See

correction [YSC^{+}21].

You:2021:CWS

[YSC^{+}21] Zongbing You, Daniel Saims, Shaoqiong Chen, Zhaocheng Zhang, Denis C. Guttridge, Kun liang Guan, Ormond A. MacDougald, Anthony M. C. Brown, Gerard Evan, Jan Kitajewski, and Cun-Yu Wang. Correction: Wnt signaling promotes oncogenic transformation by inhibiting c-Myc-induced apoptosis. Journal of Cell Biology, 220(5):??, May 3, 2021. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL https://rupress.org/jcb/article/220/5/
jcb.20020111004052021c/211961/Correction-Wnt-signaling-
promotes-oncogenic.

Yang:2000:NBC

[YSGS00] Bin Yang, Martin Steegmaier, Lino C. Gonzalez, and Richard H. Scheller. Nsec1 Binds a Closed Conformation of Syntaxin1a. Journal of Cell Biology, 148(2):247–??, January 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/148/
2/247.

Yamaji:2004:AIA

[YSK^{+}04] Satoshi Yamaji, Atsushi Suzuki, Heiwa Kanamori, Wataru Mishima, Ryusuke Yoshimi, Hirotaka Takasaki, Maki Takabayashi, Katsumichi Fujimaki, Shin Fujisawa, Shigeo Ohno, and Yoshiaki Ishigatsubo. Affixin interacts with α-actinin and mediates integrin signaling for reorganization of F-actin induced by initial cell–substrate interaction. Journal of Cell Biology, 165(4):539–??, May 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/165/4/539.

Yu:2001:IMU

[YSN^{+}01] Yi-Tao Yu, Mei-Di Shu, Aarthi Narayanan, Rebecca M. Terns, Michael P. Terns, and Joan A. Steitz. Internal Modification of U2 Small Nuclear (Snrna) Occurs in Nucleoli of Xenopus Oocytes. Journal of Cell Biology, 152(6): 1279–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/6/1279.

Yamaji:2001:NIL

[YSS^{+}01] Satoshi Yamaji, Atsushi Suzuki, Yuki Sugiyama, Yu ichi Koide, Michihiko Yoshida, Heiwa Kanamori, Hiroshi Mohri, Shigeo Ohno, and Yoshiaki Ishigatsubo. A Novel Integrin– Linked Kinase–Binding Protein, Affixin, Is Involved in the Early Stage of Cell–Substrate Interaction. Journal of Cell Biology, 153(6):1251–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/153/6/1251.

Yan:2002:MKA

[YSW02] Yibing Yan, Kyoko Shirakabe, and Zena Werb. The metalloprotease Kuzbanian (ADAM10) mediates the transactivation of EGF receptor by G protein–coupled receptors. Journal of Cell Biology, 158(2):221–??, July 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/158/2/221.

Yano:2003:AMI

[YTM03] Masato Yano, Kazutoyo Terada, and Masataka Mori. AIP is a mitochondrial import mediator that binds to both import receptor Tom20 and preproteins. Journal of Cell Biology, 163 (1):45–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/163/1/45.

Yumura:2001:MID

[Yum01] Shigehiko Yumura. Myosin II dynamics and cortical flow during contractile ring formation in Dictyostelium cells. Journal of Cell Biology, 154(1):137–??, July 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/154/1/137.

Yvon:2000:RSM

[YW00] Anne-Marie C. Yvon and Patricia Wadsworth. Regionspecific Microtubule Transport in Motile Cells. Journal of Cell Biology, 151(5):1003–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/5/1003.

Yi:2004:CMM

[YWH04] Muqing Yi, David Weaver, and György Hajnóczky. Control of mitochondrial motility and distribution by the calcium signal. Journal of Cell Biology, 167(4):661–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/167/
4/661.

Yago:2004:CBG

[YWW^{+}04] Tadayuki Yago, Jianhua Wu, C. Diana Wey, Arkadiusz G. Klopocki, Cheng Zhu, and Rodger P. McEver. Catch bonds govern adhesion through L-selectin at threshold shear. Journal of Cell Biology, 166(6):913–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/166/6/913.

Yoon:2001:IDP

[YYM^{+}01] Kyeong Han Yoon, Miri Yoon, Robert D. Moir, Satya Khuon, Frederick W. Flitney, and Robert D. Goldman. Insights into the Dynamic Properties of Keratin Intermediate Filaments in Living Epithelial Cells. Journal of Cell Biology, 153(3): 503–??, April 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/153/3/503.

Zhang:2001:IRN

[yZCKA01] Bao yan Zhang, Amy Chang, Thomas B. Kjeldsen, and Peter Arvan. Intracellular Retention of Newly Synthesized Insulin in Yeast Is Caused by Endoproteolytic Processing in the Golgi Complex. Journal of Cell Biology, 153(6):1187–??, June 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
153/6/1187.

Zink:2004:TDS

[ZAE^{+}04] Daniele Zink, Margarida D. Amaral, Andreas Englmann, Susanne Lang, Luka A. Clarke, Carsten Rudolph, Felix Alt, Kathrin Luther, Carla Braz, Nicolas Sadoni, Joseph Rosenecker, and Dirk Schindelhauer. Transcription-dependent spatial arrangements of CFTR and adjacent genes in human cell nuclei. Journal of Cell Biology, 166(6):815–??, September 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/166/
6/815.

Zambryski:2004:CCT

[Zam04] Patricia Zambryski. Cell-to-cell transport of proteins and fluorescent tracers via plasmodesmata during plant development. Journal of Cell Biology, 164(2):165–??, January 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
2/165.

Zebda:2000:PAC

[ZBB^{+}00] Noureddine Zebda, Ora Bernard, Maryse Bailly, Susan Welti, David S. Lawrence, and John S. Condeelis. Phosphorylation of Adf/cofilin abolishes Egf-induced actin nucleation at the leading edge and subsequent lamellipod extension. Journal of Cell Biology, 151(5):1119–??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/5/1119.

Zhou:2001:ACP

[ZCH^{+}01] Qiang Zhou, Po-Hsien Chu, Chenqun Huang, Ching-Feng Cheng, Maryann E. Martone, Gudrun Knoll, G. Diane Shelton, Sylvia Evans, and Ju Chen. Ablation of Cypher, a PDZ– LIM domain Z-line protein, causes a severe form of congenital myopathy. Journal of Cell Biology, 155(4):605–??, November 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/155/
4/605.

Zallen:2002:SPR

[ZCH^{+}02] Jennifer A. Zallen, Yehudit Cohen, Andrew M. Hudson, Lynn Cooley, Eric Wieschaus, and Eyal D. Schejter. SCAR is a primary regulator of Arp2/3-dependent morphological events in Drosophila. Journal of Cell Biology, 156(4):689–??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/156/
4/689.

Zhou:2003:HFT

[ZCW^{+}03] Hui Zhou, Fengli Cao, Zhishan Wang, Zhao-Xue Yu, Huu-Phuc Nguyen, Joy Evans, Shi-Hua Li, and Xiao-Jiang Li. Huntingtin forms toxic NH_{2}-terminal fragment complexes that are promoted by the age-dependent decrease in proteasome activity. Journal of Cell Biology, 163(1):109–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
163/1/109.

Zetter:2001:HL

[Zet01] Bruce R. Zetter. Hold That Line. Journal of Cell Biology, 152(6):0–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/152/6/F35.

Zondag:2000:ORD

[ZEtK^{+}00] Gerben C. M. Zondag, Eva E. Evers, Jean P. ten Klooster, Lennert Janssen, Rob A. van der Kammen, and John G. Collard. Oncogenic Ras Downregulates Rac Activity, Which Leads to Increased Rho Activity and Epithelial– Mesenchymal Transition. Journal of Cell Biology, 149(4): 775–??, May 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/149/4/775.

Zanazzi:2001:GGF

[ZEW^{+}01] George Zanazzi, Steven Einheber, Richard Westreich, Melanie-Jane Hannocks, Debra Bedell-Hogan, Mark A. Marchionni, and James L. Salzer. Glial growth factor/neuregulin inhibits Schwann cell myelination and induces demyelination. Journal of Cell Biology, 152(6):1289–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/152/6/
1289.

Zippin:2004:BRS

[ZFH^{+}04] Jonathan H. Zippin, Jeanne Farrell, David Huron, Margarita Kamenetsky, Kenneth C. Hess, Donald A. Fischman, Lonny R. Levin, and Jochen Buck. Bicarbonate-responsive “soluble” adenylyl cyclase defines a nuclear cAMP microdomain. Journal of Cell Biology, 164(4):527–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
4/527.

Zervas:2001:DIL

[ZGB01] Christos G. Zervas, Stephen L. Gregory, and Nicholas H. Brown. Drosophila Integrin–Linked Kinase Is Required at Sites of Integrin Adhesion to Link the Cytoskeleton to the Plasma Membrane. Journal of Cell Biology, 152(5):1007–??, March 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.org/

content/152/5/1007.

Zammit:2004:MSC

[ZGN^{+}04] Peter S. Zammit, Jon P. Golding, Yosuke Nagata, Valérie Hudon, Terence A. Partridge, and Jonathan R. Beauchamp. Muscle satellite cells adopt divergent fates. Journal of Cell Biology, 166(3):347–??, August 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/166/3/347.

Zelhof:2004:WRC

[ZH04] Andrew C. Zelhof and Robert W. Hardy. WASp is required for the correct temporal morphogenesis of rhabdomere microvilli. Journal of Cell Biology, 164(3):417–??, February 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/164/
3/417.

Zeitler:2004:DCC

[ZHDB04] Jennifer Zeitler, Cynthia P. Hsu, Heather Dionne, and David Bilder. Domains controlling cell polarity and proliferation in the Drosophila tumor suppressor Scribble. Journal of Cell Biology, 167(6):1137–??, December 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/6/1137.

Zhao:2002:BMP

[ZHH^{+}02] Ming Zhao, Stephen E. Harris, Diane Horn, Zhaopo Geng, Riko Nishimura, Gregory R. Mundy, and Di Chen. Bone morphogenetic protein receptor signaling is necessary for normal murine postnatal bone formation. Journal of Cell Biology, 157(6):1049–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/6/1049.

Zigmond:2000:HWR

[Zig00] Sally H. Zigmond. How Wasp Regulates Actin Polymerization. Journal of Cell Biology, 150(6):0–??, September 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/150/6/
F117.

Zhao:2002:GAR

[ZJM^{+}02] Min Zhao, Tian Jin, Colin D. McCaig, John V. Forrester, and Peter N. Devreotes. Genetic analysis of the role of G protein–coupled receptor signaling in electrotaxis. Journal of Cell Biology, 157(6):921–??, June 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/157/6/921.

Zhang:2000:RXT

[ZKW^{+}00] Lijun Zhang, Thomas J. Keating, Andrew Wilde, Gary G. Borisy, and Yixian Zheng. The Role of Xgrip210 in γ-Tubulin Ring Complex Assembly and Centrosome Recruitment. Journal of Cell Biology, 151(7):1525–??, December 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/151/7/1525.

Zahraoui:2000:TJP

[ZLG00] Ahmed Zahraoui, Daniel Louvard, and Thierry Galli. Tight Junction, a Platform for Trafficking and Signaling Protein Complexes. Journal of Cell Biology, 151(5):??, November 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/151/
5/F31.

Zong:2003:BBC

[ZLH^{+}03] Wei-Xing Zong, Chi Li, Georgia Hatzivassiliou, Tullia Lindsten, Qian-Chun Yu, Junying Yuan, and Craig B. Thompson. Bax and Bak can localize to the endoplasmic reticulum to initiate apoptosis. Journal of Cell Biology, 162(1):59–??, July 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/162/
1/59.

Zhou:2000:ARC

[ZLMP00] Honglin Zhou, Xin-Ming Li, Judy Meinkoth, and Randall N. Pittman. Akt Regulates Cell Survival and Apoptosis at a Postmitochondrial Level. Journal of Cell Biology, 151(3): 483–??, October 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/151/3/483.

Zahir:2003:ALL

[ZLR^{+}03] Nastaran Zahir, Johnathon N. Lakins, Alan Russell, WenYu Ming, Chandrima Chatterjee, Gabriela I. Rozenberg, M. Peter Marinkovich, and Valerie M. Weaver. Autocrine laminin-5 ligates α6β4 integrin and activates RAC and NFκB to mediate anchorage-independent survival of mammary tumors. Journal of Cell Biology, 163(6):1397–??, December 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/163/6/
1397.

Zhang:2002:PAK

[ZLVS02] Hongquan Zhang, Zhilun Li, Eva-Karin Viklund, and Staffan Strömblad. p21-activated kinase 4 interacts with integrin αvβ5 and regulates αvβ5-mediated cell migration. Journal of Cell Biology, 158(7):1287–??, September 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/158/7/1287.

Zhang:2002:SCI

[ZMGL02] Yan Zhang, Richard McLaughlin, Cynthia Goodyer, and Andréa LeBlanc. Selective cytotoxicity of intracellular amyloid β peptide_{1−−42} through p53 and Bax in cultured primary human neurons. Journal of Cell Biology, 156(3):519– ??, February 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/156/3/519.

Zhao:2001:IAR

[ZN01] Tieming Zhao and Peter J. Newman. Integrin Activation by Regulated Dimerization and Oligomerization of Platelet Endothelial Cell Adhesion Molecule (Pecam)-1 from within the Cell. Journal of Cell Biology, 152(1):65–??, January 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140

(electronic). URL http://jcb.rupress.org/content/152/
1/65.

Zimmermann:2002:RAS

[ZRDG02] Katja C. Zimmermann, Jean-Ehrland Ricci, Nathalie M. Droin, and Douglas R. Green. The role of ARK in stressinduced apoptosis in Drosophila cells. Journal of Cell Biology, 156(6):1077–??, March 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/156/6/1077.

Zeitlin:2001:CPA

[ZSS01] Samantha G. Zeitlin, Richard D. Shelby, and Kevin F. Sullivan. CENP-A is phosphorylated by Aurora B kinase and plays an unexpected role in completion of cytokinesis. Journal of Cell Biology, 155(7):1147–??, December 2001. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic).

URL http://jcb.rupress.org/content/155/7/1147.

Zeng:2003:PRI

[ZSY^{+}03] Li Zeng, Xiaoning Si, Wei-Ping Yu, Hoa Thi Le, Kwok Peng Ng, Raymond M. H. Teng, Kenneth Ryan, Dennis Z.-M. Wang, Sathivel Ponniah, and Catherine J. Pallen. PTPα regulates integrin-stimulated FAK autophosphorylation and cytoskeletal rearrangement in cell spreading and migration. Journal of Cell Biology, 160(1):137–??, January 2003. CO-DEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (elec-

tronic). URL http://jcb.rupress.org/content/160/1/
137.

Zhang:2003:DLB

[ZTK^{+}03] Feng Zhang, Clifford C. Tom, Matthias C. Kugler, Tsui-Ting Ching, Jordan A. Kreidberg, Ying Wei, and Harold A. Chapman. Distinct ligand binding sites in integrin α3β1 regulate matrix adhesion and cell–cell contact. Journal of Cell Biology, 163(1):177–??, October 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/1/177.

Zhang:2003:VCF

[ZVPK03] Huanxiang Zhang, Laszlo Vutskits, Michael S. Pepper, and Jozsef Z. Kiss. VEGF is a chemoattractant for FGF-2– stimulated neural progenitors. Journal of Cell Biology, 163 (6):1375–??, December 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/163/6/1375.

Zeitler:2004:FRA

[ZW04] Bryan Zeitler and Karsten Weis. The FG-repeat asymmetry of the nuclear pore complex is dispensable for bulk nucleocytoplasmic transport in vivo. Journal of Cell Biology, 167(4): 583–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.rupress.

org/content/167/4/583.

Zhang:2003:SFR

[ZWAH03] Huaye Zhang, Donna J. Webb, Hannelore Asmussen, and Alan F. Horwitz. Synapse formation is regulated by the signaling adaptor GIT1. Journal of Cell Biology, 161(1):131–??, April 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
161/1/131.

Zhao:2004:MDA

[ZWB04] Xiaolan Zhao, Chia-Yung Wu, and Günter Blobel. Mlpdependent anchorage and stabilization of a desumoylating enzyme is required to prevent clonal lethality. Journal of Cell Biology, 167(4):605–??, November 2004. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/167/4/605.

Zweifach:2000:TCC

[Zwe00] Adam Zweifach. Target-cell Contact Activates a Highly Selective Capacitative Calcium Entry Pathway in Cytotoxic T Lymphocytes. Journal of Cell Biology, 148(3):603–??, February 2000. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-

8140 (electronic). URL http://jcb.rupress.org/content/
148/3/603.

Zhou:2002:FLA

[ZWSC02] Feng-Quan Zhou, Clare M. Waterman-Storer, and Christopher S. Cohan. Focal loss of actin bundles causes microtubule redistribution and growth cone turning. Journal of Cell Biology, 157(5):839–??, May 2002. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http://jcb.

rupress.org/content/157/5/839.

Zheng:2003:TXC

[ZZM^{+}03] Qiping Zheng, Guang Zhou, Roy Morello, Yuqing Chen, Xavier Garcia-Rojas, and Brendan Lee. Type X collagen gene regulation by Runx2 contributes directly to its hypertrophic chondrocyte–specific expression in vivo. Journal of Cell Biology, 162(5):833–??, September 2003. CODEN JCLBA3. ISSN 0021-9525 (print), 1540-8140 (electronic). URL http:

//jcb.rupress.org/content/162/5/833.